A Complete Bibliography of Publications in Marine Policy (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 October 2023 Version 1.08

Title word cross-reference

2 [GWWH17, GZW^{+}19a, GZW^{+}19b, Mil14, ZEP^{+}19, dlFRG17]. e [DDL^{+}17]. i [COG^{+}19, IOG18].

-13 [Rah19].

1-116 [Ano15x]. 100 [BB19b]. 1026 [Vat17]. 1026/2012 [Vat17]. 104 [HS19]. 10m [GKSJ11]. 1117-1388 [Ano10h]. 1193-1314 [Ano12h]. 121 [KS18a]. 1379/2013 [DAGG16]. 14 [Vir18]. 14.5 [DJR^{+}18]. 140 [PDC19]. 14th [KSH^{+}15]. 19-3 [PQLH15]. 1940s [Nag16]. 1950s [WP13]. 197-352 [Ano10i]. 1970s [JKS13]. 1992 [Cho10]. 1995 [HFF11, PDC10]. 19th [BGG19].

20 [FQ13]. 2000 [Foc11, PW16]. 2000s [WP13]. 2009 [BE15, Rei10]. 2010’ [PF12a]. 2010s [JKS13]. 2012 [Vat17]. 2013 [Ano13g, Ano13j, Ano13i, Ano13h, LLS^{+}15]. 2014 [Ano14o, Ano14n, Ano14i, Ano14m, Ano14k, Ano14j, Ano14l]. 2015 [Ano15p, Ano15m, Ano15u, Ano15q, Ano15w, Ano15n, Ano15r, Ano15s, Ano15v, Ano15t, Ano15x, Ano15o]. 2016 [Ano16n, Ano16r, Ano16x, Ano16o, Ano16q, Ano16t, Ano16s, Ano16m, Ano16p, Ano16u, Ano16v, Ano16w]. 2017 [Ano17n, Ano17p, Ano17r, Ano17o, Ano17t, Ano17v, Ano17q, Ano17w, Ano17m, Ano17u, Ano17s]. 2018 [Ano18p, Ano18q, Ano18u, Ano18v, Ano18n, Ano18w, Ano18m, Ano18s, Ano18o, Ano18t, Ano18r, Ano18x, PZ19]. 2019 [Ano19a, Ano19b, Ano19o, Ano19p, Ano19w, Ano19u, Ano19v, Ano19s, Ano19q, Ano19t, Ano19r, Ano19y]. 2030 [KS18a, WMVR18]. 21st [HC19, SH10]. 23^{o} [GCP19]. 271-426 [Ano11h].

3 [PRO16, Rah19]. 34 [RLM12]. 35 [HD11a]. 353-738 [Ano10j]. 36 [JS13, LGH^{+}12a]. 39 [Thi13a].

400 [McL15]. 43C [BFGR14a].

5-year [VGA15]. 51 [McL15]. 55 [Wit15b]. 559-736 [Ano11i]. 567-844 [Ano12j].

68 [PBW17].

73 [KS18a, WMVR18]. 75 [GDD^{+}19].

800-Pound [HTF^{+}15]. 821-1116 [Ano10k]. 845-946 [Ano12k]. 85 [PDC19]. 86 [Hey18].

947-1192 [Ano12l]. 95-270 [Ano11j]. 978 [Rah19]. 978-1-59114-693-3 [Rah19]. 978-1-78643-220-9 [Tec19].

A.P [Tec19]. abalone [GDD11, GDD13, LRJ^{+}19, SARF16]. abandoned [Gil15]. Abatement [KLN17, WMWH18]. Abbot [LBN15]. ability [CETQ16, RCM18]. ABMTs [De 18, JFK18]. ABNJ [ARGW14, Has18, JSG^{+}18]. above [LGDB17]. abroad [MM15b]. Abrolhos [GTL^{+}18]. absent [MJS16]. abundance [HC18, Hor10, MACM14, MRSL16, MMDS12, RW15]. abyss [CSF^{+}18]. academic [Cha17b, ETEW15]. acanthias [HKR19]. accede [Smi17]. acceptability [GR16, VGG15, VGBG15]. Acceptable [BLCD16, BRR18a]. Acceptance [DZLL15, BRR18b]. Access [BE15, Gau10, TLB19, Wos19, AJ12, BMW17, BAA^{+}18, BKHA^{+}18, CBL18, CS17, Cou11, CE13, Hav10, NYJW15, Nol19, PM14, RAG18, RW15, Sch17, TTS14, VPL^{+}13, WSB19, WCCO18]. accessibility [BTOM13, ZPT^{+}16]. accessible [RHB17]. accident [BA19, FW15, KH11]. accidental [GLSA10, WIC18]. accidents [CL16b, MÁSSI13, NPS^{+}19]. account [SL15a, dBMA13]. accountabilities [Ben12]. Accountability [BMAG17, RIK16]. Accounting [Bre15, CMAA13, HK11, RLM10, RLM12, DRGP19, Dva19, MBBDS19]. accumulation [Gat18, RRCS^{+}19]. ACE [SM17b]. Aceh [KH15]. achieve [LGD^{+}19, LT17, Liu13, RWG^{+}19, Sea17b, Sol14]. achieved [ARGW14]. achievement [MB15]. achievements [SWG18, LSK18]. Achieving [DSWG13, RRMSSE16, BRR^{+}12, CBE^{+}19, FJ12, Gib10, HBP16, KKSC19, NALO19, NLP11, SI18, SPRH19]. acidification [FMKT15, HKHG19, RvdBG13, SPP18]. Acknowledging [CCR16]. acoustic [BFL18]. acoustics [HVB^{+}17]. across [BHM16, CBE^{+}19, KJS14, Lew16, MNA15, RRRM^{+}15, WSG18, ZSZPR19]. Act [BE15, TLB19, AJ12, AMR14, DSR12, BPMY18, RBC^{+}13, SC13a]. Action [DW13, DBE^{+}16, Elo10, FBP^{+}18, KBB16, ADH18, ATB^{+}18, BAY^{+}19, DB19b, FKWV19, HGG^{+}18, Lew16, MFB^{+}14, PMN^{+}15, SVA^{+}12, SJ19c, VBSD17]. Actions [HSRH16, MSGM^{+}12, GBK^{+}15, Hua11]. active [Tan17, VAHG^{+}18]. Activities [Smi18b, BS18, CCD^{+}13, CSBE14, CKBS14, DSP^{+}18, FBP^{+}18, PWH^{+}16, Ped13, Ray19, SSW^{+}18, SvHdB^{+}15, THNR12, UEMRG^{+}18, WL18, WTA^{+}16]. activity [AdV18, FTG^{+}18, KMC^{+}17, Kai15a, Kai15b, Kai16, MS14, MOH11, RBS^{+}19, SPS14, TPS15]. actor [PDMÖ19]. actors [BDHSJ19, GSW18, Jet18, NJ18, DB19a]. actual [Edd14, RRRM^{+}16]. Acu [DR14]. ad [SRSN17]. adapt [Kee18, KO11, SNDB^{+}16]. adaptability [San10, SS19a, SNDB^{+}16]. Adaptation [CLLS13, Gra10, MI15, PKM^{+}19, BvPS15, CRCD13, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, FOG^{+}11, GOO^{+}19, ISHP14, JOE^{+}19, LDF^{+}18, LSTD12, MAJM^{+}18, ODF^{+}16, RHW^{+}19, RDVSC16, RCM18, RHB17, SSD17b, SSD^{+}16, SCK15, WCMHO18, KW18, Var11]. Adaptations [BCMH^{+}18, HIBG15]. Adaptative [GLVLGG19]. Adapting [NBMB^{+}18, SSD17a, MBD13]. Adaptive [BHAHW17, FBP^{+}18, GHP^{+}15, MOR15, MLV^{+}11, BLC16, BDB19, Boh19, BGS13, BMGH11, Cra19, CBP^{+}18, CHvKM15, DA11, DDS^{+}11, DSH^{+}19, GSB^{+}17, Hey17, Hey18, HLL^{+}18, Jae16, Jet18, KNA10, MHA^{+}17, MG14, NBFO18, PS19, SBMD13, SBSK19, TR18, TKN19, GSB^{+}17]. added [RUD^{+}12, SDP^{+}16]. Adding [GZW^{+}19b, ZEP^{+}19]. additional [Maz17]. additive [KJP14]. address [ARS16, EGGT12, Gil14, HFM^{+}13, JHN^{+}19, SM18b, Stø19]. Addressing [BS18, CR11, GZW^{+}19b, GMM15, HCBM15, AdSC11, CKN19, HA13, dlTCL10]. Adelaide [KC13a, MAR^{+}15]. Adélie [TRCR^{+}19]. adequate [MNHN18, WMSB15]. adequately [dMAH^{+}18]. adjacent [OIS^{+}18]. adjustment [BB19d]. adjustments [Bro10]. Administration [CL19]. administrative [KH11, SCMS15]. adopting [DBD19, HFF11]. Adoption [KPF12, DP17, DPB17, ETEW15, KCKA^{+}18, SSV18]. Adriatic [BMF^{+}16, LBD^{+}18, RGB^{+}19]. Adriatic-Ionian [RGB^{+}19]. advance [BSMS15, BTC18, Cha11, GSR19, HD16a, RHW^{+}19]. advances [CPBQ19, BAA10]. Advancing [MAAB15, WAG^{+}15, DTMJH16, NWW^{+}17, OGM^{+}14, Cha16, HR14]. advantage [NBD15, PT10]. advantages [XJT18]. Adventures [Bai16]. advice [GNHM18, Kno10, Pol12b, RMRN^{+}16a, RS17, Wen16, Wen17, vHK17]. advise [Far19]. Advisory [CRO12, OR19, HG14]. advocates [dO13]. Aegean [KLN17]. affect [KHD12, LQC10, SCC^{+}17]. affecting [AK15, Bre13, DP17, LCC19, RVdCGN11, SOUN16]. affiliation [MLHP15]. Africa [Cro15, Wit15b, BSD^{+}17, BF15, BDvTvKT17, BTS^{+}17, BSL19, BDHSJ19, CN17, COS14, CR11, Gre15, GCD^{+}11, IW19, JHB^{+}18, Nei15, Nij17, NW10, Nor15, OPMB12, PP10, SC10, SSRS14, TCC^{+}19, TGBL11, TKN19, VKA^{+}19, Wit15a, dCW15]. African [Wit15b, BLC16, GCG^{+}10, PIH^{+}10, SJ14, Sun12, TvN13, Wit15a, vAS19]. after [AT14, CdLR^{+}15, CC17a, Che12b, CCS16, FGW15, MMJM16, McN18, NY13, Her18b]. against [JOE^{+}19, Kim19, LGC16, RWG^{+}19]. Agardy [WFK^{+}12b]. age [Bro11, NKHL18, SS11]. age-selectivity [SS11]. agencies [FSMG10]. agency [GDD13, HGG^{+}18, MO14, WS15b]. agenda [BJS18, CM17, GRMAC18, GSS^{+}19, LPR^{+}17, MF19c, Men18, RFG^{+}13, WOdG^{+}16, dS17a, vHFJ^{+}19]. Aggregate [ACSK^{+}10]. aggregates [CG18]. aggregating [BAA^{+}15a, DMMG14]. aggregations [HGB^{+}19]. Agnew [FP13]. agreed [Lea19]. Agreeing [Lea19, PZ19]. agreement [BDB19, BY16, CMS19, DG14, Has18, KPF12, KLMD15, Mar17, WTD^{+}15]. agreements [Ant16, ARGW14, CO16, GvdB13, HTT10, Hav10, LAH^{+}13, NYJW15, Sch17, Tei18, WM19]. ahead [QSHI17, STG11]. AHP [Mor17]. Aichi [LGD^{+}19, MSdAM16, MB15, RFL^{+}18]. aid [BW18a, MTF^{+}15]. ails [PRG^{+}17]. aims [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l]. air [KLN17]. Aires [Bai12]. AIS [BA19, EGD18, FTG^{+}18, JMJ^{+}18, Kai16, Mul19, She15, SSH^{+}19]. al [DDL^{+}17, FP13, GZW^{+}19b, Han19, HAK^{+}19, HCF^{+}17, HS19, Kep14, PC13a, PZ17a, Sch15a, SJ19c, Tur10]. al. [Lon12]. Alaska [CL15, Car15, CBL18, CS13, DSH14, Don16, FFKO15, FMKT15, HK18, Jen15, Kna11, KCCO18, LL15, LCC19, Man10, RYRYM17, RYD18, RCD^{+}18, Seu16, TH15, VH16]. Alaskan [Lan15, PC13b, SRK14]. albacares [BB19b, BB19c]. albacore [SPRH19]. Alentejano [SFCI14]. Aleutian [BFL18]. algae [RU12]. algal [MCB^{+}19, WPRA18]. Algarve [RPG^{+}14, RPG^{+}15, RPdO^{+}19]. Aligning [Ben17a, BW18a, KH10, SC19, WF18, vdPLMS16]. alive [KCM^{+}18]. alleviation [DMW19, MSS19, SSRS14]. Alliance [CR10]. alliances [FGPP18]. allocated [Lyn14]. Allocating [GNW19, HMK19]. Allocation [AM15, BIPS13, CMFG15, COS14, DRH^{+}18, HK11, NMvPNL15, NM12, SH14b, SF16, TI13, TZBBL18]. allocations [Don16]. allowable [Bor18, MKSS18, PDS15, WTA^{+}16]. allowed [Kim17]. along [Asa12, ATB^{+}18, CJW^{+}16, DPB17, FC19b, LRJ^{+}19, MB15, MGB^{+}18, PRBX19, RRCS^{+}19, SGN^{+}17]. alter [MHK^{+}10]. alternate [BMAG17]. Alternative [TII^{+}19, BRM^{+}17, BBE^{+}16, Bro18, COS14, CP15, DRH^{+}18, HK10, Kat16, LH12, JCBM10, WS17]. Alternatives [Pin13, CH16, THA15]. Amazon [CMAA13, MLF^{+}17]. ambassador [PMM^{+}17]. ambiguity [vLvHvT12]. ambitions [Kid13]. amendment [CL16a]. America [BDSB16, Fab16a, FMS17]. American [Jet18, KBBR16, LR15, PD15, SD18, SCOS17]. among [BK16, BOF^{+}15, Bre13, CLS16, CKC^{+}13, CvPH^{+}18, DFSB15, FC19a, HM18, JBJ12, KVA18, LCH^{+}13, OPGL16, Sch15b, SCMS^{+}18, WBC14]. amount [LBHS16]. amounts [OIS^{+}18]. AMP [GSB^{+}17]. Ampat [ALM17, GAW^{+}13, VAP^{+}10]. anadromous [SMQ^{+}16]. analyses [BBN^{+}13, GAW^{+}13]. Analysing [San18, San19, SI18, RPG^{+}14, CF17]. Analysis [BAO^{+}12, BH14a, DRG^{+}16, MPMGGAHR18, NBM15, PL12, Asa12, AdV18, ATDP16, BDIS18, BBA15, BDG^{+}14, BLR15, BCE^{+}17, BDP19, BHM16, CMFG15, CP13, CS14, CCD^{+}13, CDO16, CCS16, CGM^{+}18, CC18, DFP^{+}16, DSP^{+}18, DB19a, DL19a, DRH^{+}18, DL19b, DV17, ERD^{+}13, EG15, FvRA11, FFH18b, FP19, FLH11, FVH18, dHJTG17, GBC17, GW11, GRLJ15, HF16, HHS16, Har10, HPP^{+}17a, HCJ15, IH18, IMA16, JTP18, JHW13, Jon13, JGMS16, KFB19, Kru16, KNY14, LY14, LBD^{+}18, MVRG13, MPF12, Man17, MPPSdVSOHR14, MO13, MUHH18, MA15, MCBM18, MLHP15, NACS17, NKBS17, NKCR13, OO19, PDC10, PHM16, PML10, PVK18, PPCM16, PFP17, PMA^{+}17, PP19, PCLE17, Qiu13, QG17, RWJ18, RJ18a, RL19, RPC^{+}17, RRC^{+}17, SCV^{+}15, SGNM19, Seu16, SK18, SG15b, SC16, Smi17, SSG^{+}14, SES^{+}19]. analysis [SG12, TGT^{+}19, TRM19, Tam15, TTM18a, Tei18, TINL^{+}14, THNR12, TII^{+}19, TRC^{+}17, TCC^{+}19, VRT13, VGA15, VBMM17, WM18a, WW19, Won19, dlTCFB^{+}17, KSS17, MÁSSI13, PCC15, SS16, TSW19, Tyn17, Tyn18, WSGID17]. Analytic [Mor17]. analytical [Zag18]. analyzing [dlFFVRC16, HDJD15, CR11, KS18b]. anatomy [Bes19]. ancestral [Cap10]. anchor [DBG^{+}16]. anchored [MBvZ15, MKB17]. Anchors [DBG^{+}16]. Anchovy [KSC^{+}19, MCAM13, RRC^{+}17, SCMS15, Sch12]. ancient [KB15]. Andalusia [MJRdC16, PSP10]. Andrew [Rah19]. Angeles [SSG12]. Angler [CACL18, CJB15, DS17b]. anglerfish [MFF^{+}18]. anglers [FLT^{+}19, HHA17, LL15, SPB19, SSG12, GCHO19, BJO19, VPL^{+}13]. angling [BJO19, GCHO19, HSRH16, RMR^{+}17]. Angola [SC10]. Anguilla [FLWG14, SBC^{+}16]. Anilao [BPYC19]. animal [BR13b, BR14, DHSC14, NSB18b, PLL17]. Annapolis [Rah19]. annual [Das16, SL15b, SL16, SC11]. Anping [CL16a]. Antarctic [CHKG16, GRLJ15, JBPBA16, LB18, NVK^{+}17, Sco18]. Antarctica [NFHJ16, SRSN17, TRCR^{+}19]. Anthropogenic [RHP^{+}17, ECR^{+}16, WWW10]. anticipate [Far19]. Antifouling [McN18].

Aotearoa [DMKL18, LLA^{+}19]. apex [BSP19]. Apo [HHG10]. Apparent [Mul19, HJN16]. appeal [BTO^{+}17]. Appendix [PC11]. appetite [CAGH16]. Applicability [YCA^{+}12, LPM^{+}14, MJS16]. Application [Cha11, MMC^{+}14b, RCF^{+}12, Sch10a, BBR^{+}18, BvRH11, BMGF^{+}17, CKdS14, Das16, Dva19, DSWG13, FÓ12, dlFFVRC16, GNIL14, GMM^{+}13, GSDB14, HB19, HDTS14, HTH13, KPF17, Man10, ND17, PT10, RRA^{+}10, SCMS15, STG11, SSB^{+}18, SW10, SRGD18, SBS^{+}13, SH19, SCS14, Tei18, Thu10, THRS15, ZZ15b, CCD^{+}13, CdL17, MC14, PVK18]. applications [PGB^{+}17, VRA^{+}12]. applied [AYR^{+}14, ASWM18, MM14, SGMG18]. Applying [FMM13, GRW^{+}16, DGC^{+}17, Sea17b, Goo17]. Approach [GNIL14, AWJ19, ASP16, ACA^{+}13, ACP^{+}14, BBE19, Bla17, BG15, BDP19, BC13, BB19d, CAGH16, Chu10, CGM^{+}18, CHL^{+}19, CSMR18, DMdJ^{+}15, Dix16, DDS^{+}11, DFK13, EGP^{+}18, ECRL17, FGW17, FR16, FW19, FVT19, GFMG^{+}14, GJ15, GMD15, GHM^{+}18, GWWH17, GMP^{+}16, GHO^{+}17, Gri17, GNW19, GCD^{+}11, HHL^{+}15, HTT10, HR16, HP13, HPB^{+}15, HDJD15, HNVZ12, KS12b, KCKA^{+}18, KS18b, Kir19, LLS^{+}15, LDF^{+}18, LLN17, LJL^{+}14, MM18, MB10, Mat14, MJRdC16, MCHO11, MBBDS19, MSC18, MYA^{+}18, Muk15, MM14, NMvPNL15, NBF^{+}18, ÖGB^{+}10, ODRRM12, PH16, Par15, PIH^{+}10, PWR17, PVVB17, PGP18, PVA^{+}18, PPR^{+}15, PLP14, RMRN^{+}16a, RYD18, RUH^{+}14, STIC19, SCV^{+}15, SBC^{+}15b, Sea17a, SSLR14, SS15a, Son15, SJ19c, SA18, SLDHS18, TTJ18, TJJAM19, WS14, WK14, WTD^{+}15, WSA^{+}15, WGMG^{+}16]. approach [WB18, Wen17, WTA^{+}16, YFC13, YCHM14, ZS18, dBMA13, dVM12, vH15, DTMJH16, DTM18, FSWFG13, KAPF^{+}19, OPMB12, PGP^{+}19, RMRN^{+}16b]. approaches [BIPS13, BMD^{+}14, BS10, Boh18, BD13, CACL18, CKY^{+}13, CGF14, FSCJM19, JDM^{+}19, LM14, LRS^{+}13, LLX^{+}15, MDM16, ODF^{+}16, Rei14, RSS14, SSBL^{+}15, TL16, WS14, WNB^{+}15]. Approaching [Has18]. Appropriate [Wes18, WMSB15, KBF12]. April [Ano15p, Ano16n, Ano17n, Ano18p, Ano19w]. Aquaculture [GAC^{+}19, MNLHRMP17, PF12a, PZ17b, AG16, AA17, ARS16, BBAP^{+}17, Bra12, BBCPC11, Car19, CQ14, CSP^{+}14, CGM^{+}18, Cra19, CKFLO19, DAGG16, DJTK17, DCF^{+}19, EZBL19, ET19, EOC15, GB15, GKG16, GVSZ19, GGVRVL16, GGSRVL17, GSG^{+}15, GHO^{+}17, HBL14, HDK^{+}15, HBK17, HAC19, HCC^{+}19, HUHA18a, HUHA18b, HSR^{+}18, KYS10, KP13, KBK19, KEM^{+}18, KLM^{+}13, KN17, KGS^{+}19, LLS^{+}15, LY14, LJS^{+}18, Lim16, MHC19, MF19c, MFT19, MS19, ML19, MD15, MG16, NF17, NHFV13, Ola18, OO17a, OO17b, PZ17a, PWN18, Ren18, RDBB17, SS19a, SG15b, SBSK19, SS19c, SF19, SNTL18, SLBC19, TSK18, TBB12, TRC^{+}17, VM18, WB18, WKB15, WCK^{+}17, YBB^{+}17, YBD^{+}19, ZMZ18, vOHO^{+}17, vdBKB^{+}17]. Aquarium [HTF^{+}15, MDE14, MKSS18, PKR^{+}17]. Aquatic [GRW^{+}16, Coc15, LCC14, WLV15]. arawana [MLF^{+}17]. arbitrage [SL16]. Arc [FMMA^{+}15]. archaeology [Hua14]. Archipelago [Cro15, MFRvP16, GFC^{+}15, PGC16, dOSL18, YYN^{+}15]. Archipelagos [SSM15]. architecture [FOG^{+}11]. Arctic [GDD^{+}19, PH16, PS19, Ray19, SDHS19, VHRM17, ZH18, APY^{+}17, BG16, Cha19, DA11, DJS17, De 17, ERD^{+}13, FKMV16, FC19b, GDN19, GDD^{+}17, HTD^{+}18, Ho10a, KBB^{+}15, KPF12, Kim15, KFM17, MGGS18, NRB^{+}18, NKAP13, PDR11, Pla19, RRG^{+}12, REA^{+}14, RRCS^{+}19, RTM19a, SDP^{+}16, SGMG18, SS15b, SDHS19, SH16, TSH^{+}19, TRM19, TL16, Weg15, Xin19, You16b]. Arctophoca [CKBS14]. Area [CG19, RAZ16, SM18b, TTM18a, ALSC13, BMPV10, BD14, BGS13, Bur16, CL18, CSS15, CvPH^{+}18, DMKL18, DD13, DJM11, De 13, De 18, DJR^{+}18, Eli14, EMP^{+}14, FSM^{+}12, GBF^{+}10, GA19, GAW^{+}13, HF16, HBA^{+}14, HPP^{+}17a, HHG10, HDJD15, Hun13, KAF^{+}15, KRM14, KR14, KLMD15, LM16, LF17, LJ16, MAR^{+}15, MMJM16, MLF14, MCH18, MBTV14, Ngo18, ND17, RR17, RWK15, RRRM^{+}15, RL14, SMBL13, SRA15, SWG18, SSG^{+}14, SSG12, SLDHS18, SRGGVL14, TBKH13, TTM18b, TCCV16, WK17, WP12, Wis14, YYN^{+}15, ZZ16, DLM^{+}18, ARJ18, CDR^{+}16, DFP^{+}16, GLJdlTC14, Han19, HMS^{+}17, HBP16, KR18, MVRG13, MBCR16, MSTB19, MKMS12, ND17, Oya18, RMH^{+}15, RBS^{+}19, SLDHS18, TRCR^{+}19, Vau17, VGG12, WS15a, WFK^{+}12b]. area-based [De 18, DJR^{+}18, SLDHS18]. Areas [CS19, CD15, DBB^{+}14, HCF^{+}16, LAP^{+}12, LH16b, MSA^{+}18, AFR^{+}19, AdSC11, ÁFFSCF17, AVH10, Asa12, AMR14, ADL^{+}18, BF18, BMF^{+}16, BY16, BDP^{+}17, BAHCG14, BBE^{+}16, CMA15, CG19, CCC^{+}14, CBG^{+}17, CPFS^{+}13, CHDJ^{+}19, CSMR18, DDL^{+}17, Dva19, EKSZ14, FP19, FJA^{+}14, GCP19, GPB^{+}18, GWM^{+}10, GPA18, GMP^{+}14, Gor19, GAP12, GOdFC16, GCG^{+}10, GCD^{+}11, HTD^{+}18, HM18, HCF^{+}17, HR14, Hou14, HBH^{+}19, JQD13, JDQV13, JD16, JLCB16, KMD15, Kat16, KBF12, KBGF12, KFBG13, KJWP14, KWT17, KH15, Lea19, LMH^{+}12, LB18, LSC14, LEFC19, MMM19, MD16, MKR12, MMJM15, MWR^{+}17, MJS16, MSZS^{+}11, MB15, MSH13, MC14, MS14, NCH^{+}13, NWD18, NKBS17, OBJ^{+}12, PGR14, PCK^{+}19, PTP13, PVS^{+}19, PBJ^{+}14, RLKH19, RAF14, RBWL11, RUH^{+}14, RRRM^{+}16, Ros18, RL14, Sen10, SSLR14, SL15a]. areas [SBC^{+}16, dOSL18, SBS^{+}13, SFC^{+}15b, ST13, SJT10, Tan17, TTP12, Thu10, TDMN19, THK^{+}18, TAGP15, WH18, WGB^{+}16, WAG^{+}15, XJT18, YTTK10, YYN^{+}15, YSJU^{+}17, ZMZ18, dAdOS17, dGB10, dO13, AAT12, DAB^{+}14, FSS17, GR16, LV18, PGL^{+}16, RFG^{+}14a, RS17]. Arena [UEMRG^{+}18, Stø19]. aren’t [GNN^{+}17]. Argentina [BBG^{+}16, GCMA^{+}12, PRLCL12, VC17]. Argopecten [OTOL17]. argument [WLV15]. arising [HO12, Tei18]. arrangements [Bes12, CBF^{+}12, FE12, LP16, MF19b, PGPS10, dVvT11]. array [WWM^{+}19]. art [AMdB12, QSHI17]. Article [GDN19, Hee14, Kep14, RJ18a, She14, KSC^{+}15]. Artificial [RR14, TC15a]. Artisanal [OCW^{+}17, PSP10, TJ13, BLC16, Cam15, CGM^{+}18, CdL17, DPH^{+}15, DZ12, GFMG^{+}14, dlFFVRC16, JSD19, KNA10, KVMM10, LPU^{+}16, MG19, MFRvP16, MSC18, NSB^{+}18a, NMvPNL15, OPGL16, OCG14, OCG15, RBTR^{+}19, RWBF19, STM16, San15, Sch15b, SBMD13, SCR^{+}19, dAdOS17, DRMA18, Mat16]. ASC [LJT19]. Ascension [RBS^{+}19]. ascidian [SHS^{+}14]. Asia [BHI^{+}17, Chu10, CL16b, DHM^{+}17, GMP^{+}16, Kim10, PDT^{+}14, Pom12, PPCM16, YSJU^{+}17]. AsiaFish [TRC^{+}17]. Asian [TC13, PGP^{+}19]. aspects [AARK15, ERB14, Gil11a, MBT^{+}19, THM15, vD16]. aspiration [Gor19]. aspirations [TJ13, SFG15]. assertion [JRP17]. assess [BTC18, DGC^{+}17, FMGV16, GGSRVL17, KTD^{+}15, LR15, OC17, PCK^{+}19, TTM18b, TPKL16]. assessed [WGB^{+}16]. Assessing [ATDP16, BG15, BHS19, BC13, BHT^{+}18, De 16, EAG18, FWK16, GRMAC18, GA19, GBF^{+}12, JS18, JWM^{+}16, KV18, LJS^{+}18, MSdAM16, MLA16, PGR14, RDP^{+}18, RCdLS13, RMR^{+}17, SLGC^{+}18, SDP^{+}16, SFC^{+}15b, TCCV16, ASP16, BPF^{+}16, BSP^{+}15, DSH^{+}19, LPS12, RRRM^{+}16, BH14a, Cha17b, DWCS18, DTM18, FVM14, KCM^{+}18, KCCO18, MLF^{+}17, SPS14, TSK18, TMR^{+}14, TE16]. Assessment [CLDP17, ES18, KBK19, KVH14, MML^{+}19, MHM16, OHQ^{+}16, Pol12a, QSICTdV19, vdBAMJT19, AGSP^{+}13, Ard18, AAL^{+}18, BPYC19, BS18, BRR^{+}12, BRR15, CLV^{+}16, CBG14, COS14, CKC^{+}13, CPB19, DRPSA17, DHC11, DMdJ^{+}15, DDP16, DPB17, EBS18, Fen17, FP12, FP13, GRLM15, GOK14, GPA18, GCM11, HGB^{+}19, HNVZ12, KC13a, KJH^{+}10, LM16, LEKB10, LJB^{+}14, MKX17, Mal16, MMJM15, MSB^{+}18, MFAA16, MMA16, NBF^{+}18, NKAP13, OGM^{+}14, PRG^{+}17, PWB19, RR14, RHP^{+}17, RR18, RLHM14, RFHEJK13, STIC19, SSCC19, SRSN17, SNIN17, SCMS^{+}18, SKS^{+}19, SGP^{+}17, TRM19, UNS^{+}18, VVS18, VRMF12, Vil10, Vir18, VGG12, ZZ15b, dLS10, AZAMG13, CCC^{+}13, DLM^{+}18, MHM16, VRB14, WFK^{+}12a]. assessments [BDF^{+}10, CMP^{+}15, CdLH^{+}18, CCK19, CPBQ19, DDP16, FVT19, HBA^{+}14, Kai19, KFB19, LCR16, NWW^{+}17, SV15, TMG^{+}18, WBGJ18, SWP13]. asset [Gla19]. assets [MH13, Por10]. assignment [EGHC16]. Assistance [OC17, CMH19, HBS^{+}19, MCY^{+}19, NPAR19, QZH^{+}18, RPA17, MHC19]. associated [BAO^{+}12, GBF^{+}12, HGB^{+}19, JHW13, MCH18, MGG^{+}19, NWW^{+}17, RHP^{+}17, SHS^{+}14, SFC^{+}15b, VH16, WHB^{+}18, WP12]. Association [Rei10]. associations [RPdO^{+}19, MV17]. assurance [Che11]. assured [Han14a]. Asturias [dlFFVRC16, RGGFA17]. Asymmetric [BB19d, JFA17]. Asymptotic [HTF^{+}15]. at-sea [SSW^{+}18, SF16]. at-vessel [BW19]. Atlantic [ÁFFSCF17, CLF^{+}19, DPH^{+}15, DFV15, FMMA^{+}15, GPB^{+}18, HD11a, KAPF^{+}19, Rei10, dOSL18, dVM14, ADGGVL12, Asa12, BK16, BGC^{+}17, BHM16, BWRT17, BMGH11, CETQ16, CRO12, DB19b, DT10, Dru10, FMGV16, FBAK14, GBW16, GFC^{+}15, GTL^{+}18, GSC^{+}16, HJN16, HD11b, HSB^{+}16, JKS16, JFK18, LLH^{+}14, LBHS16, MJRdC16, MG14, MPT13, MSG11, MSGM^{+}12, NAYSN18, OHQ^{+}16, PCC15, PGC16, PAD^{+}17, RWG^{+}19, SGMG18, SJ10, TEF^{+}15, TJJAM19, Web11, Yia18, ZMM^{+}14, dS17a]. atlas [MV19]. attachment [UA13]. attacks [AT14, TGT^{+}19]. Attainment [MCHO11]. attempt [Gre15, Vat17]. attempted [SHS^{+}14]. attention [ZPT^{+}16]. Attitudes [AWH13, DS17b, MCH18, ROD15, AMdlCMS^{+}18, ATB^{+}18, Car13b, CST10, EMP^{+}14, FV10, GGADSR17, GAKP^{+}19, HNC14, HSR^{+}18, KR14, KGS^{+}19, MMJM16, MK15, MPCB14, MGG^{+}19, NY13, PPOM16, Roy16, SCD^{+}17, SPB19, TMG15, Alf17, JBJ12, WK19]. attractive [MA15, ZMZ16]. attributes [SGLY13]. auction [IT12, SGZI19, ZM15]. auctioned [Mox12]. auctions [IT15, TI13]. August [Rei10, Ano15m, Ano16r, Ano17p, Ano18q, Ano19a]. Australia [BS10, BJS^{+}16, BvPS15, CdLR^{+}15, CKHJ^{+}17, CvPH^{+}18, DD13, DDP16, DCC19, DDS^{+}11, EGHC17b, FBM18, GW15, Har15, HR17, Hob14, Hun13, KC13a, KBF12, KBGF12, KH13, Lal17, LLS^{+}15, LSMW10, MAR^{+}15, MPPB10, MC19, NBFO18, PR13, PRS10, RCdLS13, RG19, RHP^{+}17, RBB^{+}16, RTM^{+}19b, SSCC19, SBC^{+}14, TK13, Vin13, VSS^{+}15, VBMM17, WM18b]. Australian [BP13, Bro10, BSP^{+}15, CHvKM15, DK15, DDvP^{+}16, ENWL19, FLT^{+}19, GDD11, GDD13, HC19, Har19, Hea13, HCP^{+}17, Kay15, KKWS15, MTP^{+}15, ML19, ODF^{+}16, SSD^{+}16, SHM12, WNB^{+}15]. australis [CKBS14, FSS17]. authority [Mso16, Ard18, Bro15, Jae16, LJL^{+}14]. authors [SGW^{+}15]. availability [HTA^{+}15a]. available [Mul19]. avenue [HSRH16]. averages [CG18]. Averting [CCTG^{+}13]. Avery [SJ19c]. Avery-Gomm [SJ19c]. avoid [HdF^{+}19, SNDB^{+}16]. avoidable [SCC10]. avoidance [BSKH17, Eli14]. avoided [HBH^{+}19]. awareness [EAG18, GFC^{+}15, KSK^{+}19, MDK^{+}16]. away [DBG^{+}16, THA15]. Azorean [NM19]. Azores [ALSC13, DPS16, TdCdSR16].

B [KS18a, WMVR18]. Babel [VvDM10]. Back [SH14b, SS16]. background [GCW^{+}18, YTTK10]. backs [GBC17]. bad [EB16]. bag [KSC^{+}15, LL15]. bags [SSV18]. Bahamas [Wis14]. Bahia [NACS17, CST10]. bait [RG19]. baitfish [Gil11a]. Baja [RROR12, SARF16]. balance [Bes10, Goo17, PHC14, SC13a, SS19a, CR10]. Balanced [MFB^{+}14, PFH16]. Balancing [DTN10, NJ18, RFKEJ^{+}15, KBBR16, WBH^{+}15, BTOM13]. Balearic [DTN10]. Baleia [MVRG13]. Ballast [SDLT15, DMC^{+}17, FSV^{+}16, LWS19, RRCS^{+}19, SLDHS18, VH16, WCS18, LCC14]. ballast-water [FSV^{+}16, WCS18]. Balloon [Bla15]. Baltic [WPJ10, AAC^{+}13, Bac18, Boh18, EKV12, Elo10, Eng18, GE17, Grö18, Gus18a, HK10, HGG^{+}18, HLRO19, IH18, JKK13, Jet18, JR14, Kee18, KS18b, KR18, Lan18, LG19, LE12, LGM14, LSAL19, MOE11, MKU19, NALO19, PSH18, RWG^{+}19, RJ18a, RJ18b, Rin18, SFC^{+}15a, SJT10, TK14, TPKL16, Tyn17, Tyn18, Val13, Var11, Wes18, WBA18]. bamboozled [PP19]. Ban [GBBB15, CGC14, Eli14, GRLM15, HS19, LB17, MdMGV^{+}18, MMBK17, Nij17, PDC19, SH17, SS16]. Bangladesh [ATAM10, AOAM13, ACTG17, AF14, AA17, Deb18, IMA16, ISM^{+}17, PMA^{+}17, QG17, RM16, Roy16, SI18]. Bangladeshi [ISHP14]. bank [MSG11, GTL^{+}18]. Banking [DR10]. banks [SCD^{+}16, DR10, EE17]. banner [SJ15]. bans [FKWV19, Nij15]. Barbados [BBR^{+}18]. barcoding [ELAC19]. Barents [Han16, Kno10, Kno11, OHH^{+}16, SS13]. Bargaining [KLMD15, PG19, WSWM17]. barnacle [CVP^{+}17, RGGFA17]. barnacles [SFCI14]. Barrier [DD13, DSM^{+}13, DDS^{+}11, FSB14, FAPP17, HMS^{+}17, LBN15, LSTD12, PRS10, WS15a]. Barriers [LK15, CHvKM15, FMD^{+}18, ISHP14, RPSS17, Rei19, RJ13, Tyn17, WCS18]. base [GRYA^{+}15, GSDB14, HTA^{+}15b]. Based [FSMG10, Man10, AMLK^{+}12, AH19, ABBD17, ABB14, ACM^{+}12, AAL17, AFCMET18, BAHD10, BPYC19, BK16, BCE^{+}17, Bro16a, BMGF^{+}17, CdLH^{+}18, CR10, CC13, CHKG16, CAGH16, Che12c, COG^{+}19, CSP^{+}14, CSS15, CWP16, CP10, DL19a, De 18, DAJM15, DFSB15, DJR^{+}18, DDS^{+}11, Dup14, EGGT12, EAA^{+}18, EGD18, ERCMD11, FÓ12, FLH11, FDS15, GGK17, GMD^{+}12, GRMAC18, Gib10, GAB^{+}17, GLJdlTC14, Han11, HGS12, HR17, HDM13, Hav13, HHG10, HW10, HvPN^{+}19, HCF^{+}16, HCF^{+}17, IOG18, Iri18, JHB^{+}18, KYS10, KARR^{+}19, KEG12, KCKA^{+}18, Kir19, KJH^{+}10, KRK13, KC13b, Lan15, LBK^{+}13, LR15, LCS15, LTM19, MD16, MRC17, MMdJ^{+}14, MSB^{+}18, MPCB14, MRG15, MHA^{+}17, MCHO11, MDR^{+}16, MMA16, NHA15, NBD15, Ols11, OOvdM^{+}11, ODRRM12, PL15, Pla19, PGPS10, PFF^{+}18, PM10, PT10, QSICTdV19, RvLvTH14, RWL14]. based [RAF14, RLM10, RLM12, RvLG^{+}15, RO13, STIC19, SLJ^{+}11, SLH^{+}13, SBC15a, SBC^{+}15b, SB14, SVL^{+}16, SPO^{+}18, SNIN17, SSLR14, Shi16, SS15a, SCK15, SW10, SGN^{+}17, SS19c, SBS^{+}13, SSBL^{+}15, SLDHS18, SSSO17, TLR^{+}10, TK13, THA15, TML^{+}19, TGBL11, VMMB^{+}19, VAW14, WSA^{+}15, WGMG^{+}16, YCA^{+}12, YP17, vdBAMJT19, DTMJH16, DTM18]. Basel [RM18]. baseline [DRG^{+}16, FSS17, KDS^{+}15, VBHC^{+}17]. baselines [RCD^{+}18, RJPS11]. basic [CAP15]. basis [JSD19, KPF12, RGWH18, RSMM19, Xin19]. Basque [AM15, MMA16]. bass [GCHO19, Mat14, WCCO18, MSG18]. Batangas [BPYC19]. Bathing [HTH13]. Bay [AFGRF^{+}17, BGA^{+}15, BSM^{+}14, DTJ10, GCP19, GA14, LU13, MMS^{+}14, MG14, PW16, SOUN16, SMW13, vdGMA^{+}13, FGPP18, OTOL17, AK15, CSS15, CKBS14, FJ12, GMM^{+}13, MRHP19, QG17, RDP^{+}18, RBW18, SMBL13, Sch15b, SSG^{+}14, TKAO^{+}19, VM18, vS15]. Bayesian [CGM^{+}18, NMvPNL15, vPLB^{+}13]. bays [Dva19]. BBNJ [BDB19, De 17, De 18, Lea19]. BC [Pin13]. BCLME [SC10]. be [ARGW14, CCC17, DDP16, FSS17, GMP^{+}14, HSH14, HBH^{+}19, Kim17, LVRM^{+}18, NS18, PZ17a, PS11b, SGMG18, SSC^{+}15, SJ19c, TGS18, dMAH^{+}18]. beach [CT16, DAJM15, DV17, GJA16, KWT17, MDK^{+}16, PNW18, RPG^{+}15, VLKN17]. bear [Mee11, PC11, Wen11]. bears [CMC^{+}13, PC11, PC13a]. Beaufort [APC16]. became [ØR18]. beche [PWN18]. beche-de-mer [PWN18]. become [Cha15, JSG^{+}18]. been [EZBL19, Lyn14, RUD^{+}12]. before [McN18, NY13, PDB^{+}13]. begins [DL19a]. behavior [ASB10, JFA17, MS14, MMDS12]. behavioral [SOUN16, SFDB18]. behaviors [BRM^{+}17, WS15b]. behaviour [ADH18, FFH18a, MWR^{+}17, MK15, MFRvP16, PRS10, SAGG13, SB18, WWH^{+}17, vHCT^{+}16]. behavioural [BvRH11, WSO^{+}14]. behaviours [NPS^{+}19]. behind [CGSP13, MKR12, MSS18, SHM13, KC13a]. Being [Rog15, AYA^{+}19, Car15, HHP17, HM18, JGG^{+}16, MD16, MKR12, Ngo18, PRS^{+}19, WPRA18]. Belgian [PHM16]. Belgium [LDVM10]. belief [CGM^{+}18]. Belize [CS19, CAGH16, MHA^{+}17, WSB19]. Belt [YC18]. beluga [CN11]. benefit [BDG^{+}14, CHDJ^{+}19, Dav10, DPV^{+}15, EGHC17a, HD16b, LH12, OC17, VKGN^{+}14b, WM18a]. Benefits [CETQ17, Hun13, AVH10, Ásg16, BBR^{+}18, BSH^{+}19, BSMS15, BAO^{+}12, BAHCG14, BDH^{+}18, CPKF19, CMMBV^{+}19, DHFL^{+}16, Don17, ECSE^{+}17, FSV^{+}16, GSC^{+}16, HLC^{+}10, HHS19, JAG16, KM15, LMH^{+}12, MD16, MCY^{+}19, NALO19, PWH^{+}16, RFG^{+}14a, Ros18, RDG^{+}13, SCMS^{+}18, SCC14, WSGID17, WPdR^{+}14, ZKR^{+}19, CCS14, Cha11, WCMHO18]. Bengal [QG17]. Bengis [Gla19]. Benguela [Nei15, SC10]. benthic [CCRT16, EGP^{+}18, RWG13, SNP^{+}18, SSG^{+}14]. Bering [ZH18, HDH^{+}15, HBH^{+}19, Man10, PAC^{+}19, RYD18, SSL16]. Bermuda [LRMM17]. beset [PHAC15]. best [BSKH17, DS10, Gil11b, HHL^{+}15, HSRH16, KFM17, PZ17a, PSV^{+}14, TFBS19, TC15a]. Better [BMD^{+}14, BF19, KEM^{+}18, MLHP15, Pas14, She15, dlTCFB^{+}17]. between [BFTS^{+}16, Ben12, Bes10, BDHSJ19, BRM^{+}17, BAHCG14, Bro11, BSP19, CR10, CMW10, CBS^{+}16, DSK^{+}19, DDV^{+}13, DZ12, Edd14, ERB14, GCP19, Gor19, HBS^{+}19, Hut19, JA10, JS18, KH12, Kir19, KGS^{+}19, KLMD15, LPU^{+}16, LWvdBL16, LH16a, LJT19, MP12, MJRdC16, Mee11, MC16, NMvP19, NHFV13, NM19, ØR18, ONF15, PL12, PHC14, PVS^{+}19, QG17, RST17, Ros17, San18, SS19a, SCKP19, Sev13, TÁB^{+}10, VBMM17, WS14, ZVP^{+}13, dVvT11, vH19, LE12]. Bevan [KSH^{+}15]. Beyond [EGHC17a, GDD11, OBH15, RCUR13, ACA^{+}13, ARJ18, BIPS13, BY16, BDP^{+}17, BAHCG14, CC17b, CHDJ^{+}19, De 18, DG14, EKSZ14, FDP18, FJA^{+}14, HTT10, HD16a, HR14, Hou14, Kay15, KTA^{+}13, Lea19, LGD^{+}19, MLF14, MFT19, MDNT19, PVS^{+}19, RMB^{+}17, RUH^{+}14, Sco16, TDMN19, WAG^{+}15, Mac14, PS19]. bias [BDF^{+}10, WWW10, ZP18]. Bibliometric [CC18]. bicirrhosum [MLF^{+}17]. Bidding [IT15]. Big [LG19, DK16, RHB17]. Bigeye [AHB18, Han12]. bigger [IW19]. Bight [BSP19]. Bijagós [Cro15]. bilateral [Ant16]. Bill [Gau10]. Billions [RD12]. binding [Sco16]. Bio [KCJ15, McL15, ERD^{+}13, PAA^{+}12, SLDHS18, SRA15]. bio-economic [ERD^{+}13, PAA^{+}12, SRA15]. bio-invasion [SLDHS18]. Bio-logging [KCJ15, McL15]. biocides [McN18]. biodiesel [LH12, Lin13]. Biodiversity [MSdAM16, RFL^{+}18, VLPC14, ARGW14, ACR18, BM13, BY16, BDP^{+}17, De 18, DJR^{+}18, DG14, FLWG11, FJA^{+}14, FGR18, HD16a, Has18, HDL^{+}13, JTP18, LGD^{+}19, LTM19, LSAL19, MSS19, MRHP19, MGJB^{+}17, MDNT19, NMJS17, RRA^{+}10, RUH^{+}14, RRMSSE16, RRRM^{+}16, RFHEJK13, SRGD18, SSG^{+}14, TPKL16, TDMN19, VGL^{+}17, WWH^{+}17, WAG^{+}15]. bioeconomic [CSMR18, OLF19, TII^{+}19]. Biogeographic [CMP^{+}15]. Biological [RRC^{+}17, DGC^{+}17, HLRO19, Lon11, MJS16, ONF15, Sco16, WNB^{+}15, DAB^{+}14, JGMS16]. biologically [GAP12, YYN^{+}15, YSJU^{+}17, DAB^{+}14]. biologists [Sch10b]. Biome [BH14a]. bioregional [Vin13]. biosecurity [DHC11, SS15a]. Biosphere [VAW14]. biotechnology [ABSB15, ABSGB18]. bird’s [BPBM17, GOHDB17, Han13b, MMOL15, RTD^{+}17]. bird’s-eye [BPBM17]. Biscay [BGA^{+}15, GMM^{+}13, MRHP19, MMS^{+}14, SMBL13].

Bissau [OY15]. bite [NY13]. bites [NY13]. bivalve [BBCPC11, OCW^{+}17, SF19]. black [NSB^{+}18a, ACG^{+}17, AvTT17, CBP19, CCRT16, GSB^{+}17, HP12, LLO16]. blacodes [HKR19]. blame [CKVO16]. bleaching [COG^{+}19]. blooms [BBAP^{+}17, MCB^{+}19, WPRA18]. Blue [DRGP19, How18, MM17b, AG16, ACTG17, BCM^{+}18, Cha17b, ESGC13, EMD^{+}18, GSR19, Had18, HC18, KARR^{+}19, KED^{+}18, Laf14, MNG^{+}18, Pau18, PCC15, RWG^{+}19, TEVN19, VMMB^{+}19, Vil12, VKGN^{+}14a, WSGM16, YYYL15, BVB18, CMMBV^{+}19, DGC14, KSWS18, LPR^{+}17, NRB^{+}18, RRMSSE16, SvdBH^{+}18, vdBAMJT19]. Bluefin [Pol12a, STM16, Dru10, FBAK14, LLH^{+}14, NIFA17, Web11, SH12]. blueprint [MBD13]. Bluff [YFR10, YFC13, YCHM14]. board [NPAR19, TS14, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19c, Ano19d, Ano19e, Ano19f]. Board [Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n]. Board/aims [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l]. Boat [SJW^{+}16, CC17a, HR17, Knu17, LL15, MPCB14, ST17, SCC10, Wea16]. boat-based [HR17]. boaters [VGL^{+}17]. boating [FSS14, GCK^{+}11]. boats [LH12, SWB^{+}14, SHM12]. Bocas [SJW^{+}16, SSM15]. Bodies [MHM16]. Body [HTF^{+}15]. Bohai [WZF^{+}18, PQLH15]. Bonaire [Thu10]. bonding [Por17]. bonds [CCS16]. Book [Rah19, Tec19]. boom [KS16, KS18a, SMH17, WMVR18]. boost [PMN^{+}15]. Border [Gus18a, SHS^{+}14]. borders [CBE^{+}19]. both [DYN^{+}19]. Bothnia [KARR^{+}19]. bottlenose [PRT10]. bottom [BOF^{+}15, CdLH^{+}18, Cho12, DPH^{+}15, FTG^{+}18, GOM17, MMdJ^{+}14, MMN18, MV17, MMDS12, SC13a, SMR^{+}15, SKT^{+}10, VKA^{+}19, WTD^{+}15, WAG^{+}15]. bottom-up [BOF^{+}15, MV17]. bottomfish [HB14]. boundaries [Árn17, Ásg16, DG14, ItDD^{+}18, Pol12b, vdBAMJT19]. boundary [Nag16, OY15, VGCF18, Yia18]. bounty [TLR^{+}12]. bowhead [RRG^{+}12]. bowl [ST17]. box [vD16]. BPA [BH14a]. Brannstrom [NGS16]. Brave [Cla15]. Brazil [ATB^{+}18, BBM^{+}17, DRMA18, EMP^{+}14, GLTB^{+}15, HIBG15, KNA10, LV18, MVRG13, MMM19, MSdAM16, MMN18, NACS17, PV16, RAG18, SCV^{+}15, San15, TTM18b, TJ13, ZVP^{+}13, dS17a]. Brazil. [FPG^{+}17]. Brazilian [BFTS^{+}16, DBD19, GC18, GMM15, JLCB16, KL14, SO13, SBMD13]. bread [AJ17]. Breaking [FVH18, BD10]. breeding [OBC^{+}18]. BREXIT [OR19, AWJ19, FDP19, McN18, OO19, PS18, SP19b]. BRIC [dVM10a]. Bridge [GGLP16]. Bridges [BSKH17]. Bridging [BAY^{+}19, RHW^{+}19, Sea17a, BAHCG14]. brief [RvLvTH14]. Bringing [KCM^{+}18, dLS10]. Bristol [AK15]. British [EE17, EP19, Gek10, GS17, GBF^{+}12, HES16, HCC^{+}19, O’R15, PS10, Pin15, Rei19, WEKH19]. broad [GCM11, YYN^{+}15]. broad-scale [GCM11]. broader [HR17, HvPN^{+}19]. brown [DHR16]. Brunswick [RBW18, KN17]. BSRBCC [Gus18a]. budget [GNW19]. Buen [CWP16]. Buenos [Bai12]. build [LRJ^{+}19]. Building [BSAT18, CvPH^{+}18, GRYA^{+}15, KNA10, LEM^{+}18, LBP^{+}19, TDI^{+}14, Tip16, ZQZ19, ABSB15, CMH19, Jet18, JLY18, LLF^{+}16, LEFC19, OPMB12, Sch15a, SH15, dVM10a, JSG^{+}18]. bulk [RPM13]. bullet [ARS16]. bumpy [CQ14]. buoy [LHJ^{+}15]. burden [ADHV16, HEJK10, Han12, HCBM15, LJ13, HK11]. Burrunan [FSS17]. Business [vdBKB^{+}17, Che12c, Cla15, LFGM15, SS15c]. businesses [CLDP17]. Busting [CMS19]. buy [CLDP17]. buyback [HSW17, QSMGVA18]. buybacks [HSW17, ZM15]. buyers [SGZI19]. by-catch [MRCG11, MFF^{+}18]. by-products [SNTL18]. Bycatch [Gil11b, Hua11, ATB^{+}19, BPYC19, BKKN10, BSKH17, GCMA^{+}12, GMM15, HKR19, HSB^{+}16, JG13, KBG^{+}18, KW14, Muk15, OBNH15, PB13, PMN^{+}15, PBSG12, SJP17, SRGD18, WHB^{+}18, DBE^{+}16]. bycatches [ES18, WLV15].

C [KS18a, WMVR18]. C. [Kep14]. ca [vS15, HvD13]. Cabo [SGNM19]. Cadastralizing [PS11b]. cadets [ZZZ^{+}17]. Cádiz [RRC^{+}17]. cage [ADH18, MJQMVÁ17]. cage-dive [ADH18]. CAGs [FBP^{+}18]. Cahuita [HGS13]. Caicos [RNS19]. cake [TN15]. California [CST10, CSCBR10, MLTBMB14, RROR12, RRAH13, SARF16, AMLK^{+}12, AFMCMP12, BSSS17, CJB15, HLM^{+}13, HPPH16, JTP18, MKMS12, RCF^{+}12, SC13a, SPB19, SSG12, VGL^{+}17, WSA^{+}15]. californianus [CJB15]. call [Had18, MF12]. calls [EFK^{+}19, RvLvTH14]. calming [Sea17a]. Cameroon [Bes19]. camp [KHT17]. campaigns [WMWH18, CW15]. Campeche [UEMRG^{+}18]. Can [BAT^{+}18, FSS17, FJA^{+}14, GB15, GKG16, GGK17, GMP^{+}14, HPP^{+}17b, HSH14, HSW17, HTA^{+}15b, PKOF18, RM18, VSCH16, WSWM17, ARGW14, BF19, DDP16, DPS16, DMB^{+}18, Gil14, He18, JOE^{+}19, JE11, KVA18, Kee18, KWT17, LPR^{+}17, MMdJ^{+}14, MWR^{+}17, PZ17a, RM17, She15, SSC^{+}15, SM18b, SJ19c, SOS12, TCS^{+}18, TBB^{+}10, ADH18, HD16a, Lew16, ÖGB^{+}10, PS11b, SNDB^{+}16, SGMG18]. Canada [WFK^{+}12b, BFO^{+}16, BAA10, BBN^{+}13, BKHA^{+}18, BWRT17, DHFL^{+}16, DA11, DB19b, DN12, DJS17, DS18, Doy15, Dup14, FC19b, GW10, GLW15, KR14, LRJ^{+}19, LGD^{+}19, Liv10, LM11, MRSL16, MF10, MS19, MMDS12, MAAB15, PDR11, RBW18, RBWL11, RTM19a, SDC13, SDLT15, Syb15, TSH^{+}19, TÁSLC19, WBB^{+}15, WTD^{+}15, WH18, WB18, vdPLMS16].

Canadian [Cha19, PBW17, APC16, BAO^{+}12, BR13b, BR14, DHSC14, FMN15, HTD^{+}18, HAW19, MI15, PBW16, TL16, WRP^{+}10]. Canary [CdL17]. candidate [KJWP14, SPRH19]. candidates [MMBK17]. canned [BSH^{+}19, Cam16, dHJTG17, LHN18, SCOS17]. cannery [SK15]. cannot [dMAH^{+}18, vHHB14]. Cantabrian [AdV18]. Canute [GNN^{+}17]. canyon [SMBL13]. Capabilities [TPKL16]. capability [SHS^{+}14]. capacities [PL12]. Capacity [SHS^{+}14, ABSB15, AMdB12, ABB14, BDB19, BSAT18, BBCPC11, CMH19, CT16, DA11, DSH^{+}19, GPA18, HWY^{+}18, HPW^{+}18, HC10, Jet18, JCD17, KNA10, KFB19, LEFC19, MOR15, NBFO18, PHC14, RYJ^{+}17, SLJ^{+}11, SBSK19, TR18, WHMG^{+}16, ZSH18, OPMB12]. capacity-limited [HPW^{+}18]. Capbreton [SMBL13]. Cape [DB19a, Nor15]. capital [BdKvT19, Dra15, FGW17, KS12b, LJB^{+}14, PCC15, Por17, VG15]. capitalist [STM16]. capitalization [Dav15]. captains’ [CBL18]. capture [BYK15a, BVB18, Gar12, Gra10, HWL^{+}19, LY14, SW10, ZZ17c, NSB^{+}18a]. captured [BW19, VRMF12]. carbon [AG16, ACTG17, GNN^{+}17, GZW^{+}19a, GZW^{+}19b, HK11, MSBB14, MM15a, Man17, PR19, RPSS17, SGMWE14, VMMB^{+}19, WML17, WSGM16, ZEP^{+}19, ZZ17c]. care [JRS^{+}16]. cares [FPG^{+}17, PPOM16]. cargo [GCP19]. Caribbean [FMM13, AFCMET18, DFP15, ELAC19, FMS17, FLWG11, KAF^{+}15, KDS^{+}15, PAA^{+}15, RBC^{+}13, RGC^{+}19, SBC^{+}16, TCCV16]. CARICOM [Has18]. Carolina [BHG11, MRCG11]. carriage [KP17]. carriers [RPM13]. carrying [BBCPC11, CT16, GPA18, HWY^{+}18, JCD17, KFB19]. cart [PDB^{+}13]. Cartagena [ASP16]. cartography [KP17]. case [ASP16, AMLK^{+}12, ASBAM13, ALMS14, ADL^{+}18, BAA10, BPB19, BJS^{+}16, Ben17b, BDG^{+}14, BPMY18, Bol17, BB19b, BB19c, BSMS15, BAMAH17, BM15, BDH^{+}18, Bro18, BAY^{+}19, BHG11, CMFG15, CSH17, CMAA13, CBE^{+}19, CL16a, CL18, CL16b, CSP^{+}14, CE13, Cro15, CKFLO19, DFP^{+}16, DJS17, DP17, DR14, Don17, DT10, FD11, FSS17, Fla10, FWHS17, FLWG14, FDGM16, FBAK14, GLVLGG19, dlFFVRC16, GGVRVL16, GGSRVL17, GGD16, GLTB^{+}15, GHM^{+}18, GA14, Gla19, GCMA^{+}12, Gor19, GSA10, GKSJ11, GCK^{+}11, GFA^{+}18, Gus18a, HvD13, HF16, HHL^{+}18, HAH^{+}16, HDK^{+}15, HGS12, HHP17, HBA^{+}14, HVB^{+}17, HGDB10, HMH18, HdF^{+}19, HP12, HK18, IH18, IW19, IHF15, JCD17, JSG^{+}18, JLQ16, KNA10, Kat16, KLC^{+}15, KBG^{+}18, KEG12, KH10, KS19, KJH^{+}10, KFM17, KHP14, KV18, LLS^{+}15, LQC10, LJB^{+}14, LC19]. case [LU13, LEFC19, LT19, LSAL19, MKX17, MOE11, MCCC16, MSS19, MCMB^{+}18, MB10, MMOL15, MKMS12, Mat14, MHK^{+}10, MBB18, MPPSdVSOHR14, MG14, Mso16, MCAM13, MM14, MSG11, MSGM^{+}12, Nik19, OP15b, OCG14, OCW^{+}17, PRT10, PIH^{+}10, PSH18, PBSG12, PW16, Pol12b, PV16, RPdO^{+}19, RBW18, Rei19, Ren18, RNS19, RNCGB15, RJPS11, RBS^{+}15, SC19, SLJ^{+}11, SBC15a, San18, SOCA15, Sea17b, Sev13, SPMW12, SJ14, Smi13, SGZI19, SAdC15, SY18, SRGGVL14, TFBS19, TvN13, TEF^{+}15, TMG15, TMDSF19, Tos11, THK^{+}18, VDB16, VGL^{+}17, Vil12, WSB19, WL18, WM17, WBC14, WM18b, WSGID17, Wen16, Wen17, WTA^{+}16, WPRA18, Wri15, WvLvK12, WSGM16, XDG^{+}16, YPS13, Zer18, ZTAO13, ZZZ^{+}17, ZLSL17, dMAH^{+}18, dO13, dVM10b, vdBKB^{+}17, CVFP^{+}12, CPRP19, GGK17, KARR^{+}19, LSC14, PW15, PHC14, SOUN16]. Case [SM18c, SMW13, Yos17]. case-study [Cro15, HdF^{+}19, WTA^{+}16]. Caspian [ES18, MGG^{+}19]. Cassurubá [NACS17]. Casting [MLHP15]. Casualties [RPM13, RMJ12]. Catalan [VLKN17]. Catalina [Mat16]. Catch [All14, DPS16, HB14, JS12, JS13, KBBR16, MMC^{+}14a, SS19c, WBH^{+}15, AGB13, ACR18, AHB18, BIPS13, BKKN10, BMAG17, CL15, CBP19, CMCMHP13, CKdS14, CS13, CMB^{+}15, DES10, DDV^{+}13, FC19a, Fla13, Gar12, GBF^{+}12, HTA^{+}15a, HK18, HSB^{+}16, JHN^{+}19, Lyn14, Mac14, MKB17, MJQMVÁ17, MC19, MRCG11, MKSS18, MFF^{+}18, MA15, OIS^{+}18, Pan14, PDS15, PZ17a, PS12, PS14, RLAT19, RWG13, SRA15, Sev14, SC11, SL15b, SL16, TWA^{+}15, VAP^{+}10, ZPT^{+}16, vHCT^{+}16, PCS14, Sev14]. Catch-based [SS19c]. Catch-quota [WBH^{+}15, KBBR16]. catchability [SV15]. catches [Bor18, CGP^{+}18, ES18, EYN^{+}15, JLW17, LBHS16, NMvP19, SBM^{+}17, WDK^{+}13]. Catching [WSGID17, Che11, Sev14, ZPT^{+}16]. categories [BPYC19, ND17]. categorization [HJN16]. Categorizing [BDP19, HdF^{+}19]. Caught [GOO^{+}19, BKKN10, CBP19, GTL^{+}18, STM16, SGLY13]. causal [VvD10]. causative [LEG19]. caused [DB19b]. causes [DL19b, MMS^{+}14, RGWH18, THM15]. CBD [JGMS16]. CBFM [YP17]. CBSS [KR18]. CCAMLR [ND17]. CCS [MKX17]. ceasing [dMAH^{+}18]. Cedeira [FVM14]. Celtic [CE17]. census [TSW19]. Central [SDB14, ZH18, EYN^{+}15, Gil11a, Han12, HCC^{+}19, RNEBG19, SDP^{+}16, ZMZ18, BDvTvKT17, Hav10, Hav13, MBM14, NW10, PH16, Ray19, SDHS19, VHRM17, WFK^{+}12a, YBRBB16, ZH18]. centralised [HHG10]. Centralization [BGG19]. centre [Dup14]. century [BGG19, SH10, WT18, HC19]. certainty [Ásg16]. certification [BHI^{+}17, BTOM13, DK16, GSA10, KNY14, LSK18, MM15a, PRLCL12, PRPLBLC12, SJ15, SSSO17, TM13, WW17, MHM16]. certified [AGSP^{+}13, BB19b, BB19c, FP12, FP13, JGG^{+}16, OHQ^{+}16]. Cetacean [ACG^{+}17, MRHP19, ATB^{+}19, DS10, ItDD^{+}18, LLZ^{+}19, RJPS11, DBE^{+}16]. cetaceans [ASB10, BRR15, REA^{+}14, RBJ^{+}11]. CFP [Pas14, CRO12]. Chagos [Har14, She14, DJM11]. chain [DPB17, GKG16, HHS16, HHS19, KKG16, Lim16, MM17a, NJ18, NAYSN18, PMA^{+}17, PP19, PCLE17, RPC^{+}17, SP19b, WSC19, vdVD16, dOSL18]. chains [Fab16a, FDP18, RV16b, TCN^{+}13]. challenge [AH19, EG15, Maz17, MHM16, SBP^{+}19, dVMdC^{+}19]. Challenges [Bod17, DBD19, FJ12, JHB^{+}18, Kai19, KMD15, MKX17, PS19, RS17, SD13, SOCA15, VC15, WM18b, Wit15b, vAS19, AF14, ATB^{+}19, BS18, Bro16b, CRSM^{+}17, CSF^{+}19, CWP16, De 18, DGC14, DK16, FEF^{+}12, GH18, Gla19, Han16, Har15, Hua11, HDTS14, Jae16, KLC^{+}15, KAPF^{+}19, KW18, LPC18, LJ16, LWS19, LZ16, MRC17, MV16, MUHH18, MMC^{+}14b, MNG^{+}18, NWW^{+}17, NFHJ16, OBJ^{+}12, Pan14, PRG^{+}17, RMRN^{+}16a, Ren18, RPA17, SHS^{+}14, STJ^{+}12, SFG15, SS19c, SBT16, TGBL11, VAW14, VKGN^{+}14b, WRP^{+}10, Wis14, WBSW13, YBD^{+}19, ZZ17c, dVvT11, ATAM10, BB19a, EBR^{+}18, FOO^{+}15, GAFAV16, GdV16, HCC^{+}19, LF17, PKR^{+}17, RFG^{+}14b, SSSO17, SH10, Wit15a, YY19]. Challenging [HD10a, HID12]. Cham [Ngo18]. chances [MHC19]. Chandler [Tec19]. change [AOAM13, ACTG17, ACR18, BAHD10, BLC16, BvPS15, BF19, BSM^{+}14, CMVLGG14, CRCD13, CJW^{+}16, CPBQ19, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, DCHC17, DCY^{+}19, DSH^{+}19, FGW15, FTG^{+}18, FMB^{+}18, FC19b, FLWG14, GM19, Gil14, Gra10, Gug19, HWO^{+}18, HIBG15, Hav13, HDL^{+}13, ISHP14, JA10, JOE^{+}19, JWM^{+}16, JFK18, Kee18, KEF18, KBK19, KO11, LDF^{+}18, LB13, LM11, MKB17, Mal18, MC15, MWR^{+}17, MOR15, MHK^{+}10, MLTBMB14, MSC18, NW10, NBPF^{+}12, NBMB^{+}18, NBFO18, ODF^{+}16, OPMB12, PS11a, PKOF18, PO10, PA10, PAA^{+}15, PWB19, Pri10, RRR17, RDVSC16, RCM18, RDG^{+}13, SG17, SSD17a, SSD17b, SSD^{+}16, SV13, SOÖ12, SMH17, SF19, SLBC19, TÁSLC19, Val13, VSRD16, VvDM10, WCMHO18, WFSV15, WHMG^{+}16, WS15b, Zer18, vPJH^{+}17, vdMCJ17, vdPLMS16, BLCD16, GDG^{+}18]. change-related [WS15b]. changed [PTML16]. Changes [ENWL19, LC19, vHCT^{+}16, Árn17, BGG19, CN11, CLLS13, DSM^{+}13, FFKO15, GLTB^{+}15, Kim10, KV18, LSTD12, LBHS16, LCC^{+}18, MG19, MKMS12, PFMA^{+}19, SSW^{+}18, SS11, SAdC15, SPS14, SMS^{+}18, TWA^{+}15, TLA16, THE^{+}18, TLA18, BG16]. Changing [dVM14, dVM10b, AAvH12, CCC17, GAFAV16, GHP^{+}15, KH12, KFM17, LSW17, LB18, MLV^{+}11, MI15, PC13b, PS10, PKM^{+}19, RJ18a, TR18, WB19, WBB^{+}15, WLT^{+}19a, WP13, vPG10]. Channel [RDP^{+}18, VGL^{+}17, CSvH^{+}18, ECR^{+}16, RRRM^{+}15]. channels [Seu16]. Chapala [PG19]. characterisation [CVP^{+}17]. characteristics [BTS^{+}17, CMW10, MFMP10, Pan14, PKA^{+}16, THPJ17, WSG18]. Characterization [SRS14, FLT^{+}19, GCK^{+}11]. Characterizing [SSW^{+}18]. Charr [TRM19]. chart [KP17]. Charter [RL19, LSTD12, LL15, RTM^{+}19b, CBL18]. Charting [MPPB10, TTUR14, WML17]. Charts [KP17]. check [RCM18]. Cheilinus [PF12b]. Cheltenham [Tec19]. chemosynthetic [VSA^{+}12]. Chesapeake [FVH18, FVT19]. Chile [GCP19, MGA^{+}10, NVK^{+}17, NSB^{+}18a, OP15b, PFF^{+}18, RBTR^{+}19, SCMS15, Sch10a, Sch10b]. Chilean [HAC19, LQC10, LFGM15, MB10]. chilled [SGLY13]. China [LPP17, LLZ^{+}19, Lyo14, OS12, STG11, Sco18, Zha18, ZZ17c, BGG19, BG16, CL19, CLLL17, DGC14, DZLL15, Fab16b, FFSF17b, HWY^{+}18, JCD17, LF17, LZL19, LXWX17, LB18, LLX^{+}15, LZ16, Mal13, Mal16, PQLH15, PH16, Qiu13, RWJ18, SH14a, SHM13, SY18, Tan17, TO10, WCS18, WLV15, WSC19, WW19, Xin19, XDG^{+}16, YYYL15, YLZ^{+}19, YCY^{+}15, YC18, YY19, ZZ15a, ZZ16, ZS18, ZHH14, ZMZ18, ZLL19, ZLSL17, Rah19]. Chinese [Fab16a, FFSF17a, FFSF18, Gol13, He15, Kim19, LHH15, PWN18, WZF^{+}18, ZZ15a, Zha16, ZZ17a, ZZZL18, ZZZ^{+}17]. Chionoecetes [Han16]. Chiuta [Don17]. choice [BL18, DFV15, HTH13, HGBJ16, KM15, NMvPNL15, WK17]. choices [LJ17, ODL^{+}13, SU11]. choke [MUHH18]. chokka [COS14]. Choosing [DPJ^{+}13, Kir19]. Choppy [STG11, Bes19]. Christie [WFK^{+}12b]. Chumbe [NKCR13]. circumstances [De 17]. Citation [PDC10]. CITES [CMH19, CMC^{+}13, Coc15, FKWV19, KV18, Nij15, Nij17, PRT10, PC11]. CITES-listings [Nij15]. cities [Cha17b, Fen17, KCM^{+}18, PRS^{+}19, YCY^{+}15]. Citizen [JBKB15, HTA^{+}15b, MRHP19, VSCH16, WLT^{+}19b]. citizens [BYK^{+}15b]. citizenship [MF12]. City [MBvZ15, MKB17, DDvP^{+}16, Joh19, PRS^{+}19]. Civilization [ZQZ19]. claims [Bro16a, TBB12]. clam [GMD15, MOR15, PS11b]. clan [BRR18b]. clarification [SGW^{+}15]. clarifications [Bro16a]. Clarion [LJL^{+}14, VVS18]. clarity [GNHM18]. Clark [PC13a]. class [EP19, KS12b]. classification [HCF^{+}16, HCF^{+}17]. classifications [GAP12]. Classifying [AAT12]. clauses [Tei18]. CLC [Cho10]. CLC/FC [Cho10]. clean [Kai15a, WvLvK12]. clean-up [Kai15a]. clear [CJ16]. client [MFRvP16]. Climate [BGL^{+}16, Fen17, GM19, GDG^{+}18, HDL^{+}13, JFK18, SSD17b, TÁSLC19, AOAM13, ACTG17, ASB10, ACR18, BAHD10, BLC16, BvPS15, BF19, BSM^{+}14, CRCD13, CLLS13, CJW^{+}16, CPBQ19, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, DCHC17, DMB^{+}18, FOG^{+}11, FMB^{+}18, FC19b, Gil14, GHP^{+}15, Gra10, Gug19, HWO^{+}18, HPB^{+}15, IHS13, ISHP14, JOE^{+}19, JWM^{+}16, KNA10, Kee18, Kim10, KBK19, KO11, LDF^{+}18, LB13, Mal18, MHK^{+}10, MAJM^{+}18, MLTBMB14, NBPF^{+}12, ODF^{+}16, PKOF18, PAA^{+}15, RHW^{+}19, RRR17, RDVSC16, RCM18, RDG^{+}13, SG17, SSD17a, SSD^{+}16, SF19, VSRD16, WCMHO18, WF18, WS15b, Zer18, ZFG12, HDK^{+}15, NBPF^{+}12]. climate-driven [ASB10]. climate-informed [HPB^{+}15]. climates [LSW17, WLT^{+}19a]. Clipperton [LJL^{+}14, VVS18]. closed [CA17, HMH18, KR14, RNN^{+}18, SSLR14]. closing [Bro16c, Nol19, SvZdJ15]. closure [FJ12, GH15, HMGR14, KR14, RLH18]. closures [CBP^{+}16, MCB^{+}19, PMC^{+}15, RBWL11, SRA15, WAG^{+}15]. cluster [BDP19, DSF16, Dol17, KPR^{+}17, SAM13]. clustered [MJRdC16]. clusters [DCC19, DS18, DFK13, KPR^{+}18, PCC15]. Clyde [FÓ12, Wen11]. CMSP [HDG^{+}12]. CO [GZW^{+}19b, ZEP^{+}19, BGS13, CKY^{+}13, CSBE14, DK15, GSG^{+}15, Var11, WCM12, APC16, BHS19, BM15, BMGH11, CVP^{+}17, CC13, CCC^{+}14, DA11, Don17, EHR15, FVM14, FB15, GB15, GMD15, GHP^{+}15, Gor19, GJA16, HRC15, HA14, HAA15, HLL^{+}18, KMF^{+}18, LR15, LH16b, MB10, MNLHRMP17, MMPD^{+}14, MOR15, MMA16, NBFO18, PIH^{+}10, RGGFA17, RLHM14, Sch10b, SBMD13, SCR^{+}19, SCMS^{+}18, SMH17, TCS^{+}18, VC15, dVvT11, vdBAMJT19, GWWH17, GZW^{+}19a, Mil14, dlFRG17]. co-benefits [SCMS^{+}18]. co-developed [SMH17]. Co-existence [Var11]. Co-location [CSBE14, GSG^{+}15, BHS19, HA14, HAA15]. co-managed [GMD15, TCS^{+}18]. Co-management [BGS13, CKY^{+}13, WCM12, APC16, BM15, BMGH11, CVP^{+}17, CC13, CCC^{+}14, DA11, Don17, EHR15, FVM14, FB15, GB15, GJA16, HRC15, HLL^{+}18, KMF^{+}18, LR15, LH16b, MB10, MNLHRMP17, MMPD^{+}14, MOR15, MMA16, NBFO18, RGGFA17, RLHM14, Sch10b, SBMD13, SCR^{+}19, VC15, dVvT11]. Co-managing [DK15, Gor19]. co-operation [GHP^{+}15, KMF^{+}18, PIH^{+}10]. co-production [DA11]. co-use [vdBAMJT19]. coast [Asa12, CMW10, CSMR18, DHFL^{+}16, DDvP^{+}16, FC19b, GLTB^{+}15, GJA16, HCC^{+}19, HK18, KL14, KSC^{+}15, LRJ^{+}19, MMPD^{+}14, PRBX19, RROR12, RBWL11, Sat18, VLKN17, ZVP^{+}13, DL19a, DSH14, JWBJ10, KL12, KBG^{+}18, MC15, MCB^{+}19, RLH18, SBP^{+}19, WSG18]. Coastal [AG16, BKHA^{+}18, BM13, BE15, BH14a, Coo11, FSGB^{+}15, GDG^{+}18, GSDB14, HPPH16, MA18, RSS14, RCM18, RHB17, TLB19, TBB12, WRP^{+}10, ATAM10, AOAM13, ACTG17, ATB^{+}19, AJ12, AMvT11, APD^{+}19, BBR^{+}18, Bar10, BvRH11, BPB19, BAMAH17, BvPS15, BRM^{+}17, BY19, BRRG^{+}18, Bur15, CSH17, CRCD13, Cha14, Che12c, Cho12, CPBQ19, CFC^{+}18, CNFJ10, DSK^{+}19, DJTK17, Deb18, DFSB15, DRMA18, DTN10, DRH^{+}18, DSF16, DTMJH16, DMB^{+}18, DDvP^{+}16, Dva19, ESGC13, EWG^{+}18, EGF^{+}17, FOG^{+}11, FSWFG13, FJG^{+}14a, FMN15, GGK17, GLTB^{+}15, GBK^{+}15, GPA18, GBC17, Goo17, GOO^{+}19, GCM11, GCD^{+}15, Gug19, HBG12, HC19, HGS12, HVB^{+}17, HM18, HP12, HF12, HTH13, ISHP14, ISM^{+}17, JLW17, JSD19, JKS13, KNB^{+}18, Kel18, KCM^{+}18, KLY14, KO11, KM15, KHP14, KWT17, KVMM10, Lan15, LEKB10, LRJ^{+}19, LGFB^{+}19]. coastal [LWvdBL16, LZL19, Liu13, LCC^{+}18, LRS^{+}13, LTM19, MKX17, MAR^{+}15, MFB^{+}14, MVRG13, MB10, MGA^{+}10, MSdAM16, MTP^{+}15, MMGM16, MRSL16, MB18a, MTP^{+}17, NCH^{+}13, Nol19, PTDH17, PTML16, PCCR15, PGL^{+}16, PA10, PAA^{+}15, PRS^{+}19, PVS^{+}19, PPOM16, RCF^{+}12, RFG^{+}13, RFL^{+}18, RHW^{+}19, Rei16, Rei14, RWJ18, RR14, RWG^{+}19, RFG^{+}14b, RGSF19, RL13, RFHEJK13, SOUN16, SBC15a, SSCC19, SB14, SSB^{+}18, SBC^{+}14, SSD^{+}16, SCK15, SGN^{+}17, SBSK19, SMW13, SSA16, SCV^{+}17, SJ10, TC13, Tho13, TH17, TML^{+}19, TJ13, TE16, VPC19, Var11, VKA^{+}19, VGBG15, VBMM17, WBH^{+}13, WS15a, WM18b, WWH^{+}17, WBSW13, YP17, YCY^{+}15, ZVP^{+}13, ZMZ18, dlTCFB^{+}17, BGG19, FBP^{+}18, Gau10, HDG^{+}12, KW18, TE16]. coastal-marine [MVRG13, PAA^{+}15]. coastlines [GNW19]. Coasts [VKGN^{+}14b, CBP^{+}18, Dup14, MB15, SCK15, SF13, VKGN^{+}14a, CJW^{+}16]. cobalt [MP12]. Cod [Eli14, XDG^{+}16, BD19, BHM16, BB19d, DR10, EKV12, ERD^{+}13, HBG12, HJN16, HHL14, KBC^{+}13, MOE11, MG14, SS15b, SH15, SJT10]. Code [HFF11, KSC^{+}15, PDC10, WIC18, BTO^{+}17, vD16, Bha12]. codify [Bro11]. Codland [Fla10]. coexistence [AMdlCMS^{+}18]. cofradías’ [AdV18, GLVLGG19]. cognitive [dlTCL10]. coherence [BWD^{+}19, HGG^{+}18, JA10, Mal16, MM15b, Pla19, SGMG18]. coherent [CETQ16, LJ16]. cold [MSG11, BG16]. cold-water [MSG11]. collaboration [ACG^{+}17, vdVD16]. Collaborative [PSV^{+}14, AvTT17, DM15, KBB16, GB15]. collaboratively [KC13b]. collapse [DES10, OTOL17]. collateral [SVL^{+}16]. collation [BBN^{+}13]. collection [Das16]. Collective [HGG^{+}18, VG15, FMD^{+}18, HTT10, ODL^{+}13]. Collision [GCP19]. collisions [SCKP19, WS14]. Colombia [CSMR18, JSD19, Mat16, Sat18, WZCHS10]. Colombian [ELAC19, SDRML15]. color [KHP14]. colours [BVB18]. Columbia [EP19, EE17, GS17, HES16, HCC^{+}19, O’R15, Pin15, Rei19, WEKH19]. Columbian [GBF^{+}12]. combat [Boh18, You16a]. Combating [CL16b]. combinatorial [IT12, IT15, TI13]. combine [HL15b]. Combining [SRA15, RTM19a]. combustion [GZW^{+}19a, GZW^{+}19b, ZEP^{+}19]. come [PML10, RCD^{+}18]. comes [MJvP^{+}18]. Coming [GFC^{+}12, KS16, KS18a, WMVR18]. commensurate [RLM10, RLM12]. Comment [Hee14, KBB^{+}15, Kep14]. commentary [DJM11, McL15]. Comments [HBC^{+}13, PF12a, PZ17b, GPB^{+}18, PZ19, YBB^{+}17]. Commercial [Har15, KDW10, RBWL11, APC16, BDH^{+}18, Bro16c, BR13b, BR14, CHSW17, DHSC14, Dav12, ENWL19, FGW17, HLRO19, HSB^{+}16, JHW13, KS16, KS18a, LSTD12, Liv10, MACM14, MC19, MRSL16, MRCG11, MOH11, MM14, PB13, PM14, PJS13, REA^{+}14, RCD^{+}18, RTM^{+}19b, SLBC19, TLH18, TSW19, Uhr16, WFSV15, WEKH19, WMVR18, vdGMA^{+}13, BSSS17]. commercialization [RVdCGN11]. commercially [Coc15, FDS15]. commercially-exploited [Coc15]. Commission [HGDB10, AvTT17, Bai12, FÓ12, HD16a, Ili10a, Ili10b, MLHP15, VBHC^{+}17]. commitments [BP13, GCCSS^{+}19, HBP16, YPS13]. commodifying [BMW17]. commodities [RCUR13]. commodity [FDP18, Nol19]. Common [Foc11, GGADSR17, HG14, MMBK17, Pas14, Boh19, Bor18, BTC18, FTF^{+}17, JAG16, JGA17, KJH^{+}10, KB15, Lax10, LR15, LLO17, SCMS15, TCS^{+}18, Tho13, VRMF12, vD16, DCV12, EHR15, FQ13, GAFD^{+}18, KFPR10, KMTS16, Mar10, MFAA16, MPMK16, RDSB10, SMD14]. common-pool [LR15]. commons [FEF^{+}12, LKC15, OO19, PS11b, PHAC15]. Commonwealth [ENWL19, MCY^{+}19]. communal [TH15]. Communicating [MSBR15, NBPF^{+}12, SMW13, TH17, Tam15, Laf14]. communication [DFK13, Har10, LSAL19, Sin19, TS14, TTFC19]. communicative [FFSF17a, FFSF18, LJ13]. communities [ASP16, ABBD17, BdKvT19, Bar10, Ben17b, BvPS15, BRM^{+}17, CPKF19, CKN19, CJW^{+}16, CNFJ10, EJB19, EA19, ELAC19, FMN15, HGS13, HA13, ISHP14, JWBJ10, JWBJ13, JKS13, KMD15, KCJW15, KRM14, KJH^{+}10, KW16, KGS^{+}19, LM15, LBC^{+}16, LCC19, MSS19, MDE14, MGA^{+}10, Mat16, MA18, MCB^{+}19, MLTBMB14, NCH^{+}13, NW10, OPGL16, PA10, PFP17, PVS^{+}19, SG17, SWB^{+}14, VG15, VBMM17, Yos17]. Community [CSS15, LBK^{+}13, MK15, MSA^{+}18, NC10, RNS19, AAL17, AFCMET18, BRR18b, BKHA^{+}18, BCE^{+}17, BDSB16, BLR11, BRW^{+}13, CH16, CD15, CSCBR10, CCRT16, CvPH^{+}18, DAJM15, Don16, FBP^{+}18, FDGM16, FVH18, GGK17, GTM^{+}12, GFH^{+}17, GLJdlTC14, HL15a, HR17, HDM13, HD10a, HHG10, HRA^{+}15, JLW17, Lan15, LMH^{+}12, LR15, LTM19, LT19, LCR16, LCC19, MD16, Nak16, PS11b, PRS^{+}19, RJ13, Ros13, Roy16, SRK14, SSM15, THRS15, THA15, VAW14, WRP^{+}10, YP17, Yat14, vBvTM16, vdVD16, BD14, SM18a]. Community-based [LBK^{+}13, AAL17, AFCMET18, BCE^{+}17, DAJM15, GGK17, GLJdlTC14, HDM13, HHG10, Lan15, LR15, MD16, THA15, VAW14, YP17]. Comoros [FDGM16, HDM13]. companies [KS19, RMJ12, ZMZ16, LK15]. companion [KS12d]. company [WLX19]. Comparative [NKBS17, WRW^{+}17, CDO16, HSR^{+}18, JGMS16, Sev13]. compare [HK10]. compared [HBG12]. Comparing [APD^{+}19, FC19a, HLRO19, MLR^{+}16, Por17, LM11, ODF^{+}16, TPS15]. Comparison [IT12, PVK18, AAC^{+}13, BS10, GAW^{+}13, HM18, KR18, KH15, Lon11, LLX^{+}15, MP12, NS18, TL16, YCA^{+}12, YSJU^{+}17, LJT19]. Comparisons [TÁB^{+}10]. compatibility [KDW10, You16a]. Compatible [HDK^{+}15, THNR12]. compelling [WLV15]. Compensating [FW15, ZZ17b]. compensation [CLV^{+}16, DZLL15, KLMD15]. Compensatory [LPS12]. competence [FFSF17a, FFSF18]. competences [KP17]. competing [PS11b, TBB12]. competition [DGC14, PPML16]. competitive [NBD15, PT10, dVM10a]. competitiveness [YYYL15]. Complementary [RDSB10]. completeness [EGD18]. complex [Ben17a, Big15, DD13, DR14, IOG18, JE11, LPP17, LCC19, MRCG11, NMvP19, San11, Vil12, YTSO15]. Complexities [Eng18, OY19]. complexity [BE15, DJS17, LT17, Sin19]. Compliance [KSC^{+}15, AMLK^{+}12, ASBAM13, AA17, BR18, BD10, Bro15, CN17, DPS16, EKSZ14, FSS15, HvD13, HR16, JFA17, KL14, KH13, MJvP^{+}18, MC16, PJS13, RM16, RWK15, Sch15b, SJW^{+}16, Sun12, WP12, MFT19]. compliant [TZBBL18]. comply [JBJ12]. component [WWH^{+}17]. components [HLM^{+}13]. composition [BKH16, FPP^{+}15, GJA16, LBHS16]. comprehending [Won19]. comprehensive [CM17, LM14, RWG^{+}19, WCS18]. computer [RWL14]. Concentrated [CW15]. concentration [HES16, SC11]. concept [KML^{+}16, MV17, PRS^{+}19, RMB^{+}17, Stø19, SH19, ZP16]. conceptions [GLTB^{+}15]. concepts [KFBG13, Kir19, Ros13, SDF^{+}14a, SF13]. Conceptual [CPBQ19, DSK^{+}19, DMJ^{+}17, FVT19, GDD13, HdVB13, KC13b, Par15]. Conceptualisations [BDHSJ19]. Conceptualising [KEF18]. conceptualization [Gat18]. Conceptualizing [ET19, MHL^{+}15]. concern [MS14, OP15a]. concerning [vAS19]. concerns [GPB^{+}18, GZW^{+}19b, HCBM15, HFM^{+}13, KSK^{+}19]. Concessions [LPM^{+}14, IHF15, MMPD^{+}14]. Concluding [RJ18a]. conclusions [WWW10]. Concordat [TE16]. Concrete [AAL^{+}18]. concurs [SARF16]. condition [AAC^{+}13, KAF^{+}15, KJWP14]. conditions [BJGMG18, FOO^{+}15, HB14, KP13, LPC18, LGDB17, Næv17, PKM^{+}19, TTP12, VPC19, WPdR^{+}14]. Conduct [HFF11, KSC^{+}15, PDC10, BTO^{+}17]. conducted [DDP16]. Conference [Rei10, ZQZ19, MDNT19, CGF14, Cha19, FCdS^{+}15, WC16]. confidence [SK18]. Conflict [CGM^{+}18, JBV^{+}18, Mso16, WWW10, BF19, BMGH11, CKFLO19, GBW16, Gue19, MHL13, PPML16, ST13, VBMM17, WSWM17, ZVP^{+}13]. conflicted [Car19, Tor16]. Conflicting [HHL14, BHA^{+}14]. Conflicts [DZ12, EOC15, FWK16, LV18, PV16, Rei19, SDRML15, TBB12]. Conformity [MHM16]. confrontations [Lew16]. Confronting [MRC17, Pin15]. confusion [LBB^{+}12]. Conjoined [RD13]. conjunction [Kim10]. connect [GGLP16]. Connecting [NM18, PMM^{+}17]. Connections [VBMM17, Ben12]. connectivity [JS18, NRB^{+}18, PVS^{+}19, SM18b]. cons [Cou11]. Consensus [NFHJ16, Car19, CKC^{+}13, JBV^{+}18, MRG15, HJN16]. consensus-based [MRG15]. consent [BN19]. consequences [AKV15, ASB10, CCS14, CSCBR10, DJS17, GLTB^{+}15, GMM^{+}13, HdV17, NPAR19, Sat18, SSFL10, Vau17, VvD10, CS13]. consequential [VRB14]. Conservation [AFMCMP12, CS19, De 16, DFV15, HGDB10, LJ16, MKMS12, MPCB14, MRCG11, MLA16, ODL^{+}13, SLH^{+}13, SALS17, SCS14, ZH18, ALSC13, AG18, AFR^{+}19, AAGMH17, AL17, ADH18, ARGW14, ALM17, ACG^{+}17, ATB^{+}18, BAA10, BFTS^{+}16, BP13, BJS^{+}16, BD14, BTO^{+}17, BCP19, Ber19, BB16, BRM^{+}17, CSVHS14, CEB^{+}12, Cli13, CF17, CSS15, CBP^{+}16, Cro11, CP18, DJM11, De 13, DBB^{+}14, DJR^{+}18, DS10, DS17b, EAG18, EJB19, EG15, FJG14b, GRYA^{+}15, GJ15, GTL^{+}18, GW10, GHP^{+}15, GCHO19, Han12, HCBM15, Han19, HIN^{+}17, HDL^{+}13, HD11a, HD11b, Hob14, HB19, HBP16, HdF^{+}19, ItDD^{+}18, ISM^{+}17, JTBP17, JRP17, KVA18, Kam18, KLC^{+}15, KBF12, Kep14, KR14, KDW10, KJP14, KDS^{+}15, KMR^{+}13, LA14, LDS11, LGD^{+}19, LMP^{+}16, LB13, LLZ^{+}19, LTM19, LSAL19, MHL13, MHF13, MSS19, MSdAM16]. conservation [MGH14, MK15, Maz17, MGJB^{+}17, MRSH13, MGG^{+}19, Nei18, NGS16, Nor15, OP15a, OPGL16, OBC^{+}18, PCK^{+}19, PdSA^{+}13, PKR^{+}17, RR17, RFL^{+}18, RBC^{+}13, RGGFA17, RJ13, RWW^{+}14, RUH^{+}14, RRMSSE16, Rog15, Roy16, RFKEJ^{+}15, RS17, SOUN16, SCD^{+}16, SB15, SBTF15, Sch15a, Sco16, SNA^{+}17, SKL16, SMQ^{+}16, SSBL^{+}15, SB18, Syb15, TTM18b, TK11, TTJ18, TCCV16, TKN19, VBHC^{+}17, VMR^{+}18, WM19, WSO^{+}14, WMSB15, Wis14, YSJU^{+}17, YPS13, MSTB19, MRSH13]. Conservatory [SRGD18]. conserve [CYI13, GBK^{+}15, HD11a, HD11b, PMM^{+}17, SO13, Vir18]. Conserving [JGA17, Laf14]. consider [HD11a, HD11b, Shi16]. consideration [TGS18]. Considerations [GDG^{+}18, GMP^{+}14, KMvT15, LSK18, SM18a, WCK^{+}17]. Considering [Don16, LBC^{+}16, MMP15, Nei18, WLX19]. consistency [SAGG13]. Consolidation [AMG16, Bod17]. Consortium [MMB^{+}19]. Constraints [Nak16, PLP14, VMMB^{+}19, BKH16, CEB^{+}12, FLWG11, MMM19, Mor17, PS11a, SBC^{+}14, Tyn18, WKB15, HA14]. Constructing [Cha14, UA13, KMvT15, SM17a]. construction [CF17, Kai17b, Kai17c]. consultation [GC17, HD10b]. consumer [LBB^{+}12, SCOS17, TJ18, JGG^{+}16]. consumers [Far19, GGVRVL16, SCOS17, LHN18]. consumption [Fab16b, GTL^{+}18, SCD^{+}17, TJ18, WCMHO18, dMAH^{+}18]. Consumptive [HGS13, BRM^{+}17, FSB14, THNR12]. contamination [AYR^{+}14]. contemplating [Won19]. Contemporary [Ili10a, Ben17b, dS17a]. contestation [LLA^{+}19]. contestations [And15]. contested [Bro16a, ŠST^{+}18]. Contesting [HRJH15, BMW17]. context [AOY16, BDG^{+}14, BD13, CSP^{+}14, DYN^{+}19, FMB^{+}18, GAFAV16, GCG^{+}10, Har14, LMG^{+}16, Mee11, NNP^{+}14, OL10, PC18a, PAA^{+}15, PC18b, RL14, SVL^{+}16, SS19b, SF19, SH19, VSCH16, vAS19, She14]. contexts [ACM^{+}12, LCR16]. Contextual [ETEWZD11, FEF^{+}12]. Contextualising [WM17]. Continental [KS12a, KS12b, KS12c, Kay15, ZMM^{+}14]. Contingent [DV17, BvRH11, FFH18a, LKL15, PRS10]. continued [PS14]. continuous [BTOM13]. contours [BOF^{+}15]. contracts [HRJH15]. contradictions [SS19b]. Contrasting [GLW17, SPB19, BB16]. contribute [BDH^{+}18, CS19, HTA^{+}15b, MMdJ^{+}14, SSBL^{+}15]. contributed [WLT^{+}19b]. contributing [CN11, LYB17, QSMGVA18]. Contribution [RBTR^{+}19, DHM^{+}17, EZBL19, HRA^{+}15, JKK13, KM10, RMR^{+}17, Sch17, TSK18, HZH^{+}13, LGFB^{+}19, STJ^{+}12, TGdCR11]. Contributions [FCdS^{+}15, BCMH^{+}18, LZL19, RLAT19, RL19, HHS16, SS19b]. contributor [BFGR14a, BFGR14b]. Control [Gus18a, Boh19, CdLR^{+}15, CL13, DV17, EE17, ERD^{+}13, EKSZ14, GCW^{+}18, GDD^{+}17, GDD^{+}19, HES16, LEG19, MÁSSI13, MSH13, dMAH^{+}18, BB19a, CW15, CZX^{+}19]. controlled [dMAH^{+}18]. controlling [ZZ17b]. controls [CBL18, CYI13, EGGT12]. controversies [VvDM10]. controversy [OO17b]. convenience [CLLL17, FW19, KL12]. convenient [AGWMP19, McC17, SJ19b]. Convention [ND17, LWS19, ZP16, Dro13, DG14, DAB^{+}14, EZPRZ18, GDN19, JGMS16, LCC14, ZZ15a, ZH18]. Conventions [RM18, RB12, dGNC18]. convergence [DK15]. conversations [MUHH18]. conversion [FMD^{+}12, PML10]. Cook [CN11]. Cooke [Lon12]. COOL [He18]. Cooperation [EKSZ14, Gus18a, MGGS18, PCC15, Ros17, GDD11, Han14a, LH16a, MLF14, STG11, SS13, Zha18, vHHB14, vTRvLvH14]. Cooperative [HGB^{+}19, MV17, BFL18, CP15, EBR^{+}18, IHS13, ODL^{+}13, SCS14]. Cooperatives [MMPD^{+}14, Ben17b, DSH14, GKG16, MFP18, SCS14, WPdR^{+}14]. coordinate [MNA15]. coordinated [CETQ17, VKGN^{+}14b, WS15a]. coordinating [MF19a, PS11b]. Coordination [Hut19, CETQ16, HGG^{+}18, RvLvTH14, VHRM17, Wen16, Wen17]. Coping [PO10]. Coral [Hun13, NKCR13, Qiu13, TINL^{+}14, TCCV16, AOA^{+}15, Bre13, BM15, CG18, Che12a, COG^{+}19, DRG^{+}16, FMP^{+}18, HO10b, HLL^{+}18, KJWP14, LSP^{+}17, LGDB17, MSG11, SBW^{+}17, TTJ18, TCN^{+}13, TBB^{+}10, VSRD16, WCM12, Wea16, WGMG^{+}16, WS15b, CCTG^{+}13, CF17, FEF^{+}12, FE12, FAA^{+}13, GMP^{+}16, HBZH16, PPR^{+}15, RO13]. Coralliidae [CYI13]. Corallium [TRBS13]. corals [CYI13, Cha15]. Cornick [CMC^{+}13]. Cornwall [KW18, WPRA18]. Corporate [CP15, HES16, Bur16, EE17, Knu17, WLX19]. Corporate-cooperative [CP15]. Corporation [Bro18]. corporations’ [RIK16]. Correct [ND17]. Correcting [Bra12]. Corrigendum [BB19b, BFGR14a, GDD^{+}19, JS13, PBW17]. Corruption [Sun12]. Cost [BDG^{+}14, CLC^{+}16, DPV^{+}15, Elo10, LH12, AAL^{+}18, DRGCPS17, Das16, Dav10, DDV^{+}13, GBK^{+}15, GOTS19, HKR19, KS12a, KS12b, Kai17a, Kai17b, KTD^{+}15, MCMB^{+}18, MSZS^{+}11, Sea17b, SRGD18, SF16, WM18a, Ásg16]. Cost-benefit [BDG^{+}14, Dav10]. cost-effective [AAL^{+}18, GBK^{+}15, SF16]. Cost-effectiveness [Elo10, KTD^{+}15]. Costa [HGS13, SFCI14, DDL^{+}17, JSG^{+}18, LH16b, MBSL13, OPGL16]. Costly [LHB^{+}17]. Costs [FSV^{+}16, BMGH11, CLV^{+}16, CCS14, Hun13, LMH^{+}12, LJB^{+}14, RAG18, RNCGB15, SvZdJ15, SCMS15, SDC13, SS15c, Tur10]. could [CP18, DES10, EFK^{+}19, FG18, LGD^{+}19, PMN^{+}15]. Council [LSK18, CRO12, GSA10, Hol12, LHN18, MHM16, OR19, TZBBL18]. councils [BK16, HG14]. count [HRA^{+}15]. counter [AL18]. counteraction [SSJ17]. countermeasures [HO12]. countries [ACSK^{+}10, ACR18, Bar10, BAA^{+}15b, BSH^{+}19, BS18, BHM16, DRG^{+}16, FW19, GvdB13, GHM^{+}18, HM18, JS12, JS13, KVMM10, MBD13, NC10, NAB^{+}18, OY19, PHLM19, PRPLBLC12, PVS^{+}19, RGSF19, Sha10, SC10, VSRD16, ZT19, dVM10a, dVDA^{+}16, LH16a]. country [AYA^{+}19, BSAT18, CPB19, ETEW15, He18, KCJW15, KM15, PW16, RRRM^{+}15, Zag18]. country-level [CPB19]. country-of-origin [He18]. counts [GPA18]. County [LSC14, SSG12]. coupled [TRM19]. course [Kim19]. courses [TTUR14]. Court [Cla15, Hea13]. courts [CMFG15]. cover [GDN19]. coverage [RRRM^{+}16, WHB^{+}18]. CQM [MA15]. Crab [ØR18, Dav15, DHR16, Han16, MI15, SH16, MMDS12]. crabs [MOR15]. create [Kar17, NRB^{+}18, OPMB12]. created [HR17, OHH^{+}16]. creates [LHB^{+}17]. Creating [WP12, CSBE14, GRW^{+}16, Lan15, LPC18]. creation [ESGC13, Fla10, GC17]. credibility [BTOM13]. credible [HAK^{+}19]. Credit [GE17, HIBG15, PRBX19]. Credulity [PP19]. creels [ECF^{+}18]. Creeping [Lal17]. Cres [MHF13]. Crespo [McL15]. crew [CG13, Lew16, PGP18]. crews [SPMW12, TS14]. crime [Bes19, CSO17, Gla19, IW19, Stø19, SH19, VWG19, Vra19, Wit15b, vAS19, Wit15a, dCW15]. criminalizing [ZZ17b]. crises [Val13, VDB16]. Crisis [CCTG^{+}13, CGDB13, GA14, KLC^{+}15, LGH^{+}12a, LGH^{+}12b, PFMA^{+}19, VS18]. criteria [ASP16, CCL^{+}19, CE13, EG15, GCD^{+}15, HF16, HPP^{+}17a, PVVB17, RBS^{+}15, WCCO18, YYN^{+}15, YSJU^{+}17]. critical [BSM^{+}14, CBP^{+}18, CNFJ10, DL19a, EZPRZ18, FFSF17b, FÓ12, GvdB13, GBC17, KNY14, MTF^{+}15, MGJB^{+}17, MPT13, SB14, WCM12, AAL17, WS15a, Wea16]. critically [LHH15, SS16]. critique [BR14, DHSC14, KFPR10]. Croatia [LT19]. Cross [RRRM^{+}15, Bai16, DD13, ECR^{+}16, FP19, NRB^{+}18, SD13, SHS^{+}14, WP12, YFB16]. cross-border [SHS^{+}14]. Cross-country [RRRM^{+}15]. cross-cultural [YFB16]. cross-disciplinary [Bai16]. cross-jurisdictional [DD13]. cross-national [FP19]. cross-roads [ECR^{+}16]. cross-scale [NRB^{+}18]. cross-sectional [WP12]. cross-sectoral [SD13]. Crossing [ItDD^{+}18, Pol12b, Por15]. crossroad [BFO^{+}16]. crossroads [Her18b]. crowdsourced [CRB^{+}19]. cruise [DJS17, KMvT15, Nik19, WLX19, Won19]. cruisers [Bro11]. crust [MP12]. crustacean [SMR^{+}15]. crustaceans [AYB18]. Cu [Ngo18]. Cuba [BPB19, BRM^{+}17, GAFAV16, WPdR^{+}14]. Cuban [BRM^{+}17]. cucumber [ECL^{+}15, O’R15, PRPS17, PLP14]. culinary [SK15]. cultivation [ACTG17]. Cultural [Dro13, Cli13, DFV15, Gus18b, IH18, LZ16, MMGM16, NMvP19, dlÁPAVS19, RCUR13, RFHEJK13, TCS^{+}18, UAZ13, WPRA18, YFB16, dlTCL10]. cultural-cognitive [dlTCL10]. culture [BR18, ERB14, LC19, Næv17, NPS^{+}19, Soo17]. Cumulative [BAA10, BDF^{+}10, GCM11, MAAB15, RDP^{+}18, WTA^{+}16, WBGJ18]. Curacao [dGB10]. Current [MGB^{+}18, PM10, SC10, AMR14, COS14, DAB^{+}14, HAC19, Lea19, MLF^{+}17, MM17a, NFHJ16, RMRN^{+}16a, RCMBG^{+}19, Smi17, SJ19c, TÁSLC19, TC15a, VLPC14, YLZ^{+}19, CL19, CEB^{+}12, CJHR14, DGC14, HLM^{+}13, LPP17, MMGM16, Oan18, PR19, RMRN^{+}16b, SSM15, TBJC18, VSRD16]. currently [Lon11]. customary [CBF^{+}12, Cou11, Kar17, Ros17]. Cycle [VRB14, GWWH17, HBG12, HNVZ12]. cyclones [Uhr16]. Cyprus [HBL14].

D [PDC19, Rah19]. Daily [SKFL15]. damage [SGN^{+}17]. damages [PQLH15]. dance [KHvH^{+}14]. Danish [MNHN18, MA15, SKL16]. dark [FW19]. Darwin [De 13]. Data [HFM^{+}13, MNG^{+}18, AGB13, BBN^{+}13, BPYC19, BO12, BA19, CMA15, CBP19, Cha14, Cha16, CCK19, Das16, DMdJ^{+}15, DPJ^{+}13, Dru10, FTG^{+}18, FSS17, GOTS19, GA19, GW11, HA13, HTA^{+}15a, JMJ^{+}18, JHN^{+}19, JS18, Kai16, LSP^{+}17, LG19, LDVM10, LCR16, MBCR16, MMdJ^{+}14, MRHP19, MJS16, MHA^{+}17, MUHH18, MK14, Mul19, NWW^{+}15, NWW^{+}17, OCW^{+}17, PC11, PGC19, PZ17a, PGR14, PFCP^{+}14, PVPF19, PFMA^{+}19, RWK15, RCMBG^{+}19, She15, SS16, SCOS17, SGGM14, TRM19, TTP12, TLA16, THE^{+}18, TLA18, TMDSF19, VRT13, VGA15, WMJF16, YBB^{+}17, ZPT^{+}16, ZP18, ZMZ16, RHB17]. data-limited [CMA15, CCK19, MHA^{+}17, TRM19, TLA16, THE^{+}18, TLA18]. data-poor [GOTS19, HA13, NWW^{+}15, NWW^{+}17, PFMA^{+}19]. Database [PSCY12, Gar12, KP17, ZMM^{+}14]. date [BG15]. David [McL15]. dawn [Man17]. Day [Hav10, Hav13, PSP10, Sha10]. DEA [GW11, LVRM^{+}18, VRT13]. deal [Tho13]. debate [CNJ^{+}10, MF12, Mee11, VvD10, LHH15]. debris [CSF^{+}18, KWT17, LHJ^{+}15, LKC15, RM17, RBW18, VS18]. decade [BHT^{+}18, CSK^{+}18, Gar12, LPU^{+}16]. decades [Car11, MMC^{+}14a, SJ19a]. decapod [HA14, HAA15]. decay [AFR^{+}19]. December [Ano15u, Ano16x, Ano17r, Ano18u, Ano19b]. Decentralising [EHR15]. decentralization [BGG19]. decentralized [HW10]. Decision [BDF^{+}10, CPRP19, PGB^{+}17, BHS19, BH14b, De 10, De 16, ERCMD11, EG15, GGLP16, HCJ15, JWM^{+}16, LQC10, LM18, LLO16, LLO17, PRT10, PD17, PGB19, PPT10, PSCY12, QZH^{+}18, RBS^{+}15, SCKP19, Sev13, SM18b, KNB^{+}18, KJP14]. Decision-making [CPRP19, De 10, De 16, ERCMD11, LM18, LLO17, PPT10, PSCY12, RBS^{+}15, SCKP19]. decision-support [JWM^{+}16]. Decisions [MMC^{+}14b, GCD^{+}15, HVB^{+}17, KEM^{+}18, NYJW15, PLL17, RCF^{+}12, SFDB18, BJvH18]. declaration [SDP^{+}16]. decline [CN11, Ho10a, LSMW10, PS12, PS14, RNS19, WEKH19]. declines [BSM^{+}14]. Declining [GDD^{+}17, GDD^{+}19, BHAHW17, JLW17, MKB17, VGA15]. decommissioning [KS12a, KS12c, Kai17a, MNG^{+}18, TC15a]. Decomposing [FFKO15]. deconstructing [JDQV13]. decreased [BE15]. decree [BPB19]. dedicated [LLO16, PM14]. Deemed [SL16]. Deep [FFH18a, HL15a, HBT^{+}10, Jae16, LACG18, TTFC19, ARJ18, BSAT18, BTC18, BS18, CSF^{+}18, Cho11, CKC^{+}13, CCRT16, DDM^{+}19, DMJ^{+}17, DLM^{+}18, FG18, FFH18b, JAG16, JGA17, JML16, JFK18, JDM^{+}19, Kim17, LM16, LMG^{+}16, Lil16, MFF^{+}18, MB18b, NWW^{+}15, NWW^{+}17, NBC^{+}12, Oan18, OHB^{+}18, PKA^{+}16, SSA16, SKT^{+}10, VSA^{+}12, VAP^{+}14, VAHG^{+}18, WS14, WL18, WM18a, ZCD^{+}15, Smi18b]. Deep-sea [HBT^{+}10, LACG18, TTFC19, CSF^{+}18, CKC^{+}13, DMJ^{+}17, DLM^{+}18, JFK18, JDM^{+}19, Lil16, NBC^{+}12, Oan18, PKA^{+}16, SSA16, SKT^{+}10, VAHG^{+}18, ZCD^{+}15]. deep-seabed [ARJ18, JGA17, LMG^{+}16]. deep-water [MFF^{+}18, NWW^{+}15, NWW^{+}17]. deeper [CG18]. deepwater [HPW^{+}18, Kai17c, FW15, GSS^{+}18, HB19]. defence [EGF^{+}17, PP19]. Defiance [Sch15b]. Defiant [Bur16]. deficiencies [KH11]. defies [FKWV19]. define [GFMG^{+}14]. defined [UNS^{+}18]. defines [vdVD16]. Defining [ADHV16, DDL^{+}17, HF12, KARR^{+}19, LMG^{+}16, RFL^{+}18, RWBF19, ZHH14, FB15, KSWS18, SSLR14, LLO16, MHL^{+}15, Par10]. definition [HWL^{+}19, SRGGVL13, WNB^{+}15, CKdS14]. definitions [MS14]. deforestation [AG16, ACTG17]. degradation [FSGB^{+}15, LU13]. degree [GS17]. degrowth [Had18]. del [SJW^{+}16, SSM15]. delay [MCMB^{+}18]. Delegating [NHA15]. delegation [Mso16]. delimitation [Lan18, QG17]. delimitations [Nag16]. delineation [RBJ^{+}11]. deliver [PBJ^{+}14]. Delivering [KCKA^{+}18, CBP^{+}18, MSS18]. delivery [HID12]. Delta [HvD13, SNIN17]. demand [AFGRF^{+}17, dHJTG17, GGVRVL16, JLW17, PR19, SCOS17, TRC^{+}17, TCC^{+}19, VLKN17, ZMZ18]. demands [PHAC15]. demersal [BW19, KKG16, MUHH18, RBS^{+}15, SKT^{+}10, WBH^{+}15, ZH14]. Democracy [SJ14]. Demonstration [MKX17]. density [LGM16]. dependence [JWBJ10, Jac19, SRS14]. dependency [Ros13]. dependent [CJW^{+}16, CA17, FLWG14, LT19, NCH^{+}13, PFCP^{+}14, Wit19]. depleted [FP13]. deployment [GOHDB17]. deposits [CCC^{+}13]. depredation [CJB15, RTM^{+}19b]. derbies [MCCC16]. derelict [Cho11, KLY14, Uhr16]. derived [AVH10, DHFL^{+}16, Dru10, MRHP19]. Deriving [SG17]. desalination [HPPH16, HPP^{+}17a, KC13a]. Describing [BDSB16]. Description [CDR^{+}16]. Descriptor [PRO16]. Desiderata [VAP^{+}14]. design [CR10, CS13, GRW^{+}16, GMP^{+}14, HW10, KW16, LR15, NWD18, SWP13, SM18b, WK17, SSG^{+}14]. Designating [VSA^{+}12, MPT13]. designation [LHH15, MDR^{+}16, RRRM^{+}15]. designations [CSBE14]. Designed [KFPR10, BO12, DTM18]. Designing [CMA15, GMM15, MMTC16]. designs [CBF^{+}12, IT12]. destruction [Han14a, ŠST^{+}18]. Destructive [WHMG^{+}16, BD19]. detail [She15]. detailed [LCR16, MMBK17]. Detecting [JL11]. detection [VDB16]. detention [CZX^{+}19]. Determinants [AFGRF^{+}17, PW16, FSS14, FP19, NBM15]. Determination [WZF^{+}18]. determine [CA16, WCCO18]. determines [TMG15]. Determining [AVC18, GS17, GW11, KS12d]. Devastation [MGA^{+}10]. develop [FSCJM19]. developed [CH16, FSS17, SMH17]. developer [KLMD15]. developers [HAA15]. Developing [CCC^{+}14, DDS^{+}11, DMB^{+}18, HWL^{+}19, NWD18, PO14, RPG^{+}14, SOBS18, Sen10, SK18, Ant16, BBR^{+}18, BPB19, CC17b, DGC^{+}17, DS10, ETEW15, GvdB13, GSB^{+}17, GHM^{+}18, HP13, JS12, JS13, MKR12, MJS16, MGGS18, Oya18, PRPLBLC12, PVS^{+}19, TLH18, WNLS17, ZT19, CL13, CCL^{+}19, CKC^{+}13, LAH^{+}13, MBD13]. Development [ABB14, APC16, JWBJ13, KFM17, MHC19, NM18, RFL^{+}18, SI18, SMR^{+}15, SCMS^{+}18, Vir18, VKGN^{+}14b, ZQZ19, ABSB15, BDHSJ19, BW18a, Bro10, Car11, CLS10, CQ14, CLLL17, CGS^{+}17, DJTK17, DK12, DSF16, DAB^{+}14, EFK^{+}19, FWHS17, FSM^{+}12, GS16, HL15a, HBS^{+}19, HCC^{+}19, HRA^{+}15, KS12a, Kim10, KAPF^{+}19, KMvT15, KKSC19, LPC18, LBVN19, Liu13, LT19, LCC19, MDG^{+}14, MHA^{+}17, MCY^{+}19, Men18, MTP^{+}17, MO12, Nik19, NKAP13, OL11, OY19, PHLM19, PWB19, PHAC15, RDP^{+}18, REA^{+}14, ROD15, RWG^{+}19, RNS19, RCMBG^{+}19, SG15a, SC19, SPGP12, SSCC19, San18, STG11, Smi13, SHM13, SS19c, SSA16, SSFL10, SF13, Stø19, SH14c, TBKH13, VKGN^{+}14a, WL18, WCS18, YBD^{+}19, YY19, AMdB12, EMD^{+}18, HDK^{+}15, MB15, PSV^{+}14, Vir18]. developments [ALMS14, CL19, EGF^{+}17, GDD^{+}17, GDD^{+}19, LPC18, MMOL15, She15, SAGG13, WMJF16, WMK15]. devices [BAA^{+}15a, DMMG14, DBD19]. devils [BTOM13]. devolution [CMFDCLV10]. Devon [DHR16]. dexterity [PA10]. Dialectics [CHSW17]. dialogues [WKB15]. dichotomy [KTA^{+}13, VBHC^{+}17, vAS19]. Didemnum [SHS^{+}14]. differ [KRM14]. difference [Edd14, Kir19, KGS^{+}19, Ols11, PHLM19]. Differences [CMW10, HR17, KMF^{+}18, KWT17, VvD10, RRRM^{+}15, Sev13, CA17]. Different [KKWS15, Gus18b, HK11, KCM^{+}18, MMJM15, RAZ16, Rog15, SL15a, TII^{+}19, WSG18, WNB^{+}15, vHCT^{+}16, Fab16a, VvDM10]. differential [AHB18]. differentiated [LHN18]. differentiating [BKKN10]. Differing [Por17]. difficulties [SG12, DAGG16]. dignity [Son15]. digraphs [KEM^{+}18]. dimension [DYN^{+}19, Ger15, HvPN^{+}19, LJ14, SMKBM16, dVM14, WGMG^{+}16]. dimensional [BC13]. Dimensions [Pla19, CBG^{+}17, DTJ10, EAA^{+}18, JAO^{+}16, KRK13, LM11, MLV^{+}11, MAHP16, MM14]. diminishing [Bro11]. dioxide [GZW^{+}19a, GZW^{+}19b, HK11, MSBB14, ZEP^{+}19]. diplomacy [BDIS18, Gre15]. diplomatic [ØR18, GCG^{+}10]. Direct [BLR11, Knu17]. directed [Ber19]. directing [RWW^{+}14]. direction [SMQ^{+}16]. directions [GAFD^{+}18]. Directive [BFGR14a, PRO16, BLG^{+}18, BFGR14b, FGKMP14, HWN11, LLO16, ODRRM12, Thi13a, Thi13b, ZPB^{+}13, AGB^{+}18, ACG^{+}17, BJGMG18, BR13a, BDG^{+}14, CETQ16, CBE^{+}19, HTH13, KTD^{+}15, LJB^{+}14, LLO17, MKU19, RTD^{+}17, RDSB10, SGGM14, dVM12, vHHB14, vLvHvT12, vLRvH^{+}14]. dis [Mat16]. dis- [Mat16]. disagree [Lea19]. disappearance [CL19]. disaster [GC18, MGGS18]. Discard [GBBB15, PVVB17, SOS12, Eli14, GRLM15, LB17, MMS^{+}14, MMBK17, VPP^{+}16]. discarded [Gil15]. discarding [GKSJ11]. Discards [FPP^{+}15, CGC14, CSMB^{+}12, JE11, MdMGV^{+}18, MMS^{+}14, SSC^{+}15, SF16, VRMF11]. discards-proportional [SF16]. Discerning [BR18]. discharge [TML^{+}19]. discharges [Kno11]. disciplinary [AARK15, Bai16]. disciplines [Tip16]. disconnect [Mee11]. discontinuity [MSC18]. discourse [BMPV10, HG14]. Discourses [FHJ16, Bre13, Mis19, OCCvT^{+}18, LJ17]. discovery [Bes11]. Discrete [HGBJ16, DFV15]. discussion [HO12, JLQ16, VHA^{+}19]. disease [MSBR15, ML19, MG16]. disenchantment [Kam18]. Disentangling [RV16b]. disincentives [DPB17]. disjuncture [HBS^{+}19]. disparate [LCC19]. disparities [CSCBR10, LXWX17]. displaced [Sen10]. displacement [Vau17]. display [SJW^{+}16]. disposal [LBN15, LEKB10]. Dispossession [Kam18]. disproportionate [ADHV16]. dispute [DZ12, OY15, ØR18]. Disputed [OS12, Zha16]. disputes [HJN16, HO12]. disrupted [PTN^{+}17]. dissonances [Rog15]. Dissostichus [GRLJ15]. distal [FTF^{+}17]. distance [FGW17]. distances [CACL18]. Distant [CLS10, Mal13, NYJW15, TOL12]. distracting [AGWMP19, SJ19b]. distraction [SJ19c]. Distributing [Han12]. Distribution [HEJK10, PCLE17, REA^{+}14, AMG16, DPH^{+}15, Fla10, HD10a, HID12, MGH14, OC17, OHB^{+}18, SBW^{+}17, SPS14, TKAO^{+}19, VPC19, WBC14, MMS^{+}14]. distributional [HDJD15, MBBDS19]. distributions [Bre13, NKBS17, TPS15]. distributive [GKSJ11, GLJdlTC14]. district [SY18]. Disturbing [Pin17]. dive [ADH18, DHM^{+}17, LTM19, MMTC16, PDT^{+}14, RPG^{+}14, dGB10]. diver [TCCV16]. divergence [DK15]. divergences [San11]. Divergent [dAdOS17]. divers [KJH^{+}10, LGDB17, dAdOS17]. diversification [Mat16, MPMGGAHR18, Mor17, SPGP12, SLBC19]. Diversifying [BAA^{+}15b, WBSW13]. Diversity [CJ17, DAB^{+}14, JGMS16, LLA^{+}19, BHI^{+}17, BDSB16, HD16b, JQD13, MC14, MTP^{+}17, PAA^{+}12, RLKH19, RWW^{+}14, Sco16, TC15b]. divestiture [WBB^{+}15]. divide [HEJK10, KJS14, Por15]. diving [GH18, KC13b, VMR^{+}18, WS14, ZKR^{+}19, CG18, DM15]. division [PG19]. Djerba [DV17]. DNA [ELAC19]. Do [PBJ^{+}14, PHS18, RPdO^{+}19, BD14, ELAC19, GJ15, Kee18, LGFB^{+}19, MMS^{+}14, PML10, SFCI14, VPP^{+}16, YFB16, GCW^{+}18, ISM^{+}17, TC15a]. Docking [BN19]. doctrine [Bes11]. documented [vHCT^{+}16]. documents [WLV15]. Does [DWCS18, Han14a, MC16, NBFO18, PD17, RWG^{+}19, AAL17, DW13, DRH^{+}18, GFA^{+}18, MFT19, Sco18, KHD12, Pas14, PVA^{+}18, WWW10]. dogfish [HKR19]. Doing [RCM18, MWR^{+}17]. doldrums [LJ16]. dolphin [Che11, FSS15, LHH15, MJS16, SC13b]. dolphin-swim [FSS15]. dolphins [FSS17, HVB^{+}17, PRT10, SHM12, WIC18]. domain [MSH13]. Dome [JSG^{+}18]. Domestic [BDIS18, AF14, CGP^{+}18, CO16, Sea17b, AA17]. dominance [BYK15a]. Dominant [Bre13, BFGR14a, BFGR14b]. Dominican [CWP16]. Dominion [Rei10]. Donggang [LYB17]. doomed [FV10]. dormant [Tei18]. Dose [LOD^{+}15]. double [MMN18]. double-rig [MMN18]. down [ACA^{+}13, Che10, Fol13, GBF^{+}10, HD16b, JD16, LRS^{+}13, MMBK17, SC13a]. downlisting [BJS^{+}16]. Drafting [GDG^{+}18]. Drawing [LRJ^{+}19]. dreams [Par10]. Dredge [OCW^{+}17, LBN15, OCG14, OCG15]. dredging [FJ12, LEKB10]. drifting [DMMG14, VGCF18]. driftnet [Sal16]. driftnets [SLS18]. drilling [KS13]. drive [CSH17, OHH^{+}16]. driven [ASB10, HLC^{+}10, RRR17, SJ10]. driver [BSM^{+}14]. Drivers [PPML16, ACR18, BSD^{+}17, Bre15, FTF^{+}17, GAKP^{+}19, KVMM10, NMvPNL15, SMS^{+}18, TINL^{+}14, Tyn18, WM17, vHHB14, WWT^{+}16]. driving [CLS10, RGG^{+}16, RV16b, AKV15]. Dudley [HCF^{+}17]. due [MCB^{+}19, SG17]. dugongs [HIN^{+}17]. dumping [HL15b]. during [AAL^{+}18, BHT^{+}18, GWWH17, HD10b, MPMGGAHR18, PKOF18]. Dutch [HdV17, dVM10b, dVvT11, vLvT10, vdVD16]. Dynamic [MHL^{+}15, DRPSA17, Hey17, Hey18, SSL16, TRCR^{+}19]. dynamics [BD10, BY19, CMVLGG14, LSP^{+}17, MFAA16, ST17, SPS14, TTM18b, VvD10]. dynamism [MTP^{+}17]. Dysfunction [CS13]. Dysfunctions [Lax10].

E. [Tur10]. Each [SL15a]. EAF [OPMB12, PP10]. EAFM [RMRN^{+}16b, MJS16, RMRN^{+}16a]. Early [AMvT11, EOC15, LWS19, VC15, GFC^{+}12]. Early-stage [AMvT11]. earnings [EA19]. Earth [HBZH16]. East [BTS^{+}17, CN17, CRO12, PDT^{+}14, RJ13, GCG^{+}10, JRS^{+}16, UA13, CETQ16, Cho11, CR11, GCD^{+}11, LPP17, Por15, STG11, TC13, TKN19, YSJU^{+}17, Yia18, ÁFFSCF17]. Easter [AG18]. Eastern [VAP^{+}10, BAMAH17, KGMP18, MOE11, San15, BDHSJ19, CJW^{+}16, EKV12, LLH^{+}14, LLO16, WFK^{+}12b]. easy [CSH17]. eat [HPP^{+}17b]. eating [TN15]. EBFM [RLM12, DL19a, RLM10]. EBM [AH19]. EBSA [JFT^{+}18]. EBSAs [YSJU^{+}17]. ECDIS [KP17]. echo [vS15]. echo-sounding [vS15]. Eco [HA13, KHD12, RPG^{+}15, KCM^{+}18, KVH14, MM15a, PHS18, SVL^{+}16, SSV18, SCOS17, TK13, WC17]. eco-engineered [KCM^{+}18]. eco-friendly [SSV18, SCOS17]. Eco-labeling [HA13, KHD12, SVL^{+}16]. eco-labelling [KVH14, WC17]. eco-labels [MM15a]. eco-rating [PHS18]. eco-tourism [TK13]. Eco-touristic [RPG^{+}15]. ecolabel [LHN18]. ecolabeled [LHN18]. ecolabelling [BHI^{+}17]. ecolabels [WB18]. Ecological [Árn17, GOK14, KMvT15, PVS^{+}19, VAP^{+}14, WFK^{+}12a, ZQZ19, AG18, AOAM13, APD^{+}19, Ben17a, Big15, Big16, BDB19, BCE^{+}17, BD13, CHBRJ19, CN11, CCC17, CNFJ10, DDM^{+}19, ECSE^{+}17, FSWFG13, FHM^{+}10, GGK17, Gat18, GOM17, IOG18, Iri18, JG13, JQD13, KVA18, KL12, KAF^{+}15, KC15, LLF^{+}16, LSP^{+}17, LPS12, LC19, LLX^{+}15, MSC18, MI15, MD15, MDM16, NMvPNL15, NHPD19, ÖGB^{+}10, ÖSS^{+}11, Par15, PO10, RR17, RDP^{+}18, RR18, SARF16, SPO^{+}18, TDI^{+}14, TMG^{+}18, Vil12, WLP10, YBD^{+}19, dlTCFB^{+}17, BBN^{+}13, BBCPC11, LCC^{+}18, SWP13]. ecological- [ÖGB^{+}10]. ecologically [GAP12, LJ16, YYN^{+}15, YSJU^{+}17, DAB^{+}14]. ecology [BMD^{+}14, HBL14, LE12, MOE11, Nol19, PWR17, RFG^{+}14a]. Economic [DGVS^{+}16a, DGVS^{+}16b, GVSZ19, GSDB14, HHS19, IMA16, KYS10, Lee10, LKL15, MMN18, Pan14, RDVSC16, RBS^{+}19, SHL^{+}18, TEF^{+}15, VMO14, AYR^{+}14, ALMS14, AdV18, BSD^{+}17, BTS^{+}17, BDF^{+}10, BLR15, BAO^{+}12, Bro16c, BF19, CS14, CdLR^{+}15, CVP^{+}17, CGDB13, CKdS14, CSP^{+}14, CA17, CE13, DRPSA17, DR10, DHM^{+}17, DRG^{+}16, DSF16, ERD^{+}13, EGHC17a, EJB19, FvRA11, FFH18b, GFC^{+}15, GAFAV16, GAKP^{+}19, GFH^{+}17, GA19, GHO^{+}17, HCP^{+}17, HdV17, HLRO19, HCJ15, JLH14, JHW13, KL12, KLC^{+}15, KSS17, KJH^{+}10, KVMM10, LB17, LJS^{+}18, LCC^{+}18, LBP^{+}19, MKB17, MPPSdVSOHR14, MLA16, Muk15, NALO19, NLP11, PQLH15, Par10, PGC19, PB13, PHC14, PC18a, PGC16, PVK18, PAA^{+}12, PCLE17, RMH^{+}15, RCdLS13, RWJ18, Ren18, RLH18, RvdBG13, RL19, Ros18, RFHEJK13, RRC^{+}17, SCD^{+}16, SSCC19, SRA15, SCC^{+}17, SBP^{+}19, SG17, Seu16]. economic [SK18, SAM13, SWS12, SMS^{+}18, SRGGVL14, TTJ18, TDI^{+}14, TINL^{+}14, TH17, VPC19, VVS18, VLKN17, VBMM17, WHP^{+}16, WZCHS10, WCCO18, YYYL15, YBRBB16, ZKR^{+}19, Hav13]. economically [RVdCGN11]. economics [EGHC17b, Gat18, GA14, Liv10, MP12, SCS16, WPJ10, PRPS17]. economies [dlFFVRC16, HZH^{+}13, JTK18, Jen15, KM10, WNLS17]. economy [BRW^{+}13, FMMA^{+}15, FMGV16, GGSRVL17, Han19, JBBV^{+}19, KARR^{+}19, LY14, LE12, LXWX17, Mal18, MSTB19, MLA16, MO12, MO13, Pau18, Pin15, RR14, RMR^{+}17, RSMM19, SAM13, SHM13, SRGGVL13, VG15, WW19, ZHH14, CMMBV^{+}19, KSWS18]. Ecoplata [PCCR15]. ecoregions [TJJAM19]. EcoSpatial [ZMM^{+}14]. Ecosystem [CLV^{+}16, CdLH^{+}18, CR10, CCRT16, HvPN^{+}19, KM15, PSCY12, PGPS10, RvLvTH14, RMRN^{+}16b, SCK15, SC10, AMLK^{+}12, AH19, ACR18, ACM^{+}12, AAL^{+}18, ATDP16, BLG^{+}18, BK16, BYK^{+}15b, BBP^{+}14, BDH^{+}18, BMGF^{+}17, BD13, BPBM17, BSM^{+}14, CC13, CHKG16, CSF^{+}19, CP10, DL19a, DFSB15, DMdJ^{+}15, DDS^{+}11, DFK13, Dva19, EGP^{+}18, EGGT12, EAA^{+}18, ERCMD11, ECR^{+}16, FMM13, FR16, FLH11, GMD^{+}12, GRMAC18, GMD15, GHM^{+}18, Gib10, GRW^{+}16, GMP^{+}16, GAB^{+}17, HAH^{+}16, Han13b, HPB^{+}15, HM18, Hor10, HDTS14, HRA^{+}15, IOG18, Iri18, KCKA^{+}18, KS18b, Kir19, KRK13, Lee10, LCH^{+}13, LEM^{+}18, LPS12, LWvdBL16, LPR^{+}17, LCS15, MOE11, MM18, MD16, MRC17, MSdAM16, MTP^{+}15, MMdJ^{+}14, MMGM16, MBB18, MCHO11, MDR^{+}16, MYA^{+}18, MM17b, NVK^{+}17, Nei15, Ngo18, Ols11, ÖGB^{+}10, OOvdM^{+}11, ODRRM12, OCCvT^{+}18, PIH^{+}10, PL15]. ecosystem [PTML16, PVA^{+}18, Pla19, PPR^{+}15, PMA^{+}17, PBJ^{+}14, PM10, PLP14, RMH^{+}15, RWL14, RLM10, RLM12, RV16a, RvLG^{+}15, RBB^{+}16, RO13, RFHEJK13, SLJ^{+}11, SLH^{+}13, SBC15a, SSCC19, San19, SVL^{+}16, SPO^{+}18, SBS^{+}13, SGMWE14, TLR^{+}10, TÁSLC19, TJJAM19, TH17, TML^{+}19, THF^{+}15, TTFC19, VSA^{+}12, WSO^{+}14, WTD^{+}15, WSA^{+}15, WGMG^{+}16, Wen17, WPRA18, YWC^{+}16, dLS10, vH15, DTMJH16, DTM18, FSMG10, KAPF^{+}19, Man10, OPMB12, PGP^{+}19, PFF^{+}18]. Ecosystem-based [CdLH^{+}18, CR10, HvPN^{+}19, PGPS10, RvLvTH14, SCK15, AMLK^{+}12, AH19, ACM^{+}12, BK16, BMGF^{+}17, CC13, CHKG16, DL19a, DFSB15, DDS^{+}11, EAA^{+}18, ERCMD11, FLH11, GAB^{+}17, IOG18, Iri18, Kir19, KRK13, LCS15, MRC17, MMdJ^{+}14, MDR^{+}16, Ols11, OOvdM^{+}11, ODRRM12, PL15, Pla19, RWL14, RO13, SLJ^{+}11, SLH^{+}13, SBC15a, SVL^{+}16, SPO^{+}18, TLR^{+}10, WSA^{+}15, WGMG^{+}16, DTMJH16, DTM18, PFF^{+}18]. ecosystems [ACSK^{+}10, ABBD17, ACP^{+}14, DGC^{+}17, HPPH16, HA13, LPS12, LHB^{+}17, MFMP10, MSGM^{+}12, OGM^{+}14, PS19, RWG13, SSD^{+}16, TH15, VAHG^{+}18, VMMB^{+}19, WHP^{+}16, WF18, ZCD^{+}15, JL11]. Ecotourism [WCMHO18, PGC16, SB18]. Ecuador [ATB^{+}19]. Ecuadorian [GOdFC16]. Ed [Tec19]. edge [KJS14]. Editorial [CK11, Han19, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19c, Ano19d, Ano19e, Ano19f]. Editorial [Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n]. Editors’ [PCS14]. Eds [Rah19]. education [FFSF17b, GdV16, HP12, JJ19, LA14, LMH^{+}12, LEFC19, RPG^{+}15, SB14]. educational [WMSB15]. educators [HPH^{+}18]. Edward [Tec19]. edwardsii [LSMW10]. EEA [Dva19]. eel [Doy15, Kai19, Nij15, Nij17, SG12]. eels [Nij15]. EEZ [Árn17, GSG^{+}15, OIS^{+}18, VRMF12]. Effect [She16, Bes12, BB16, FSCJM19, KP17, VRT13]. Effective [AMLK^{+}12, DD13, FMB^{+}13, GDG^{+}18, AAL17, AAL^{+}18, BFTS^{+}16, BS18, CG19, DJR^{+}18, EGF^{+}17, FOG^{+}11, FSCJM19, GBK^{+}15, KP13, Kar17, KW14, LEKB10, LGD^{+}19, LEM^{+}18, LT17, LYB17, PGR14, PMC^{+}15, SLGC^{+}18, SHM12, SF16, YCY^{+}15, vLRvH^{+}14, MLF14]. effectively [FSS17, PKOF18]. effectiveness [BR13a, Bha12, BRRG^{+}18, CKY^{+}13, CH16, DSWG13, Elo10, HR16, HDM13, HD11a, HD11b, JFK18, KTD^{+}15, KKSC19, Mal16, MMJM15, Pin13, RR18, SVA^{+}12, TAGP15, TE16]. Effects [CGDB13, ECF^{+}18, SMH17, Bla15, BW18b, Bro16c, BHT^{+}18, CD15, CA17, CHL^{+}19, CKBS14, DFL15, FFH18a, GGSRVL17, GA14, GBF^{+}12, HDJD15, JLH14, KMF^{+}18, KL12, KKG16, LEFC19, LEG19, LSAL19, MSBR15, MAAB15, NLP11, Por17, RMH^{+}15, RCdLS13, RLM10, RLM12, San18, SJP17, SWP13, ŠST^{+}18, TCN^{+}13, VS10, WTA^{+}16, WBGJ18, BSKH17]. efficacy [FCC^{+}12, PGL^{+}16]. Efficiency [SL15b, CG13, CGDB13, FFKO15, GWWH17, GOM17, HCP^{+}17, LVRM^{+}18, MFP18, PHC14, Pin13, QSMGVA18, RST17, RYJ^{+}17, SS15b, VPC19, VS10, WLX19, Zag18, ZH14, AVC18]. efficient [FSMG10, GW11, NWD18, ZPB^{+}13]. Effort [MKMS12, BRM^{+}17, BAT^{+}18, CdLR^{+}15, Cha16, DMdJ^{+}15, DPH^{+}15, ECF^{+}18, Gar12, GMP^{+}14, GZW^{+}19b, HK18, JBK15, LSMW10, MÁSSI13, MNA15, MCBM18, Mul19, Pin13, RCdLS13, SKFL15, Sen10, TKAO^{+}19, Vau17, TZBBL18]. efforts [EAG18, EJB19, Fen17, GBBB15, GAB^{+}17, RFL^{+}18]. EFZ [Árn17]. Egadi [DFP^{+}16]. egg [MBSL13]. Egypt [LKMN16, SNIN17]. elasmobranch [BTS^{+}17, PSV^{+}14, TdCdSR16]. elasticity [RW15]. electric [LBVN19]. electricity [KLN17]. electronic [ENWL19, PHEM^{+}17, KP17]. eleginoides [GRLJ15]. elements [CAP15, RFL^{+}18, dlTCL10]. elephant [Tur10]. elephants [vH15]. eleven [RWJ18]. Elgar [Tec19]. elicitation [WML^{+}16]. Elicited [HLM^{+}13]. Eliciting [CCK19]. Eligibility [AVC18]. Eliminating [Cho12]. embeddedness [FMN15]. Embracing [BJS^{+}16]. emerge [GB15]. Emergence [RLKH19, CHSW17, JTP18, MO14, Mes12, PCC15, RO13]. emergencies [WZF^{+}18]. emergent [BAT^{+}18, GBF^{+}10]. Emerging [BHG11, GH18, APC16, DMC^{+}17, JHB^{+}18, LKMN16, QJ13, RMRN^{+}16b, SJ15, Weg15, Wri15, dVM10a, PGPS10]. EMFF [UNS^{+}18]. emission [DT10, ZLSL17]. emissions [AG16, ACTG17, Fen17, GWWH17, GZW^{+}19a, GZW^{+}19b, HK11, KS19, PR19, RIK16, Shi16, ZEP^{+}19, dlFRG17, Gri17, GZW^{+}19b]. empathy [NM19]. Empirical [MPPSdVSOHR14, BJO19, DPH^{+}15, DCC19, FFSF18, FP19, GGADSR17, JHW13, JSY^{+}19, JDQV13, LSP^{+}17, MK15, MSA^{+}18, MLR^{+}16, RWJ18, VRA^{+}12, Che10, MBSL13]. employ [CYI13]. employability [FFSF17a]. Empowering [CLDP17, HPH^{+}18]. Empty [GCCSS^{+}19]. Emptying [Cus17]. enable [ERB14, JOE^{+}19]. Enabling [OPMB12, LPC18, CRCD13]. Enacting [BMW17, Ols11]. enactment [Cho10]. enclaves [MJRdC16]. enclosure [Don16, FMN15, PD15, WK19]. encounter [GW15]. encountered [WHB^{+}18]. encounters [BSP19]. encourage [AMLK^{+}12]. encrasicolus [RRC^{+}17]. End [PGB19, CMS19, JLW17]. Endangered [HAGB15, DB19b, LHH15, MPT13, RBJ^{+}11, SBTF15, SDC13, SS16]. endemic [REA^{+}14]. ending [Liv10]. Energy [AVC18, Her18a, AWH13, ALMS14, CSVHS14, CGDB13, DS10, FMD^{+}12, GSR19, GCD^{+}15, Gue18, KCJW15, KLN17, KLMD15, LPC18, LGM16, MMOL15, NKAP13, NNP^{+}14, OL11, RST17, ROD15, RRMSSE16, SM17a, San18, SMGL17, SAGG13, Tod12, WMJF16, WWM^{+}19, Wri15, WOdG^{+}16, Wri16, YLZ^{+}19, vH19]. Enforcement [PJS13, BAMAH17, CMFG15, DCV12, Gol13, JBK15, KAF^{+}15, Kim19, MMTC16, Nor15, Por10, WMSB15]. enforcers [MC16]. engage [BRM^{+}17]. engagement [BdKvT19, CHL^{+}19, CBP^{+}18, FDGM16, GFC^{+}12, RJ13, SBTF15, Tan17, YY19]. Engaging [MWGD11, SSG12, WS15b, BYK^{+}15b]. engineered [KCM^{+}18]. England [DT10, BK16, BE15, BMAG17, DWCS18, EWG^{+}18, FBP^{+}18, FMB^{+}13, FJG^{+}14a, GKSJ11, HW10, LJ16, Por17, RCUR13, SF16, TC15b, THA15, TE16, TSW19, UA13, ZTAO13]. English [RDP^{+}18, CSvH^{+}18, CHSW17, FFSF17b, PO14, RMR^{+}17, RRRM^{+}15, RS17, Vau17, WCCO18]. Engraulis [RRC^{+}17]. Enhance [BSMS15, Bro16c, CS19, CCC^{+}14, Gau10, KEM^{+}18, RR14, SPO^{+}18, WF18, dlTCL10]. enhanced [PRS^{+}19]. enhancement [BSMS15, Har15, MPPSdVSOHR14, TFBS19]. Enhancing [EBS18, FDGM16, GDD13, OCG14, GCD^{+}11, CHDJ^{+}19]. enough [HAGB15, VGG12]. enquiry [Bha12]. ensure [BSAT18, Nik19]. ensuring [MSS18, OHB^{+}18]. entanglements [BGC^{+}17, MMB^{+}19, NM19]. Entering [Mar10]. entitlement [SC11, SL15b, SL16]. entrance [ALM17]. entrepreneurial [dGB10]. entrepreneurs [RO13]. entrepreneurship [GGK17]. entry [BH14b, Cro15, PDT^{+}14]. entry-strategies [Cro15]. envelope [SNDB^{+}16]. envelopment [GW11, VRT13]. environment [AAvH12, Bes10, Chu10, De 17, EBS18, GKG16, GGVRVL16, GSJ18, GGD16, HNC14, JBdL^{+}14, JSY^{+}19, KML^{+}16, LMG^{+}16, LGM14, MRCG11, OOvdM^{+}11, PWH^{+}16, PO14, PR13, RFL^{+}18, RFHEJK13, SJ14, TC15a, Tyn17, Tyn18, WMWH18]. Environmental [CSO17, De 11, DLM^{+}18, ECRL17, GFA^{+}18, JML16, LM16, MVRG13, MR10, NALO19, Ola18, UNS^{+}18, VRMF12, vD16, AVC18, ARS16, BDvTvKT17, Ben17a, BR13a, BS18, BRR^{+}12, Bre15, Bro16c, CLLS13, CGF14, CKC^{+}13, CE13, CKFLO19, De 10, De 16, DS10, DMJ^{+}17, DSH^{+}19, EAG18, ESM^{+}19, FE12, FLWG11, FLWG14, FDS15, Gau10, GC17, GLTB^{+}15, GC18, GCD^{+}11, GDD^{+}17, GDD^{+}19, Hav13, He15, HNVZ12, HCJ15, JTP18, JDM^{+}19, KC13a, KS18b, LEKB10, LJL^{+}14, MC15, MWR^{+}17, MK15, MRHP19, MOR15, MF12, MNG^{+}18, NKAP13, NBFO18, PHS18, RNS19, SO13, San18, SBMD13, SL15a, Sin19, SMW13, SPP18, SJ19c, SMS^{+}18, TGT^{+}19, TPKL16, WLX19, WCCO18, WCK^{+}17, WS15b, vLvT10, AVC18, CCC^{+}13, Gol13, RPG^{+}15]. environmentally [GNW19]. environments [BOF^{+}15, EMD^{+}18, EWG^{+}18, Gat18, GLTB^{+}15, LLF^{+}16, PPOM16, TR18, WRW^{+}17]. epistemological [Rog15]. equal [WBH^{+}15]. equality [SSRS14]. equilibrium [DRPSA17]. equitable [BAHCG14, CG19, LAH^{+}13]. equity [BTC18, CMMBV^{+}19, HCBM15, KVA18, OHB^{+}18, Pin13, YP17, Zag18].

Equivalency [WSGID17]. era [DMC^{+}17, KLC^{+}15, MB18a, MNG^{+}18, RLHM14, RFG^{+}14b, MNA15]. eradication [SHS^{+}14]. Erickson [Rah19]. Eritrea [HF16]. Erosion [GC17, DV17]. Erratum [HD11a, LGH^{+}12a, RLM12, Thi13a]. Escape [MJQMVÁ17, THM15]. Escaping [Knu17]. essential [HAGB15, MDR^{+}16]. establish [SvHdB^{+}15]. established [KGS^{+}19]. Establishing [GCG^{+}10, KDS^{+}15, MPCB14, WOdG^{+}16, DTN10, MSZS^{+}11, MHL13, TM13]. establishment [ALSC13, FGPP18, RNCGB15, Tan17, YYN^{+}15]. estimate [HAK^{+}19, SLL^{+}16, WS14, WML^{+}16]. Estimates [GPM19, PNPD14, PPM19, NWW^{+}15, SES^{+}19, WHB^{+}18, PPM17]. Estimating [BBR^{+}18, GMM^{+}13, PQLH15, PDT^{+}14, PB13, VRMF11, VS10, Wit19, vOHO^{+}17, KM10, PGP18, WMJF16, GSC^{+}16, SCOS17]. Estimation [WZCHS10, Cha16, Das16, DRG^{+}16, KS12b, KLY14, LGM16, OIS^{+}18]. estuaries [FMS17, HWN11, SBC^{+}14]. estuarine [SSD^{+}16, SGN^{+}17]. Estuary [BPMY18, GSS^{+}18, MPCB14, GOHDB17, GSDB14, KNA10, PV16, WSA^{+}15, BAO^{+}12, HMH18, WHP^{+}16]. Ethical [RPA17, HBK17, NPAR19]. Ethnoclimatology [DRMA18]. ethnographic [LCR16]. EU [BFGR14a, DAGG16, BR13a, BDG^{+}14, Big15, Boh19, Bor18, BFGR14b, CDA^{+}13, CKVO16, CRSM^{+}17, CBE^{+}19, Cor14, CK18, DBE^{+}16, Eli14, EFK^{+}19, GSC^{+}16, HTH13, JE11, KTD^{+}15, KBC^{+}13, LAH^{+}13, LGC16, LJ14, MML^{+}19, MM17a, MBM14, MMBK17, MFF^{+}18, MM14, NCH^{+}13, Nij15, Nij17, ØR18, PAA^{+}12, Pri10, PKM^{+}19, RMRN^{+}16b, RTD^{+}17, RST17, RWG^{+}19, SC19, Sal16, SMD14, SBSK19, SG12, SSWP10, Vat17, VPR^{+}16, dVM14, Wit19, ZT19, dVM12, vH19, vHHB14]. Eubalaena [HD11a, HD11b]. Eunicella [PGL^{+}16]. Europe [JKS16, Thi13a, AAvH12, BRR^{+}12, BMGF^{+}17, EOC15, FMD^{+}18, FQ13, GFMG^{+}14, HEJK10, LBVN19, LBB^{+}12, Mes12, MRSH13, MBM14, MNA15, NSB18b, PAD^{+}17, PVA^{+}18, QJ13, RSS14, SD13, SHS^{+}14, SBC^{+}15b, SH10, Thi13b, VLPC14, dGNC18]. European [CHSW17, ARDP17, Ant16, ACG^{+}17, BB19a, BH14a, CETQ17, CGDB13, CGC14, CE13, DRGP19, DSP^{+}18, De 10, De 11, DFV15, ECR^{+}16, FW15, FCC^{+}12, FSV^{+}16, FMMA^{+}15, FMGV16, FÓ12, Foc11, FGKMP14, Fri10, FH15, FWC^{+}18, GRLM15, GSR19, GLSA10, GGADSR17, GMD^{+}12, GGD16, GSS^{+}19, GCW^{+}18, GAC^{+}19, Had18, HPH^{+}18, HG14, HJKR14, KFPR10, LMA^{+}15, LP16, LJ13, LCC^{+}18, MM18, MDK^{+}16, MdMGV^{+}18, MCHO11, MCAM13, MM15b, Nij17, ÖSS^{+}11, Pas14, Pen18, PL12, PW16, PPOM16, RvLvTH14, RvT14, RDSB10, RWG^{+}19, RJ13, Ros17, RRC^{+}17, SLS18, STJ^{+}12, SFC^{+}15a, SSC^{+}15, SS16, SAH14, SSFL10, SRGGVL13, SRGGVL14, Tei18, Tyn17, VPR^{+}16, VS10, Vil10, VPP^{+}16, dVMdC^{+}19, Weg11, ZPB^{+}13, dVDA^{+}16, vLvHvT12, vTRvLvH14]. eutrophicated [KNB^{+}18]. eutrophication [Boh18, GNIL14, Jet18]. evaluate [ASWM18, GSG^{+}15, QSICTdV19]. evaluated [HNVZ12]. Evaluating [BL18, BPMY18, BRRG^{+}18, CBL18, CAP15, FOO^{+}15, FBM18, GR16, JSD19, KJWP14, LCH^{+}13, OPGL16, OLF19, SVL^{+}16, ST13, SLBC19, TSH^{+}19, TAGP15, ZMZ18, HD11a, HD11b, SKS^{+}19, YYN^{+}15, BJS^{+}16, SBSK19, WGB^{+}16, Zha18]. Evaluation [BO12, Car13a, FP12, JCD17, PCCR15, PBSG12, SFCI14, Cha16, CCL^{+}19, CGS^{+}17, COS14, CBP^{+}18, CSP10, FMB^{+}13, HWY^{+}18, HHP17, HSW17, JWBJ13, KGST14, LKL15, LH12, SB14, SBS^{+}13, SFC^{+}15b, TL16, WCM12, YFC13, YCHM14, AGSP^{+}13, FP13]. evaluations [DJTK17]. evangelism [Mac14]. events [RTM^{+}19b]. Evidence [FTF^{+}17, ITNL^{+}14, ZKR^{+}19, AMR14, Car11, Che10, GRYA^{+}15, GBF^{+}12, HL10, HTA^{+}15b, MBSL13, AYB18, AYA^{+}19, BG15, EAE15, Lim16, MNHN18, NAB^{+}18, SB18, ZLL19]. evidence-base [GRYA^{+}15, HTA^{+}15b]. Evolution [SY18, Bur15, DMKL18, DSH14, NWW^{+}15, Sch12, SRSN17, SK15, Sin19, SSM15, YFR10, GAC^{+}19]. Evolving [MM18, FJG^{+}14a, MBB18, Mal13]. ex [GM14, SKFL15]. ex-vessel [GM14, SKFL15]. Examination [KSC^{+}19, FFSF17b, FÓ12, SDHS19, VKA^{+}19, YFB16]. examine [RFG^{+}14a]. Examining [Gus18b, WM19, MC19, SMH17]. example [DTJ10, EE17, GS17, JG13, LB13, MR10, Ols11, SARF16, Sch10a, WNB^{+}15, Wit15b, Wit15a, dCW15, SMBL13]. Examples [Har10, VPC19]. exception [RBWL11]. Excess [RYJ^{+}17]. excessive [BE15]. exchange [CHvKM15, PMM^{+}17, TWJ^{+}17, Cha19]. exchanges [BRM^{+}17, GLR17, JTBP17, THPJ17]. excluder [DBD19]. Exclusion [FHL18, Bod17, MMDS12, Tam15]. Exclusive [Mul19, RBS^{+}19, SWS12]. exclusivity [MP12, Tei18, Web11]. existence [OL10, Var11]. Existing [JDM^{+}19, ARGW14, ACM^{+}12, CDR^{+}16, FJA^{+}14, OL11, SLJ^{+}11, SvHdB^{+}15, SGMWE14, TvN13, YSJU^{+}17]. exit [BH14b]. Expanding [WS15a]. expansion [BBRT16, CG19, GHO^{+}17, PS12, TBB12]. expectations [HMK19, JGG^{+}16, KFBG13, KH10]. expected [KMTS16]. expenditure [CA17]. experience [AT14, Cap10, Che12c, CC17b, FMB^{+}13, Han13a, Han14b, Hee14, KGAB14, LJS^{+}18, LWS19, MM18, MM19, Pap16, PO14, PFF^{+}18, SDF^{+}14b, Shi13, TGdCR11, Wis14]. experiences [ADH18, BPF^{+}16, Eli14, HD10a, JE11, SGN^{+}17, VM18, FGPP18, FJA^{+}14, GBBB15, Har15, HID12, HHA17, HC10, JKS13, NC10, NW10, SCS16, TK14]. experiment [DFV15, KM15, Man10, PNW18]. Experimental [Dva19, HR16, Mox12, Sun12]. experiments [BL18, EA19, HTH13, LLF^{+}16, LPU^{+}16, TI13]. expert [CCK19, HK10, Kep14, KTD^{+}15, LM16, WML^{+}16, BB19a]. expertise [BSAT18, Dix16]. Explaining [FSM^{+}12, San10, San11, Sev13]. explanations [AOY16]. explanatory [BRR18b]. explicit [CGM^{+}18, RFHEJK13]. Exploitation [ISM^{+}17, GBBB15, HHS16, HBA^{+}14, Hor10, ONF15, RCMBG^{+}19, WLP10, WDK^{+}13, ZZ16, BSL19]. Exploited [SMQ^{+}16, Coc15, FDS15, FBAK14, Hil10, HAGB15]. Exploration [MBCR16, BD13, GSW18, MPPB10, SGLY13, ZZ16]. explore [LGM16]. Explorer [GRW^{+}16]. Exploring [BFL18, CN11, DL19b, FMB^{+}18, GOHDB17, GdV16, JWBJ10, KGS^{+}19, LSP^{+}17, LRMM17, LJ14, MMJM16, MJS16, MBBDS19, Ros13, TKN19, YY19, BDB19, DR10, DMdJ^{+}15, GCK^{+}11, Has18, HJN16, KP13, KL14, KR14, LLN17, MPPB10, NPS^{+}19, TR18, Tos11]. Exponential [Pas14]. export [FKWV19, Nij15]. Exporting [CBP^{+}16]. exports [She16]. expression [ZPT^{+}16]. extant [HAGB15]. extensive [She16, TRCR^{+}19]. Extent [CMCMHP13]. external [MM14, SCC14]. externalities [ARS16]. extinction [CSH17, Sch15a, SGW^{+}15]. Extracting [CF17, PGC19]. Extraction [EA19]. Extractive [NACS17, NGS16, RFHEJK13, RFKEJ^{+}15, SB15]. extreme [CLLS13]. eye [AGB13, BPBM17, vS15]. eyes [LMP^{+}16]. Eyre [LLS^{+}15, PR13].

F [HS19]. face [BF19, DCY^{+}19, EGHC17a, MKB17, MOR15, OR19, PS10, RDG^{+}13, SLBC19, WP13]. faces [TLR^{+}10]. facilitate [FB15, MF10, NBFO18, ÖGB^{+}10]. facilitating [BHS19, EBR^{+}18, VAW14]. facilities [JA10]. facility [HHL^{+}18, LGM16]. facing [Bro16b, HCC^{+}19, KMD15]. Fact [FVM14]. factor [BRR18b, DAJM15, SAdC15, TWA^{+}15]. Factors [AK15, BH14b, Che10, DP17, DCC19, HD11a, HD11b, LYB17, MGG^{+}19, OM19, QSMGVA18, RGG^{+}16, SOUN16, Bre13, CN11, CZX^{+}19, ETEWZD11, GM14, KH15, LQC10, MKB17, NKBS17, VAW14, WHMG^{+}16, WWT^{+}16, WP12, vPLB^{+}13, ALSC13]. factory [FFKO15]. FAD [HHS19, MBvZ15, MKB17, MHM16, TZBBL18]. FADs [DMMG14]. fail [SBSK19]. failed [Vat17]. Failing [CSP10, SAH14]. failure [Ben17b, DA18, KFPR10, PE10, Tam15, DSR12]. Fair [BB19b, BAHCG14, BB19c]. fairer [FJ12]. fairness [HCBM15]. Falkland [ADL^{+}18]. fall [FvRA11]. Falmouth [KMvT15]. false [HdF^{+}19]. Family [vdVD16]. fans [PGL^{+}16]. FAO [HFF11, KSC^{+}15, PDC10, PZ17a, Bra12, Gar12, PF12a, PZ17b, PZ19, YBB^{+}17]. far [CMW10, He18, ZZ15a]. farm [BN19, RDP^{+}18, VLPC14]. farmed [KTA^{+}13, LOS11, PS11b]. farmers [GGK17, KKWS15, ZMZ18]. farming [AYR^{+}14, CS14, CJ17, FHJ16, GA14, HBG12, HMK19, LFGM15, LKC15, LJT19, MBT^{+}19, MR16, PR13, RVdCGN11, SMH17]. farms [BDF^{+}10, CLV^{+}16, GOHDB17, GSG^{+}15, HAA15, HHA17, SS19a, THM15, VLKN17, WKB15, vdBKB^{+}17]. Faroe [DA18]. Faroes [TÁB^{+}10]. Faroese [Jac19]. fatal [MSB15]. Fatalities [HUHA18b]. fate [FSCJM19, FVT19]. fauna [RHP^{+}17]. favorable [Has18]. favour [MCH19]. FC [Cho10]. FCM [GOO^{+}19]. fear [PNW18]. Feasibility [HBG12, BGA^{+}15, KYS10, WPdR^{+}14, ZZ15b]. features [GFH^{+}17, Lyo14]. February [Ano14o, Ano15q, Ano16o, Ano17o, Ano18v, Ano18n, Ano19u]. federal [CR10, DB19b, DFSB15, RYRYM17, SJ15, GBF^{+}12]. Feedback [Bur15]. feedbacks [IT15, ÖSS^{+}11]. fees [ALM17, PDT^{+}14, Thu10]. feet [CJ16]. FERC [Kai17a]. ferries [RST17]. fertilization [GPO^{+}10, MR10]. few [HBL14]. fiction [FVM14]. field [EA19, SJ17]. fifty [KJH^{+}10]. fifty-year [KJH^{+}10]. Fiji [Bol17, DK12, DGVS^{+}16a, FFH18a, GAKP^{+}19, HGS12, KJWP14, KJP14, PR19, Sea17a, Sea17b]. Fijian [SC16]. Filipino [AT14, MM19, TZBBL18]. Fill [CL13]. Filling [WFK^{+}12b]. Fin [TINL^{+}14, GCP19, HS19, JLW17, PDC19, ST17, SH17, dMAH^{+}18]. Fin-Fish [TINL^{+}14]. finance [VMMB^{+}19, WB19]. Financial [KLM^{+}13, Bro18, CCC^{+}14, FDP18, GFH^{+}17, GS16, PVK18]. financialization [KN17]. Financing [ALM17, CCC^{+}14, MMTC16, Thu10, VMR^{+}18, ZM15, RD12]. Finding [LHJ^{+}15, PHC14]. findings [JLQ16, Sun12]. fine [DMdJ^{+}15]. fine-scale [DMdJ^{+}15]. finfish [BBAP^{+}17, EOC15]. Finland [HHL^{+}15, LM11, Sev13, SBSK19]. finning [JLW17]. Finnish [NALO19, PVK18]. fire [vS15]. First [NM12, BHT^{+}18, CCS16, FG18, FMP^{+}18, NY13, OBJ^{+}12, SJ19a, WTD^{+}15, ZP16]. Firth [BMGH11, JKS16, San18]. Fiscal [MSS18, MB18b]. Fish [AGB13, BPYC19, BBRT16, HB19, KKWS15, SCD^{+}16, TINL^{+}14, TCC^{+}19, And15, BIPS13, BKKN10, BAA^{+}15a, BSH^{+}19, BTS16, Bre13, CDA^{+}13, CSH17, CA16, Cha14, CRO12, Cou11, Cro15, DMMG14, FD11, FBAK14, GBW16, GRLJ15, HH10, Han13b, HPP^{+}17b, HD10a, HID12, HGB^{+}19, HC18, HO12, IH18, IW19, KGAB14, KKG16, KW16, MACM14, MC19, MRSL16, MDR^{+}16, MM15b, MMDS12, NMvP19, OL10, OHQ^{+}16, OSS14, PTN^{+}17, PG19, PML10, PS12, PS14, PJS13, PNPD14, RCUR13, RL14, RW15, SO13, San10, San11, SBM^{+}17, Sev13, SS16, SALS17, SGLY13, SGZI19, SMQ^{+}16, SKT^{+}10, THM15, TCN^{+}13, TRC^{+}17, VvDM10, WCMHO18, WCCO18, YFB16, ZT19, ZMZ18, vdVD16, Hor10, LLN17]. FISHDA [MMC^{+}14b]. fished [KTA^{+}13]. Fisher [ASBAM13, BR18, Bod17, CH16, Cou11, Gre15, GJA16, HIN^{+}17, HB14, MGG^{+}19, MMC^{+}14a, NAB^{+}18, RPdO^{+}19, SC11, SM17b, Wit19]. fisher-developed [CH16]. Fisheries [ABCDMG17, BBE19, BJS18, BRM^{+}17, Bro16b, Bro18, BGL^{+}16, Car15, CGP^{+}18, CMFDCLV10, DCV12, DA18, Das16, EHR15, ETEW15, Fla10, FSMG10, Foc11, FQ13, GGADSR17, GAFD^{+}18, GDG^{+}18, Gre15, HG14, Hol12, IW19, JFA17, KFPR10, Kim10, KAPF^{+}19, LT19, Mal16, MV17, MMBK17, MPMK16, ODF^{+}16, Pas14, PF12a, PZ17b, Por10, QSHI17, RDSB10, Rei10, RLHM14, SMD14, San15, SRGD18, SH10, TN15, TI13, VSRD16, Zha18, vH15, AAvH12, ARSHH16, AARK15, AMG16, ASP16, ATB^{+}19, AL17, ADGGVL12, ASWM18, AMdB12, ACA^{+}13, ACR18, ABB14, APC16, ADHV16, BSD^{+}17, Bai16, BdKvT19, BRR18b, BFTS^{+}16, BDIS18, BMW17, BLG^{+}18, BLC16, BSL19, BCMH^{+}18, BAA^{+}18, BSSS17, BBA15, BHAHW17, Ben12, BDF^{+}10, Bes19, Bes12, BK16, BLR15, Bla17, BDHSJ19, Bla15, BYK15a, BY16, BW18a, BWD^{+}19]. fisheries [BW18b, BCE^{+}17, Boh19, BD10, BVB18, Bor10, Bor18, BAMAH17, Bra12, BM15, BAT^{+}18, Bro16a, BSP^{+}15, Bro16c, BAY^{+}19, BHT^{+}18, BF19, BBvZ^{+}17, BMGH11, BHG11, CCR16, CACL18, CSVHS14, CLF^{+}19, CdLH^{+}18, CPR14, CDA^{+}13, Car14, CBP19, CMAA13, CC13, CRSM^{+}17, CN17, CSF^{+}19, CCS14, CLS10, Cha11, CLLS13, CYI13, Cha14, Che12b, Che12c, CC17b, CdlPS^{+}14, CGS^{+}17, CRB^{+}19, CBF^{+}12, CST10, CSCBR10, CMCMHP13, CMS19, CBG14, CAP15, CHSW17, CP15, CF13, CGC14, CdL17, Cor14, CE13, CBS^{+}16, CSMR18, CSP10, DR10, DWCS18, Dav15, DL19a, DK12, DAJM15, DMdJ^{+}15, DK15, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, DYN^{+}19, DCHC17, DMW19, Don17, DDV^{+}13, DT10, Dru10, DS17b, DBD19, DCY^{+}19, DFK13, DK16, DMB^{+}18, DFL15, EBR^{+}18, EE17, EGGT12, EGHC16, EGHC17a, EGHC17b, ENWL19, ECL^{+}15, ELAC19, ETEWZD11]. fisheries [EYN^{+}15, FR16, FPG^{+}17, FB15, FMB^{+}18, FDP19, FV10, FDGM16, FGR18, FWC^{+}18, GNHM18, GFMG^{+}14, GLR17, GMD^{+}12, GMD15, GHM^{+}18, Gib10, GS17, GTL^{+}18, GOTS19, Gil11b, GDD11, GDD13, Gla19, GTM^{+}12, GH15, GFH^{+}17, Gol13, Goo17, GMP^{+}16, GA19, Gra10, GGL17, GZW^{+}19a, GZW^{+}19b, GW11, GBF^{+}12, GMM15, Gug19, GMM^{+}13, GSC^{+}16, GBBB15, GAB^{+}17, GJA16, HES16, HEJK10, Had18, HWL^{+}19, HTT10, Han12, HCBM15, HWO^{+}18, HZH^{+}13, Har15, Har10, HMGR14, HDM13, He15, HHL14, HPB^{+}15, HSH14, HGB^{+}19, HPW^{+}18, HRC15, HW10, HRJH15, HLRO19, HL10, HA14, HAA15, HNVZ12, HvPN^{+}19, HK18, HO10b, Hua11, HGBJ16, IT12, IH18, IHF15, ISM^{+}17, Jac19, Jac15, JGG^{+}16, JA10, JMJ^{+}18, JHN^{+}19, JTBP17, JCBM10, JSD19, JE11, KP13, KNA10, KL12, Kar17, KL14, KML^{+}16]. fisheries [KCKA^{+}18, KLY14, KSS17, KR14, KVH14, KW14, KLM^{+}13, Kna11, KN17, KKG16, KBC^{+}13, KVMM10, KCCO18, KKSC19, KBBR16, KNY14, Lan15, LDF^{+}18, LMA^{+}15, LQC10, LM14, LY14, LRJ^{+}19, LB17, LP16, LBK^{+}13, LPU^{+}16, LDVM10, LR15, LLN17, LBL^{+}17, LH16a, LZL19, LJ13, LJ14, Liu13, LCC^{+}18, LRS^{+}13, LV18, LH16b, LKMN16, LBC^{+}16, MBD13, MHC19, Mac14, MWGD11, MM18, MFP18, MLF^{+}17, MG19, MML^{+}19, MCMB^{+}18, Mar17, MB10, Mar10, MV16, MMdJ^{+}14, MNLHRMP17, Mat14, MJQMVÁ17, MdMGV^{+}18, MRCG11, MMTC16, MHA^{+}17, MCY^{+}19, MCH19, MNHN18, Mes12, MBM14, MPPSdVSOHR14, MPMGGAHR18, MSS18, MLA16, MDR^{+}16, MCB^{+}19, MMS^{+}14, MHM16, Mor17, MN18, MAHP16, MFAA16, MK14, Mso16, MMC^{+}14a, MMC^{+}14b, MYA^{+}18, MM14, MCBM18, MLHP15, MMA16, NSB^{+}18a, Nak16, NC10, NHFV13, NCH^{+}13, NWW^{+}15, NRN14]. fisheries [NHA15, NRB^{+}18, Nol19, NKBS17, NLP11, NBC^{+}12, NBFO18, OC17, Oan18, OCG14, Ols11, ÖSS^{+}11, ONF15, ODL^{+}13, Pan14, PQLH15, PH16, PRG^{+}17, PDS15, Par15, PDB^{+}13, PIH^{+}10, Pat15, PB13, PL15, PZ17a, Pau18, PMM^{+}17, PS11a, PMN^{+}15, PKOF18, PRPLBLC12, PC13b, PDMÖ19, PCK^{+}19, PHC14, PS10, PS18, PSH18, dlÁPAVS19, PGP18, PS11b, PD15, Pin15, Pin17, PPT10, PAD^{+}17, PVA^{+}18, PVPF19, PFMA^{+}19, PWB19, PSV^{+}14, PVK18, PGPS10, Pom12, PPR^{+}15, PPML16, PPCM16, Pom16, PC18b, PGP^{+}19, Por17, PMA^{+}17, PM10, PAA^{+}12, Pri10, PKM^{+}19, PLP14, PCLE17, PWN18, RLAT19, RMRN^{+}16a, RROR12, RD12, RYRYM17, RCUR13, RGWH18, RWG13, RMB^{+}17, RCD^{+}18, RBWL11, RBTR^{+}19, RUD^{+}12, RV16b, RRRM^{+}16, RGSF19, RPC^{+}17, Ros13, RBS^{+}15, RWBF19, RTM19a, RCMBG^{+}19, RD13, SDRML15]. fisheries [SvZdJ15, SC19, Sal16, STIC19, SMBL13, SBC^{+}15b, SRA15, Sat18, SSW^{+}18, Sch17, SCC^{+}17, SBP^{+}19, SVL^{+}16, SGNM19, SBM^{+}17, Seu16, Sha17, SSLR14, SH14a, SH17, SAM13, SSC^{+}15, SMR^{+}15, SNA^{+}17, SS16, SW10, SJ14, Smi13, SKS^{+}19, SNP^{+}18, SCJ13, Son15, SSS^{+}17, SWM11, Soo17, SC10, SSRS14, SV13, SCC14, SWS12, SS15b, SH15, SAH14, Stø19, SH19, SDB14, SLL^{+}16, SES^{+}19, Sun12, SH14c, SRS14, SJ10, SG12, SCS14, TSH^{+}19, TFBS19, TRM19, TCS^{+}18, TTS14, Tei18, TSK18, TLB19, TBJC18, TEF^{+}15, TWJ^{+}17, THPJ17, TH15, TWA^{+}15, TLA16, THE^{+}18, TLA18, TÁB^{+}10, TMDSF19, Tip16, Tir17, TJJAM19, TII^{+}19, THA15, TM13, TdCdSR16, TO10, THK^{+}18, TPS15, UAZ13, Var11, VAP^{+}10, VRT13, VPR^{+}16, VC17, VvD10, VCdR^{+}18, VPP^{+}16, VWG19, Vra19, WSB19, WW17]. fisheries [WFSV15, WPJ10, WBH^{+}13, WCM12, WBC14, WP17, WIC18, WP13, WFK^{+}12a, WNB^{+}15, WZCHS10, Wit15b, Wit15a, WBSW13, WAG^{+}15, YCA^{+}12, YP17, YBB^{+}17, YTSO15, ZVP^{+}13, ZP18, ZTAO13, ZH14, ZEP^{+}19, ZM15, ZH18, dBMA13, dVDA^{+}16, dlTCL10, vAS19, vdMCJ17, Bro15, CPB19, Doy15, HFF11, KAPF^{+}19, KSC^{+}15, KSH^{+}15, OPMB12, PCS14, PDC10, PHAC15, PFF^{+}18, RMRN^{+}16b, SALS17, SH14b, STB16, vHK17, MMM19, HS19, Par10, PDC19]. Fisheries-dependent [LT19]. fisherman [DHR16, SL15a]. Fishermen [SJT10, ADGGVL12, AYA^{+}19, ATB^{+}18, CC17a, CSMB^{+}12, Dav12, DRMA18, FC19a, GLVLGG19, HRLFMP15, KHD12, MMP18, MFRvP16, MC16, MMS^{+}14, MMA16, PC13b, PBSG12, ROD15, STM16, SSD17a, SSD17b, SL15a, Tho13, ZFG12, Zha16, vdGMA^{+}13, AM15, AdV18]. Fishers [EB16, GC18, dO13, AWH13, BAA^{+}18, Bre13, CLS16, CMW10, DSM^{+}13, Deb18, DP17, DZ12, ENWL19, GDD13, HR17, HAA15, JBJ12, JTK18, KGAB14, KR14, KKWS15, LSTD12, LBP^{+}19, MBvZ15, MKB17, MCH19, MA15, NM19, Nei15, NBPF^{+}12, PRBX19, PMM^{+}17, PM14, PTP13, RAZ16, RTM19a, RCMBG^{+}19, SD18, Sch15b, TTS14, TSW19, WBC14, WWT^{+}16, YFB16, dAdOS17, dVvT11, AWJ19, CST10, FV10, PPT10, UA13, FB17, KRM14, LAP^{+}12, LBHS16, MACM14, MdMGV^{+}18, SGNM19, VAW14, vS15, Che12b, CR11, GGADSR17, GSJ18, GAKP^{+}19, HIBG15, MMM19, MMJM16, NKHL18, PHEM^{+}17, SCR^{+}19, TMG15, TKM^{+}18, TJ13, NSB^{+}18a]. Fisherwomen [RLAT19]. Fishery [CSK^{+}18, CJHR14, HvD13, Har10, KMTS16, Kru16, LM15, Man10, MLHP15, SKS^{+}19, SCS14, TGBL11, WPdR^{+}14, AZAMG13, AAL^{+}18, BAHD10, BFL18, BTS^{+}17, BTS16, BSSS17, BHAHW17, BV19, BSKH17, BK16, BDSB16, BB19b, BB19c, BLR11, Bro10, BH14b, BMGH11, BHG11, CVFP^{+}12, CdLR^{+}15, CH12, CH16, CHBRJ19, CPKF19, CVP^{+}17, CHKG16, CBL18, Cha16, Che10, Che12a, CCL^{+}19, CO16, CMSM^{+}16, CKdS14, COS14, CSS15, CWP16, CS13, DAGG16, DRPSA17, DRGCPS17, DMMG15, DJ15, DSR12, DPV^{+}15, DPJ^{+}13, DRH^{+}18, DPH^{+}15, Dra15, DT10, DCY^{+}19, DHR16, EAE15, EBR^{+}18, EP19, ECF^{+}18, ERD^{+}13, ES18, FGW17, FC19a, FvRA11, FMN15, FV10, GMD15, GOK14, GOM17, GMM^{+}13, Han14a, Hav13, HCP^{+}17, HPW^{+}18, HB14, HKR19, HK18, HSB^{+}16, HLL^{+}18, IT15, IHS13, ISM^{+}17, KSB15, KMF^{+}18, KBG^{+}18, KGMP18]. fishery [KS12d, KSC^{+}15, LEM^{+}18, LSMW10, LYB17, MBvZ15, MC15, MSB^{+}18, MOR15, MHK^{+}10, MKSS18, MFF^{+}18, MSG18, MI15, NMvPNL15, NBF^{+}18, O’R15, PGC19, PC19, PIH^{+}10, PWR17, PHC14, PM14, PFCP^{+}14, Pin13, PC18a, PHEM^{+}17, PJS13, PMC^{+}15, PRPS17, QSMGVA18, RGG^{+}16, RRAH13, RPdO^{+}19, RCdLS13, Rei19, RLM10, RLM12, RLH18, RNEBG19, SvZdJ15, SARF16, STM16, SCMS15, SKFL15, SRA15, SD18, Sch12, SS11, SK18, SBMD13, SMR^{+}15, SPRH19, SAdC15, SGP^{+}17, SC11, SM17b, SF16, SH16, SJT10, SJ10, TINL^{+}14, TC15b, TOL12, TMDSF19, Uhr16, VHRM17, Var11, VRB14, VPL^{+}13, Vil12, VKA^{+}19, WK14, WHMG^{+}16, WEKH19, WHP^{+}16, WC17, WSG18, WRP^{+}10, WCCO18, WBH^{+}15, Wos19, YFR10, YFC13, YCHM14, vPLB^{+}13, vS15, BSMS15, EAA^{+}18, NBMB^{+}18, PW15, RG19, RYJ^{+}17, SSFL10, TINL^{+}14]. Fishery-based [TGBL11, BAHD10]. fishery-dependent [PFCP^{+}14]. fishery-independent [DPV^{+}15]. Fishery-seal [CJHR14]. Fishes [HTF^{+}15, BB19a, BSW19, Hil10, KV18, RLAT19, SCV^{+}15, SDC13]. Fishing [BGA^{+}15, BSSS17, BJO19, CSMB^{+}12, DJ15, FR16, GAKP^{+}19, HJKR14, HC10, JG13, KP13, KB15, LPP17, LDVM10, LH16b, MDE14, MÁSSI13, MK14, MMC^{+}14b, NYJW15, RR18, RGB^{+}19, SSRS14, SSA16, SM17b, SBT16, UEMRG^{+}18, Vau17, ZSH18, dlTCL10, ASP16, ATAM10, AOAM13, Ant16, AMdB12, AM15, AG12, ABB14, Ben17b, BO12, BC13, BAT^{+}18, BDH^{+}18, BRR15, Bro16c, BH14b, BHT^{+}18, BRW^{+}13, CMVLGG14, Cam15, Cap10, CdLH^{+}18, CSF^{+}13, CPKF19, CBL18, CKN19, CGDB13, CL13, CC17b, Cho11, Cho12, CSCBR10, CGM^{+}18, CJW^{+}16, CA17, CGC14, CBP^{+}16, CJ16, DSH14, DMdJ^{+}15, DRMA18, DP17, ECF^{+}18, EB16, EJB19, EA19, FSS14, Fla13, Fol13, GvdB13, dlFFVRC16, GGSRVL17, Gil11a, Gil15, GKSJ11, GZW^{+}19b, GS16, Gus18b, HBG12, HAH^{+}16, HHS16, Han11, HSRH16]. fishing [HD16b, HSW17, ISHP14, IMA16, JWBJ10, JWBJ13, JTK18, Jen15, JHW13, KMC^{+}17, Kar17, KBF12, KGMP18, KLY14, Kim19, KR14, KDW10, Knu17, LVRM^{+}18, Lax10, LGH^{+}12a, LGH^{+}12b, LAH^{+}13, LM15, LGC16, LH12, LT17, LCR16, MJvP^{+}18, MKB17, Mal13, MML^{+}19, MGA^{+}10, Mat16, MJRdC16, MC19, MS14, MFRvP16, MPPSdVSOHR14, MGG^{+}19, MCB^{+}19, MLTBMB14, MSC18, Mul19, MSGM^{+}12, MI15, MMB^{+}19, NSB^{+}18a, NW10, OPGL16, OY19, OIS^{+}18, OLF19, PC19, PS11a, Ped13, PL12, PC13b, PSP10, PVK18, Pol12a, PFP17, PRS10, Ray19, RRG^{+}12, RAG18, RNN^{+}18, RL19, Ros17, RBS^{+}19, RTM^{+}19b, SARF16, SKFL15, SMBL13, Sch15a, Sen10, SG17, SRK14, Sha10, SNIN17, SDP^{+}16, SGW^{+}15, SCC10, SNA^{+}17, SWP13, SS19b, SPS14, SWB^{+}14, SMS^{+}18, SLBC19, SRGGVL14, TCS^{+}18, TS14, TC15b, TMHJ12, TJ13, TLH18]. fishing [TPS15, Tur10, UL17, VRMF11, VRMF12, VG15, VPL^{+}13, VCdR^{+}18, VS10, Vil10, VKA^{+}19, VBMM17, WT18, Yat14, Yos17, You16a, YFB16, dVM10b, vHCT^{+}16, vPG10, vPJH^{+}17, FVM14, LPM^{+}14, NM19, Sat18]. fishmeal [PML10]. fishy [TMR^{+}14]. fit [AAL17, DRH^{+}18, GSS^{+}19, PAA^{+}15]. Five [BCM^{+}18, BG15, FH15, GDD11, KS13, MMC^{+}14a, YBD^{+}19]. fix [RR17]. fixed [CLC^{+}16, Das16, RRC^{+}17]. fixed-point [CLC^{+}16]. fixing [LRS^{+}13]. Flag [MS14, PKPdP10, Sch17]. flags [CLLL17, FW19, RMJ12]. flatfish [dVvT11]. flawed [GPB^{+}18]. fleeing [GOO^{+}19]. fleet [ABB14, BH14b, CMVLGG14, CFC^{+}18, CE13, FGW15, FFKO15, GMD^{+}12, GKSJ11, GMM^{+}13, HCP^{+}17, MMdJ^{+}14, MS14, MFAA16, OCG14, OCG15, PA10, PVK18, RGBT16, SS15c, SSA16, TC15b, VRMF11, VS10, Vil10, ZSH18, OCW^{+}17]. fleet-based [GMD^{+}12]. fleets [CGDB13, DRGP19, FPP^{+}15, dlFFVRC16, LVRM^{+}18, PL12, RAG18, SBM^{+}17, SKT^{+}10, vHCT^{+}16]. Flettner [Sea17b]. Flexibility [GH15, MH13]. flexible [EFK^{+}19]. Floating [KNB^{+}18, BGA^{+}15]. flood [MSA^{+}18]. Florida [SBP^{+}19, Uhr16]. flounder [BHG11]. flow [Sea17a]. flows [BCP19, PBJ^{+}14, WNLS17]. fluids [dlFRG17]. flying [HO12]. flyingfish [LC19]. focus [LG19, AARK15]. focusing [BPMY18]. fold [Edd14]. follow [MBSL13, PVA^{+}18]. following [Cla15, CSMB^{+}12, ENWL19, GSS^{+}18, Nij17, SMH17]. Food [FAA^{+}13, PSH18, dVMdC^{+}19, vHFJ^{+}19, BAA^{+}15b, BAA^{+}15a, BCMH^{+}18, BAA^{+}18, BSH^{+}19, BF19, CG18, CSP^{+}14, DGVS^{+}16b, DCHC17, EZBL19, FD11, Fol13, GAKP^{+}19, HZH^{+}13, He15, KMTS16, LGH^{+}12a, LGH^{+}12b, LLN17, OCdS14, RLAT19, SC10, SNTL18, TSK18, TCC^{+}19]. footing [ODRRM12]. footprint [CSF^{+}18]. footprints [MM15a]. forage [HH10, Han13b]. foraging [SAGG13, SBC^{+}16]. force [Kim19]. Forced [CLS16]. forces [AKV15, AL18, CLS10, AL18]. fore [KSH^{+}15]. forecast [Kai15b]. Forecasting [ASB10, OCW^{+}17, dVM17, KL12]. forecasts [SJ10]. Foregone [Lan15]. foreign [CGP^{+}18, Hea13, RMJ12, SBM^{+}17]. forerunner [Bac18]. foreshore [Bes11]. forest [ATDP16, Roy16]. form [BAT^{+}18, Fri10, HW10]. Formal [EA19, PRBX19, CSCBR10, HIBG15, JSD19]. Formalizing [HK10]. formation [GM14]. Formosa [GNIL14]. formulation [KPR^{+}17]. forsteri [CKBS14]. Fortress [DJM11, Har14, She14, Gam11]. forward [ACM^{+}12, CBN^{+}12, DCF^{+}19, GJ15, Sch15a, ZZ16]. fossil [ABBD17]. foster [MWR^{+}17, Pla19]. Fostering [Cra19, BPF^{+}16, ACG^{+}17, JAO^{+}16]. fouling [FSV^{+}16]. four [LWvdBL16, MDK^{+}16, SJ19a, Yia18, Zag18, dVDA^{+}16, BGC^{+}17]. four-step [LWvdBL16]. Fracture [VVS18]. Fragmentation [Bla17, CWP16]. Fragmented [TK11, Chu10]. framed [EA19]. frames [MSBR15]. Framework [AGB^{+}18, ADL^{+}18, ACG^{+}17, BJGMG18, BR13a, BDG^{+}14, BFGR14a, CETQ16, CBE^{+}19, GNIL14, KTD^{+}15, LJB^{+}14, LLO17, MKU19, PRO16, QSICTdV19, RDSB10, SGGM14, dVM12, vHHB14, vLvHvT12, vLRvH^{+}14, AF14, APC16, BDIS18, BLG^{+}18, BSW16, Big15, Big16, BCE^{+}17, BHS19, BO12, BFGR14b, BMGF^{+}17, CMP^{+}15, CGM^{+}18, CPB19, DM15, DDS^{+}11, DMJ^{+}17, DSH^{+}19, Eng18, FMM13, FSGB^{+}15, FSMG10, FJG^{+}14a, FGKMP14, GSG^{+}15, GMM15, HKHG19, HWN11, HHP17, HBK17, HMS^{+}17, HdVB13, HDTS14, IMA16, JJ19, JDQV13, KCKA^{+}18, LAH^{+}13, LLO16, LCR16, MDG^{+}14, MNM10, MHA^{+}17, MCB^{+}19, NWD18, OL11, ODRRM12, Par15, PL15, PBW16, PBW17, PCK^{+}19, PM14, Pom16, RM18, SOCA15, SCKP19, Sen10, SI18, SBS^{+}13, Thi13a, Thi13b, Wri16, YTTK10, YFR10, YCY^{+}15, ZPB^{+}13, ZZ17b]. frameworks [DHC11, ECRL17, EGHC17b, TC15a, WL18]. framing [LEG19, OO17a]. Franca [MVRG13]. France [GA14, FB17, LJB^{+}14]. fraud [DL19b]. Free [HW10, TZBBL18]. French [TFBS19, BLG^{+}18, CBG14, CHL^{+}19, DBB^{+}14, LACG18, MB15, TGdCR11]. Frequent [MC14]. friendly [SSV18, SCOS17]. friendship [STG11]. Froese [AGSP^{+}13, Bra12, FP13]. frontier [DDM^{+}19, MDE14, Nol19]. fronts [MC14]. Frøya [THRS15]. frozen [VRMF12]. fry [IW19]. Fuel [DT10, PR19, ABBD17, CGDB13, GZW^{+}19a, GZW^{+}19b, HBC^{+}13, IHF15, LH12, RU12, Sea17b, ZH14, ZEP^{+}19]. fulfill [ZPB^{+}13]. fulfilment [KP17]. Full [CCS14, HBG12, RMRN^{+}16a]. full-cycle [HBG12]. fully [vHCT^{+}16]. function [Cha16, FGW17, SBC^{+}14]. functional [BM15, DFK13, EGF^{+}17, HWL^{+}19, LLX^{+}15]. functions [LPS12]. fundamental [Son15]. funded [HSW17]. Funding [Gug19, BCP19, BWD^{+}19, MCY^{+}19, PWB19, WWW10]. funds [Ber19, MPMGGAHR18]. Fundy [RBW18, SMW13]. Further [PMN^{+}15, GPO^{+}10, LT19]. Future [ACR18, WWM^{+}19, APY^{+}17, CRO12, CJHR14, DMMG14, Dav15, ET19, FCC^{+}12, FFH18a, Gle14, GAC^{+}19, HK10, HO10b, Ili10a, Ili10b, JY19, LB18, LBP^{+}19, MMGM16, MGB^{+}18, NBMB^{+}18, Oan18, OR19, PRG^{+}17, PGB^{+}17, PM10, RRG^{+}12, Sco18, SDLT15, SMGL17, SvdBH^{+}18, SMKBM16, SBT16, SH14c, TSH^{+}19, TSK18, TDMN19, Uhr16, VSS^{+}15, WML17, Won19, Laf14, Weg15]. futures [BMAG17, FLH11, GNN^{+}17, TR18, TRC^{+}17, LLF^{+}16]. fuzz [GOO^{+}19].

Gabes [HAH^{+}16]. Gadus [HBG12, MG14]. gain [CH12, KLMD15]. gains [TEF^{+}15]. Galapagos [BH14b, BHT^{+}18, BRRG^{+}18, CC13, GOdFC16, CSF^{+}13, Jon13]. Galicia [GLVLGG19, GGSRVL17, ONF15, dlÁPAVS19, PFMA^{+}19, SRS14, dO13]. Galician [BRR18a, RVdCGN11, VRMF11, Vil12]. game [Hea13]. Gandía [HRLFMP15]. Gannets [SAGG13]. gap [AdSC11, Bro11, BAY^{+}19, CL13, Gor19]. gaps [BAHCG14, CR11, HAC19, JHN^{+}19, LKMN16, PGB^{+}17, VHRM17, WFK^{+}12b, WKB15, KC15]. garbage [CL13]. gas [DT10, Fen17, GDD^{+}17, GDD^{+}19, KYS10, Kai16, MNG^{+}18, RIK16, Shi16, vBvTM16]. gateway [JKS16]. Gathering [FJG14b, PSCY12]. Gauging [FMKT15, WTA^{+}16]. Gear [CSVHS14, CWP16, SS15b, AG12, Cho11, EFK^{+}19, FSCJM19, Gil15, JG13, KLY14, KRM14, MMS^{+}14, MFAA16, NBF^{+}18, RRG^{+}12, RGWH18, WHMG^{+}16]. Gear-based [CWP16]. Gear-specific [CSVHS14]. gears [BRR15]. GecoServ [PSCY12]. gelatinous [AAL^{+}18]. Gender [dlTCFB^{+}17, PG19, UEMRG^{+}18]. gendered [Wos19, San15]. gendering [GPH18]. general [CMFDCLV10, KGS^{+}19, MBvZ15, MKB17]. generated [RWL14, RWG13]. Generating [HL10]. generation [LDS11]. generations [RCD^{+}18]. generic [BMGF^{+}17, SBS^{+}13, ASWM18]. genetic [BAHCG14, CMH19, CHDJ^{+}19, DGC^{+}17, HJN16, JS18, Lea19, LSAL19]. Genypterus [HKR19]. geoduck [SK15]. geographic [DCY^{+}19]. geographical [vdBAMJT19]. Geological [PKA^{+}16]. geopolitical [Ger15, KLC^{+}15, dVM14, dVM17, dVM10a]. Geospatial [Asa12, RHB17]. German [GSG^{+}15, SFC^{+}15a]. Germany [Thi13a, AYB18, BDG^{+}14, BB19d, SB14, Thi13b, Tos11]. GES [UNS^{+}18]. get [FG18]. gets [LAH^{+}13, WCCO18]. Getting [TJ18]. Ghana [Mul19, Nol19]. ghanaian [PWR17]. GHG [Gri17]. ghost [Gil15, SNA^{+}17]. Giang [HRC15]. giant [O’R15]. Gilbert [MG14]. gill [KLY14]. gill-net [KLY14]. gillnet [BHG11, PHC14]. GIS [SMKBM16, GSG^{+}15, NHPD19, RBW18, TMDSF19]. Gisborne [Bre14]. give [LGFB^{+}19]. GKI [ASWM18]. glacialis [HD11a, HD11b]. glaring [Hob14]. Global [BTS16, FKWV19, GZW^{+}19a, GZW^{+}19b, MKR12, NF17, NRB^{+}18, OBNH15, SMGL17, SP19a, SWS12, SLL^{+}16, Vil10, VKGN^{+}14b, WNLS17, WDK^{+}13, ZEP^{+}19, AAGMH17, ABBD17, ARS16, Bac18, Big16, Bla17, Bla15, BDP19, BOF^{+}15, CP13, Cam16, CMH19, CRB^{+}19, CMSKP10, CHDJ^{+}19, CPBQ19, CGC14, CBS^{+}16, CSP10, FSM^{+}12, Gar12, GLTB^{+}15, Gil11b, GRLJ15, Gri17, HL15b, JLH14, Jac15, KFB19, MF19b, MDG^{+}14, MTF^{+}15, Maz17, Nag16, NMJS17, NW10, Pau18, Pen18, PTML16, PML10, PO10, PWB19, PS12, RIK16, RM17, RM18, RFL^{+}18, RBC^{+}13, RWW^{+}14, RO13, RWBF19, RL13, SK15, SPMW12, SBW^{+}17, SES^{+}19, THNR12, TTUR14, VRMF11, WTD^{+}15, WP17, WP13, WT18, Wea16, WSGM16, ZPT^{+}16, vdMCJ17, BVB18, VRA^{+}12]. globalized [AL18]. globally [VKGN^{+}14b]. globe [Lew16]. go [He18]. Goal [Vir18, VKGN^{+}14b, LPS12, RWG^{+}19, VKGN^{+}14a, RFL^{+}18, SI18]. goals [BW18a, CG19, DFP15, DTM18, FWHS17, HHL14, KKSC19, OY19, SC19, Xin19, NM18, SCMS^{+}18]. going [DBG^{+}16, FJA^{+}14, RNCGB15]. Gomm [SJ19c]. Gone [TCS^{+}18]. Good [CE17, Smi18a, BSAT18, BRR^{+}12, FDS15, OHB^{+}18, Hil10, NALO19, UNS^{+}18]. good- [OHB^{+}18]. goods [SCC14]. Google [HBZH16]. gooseneck [CVP^{+}17, RGGFA17]. Governance [DFP^{+}16, DCF^{+}19, DSH^{+}19, ECL^{+}15, FGKMP14, LPC18, MFMP10, NACS17, PHM16, PAA^{+}15, RGC^{+}19, SGRV16, TRM19, ZLL19, ARJ18, AMvT11, AvTT17, BD14, Bla17, BSAT18, BGS13, BE15, BAT^{+}18, CMVLGG14, CACL18, CGF14, CBG^{+}17, Chu10, CE17, Cli13, CP15, CWP16, Cro15, CKHJ^{+}17, DB19a, DD13, De 13, DPJ^{+}13, DSWG13, FOG^{+}11, FMM13, FVM14, FKMV16, FJG^{+}14a, GGADSR17, GSJ18, GGD16, Gil11b, GPH18, GBC17, Goo17, Grö18, Gue18, GDD^{+}17, GDD^{+}19, HD16a, Har10, HDM13, Her18a, HHG10, HdVB13, HR14, Hou14, HP12, HBH^{+}19, JTP18, JAO^{+}16, JCBM10, Jet18, Jon13, JGMS16, JR14, KBB^{+}15, KP13, Kar17, KSWS18, KEF18, KS18b, KR18, KBB16, KHvH^{+}14, Lan18, LM15, LRJ^{+}19, LJ13, LJ14, LB18, MVRG13, MHF13, MF19a, MF19b, Mar17, MB10, MHK^{+}10, MLV^{+}11, MCHO11, MRSH13]. governance [MTP^{+}17, MSS18, MSC18, MCAM13, NKCR13, Nym19, OHH^{+}16, ÖGB^{+}10, ONF15, PSH18, PW16, Pla19, PGPS10, Qiu13, RvLvTH14, RvT14, RMRN^{+}16a, RBM^{+}15, RGSF19, SC13a, SC19, SNDB^{+}16, SDP^{+}16, Smi18a, SM18c, SCJ13, SKL16, TK11, Thi13a, Thi13b, TTUR14, TJ13, TGBL11, TKN19, Tyn17, Tyn18, VBSD17, WBB^{+}15, WPdR^{+}14, Wri15, Wri16, WvLvK12, YFR10, YFC13, YCHM14, Zer18, dVM10a, dVDA^{+}16, vHK17, vLvT10, vdMCJ17, ATB^{+}19, LACG18, SBC15a, TBKH13, vTRvLvH14, Fri10]. Governing [DCY^{+}19, FEF^{+}12, FMN15, JQD13, LJT19, MVRG13, OCCvT^{+}18, You16b, Gus18b, HKHG19, JDQV13, ALM17, LKC15, VG15]. Government [DFL15, HR16, EGHC16, EGHC17a, GC17, KH10, Nik19, PS11b, RCM18, WMWH18, dVvT11, LWJH11]. governmental [Cus17, MDNT19]. governments [CR10, Tec19]. grabbing [BGS15, KN17, Mal18]. Grand [Bai16, CMVLGG14, DR10]. Grande [RAG18]. graphs [MJRdC16]. gravity [NBM15]. Gray [Rah19]. Graying [CFC^{+}18]. Great [DD13, Jet18, Weg11, DSM^{+}13, DDS^{+}11, FSB14, FAPP17, GRW^{+}16, HMS^{+}17, HMK19, LBN15, LSTD12, LU13, MLHP15, PRS10, SHM13, WS15a]. greater [BWD^{+}19, FJ12, GGADSR17, HBH^{+}19, KAF^{+}15, EFK^{+}19]. Greece [TÁB^{+}10, MK14, NPS^{+}19]. Greek [MPMK16, Pap16]. Green [ARSHH16, NRN14, AJ17, BDvTvKT17, CCL^{+}19, CJ17, ESGC13, FHJ16, MOR15]. greenhouse [DT10, Fen17, RIK16, Shi16]. Greening [WvLvK12]. Greenland [HRJH15]. grey [SWM11, Var11, WNB^{+}15]. ground [Doy15, FAPP17, KB15]. groundfish [BMAG17, CKdS14, EE17, GLW17, HW10, KBG^{+}18, KS10, MC15, Por17, SvZdJ15, SF16, TC15b, WTD^{+}15, WSG18]. grounds [CMVLGG14, DMdJ^{+}15, KR14]. group [CA17, CBP^{+}18, GSJ18, Hol12, Bai12, Mso16]. Grouper [HTF^{+}15, MSB^{+}18, ASWM18, TMHJ12]. groupers [GTL^{+}18, NWW^{+}15, NWW^{+}17, SNA^{+}17]. Groups [FBP^{+}18, KWT17, MJRdC16, dVM10b]. growing [HBG12, RS17]. growth [BCM^{+}18, Cha17b, EMD^{+}18, GAC^{+}19, HC18, How18, KS16, KS18a, KARR^{+}19, KED^{+}18, LJS^{+}18, MM17b, MNG^{+}18, NF17, NRN14, Pas14, PCC15, RRG^{+}12, RWJ18, Ren18, RWG^{+}19, SvZdJ15, SvdBH^{+}18, SV15, TEVN19, VMO14, WMVR18, YBD^{+}19, BVB18, DRGP19, LPR^{+}17, NRB^{+}18, RRMSSE16, vdBAMJT19]. Guadiana [GSDB14]. Guam [WGMG^{+}16]. Guangxi [HWY^{+}18]. guarantee [MM15b]. Guatemala [GBC17]. Gueishan [LSC14]. guessing [DMMG15]. Guest [PCS14]. Guiana [CBG14]. guide [ItDD^{+}18]. guided [CBL18, RPG^{+}14]. Guidelines [TWJ^{+}17, DJ15, HPW^{+}18, SC19, Bro11, CPB19, CMB^{+}15, SALS17]. Guiding [FHM^{+}10, BGC^{+}17, FÓ12, RBJ^{+}11, Son15]. Guillermo [McL15]. Guinea [Ben12, Bro15, BSM^{+}14, CBF^{+}12, MDE14, OY15, OY19, TGBL11]. Guinea-Bissau [OY15]. Gulf [AFMCMP12, CST10, HGB^{+}19, Kai15b, Kai16, Kai17a, Kai17b, Kai17c, KARR^{+}19, Man10, MLTBMB14, RHP^{+}17, SBP^{+}19, AMLK^{+}12, CS14, CR10, CSF^{+}19, CSCBR10, CJW^{+}16, CSMB^{+}12, Don16, GSS^{+}18, HHL^{+}15, HAH^{+}16, Har10, JWBJ10, JWBJ13, KYS10, Kai15a, Lee10, LH16b, NC10, OY19, PSCY12, PJS13, RRAH13, RRC^{+}17, SAdC15, SMW13, TMHJ12, YWC^{+}16]. Gully [WFK^{+}12b]. gut [FFKO15]. Gwaii [JRP17]. Gyeong [KSS17].

Gyeong-Nam [KSS17].

Ha’apai [KH12]. Habitat [Dru10, GSS^{+}19, GRW^{+}16, CN11, ESGC13, EGP^{+}18, GRYA^{+}15, HIN^{+}17, HBZH16, JG13, LGM16, MTF^{+}15, MRSL16, MDR^{+}16, MPT13, RBJ^{+}11, SGMG18, WTD^{+}15, WSGID17]. habitats [DSP^{+}18, GOHDB17, GBK^{+}15, HA13, KM15, Lyo14, MRSL16, OTOL17, SSLR14, vH19]. Haida [JRP17]. Hainan [LLZ^{+}19]. hake [AG12, Dra15, Pat15]. half [WT18]. Halfway [AFCMET18]. halibut [CL15, Car13b, CBL18, EP19, FC19a, PC13b, Pin13]. Hamble [MPCB14]. Hammerfest [vBvTM16]. Hana [VC15]. handline [BB19b, BB19c]. Handling [MMC^{+}14b, PSV^{+}14]. Hannesson [Hee14]. happier [AYA^{+}19]. harbors [CCL^{+}19]. Harbour [WBB^{+}15, ItDD^{+}18, SKL16]. harbours [KCM^{+}18]. hard [KPF12, MDNT19, SU11]. hardback [Tec19]. hardened [SGN^{+}17]. Hare [JR14]. harm [WTA^{+}16, LMG^{+}16]. Harmful [BHA^{+}14, WPRA18, CMS19, MCB^{+}19]. Harmonization [TML^{+}19]. Harnessing [Sea17b]. harp [Liv10]. Harris [She14]. harvest [BKH16, CBL18, ERD^{+}13, HPW^{+}18, JBJ12]. harvested [CF13]. Harvester [O’R15, MMDS12]. harvesters [HK18, MBSL13, BD13]. Harvesting [RAG18, CVP^{+}17, KHD12, Kas16, NYJW15, PFH16, TRBS13]. harvests [Lan15]. Hastings [UA13]. Hatton [MSG11]. haven [Gam11, Har14, She14]. Having [TN15]. Hawai’i [FMP^{+}18, LR15, RL14, TLH18, AHB18, BKKN10, Pan14, PW15, RL19, ST13, SB18]. Hawaiian [KDW10, HVB^{+}17, HB14]. Hawai‘ [COG^{+}19, IOG18, TSK18]. hazard [GW15]. hazardous [CPRP19]. head [FFKO15]. headache [ØR18]. heading [Kim15]. health [BAA^{+}15b, BPF^{+}16, FTF^{+}17, HLM^{+}13, KKWS15, MF12, PWH^{+}16, SFDB18, TSW19, ZZ17a]. healthcare [NPAR19]. healthy [BWD^{+}19]. heard [Rog15]. heart [Laf14]. heat [dlFRG17]. Hebei [Cho10, CL16b]. Hebrides [Mso16]. Hegemony [Pin17, BYK15a]. HELCOM [VDB16]. held [KSB15]. Helena [Sch15b, vS15]. help [BF19, MWR^{+}17, Pan14, RM17]. Heritage [HMS^{+}17, WS15a, BTC18, DFV15, JAG16, JGA17, LZ16, vD16, Dro13, JSG^{+}18, PO14]. Herring [Lee10, BB19d, CE17, DT10, HES16, JRP17, KN17, PSH18, TEF^{+}15, TH15, WBH^{+}13, vdPLMS16]. Hesitant [SA18, Mes12]. hesitates [PdSA^{+}13]. heterogeneity [CMAA13]. Heterogeneous [RGGFA17, BL18]. heuristic [Par15]. Hidden [PL15, PFMA^{+}19, Tur10]. Hierarchy [Mor17, SRGD18]. High [APY^{+}17, GCCSS^{+}19, HD16b, Bla17, BY16, CT16, CSP10, DDV^{+}13, ECSE^{+}17, Han11, Hof19, Hua11, KSK^{+}19, MSG11, MSGM^{+}12, NKHL18, NBD15, NFHJ16, OBJ^{+}12, SKT^{+}10, VCdR^{+}18]. High-profile [GCCSS^{+}19]. High-resolution [APY^{+}17]. high-use [CT16]. high-value [DDV^{+}13]. higher [JJ19]. highly [GOHDB17, dMAH^{+}18]. hijacking [SCC10]. hilsa [IMA16, PMA^{+}17]. Himmerfjärden [KHP14]. Hinase [THK^{+}18]. hinder [CMS19]. historic [Hua14, PO14]. Historical [MOE11, MRSL16, Nag16, SDLT15, WM17, LDVM10, Mis19]. History [DPH^{+}15, Rei10, WSG18, BD10, NWW^{+}15, NWW^{+}17, Sch15a, TRM19, WLP10, CLS10]. Hoa [NJ18]. hoc [BCE^{+}17, SRSN17]. Hogarth [Ili10b]. Holbox [RCMBG^{+}19]. holders [Car13b]. holdings [AMG16]. holistic [BSW16]. Hombre [CWP16]. home [Har14, She14]. homeowner [SGN^{+}17]. Honduran [BGS13]. Hong [PC19, ST17, ZZ15b, ZZ17b]. hook [FSCJM19, HSB^{+}16]. hopeful [Pin17]. Horizon [FW15, GSS^{+}18, HB19]. Hormigas [SGNM19]. hornets’ [AJ12]. horse [PDB^{+}13]. Horta [DDL^{+}17]. host [NYJW15]. hotspots [GHP^{+}15, Iri18]. household [LBP^{+}19]. households [MBBDS19]. Hudson [HMH18]. Hueter [PDC19]. hull [FSV^{+}16]. hull-fouling [FSV^{+}16]. Human [CSF^{+}18, DSP^{+}18, EAA^{+}18, GTL^{+}18, LGM14, PT10, RD13, Son15, SH19, WSO^{+}14, Zer18, AMdlCMS^{+}18, AH19, ASB10, Ant16, BSW16, BSM^{+}14, Car11, CBG^{+}17, DTJ10, DB19b, FBM18, Gue19, HM18, HvPN^{+}19, HBZH16, IW19, Kel18, KRK13, LM11, LEG19, MD16, MSB15, MLV^{+}11, Ngo18, PCC15, PBJ^{+}14, RPG^{+}15, RFG^{+}14a, SS19b, SMKBM16, WGMG^{+}16, WTA^{+}16, WPRA18]. human-caused [DB19b]. human-wildlife [Gue19]. humankind [JGA17, vD16]. humpback [BJS^{+}16, BSP19, GCP19]. Humphead [PF12b]. hundred [Edd14]. hundred-fold [Edd14]. hungry [LGH^{+}12a, LGH^{+}12b]. hunt [BR13b, BR14, DHSC14, Liv10]. hunting [Wen11]. hurdles [TRBS13]. Hurricane [Kai15a, SGN^{+}17, CCS16]. hybrid [BRW^{+}13, NM19, Shi13]. hydrocarbon [REA^{+}14]. hydrocarbons [OS12]. hydrothermal [Joh19, MV19, TTFC19, VAHG^{+}18].

I4E [AVC18]. Iberian [FPP^{+}15]. ICCAT [Che12b]. ICCMS [HGDB10]. ice [APY^{+}17, GDN19, Ho10a]. icebreaking [PDR11]. Iceland [HRJH15, TÁB^{+}10, YCA^{+}12]. Icelandic [AMG16, CC17a, GS16, KKG16, KNY14, SG15a, SAM13, WBH^{+}15]. ICES [SOÖ12, Wen17]. Ichthyoplankton [KW16]. ICZM [BH14a, Coo11, TBB12, GAFAV16, MFB^{+}14, MDK^{+}16, RB12]. Identification [CZX^{+}19, LLH^{+}14, YYN^{+}15, YSJU^{+}17, ACP^{+}14, Hua14, Kru16, LLO17, MTF^{+}15, MSGM^{+}12, TÁB^{+}10, TTFC19, Xin19]. identify [FMD^{+}18, PGR14]. Identifying [CW15, DFK13, Hou14, Iri18, MUHH18, NCH^{+}13, PV16, Seu16, THF^{+}15, VRT13, BSW16, FW19, MS14, MCB^{+}19, JLCB16]. identities [GSJ18, San15]. identity [SPP18, UA13]. ideology [CF17]. if [DPS16, FP13]. IFQ [Car13b, PJS13]. IFQs [SAdC15]. Ignoring [PE10, AAL^{+}18]. II [SDF^{+}14b]. III [PSP10]. Ilha [EMP^{+}14]. Illegal [CSF^{+}13, OY19, VAP^{+}10, BO12, Cho12, GPM19, Kim19, LT17, MBM14, MGG^{+}19, Mul19, NSB^{+}18a, PF12b, PNPD14, PPM17, PPM19, TJ18, You16a, CKN19]. illegalities [CN17]. illustrated [FJ12, MV17]. illustration [FB15]. illustrative [FBAK14, Wit15b, Wit15a, dCW15]. I’m [FVT19]. imagery [HBZH16, JL11, Lyo14]. images [JCBM10, OIS^{+}18, SCJ13, VGBG15, vdVD16, NSB^{+}18a]. imaginary [Fla10]. IMO [Bro11, GJ15, KP17, KH11]. Impact [CGS^{+}17, CCC^{+}13, KGMP18, RNN^{+}18, UNS^{+}18, AOAM13, BAA10, BG15, BS18, BKH16, BHT^{+}18, CKC^{+}13, CK18, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, DSF16, DTM18, dlFFVRC16, GA19, GHO^{+}17, GAC^{+}19, HHP17, HK18, JTK18, KC13a, KB15, LM16, MKX17, MDE14, MLF14, MC16, MLA16, MBBDS19, MPT13, MMA16, OCG15, PDT^{+}14, Pin17, RM18, SFDB18, TÁSLC19, TLB19, VGA15, VLKN17, WP17, WZCHS10, Wit19, YP17, YC18, DLM^{+}18, MML^{+}19]. impact-assessment [MMA16]. impacting [BAY^{+}19, MMP18]. Impacts [BAHD10, Bar10, BSSS17, BYK15a, Jen15, LSMW10, MM17a, MMDS12, Ngo18, vdGMA^{+}13, ALMS14, AHB18, BD14, BDF^{+}10, BOF^{+}15, Bor18, BMGH11, Car13b, CA17, CFC^{+}18, DBG^{+}16, DCHC17, ESM^{+}19, EJB19, FC19b, GOdFC16, HMGR14, HPPH16, HA13, JG13, KSS17, KMTS16, LDS11, LC19, MHC19, MC15, MCH18, PB13, PPML16, RPG^{+}15, RDP^{+}18, RRG^{+}12, RLH18, RvdBG13, RRR17, RDVSC16, SBP^{+}19, SVL^{+}16, Seu16, SK18, SSD17a, SNP^{+}18, SHM12, UAZ13, VMO14, WTD^{+}15, WWT^{+}16, WML^{+}16, WPRA18, BLC16, FSV^{+}16, JLW17]. impede [GC17]. Impediments [CBE^{+}19, CETQ17, HTA^{+}15a, vLRvH^{+}14]. imperative [BFO^{+}16, BKHA^{+}18]. imperilled [SDC13]. imperishable [OO17b]. implement [CMH19, FGKMP14, RCM18]. Implementation [De 18, SFC^{+}15a, TMKK^{+}15, AF14, BPMY18, Boh19, CRCD13, CLS10, CPB19, CSMB^{+}12, DAGG16, DW13, DSR12, Eli14, EHR15, ENWL19, ETEW15, EGF^{+}17, GMD^{+}12, GRMAC18, GSS^{+}18, HJKR14, HAC19, KBGF12, KH11, KAPF^{+}19, LSK18, MG19, MKR12, MMdJ^{+}14, MMJM16, MPMK16, MKU19, OPMB12, PRT10, PP10, RGG^{+}16, RAZ16, SALS17, SWP13, SBS^{+}13, SGMWE14, Thi13a, Thi13b, Vin13, Wis14, YPS13, dGNC18, vLRvH^{+}14, vTRvLvH14, MML^{+}19, dVDA^{+}16]. implemented [Asa12, BSSS17, CCC17, VPL^{+}13]. Implementing [BS10, LWS19, PFF^{+}18, vLRvH^{+}14, AH19, DG14, Has18, HPW^{+}18, HFF11, JSD19, JGMS16, MBD13, Mar17, PRPLBLC12, RTD^{+}17, RSS14, SOCA15, Vau17, WM18b, vHHB14, vLvHvT12, CETQ16]. Implication [DGVS^{+}16b, RAG18]. Implications [RRG^{+}12, SSLR14, VPC19, YBRBB16, ZH18, BMPV10, BH14a, BR13b, BR14, CCR16, CMCMHP13, CMMBV^{+}19, CDR^{+}16, DHSC14, De 11, DZLL15, DT10, Fol13, GOHDB17, Gek10, Gol13, GM14, GJA16, HWN11, Ho10a, JSLL19, JSY^{+}19, KCJW15, KH12, KH13, MSBB14, Mal13, MR10, NY13, OTOL17, PCC15, SRA15, SDLT15, SMGL17, ST17, SCS16, TJ13, VH16, WWM^{+}19, Bes11, BMGH11, CACL18, De 13, KS12d, LB18, LLZ^{+}19, MRSL16, MLV^{+}11, MLTBMB14, NGS16, NSB18b, PWN18, QSMGVA18, SKFL15, SB15, SAGG13, SBC^{+}16, SS15b, TS14, TKAO^{+}19, WBB^{+}15, ZP18, dlFRG17, vHK17]. import [He18, PP19]. Importance [GM14, AG12, DBB^{+}14, FMGV16, FC19b, GNHM18, Kep14, KCM^{+}18, KM10, MD15, Nei18, PWH^{+}16, dlÁPAVS19, PC18a, RSMM19, SAM13, SPP18, Tho13, VRT13, JBBV^{+}19]. Important [SBC^{+}16, CG18, DDvP^{+}16, HTD^{+}18, SBSK19, WWH^{+}17, YYN^{+}15, YSJU^{+}17]. imported [WSC19]. imports [GPM19, PNPD14, PPM17, PPM19]. imposed [MPMK16, SJP17]. impoverished [JOE^{+}19]. Imprecise [WGB^{+}16]. improve [BAA^{+}15b, BAA^{+}18, BF19, CKY^{+}13, CST10, DSK^{+}19, DMB^{+}18, ELAC19, FSS15, GDD^{+}17, GDD^{+}19, GBBB15, KL14, NWW^{+}15, NPAR19, PHS18, RWK15, SBTF15, SCJ13, SG12, VCdR^{+}18]. improved [LTM19, PKPdP10]. improvement [BTOM13, CSK^{+}18, CEB^{+}12, DJ15, Yat14, BBE19]. improvements [HTH13]. Improving [BvRH11, BPB19, CC13, CL16a, GOM17, HCP^{+}17, HPW^{+}18, Kat16, MF12, MDR^{+}16, PPCM16, SWB^{+}14, AAGMH17, CSP^{+}14, SCK15, BSW19, Cli13, JFT^{+}18]. IMTA [ET19]. in-country [BSAT18]. In-situ [JOE^{+}19, MMJM15]. Inadequate [HBA^{+}14]. inappropriate [KBF12]. Incentive [ZLSL17]. Incentives [ÖSS^{+}11, AM15, CKY^{+}13, DR10, DPB17, DSWG13, EAE15, IMA16, MMA16, ODL^{+}13, RWG13, SHL^{+}18]. Incentivising [CGC14]. inception [CPFS^{+}13]. incident [Cho10, CPRP19]. incidental [FC19a]. include [PLL17]. including [Bor10]. inclusion [PSCY12]. Inclusive [CHDJ^{+}19, Ben18, ZS18]. inclusiveness [BHI^{+}17]. incoherent [Tor16]. income [AYA^{+}19, CA17, WBC14]. incompatibilities [PFH16]. incompatible [GAFD^{+}18]. incomplete [JBK15]. incorporate [MG14]. Incorporating [ARJ18, BBP^{+}14, EGHC17b, GCD^{+}15, MM15a, SGMWE14, WMJF16, LB13, MDM16, WSGID17, CBG^{+}17]. incorporation [RYRYM17]. increase [BF19, EGHC17a, Nij15]. Increased [CP18, OP15a, vH19, OHH^{+}16, ZKR^{+}19]. increases [NKHL18, OP15a]. increasing [AFR^{+}19, HRA^{+}15, RCMBG^{+}19, SBSK19, TMR^{+}14, SNTL18, dlTCFB^{+}17]. increasingly [Gue19]. incremental [CL16a]. independent [DPV^{+}15, FC19a, Pol12b, Sch10b, LM16]. index [FMGV16, MCB^{+}19, WK14, AVC18, VRA^{+}12]. India [CA17, Kar17, KSC^{+}15, PRG^{+}17, PRBX19, PKR^{+}17]. Indian [AARK15, DMMG14, DMMG15, ECL^{+}15, NM12, VGA15]. indicates [LBHS16]. Indicator [ASWM18, Del17, FDS15, MACM14, MHA^{+}17, MC14, NBF^{+}18, YCHM14]. indicator-based [MHA^{+}17]. Indicators [CJW^{+}16, AARK15, ASP16, APD^{+}19, BJGMG18, BSW16, CCTG^{+}13, DTN10, EGP^{+}18, HvPN^{+}19, JWBJ13, JLCB16, KVMM10, LAP^{+}12, LPS12, Liu13, MLTBMB14, OP15b, PGL^{+}16, PM10, SG17, SGMG18, THF^{+}15, Vir18]. indices [MSdAM16]. Indigenous [BF18, DAJM15, Bro18, BRW^{+}13, Cap10, FSV^{+}16, GRYA^{+}15, KCJW15, Lal17, LRJ^{+}19, LOD^{+}15, LCC14, OGM^{+}14, vPLB^{+}13, BKHA^{+}18, RYD18, vdPLMS16]. indirect [RLM10, RLM12]. indispensability [BYK^{+}15b]. indispensable [DAJM15]. Individual [BBG^{+}16, CD15, DRH^{+}18, KSC^{+}19, Mox12, MCBM18, SAGG13, Bod17, EGGT12, EGHC16, FGW15, GDD11, GS16, MNHN18, MUHH18, Sol14, SCS16, TMHJ12, Tur10, CSMB^{+}12, PD17, YCA^{+}12]. individuals [CP18, LGFB^{+}19]. Indo [LB13, PRT10]. Indo-Pacific [PRT10]. Indo-west [LB13]. Indonesia [MFRvP16, ARSHH16, ALM17, CKY^{+}13, CM17, CKN19, CBF^{+}12, Cli13, GBF^{+}10, GAW^{+}13, HWL^{+}19, KGMP18, KH15, MBT^{+}19, MYA^{+}18, Nij15, RSMM19, SSW^{+}18, SH14c, TRC^{+}17, VAP^{+}10]. Indonesian [DK16, EJB19, FSGB^{+}15, PRPS17, Tam15, TRC^{+}17]. Industrial [RDBB17, BLC16, DYN^{+}19, DR14, DZ12, JTK18, JCD17, LPU^{+}16, MMN18, MHM16, RRG^{+}12, RAG18, SCMS15, SCS14, WS15a, WZCHS10, YYYL15, BHT^{+}18]. industrialised [Wri15]. industrialization [SH15]. Industrie [MMC^{+}14b]. industries [Bar10, GLW17, MTF^{+}15, MMTC16, NS18, She15, UL17]. Industry [SC13b, AT14, AYR^{+}14, AF14, AL18, BD19, BBRT16, BSAT18, BDH^{+}18, CJ17, CKBS14, DHM^{+}17, EFK^{+}19, ET19, EGHC16, EGHC17a, FHJ16, FFH18a, GS16, HHS16, HHP17, HSW17, JBBV^{+}19, JKS13, Kai16, KH10, KS19, Knu17, KGS^{+}19, Lim16, LZL19, LFGM15, LJT19, Mal13, MC19, MFT19, MPPSdVSOHR14, ML19, OP15b, Ola18, OTOL17, PHLM19, PKPdP10, PML10, PT10, RRA^{+}10, RPSS17, Ren18, Sea17b, SS19c, THRS15, VBMM17, WW19, Wri15, Wri16, ZSZPR19, dVM10b, dVvT11, vPG10, HLC^{+}10]. Industry-driven [HLC^{+}10]. ineffective [ONF15]. inefficiency [VRT13]. inefficient [Gib10]. inequalities [PL12]. inequality [WBC14, Zag18]. inertia [KBB16, LHB^{+}17]. inference [Man10]. Influence [HSB^{+}16, ADH18, Ben17b, BH14b, CKVO16, DS17b, GGK17, GSW18, KJWP14, KWT17, MKB17, PTDH17, PNW18, SWM11, Soo17, THM15, TKAO^{+}19, VGBG15]. influenced [Che12b]. Influences [ATB^{+}18, NPS^{+}19, PD17, SCV^{+}17, VGG15]. influencing [ALSC13, AL18, Che10, DCC19, ETEWZD11, NKBS17, OM19, WHMG^{+}16, Weg11, vPLB^{+}13]. inform [ADL^{+}18, BPMY18, CCK19, Cro15, DMdJ^{+}15, GOTS19, GAW^{+}13, JMJ^{+}18, JG13, JS18, KMC^{+}17, LCH^{+}13, LGDB17, MAJM^{+}18, NACS17, OBH15, SSCC19, SBTF15, She15, SMKBM16, WK17]. informal [EA19, HIBG15, JSD19, PRBX19, Ped13]. Information [ZMM^{+}14, BBvZ^{+}17, CMP^{+}15, CRB^{+}19, CP18, EB16, HFM^{+}13, HGBJ16, IT15, JFA17, JBK15, MMJM15, OCG14, PGC19, SWM11, SMW13, Soo17, TEVN19, GKG16, LDS11]. informational [HdVB13]. informed [BM15, HPB^{+}15, WLT^{+}19b]. informing [AAC^{+}13, CPBQ19, WHB^{+}18]. informs [Cro11]. infrastructure [BBR^{+}18, ESGC13, GCD^{+}15, KCM^{+}18, KMvT15, MNM10, WMK15]. inherent [vdPLMS16]. initial [Ard18, LJB^{+}14, SJP17, VKA^{+}19]. initiative [DS18, Ili10b, WC16, Fen17, FAA^{+}13, HBZH16, RO13, YC18]. initiatives [HAW19, Ili10a, MV17, TC13, VKGN^{+}14b]. injuries [Næv17, WLV15, HUHA18a]. injury [ML19]. Inlet [CN11]. inner [San18]. innovate [Vat17]. Innovating [vdMCJ17]. Innovation [LM18, ABSB15, ABSGB18, CHDJ^{+}19, Cra19, DCC19, MO14, PCC15, SvdBH^{+}18]. innovations [Goo17, GMP^{+}16, JA10, Jet18]. innovative [LLO16, ZMM^{+}14, vH19, LLO17]. input [EGGT12, GGSRVL17, GHO^{+}17, GSW18, LY14, MO13, NBD15, WW19]. input-output [GHO^{+}17]. inputs [SHL^{+}18, CDR^{+}16, MFAA16]. Inshore [Mso16, Mul19, DHR16, KMC^{+}17, PC19, PS10, PTP13, RCUR13, SvZdJ15, SM17b, TLB19, TPS15]. Insights [CST10, MHK^{+}10, AAL17, FSCJM19, MOE11, MTP^{+}17, PHEM^{+}17, ZMZ16, AM15, CVP^{+}17, CKHJ^{+}17, CvPH^{+}18, Gat18, GRMAC18, GPA18, HES16, LLF^{+}16, LPU^{+}16, MMC^{+}14b, NBFO18, Ola18, PRBX19, SDRML15, Soo17, WNLS17, WMK15]. Inspection [CW15, vAS19]. inspection/investigation [vAS19]. inspections [CW15, GCW^{+}18]. inspector [GCW^{+}18]. inspectors’ [PHEM^{+}17]. INSPIRE [AGB^{+}18]. instability [KLC^{+}15, dVM17]. installations [TC15a]. installing [CLV^{+}16]. Institute [Rah19, Wen16]. institution [LCC19]. Institutional [CMVLGG14, CMFG15, CBF^{+}12, Fla13, RMRN^{+}16a, RO13, San10, Val13, WCS18, vLvHvT12, Bes12, CR10, DSH14, GCMA^{+}12, HRLFMP15, IMA16, JQD13, Mat16, MM14, NF17, PFH16, TTM18b, vLRvH^{+}14, SS19c, WHMG^{+}16, YTTK10]. institutionalisation [SRSN17]. institutionalised [KR18]. Institutionalizing [BDvTvKT17]. Institutions [GGADSR17, JA10, EA19, FJA^{+}14, PDS15, RDSB10, RO13, Tir17, Wos19, dlTCL10]. instrument [FR16, Sco16]. instruments [GCD^{+}11, HR16, Sch12, YBRBB16]. insular [KDS^{+}15, Lyo14]. insurance [PRG^{+}17, ZMZ18]. intake [VLKN17]. integrate [GKG16]. Integrated [FH15, HNVZ12, TE16, dVM12, APC16, BLG^{+}18, CETQ17, CBE^{+}19, CRCD13, Cha15, Chu10, DTN10, ET19, GMD^{+}12, HP13, HMS^{+}17, HID12, Kno10, KHP14, LTM19, MFB^{+}14, MSC18, MCAM13, OOvdM^{+}11, PCCR15, RSMM19, WHP^{+}16, WM18b, YCY^{+}15, ZPB^{+}13, dLS10, vOHO^{+}17, BH14a, CBG14, Coo11, FOG^{+}11, GSDB14, OHH^{+}16, RSS14, RCM18, TBB12]. Integrating [BBCPC11, EGP^{+}18, HP12, VCdR^{+}18, WSB19, WNB^{+}15, EG15, LCR16, MRC17, KJS14, MF19a]. Integration [PFCP^{+}14, AYB18, BMD^{+}14, CETQ16, Kid13, KMR^{+}13, MMdJ^{+}14, SvHdB^{+}15, WBC14, XJT18, vH19]. intelligence [FMD^{+}18]. Intended [GLR17]. intensity [BH14b]. intensive [She16]. Intent [FAPP17]. intentions [SOUN16, WSO^{+}14]. Inter [FGPP18, SGMG18, JAO^{+}16, MDNT19, Seu16, TE16, DZ12]. inter-governmental [MDNT19]. inter-jurisdictional [JAO^{+}16]. Inter-local [FGPP18]. Inter-regional [SGMG18, Seu16]. Inter-sectoral [DZ12]. inter-tidal [TE16]. interacting [KED^{+}18]. interaction [LGM14, RvLG^{+}15]. Interactions [BFTS^{+}16, LPU^{+}16, NHFV13, BHS19, BHG11, CBS^{+}16, CJHR14, DZ12, FBM18, IOG18, KH12, MSB15, MD15, NRB^{+}18, PC13b, RvLG^{+}15, RNEBG19, Ros17, SAGG13, WWM^{+}19, WFK^{+}12a]. interconnections [VBMM17]. interdependence [Car14]. interdisciplinary [VS18, RCMBG^{+}19]. interest [CMFDCLV10, HDJD15, Hol12, WWW10, FVM14]. interests [NJ18, PVS^{+}19]. interface [BD13, Soo17]. interfaces [RMRN^{+}16b]. Interference [DRMA18]. Intergenerational [TCS^{+}18]. Intergovernmental [HD16a]. interlinked [BMD^{+}14]. International [NWW^{+}15, NWW^{+}17, Oan18, YPS13, You16a, AAC^{+}13, AF14, ARGW14, ADHV16, BDIS18, Big16, BY16, BW18a, CG19, Che12b, CO16, CDR^{+}16, DHC11, DZLL15, FFSF17a, FE12, Gil15, Goo17, GHP^{+}15, GCCSS^{+}19, GCD^{+}11, HKHG19, HK11, KV18, MHL13, MHK^{+}10, MRSH13, MS14, MN18, Muk15, NBM15, PCLE17, RM16, Ros17, RO13, SJP17, Shi16, SOÖ12, Stø19, TEF^{+}15, VAHG^{+}18, WSWM17, WM19, WF18, WvLvK12, Xin19, ZZ16, dMAH^{+}18, Ard18, Bai12, Cla15, DW13, HGDB10, Ili10a, Ili10b, Jae16, KHT17, LJL^{+}14, SVA^{+}12, SALS17, VBHC^{+}17]. internationally [BIPS13]. Internet [SB14]. interplay’ [AH19]. interpretation [ERB14, KBGF12]. Interpreting [Men19]. INTERREG [ECR^{+}16]. interrelationships [LWvdBL16]. Interrupting [AJ17]. intervals [SK18]. intervention [ARS16, Sol14]. interventions [Kat16, NGS16, SB15]. interviews [MSB^{+}18]. introduced [SH16]. Introducing [Sha10, SJ10, TEF^{+}15]. Introduction [FKMV16, HR14, JDQV13, Nei15, PO10, Rei14, RJ18b, Smi18b, UAZ13, VWG19, BE15, LCC14, MDE14, NPAR19, PCS14, WFSV15, vdMCJ17, JML16, JLQ16, RvT14]. Inuit [Wen11]. invariant [TRM19]. invasion [CAGH16, SLDHS18]. invasions [DGC^{+}17]. Invasive [KSK^{+}19, MCCC16, MOR15, SHS^{+}14]. Invasiveness [HTF^{+}15]. invention [vS15]. inventory [PVK18]. invertebrate [Har15]. investigate [Vra19]. Investigating [Hol12, HNC14, BOF^{+}15]. investigation [BJO19, MK15, MSA^{+}18, Mor17, Mox12, Sch10a, ZMZ18, vAS19]. Investigations [KB15]. Investing [ESGC13, BBR^{+}18, TC13]. investment [HHL^{+}18, Kru16, MBBDS19]. investments [CPBQ19]. investor [EP19]. invisible [ZTAO13]. involvement [CBN^{+}12, CS14, EFK^{+}19, FJ12, KHvH^{+}14, MM18, NC10]. Involving [CKN19, PLL17]. Ionian [RGB^{+}19]. Iranian [BSMS15]. Ireland [BvRH11, BC13, Car19, Coo11, CK11, FOG^{+}11, FOO^{+}15, GCHO19, HBC^{+}13, HF12, HNC14, KO11, MCHO11, MLA16, MOH11, OL11, ROD15, RU12, VMO14, YPS13]. Irish [Cro11, CJHR14, FvRA11, HWN11, HSR^{+}18, Kid13, MLA16, MO12, MO13, OO19, Ren18, vOHO^{+}17]. iron [GPO^{+}10, SS15c]. irony [Web11]. irradians [OTOL17]. ISBN [Tec19]. Isla [UEMRG^{+}18]. Island [BTS^{+}17, DB19a, DMB^{+}18, HHG10, KRM14, LH16a, MBT^{+}19, NKCR13, RBS^{+}19, SOBS18, ALSC13, BBR^{+}18, BPB19, BS18, FOO^{+}15, GOTS19, HDM13, KVA18, KLN17, KVMM10, ROD15, RNS19, THRS15, WF18, AG18, ACR18, Bar10, BAA^{+}15b, BSH^{+}19, DTJ10, DJTK17, Dva19, FMB^{+}18, HDJD15, JBV^{+}18, KH15, LLZ^{+}19, LSC14, MAJM^{+}18, OMF14, Par10, PTDH17, PLP14, RGSF19, RCMBG^{+}19, Sha10, TGT^{+}19]. Island/Massachusetts [HDJD15]. Islands [Bro16b, DFP^{+}16, Han19, MSTB19, PCS14, VGL^{+}17, BMPV10, CdL17, HTT10, JOE^{+}19, KVA18, Men19, Sea17a, SNA^{+}17, AOA^{+}15, ADL^{+}18, BFL18, BAA^{+}15a, BCMH^{+}18, BAA^{+}18, Bol17, Bre13, CSP^{+}14, DA18, DGVS^{+}16b, DTN10, GOK14, GNN^{+}17, GOdFC16, HWO^{+}18, HB14, KGST14, KDW10, LDF^{+}18, Lil16, PRT10, RLAT19, RB15, RNS19, Sea17a, SSM15, VSRD16, WL18]. Islas [SGNM19]. ISM [Bha12]. ISSF [CRB^{+}19]. Issue [Tyn18, DMW19, FKMV16, JHB^{+}18, Nei15, WFSV15]. Issues [HDTS14, KSH^{+}15, EZPRZ18, GAB^{+}17, KBB^{+}15, LJS^{+}18, Nol19, PRG^{+}17, RU12, SBMD13, SJ19c, Val13, VSRD16, Vin13, CBP^{+}16, HBC^{+}13]. Italian [MCBM18, PC18a]. Italy [UNS^{+}18]. ITQ [ITNL^{+}14, SG15a, SCMS15, SA18, TINL^{+}14, WBH^{+}13]. ITQ-managed [TINL^{+}14]. ITQs [YCA^{+}12, BW18b, Gib10, Han13a, Han14b, Hee14, Pin13, Sol14, TEF^{+}15]. IUCN [ND17, PC11]. IUU [CMCMHP13, ES18, HAK^{+}19, Hut19, LGC16, MS14, OIS^{+}18, Ped13, Pol12a, Ros17, You16a]. IV [ECR^{+}16]. IVQ [KSC^{+}19].

J [JS13, Kep14]. Jamaica [KMvT15]. James [McL15]. January [Ano10g, Ano11g, Ano12g, Ano14n, Ano15w, Ano16q, Ano17t, Ano18w, Ano19v]. Japan [PPM19, Cla15, GPM19, NKBS17, PPM17, SOUN16, SSWP10, SSSO17, THK^{+}18, VM18, WSO^{+}14, YTTK10, YCA^{+}12, YTSO15, Yos17]. Japanese [BM13, Kai19, MV17, OIS^{+}18, TII^{+}19, YCA^{+}12, YYN^{+}15]. Jasus [LSMW10]. Jeju [KJH^{+}10]. Jellyfish [BBAP^{+}17, TMR^{+}14]. jeopardize [LGD^{+}19]. Jiangsu [BGG19]. JMPO [BFGR14a]. job [ESGC13]. John [Tec19]. Johnston [McL15]. Johor [HIN^{+}17]. joint [LLO16, STG11, SvHdB^{+}15]. Jones [AGWMP19]. journal [Won19]. journey [CQ14, WSB19]. judgements [MC19]. judgment [KTD^{+}15]. judgments [EG15]. judicial [CSO17]. July [Ano12k, Ano13g, Ano14i, Ano15r, Ano16s, Ano17v, Ano18s, Ano19q]. jumbo [RRAH13]. June [Ano15n, Ano16t, Ano18m, Ano19s]. jurisdiction [ARJ18, BY16, BDP^{+}17, BAHCG14, CHDJ^{+}19, De 18, DG14, EKSZ14, FJA^{+}14, HD16a, HR14, Hou14, Lea19, MLF14, MDNT19, PVS^{+}19, RUH^{+}14, Sco16, TDMN19, Vra19, WAG^{+}15]. jurisdictional [DD13, JAO^{+}16, WNB^{+}15]. jurisdictions [CS17]. jurisprudence [Goo17]. just [Ben18, Cou11, SWB^{+}14]. justice [De 11, GKSJ11, GLJdlTC14, KSB15, Cla15]. Justification [IH18, Lon12, SWP13]. Justifying [CN17, Árn17]. juvenile [MRSL16].

Ka’an [VAW14]. kaiar} [Lal17]. Kali [LT19]. Kalpitiya [MJS16]. Kaplan [WMVR18]. Karimunjawa [CKY^{+}13]. Kattegat [Eli14]. Keeping [GW10, GAB^{+}17]. Kenai [Kru16]. Kent [RJ13]. Kenting [KC13b]. Kenya [CNFJ10, MD16, TKAO^{+}19]. Kenyan [CMW10, WBC14]. kernel [LGM16]. Kesennuma [VM18]. Kettle [KKWS15]. Key [JY19, KH15, LCS15, Par15, THPJ17, TGBL11, ADL^{+}18, CZX^{+}19, DSP^{+}18, KO11, LJS^{+}18, TGS18, WCK^{+}17, ZZZL18, GFH^{+}17]. Keys [WSGM16, Uhr16]. Khanh [NJ18]. Killing [MSB15, NSB18b]. king [SH16, GNN^{+}17]. Kingdom [FDP19, MPF12, Mil14, RFG^{+}13, SHS^{+}14]. Kino [CST10]. Kiribati [Mal18]. Klein [Kep14]. know [TJ18]. Knowing [BD13, And15]. Knowledge [ASWM18, ATB^{+}18, Ben17a, BAY^{+}19, BD13, CHBRJ19, CN11, CCK19, CRB^{+}19, CMW10, CST10, CHL^{+}19, CHvKM15, DAJM15, Dra15, DHR16, ELAC19, FV10, GTL^{+}18, Gre15, HK10, HIN^{+}17, HJN16, Kep14, LLF^{+}16, LBHS16, LJ14, LC19, LSAL19, MACM14, MBTV14, MSG18, Nei15, OP15a, RYRYM17, Rei16, RTM19a, SARF16, SCR^{+}19, TÁSLC19, TMG15, TEVN19, VG15, XJT18, vS15, DA11, Pat15]. knowledge-action [BAY^{+}19]. Kodiak [Car15, RCD^{+}18]. Kong [PC19, ST17, ZZ15b, ZZ17b]. Korea [Cho10, Cho12, CPRP19, CL18, JLY18, KLY14, Kim15, KSS17, Kim19, KJH^{+}10, LEKB10, LM14, LM15, SK18, TJ18]. Korean [KBK19, KS19, LY14]. Kotor [Nik19]. Kraska [McL15]. krill [CHKG16, KL12]. Kutum [BSMS15]. Kuwait [AZAMG13].

labeled [SCOS17]. labeling [HA13, KHD12, SVL^{+}16, XDG^{+}16]. labelling [BHM16, DAGG16, He18, KVH14, WC17]. labels [MM15a]. labor [PG19, THRS15]. Laboratory [TI13]. Labour [EZPRZ18, CLS16, CS17, FFSF17a, MV16, ZP16, BW18b, ZZ15a]. Labrador [MG14, CPKF19, KR14, MMDS12]. lack [CPR14, San10]. Lagoon [HRC15, KNA10, AYR^{+}14, HDK^{+}15, RLAT19]. lagoons [APD^{+}19, ESM^{+}19, KNB^{+}18, PTDH17]. Lake [CN17, Don17, PG19]. Lakes [GRW^{+}16, MLHP15, Jet18]. land [BG16, CSP^{+}14, Fen17, JHB^{+}18, KJS14, KSH^{+}15, TML^{+}19, VPP^{+}16]. land-based [JHB^{+}18, TML^{+}19]. landfall [CCS16]. Landing [CKVO16, Boh19, CRSM^{+}17, EB16, EFK^{+}19, PHEM^{+}17, VPR^{+}16, Wit19, dVDA^{+}16]. landings [CVFP^{+}12, DPH^{+}15, HD10a, HID12, LDVM10, MMdJ^{+}14, OCW^{+}17, SS16]. lands [LU13]. Landsat [HBZH16]. landscape [KW18, MBB18, NMvP19, QJ13, RJ18a]. landscapes [KW18, SCV^{+}17]. Lanka [HDK^{+}15, DAJM15, GKG16, GGK17, GJA16, MJS16]. Lankan [GB15]. Lao [Ngo18]. Large [BMF^{+}16, GPB^{+}18, SGZI19, SC10, ABCDMG17, AdSC11, AYA^{+}19, BSW19, BGC^{+}17, BHG11, CBG^{+}17, CE13, DD13, FMM13, FEF^{+}12, GCP19, JLY18, JD16, MFMP10, MMB^{+}19, NWD18, dOSL18, SMKBM16, WK17, WMK15]. large-mesh [BHG11]. large-scale [AYA^{+}19, CBG^{+}17, CE13, FEF^{+}12, WMK15]. largest [WLT^{+}19b]. last [MDE14, MMC^{+}14a]. late [BGG19]. later [GSC^{+}16, MCMB^{+}18]. Latin [FMS17]. Lau [Sea17a]. laundering [PP19]. Law [Cho10, WC16, ADHV16, BPB19, BDB19, CDO16, CR11, FPG^{+}17, GDN19, KPF12, Kar17, Kim19, KCJ15, McL15, Men19, MRSH13, Nor15, OL11, PS19, Ray19, RHW^{+}19, Ros17, Smi17, TK13, Xin19, You16a, YC18, DG14, Mil14]. Lawrence [BAO^{+}12, Doy15]. laws [BD10, HAGB15, ISM^{+}17, MYA^{+}18]. lay [Pin13, AOY16]. lay-up [Pin13]. LCA [LVRM^{+}18]. lead [AHB18, GJ15, GW10, KVA18]. leadership [Men19, VM18]. learn [JE11, KGAB14, SOS12]. learned [BN19, BMGF^{+}17, CNJ^{+}10, DN12, De 16, ECR^{+}16, FGPP18, Kai17c, PRLCL12, SM18c, TC13, WSA^{+}15, WSGM16]. Learning [HJN16, LLS^{+}15, Wis14, BRM^{+}17, BMAG17, DA11, GLR17, JTBP17, KGAB14, PMM^{+}17, TWJ^{+}17, THPJ17, GW15, KEF18, XJT18]. learnt [PP10, RUD^{+}12, SMR^{+}15]. Lease [vPG10]. leasing [CBL18, Dav10, PE10, Tur10]. least [SRGD18]. least-cost [SRGD18]. Leatherback [RBC^{+}13]. led [GRYA^{+}15, KFBG13, KBBR16, LZ16, dGB10]. ledger [Dav10]. left [MKR12, MSS18]. legacy [Rei14]. Legal [AF14, Gla19, GCMA^{+}12, ZZ17c, AGSP^{+}13, Ásg16, BW15, Big15, Big16, CVFP^{+}12, DSH14, EZPRZ18, FP12, FP13, Hou14, Hua14, IMA16, Lan18, LPP17, MR10, Pla19, RGSF19, SOCA15, Sch12, SI18, SBSK19, TvN13, TK11, TC15a, Tei18, WPdR^{+}14, WOdG^{+}16, ZZ17b, vdPLMS16]. Legislation [GDG^{+}18, CDO16, FKWV19, GH15, HC19, Kar17, KH11, MB18a, RBB^{+}16, SO13, Syb15, ZZ16]. legislative [LOD^{+}15, PV16]. legitimacy [HRLFMP15, KL14, PDS15, SS15b]. legitimate [PW16]. leisure [RRA^{+}10, RMH^{+}15]. LEK [HJN16]. length [CVFP^{+}12, DWCS18]. lens [Cli13]. Leone [SBM^{+}17]. Lepidorhombus [MACM14]. less [CP18, AGB13]. Lessons [Boh18, KBC^{+}13, LH16b, SM18c, TC13, WSA^{+}15, WKB15, ACA^{+}13, AFCMET18, BN19, CL15, CGC14, EOC15, FGPP18, Kai17c, KSWS18, MM18, MUHH18, PRLCL12, PVPF19, Sco18, TGBL11, VC15, WM17, ADHV16, BRR18a, BY16, BMGF^{+}17, CNJ^{+}10, CMFDCLV10, DN12, De 16, Dup14, EAA^{+}18, ECR^{+}16, GLR17, GFA^{+}18, HL15a, HG14, Hav10, HPW^{+}18, HRJH15, JR14, KML^{+}16, LSMW10, PP10, PFF^{+}18, Por10, Roy16, Sat18, SMR^{+}15, SCS14, TTM18b, TK13, WSGM16]. Leste [MTP^{+}17, RDVSC16]. Let [MC19]. lethal [WTA^{+}16]. letter [SGW^{+}15]. level [CD15, CDO16, CPB19, DS18, FSMG10, Hea13, Her18a, HD16b, Jet18, KVA18, Mal18, Man17, PHLM19, RCdLS13, RJ18b, Sev13, SCK15, SPS14, SWS12, Wes18]. Levelling [SJ17]. levels [AAT12, Edd14, SJW^{+}16, WHB^{+}18]. liabilities [CMFG15]. liability [Oya18]. liberalism [RD13]. liberalization [SS15b]. lice [Mis19]. Licence [EE17, CKHJ^{+}17, FG18, KPF17, MF19c, VGG15]. licences [SS19a]. license [TLH18]. licenses [HMK19]. licensing [Ren18, WCK^{+}17]. lies [CGSP13]. Life [SC13a, DG14, GWWH17, HNVZ12, NWW^{+}15, NWW^{+}17, RMJ12, RPM13, San15, TRM19, TJ13, ZSZPR19, VRB14]. light [FW19]. like [Cha14, MFT19]. likely [JFK18]. limit [CMB^{+}15, JFK18, NIFA17, PS12]. Limitations [Cho10, YBD^{+}19, BAA10, CS13, RWW^{+}14]. limited [CMA15, CBL18, CCK19, GA19, GNW19, HPW^{+}18, MHA^{+}17, TRM19, TLA16, THE^{+}18, TLA18]. Limits [ISHP14, AHB18, GMM^{+}13, HHS19, LL15, TML^{+}19, WTA^{+}16, BLCD16]. limpet [DPS16]. line [CdLH^{+}18, Gil11a, HPW^{+}18, KGMP18, LLX^{+}15, SGLY13]. line-caught [SGLY13]. lines [Bor10, Pat15]. ling [HKR19]. link [CNFJ10, MA18, NMvP19, RNS19]. linkages [Ped13, RL19]. linked [CHDJ^{+}19]. Linking [KAF^{+}15, LTM19, MSS19, MN18, NHPD19, BLCD16, WNLS17, HB14, RFG^{+}14a, SDF^{+}14a, WHP^{+}16]. links [DSK^{+}19]. lion [CJB15, MJQMVÁ17]. lionfish [CAGH16, KSK^{+}19, MCCC16]. lions [Man10, SPB19]. Lipstick [Mac14]. Liquefied [vBvTM16]. liquidation [KS10]. Lira [FVM14]. listing [DN12, PC11]. listings [CMH19, Nij15]. lists [Hut19]. literacy [CC18, GLW15, HPPH16]. literature [CC18, CP10, JS12, JS13, JFA17, WCM12]. litter [BGA^{+}15, HPH^{+}18, HP13, Pen18, RM18, RGB^{+}19]. live [FD11, PRT10, UL17]. Livelihood [MBT^{+}19, Mat16, MTP^{+}17, SB15, BD14, FSGB^{+}15, Kat16, PFP17, SPGP12, SMH17, NGS16]. livelihoods [ATAM10, BAHD10, CMW10, FLWG14, HvD13, JLW17, LTM19, MMP18, PRPS17, San15, SWB^{+}14, SCV^{+}17, WBSW13, FSWFG13]. lividus [ONF15]. living [CMW10, Hof19, KJH^{+}10, PRS^{+}19, SI18, Weg15, GBC17]. LME [FMM13]. LMPAs [BMF^{+}16]. Lobitos [MJQMVÁ17]. Lobos [EMP^{+}14]. Lobster [RG19, RNEBG19, RYJ^{+}17, BMW17, BH14b, BRRG^{+}18, CdLR^{+}15, CdLH^{+}18, EBR^{+}18, ECF^{+}18, GSA10, Lal17, LSMW10, NBPF^{+}12, RCdLS13, Uhr16, vPG10]. Local [BPF^{+}16, GSS^{+}18, Kna11, Roy16, SARF16, ABBD17, AK15, BFL18, Ben17a, BD14, BD13, CSH17, CR10, CHBRJ19, CN11, CRCD13, CKN19, CWP16, CFC^{+}18, ELAC19, FGPP18, GFH^{+}17, HZH^{+}13, JRP17, KMD15, KCJW15, KGS^{+}19, LWJH11, LC19, MDE14, MBTV14, Nik19, OM19, RBC^{+}13, RV16b, RCM18, San11, Sha17, SMH17, WM18b, WMSB15, WBGJ18, WF18, YPS13, vdVD16, Gri17]. locally [BH14a, KMD15]. locally-managed [KMD15]. location [BHS19, CSBE14, GSG^{+}15, HA14, HAA15]. locations [EB16, MC14]. locked [BD19]. Lofoten [HBA^{+}14]. logbook [Cha14, ENWL19, WHB^{+}18]. logbook-like [Cha14]. logbooks [And15]. loggerheads [SJP17]. logging [KCJ15, McL15]. logic [BD19]. Lomaiviti [Bol17]. Lonergan [CLW^{+}12]. long [CH12, CQ14, DDS^{+}11, DCY^{+}19, HGH^{+}18, RW15, SS11, SFDB18, Dva19]. long-run [RW15]. long-term [DDS^{+}11, DCY^{+}19, HGH^{+}18, SS11, SFDB18]. Longline [PW15, BKKN10, DPH^{+}15, FPG^{+}17, GOK14, HKR19, Hua11, Pan14, SSW^{+}18, SPRH19, WFK^{+}12a]. longline-caught [BKKN10]. longliners [AHB18]. longlines [BW19]. look [MFT19]. Loreto [AFGRF^{+}17]. Losing [Doy15]. Losinj [MHF13]. loss [CH12, HRLFMP15, KWT17, RGWH18, RMJ12, RPM13, LPS12]. losses [SWS12]. lost [BSW19, Gil15, Joh19, PdSA^{+}13, SBC^{+}14, MS19]. louder [HSRH16]. low [CP15, DDV^{+}13, GNN^{+}17, HD16b, Man17, PR19, RPSS17, SJW^{+}16, WML17, GOTS19, KSK^{+}19]. low- [DDV^{+}13]. low-carbon [PR19]. Lower [LLS^{+}15, AYA^{+}19, Sea17a]. lower-middle-income [AYA^{+}19]. Lucas [GPB^{+}18]. lucens [LYB17]. luck [CE17]. lugworm [TMDSF19]. luxury [Fab16a, SK15, dMAH^{+}18]. Lyme [FJ12, SSG^{+}14].

M [Cho10]. M. [KS18a, WMVR18]. M/V [Cho10]. Māori [Bod17]. MA [JBV^{+}18]. Maasim [DP17]. machines [SWB^{+}14]. mackerel [Han14a, TEF^{+}15, WNB^{+}15]. macro [RU12]. macro-algae [RU12]. macroalgae [HBC^{+}13, WCK^{+}17]. macroregion [RGB^{+}19]. Madagascar [LGH^{+}12a, GLR17, HAGB15, JHN^{+}19, LGH^{+}12b, RLKH19]. Madeira [CLF^{+}19, HD16b]. Mafia [KRM14]. Magnuson [DSR12, KSH^{+}15]. Maharashtra [KSC^{+}15]. main [HB14]. Maine [SMW13, EBR^{+}18, GSA10, Har10, Lee10, MOR15, SLBC19]. Maine/Bay [SMW13]. Mainstreaming [DJR^{+}18, FGR18]. maintain [BCMH^{+}18]. Maintaining [Bes10]. maintenance [LJB^{+}14]. Maio [DB19a]. major [BOF^{+}15, CSCBR10, Dro13, Han16, PAC^{+}19, PHLM19, PGC16, SSWP10]. make [Bro11, vHHB14]. makers [GSB^{+}17, SJ10, VRB14]. makes [Kir19, MBSL13, RL14]. Making [GOTS19, HRA^{+}15, KEM^{+}18, ÖGB^{+}10, SvHdB^{+}15, BDvTvKT17, CKN19, CPRP19, DFP15, De 10, De 16, ERCMD11, GGLP16, HHL^{+}15, LQC10, LM18, LLO16, LLO17, MNM10, PD17, PPT10, PSCY12, RMRN^{+}16a, RBS^{+}15, SCKP19, Sev13, SMW13, TLR^{+}10]. mako [FLT^{+}19]. Malacca [Rah14]. Malawi [Don17]. Malaysia [ATDP16, HIN^{+}17, MK15, MSA^{+}18, PF12b, SRSN17, SSD17a, TTJ18, VMR^{+}18]. Malaysian [AOY16]. Maldives [RAZ16, SW10, ZKR^{+}19]. Malin [Han19]. Malindi [TKAO^{+}19]. Malta [STM16]. Maluku [BB19b, BB19c]. mammal [BAO^{+}12, CLW^{+}12, DA11, Lon11, Lon12, LM11, MRG15, MLV^{+}11, Rei19]. mammal-fishery [Rei19]. mammals [PLL17, SSJ17, WWW10, WLV15]. manage [CAGH16, DJTK17, DGC^{+}17, DWCS18, PKOF18, SL15a, TFBS19, Tec19, TN15]. managed [GMD15, GLJdlTC14, KMD15, MSH13, OCG14, RYJ^{+}17, SBS^{+}13, SFC^{+}15b, TCS^{+}18, TTS14, TINL^{+}14, WSB19, WEKH19]. Management [BH14a, CSF^{+}19, Che12a, CT16, CLW^{+}12, Coo11, FMS17, Fol13, HPP^{+}17a, HGDB10, Hol12, KH12, LCC14, MFF^{+}18, PDB^{+}13, RMRN^{+}16b, TTM18a, Tec19, TBB12, TRBS13, VBHC^{+}17, ZH18, AOY16, AG18, AMLK^{+}12, AH19, ÁFFSCF17, ADGGVL12, AMdB12, ACA^{+}13, Asa12, ACM^{+}12, AAL17, AAL^{+}18, APC16, Bai16, BAA10, BMD^{+}14, BGG19, BFL18, BL18, BPB19, BW15, BHAHW17, BD14, BM13, Bes12, BSKH17, BK16, Big15, Big16, BAB^{+}14, BG15, BCE^{+}17, BJvH18, Bor10, BGS13, BS18, BRR^{+}12, Bre15, BM15, BSP^{+}15, BBE^{+}16, BMGF^{+}17, BD13, BF19, Bur15, BMGH11, BBCPC11, CCR16, CKY^{+}13, CG19, CdLH^{+}18, CSH17, Car11, CR10, CSvH^{+}18, CH16, CVP^{+}17, CMAA13, CC13, CETQ16, CETQ17, CBE^{+}19, CHKG16, CRCD13, CBL18, Cha11, CLLS13, CYI13, Cha16, Che10, Che12b, Che12c, CL13, CCC^{+}14, CQ14, CCD^{+}13]. management [COG^{+}19, CBF^{+}12, CAP15, CHSW17, Coc15, CRO12, CP15, CSS15, CDR^{+}16, CdL17, CHL^{+}19, CBP^{+}18, CSP10, CP10, CHvKM15, DA11, DSK^{+}19, DTJ10, DMC^{+}17, DMMG15, DL19a, De 18, DAJM15, DSR12, DFSB15, DYN^{+}19, DTN10, DM15, DDS^{+}11, DTMJH16, DTM18, Don17, DDV^{+}13, DT10, Dru10, DFK13, Dup14, DMJ^{+}17, EHR15, ECRL17, EGGT12, EGHC17b, EAA^{+}18, ELAC19, ERCMD11, EG15, EYN^{+}15, FOG^{+}11, FBP^{+}18, Far19, FGW15, FR16, FVM14, FB15, FSMG10, FLWG11, FLH11, FAPP17, FDGM16, FBAK14, GB15, GKG16, GNHM18, GNIL14, GMD^{+}12, GRMAC18, Gib10, Gil15, GDD13, GBF^{+}10, GW10, GPA18, GRW^{+}16, GMP^{+}14, GHP^{+}15, GMP^{+}16, GA19, GW11, GCG^{+}10, GM14, GAB^{+}17, GJA16, Gus18a, HK10, HWL^{+}19, HTD^{+}18, HAH^{+}16, HTT10, Har10, HGS12, Hav13, HHL14, HPB^{+}15, HVB^{+}17]. management [HSH14, Her18b, Hey17, Hey18, HGB^{+}19, HD11a, HD11b, HHG10, HRC15, HdV17, HW10, HLRO19, HL10, HNVZ12, HvPN^{+}19, Hor10, HC10, HO10b, HO12, HBC^{+}13, HLL^{+}18, HRA^{+}15, ItDD^{+}18, IOG18, Iri18, IHS13, JBV^{+}18, Jac19, Jae16, JMJ^{+}18, JSD19, JE11, JRP17, JDM^{+}19, KMF^{+}18, KL12, KL14, KBF12, KFBG13, KCKA^{+}18, KS18b, KH13, KW14, Kir19, Kno10, KKG16, KJH^{+}10, KW18, KBC^{+}13, KHP14, KVMM10, KB15, KH15, KDG^{+}15, Lal17, LDF^{+}18, LMA^{+}15, LLH^{+}14, LP16, LBK^{+}13, LRMM17, LR15, LBL^{+}17, LSMW10, LFGM15, LWJH11, LC19, LWS19, LLZ^{+}19, LJL^{+}14, LGDB17, Lon11, Lon12, LCS15, LRS^{+}13, LM11, LTM19, LH16b, LYB17, LKMN16, LBC^{+}16, MOE11, MFB^{+}14, MMM19, MBD13, MBCR16, MLF^{+}17, MRC17, MCMB^{+}18, MB10, MTP^{+}15, MMdJ^{+}14, MNLHRMP17, MSA^{+}18, Mat16, MLF14, MHL^{+}15, MRSL16]. management [MMPD^{+}14, MRCG11, MOR15, McC17, MF10, MHA^{+}17, MB18a, MCH19, MGJB^{+}17, MLV^{+}11, MBTV14, Mis19, MDR^{+}16, MG14, MAHP16, MFAA16, MA15, Mso16, MYA^{+}18, MCAM13, MLHP15, MMA16, MSG18, MDM16, MG16, MNG^{+}18, Nak16, NC10, NM19, NWW^{+}15, ND17, NHA15, NWD18, NFHJ16, NW10, NHPD19, NBMB^{+}18, NBFO18, O’R15, ODF^{+}16, OGM^{+}14, Ols11, OBC^{+}18, OTOL17, OOvdM^{+}11, ODRRM12, ONF15, Pan14, PLL17, PDS15, PD17, Pas14, PO14, PL15, PVVB17, PKOF18, PCCR15, PDMÖ19, PS10, PFCP^{+}14, PPT10, PAD^{+}17, PVA^{+}18, PGPS10, PPR^{+}15, PPCM16, PGP^{+}19, PJS13, PM10, PMC^{+}15, PT10, RvLvTH14, RROR12, RAZ16, RYRYM17, RWK15, RWL14, Rei16, Rei14, RGWH18, RLM10, RLM12, RMB^{+}17, RGGFA17, RU12, RUD^{+}12, RvLG^{+}15, RRRM^{+}15, RRR17, RLHM14, RNCGB15, RPC^{+}17, RO13]. management [RBS^{+}15, RTM19a, SDRML15, SG15a, STIC19, SLJ^{+}11, SKFL15, SMBL13, SLH^{+}13, SBC15a, SBC^{+}15b, SBTF15, SB14, SSB^{+}18, Sch10b, SS19a, SDLT15, Sea17a, SBP^{+}19, SVL^{+}16, SPO^{+}18, Sha17, SI18, Sha10, ST17, SH14a, Shi13, SS15a, SSL16, SBMD13, SMR^{+}15, SCR^{+}19, SM18a, SW10, Smi13, SKS^{+}19, SNP^{+}18, SGP^{+}17, SWM11, Soo17, SHM12, SBS^{+}13, SFCI14, SMQ^{+}16, SLDHS18, SH16, SGMWE14, SJ10, SG12, TLR^{+}10, TK11, TC13, TLB19, TKAO^{+}19, TLA16, THE^{+}18, TLA18, TOL12, Tir17, TJJAM19, TII^{+}19, Tor16, THF^{+}15, TAGP15, TGS18, VSCH16, VC15, Vau17, VH16, VCdR^{+}18, Vil12, WSB19, WL18, WM19, WCM12, WCS18, WHP^{+}16, WSG18, WM18b, WSA^{+}15, Web11, Weg15, Wen11, Wen16, Wen17, WGB^{+}16, WRP^{+}10, WBA18, WBGJ18, WAG^{+}15, WS15b, XJT18, YTTK10, YCA^{+}12, YP17]. management [Yat14, YCY^{+}15, YBRBB16, dBMA13, dVvT11, dlTCL10, dlTCFB^{+}17, vHCT^{+}16, vH15, AFCMET18, ETEW15, FSMG10, GSDB14, KEF18, KED^{+}18, Man10, PFF^{+}18, RSS14, RCM18, Wes18, vHK17]. manager [KLMD15]. managers [CvPH^{+}18, VRB14, dLS10, LSAL19]. Managing [AMdB12, BGC^{+}17, CMB^{+}15, FMD^{+}13, Gue19, Liu13, MCCC16, OHB^{+}18, PWR17, PVPF19, Pom12, SOÖ12, SH12, ACSK^{+}10, BRR18a, Bes12, DK15, EAA^{+}18, FDP19, Gor19, Han16, KC13b, LDS11, PLP14, ST13, TLR^{+}12, DA18]. Manche [RRRM^{+}15]. mandatory [Bro11]. Maneuvering [HLL^{+}18]. manganese [MP12]. mangrove [AG16, ACTG17, ATDP16, CHBRJ19, RBB^{+}16, Roy16, SCK15, SP19a]. mangrove-fishery [CHBRJ19]. Manifestation [Maz17]. mankind [BTC18, JAG16]. mantras [EZBL19]. manual [PSV^{+}14]. manufacture [GWWH17]. many [BVB18, DDP16, RCMBG^{+}19, TLR^{+}10]. map [FHN^{+}18, MTF^{+}15, TM13]. mapped [DSP^{+}18]. Mapping [CSVHS14, DTJ10, FTG^{+}18, FE12, Gue18, NVK^{+}17, SBW^{+}17, TPS15, WSC19, WT18, ADL^{+}18, ATDP16, BAA10, Dru10, DSH^{+}19, GSS^{+}19, GOTS19, HIN^{+}17, HBZH16, KMC^{+}17, LGDB17, MAAB15, TEVN19]. maps [GBK^{+}15, ZZZL18]. Mar [GDD^{+}19, HD11a, JS13, KS18a, LGH^{+}12a, RLM12, Wit15b, WMVR18]. March [Ano10i, Ano11j, Ano12i, Ano13j, Ano14m, Ano15s, Ano16m, Ano17q, Ano19t]. Marennes [GA14]. marginalization [Doy15]. margins [She16]. Mariculture [CP13, SPGP12, CLLS13, RGC^{+}19, YY19, ZLL19]. marina [AFMCMP12, WMK15]. Marine [ALSC13, ACG^{+}17, BR13a, BDG^{+}14, BYK^{+}15b, BB19b, BFGR14a, Bro16b, BHT^{+}18, BRRG^{+}18, BGL^{+}16, CNJ^{+}10, Car11, CSF^{+}13, CC13, CGF14, Cha17b, CPFS^{+}13, CHSW17, CEB^{+}12, DFP^{+}16, DA11, DSM^{+}13, De 16, DMW19, DTMJH16, FLWG11, FLWG14, FAPP17, GSB^{+}17, GGVRVL16, GWM^{+}10, HD16a, HP13, HdVB13, HTF^{+}15, Hor10, HBZH16, HS19, JTP18, JSLL19, JKS13, JL11, Jon13, JLQ16, KLC^{+}15, KTD^{+}15, LBN15, LSTD12, LF17, LLO16, LLO17, LV18, LH16b, MRSH13, Pap16, PRS^{+}19, PDC19, Por15, PKR^{+}17, Qiu13, RvT14, RAF14, RSMM19, RRRM^{+}16, RFG^{+}14b, RNCGB15, RS17, SC13a, SOBS18, San19, STJ^{+}12, SNDB^{+}16, SJ17, SC10, SSFL10, SMKBM16, SRGGVL13, Tod12, THK^{+}18, VGL^{+}17, VGCF18, Vin13, WFK^{+}12b, WWH^{+}17, Wri15, YTTK10, ZPB^{+}13, ZQZ19, vHHB14, vLvHvT12, vdGMA^{+}13, AOY16, AG18]. marine [AARK15, AFR^{+}19, AdSC11, AAC^{+}13, AAGMH17, ABSB15, ABSGB18, ACSK^{+}10, AWH13, AH19, ALMS14, ÁFFSCF17, AVH10, AJ12, ACP^{+}14, AMvT11, ACR18, AMR14, ALM17, AAL^{+}18, ADL^{+}18, AFCMET18, BdKvT19, BAA10, BBN^{+}13, BF18, BS10, BP13, BMF^{+}16, BGA^{+}15, BLG^{+}18, BW15, Ben17a, BD14, BTO^{+}17, BKHA^{+}18, BLCD16, BCP19, BDF^{+}10, Bes10, BSW16, Big15, BB19a, BYK15a, BY16, BDP^{+}17, BWD^{+}19, BJvH18, BD10, BOF^{+}15, BBP^{+}14, BGS13, BAO^{+}12, BE15, BFGR14b, BAHCG14, BDH^{+}18, BBE^{+}16, BD13, BPBM17, Bur16, CMA15, CBN^{+}12, CMP^{+}15, CACL18, CSVHS14, CG19, Car13a, CS19, CETQ16, CBE^{+}19, CL16a, CCC^{+}14, CSBE14, CBG^{+}17, Chu10, CMW10, CMCMHP13, CGM^{+}18, CF13, CHDJ^{+}19, CJB15, CLW^{+}12, Cou11, Cra19, CSMR18, CKHJ^{+}17, CP10, CMB^{+}15, CHvKM15, CvPH^{+}18, DSP^{+}18, DTJ10, DB19a, DGC^{+}17]. marine [DMC^{+}17, DMKL18, DCF^{+}19, DD13, De 10, DJM11, De 13, DFSB15, DCHC17, Dix16, DJR^{+}18, DS10, DG14, DDL^{+}17, DSH^{+}19, Dva19, EAG18, Edd14, EMD^{+}18, EWG^{+}18, EMP^{+}14, EJB19, EG15, EBS18, FMM13, FSB14, FCC^{+}12, FKMV16, FSWFG13, FSGB^{+}15, FEF^{+}12, FÓ12, FJ12, FMB^{+}13, FJG^{+}14a, FHM^{+}10, Fol13, FP19, FSM^{+}12, FLH11, FWK16, FGKMP14, FDS15, GLSA10, Gat18, Gau10, GRMAC18, GSS^{+}19, GLTB^{+}15, GPB^{+}18, GSG^{+}15, GdV16, GBF^{+}10, Gle14, GBC17, GCMA^{+}12, GFC^{+}12, GMP^{+}14, GHP^{+}15, Gor19, GA19, Gra10, GAW^{+}13, GCK^{+}11, GZW^{+}19a, GZW^{+}19b, GCG^{+}10, GCD^{+}11, Gug19, HF16, HBL14, HWL^{+}19, HWY^{+}18, HR16, HPH^{+}18, HGG^{+}18, HDL^{+}13, HPPH16, HPP^{+}17a, Hey17, Hey18, HC18, HHG10, Hob14, HL10, HBP16, HCF^{+}17, HCJ15, HBC^{+}13, HAGB15, Hun13, HTA^{+}15b, HNC14, ISM^{+}17, JLH14, JMJ^{+}18]. marine [JBKB15, JY19, JAO^{+}16, JJ19, JBdL^{+}14, JCD17, JSD19, JSY^{+}19, JQD13, JDQV13, JGMS16, JD16, JLCB16, KMC^{+}17, KBB^{+}15, Kam18, KAF^{+}15, KHT17, KMD15, Kat16, KBF12, KBGF12, KFBG13, KEG12, KGST14, KPF17, KEF18, KCJW15, KRM14, KDS^{+}15, KJH^{+}10, KO11, KMTS16, KMvT15, KW16, KHvH^{+}14, KCJ15, KWT17, KV18, KH15, KMR^{+}13, KLMD15, LPC18, LA14, Lan18, LLF^{+}16, LLA^{+}19, Lea19, LEKB10, LGM16, LOD^{+}15, LMH^{+}12, LMP^{+}16, LCH^{+}13, LRMM17, LMG^{+}16, LPS12, LJB^{+}14, LB13, LJ16, LPR^{+}17, Lin13, LZL19, LWJH11, LKC15, LXWX17, LB18, Lon11, Lon12, LCS15, LM11, LSC14, LLX^{+}15, LEFC19, LHB^{+}17, Lyo14, MVRG13, MMM19, MHF13, MWGD11, MPF12, MD16, MFMP10, MMP18, MCCC16, MSS19, MDE14, MKR12, MSdAM16, MGH14, MMJM15, MTF^{+}15, MMJM16, MMGM16, MWR^{+}17, MKMS12, MSBR15, MSZS^{+}11, MF10, MRG15]. marine [MF12, MA18, McL15, MCH18, MGJB^{+}17, MLV^{+}11, MB15, MBTV14, MKSS18, MSH13, MC14, MSS18, MNA15, MN18, MOH11, MO12, MO13, MSC18, MSGM^{+}12, MAAB15, MD15, MDM16, NVK^{+}17, NKHL18, Nei18, NM19, Nei15, NGS16, Ngo18, NALO19, NWD18, NMJS17, NHPD19, NKBS17, Nor15, NM18, OBJ^{+}12, OBH15, OGM^{+}14, OMF14, OBC^{+}18, OOvdM^{+}11, ODRRM12, OCCvT^{+}18, PLL17, PWH^{+}16, PDT^{+}14, PO14, Pau18, PGR14, Pen18, PTML16, PHLM19, PO10, PBW16, PBW17, PCK^{+}19, PGB^{+}17, PGB19, PTP13, PAD^{+}17, PVA^{+}18, PAA^{+}15, PW16, PS19, PGPS10, PPCM16, PC18b, PS12, PS14, PBJ^{+}14, PPOM16, PV16, QJ13, QSICTdV19, RvLvTH14, RB15, RR17, RMRN^{+}16a, RLKH19, RM17, RM18, RWK15, RRA^{+}10, RFG^{+}13, RFG^{+}14a, RFL^{+}18, RBW18, Rei19, ROD15, RWJ18, RWG13, RV16a, RBWL11, RHP^{+}17, RU12]. marine [RJ13, RWW^{+}14, RUH^{+}14, RvLG^{+}15, RRRM^{+}15, RLHM14, Rog15, RNEBG19, Ros18, RL14, RFHEJK13, RFKEJ^{+}15, SDRML15, SCD^{+}16, SD13, SHS^{+}14, SO13, SCV^{+}15, SLH^{+}13, SBC15a, SSCC19, SM17a, SDF^{+}14b, SDF^{+}14a, SOCA15, SB15, SFG15, SFC^{+}15a, SS19a, Sco16, Sen10, SBM^{+}17, SPO^{+}18, Sha17, SI18, She15, SH14a, SG15b, SL15a, SW10, SWG18, Smi18a, SM18c, SBC^{+}16, SBSK19, SHM13, SS19c, SMW13, SSM15, SHM12, SBS^{+}13, SLS^{+}13, SSG^{+}14, SSG12, ST13, SSBL^{+}15, SSJ17, SY18, ŠST^{+}18, SJT10, TvN13, TÁSLC19, TK14, Tan17, TBKH13, TTM18b, TC15a, TPKL16, TTP12, TTJ18, TC13, TLB19, TDI^{+}14, Thi13a, Thi13b, TH15, TKAO^{+}19, Thu10, THNR12, TR18, TH17, TML^{+}19, Tos11, THF^{+}15, TGdCR11, TCCV16, TKN19, TAGP15, TE16, TGS18, UAZ13, VSCH16, VC17, VS18]. marine [Vir18, VBMM17, WWW10, WMJF16, WSO^{+}14, WK17, WHP^{+}16, WLX19, WW19, WP12, WMSB15, WT18, WH18, Weg15, Wen16, Wen17, Wes18, WKB15, WGB^{+}16, WZCHS10, WML^{+}16, WMWH18, WBGJ18, WMK15, WF18, WAG^{+}15, Wri16, XJT18, YYN^{+}15, YSJU^{+}17, YYYL15, YLZ^{+}19, YPS13, Yat14, ZEP^{+}19, dAdOS17, dGB10, dO13, dVM12, vBvTM16, vLvT10, vdBAMJT19, vdPLMS16, AGB^{+}18, AAT12, ADL^{+}18, BL18, BJGMG18, BE15, BvPS15, CETQ16, CD15, CDR^{+}16, DTM18, EOC15, FVM14, FSS17, FOO^{+}15, FHL18, GR16, Gau10, GSA10, GOdFC16, GLJdlTC14, HDG^{+}12, HR17, HBP16, HCF^{+}16, Kid13, LSK18, LAP^{+}12, LJB^{+}14, LHN18, LLX^{+}15, MHL13, MSA^{+}18, MPCB14, MO14, MLA16, MHM16, MKU19, NBMB^{+}18, PGL^{+}16, PRO16, RAZ16, RDSB10, RFG^{+}14a, RMH^{+}15, RBS^{+}19, SM18b, SGGM14, TTJ18, TLB19, TRCR^{+}19, TZBBL18, Vau17, VGG12]. Marine [dVM12, vLRvH^{+}14, SJ19a]. marine-associated [RHP^{+}17]. Marinha [RPG^{+}15]. Maritime [BMGF^{+}17, CM17, Fri10, FH15, Hua14, Lew16, Rah19, SCC10, SVA^{+}12, ZZ15a, ZZ17a, dVM12, ACSK^{+}10, AL18, Ásg16, BW15, BSL19, Bes19, BHS19, BFGR14a, BFGR14b, Bue15, BA19, CB13, CDO16, DCC19, DSF16, Dol17, DS18, Dro13, DFV15, FFSF17a, FMMA^{+}15, FMGV16, FCdS^{+}15, FW19, GSR19, Ger15, GGD16, GM19, HHL^{+}15, Had18, KPR^{+}17, KPR^{+}18, LG19, Maz17, MH13, Næv17, NPS^{+}19, Nag16, NS18, Nym19, OY15, PCC15, PPML16, PR19, QG17, RSMM19, RRMSSE16, STG11, SFDB18, SSH^{+}19, TEVN19, TGBL11, VGA15, dVM14, dVMdC^{+}19, Weg11, ZP16, dGNC18, dVM10a, De 11, EZPRZ18, vH19]. Market [AYB18, DPB17, RRR17, VRA^{+}12, WBC14, ARSHH16, AKV15, AG12, BW18b, Cha14, CAGH16, FFSF17a, dHJTG17, HB19, IT12, ITNL^{+}14, Jen15, PE10, PKM^{+}19, PWN18, RCF^{+}12, RV16b, SK15, Shi16, SPMW12, SL15b, SL16, SSSO17, THRS15, TBB^{+}10, THA15, VMMB^{+}19, WLT^{+}19b, You16a, ZT19, vdVD16, Car13b]. market-based [CAGH16, Shi16, SSSO17, VMMB^{+}19]. market-related [You16a]. marketing [BLR11, CP18, Knu17, WS17]. marketization [FB17]. markets [GE17, HTA^{+}15a, KS13, LKMN16, MCAM13, PCLE17, SCMS15, SCD^{+}17, SSWP10, ARSHH16]. Marlborough [TMG15]. Marovo [RLAT19]. marsh [RBB^{+}16]. Marshall [GOK14, WL18]. Martinson [Rah19]. Mass [DB19b]. Massachusetts [HDJD15]. massive [CCC^{+}13, HBT^{+}10]. master [Kim15, SSV18]. Masterstroke [SMD14]. matching [BP13]. Material [Dix16, GWWH17, Sea17a]. matrix [BPBM17, MBBDS19]. matter [AWJ19, CBG^{+}17, DWCS18, GPH18, GCW^{+}18, ISM^{+}17, PD17, PVA^{+}18, TMG^{+}18]. matters [CH12, CS17, JLCB16, ZH14]. mature [ITNL^{+}14]. Mau [HvD13]. Maui [SC13b]. Mauritania [Cor14, TGBL11]. Mauritanian [VRMF12]. Mauritius [MBBDS19]. mawsoni [GRLJ15]. maximise [DFK13, GH15, HCP^{+}17]. Maximising [MHC19, Par10]. Maximum [CdLR^{+}15, ACR18, GSC^{+}16, Mes12, NLP11, PA10, RCdLS13]. May [Ano10j, Ano11h, Ano12j, AFR^{+}19, BDH^{+}18, SCJ13, Ano13i, Ano14k, Ano15x, Ano16p, Ano17w, Ano18o, Ano19r]. MCDA [TK14]. MCZ [De 16]. MD [Rah19]. me [DPS16, VGBG15, Kel18]. mean [GFA^{+}18]. meaning [LGFB^{+}19]. meaningful [GCCSS^{+}19, SJ19c]. meanings [LCR16]. Means [LCR16, BTC18, CSBE14, OC17, dMAH^{+}18]. MEAs [AA17]. measure [CK18, DW13, JBK15, PKPdP10, RW15]. measured [VRT13]. Measurement [ABSGB18, LZL19, ZS18]. Measures [TC15b, BFTS^{+}16, DJR^{+}18, GA19, HL15b, Hua11, KTD^{+}15, LGC16, LLO17, MG16, She16, Shi16, SGP^{+}17, TKM^{+}18, UNS^{+}18, WGB^{+}16, You16a]. Measuring [CKdS14, FGW17, FMMA^{+}15, JLCB16, LMH^{+}12, MDM16, PW15, TWA^{+}15, WK14, WLX19, LVRM^{+}18, Ros18]. meat [TJ18]. mechanics [KED^{+}18]. mechanism [CCC^{+}14, CBP^{+}18, MA15, ODRRM12, SVL^{+}16, VMR^{+}18]. mechanisms [Bro18, CMVLGG14, HB19, MF19a, RBM^{+}15, SKL16, Thu10, WBH^{+}15]. Media [LBN15, OO17a, FBM18, KFB19, MSB15, SPB19]. mediating [WWT^{+}16]. mediatised [CKFLO19]. medical [NPAR19, RPA17]. Mediterranean [ACA^{+}13, ACG^{+}17, BMF^{+}16, BBAP^{+}17, CBP19, Cus17, DFP^{+}16, DMdJ^{+}15, FTG^{+}18, GRLM15, GSB^{+}17, GOM17, HP12, JR14, KLC^{+}15, KSK^{+}19, LLO16, LLO17, LPM^{+}14, LBD^{+}18, MMdJ^{+}14, MDK^{+}16, MSB^{+}18, MJRdC16, MB15, PdSA^{+}13, PMC^{+}15, RvdBG13, RRRM^{+}16, RGB^{+}19, RBS^{+}15, SKFL15, SNIN17, SH12, TK14, ZCD^{+}15]. medium [KLM^{+}13]. medium-scale [KLM^{+}13]. meet [YPS13]. meeting [MMM19]. meetings [HL15a]. meets [AGB13]. megafauna [ADL^{+}18, CMB^{+}15, HAGB15, TTJ18]. Megaptera [BSP19]. Megrim [MACM14]. Mejillones [GCP19]. Mekong [HvD13]. Melanesia [AAL17]. meltdown [CCTG^{+}13]. Member [CKVO16]. members [CvPH^{+}18, RPdO^{+}19]. membership [KGAB14, PD17]. Memoranda [MF10]. Menai [CSS15]. Mental [KKWS15]. menu [SS16, dMAH^{+}18]. mer [PWN18]. merchant [MM19, RMJ12]. mesh [BHG11]. message [MSBR15]. meta [DL19b, FFH18b, Won19]. meta-analysis [DL19b, FFH18b, Won19]. metaphors [Nei18]. method [AG12, JBK15, Kru16, LVRM^{+}18, LB13, LWvdBL16, LLO17, MMP15, PVK18, TML^{+}19]. methodological [NWW^{+}15]. Methodologies [SSB^{+}18]. methodology [AAL^{+}18, CW15, CK18, Das16, dlFFVRC16, PKPdP10, ZCD^{+}15]. Methods [VGG12, DPJ^{+}13, GGSRVL17, GRW^{+}16, HAK^{+}19, MDM16, NWW^{+}17, NHPD19, SSC^{+}15, WS14, WMSB15]. métier [STIC19]. métier-based [STIC19]. métiers [MG19]. Metrics [EGD18, CKdS14, RWW^{+}14, TLR^{+}12]. metropolitan [CL18, MAR^{+}15]. Mexican [AFCMET18, MCMB^{+}18]. Mexico [CR10, PSCY12, AJ17, Ben17b, BRM^{+}17, CSF^{+}19, CBF^{+}12, CST10, CSCBR10, CMCMHP13, CSMB^{+}12, ETEWZD11, FB15, GLR17, GSS^{+}18, HGB^{+}19, JWBJ13, KYS10, Kai15a, Kai15b, Kai16, Kai17a, Kai17b, Kai17c, MMPD^{+}14, MLTBMB14, PS11a, Ped13, PJS13, RROR12, RRAH13, SARF16, STIC19, SAdC15, TMHJ12, UEMRG^{+}18, VAW14, YWC^{+}16]. MEY [GMM^{+}13]. Micro [TINL^{+}14]. Micro-economic [TINL^{+}14]. microbiological [AYR^{+}14]. Microplastics [Far19, CL18]. micropolitics [Kam18]. microscope [MGJB^{+}17]. mid [BSKH17, BK16]. Mid-Atlantic [BK16]. mid-water [BSKH17]. middle [AYA^{+}19, VHRM17, Por15]. Middlemen [CNFJ10, Ped13, Bre13]. Migrant [WWT^{+}16]. migrants [Cus17, THRS15]. Migration [NW10, CR11, TRCR^{+}19]. migratory [GJ15, KCJ15, McL15]. miles [Kay15]. militarisation [Nor15]. Milne [BSM^{+}14]. Mind [AdSC11, Gor19, CCL^{+}19]. Mindanao [FGPP18]. mine [FG18]. mineral [PKA^{+}16, WL18]. minerals [BSAT18, BS18, Lil16, WM18a, FG18]. Minimising [SAH14, SHM12]. minimization [SOS12]. minimum [CVFP^{+}12, WHB^{+}18]. Mining [Smi18b, ARJ18, BTC18, CCC^{+}13, CKC^{+}13, DMJ^{+}17, DLM^{+}18, ECRL17, FFH18a, HBT^{+}10, Jae16, JAG16, JGA17, JML16, JDM^{+}19, Kim17, LM16, LACG18, LMG^{+}16, LJL^{+}14, MPPB10, MB18b, OCW^{+}17, OHB^{+}18, VAHG^{+}18, VVS18, WM18a, vD16]. Ministry [Bro16b]. mis [Jac19]. mislabeling [DL19b, WLT^{+}19b]. Mismatch [ONF15]. mismatches [AG18]. misrepresentation [SWG18]. misrepresentations [Bra12]. missed [Gau10, MHF13]. Missing [Wea16, MA18, VHRM17]. Misunderstandings [EZBL19]. misuse [Coc15]. mitigate [EE17, GCMA^{+}12, PBW16, PBW17, RRG^{+}12]. mitigated [SSC^{+}15]. Mitigating [ATB^{+}19, RGBT16, PAC^{+}19, RBW18]. mitigation [BB16, FSCJM19, Gil11b, LPS12, MG16, PC13b, SSC^{+}15, SRGD18, SOS12, TJ18, Zer18]. mixed [GMM^{+}13, MDM16, QSMGVA18, STIC19]. Miyagi [SOUN16]. Mobilising [Bre14, DHR16]. mobilization [GMP^{+}16]. mode [SFG15]. model [DRPSA17, DSK^{+}19, DRH^{+}18, DMJ^{+}17, GDD13, GSDB14, HAH^{+}16, KC13b, LZ16, MMC^{+}14b, NBM15, PRS10, QZH^{+}18, RM17, SCD^{+}16, Shi13, TRC^{+}17, WCS18, WHP^{+}16, Zag18, vHK17, vPLB^{+}13, BLCD16, DRG^{+}16]. Modeling [AYR^{+}14, KS12a, KTD^{+}15, FVT19, IOG18, Kai15b, KEM^{+}18, LSP^{+}17, NMvPNL15]. modelling [CCD^{+}13, CSMR18, GSG^{+}15, HRA^{+}15, HGBJ16, KPR^{+}18, LB13, MGH14, RUD^{+}12, SPO^{+}18, TMDSF19, YFC13]. models [CSF^{+}19, HRA^{+}15, MGB^{+}18, PGPS10, PAA^{+}12, TRM19, TPKL16, TDI^{+}14, WK17, vTRvLvH14]. moderator [LEG19]. modern [VPL^{+}13]. modes [BGS13, dVDA^{+}16]. modifications [EFK^{+}19]. modified [BBE19]. money [DPV^{+}15, PRPS17, SSV18]. monitor [KW16]. Monitoring [BB19a, HGH^{+}18, LSAL19, MSH13, SvHdB^{+}15, SBS^{+}13, SSG^{+}14, BLG^{+}18, CVFP^{+}12, CSVHS14, CLS10, Cha11, ENWL19, Gil15, HWN11, He18, HMS^{+}17, HGB^{+}19, LLO16, Nym19, PHEM^{+}17, RRRM^{+}15, San18, ST17, SFC^{+}15b, SF16, TKAO^{+}19, ZPB^{+}13, LOD^{+}15]. Montreal [RM17]. moral [Pin15]. morality [CN17]. Morals [SSV18]. moratorium [ECSE^{+}17, KGMP18]. Moray [BMGH11]. Moreton [vdGMA^{+}13]. morhua [HBG12, MG14]. Morone [MSG18]. mortality [BW19, DB19b, GA14, MPT13, WIC18]. most [GOTS19]. motivations [ASBAM13, YFB16]. motives [GGK17]. Mounds [De 13]. mountain [DL19a]. move [RMRN^{+}16a]. movement [And15]. Moving [ACA^{+}13, BIPS13, CBP^{+}18, DL19a, DS18, FDP18, KTA^{+}13, PPR^{+}15, RMB^{+}17, CdLR^{+}15, RCdLS13, Che11, HTT10, JWM^{+}16]. Mozambique [LEFC19]. MP13 [BR14]. MPA [RAZ16, SGNM19, She14, BHT^{+}18, CGSP13, DSWG13, FSM^{+}12, Har14, LHH15, MG19, MG14, RRRM^{+}15, SC13a, Tam15]. MPAn [SM18b]. MPAs [CD15, HCF^{+}16, LAP^{+}12, MSA^{+}18, BDP19, CHL^{+}19, DFP15, GJ15, JS18, KH15, MK15, OBJ^{+}12, PHM16, VGG15, VGBG15]. MSC [LSK18, BTOM13, OHQ^{+}16, PRLCL12, TM13, WC17]. MSC-certified [OHQ^{+}16]. MSFD [BFGR14a, PRO16, BFGR14b, HJKR14, SvHdB^{+}15, vTRvLvH14]. MSP [BFGR14a, FÓ12, BFGR14b, FCdS^{+}15]. MSY [GMM^{+}13, KML^{+}16, Mes12, RMB^{+}17]. much [BRR18b]. muck [DHM^{+}17]. muddy [KP13]. Multi [Jet18, KSS17, RBS^{+}15, SWM11, AARK15, ASP16, CCL^{+}19, CDO16, CE13, ET19, EG15, EGF^{+}17, FDGM16, GCD^{+}15, GMM^{+}13, HF16, HPP^{+}17a, Her18a, ITNL^{+}14, KML^{+}16, KS12d, LLF^{+}16, LLA^{+}19, MMP18, Man17, MOH11, MFAA16, NMvPNL15, NBF^{+}18, PVVB17, PV16, RL19, SKFL15, San18, SK18, TvN13, TWA^{+}15, VvDM10, WCS18, WB18, WNB^{+}15, WBH^{+}15, Zag18, vOHO^{+}17, RJ18b]. multi-community [FDGM16]. multi-country [Zag18]. Multi-criteria [RBS^{+}15, ASP16, CCL^{+}19, EG15, GCD^{+}15, HF16, HPP^{+}17a, PVVB17]. multi-disciplinary [AARK15]. multi-factor [TWA^{+}15]. multi-fleet [CE13, GMM^{+}13]. multi-functional [EGF^{+}17]. multi-gear [MFAA16, NBF^{+}18]. multi-jurisdictional [WNB^{+}15]. multi-legislative [PV16]. Multi-level [Jet18, CDO16, Her18a, Man17, RJ18b]. multi-objective [KML^{+}16]. multi-region [RL19]. Multi-regional [KSS17, SK18]. multi-sector [NBF^{+}18]. multi-sectoral [MOH11]. multi-species [GMM^{+}13, ITNL^{+}14, KS12d, NMvPNL15, NBF^{+}18, SKFL15, WBH^{+}15]. Multi-stakeholder [SWM11, VvDM10, WB18]. multi-stream [WCS18]. multi-stresses [MMP18]. multi-trophic [ET19, vOHO^{+}17]. multi-use [LLF^{+}16, LLA^{+}19, San18, TvN13]. multi-use/user [LLA^{+}19]. multidimensional [HAC19, RMB^{+}17]. multidisciplinary [CLC^{+}16]. Multifactor [MC15]. multilateral [AAGMH17]. Multilateralism [Rah14]. Multilevel [MCAM13]. Multilingual [TS14]. multiple [HF16, HO12, IT15, KRM14, KED^{+}18, MMM19, MMJM15, MMJM16, OCW^{+}17, PV16, RRRM^{+}16]. multiple-use [HF16, KRM14, MMM19, MMJM15, MMJM16, PV16]. multiplier [JLH14]. multiregional [HK18]. multispecies [BKH16, Kas16, KBBR16, PM14, WSG18]. Mumbai [SG17]. municipal [DP17, Fen17]. muses [GNN^{+}17]. mussel [RRR17, RVdCGN11, vdBKB^{+}17, CS14]. Mussel-farming [CS14]. must [AMLK^{+}12]. mutual [HDJD15, Hut19, MP12]. mutually [Han14a]. my [CJ16]. Myanmar [QG17, TBJC18]. myth [BJvH18, NAB^{+}18]. myths [CMS19, EZBL19].

N [RGGFA17, DAGG16]. N.S.W. [FBM18]. Nadu [CA17]. Nam [HvD13, KSS17]. Namibia [SC10]. Namibian [Dra15, Pat15]. Nang [PHC14]. Nantucket [JBV^{+}18]. NARDL [BB19d]. Narragansett [DTJ10]. narrative [AWJ19, MSC18]. narratives [CJ17, Kel18]. Narrowing [MMBK17]. National [AFGRF^{+}17, BL18, BGL^{+}16, CKY^{+}13, Cli13, HF16, HGS13, KC13b, KNY14, LM14, Qiu13, San19, Thu10, TMKK^{+}15, VGL^{+}17, ARJ18, BW15, Ben12, BY16, BDP^{+}17, BE15, BAHCG14, Bro15, CG18, CRCD13, CHDJ^{+}19, DHC11, DW13, De 18, DG14, EKSZ14, EBS18, FSGB^{+}15, FKWV19, FP19, FJA^{+}14, GRMAC18, GSS^{+}18, HD16a, HHG10, HR14, Hou14, JBBV^{+}19, KM10, KH11, Lea19, LY14, LJT19, MLF14, MDNT19, MO13, MK14, MYA^{+}18, NPS^{+}19, PVS^{+}19, RDSB10, RUH^{+}14, Sco16, SJW^{+}16, SWG18, TvN13, TDMN19, WL18, WW19, WAG^{+}15, YPS13, Bro18, GSS^{+}18, dVM10b]. Nations [DG14, GDN19, RFL^{+}18, Fla13, MSGM^{+}12, NYJW15, YBD^{+}19, ZS18, Smi17]. nationwide [WLV15]. native [FSV^{+}16, Han16, Mat14, SHS^{+}14, SS15a]. Natura [Foc11, PW16]. Natural [SFCI14, KED^{+}18, LJB^{+}14, Mis19, PD17, Sch10a, SGN^{+}17, SOS12, SGMWE14, vBvTM16, vHFJ^{+}19]. Nature [Qiu13, CHSW17, Deb18, ESGC13, KEG12, KMR^{+}13, Ols11, SGN^{+}17]. nature-based [KEG12, SGN^{+}17]. Nautical [KP17, Kay15]. Nautilus [FG18]. Naval [Rah19]. navigability [APY^{+}17]. Navigating [Ben18, Kar17, San11, SSM15, ZSZPR19, LJ16]. navigation [LG19]. Navigational [KP17]. Near [HDG^{+}12, CMW10, HGS13, LBK^{+}13, LBP^{+}19, WF18]. near-shore [LBK^{+}13, WF18]. Near-term [HDG^{+}12]. nearly [WT18]. nearshore [BAA^{+}15a]. necessary [Bro16a, PdSA^{+}13]. Need [Kel18, BFTS^{+}16, BBE19, Cha15, Cor14, DMdJ^{+}15, NNPV17, QZH^{+}18, SSJ17, VKGN^{+}14a, WS15a, vHK17]. needed [GPO^{+}10, Pla19]. Needs [Kel18, FCC^{+}12, Gle14, LDS11, RPdO^{+}19]. neglected [Wea16]. negotiated [MLR^{+}16]. negotiating [BDP^{+}17, DHR16]. negotiations [De 17, Lea19, Sco16, Yia18]. Negros [Gor19, MSS19]. neo [RD13, TH15]. neo-communal [TH15]. neo-liberalism [RD13]. Neoliberal [TH15, Jen15, Lan15, Pin15, Pin17]. neoliberalera [WS17]. Neoliberalism [PD15, AJ17, BMW17, Bre15, Gek10, Lal17]. neoliberalization [Ben17b]. Nephrops [HNVZ12]. nest [AJ12]. nested [RvLvTH14]. Net [NLP11, Car14, KLY14, Knu17, KSC^{+}15, KLMD15, LPS12, MLHP15, STM16, SC13b]. Network [MB10, SBC15a, CMA15, CGM^{+}18, DFSB15, FVH18, GJ15, GMP^{+}14, GAW^{+}13, GCG^{+}10, Har10, HBP16, KMF^{+}18, LJ16, MLF14, MLHP15, NMvPNL15, OBJ^{+}12, RFG^{+}14a, RRRM^{+}15, RS17, SOÖ12, HGH^{+}18, RR18]. Networks [Dra15, SM18b, BS10, CGF14, Fen17, Grö18, LF17, MJRdC16, MJQMVÁ17, OC17, RO13, VSA^{+}12]. new’fisheries [dCW15]. Newfoundland [CPKF19, Dav15, EAA^{+}18, KSB15, SvZdJ15]. newly [HR17, VPL^{+}13]. News [PKA^{+}16]. next [DS18, FH15, MGH14, OBJ^{+}12]. nexus [CMMBV^{+}19, HP12]. Ngazidja [HDM13]. NGO [CBN^{+}12, DJ15, HGS12]. NGOs [BDP^{+}17, SALS17]. Nicoya [LH16b]. Nigerian [CDO16]. Nile [SNIN17]. Ningaloo [CvPH^{+}18]. No [AAL^{+}18, AFCMET18, BRRG^{+}18, CS19, KBF12, LPS12, MSS18, NKHL18, TM13, VAW14, Bre14, KBGF12]. no-take [AFCMET18, BRRG^{+}18, KBF12, NKHL18, TM13, VAW14, KBGF12]. NOAA [HGH^{+}18, SD18]. NOAA/NPS [HGH^{+}18]. nodules [MBCR16, MP12, VVS18]. Noise [HGH^{+}18, KS16, KS18a, PTN^{+}17, WWW10, WMVR18]. nomenclature [BF15]. Non [LLF^{+}16, BW18b, BD10, BRM^{+}17, CN17, CMW10, FSB14, FSV^{+}16, FHL18, HvD13, Han16, HGS13, Jen15, LOD^{+}15, LCC14, Mat14, OGM^{+}14, PDMÖ19, RCF^{+}12, RLM10, RLM12, RFKEJ^{+}15, SHS^{+}14, She16, SS15a, SCS14, THNR12, TLH18, Cus17, DPS16, RFHEJK13, Wis14, UA13]. non- [HvD13]. non-commensurate [RLM10, RLM12]. non-commercial [TLH18]. non-compliance [BD10, CN17, DPS16]. non-consumptive [BRM^{+}17, FSB14, HGS13, THNR12]. non-extractive [RFKEJ^{+}15, RFHEJK13]. non-fishers [CMW10, UA13]. Non-governmental [Cus17]. Non-implementation [Wis14]. non-indigenous [FSV^{+}16, LOD^{+}15, LCC14, OGM^{+}14]. non-industrial [SCS14]. non-market [Jen15, RCF^{+}12]. non-native [Han16, Mat14, SHS^{+}14, SS15a]. non-participation [FHL18]. Non-sectarian [LLF^{+}16]. non-state [PDMÖ19]. non-targeted [BW18b]. non-tariff [She16]. noncompliance [KS10]. Nordic [JE11, NAB^{+}18, TKM^{+}18]. Norfolk [Rei10]. normal [Bre14]. Normalization [BA19]. normative [APD^{+}19, DJTK17, dlTCL10]. norms [AL17, LJ17]. North [BHG11, HD11a, Rei10, CDR^{+}16, HEJK10, HCC^{+}19, Kim15, KR18, ÁFFSCF17, Asa12, BDSB16, BVB18, BHM16, BWRT17, BBE^{+}16, CETQ16, CRO12, DR10, DB19b, DPH^{+}15, Fab16a, GJA16, HG14, HD11b, HM18, Jet18, JFK18, KML^{+}16, KBBR16, MACM14, MRCG11, MLV^{+}11, MRSH13, MUHH18, MPT13, MNG^{+}18, Nij17, PHM16, PD15, Por10, RWG^{+}19, RJ13, SRA15, TM13, VvD10, VvDM10, Yia18, vdBKB^{+}17]. North-East [CETQ16, Yia18, ÁFFSCF17]. North-Western [GJA16]. north/south [HEJK10]. Northeast [CKdS14, GSC^{+}16, SGMG18, PQLH15, Chu10, CL16b, Das16, ERD^{+}13, GBW16, GLW17, KS10, OHQ^{+}16, SS15b, TEF^{+}15]. northeastern [Kim10, Knu17, MSB^{+}18, Ols11]. Northern [PDR11, Asa12, CPKF19, GCP19, GLTB^{+}15, LSMW10, MACM14, NAYSN18, PAC^{+}19, TMHJ12, WNB^{+}15, AFMCMP12, BC13, CYI13, Coo11, EAA^{+}18, FSB14, FGPP18, LK15, MGB^{+}18, MLA16, MLTBMB14, SSCC19, SAGG13, TR18, YPS13, ZMZ16, ZTAO13]. Northumberland [TMDSF19, TPS15]. northwest [FB15, LRJ^{+}19, GCK^{+}11, KAPF^{+}19, Nak16]. northwestern [ACA^{+}13, Cha15, GKG16, MMdJ^{+}14, OIS^{+}18, KDW10]. Norway [BHA^{+}14, ECF^{+}18, GAB^{+}17, HHL^{+}18, Han13a, Han14b, Hee14, HD10a, HID12, HUHA18a, HUHA18b, HRJH15, Iri18, IHF15, JA10, KGS^{+}19, LOS11, Mis19, NPS^{+}19, Ola18, ØR18, SS19c, SS15c, SSA16, TMR^{+}14, THRS15, TR18, Weg11]. Norwegian [BD19, CJ17, ET19, GGL17, GBBB15, HHL14, HMK19, HSR^{+}18, JBBV^{+}19, KHD12, KGS^{+}19, KNY14, Næv17, NS18, OOvdM^{+}11, SGZI19, SH15, SA18, SH16, SS13, Tho13, TS14, THM15, TBB12, THRS15, Tir17, Wen16]. Notarbartolo [WFK^{+}12b]. Note [Ano11k, Han19, WPJ10, Ano17x, Ano19x]. nothing [AGWMP19]. Notice [PPM19, Hey18, TLA18, Wit15b]. Notions [KSB15, Ant16]. Nova [BMW17, GLW15, SM17a, SMW13, WB18]. novaeangliae [BSP19]. novel [CGM^{+}18, GFMG^{+}14]. November [Ano10h, Ano12h, Ano13h, Ano14j, Ano15v, Ano16u, Ano17m, Ano18t, Ano19o]. NOWPAP [CL16b]. noxious [CPRP19]. NPS [HGH^{+}18]. NSW [VBMM17]. nuanced [HGBJ16]. nudge [MJvP^{+}18]. nuisance [Kas16]. number [GCW^{+}18, MMBK17, Pas14, WK14]. Nunavut [Wen11]. nurseries [FMP^{+}18]. nutrient [GE17, KNB^{+}18]. nutrients [GSW18]. NW [FVM14, GLVLGG19, PFMA^{+}19, dO13, RVdCGN11, VRMF11].

oasis [JSG^{+}18]. obedience [Sch15b]. object [Mis19]. objective [KML^{+}16]. objectives [BB16, BSP^{+}15, DFP15, DTM18, DDvP^{+}16, EGGT12, FR16, GAFD^{+}18, HLRO19, PDB^{+}13, RMB^{+}17, San19, Sol14, THF^{+}15]. obligation [Boh19, CRSM^{+}17, EFK^{+}19, PHEM^{+}17, VPR^{+}16, Wit19, dVDA^{+}16]. obligations [EB16, HID12, MN18, VAHG^{+}18, WBGJ18, Oya18]. obliged [EGHC17a]. observation [KC15]. observatories [CLC^{+}16]. observed [TPS15]. observer [WHB^{+}18]. observers [Por10, SD18]. observing [MDG^{+}14]. Obstacle [LT19]. Obtaining [VGG15]. Occidental [Gor19, MSS19]. Occupational [HUHA18a, HUHA18b, ML19, CA17, Dav12, GSJ18, Tho13]. occurring [Mis19]. Ocean [AGWMP19, BGS15, CDO16, DS18, GGD16, GLW15, GDD^{+}19, GPO^{+}10, HGH^{+}18, LJ17, SJ19b, TR18, TMKK^{+}15, VBSD17, VKGN^{+}14b, Weg15, dVM10a, BFO^{+}16, BKHA^{+}18, BAB^{+}14, BPF^{+}16, BG15, CL18, CBG^{+}17, CC18, CLC^{+}16, DHFL^{+}16, DBG^{+}16, FMD^{+}18, FMKT15, FMD^{+}12, FMD^{+}13, GW15, GPH18, GCCSS^{+}19, Gue19, HLM^{+}13, HKHG19, HC18, HBS^{+}19, HDJD15, HL15b, HdF^{+}19, Hou14, JHW13, Kir19, Kno10, KN17, KRK13, KHP14, LEKB10, LMP^{+}16, MF19a, MF19b, MDG^{+}14, MHL^{+}15, MR10, NJ18, Nym19, OL11, PHAC15, RYD18, RLH18, RvdBG13, RDG^{+}13, SLGC^{+}18, SHL^{+}18, SSL16, TLR^{+}12, TBB^{+}10, TTUR14, VKGN^{+}14a, dVMdC^{+}19, WB19, WF18, Wri15, WOdG^{+}16, Zer18, ZS18, ZHH14, ZMM^{+}14, vH19, vHFJ^{+}19, vPJH^{+}17, Laf14, APY^{+}17, ÁFFSCF17, CYI13, DMMG14, DMMG15, DPH^{+}15, ECL^{+}15, EYN^{+}15, GDN19, GPB^{+}18, GDD^{+}17, Hav10]. Ocean [Hav13, JL11, KSWS18, KC15, LB18, MBM14, MSGM^{+}12, NM12, OHH^{+}16, PH16, Ray19, RWG^{+}19, SRSN17, SDP^{+}16, dOSL18, SPP18, SDHS19, SDB14, Tan17, TJJAM19, VHRM17, VGA15, You16b, ZH18]. ocean-derived [DHFL^{+}16]. ocean-going [DBG^{+}16]. ocean-users [GW15]. Oceania [NNP^{+}14]. Oceanic [CL19, HB19, BAA^{+}18, CLLS13, GFC^{+}15, JSG^{+}18, KED^{+}18, MC14]. Oceanographic [HD16a]. oceanography [MKU19]. Oceans [Men18, STB16, VSS^{+}15, BMD^{+}14, Ben18, Big16, Bla17, BWD^{+}19, GAP12, Han19, KM10, MSTB19, RBM^{+}15, SF13, Vin13, Vir18, dVM17, vHFJ^{+}19]. October [Ano15t, Ano16v, Ano17u, Ano18r, Ano19p]. octopus [DYN^{+}19, PFCP^{+}14, RPdO^{+}19, SCR^{+}19, SGP^{+}17, VRMF12, Wos19, DYN^{+}19, VRMF12]. ODA [dlÁPAVS19]. odds [VPR^{+}16]. off [DDV^{+}13, GJA16, HSW17, KH13, MV16, RAG18, dMAH^{+}18]. off-shore [MV16]. offer [Sco18]. Official [CBP19, MCY^{+}19, WLV15, MHC19]. offs [EAA^{+}18, FMB^{+}18, OO19, Pin13, SCMS^{+}18, SPRH19, WCMHO18, WLT^{+}19a]. offsets [SRGD18, VLPC14]. offsetting [JTP18, NMJS17]. Offshore [KS12b, KS12c, Kai15b, Kai17b, TC15a, dAdOS17, AMR14, BDF^{+}10, BN19, CSVHS14, Cha17a, Cha16, De 13, DFL15, ECRL17, GSG^{+}15, Gue18, HL15a, HHP17, HA14, HAA15, HHA17, JKS13, KYS10, KS12a, KS13, Liu13, MMOL15, Mil14, NKAP13, RDP^{+}18, SS19a, SAGG13, Tec19, VLPC14, VHA^{+}19, VLKN17, WKB15, YWC^{+}16, ZZZL18, vLvT10, vdBKB^{+}17, RRMSSE16]. Oil [Ber19, JSY^{+}19, CMFG15, CSO17, CL16b, DZLL15, GSS^{+}18, GNW19, GDD^{+}17, GDD^{+}19, JLY18, JKS13, KYS10, Kai16, MJQMVÁ17, MNG^{+}18, OHH^{+}16, PQLH15, PML10, SS13, TL16, UL17, WZF^{+}18]. oil-spill [SS13]. Old [Rei10, HFM^{+}13, Mee11]. olds [FMD^{+}18]. Oleron [GA14]. Oman [ABSB15, ABSGB18, ASBAM13, BBA15, BAMAH17]. omission [Hob14]. on-going [FJA^{+}14, RNCGB15]. on-ground [FAPP17]. on-growing [HBG12]. on-the-water [GCK^{+}11]. once [TDMN19]. One [AAL17, Edd14, YC18, DRH^{+}18, MSS18, SH14b, WFK^{+}12b, YC18]. only [Alf17]. Onshore [CPKF19]. ontological [Dix16]. ontologies [NM19]. open [ABSB15, ABSGB18, BMD^{+}14, RAG18, RW15, TTS14, VPL^{+}13]. open-access [RAG18]. opening [vD16]. operate [CKHJ^{+}17, FG18]. operating [Bro11]. operation [BHT^{+}18, FSS15, GHP^{+}15, KMF^{+}18, PIH^{+}10]. operational [CVP^{+}17, MGB^{+}18, ÖGB^{+}10, PCK^{+}19, Shi13, TE16, vdBAMJT19]. Operationalising [BAA^{+}18]. operationalize [PL15]. Operationalizing [GHM^{+}18, Par15, SSD^{+}16, SGMWE14, TJJAM19]. operations [Ard18, BHA^{+}14, KFM17, SCC10, Rah19]. operators [SJW^{+}16]. opilio [Han16]. opinions [FB17, HJKR14, HHA17, PHEM^{+}17, RPdO^{+}19]. OPOL [FW15]. Opportunities [AAGMH17, GDD^{+}19, ARSHH16, CSF^{+}19, CGM^{+}18, EBR^{+}18, EYN^{+}15, FOO^{+}15, GAFAV16, Har15, HA14, Lan15, LB17, LLO16, LKMN16, MNG^{+}18, NVK^{+}17, NKAP13, PDR11, PS19, PKR^{+}17, RFG^{+}14b, RGB^{+}19, SSC^{+}15, SBT16, SSSO17, TBJC18, VMMB^{+}19, Wit15b, Wit15a, DK16, GDD^{+}17, RPSS17, SBC^{+}14, WBSW13]. opportunity [BMF^{+}16, BWRT17, CO16, Gau10, HKR19, Hun13, LT19, MHF13, Mar17, Men19]. Optimal [Kas16, CdLH^{+}18, MMTC16]. optimisation [EB16]. Optimising [BLG^{+}18, BAA^{+}15a, EYN^{+}15]. optimism [Nym19]. optimization [OLF19, RRMSSE16]. optimum [PL15, SV15]. option [OL10, TFBS19]. options [ADGGVL12, BL18, DBG^{+}16, ESM^{+}19, GCG^{+}10, HHL^{+}18, HK11, PNW18, SVA^{+}12, WDK^{+}13, Zha18, DG14, JGA17]. orange [FvRA11]. order [CS19, RFL^{+}18]. Orectolobiformes [BFTS^{+}16]. organic [GVSZ19, LJS^{+}18, dlFRG17]. organisational [FGPP18, THM15]. organisations [AMdB12, AM15, GLVLGG19, KR18, dBMA13]. organised [VWG19, vAS19]. organism [Mis19]. Organization [KAPF^{+}19, BAT^{+}18, OR19, Wen16, SVA^{+}12]. Organizational [VDB16, NPS^{+}19, Soo17, SOÖ12]. organizations [CAP15, CSP10, KGAB14, Oan18, PKOF18, PDMÖ19]. Organizing [Wen16, Wen17, TWJ^{+}17]. orientated [BH14a]. orientation [BCP19]. origin [AG12, He18, LHN18]. origin-differentiated [LHN18]. Origins [DAB^{+}14]. Orkney [JKS16]. ornamental [BB19a, MLF^{+}17, SO13, SCV^{+}15]. Ortuño [McL15]. Oslob [LTM19]. OSPAR [MLF14]. Osteoglossum [MLF^{+}17]. Ostrom [TTM18b]. other [DJR^{+}18, Lon11, NM18, SGLY13, SOS12]. our [AJ17, GBC17, She15, dlTCFB^{+}17, vHFJ^{+}19]. Outcomes [SKS^{+}19, AM15, BP13, BW18b, Bro10, Fab16a, FAPP17, GLR17, Gib10, Has18, Lew16, MLR^{+}16, TII^{+}19, WP12, vHCT^{+}16]. Outer [KS12a, KS12b, ZMM^{+}14, KS12c, Mso16]. outfall [Bre14]. outlined [RFL^{+}18]. outlook [DDS^{+}11]. output [GGSRVL17, GHO^{+}17, LY14, MO13, WW19]. outreach [LMH^{+}12]. outweigh [MSG18]. overall [FDS15]. overarching [MNA15, GAFD^{+}18]. overcapacity [MNHN18, Pom12]. overcome [GAFD^{+}18, NWW^{+}17]. Overcoming [CPR14, CHvKM15, LT17, RJ13]. overexploitation [Hof19, PC18a]. overexploited [BAT^{+}18, OL10]. overfished [FP13]. overfishing [FP13, MNHN18, NIFA17]. overlap [WS14]. Overlapping [GSJ18]. overlaps [HJN16]. Overseas [MCY^{+}19, FLWG11, RBC^{+}13]. overuse [CSCBR10]. Overview [BLC16, BTS16, DK12, HBC^{+}13, LLZ^{+}19, LKMN16, LBD^{+}18, MCY^{+}19, MSG11, RU12, SH14a, SBT16, VWG19, WP17]. Own [RW15, TN15, vHHB14]. Own-price [RW15]. owner [EAE15]. ownership [AK15, Bro16a, CG13, KCJW15, Kna11, LU13, RWG13]. oyster [BPMY18, EAE15, FVH18, GA14, HMH18, LKC15, PR13, Ren18, SCD^{+}17, YFR10, YFC13, YCHM14]. Oysters [PR13, FVT19].

pace [BvPS15]. Pacific [CYI13, Hav10, Hav13, MBM14, Smi18b, SDB14, All14, ACR18, BAA10, BBN^{+}13, Bar10, BMPV10, BAA^{+}15b, BAA^{+}15a, BCMH^{+}18, BAA^{+}18, BSH^{+}19, BSAT18, BS18, Cha15, CSMR18, DRG^{+}16, DMB^{+}18, EP19, EYN^{+}15, FC19a, FMB^{+}18, FHN^{+}18, Gil11a, GNN^{+}17, GCK^{+}11, HTT10, Han12, HWO^{+}18, HH10, HBS^{+}19, Hol12, HM18, IHS13, JTK18, KSWS18, KVH14, Knu17, KC15, KVMM10, LDF^{+}18, LB13, LH16a, Lil16, Mac14, MRC17, MMPD^{+}14, MAJM^{+}18, MB18b, MAAB15, NIFA17, NNPV17, NNP^{+}14, OIS^{+}18, PCS14, Par10, PRT10, PC13b, Por10, PLP14, RB15, RBWL11, RGSF19, SOBS18, Sat18, Sev14, Sha10, SPRH19, SCS14, TÁSLC19, VSRD16, VBSD17, WTD^{+}15, WFK^{+}12a, WBSW13, YBRBB16]. Package [GBBB15, ITNL^{+}14]. packages [DDP16]. PACT [CS19]. Padu [Cou11]. Pages [Ano13g, Ano13j, Ano13i, Ano13h, Ano14o, Ano14n, Ano14i, Ano14m, Ano14k, Ano14j, Ano14l, Ano15p, Ano15m, Ano15u, Ano15q, Ano15w, Ano15n, Ano15r, Ano15s, Ano15v, Ano15t, Ano15x, Ano15o, Ano16n, Ano16r, Ano16x, Ano16o, Ano16q, Ano16t, Ano16s, Ano16m, Ano16p, Ano16u, Ano16v, Ano16w, Ano17n, Ano17p, Ano17r, Ano17o, Ano17t, Ano17v, Ano17q, Ano17w, Ano17m, Ano17u, Ano17s, Ano18p, Ano18q, Ano18u, Ano18v, Ano18n, Ano18w, Ano18m, Ano18s, Ano18o, Ano18t, Ano18r, Ano18x, Ano19w, Ano19u, Ano19v, Ano19s, Ano19q, Ano19t, Ano19r, BB19b, Ano10g, Ano10i, Ano10j, Ano10h, Ano10k, Ano11g, Ano11j, Ano11h, Ano11i, Ano12g, Ano12k, Ano12i, Ano12j, Ano12h, Ano12l]. Palau [WCMHO18]. Palawan [FD11]. Palos [SGNM19]. Palos-Islas [SGNM19]. Pamlico [BHG11]. Panama [PARD17, SJW^{+}16, SSM15]. panarchy [Jac15]. Pangasius [LBB^{+}12]. Panulirus [BRRG^{+}18]. paper [BR14, MLF14, SMD14]. Papua [BSM^{+}14, CBF^{+}12, Ben12, Bro15, MDE14]. Paracentrotus [ONF15]. paradigm [RCD^{+}18, Wit15b, Wit15a, dCW15]. Paradigms [BAB^{+}14, Zag18]. paradox [Bac18, KPR^{+}18, PC19]. Paralabrax [BSSS17]. parameters [NWW^{+}15]. parasitic [Mis19]. Paraty [HIBG15]. Pardo [vD16]. Park [CKY^{+}13, Cli13, CvPH^{+}18, HGS13, SHM12, AFGRF^{+}17, BL18, DSM^{+}13, FAPP17, HF16, HR17, KC13b, LBN15, LSTD12, MHL13, NKCR13, Thu10, vdGMA^{+}13]. parks [CMW10, JCD17, MLF14, PDT^{+}14, PRS^{+}19, SMKBM16, ZLL19]. Parliament [Gre15]. Parque [SFCI14]. Parsons [CMC^{+}13]. part [BAMAH17, KCKA^{+}18, PHM16, VGBG15, HUHA18a, HUHA18b, SDF^{+}14b, SDF^{+}14a]. Participant [MRG15, MLR^{+}16]. Participation [QSICTdV19, SMKBM16, WS17, Ben17a, BH14b, Car13b, Che10, CRO12, CHDJ^{+}19, Cou11, DMKL18, De 16, Dix16, EBR^{+}18, FHL18, GGADSR17, GBF^{+}10, HK18, JBKB15, KCCO18, LM16, LLA^{+}19, LHB^{+}17, MCCC16, Nak16, PDMÖ19, PPT10, SGNM19, SSFL10, Tam15, Tyn18, VGG12, vPLB^{+}13]. Participative [EG15]. Participatory [CdL17, OP15b, SGP^{+}17, ASP16, FMB^{+}13, IOG18, LHJ^{+}15, LSP^{+}17, LP16, LGDB17, LLO17, LRS^{+}13, QSICTdV19, RBW18, RUD^{+}12, Sch10a, VvD10, KMC^{+}17]. parties [Tos11]. partnership [BY19, Cor14, ELAC19, LJL^{+}14]. partnerships [FDGM16, KEG12, SG15b]. party [CMFG15]. Passenger [BSSS17]. passive [MMS^{+}14]. Past [MSC18, CSK^{+}18, DMMG14, LDVM10, RCMBG^{+}19, SNA^{+}17, Won19, HO10b, NFHJ16, VSS^{+}15]. Past-present [MSC18]. Patagonia [CCC17]. Patagonian [GRLJ15]. patent [ZZZL18]. Path [Jac19, VHA^{+}19, Ben18, LMP^{+}16, Seu16]. pathway [SvdBH^{+}18]. Pathways [CRCD13, BYK^{+}15b, DCF^{+}19, GSC^{+}16, Hou14, Soo17, SSSO17]. Patos [KNA10]. patron [MFRvP16]. patron-client [MFRvP16]. Pattern [LPP17, BGG19, RWJ18]. Patterns [GJA16, PDMÖ19, DMC^{+}17, FSM^{+}12, GBBB15, KHD12, OBNH15, Pin17, SCD^{+}17, SP19a, VCdR^{+}18, MPMK16, ST17]. Paul [dOSL18]. Paulo [GLTB^{+}15]. Pauly [Bra12, YBB^{+}17]. pay [BBE^{+}16, CL18, DV17, HSW17, LHN18, SCOS17, TCCV16, ZMZ18, vOHO^{+}17]. payao [MBvZ15]. Paying [LV18]. payments [PMA^{+}17]. pays [CSO17, ZZ15b]. PCB [SSJ17]. Peace [MHL13, STG11]. peaceful [Sco18]. Pearl [SSM15, WHP^{+}16]. peek [dGNC18]. Pelagic [FPG^{+}17, TJJAM19, BSD^{+}17, BBM^{+}17, BSW19, BHT^{+}18, CRO12, Dru10, FSCJM19, HSRH16, JA10, LBD^{+}18, MC14, Pan14, SG15a, SGMG18, SGZI19, THF^{+}15, vS15, CRO12, OR19]. pelagics [RRC^{+}17]. Pelagos [CDR^{+}16]. Pemba [WWT^{+}16]. penaeid [DBD19, SBMD13]. penaeid-trawl [DBD19, SBMD13]. penalties [MPMK16]. Penglai [PQLH15]. penguins’ [TRCR^{+}19]. penicillatus [BRRG^{+}18]. Peninsula [EAA^{+}18, LLS^{+}15, PR13, RNN^{+}18]. Pentland [JKS16, San18]. people [ARSHH16, BD14, CBG^{+}17, JRS^{+}16, KH12, LGH^{+}12a, LGH^{+}12b, OPMB12, PC13a, Zha16, LRJ^{+}19]. peoples [Cap10, BF18]. perceive [KGAB14, VPP^{+}16]. perceived [Edd14, GGVRVL16, MCH18, TPS15, WWH^{+}17, CFC^{+}18, FLWG11, WKB15]. perception [BOF^{+}15, BBAP^{+}17, Che12b, DHFL^{+}16, HR17, HMH18, KRM14, KWT17, KH15, LSAL19, PC13b, SBP^{+}19, SL15a, TÁB^{+}10]. Perceptions [BMGH11, Dav12, EMP^{+}14, HAA15, MSBB14, MD16, MMP18, MSG18, PM14, WB18, ADGGVL12, BD14, BR18, BM15, BY19, BSM^{+}14, CSvH^{+}18, CBL18, CJB15, CSMB^{+}12, DB19a, DS17b, FHJ16, FJG14b, FMKT15, GSJ18, GLTB^{+}15, GJA16, HGS13, HAC19, HB14, JBdL^{+}14, KR14, LAP^{+}12, LRMM17, LEG19, MMJM16, MdMGV^{+}18, MPCB14, MRG15, MLR^{+}16, Mis19, MD15, NBPF^{+}12, PRLCL12, PPT10, PTP13, RPG^{+}14, ROD15, RGGFA17, RNS19, RLHM14, SDRML15, SPB19, SNP^{+}18, SPP18, SF19, TKM^{+}18, TR18, TMHJ12, VPL^{+}13, VAW14, VvDM10, dAdOS17, BYK^{+}15b, VSCH16, Yat14]. Performance [MMdJ^{+}14, VRA^{+}12, ACSK^{+}10, ÁFFSCF17, AvTT17, BBG^{+}16, CGDB13, CST10, CKdS14, Del17, Fla13, GVSZ19, GFH^{+}17, GS16, HvPN^{+}19, IT15, KP17, LAP^{+}12, MSdAM16, MPPSdVSOHR14, OCG14, PGC19, PMN^{+}15, PKPdP10, PHC14, PHS18, RAG18, SKS^{+}19, ZLL19, CSK^{+}18]. Performing [GDG^{+}18]. peri [CSP^{+}14]. peri-urban [CSP^{+}14]. perils [FLH11]. period [MÁSSI13, MPMGGAHR18]. periodically [CF13]. peripatetic [RGBT16]. Periphylla [TMR^{+}14]. Permanent [PMC^{+}15]. permit [AK15, CL15, KNB^{+}18, Kna11]. Permitted [NS18]. perpetual [PVK18]. Persistent [SSJ17]. personal [MJRdC16, Næv17]. perspective [AZAMG13, BSD^{+}17, CHBRJ19, ETEW15, Gat18, HK10, HBL14, Her18a, Hua14, KS18b, KFB19, LK15, MdMGV^{+}18, MB18a, MO14, MI15, OO19, OS12, PGB19, PHAC15, RAF14, RGSF19, STJ^{+}12, SDLT15, SSD^{+}16, Vau17, VRB14, WML17, WRP^{+}10, Wos19, Wri16, ZZZL18, vAS19, vLRvH^{+}14, WCS18]. Perspectives [PRPLBLC12, TLH18, AWJ19, AOAM13, BN19, CHKG16, Che12a, CT16, CJHR14, CHvKM15, GBW16, GZW^{+}19b, How18, MB15, MMC^{+}14a, O’R15, ODRRM12, PGB^{+}17, San15, SWM11, SFCI14, TGdCR11, Wes18, ZFG12, ZEP^{+}19, CJ16, Lan18]. Peru [MJQMVÁ17]. Peruvian [CdlPS^{+}14, KSC^{+}19, Sch12]. pervading [JSG^{+}18]. Peter [dOSL18, She14]. petroleum [BHA^{+}14, Cha17a, HBA^{+}14, NS18, Tec19, Tec19]. Phalanx’ [WC16]. phase [HD10b]. phases [VC15]. phenomenon [BDvTvKT17]. philanthropy [Han19, MSTB19]. Philippine [MMC^{+}14a]. Philippines [ACA^{+}13, BPYC19, CG18, Cap10, CD15, DP17, DPB17, FD11, Fab16a, FGPP18, Gor19, HHG10, LBP^{+}19, MBvZ15, MKB17, MMC^{+}14b, PGPS10, PFP17, PC18b, RNN^{+}18, SPGP12, TAGP15, YP17]. Phoenix [Han19, MSTB19]. phronetic [LJ14]. phycotoxin [OCG15]. Physical [MKU19, PWH^{+}16]. phytoplankton [TGS18]. piece [VGCF18]. Pilot [FÓ12, Dva19, LPM^{+}14]. Pink [PGL^{+}16]. Pinkerton [Tur10]. Pioneering [DMC^{+}17, OMF14]. pipeline [Kai17a, Kai17b, Kai17c, LE12]. piracy [JSLL19, Lew16, Maz17, SFDB18, VGA15]. Piran [MHL13]. pirate [AT14]. pitfalls [GSA10, DCF^{+}19]. place [ABBD17, MD15, SPP18, UA13]. place-based [ABBD17]. placeful [GGD16]. places [CPRP19, TLR^{+}10, WZF^{+}18]. plaice [TM13]. plain [PL15]. Plan [DW13, CB13, CL15, DJTK17, Kim15, Kno10, KBC^{+}13, LJL^{+}14, MGGS18, OMF14, OOvdM^{+}11, QSICTdV19, RAZ16, SM17a, SG12, DBE^{+}16, Elo10, San19, TMKK^{+}15]. planet [Laf14]. planetary [VGCF18]. Plankton [BJGMG18, MGJB^{+}17]. planners [Gle14]. Planning [BFGR14a, HDG^{+}12, KJS14, Mil14, TGdCR11, AdSC11, BdKvT19, BBN^{+}13, BMD^{+}14, BP13, BW15, BDF^{+}10, BHS19, BBP^{+}14, Bre14, BFGR14b, BN19, BMGF^{+}17, CNJ^{+}10, CBN^{+}12, CMP^{+}15, CSVHS14, Car13a, CL16a, CSBE14, CGM^{+}18, Cra19, DTJ10, DMMG15, DMKL18, FCdS^{+}15, FÓ12, FJ12, FSMG10, FMB^{+}13, FHM^{+}10, GSR19, GSG^{+}15, GdV16, GFC^{+}12, GHP^{+}15, GAW^{+}13, GNW19, HGG^{+}18, Hav10, Hob14, HCJ15, HD10b, JMJ^{+}18, JBKB15, JAO^{+}16, JJ19, JLQ16, KVA18, KMC^{+}17, KEG12, KGST14, Kep14, KRK13, KMR^{+}13, LCH^{+}13, LB13, LWvdBL16, LWJH11, LLX^{+}15, MPF12, MGH14, MRG15, MRSH13, NVK^{+}17, NKAP13, NHPD19, NM18, Pap16, PWH^{+}16, PGB^{+}17, PGB19, Por15, PV16, QJ13, QSICTdV19, RYD18, RRA^{+}10, RFG^{+}14a, RHW^{+}19, RRMSSE16, SC13a, SDF^{+}14b, SDF^{+}14a, SOCA15, SFG15, SSB^{+}18, SS19a, SPO^{+}18, She15, SS15a]. planning [SCK15, SJ17, Smi18a, SM18c, SBC^{+}16, SLS^{+}13, SSG12, SMKBM16, TvN13, TK14, TRCR^{+}19, TKAO^{+}19, TEVN19, TE16, VGG12, WS15a, WM18b, WK19, Wes18, WBGJ18, ZMM^{+}14, dGNC18, dVM12, ADL^{+}18, CDR^{+}16, De 11, DTMJH16, DTM18, FOO^{+}15, FHL18, HDH^{+}15, JKS16, Kid13, MO14, vH19]. plans [BG15, DTM18, HK10, HAH^{+}16, LEM^{+}18, MJS16, RFKEJ^{+}15, Vin13, VBSD17]. plants [HPP^{+}17a]. plastic [AGWMP19, CSF^{+}18, Gat18, JHB^{+}18, Men18, PBW16, PBW17, RM17, RM18, SHL^{+}18, SJ19b, SJ19c, VGCF18, WML^{+}16, WMWH18, VS18]. Plastics [Pen18, JSG^{+}18]. plate [TSK18, Car14]. platform [CRB^{+}19, KYS10, KHT17]. platform-based [KYS10]. platforms [ODF^{+}16, vLvT10]. play [De 10, Lea19]. playing [SJ17, dGNC18]. Plenty [CKBS14]. plight [STM16]. pluralism [LT17, RGSF19]. PNA [TZBBL18]. poachers [BRR18a]. poaching [BRR18a, BRR18b, BR18]. point [CLC^{+}16, NIFA17, SPRH19, LBN15]. points [CMB^{+}15, SNDB^{+}16]. Pol [GDD^{+}19]. Poland [PW16]. Polar [CMC^{+}13, Wen11, Bro11, Mee11, PC11, PC13a]. pole [Gil11a, KGMP18]. pole-and-line [Gil11a, KGMP18]. policies [ARDP17, Alf17, CST10, CSCBR10, DRPSA17, GRYA^{+}15, GSB^{+}17, GRMAC18, GW10, Had18, Har19, Jen15, Lan15, LM11, MLF^{+}17, MRC17, Mal13, MÁSSI13, MLR^{+}16, MAJM^{+}18, Mes12, MNA15, Muk15, Pin15, Pin17, RMRN^{+}16a, RST17, SSCC19, San19, SSM15, SY18, TII^{+}19, WSG18, WM18b, WMWH18, ZZZL18, Zag18]. policing [Rog15]. Policy [BMPV10, BB19b, FD11, GSB^{+}17, Hey18, HD11a, HS19, JS13, KS18a, LLS^{+}15, LGH^{+}12a, McL15, Pas14, PDC19, RM16, RLM12, Thi13a, TMKK^{+}15, VLKN17, Wit15b, WMVR18, AAvH12, AAC^{+}13, BFO^{+}16, BP13, BSL19, BKHA^{+}18, BSW16, BPMY18, Boh19, Bor18, BFGR14a, BFGR14b, BAHCG14, BAY^{+}19, CCR16, CETQ16, CGSP13, CRCD13, Cha15, CKN19, CLLL17, CCD^{+}13, CHSW17, CEB^{+}12, CF17, CGC14, CJB15, CFC^{+}18, CR11, Cro11, DSK^{+}19, DA18, Dix16, DMW19, FSGB^{+}15, FAPP17, Fri10, GBW16, GRMAC18, GC17, GW15, GOTS19, GW11, HHL^{+}15, HDG^{+}12, HR16, HKHG19, HC19, HGG^{+}18, HP13, HFM^{+}13, He15, Hea13, HAC19, HL10, HTA^{+}15b, HRA^{+}15, JG13, JS18, JSY^{+}19, Kel18, Kim15, KPR^{+}17, KKSC19, LB17, Liu13, LLO16, MSBB14, Mac14, Mal16, MNM10, MV16, MSBR15, MLR^{+}16, MA18, MBB18]. policy [MGJB^{+}17, MLV^{+}11, Mee11, Men18, MRSH13, Mil14, MM14, Nei15, NMJS17, NACS17, OL11, OBH15, PC19, PO14, PL15, PNW18, PBW16, PBW17, Por15, PKR^{+}17, Pri10, QJ13, RvLvTH14, RCF^{+}12, RMRN^{+}16b, RMRN^{+}16a, RYD18, RFG^{+}13, RFG^{+}14a, RHW^{+}19, RSMM19, RLHM14, RFG^{+}14b, RBB^{+}16, RNN^{+}18, STM16, San10, San11, SSW^{+}18, SFC^{+}15a, SJP17, SRSN17, SM18a, Sol14, SMW13, Soo17, SSM15, SCS16, SCV^{+}17, SJ10, SH10, TBJC18, TBB^{+}10, Tos11, VSRD16, VSCH16, VRB14, VH16, VPP^{+}16, Vin13, dVMdC^{+}19, WWM^{+}19, WS15a, WCS18, WLT^{+}19b, Weg11, WMK15, Xin19, YLZ^{+}19, YBRBB16, ZP18, ZLSL17, ZZ17b, vLRvH^{+}14, vdPLMS16, BBG^{+}16, DCV12, EHR15, Foc11, Fri10, FH15, FQ13, GGADSR17, GAFD^{+}18, HG14, JSLL19, KFPR10, KMTS16, Mar10, MMBK17, MFAA16, MPMK16, NY13, PH16, PGP^{+}19]. Policy [RDSB10, SMD14, SSRS14, THA15, VSS^{+}15, dVM12, SJ19a]. policy-efficient [GW11]. policy-makers [GSB^{+}17]. policy-making [RMRN^{+}16a, SMW13]. policy-planning [Por15]. Political [EOC15, Tos11, CPR14, FBM18, GFA^{+}18, HBL14, Han19, Har14, JA10, LGH^{+}12a, LGH^{+}12b, Mal18, MSTB19, MSB15, MAHP16, Nol19, PWR17, She14]. Politics [PC13a, Pol12b, Bai12, Cam16, Don16, Hea13, LACG18, LBB^{+}12, PD15, RDBB17, SJ15, Weg15]. Pollicipes [CVP^{+}17, SFCI14]. Pollock [ZH18, CS13]. pollutant [Far19]. polluter [CSO17, ZZ15b]. Polluter-Pays [ZZ15b]. pollution [AVC18, AGWMP19, Bac18, CL13, DZLL15, GLSA10, Gat18, Gau10, GDD^{+}17, GDD^{+}19, KLN17, PBW16, PBW17, Rin18, SJ19b, SJ19c, SSJ17, TL16, Tos11, VGCF18, WML^{+}16, ZZ15b, ZZ17b]. polycentric [Gri17, MF19a, MF19b]. polymetallic [MBCR16, VVS18]. Polynesia [TFBS19, LACG18]. pool [KBG^{+}18, LR15]. pooled [PRS10]. pooling [MH13]. Poor [RCMBG^{+}19, CMFDCLV10, DPJ^{+}13, GOTS19, HA13, KVA18, NWW^{+}15, NWW^{+}17, NAB^{+}18, PVPF19, PFMA^{+}19, TTP12, TMDSF19]. Poorly [DTM18]. Poorly-designed [DTM18]. popular [CP18, GTL^{+}18]. population [BSM^{+}14, CN11, HIN^{+}17, PC19]. populations [BJS^{+}16, CSH17, CPBQ19, CLW^{+}12, KW16, Lon11, Lon12, WTA^{+}16]. Porcher [HS19]. porpoise [SKL16]. porpoises [ItDD^{+}18]. Port [GCW^{+}18, SPS14, CMFDCLV10, DR14, Fen17, KLN17, MNM10, MBBDS19, Ped13, WBA18, BDvTvKT17, BPYC19, CW15, CZX^{+}19]. Portfolio [BH14a]. portfolios [BKH16, SRK14]. Ports [Fen17, BDvTvKT17, CMFDCLV10, HAW19, KDG^{+}15, SNIN17]. Portugal [RPG^{+}14, RPdO^{+}19, Thi13a, AKV15, BW15, BBRT16, CVP^{+}17, FCdS^{+}15, GSDB14, HdF^{+}19, SGP^{+}17, SFCI14, SMQ^{+}16, Thi13b, VHA^{+}19]. Portuguese [BLR15, CNJ^{+}10, CB13, CLF^{+}19, LB17, MFF^{+}18, OCG14, OCG15, OCW^{+}17, STJ^{+}12, SDF^{+}14b, SOCA15, SMR^{+}15, SCR^{+}19, VPL^{+}13]. position [Dro13, LB18, Smi17]. positions [Xin19]. positive [MHC19]. possibilities [AM15, BB19a, NS18]. possible [Bor18, DDV^{+}13, HTD^{+}18, Han11, LZ16, PdSA^{+}13, vAS19]. possibly [PZ17a]. post [AT14, BG16, BCE^{+}17, Bre14, EHR15, HHL^{+}18, KR14, PW16, SS16, Yos17]. post-2013 [EHR15]. post-ban [SS16]. post-Cold [BG16]. post-hoc [BCE^{+}17]. post-normal [Bre14]. post-quake [Yos17]. post-smolt [HHL^{+}18]. post-transition [PW16]. post-trauma [AT14]. postgraduate [LA14]. postlarvae [ATAM10, AOAM13]. pot [SGP^{+}17, TPS15]. Potential [Bol17, BSP19, ECSE^{+}17, Lon11, ACR18, AMR14, AFGRF^{+}17, ATDP16, BRR15, CCTG^{+}13, CMSKP10, CP15, DCF^{+}19, DBG^{+}16, Dru10, DK16, FWK16, GFC^{+}15, GGL17, GAC^{+}19, Har10, HC18, HdF^{+}19, KS16, KS18a, KB15, LDS11, LDF^{+}18, PRG^{+}17, PDT^{+}14, PC19, PKA^{+}16, RAZ16, RFG^{+}14a, RR14, RHB17, SSC^{+}15, SSBL^{+}15, TSH^{+}19, VRMF11, VRA^{+}12, WMJF16, WS14, WMVR18, dGB10, dGNC18, vdBAMJT19, ABSB15, OIS^{+}18, WPdR^{+}14]. potentials [San18, WKB15]. Pound [HTF^{+}15]. poverty [DMW19, MSS19, SSRS14]. Power [HRC15, MBM14, Nol19, PHAC15, PMA^{+}17, De 10, EBR^{+}18, HDJD15, JBKB15, LDS11, PG19, TBB^{+}10, Weg11, WK19, ZZZL18, Don16]. power-play [De 10]. powerful [NS18]. powers [Dro13, dVM10a]. pp [Tec19]. PPGIS [SMKBM16]. practicable [BFGR14a, BFGR14b]. Practical [SLS^{+}13, Gla19, JWM^{+}16, NPAR19, TML^{+}19, GAB^{+}17, KSWS18]. practice [BMGF^{+}17, COG^{+}19, DSK^{+}19, DS10, FD11, FMB^{+}13, FAPP17, Gil11b, Gor19, HDG^{+}12, Har19, HSRH16, Iri18, Kay15, KEG12, MFF^{+}18, NSB18b, ODF^{+}16, PGP^{+}19, SM18c, WIC18, WMK15, LZ16]. practiced [MG16]. practices [DP17, DPB17, Gus18b, HHL^{+}15, KH11, KFM17, MBCR16, Pat15, PC13b, PSV^{+}14, SOS12, WF18]. pragmatic [BCM^{+}18, TTUR14]. Pragmatics [Knu17]. Pramod [HAK^{+}19]. Prawn [ATAM10, AOAM13, HCP^{+}17]. pre [WPdR^{+}14]. pre-conditions [WPdR^{+}14]. precaution [Kno11, MR16, De 10]. precautionary [BKH16, Hun13, Mat14, PH16, SCV^{+}15, SSLR14, KBGF12, dBMA13].

Precious [Cha15, HO10b, CYI13, Che12a]. preconceived [vdVD16]. predators [Hey17, Hey18]. predict [GZW^{+}19b]. predictability [SS19a]. predicted [GMP^{+}14, SSLR14]. Predicting [RLH18, THE^{+}18, TLA18, LDS11, MSG18, TLA16]. prediction [JSLL19]. predictions [DMB^{+}18]. preference [HR16, PRS10, SSB^{+}18, TTS14]. preferences [AAC^{+}13, BL18, BJO19, HLM^{+}13, JGG^{+}16, LL15, PDS15, PNW18, PM14, SCR^{+}19, SM18a]. preliminary [MA15, MCBM18, RR14, RLHM14, Vir18, Bla17, GA14]. premiums [SGLY13]. Preparedness [KO11, BvPS15]. Preparing [FC19b]. Preseason [SJ10]. presence [Kas16]. present [DMMG14, HO10b, MSC18, SWS12, TSH^{+}19, TSK18, VSS^{+}15, Won19, Gle14, PGB^{+}17]. presentist [ZP18]. preservation [MDK^{+}16]. Preserving [vD16]. Press [Rah19]. pressure [Cou11, SH14b, SMS^{+}18]. pressures [DSP^{+}18, ECR^{+}16]. Prestige [CMFG15, CSO17]. Pretty [Hil10]. prevailing [BJGMG18]. prevalence [San19]. prevent [BO12, DES10, GOO^{+}19, LCC14, MMB^{+}19, NIFA17]. Preventing [JSG^{+}18]. prevention [Bac18, JSLL19, MGGS18, TL16]. Price [NAYSN18, CGDB13, DL19b, GRLJ15, GM14, LV18, RW15, SGLY13]. prices [BB19d, PWN18, SKFL15]. pricing [Wit19]. primary [KS12d, WZCHS10]. primer [CCC^{+}13]. principle [SCV^{+}15, KBGF12, ZZ15b]. Principles [SSD^{+}16, BGC^{+}17, BH14a, FÓ12, FHM^{+}10, Hou14, LR15, LCS15, RSS14, RBJ^{+}11, Sch10a, SCJ13, TTM18b, TTFC19, VGBG15, Xin19]. prior [PV16]. priorities [BB16, CSvH^{+}18, CT16, CF17, CJHR14, EYN^{+}15, EGF^{+}17, HDG^{+}12, HPP^{+}17a, HBS^{+}19, ItDD^{+}18, JY19, KJP14, KFB19, PAD^{+}17, RV16a, RD13]. prioritisation [BRR^{+}12, NBF^{+}18, TTP12]. prioritising [GAB^{+}17, NBMB^{+}18]. prioritization [GBK^{+}15, LLH^{+}14, RL13]. prioritize [HVB^{+}17]. Prioritizing [CMH19]. Priority [RFG^{+}13, dlÁPAVS19, RBJ^{+}11]. PRISMA [OBH15]. Private [BBvZ^{+}17, HLC^{+}10, LU13, AdV18, EAE15, Har19, WvLvK12]. Privatization [KN17, Car15, CC17a, Por17, RCD^{+}18, THA15, WRP^{+}10]. privilege [PM14]. pro [MWR^{+}17, WS15b]. pro-environmental [MWR^{+}17, WS15b]. proactive [HHL^{+}15, PS11a, SSLR14]. problem [AOY16, All14, CBP^{+}16, DCV12, LHJ^{+}15, RM17, RGBT16, VDB16, VS18, WRP^{+}10]. problems [CSBE14, EGHC16, GMM15, JE11, Mee11, Ola18, Tei18, BDF^{+}10]. Procedural [GLJdlTC14, Cra19, DMJ^{+}17, Jae16]. procedure [CK18, LM16]. Process [Mor17, Bor10, CDA^{+}13, CPRP19, DMKL18, De 16, DLM^{+}18, ERB14, FBP^{+}18, JFT^{+}18, KC13a, KNB^{+}18, KHH13, Kno10, KC13b, KMR^{+}13, LQC10, LM18, LJ16, MPF12, OBJ^{+}12, PPT10, PSCY12, SSG12, TLR^{+}10, Vin13, dLS10, SVA^{+}12]. process-based [KC13b]. processes [DK15, FJA^{+}14, GLVLGG19, Pol12b, QSICTdV19, RD13, SOÖ12, Stø19, ŠST^{+}18, TTFC19, WF18]. processing [BBRT16, GLW17, PG19]. processor [HES16, Bod17]. procurement [RST17]. produce [CETQ16, Han14a]. produced [vOHO^{+}17]. producers [Fab16a, KLM^{+}13, KGAB14]. Producing [Fab16a, Ols11, PHLM19]. product [BHM16]. production [ATDP16, CP13, Cam16, Cha17b, DA11, Gar12, KEM^{+}18, KJH^{+}10, LJS^{+}18, OHH^{+}16, SBSK19, SNTL18, VMO14, WNLS17, vBvTM16]. productive [Men19]. Productivity [FGW15, SV15, WFSV15, BF19, FFKO15, MC15, PC18a, SBC^{+}14, SAdC15, TRM19, TWA^{+}15, PW15]. productivity-susceptibility [TRM19]. products [BKKN10, BTS16, DAGG16, MM15b, OM19, PML10, SNTL18, ZT19, dMAH^{+}18]. Proelss [AGSP^{+}13, FP13]. professional [MJRdC16, VBMM17]. profile [FHJ16, GCCSS^{+}19]. profit [GAKP^{+}19, HC18]. Profitability [BSD^{+}17, DFL15, TINL^{+}14, WK14, ZSH18]. profits [PMA^{+}17]. program [BSKH17, BAMAH17, GSS^{+}18, He18, LCC19, Pan14, TMHJ12, TLH18, GSS^{+}18]. programme [PCCR15, WLV15]. programmes [AL17, BW18b, MCY^{+}19, SHS^{+}14, SvHdB^{+}15]. programs [All14, BDSB16, BLR11, CL15, CKdS14, DFSB15, GFH^{+}17, HD10b, JS12, JS13, KTD^{+}15, KC15, Mat16, PFP17, QSMGVA18, Sen10, SSSO17]. Progreso [Ped13]. Progress [Coo11, KRK13, BvPS15, DFP15, EG15, GCCSS^{+}19, HBP16, SLGC^{+}18, STG11, SvHdB^{+}15, Tip16, Che12a, HP13, RTD^{+}17, RWG^{+}19, VBSD17]. prohibited [SS16]. project [CGS^{+}17, KC13a, LPC18, LLF^{+}16, MA15, Rei14, SS19a, FÓ12, HB19, MKX17, WvLvK12]. projections [APY^{+}17]. projects [BPMY18, CSK^{+}18, DJ15, DMJ^{+}17, ECR^{+}16, Gug19, MCY^{+}19, PFP17, ROD15, VLPC14, WSGM16, BBE19]. promise [FLH11, GSA10, KRK13, Pen18]. promises [GCCSS^{+}19]. promote [AMdlCMS^{+}18, BF15, HSRH16, PSCY12, SGN^{+}17]. Promoting [BHI^{+}17, KHT17, GDD11, Lin13, YCY^{+}15]. promotion [HGG^{+}18]. proof [LJ13]. propel [Men18]. properties [Har19, LKMN16]. Property [LH16b, WK19, FB17, HW10, KJH^{+}10, MCHO11, SC16, TCS^{+}18, TTM18a, WSB19]. property-based [KJH^{+}10]. proportional [SF16]. proposal [PRO16, SSL16, SRGGVL13, VRA^{+}12]. proposals [MMM19]. Proposed [BPYC19, PC11, Sol14]. propulsion [Man17]. pros [Cou11]. prose [SJ15]. prospect [MPPB10]. Prospects [Lyo14, RU12, CEB^{+}12, GA14, HC10, LACG18, LF17, Oan18, TKN19, YY19, BHI^{+}17, HBC^{+}13]. prosperity [PRS^{+}19]. protect [ABBD17, BWRT17, Chu10, DB19b, FJA^{+}14, HVB^{+}17, HD11a, HD11b, HAGB15, WS15b]. Protected [AAT12, CD15, CDR^{+}16, DFP^{+}16, FSS17, GR16, GLJdlTC14, Han19, HBP16, HCF^{+}16, LAP^{+}12, LV18, MSTB19, MSA^{+}18, PGL^{+}16, RAZ16, RFG^{+}14a, RMH^{+}15, RBS^{+}19, RS17, SM18b, TRCR^{+}19, Vau17, VGG12, WFK^{+}12b, ALSC13, AFR^{+}19, AdSC11, ÁFFSCF17, AVH10, AMR14, BF18, BMPV10, BMF^{+}16, BD14, BGS13, BBE^{+}16, Bur16, CMA15, CG19, CCC^{+}14, CBG^{+}17, CPFS^{+}13, CSMR18, CvPH^{+}18, DMKL18, DD13, DJM11, DDL^{+}17, EMP^{+}14, FP19, FSM^{+}12, GOHDB17, GPB^{+}18, GBF^{+}10, GWM^{+}10, GMP^{+}14, Gor19, GA19, GAW^{+}13, GOdFC16, GCG^{+}10, GCD^{+}11, HF16, HPP^{+}17a, HHG10, HCF^{+}17, Hun13, JQD13, JDQV13, JD16, JLCB16, KAF^{+}15, Kat16, KBF12, KBGF12, KFBG13, KRM14, KJWP14, KH15, KLMD15, LMH^{+}12, LF17, LJ16, LB18, LSC14, LEFC19, MVRG13, MMM19, MD16, MKR12, MMJM15, MMJM16, MWR^{+}17, MLF14, MSZS^{+}11]. protected [MCH18, MB15, MBTV14, MV19, MSH13, MC14, Ngo18, ND17, NWD18, NKBS17, OBJ^{+}12, PGR14, PTP13, PBJ^{+}14, RR17, RLKH19, RWK15, RAF14, RBWL11, RRRM^{+}15, RRRM^{+}16, Ros18, RL14, Sen10, SL15a, SWG18, dOSL18, SSG^{+}14, SSG12, ST13, SJT10, Tan17, TBKH13, TTM18b, TTP12, Thu10, TCCV16, THK^{+}18, TAGP15, WK17, WP12, WH18, WGB^{+}16, Wis14, YTTK10, dAdOS17, dGB10, dO13, CS19]. Protecting [Har19, Hey18, Joh19, GNW19, MPT13, RTD^{+}17, RBB^{+}16, SDC13, Hey17]. Protection [MVRG13, SC13a, TN15, AAT12, ACP^{+}14, Asa12, Bes10, BSW19, DPS16, Edd14, Gau10, GCCSS^{+}19, HLC^{+}10, HdF^{+}19, LU13, LZ16, MRHP19, MSG11, MSGM^{+}12, RWG13, RRRM^{+}16, RR18, SLGC^{+}18, SD13, San18, SGN^{+}17, SC13b, VAHG^{+}18, VMMB^{+}19, WH18, WSGID17, WAG^{+}15, vH19, Dro13]. protocol [YYN^{+}15, RM17]. Protocols [RB12]. prototype [SLS^{+}13]. provide [RM18]. Providencia [Mat16]. provider [EGD18]. provider-based [EGD18]. providers [Cus17]. provides [MOE11, RGWH18]. Providing [MMA16, MMJM15, RS17, NNP^{+}14]. Province [BGG19, DP17, HvD13, HWY^{+}18, KSS17, ZLL19, NJ18, SY18]. provincial [Ben12]. provision [BPBM17, BBvZ^{+}17]. Provisioning [AOA^{+}15]. provisions [DAGG16, LBD^{+}18]. proximate [FTF^{+}17]. proxy [MLR^{+}16]. Pterois [CAGH16]. Pteroplatytrygon [FSCJM19]. Public [AFR^{+}19, AAC^{+}13, Bes12, BBE^{+}16, CSvH^{+}18, FJG14b, GFC^{+}15, HSR^{+}18, JBdL^{+}14, LBVN19, Alf17, BAA^{+}15b, Ben17a, BBvZ^{+}17, De 16, DK15, EAG18, FBM18, HR16, He15, HLC^{+}10, JBKB15, JLY18, KSK^{+}19, KGS^{+}19, KCCO18, LM16, MCCC16, MC19, MSB15, NY13, OP15a, OHB^{+}18, RST17, RBW18, Smi18a, SPP18, SCC14, Tod12, VGG12, WM19, Wen16, WS15b, ZPT^{+}16, vOHO^{+}17, DHFL^{+}16, SMKBM16]. publication [Han19]. publications [RJPS11, Won19]. Publicised [CKFLO19]. publicly [Mul19]. Publisher [Ano11k, Ano17x, Ano19x]. Puck [PW16]. Puget [SLJ^{+}11, WLP10]. Punishment [KL14]. purchases [OM19]. Purpose [FAPP17, GSS^{+}19, YSJU^{+}17]. purse [DMMG15, Gil11a, LH16a, LBD^{+}18, MHM16, PSV^{+}14, SSW^{+}18, TZBBL18, WIC18]. purse-seine [PSV^{+}14, WIC18]. push [MJvP^{+}18]. Puttalam [HDK^{+}15]. Putting [Mee11, OCdS14, TM13, MFB^{+}14, PDB^{+}13]. puzzle [VGCF18]. Pyramids [JCBM10].

Q [MMP15, ZCD^{+}15]. Q-method [MMP15]. Q-methodology [ZCD^{+}15]. QMS [YFR10]. qoliqoli [SC16]. Quadrilateral [Yia18]. quake [Yos17]. qualitative [Bha12, GRLM15, KEM^{+}18, MSBB14, Mal16, RFL^{+}18]. quality [APD^{+}19, BJO19, Che11, Fla13, GGVRVL16, HWN11, MAR^{+}15, NF17, NBD15, PGR14, SFC^{+}15a, TML^{+}19]. quantification [MMS^{+}14, SRGGVL14]. quantified [HSH14]. quantify [DDV^{+}13]. Quantifying [MOH11, MPMK16, PAC^{+}19, RTM^{+}19b, KARR^{+}19, ZHH14]. Quantitative [QG17, GRLM15, MMGM16, Mal16]. quarantine [SCV^{+}15]. quasi [Han14a]. quasi-cooperation [Han14a]. Québec [Doy15, DSF16]. Queensland [JRS^{+}16, PDB^{+}13, TINL^{+}14, WM18b]. question [PS18]. Questionable [KBGF12]. Questionnaire [SNIN17, GPA18, GCK^{+}11]. Questionnaire-based [SNIN17]. questions [RFG^{+}13, RL13]. Quintana [VAW14]. quite [Zha16]. Quota [GKSJ11, HdVB15, KSC^{+}19, SC11, AMG16, Bod17, Bor10, BKH16, CdLR^{+}15, CKVO16, CBL18, CK18, Dav10, DRH^{+}18, EB16, EGGT12, FGW15, FB17, GS16, Her18b, HdV17, IT12, IT15, KBBR16, LCC19, McC17, MA15, NM12, OCG14, PGC19, PE10, RYJ^{+}17, SCS16, SM17b, TI13, TMHJ12, Tor16, WBH^{+}15, vHCT^{+}16, vPG10]. Quotas [YCA^{+}12, BBG^{+}16, CG13, DRH^{+}18, EGHC16, GDD11, HD10a, MNHN18, Mox12, MCBM18, Sol14, CSMB^{+}12, Tur10].

R [PDC19, KFM17]. R&D [GLSA10]. R. [Hee14]. R/V [KFM17]. R1 [Ano15x]. R1-R2 [Ano15x]. R2 [Ano15x]. race [JD16, LGD^{+}19]. RACs [HG14]. radar [Cha14]. raise [GPB^{+}18]. Raising [MAJM^{+}18, RD12]. Raja [ALM17, GAW^{+}13, VAP^{+}10]. Ranch [KSS17]. ranching [HBG12]. RAPFISH [AARK15]. Rapid [RR18, KS16, KS18a, LM11, SCMS^{+}18, SKS^{+}19, SMH17, WMVR18, vPJH^{+}17]. rapidly [BAY^{+}19, WB19]. rarely [WHB^{+}18]. rarely-encountered [WHB^{+}18]. Raritan [HMH18]. RASS [CLDP17]. rates [CA16, WDK^{+}13]. rather [MCMB^{+}18]. ratification [ZZ15a]. rating [PHS18]. Rational [KS10, JBPBA16]. rationale [CMSM^{+}16, VAHG^{+}18]. rationality [Rog15]. ratios [BKKN10, PML10]. ray [CMH19, CBP19]. rays [BW19, OBNH15]. Re [FB15, HFM^{+}13, CPFS^{+}13]. Re-defining [FB15]. Re-thinking [CPFS^{+}13]. re-use [HFM^{+}13]. reach [FR16, MDNT19]. reaction [FBP^{+}18, KWT17]. reactions [MPPB10]. reactive [PS11a, Tan17]. Ready [SCR^{+}19, JAG16, SBC^{+}15b]. Real [HHL^{+}18, EB16, MHL^{+}15, SLGC^{+}18, TLR^{+}10]. real-time [EB16, MHL^{+}15]. realigning [BFO^{+}16]. realisation [KP17]. Realising [BSH^{+}19, KMD15]. realities [JDQV13, Lal17, Par10]. reality [FAPP17, JLQ16, RCM18]. reality-check [RCM18]. reallocating [ECF^{+}18]. reallocation [CGM^{+}18, BAT^{+}18]. really [RPdO^{+}19]. Reaping [AOA^{+}15]. reason [AAL^{+}18, Bro11]. reasoning [SM18a, VvD10]. Reassessing [QG17]. reauthorizations [DSR12]. rebuild [BHAHW17, OSS14]. Rebuilding [CDA^{+}13, DRPSA17, FWC^{+}18, LM14, WDK^{+}13]. Rebuttal [AGSP^{+}13, HS19, FP13, PC13a]. received [Cha11]. recently [BSSS17]. Reception [WBA18]. recipe [LGH^{+}12a, LGH^{+}12b]. Reclaiming [Cap10]. recognition [DJ15, Hut19]. Recognizing [RB15]. Recommendations [DGC^{+}17, vdPLMS16, CKBS14, DBB^{+}14, OGM^{+}14, RM16]. reconciled [PS11b]. Reconciling [GAP12, PS19, SBP^{+}19, Rei19, ACM^{+}12, He15, TK11]. reconstructed [ZPT^{+}16]. Reconstructing [ZP16]. Reconstruction [CGP^{+}18]. Recontextualising [PS10]. record [KCCO18]. recorders [Cha16]. recording [BO12]. records [CSF^{+}18]. recover [BAT^{+}18]. recovering [BJS^{+}16, DR10, EKV12]. recovery [BAO^{+}12, CA16, COG^{+}19, CE17, CCS16, GC18, Gla19, KBC^{+}13, OL10, PF12b, Var11, Yos17, dlFRG17]. Recreation [CHL^{+}19, RRA^{+}10, RMH^{+}15, SP19a, VLKN17, NVK^{+}17]. Recreational [EWG^{+}18, FSS14, HHA17, NKHL18, VGL^{+}17, vPJH^{+}17, BvRH11, BSSS17, BR18, Bor10, Bro16c, BAY^{+}19, CACL18, CBL18, DSM^{+}13, DS17b, GPA18, GCK^{+}11, HR17, KSS17, KC13b, MJvP^{+}18, MMJM15, MMJM16, MPCB14, MCH19, MSG18, PTDH17, PAD^{+}17, PVA^{+}18, PRS10, RMR^{+}17, RFHEJK13, RTM^{+}19b, SPB19, SK18, VPL^{+}13, VC17, VBMM17]. recruitment [BSMS15, NIFA17]. rectification [CSO17]. recycling [ARDP17, HFM^{+}13]. Red [CSMB^{+}12, BRRG^{+}18, GM14, LLX^{+}15, O’R15, RVdCGN11, SAdC15, SH16]. REDD [AG16, ATDP16]. Redrawing [FHN^{+}18]. reduce [BF19, Fen17, GWWH17, JG13, RM18, SHL^{+}18, SCKP19, SVL^{+}16]. reduced [RNS19]. Reducing [KBG^{+}18, PNW18, Shi16, JWM^{+}16, RIK16, SVA^{+}12, SWG18, WIC18, WMWH18]. reduction [GDN19, JE11, KS19, MRG15, Muk15, Mul19, PBSG12, SJP17, SRGD18, VRMF11, ZLSL17]. reductionism [Gil14]. Reef [DSM^{+}13, DDS^{+}11, FAPP17, HMS^{+}17, LBN15, LSTD12, PRS10, Qiu13, TINL^{+}14, WS15a, AOA^{+}15, BPMY18, Bre13, BM15, FD11, FTF^{+}17, HLL^{+}18, KJWP14, LSP^{+}17, LGDB17, PTN^{+}17, PJS13, SBW^{+}17, TTS14, TCN^{+}13, TBB^{+}10, WCMHO18, WCM12, WBC14, DD13, FSB14]. reefs [AOA^{+}15, CG18, DSWG13, RR14, TC15a, TCCV16, Wea16, WGMG^{+}16, WS15b]. Reference [HGH^{+}18, CMB^{+}15, KS12b, KC13b, NIFA17, SPRH19, Wit15b, Wit15a]. Refining [MS14]. reflect [RMB^{+}17]. Reflections [Dro13, Bla17, Boh19, CK11, DMKL18, Kid13]. Reflective [KEG12]. reflects [LWvdBL16, SG12]. Refocusing [Cli13]. reform [BDvTvKT17, Bor18, CDO16, CMSM^{+}16, FB15, FQ13, HC19, HG14, Mar10, SMD14, SOÖ12, TBJC18, TBB^{+}10]. reformed [CRO12]. reforming [MCMB^{+}18]. reforms [DRGCPS17, MSS18, RGG^{+}16, SA18]. Reframing [NM19, Dra15, TH15]. refuge [CPRP19, EMP^{+}14, QZH^{+}18, WZF^{+}18]. Refusal [BN19]. regard [GLTB^{+}15, RFL^{+}18]. regarding [AAC^{+}13, CKC^{+}13, HAC19, LHH15]. Regency [VAP^{+}10]. regime [AvTT17, BW15, DZLL15, EB16, Gug19, HK11, JAG16, JGA17, KPF12, MCHO11, Men19, MB18b, ÖGB^{+}10, Sch10b, Sha10, TDMN19, LJ17]. regimes [BBE^{+}16, TvN13]. Region [DP17, Gus18a, PCS14, Boh18, BSAT18, CSvH^{+}18, CF17, HWO^{+}18, HBZH16, HBH^{+}19, KAF^{+}15, LSMW10, MB18b, PPR^{+}15, RYD18, RJ18b, RL19, Sev14, SSL16, SMW13, SC10, VBSD17, YSJU^{+}17, ZZ17c, CKdS14, DPB17, DFV15, FMM13, Grö18, KS18b, Smi18b]. Regional [Bac18, CRO12, HCJ15, LXWX17, MF19a, MF19b, RBM^{+}15, vTRvLvH14, AMdB12, BPF^{+}16, BRR^{+}12, BY19, CETQ16, Chu10, CAP15, CSP10, DHC11, DSF16, DDvP^{+}16, FSMG10, dlFFVRC16, GGSRVL17, HTT10, HGG^{+}18, HBS^{+}19, KVA18, KARR^{+}19, KS18b, KSS17, LMA^{+}15, Lil16, MSdAM16, MLF14, MLA16, MO12, PKOF18, PDMÖ19, PPCM16, RvLvTH14, RUH^{+}14, Seu16, SK18, SGMG18, Smi13, VPC19, VKGN^{+}14b, WBGJ18, WAG^{+}15, YPS13, dBMA13, vHHB14, RB12, dGNC18]. regional-level [FSMG10]. regional-scale [MSdAM16]. regionalisation [EHR15, SG12, vH15]. regionalised [vHK17]. regions [BK16, HA13, HF12, LZL19, MKX17, MML^{+}19, MDK^{+}16, RWJ18, YYYL15]. register [RMJ12]. registered [YSJU^{+}17]. registration [CLLL17]. Registry [PARD17]. regression [OCW^{+}17]. regulate [CO16, MKSS18, NS18]. regulated [BV19]. Regulating [Gri17, RIK16, Wri16]. Regulation [Rin18, AF14, Alf17, BHAHW17, CCC17, CHL^{+}19, EFK^{+}19, Gil14, HCF^{+}16, HCF^{+}17, JFA17, KH13, MM17a, MBM14, MYA^{+}18, Ola18, PS19, Ray19, RJ18b, Ros17, DAGG16, LBD^{+}18, McN18, Vat17]. regulation-based [HCF^{+}16, HCF^{+}17]. regulations [BSSS17, Eli14, HvD13, HEJK10, JBJ12, JA10, KS12c, KAF^{+}15, KV18, Lax10, LJT19, MG14, MFF^{+}18, SRA15, SJW^{+}16, SGMWE14, TMG15, THNR12, VPL^{+}13, VC17, WP17]. regulative [dlTCL10]. Regulatory [Boh19, Ren18, AMLK^{+}12, AF14, AA17, Boh18, BO12, CH16, DJS17, DSH14, Eng18, HR16, Kai15b, LT17, MC15, MC16, PKM^{+}19, RJ18a, Sun12, WL18, vLRvH^{+}14, Sch15b]. reinstatement [SBC^{+}14]. reinvigorated [BJS18]. reject [MCH19]. Rejoinder [WMVR18, CMC^{+}13, Han14b, KS18a, Sch15a, Tur10, AGWMP19, PZ17a]. related [BO12, EZPRZ18, HBA^{+}14, HBK17, HM18, MSdAM16, Mis19, RFHEJK13, SCV^{+}15, Smi17, WS15b, You16a]. relation [CMFDCLV10, EYN^{+}15, LBD^{+}18, MPPSdVSOHR14, MYA^{+}18, REA^{+}14, WBC14]. relational [NM19, GC18]. Relations [MJRdC16, Fri10, JAO^{+}16, LH16b, MFRvP16, PWR17, Sco18, dVM10b]. relationship [CHBRJ19, MC16, RST17, TO10, YTSO15]. Relationships [ERB14, dVvT11]. relative [HdVB15, YWC^{+}16]. release [PSV^{+}14]. Relevance [RSS14, BAA10, JKS13]. relevant [BVB18, GAB^{+}17, JDM^{+}19]. Reliable [OIS^{+}18]. relocation [JOE^{+}19]. Remote [PHEM^{+}17, ALSC13, BBE^{+}16, GPB^{+}18, JD16, KHP14, OIS^{+}18, dOSL18]. removal [KNB^{+}18, Lon11]. removals [WZCHS10]. Removing [Cho11, CL18]. remuneration [PGP18]. Renewable [MMOL15, AWH13, CSVHS14, DS10, Gue18, KCJW15, KLN17, KLMD15, LPC18, LGM16, NNP^{+}14, ROD15, SAGG13, SV15, Tod12, WMJF16, Wri15, YLZ^{+}19]. renewables [JKS13, MC14]. rent [Fla10, GGL17, SW10]. reorganisation [Thi13a, Thi13b]. Repair [SBC^{+}14]. Replacing [Gil11a]. replenishment [RL14]. Reply [GZW^{+}19b, Lon12, NGS16]. Report [Rei10, CGF14, Cha19, CKC^{+}13, FCdS^{+}15, MMB^{+}19, TDI^{+}14, WC16, ZQZ19]. reported [EZBL19, FC19a]. Reporting [Bro16b, Bro15, Bro18, DDS^{+}11, ENWL19, FBM18, LBN15, MBCR16].

Repositioning [OR19]. represent [RPdO^{+}19, SCJ13]. representation [APD^{+}19, BY19, Hol12, KVA18, LGFB^{+}19]. representations [LBN15]. reproductive [UEMRG^{+}18]. Republic [WL18]. request [QZH^{+}18]. require [Lon12, Pas14]. required [EGGT12, GH15, PGR14]. requirement [AGB^{+}18, OSS14]. requirements [BO12, CA16, Gil11a, HMK19, KP17, LOD^{+}15, RMRN^{+}16b, WHB^{+}18, ZPB^{+}13]. requires [HvPN^{+}19, WMSB15]. requiring [MMBK17]. rescue [Cus17, FC19b, HFM^{+}13, KPF12]. Research [HCBM15, PAD^{+}17, RWJ18, BJS18, BAY^{+}19, Cro11, CJHR14, CvPH^{+}18, DK15, EBR^{+}18, EYN^{+}15, GPO^{+}10, Har10, ItDD^{+}18, JY19, JG13, JL11, KRK13, KFM17, KHvH^{+}14, LLZ^{+}19, MWGD11, MMGM16, MF19c, McL15, Men18, MNA15, Par15, PSV^{+}14, Pom16, RFG^{+}13, RV16a, RJPS11, RL13, SOBS18, SSS^{+}17, SSH^{+}19, SH10, TH17, VHRM17, WWW10, Wen16, Wen17, WKB15, WOdG^{+}16, vHFJ^{+}19, EG15, Wen16]. Reserve [NACS17, Qiu13, RR18, VAW14, BS10, MHF13, MKMS12, NKHL18, NGS16, RNCGB15, RNEBG19, SB15, THNR12, BHT^{+}18, BRRG^{+}18, CSF^{+}13, CC13, Jon13, MSS19]. reserves [CF13, FCC^{+}12, NKHL18, OLF19, THNR12, VSA^{+}12, FVM14]. reshaping [Bla15]. residence [AFR^{+}19]. resident [SOUN16]. residents’ [HPPH16]. resignation [RDBB17]. Resilience [BdKvT19, PC19, PA10, BBR^{+}18, BDB19, COG^{+}19, CFC^{+}18, FLWG14, GGK17, GFA^{+}18, JQD13, LBP^{+}19, NHPD19, SSD^{+}16, TMR^{+}14, VM18]. resilience-based [COG^{+}19]. resiliency [JWBJ13]. resilient [FMD^{+}13, LRJ^{+}19, MNM10]. resistance [EGHC17a, Pin17, WS17]. resolution [APY^{+}17, DZ12, EGD18, Mso16]. resolve [MHL13, WSWM17]. Resolving [SBMD13, EGHC16]. Resource [Tec19, AAL17, Bes10, CdLH^{+}18, CS19, CSCBR10, CSS15, CA17, CHvKM15, DTM18, DSH^{+}19, EAE15, ELAC19, FKMV16, FSGB^{+}15, GGL17, HL15a, KBB^{+}15, KED^{+}18, LSP^{+}17, LR15, Mat16, MSB^{+}18, MTP^{+}17, NM19, NHA15, NW10, Nor15, PD17, PKA^{+}16, PT10, SCR^{+}19, SW10, SV15, ST13, SFCI14, SGMWE14, TDI^{+}14, TKAO^{+}19, VvD10]. resource-based [PT10]. resource-dependent [CA17]. Resources [ZH18, AZAMG13, BvRH11, BAA^{+}18, BKHA^{+}18, BAHCG14, CHDJ^{+}19, DSP^{+}18, GNW19, Hof19, HO12, KS10, KJH^{+}10, Lea19, LC19, NBF^{+}18, PKOF18, PKA^{+}16, Sha17, SI18, SOS12, Vir18, Weg15, vHFJ^{+}19, Bro16b, Tec19]. respect [Dro13, Kay15]. responding [McL15, QZH^{+}18, SSJ17]. Response [KS18a, PDC19, She14, BR14, Che12b, CLW^{+}12, DBG^{+}16, DSM^{+}13, DDL^{+}17, FMB^{+}18, FP13, GSS^{+}18, GNW19, HCF^{+}17, JLY18, LCC14, MSB15, MH13, MGGS18, SSW^{+}18, SJ19c, SC13b, SS13, WFK^{+}12b, CLLL17, HS19]. Responses [LSTD12, PS11a, AA17, BLC16, DCY^{+}19, FBM18, GRMAC18, KNY14, MV16, MSA^{+}18, SSM15, SJT10, WC17]. Responsibility [Zag18, NHA15, Sch17, WLX19]. Responsible [HFF11, KSC^{+}15, LH16b, PDC10, Ant16, Sat18]. responsibly [CLDP17]. responsive [DSWG13, SMR^{+}15]. responsiveness [VDB16]. restitution [Gla19]. Restoration [BPMY18, HB19, DDM^{+}19, FMP^{+}18, GSS^{+}19, HMH18, LU13, OCCvT^{+}18, SSG^{+}14, VAP^{+}14, VMMB^{+}19, WSA^{+}15]. Restoring [WF18, ESGC13]. restricted [PCK^{+}19, RRRM^{+}16, SMBL13]. Restriction [SHM12, CSH17]. restrictions [SC13b, VPL^{+}13]. restrictive [LGC16]. restructuring [FGW15, PRO16]. resultant [DSP^{+}18]. Results [NHA15, DRG^{+}16, EGD18, HFF11, SBC^{+}15b, PPCM16, TKM^{+}18]. results-based [SBC^{+}15b]. Resurrecting [KCCO18]. retail [BB19d]. retailer [Alf17]. retention [CCS14]. Rethinking [BBM^{+}17, LA14, WMK15]. Retracing [Won19]. RETRACTED [Hey17, PPM17, TLA16, Wit15a]. Retraction [Hey18, PPM19, TLA18, Wit15b]. return [BG16, OO19]. returns [Par10, PCLE17]. Reunion [TGT^{+}19]. reusable [SSV18]. reveal [WRW^{+}17]. revealed [Mul19, PRS10]. Revealing [IOG18]. revenue [KWT17, OCG15]. revenues [OHB^{+}18, RAG18, SRK14, Wit19]. reverse [CHL^{+}19]. Review [KGST14, LP16, MYA^{+}18, Rah19, Sal16, Syb15, Tec19, VC17, VH16, AMR14, BR13b, BR14, Car11, CMMBV^{+}19, CP10, DHC11, DHSC14, DTMJH16, ESM^{+}19, EYN^{+}15, EZPRZ18, FSWFG13, FSGB^{+}15, FFH18b, GvdB13, GVSZ19, GA14, HTA^{+}15a, Kai17c, KPF17, LM16, LOS11, MMGM16, NHPD19, NSB18b, NNP^{+}14, PDR11, PAA^{+}12, RBM^{+}15, SCK15, SLS^{+}13, WS14, WAG^{+}15, BGA^{+}15, ZZ16]. reviewed [WLV15]. Reviewing [JFT^{+}18, HBP16]. reviews [OBH15]. revision [Árn17]. revisions [HTH13]. revisited [BPBM17, FW15, Gat18, SJ14, Pom16]. Revisiting [FMD^{+}12, LVRM^{+}18]. revitalisation [SBC^{+}14]. rewards [JGG^{+}16]. rezoning [vdGMA^{+}13]. RFMO [Pol12b]. RFMOs [AMdB12, Hut19, LM18, PKOF18, dBMA13]. rhetoric [Ant16]. Rhincodon [BFTS^{+}16]. Rhincodontidae [BFTS^{+}16]. Rhode [DTJ10, OMF14, DJTK17, HDJD15, PTDH17]. Ria [GNIL14]. Rica [JSG^{+}18, HGS13, LH16b, MBSL13, OPGL16]. rig [KS13, MMN18]. right [Asa12, Bre14, BWRT17, DB19b, Goo17, HBL14, HD11a, HD11b, MWR^{+}17, MPT13, NPAR19, PR13, SALS17, ZVP^{+}13, ZSZPR19]. Rights [ABCDMG17, Bro16a, Dup14, Han11, Hav13, KCJW15, QSHI17, Ant16, ABB14, BF18, BMW17, Bod17, CMVLGG14, Cap10, Che12c, DHR16, FB17, HLC^{+}10, HW10, IW19, JRP17, MCHO11, MMA16, NKBS17, RD13, SC16, Son15, SS19b, TTM18a, Tod12, THK^{+}18, WSB19, YCA^{+}12, vdPLMS16]. Rights-based [Bro16a, Dup14, Hav13, Che12c, HW10, MMA16, YCA^{+}12]. Rio [FQ13, RAG18]. Rise [EP19, BLR11, DHR16, FvRA11, Mal18, SCK15, SNTL18]. Rising [Kid13, LHN18]. Risk [MR16, SWP13, TÁB^{+}10, BRR^{+}12, BRR15, CPBQ19, DHC11, Dav12, DN12, DDV^{+}13, FHJ16, GCP19, GOK14, GCM11, HIN^{+}17, HBA^{+}14, KBG^{+}18, KCKA^{+}18, KH11, LBN15, MSA^{+}18, NMvP19, NBF^{+}18, NBPF^{+}12, RBW18, RG19, RHB17, SCV^{+}15, SCKP19, SS15a, Tho13, WS14, CLDP17, HMH18, SLDHS18, WFK^{+}12a]. risk-based [KCKA^{+}18]. risks [HDH^{+}15, JSY^{+}19, LGFB^{+}19, NKAP13, SLDHS18, THM15, dVM17, HTD^{+}18]. rituals [Deb18]. River [Jen15, Kru16, WHP^{+}16, Wen11, MPCB14]. Road [YC18, QSHI17]. roadmap [FÓ12, KAPF^{+}19]. roads [ECR^{+}16, GJ15]. robust [DMJ^{+}17, KC13a]. robustness [EBS18]. Rock [RG19, RYJ^{+}17, CdLR^{+}15, CdLH^{+}18, Lal17, LSMW10, NBPF^{+}12, RCdLS13, vPG10]. Rockfish [MKMS12, WLP10]. rocklobster [vPLB^{+}13]. role [AL17, BJGMG18, BDP^{+}17, BBP^{+}14, BRW^{+}13, CRO12, CKHJ^{+}17, DJR^{+}18, EGHC16, Fen17, GSR19, GSJ18, GWWH17, GDD11, GCD^{+}11, HIBG15, Hou14, HBH^{+}19, LY14, LBL^{+}17, LWJH11, MPF12, MF19b, MNLHRMP17, MSB15, MFRvP16, MO13, Nik19, OCCvT^{+}18, Oya18, Pin15, PGC16, PGP^{+}19, RB15, Ray19, RCUR13, RHB17, SCD^{+}17, Sch10b, SRK14, Sha17, SAM13, SVA^{+}12, Smi18a, SB18, TK13, TRC^{+}17, VSCH16, WW19, WWH^{+}17, WBA18, WS15b, ZM15, dGNC18, IT15]. roles [Jet18, RDSB10]. Roo [VAW14]. Room [HTF^{+}15, Tur10, vH15]. Ropeless [MMB^{+}19, MMB^{+}19]. roses [JCBM10]. Rote [MBT^{+}19]. rotor [Sea17b]. roughy [FvRA11]. round [Mar10]. Route [LK15, MGB^{+}18, ZMZ16]. routes [APY^{+}17, RPG^{+}14, RPG^{+}15, RRCS^{+}19]. routine [AAL^{+}18, BHA^{+}14]. routines [BHA^{+}14]. routing [SDHS19]. rubbish [Bre14]. rubrum [TRBS13]. rug [PC11]. rule [PS19, RBWL11]. rulemaking [MLR^{+}16]. rules [BCM^{+}18, CLW^{+}12, De 17, ERD^{+}13, GBF^{+}10, He18, HDH^{+}15, JSD19, KL14, Lon11, Lon12, MBSL13, vH19, vHCT^{+}16]. ruling [Cla15]. run [RW15, SJ10]. run-size [SJ10]. rural [HD10a, Lan15, LCC19, MSS19]. Russia [BG16, LSK18, Nak16, Tyn18]. Russian [Tyn18, SS13]. Ryan [Rah19].

S [GCP19, KS18a, PDC19, Rah19, WMVR18, PFMA^{+}19]. S-fisheries [PFMA^{+}19]. S119 [WMVR18]. S121 [WMVR18]. Sabah [HBZH16, PF12b, TTJ18]. Sabang [KH15]. sablefish [PC13b]. SAF [GNIL14]. safe [Gam11, Har14, HDH^{+}15, She14]. Safeguarding [PVS^{+}19]. Safety [Næv17, NPS^{+}19, ZP16, dlFRG17, AL18, ERB14, HHL^{+}15, HUHA18a, HUHA18b, NS18, PNW18, PSH18, SCC10, TS14, TKM^{+}18]. Sagay [MSS19]. Saharan [TCC^{+}19]. saithe [SRA15]. Sakuraebi [LYB17]. Salmon [MR16, XDG^{+}16, BMGH11, Car19, CJ17, CKFLO19, ET19, FHJ16, HK10, HES16, HMK19, HLRO19, IH18, Kna11, KN17, Knu17, LFGM15, LOS11, LJT19, MRC17, MFT19, Mis19, MAHP16, Nak16, OP15b, RLH18, SJ10, THRS15, WEKH19, vOHO^{+}17]. salt [RBB^{+}16]. Saltmarshes [MBB18]. saltwater [TLH18]. Salvador [Cam15]. Same [SLDHS18, FSS14, RCM18, vHCT^{+}16]. Samoa [LR15]. sampling [KW16, SvHdB^{+}15, SSG^{+}14]. sanctuary [HD11a, HD11b, MCH19, WP17, CDR^{+}16, VGL^{+}17]. sandwiched [VBHC^{+}17]. Sandy [CCS16]. Sanriku [Yos17]. Santa [Mat16]. Santos [MKB17, MBvZ15, NGS16]. Sanya [Qiu13]. São [GLTB^{+}15]. Sarangani [DP17]. sardine [HH10, IHS13, PIH^{+}10, SCMS15]. sardines [RNN^{+}18]. Sardinian [MFP18]. Satellite [RBS^{+}19, Dru10, EGD18, Lyo14, OIS^{+}18]. satellite-based [EGD18]. satisfaction [MLR^{+}16]. Satoumi [MV17, THK^{+}18, BM13]. Saving [LHH15]. savings [Sea17b]. Savudrija [MHL13]. saxatilis [MSG18]. Scalar [Thi13a, Thi13b]. Scale [Lan18, SALS17, ARSHH16, AdSC11, ATB^{+}19, AYA^{+}19, BRR18b, BBE19, BCMH^{+}18, BAA^{+}18, Ben17b, BM15, CPR14, CMAA13, CN17, CBG^{+}17, CRB^{+}19, CBG14, CF13, CPBQ19, CPB19, CE13, CBS^{+}16, Cro15, DRGP19, Dav15, Deb18, DMdJ^{+}15, DYN^{+}19, DRH^{+}18, DK16, EA19, FEF^{+}12, FB15, FDGM16, GLVLGG19, GSJ18, GMD15, GHM^{+}18, GTL^{+}18, GA19, GCM11, GMM15, HWL^{+}19, HWO^{+}18, HIBG15, Har15, HDM13, HAC19, IW19, JMJ^{+}18, JTK18, JHN^{+}19, KL14, KLM^{+}13, Lan15, LDF^{+}18, LLN17, LCC^{+}18, LH16b, MB10, MGA^{+}10, MSdAM16, MJQMVÁ17, MMTC16, MN18, MMC^{+}14a, MMC^{+}14b, NRB^{+}18, NWD18, OC17, PRBX19, Par15, PD15, Pin17, PVPF19, PGPS10, Pom12, PC18b, PGP^{+}19, PRPS17, PKM^{+}19, PCLE17, RROR12, RBTR^{+}19, RV16b, RPC^{+}17, RTM19a, RCMBG^{+}19, RD13, SDRML15, SC19, SCD^{+}16, Sal16, SLS18]. scale [Sat18, SCC^{+}17, SGNM19, SSD17a, SL15a, SNA^{+}17, SC10, SSRS14, Sun12, TRM19, TTS14, TLA16, THE^{+}18, TLA18, THA15, VPR^{+}16, VPP^{+}16, WW17, WBC14, Wea16, Wes18, WZCHS10, WMK15, Wos19, YP17]. Scalefish [NBMB^{+}18]. scaleness’ [GS17]. scales [OBC^{+}18, ONF15, PVA^{+}18, RRRM^{+}16]. Scaling [KVA18]. scallop [FJ12, GLW17, OTOL17, Por17]. scallop-seagrass [OTOL17]. scandals [MV16]. scanner [SCOS17]. Scarcity [KPR^{+}18, PPML16]. scenario [LLF^{+}16, MP12, dS17a, DMMG15]. scenarios [ET19, GSG^{+}15, HBA^{+}14, HDL^{+}13, HK11, KL14, PVPF19, SMGL17, TDI^{+}14, TEVN19, dVM14, dVM10a, vHHB14]. scene [GOO^{+}19, MKU19]. scenes [KC13a]. scheme [CCC17, CJ16, SF16, WCS18, Hav10, Hav13]. schemes [KVH14, KNY14, PHS18]. scholarly [Car11, Gre15]. school [HPH^{+}18, TZBBL18, OM19]. Sci [ZLL19]. Sci-Tech [ZLL19]. Sciara [WFK^{+}12b]. Science [HGS12, BFO^{+}16, BJS18, BHA^{+}14, Bre14, BAHCG14, BGL^{+}16, BBCPC11, CGSP13, Cha17b, DSK^{+}19, De 10, DTN10, FCC^{+}12, FSS15, Fri10, GMP^{+}16, Gre15, HDG^{+}12, HD16a, HP12, HTA^{+}15b, JBKB15, JY19, KHT17, KJP14, LZL19, MWR^{+}17, MRHP19, MRSH13, Nei15, OBH15, PH16, RMRN^{+}16b, RHW^{+}19, RTM19a, Soo17, VHRM17, VSCH16, WLT^{+}19b, ZP18, vHK17, ELAC19]. Science-based [HGS12]. science-policy [Fri10, Soo17]. sciences [SH10]. Scientific [Kno10, Man10, VAHG^{+}18, CC18, CF16, CvPH^{+}18, DHR16, ECR^{+}16, GNHM18, GTM^{+}12, GGLP16, HD10b, KFM17, McL15, PC11, Pol12b, RJPS11, SOBS18, SLJ^{+}11, SWM11, SOÖ12]. scientists [BOF^{+}15, CHvKM15, NM19, dLS10, RL13]. Scomberomorus [WNB^{+}15]. scope [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, JTBP17, KFB19].

Scoping [LEKB10]. Scorecard [MFB^{+}14]. scoring [HFF11, KJP14]. Scotia [GLW15, WB18, BMW17, SM17a, SMW13]. Scotland [ALMS14, BDH^{+}18, BMGH11, CJ16, DS10, FDP19, HP13, HBP16, JKS16, KMC^{+}17, MSBB14, MAHP16, Mso16, Ros13, SM17a, San19, Smi18a]. Scotland. [SJ17]. ScotMap [KMC^{+}17]. Scottish [AWJ19, AWH13, Car14, MFT19, MG16, PTP13, WMJF16, WK19, Wri16]. scour [DBG^{+}16]. scrap [SS15c]. Scrapping [BW18b, CJ16]. screening [MTF^{+}15]. scrutiny [CKFLO19]. Scuba [DM15, KC13b, RPG^{+}14, LGDB17]. SDG [Vir18, DJR^{+}18, DMW19, NM18]. SDGs [DJR^{+}18, MSS18, OY19]. SE [BGA^{+}15]. Sea [AvTT17, BKKN10, Elo10, GCHO19, Grö18, Gue19, Gus18a, Kel18, KS18b, KSS17, KR18, LK15, MGB^{+}18, MMDS12, Shi13, Smi18b, ZMZ16, dGNC18, APY^{+}17, AJ17, And15, ATB^{+}18, BG16, Bol17, BHS19, BVB18, BSAT18, BTC18, BS18, BRM^{+}17, BHG11, CLS16, CSF^{+}18, CMW10, CKC^{+}13, CJB15, Cus17, DDM^{+}19, DJM11, DZ12, DMJ^{+}17, DLM^{+}18, ECL^{+}15, Fen17, FG18, FFH18a, FFH18b, GDN19, GBC17, GNN^{+}17, GSW18, GLW15, HBL14, HGS13, Har19, Ho10a, HBT^{+}10, HHA17, Hua11, JKK13, JFK18, JDM^{+}19, KARR^{+}19, KJS14, KCJW15, KCJ15, LACG18, LDVM10, Lil16, Mal18, Man10, MRCG11, McL15, Men19, MB18b, NNPV17, NBC^{+}12, NNP^{+}14, O’R15, Oan18, ONF15, OHB^{+}18, PKA^{+}16, PS18, PGL^{+}16, PRPS17, PLP14, RTD^{+}17, RPA17, RMR^{+}17, SSW^{+}18, SPB19, STG11, Sea17a, SCK15, Smi17]. sea [SRGD18, SSA16, SKT^{+}10, SF16, SP19b, TGdCR11, TTFC19, VSA^{+}12, VAP^{+}14, VAHG^{+}18, Vra19, WL18, WM18a, WCCO18, Zag18, ZCD^{+}15, dS17a, AAC^{+}13, AdV18, APC16, Bac18, BMF^{+}16, Boh18, BBE^{+}16, CBP19, Cho11, CE17, CCRT16, DR10, DG14, Eng18, ES18, FTG^{+}18, GSB^{+}17, GE17, Gus18a, Han16, HGG^{+}18, HG14, HP12, Hun13, Jet18, JR14, KLC^{+}15, Kee18, KML^{+}16, Kid13, Kim19, Kno10, Kno11, Lan18, LPP17, LG19, LE12, LLZ^{+}19, LLO16, LGM14, LPM^{+}14, LBD^{+}18, LSAL19, Lyo14, MACM14, Man10, MR10, MGG^{+}19, MLA16, MUHH18, MNG^{+}18, MKU19, NALO19, OM19, OHH^{+}16, OOvdM^{+}11, OS12, PQLH15, PAC^{+}19, PHM16, PdSA^{+}13, PMC^{+}15, PRPS17, RWG^{+}19, RJ18a, RJ18b, Rin18, RvdBG13, RGB^{+}19, SRA15, STG11, Sco18, SGRV16, SKL16, SH12, SS13, TPKL16, TM13, Tyn17, Tyn18, Val13]. Sea [VvD10, VvDM10, WPJ10, WC16, WZF^{+}18, Wes18, WBA18, Zha18, ZZ17c, vdBKB^{+}17, ZPT^{+}16, ZH18]. sea-based [KARR^{+}19]. sea-level [Mal18, SCK15]. sea-transport [NNP^{+}14]. Seabed [Ard18, LJL^{+}14, ARJ18, Bes11, Cho11, HLC^{+}10, Jae16, JAG16, JGA17, JML16, Kim17, LM16, LMG^{+}16, MSBB14, Mal18, PKA^{+}16, VVS18, ZZ16]. Seabird [LDS11, Ber19, CBP^{+}16, WFK^{+}12a]. seabirds [DBB^{+}14, LGM16, LMP^{+}16, OBC^{+}18, RTD^{+}17, SWP13, SBC^{+}16, WMJF16, WS14, WRW^{+}17]. seafarer [ZSZPR19]. seafarers [AT14, FFSF17a, FFSF18, NPAR19, ZZ15a, ZZ17a, ZZZ^{+}17, CS17, EZPRZ18, SFDB18]. seafaring [FFSF17b, Gek10, MM19]. seafloor [CCC^{+}13, HA13, HBT^{+}10, MPPB10, RWG13]. Seaflower [TBKH13]. Seafood [CLDP17, VRA^{+}12, WLT^{+}19b, Alf17, AKV15, BF15, BKKN10, BHI^{+}17, CdlPS^{+}14, CBS^{+}16, DL19b, Fab16a, Fab16b, FDP18, Far19, FP12, GPM19, GFH^{+}17, GFA^{+}18, He18, HA13, JBBV^{+}19, KHD12, MMP15, MM15a, NBM15, OM19, OC17, OCdS14, PNPD14, PPM17, PP19, PPM19, RV16b, SVL^{+}16, She16, SJ15, SSWP10, SSSO17, SP19b, VMO14, WNLS17, WS17, XDG^{+}16, ZT19, AGSP^{+}13, FP13]. seagrass [OTOL17]. seagrasses [GCM11]. seahorse [FKWV19]. seahorses [KV18]. seal [BMGH11, BR13b, BR14, CKBS14, CJHR14, DHSC14, ES18, JBV^{+}18, Liv10, NSB18b, Var11]. seal-fishery [BMGH11]. seal-swim [CKBS14]. seals [Cro11, Var11]. Seamount [dOSL18]. seamounts [CLF^{+}19]. search [All14, Cus17, FC19b, KPF12]. Searching [SSSO17]. Seas [RB12, Bla17, BY16, CETQ17, CSP10, ECSE^{+}17, Han11, Hof19, HDTS14, HRA^{+}15, LJ16, MSG11, MSGM^{+}12, NFHJ16, OBJ^{+}12, PGP^{+}19, RvLvTH14, RIK16, Sea17a, SKT^{+}10, VCdR^{+}18, Vir18, WK19, vTRvLvH14, MNA15, ACG^{+}17, GSS^{+}19, GNN^{+}17, HJKR14, LCC^{+}18, RvT14, TK14, TC13]. SEAS-ERA [MNA15]. seascape [WS15a]. seascapes [FE12, dlTCFB^{+}17]. seashore [MD15]. season [CA17, RNN^{+}18]. seasonal [HD10a, HID12, Mox12, TRCR^{+}19]. seasonality [OCG14]. SEAT [Ros18]. Seattle [WMK15]. seawater [HPPH16, HPP^{+}17a, KC13a]. seaweed [CLV^{+}16, HAC19, MBT^{+}19, SMH17]. seaworthiness [ZP16]. Second [DMMG15, MDNT19, RMJ12]. Second-guessing [DMMG15]. secrecy [Gla19]. sectarian [LLF^{+}16]. section [FSB14]. sectional [WP12]. sector [ABSGB18, BRR18a, Bes19, BLR15, CBL18, CdlPS^{+}14, Cro15, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, Gil14, GA14, GFA^{+}18, HBL14, HGS12, JA10, KPF17, LB17, MMP15, MV16, MLA16, MO13, MBBDS19, NBF^{+}18, Nik19, PS11a, dlÁPAVS19, PR19, RRR17, SG15a, SLBC19, TÁB^{+}10, TCC^{+}19, VSRD16]. sectoral [BMD^{+}14, CMMBV^{+}19, DZ12, MOH11, SD13, SM17a, AH19]. sectorial [NPS^{+}19]. sectors [BLC16, GGLP16, HW10, JLH14, KED^{+}18, KO11, LY14, WC17, vdBAMJT19, SNP^{+}18]. secure [LBP^{+}19]. Securing [RFG^{+}14a, RDG^{+}13, VKGN^{+}14a, WH18, SALS17, CPB19]. security [BSL19, BAA^{+}15b, BAA^{+}15a, BCMH^{+}18, BAA^{+}18, BSH^{+}19, Bue15, BF19, CG18, CM17, CSP^{+}14, DGVS^{+}16b, DCHC17, FAA^{+}13, Ger15, GGD16, GM19, HZH^{+}13, He15, Her18a, LGH^{+}12a, LGH^{+}12b, PSH18, PPML16, RLAT19, SC10, TSK18, TCC^{+}19, dVMdC^{+}19, WF18, Zer18]. sedentary [Han16]. sediment [CCRT16]. sediments [LEKB10]. See [Hea13]. SEEA [Dva19]. SEEA-EEA [Dva19]. Seeing [LMP^{+}16, MD15]. Seeking [Car19, HHL^{+}15]. segmentation [THRS15]. segmented [THRS15]. Seid [HF16]. Seine [RDP^{+}18, DMMG15, DAJM15, GJA16, MHM16, PSV^{+}14, SSW^{+}18, WIC18]. seiners [LH16a, TZBBL18]. seines [LBD^{+}18]. seining [Gil11a]. Seismic [PTN^{+}17]. selected [MSS19, TAGP15]. Selecting [BSP^{+}15, JS18]. Selection [ZZZL18, De 16, KNB^{+}18, KEM^{+}18, NBF^{+}18, YYN^{+}15]. selective [CH12, CGC14, MKSS18]. selectivity [FR16, SS11]. Self [YFR10, BAT^{+}18, FC19a, GDD13, LM15, RPG^{+}14, YFC13, YCHM14].

Self-governance [YFR10, LM15, YFC13, YCHM14]. self-guided [RPG^{+}14]. self-management [GDD13]. self-organization [BAT^{+}18]. self-reported [FC19a]. Selling [Alf17]. semantics [EMD^{+}18]. semifasciatus [WNB^{+}15]. Senegal [DYN^{+}19, DZ12, OY15, SSFL10]. sense [BDvTvKT17, HdF^{+}19, Tho13]. sense-making [BDvTvKT17]. sensing [KHP14, OIS^{+}18]. sensitive [GNW19]. sensitivity [LJS^{+}18, SS11, TMDSF19]. sentinel [CCTG^{+}13]. Seoul [CL18]. separate [SJ19c]. September [Ano10k, Ano11i, Ano12l, Ano14l, Ano15o, Ano16w, Ano17s, Ano18x, Ano19y]. sequestration [SGMWE14]. sergestid [WC17]. Sergia [LYB17]. series [KC15, ZP18, KSH^{+}15]. serious [LMG^{+}16]. serve [LOD^{+}15, TC15a]. Service [Kai16, PSCY12, BPBM17, BSM^{+}14, LKL15, RMH^{+}15, BGL^{+}16, Kai15b]. services [AOA^{+}15, ATDP16, BYK^{+}15b, BBP^{+}14, BDH^{+}18, CLV^{+}16, EGD18, Han13b, HM18, HDTS14, KLM^{+}13, LCH^{+}13, LPS12, LWvdBL16, LPR^{+}17, MD16, MTP^{+}15, MMGM16, MBB18, MM17b, NVK^{+}17, PDR11, PTML16, PSCY12, PMA^{+}17, PBJ^{+}14, RV16a, RBB^{+}16, RFHEJK13, SSCC19, San19, SGMWE14, TH17, TTFC19, VKA^{+}19, WSO^{+}14, WPRA18, YWC^{+}16]. SES [Kru16]. session [GGLP16]. set [HHS19, PDS15, SLJ^{+}11, SC13b]. sets [MKU19]. Setting [BBN^{+}13, Bor18, KJP14, CKVO16, CK18, CMB^{+}15, GGK17, GRMAC18, KS12d, LQC10, TML^{+}19, WPRA18, ItDD^{+}18]. settings [ALSC13, GPA18, VvDM10]. settlement [Ásg16, Ber19, MJQMVÁ17]. Settling [Ásg16]. seven [MHK^{+}10, TAGP15]. severely [JFK18]. sewage [WBA18]. Seychelles [CEB^{+}12]. shadow [CLS16]. Shandong [SY18]. Shanghai [ZZZ^{+}17]. Shantytown [GC18]. shape [Lew16, RFG^{+}13]. Shaping [BY16, SH15]. share [All14, CKdS14, HTA^{+}15a, JS12, JS13, Pan14, TWA^{+}15]. shared [BIPS13, BK16, BOF^{+}15, MCAM13, Zag18, PW15]. shares [BMAG17, CS13, DES10, HB14, HK18, KBBR16, Lyn14, Mac14, PCS14, RWG13, Sev14, GBF^{+}12]. Sharing [JAG16, Sev14, BAHCG14, CL15, CHDJ^{+}19, EB16, GKG16, HK11, HRC15]. Shark [NY13, SM18a, VMR^{+}18, ADH18, CMH19, CSF^{+}13, CBP19, DS17b, FBM18, FLT^{+}19, FJG14b, GH18, GW15, GW10, HSRH16, HS19, JLW17, MSB15, OPGL16, PDC19, RTM^{+}19b, ST17, SH17, SB18, Syb15, TGT^{+}19, TK11, TK13, WP17, ZKR^{+}19, dMAH^{+}18]. shark-based [TK13]. Shark-diving [VMR^{+}18, GH18, ZKR^{+}19]. Sharks [HGDB10, AMdlCMS^{+}18, BFTS^{+}16, BBM^{+}17, BW19, FJG14b, GFC^{+}15, GTL^{+}18, GAKP^{+}19, LEG19, MSB15, OP15a, OBNH15, PNW18, PGC16, WDK^{+}13, DW13, LTM19, MFF^{+}18]. Shedding [FW19, FSCJM19]. Sheik [HF16]. Shelf [KS12c, HRA^{+}15, Kay15, ZMM^{+}14, KS12a, KS12b]. shelf-seas [HRA^{+}15]. shell [MOR15]. Shellfish [VM18, AYR^{+}14, AL17, BRR18a, BBCPC11, DJTK17, GGVRVL16, GA14, HCC^{+}19, MD15, PHLM19, Sch10b, WSC19]. shellfisheries [PFMA^{+}19].

Shetland [CG13, KGST14, SSLR14]. Shiffman [PDC19, Sch15a]. shift [CP18, KPF12]. Shifting [Ben12, RJPS11, DL19a, RCD^{+}18, VBHC^{+}17, LJ13]. shifts [ASB10, BYK15a, DCY^{+}19, ÖGB^{+}10, Pri10, TCS^{+}18]. Ship [Del17, PARD17, ARDP17, CLLL17, CZX^{+}19, GWWH17, MÁSSI13, MPT13, QZH^{+}18, Rin18, WBA18, ZMZ16, ZZ15b, ZZ17b, dlFRG17]. ship-source [Rin18, ZZ15b, ZZ17b]. shipbreaking [AF14, RM16]. shipmates’ [MM19]. shipped [SMGL17]. Shipping [LSW17, RMJ12, AT14, Fen17, FSV^{+}16, Gil14, Gri17, HK11, Ho10a, HDH^{+}15, HBH^{+}19, JBK15, KS16, KS18a, KS19, LK15, Man17, MGB^{+}18, MGGS18, NNP^{+}14, PKPdP10, PT10, RIK16, REA^{+}14, RPSS17, RRCS^{+}19, Sea17b, Shi16, WML17, WLX19, WMVR18, WvLvK12, YC18, ZMZ16, ZSZPR19, WLT^{+}19a]. ships [AVC18, Bac18, Bro11, KS16, KS18a, NPAR19, PHS18, SCKP19, WMVR18, ZLSL17, SLDHS18]. shipwreck [KB15]. shipwrecks [Hua14]. Shizugawa [SOUN16]. Shoal [BN19]. shocks [GFA^{+}18]. shore [GOdFC16, KLN17, LBK^{+}13, LBP^{+}19, MV16, WF18]. shoreline [FBP^{+}18]. shorelines [SGN^{+}17]. short [LBHS16, CH12]. short-term [LBHS16]. Shortcomings [BH14a, AdSC11]. shortfin [FLT^{+}19]. Should [Kim17, TGS18, CCC17, LVRM^{+}18, MG14, PLL17, PDS15, SJ19c]. shove [MJvP^{+}18]. Shrimp [AA17, EAA^{+}18, HDK^{+}15, CPKF19, CSMR18, FMN15, FV10, GB15, GKG16, GGK17, GM14, MBD13, NAYSN18, WC17, ACTG17]. SIA [MML^{+}19]. Sian [VAW14]. Sicily [PBSG12]. side [AH19, AdV18, CHL^{+}19, FW19, KLN17, PRPS17, UL17]. side-effects [CHL^{+}19]. SIDS [MCY^{+}19]. Sierra [SBM^{+}17]. sight [CCL^{+}19, PL15]. signage [MMJM15]. signed [KEM^{+}18]. significance [MLA16]. Significant [DAB^{+}14, GAP12, YYN^{+}15, YSJU^{+}17]. signs [Pin17]. Sikuliaq [KFM17]. silent [TEVN19]. silver [ARS16, MLF^{+}17, MM19]. similarities [RRRM^{+}15]. Simmering [TBB12]. SIMP [He18]. simple [GFMG^{+}14]. simplicity [Sin19]. simulate [HAH^{+}16]. Simulated [RRCS^{+}19]. simulation [CA16]. simultaneous [IT12, TI13]. since [BGG19, BE15, CP13, Nag16]. Singapore [Rah14]. Singaporean [Lim16]. single [HL15b]. singular [BDF^{+}10]. sino [PWR17, Cha19]. sino-ghanaian [PWR17]. sister [KR18]. site [BJO19, De 16, JSG^{+}18, KNB^{+}18, RJ13, TTP12]. sites [KB15, PW16, TAGP15, CHSW17]. siting [GCD^{+}15, HDJD15]. situ [JOE^{+}19, MMJM15]. Situating [SS19b]. situation [BBA15, CMA15, DGC14, VLPC14]. six [Gar12, HM18, MMJM16, Par10]. six-decade [Gar12]. Size [CH12, HTF^{+}15, AAL17, DWCS18, DRH^{+}18, Hor10, LL15, SJ10, Wit19, ZH14]. size-selective [CH12]. Skagerrak [SKL16]. skates [SS16]. skin [PC11]. Skipper [And15, VRT13, vS15]. skippers [Pat15]. Slavery [MV16]. sleuthing [WLT^{+}19b]. Slow [Man17, STG11]. SMA [TTM18a]. Small [CE13, DK16, DMB^{+}18, HWO^{+}18, HIBG15, KSS17, LCC^{+}18, LBD^{+}18, RTM19a, SOBS18, SNA^{+}17, SALS17, SC10, Weg11, ALSC13, ARSHH16, ATB^{+}19, BSD^{+}17, BRR18b, BBR^{+}18, BBE19, BPB19, BCMH^{+}18, BAA^{+}18, Ben17b, BM15, CPR14, CMAA13, CN17, CC17a, CRB^{+}19, CBG14, CP15, CF13, CBS^{+}16, Cro15, DRGP19, Dav15, DL19a, Deb18, DYN^{+}19, DRH^{+}18, EA19, FB15, FDGM16, GLVLGG19, GSJ18, GMD15, GHM^{+}18, GTL^{+}18, GOTS19, GA19, GMM15, HWL^{+}19, HDM13, HAC19, IW19, JOE^{+}19, JMJ^{+}18, JTK18, JHN^{+}19, KL14, KLM^{+}13, Knu17, Lan15, LDF^{+}18, LLN17, LH16b, MB10, MGA^{+}10, MJQMVÁ17, MMTC16, MN18, MMC^{+}14a, MMC^{+}14b, OC17, PRBX19, Par15, PD15, Pin17, PVPF19, PGPS10, Pom12, PC18b, PRPS17, PKM^{+}19, PCLE17, RROR12, RNS19, RBTR^{+}19, RV16b, RPC^{+}17, RBJ^{+}11, RCMBG^{+}19, RD13, RRC^{+}17]. small [SDRML15, SC19, Sal16, SLS18, Sat18, SCC^{+}17, SGNM19, SSD17a, SAM13, SL15a, SSRS14, SC11, Sun12, TRM19, TTS14, TLA16, THE^{+}18, TLA18, VPR^{+}16, VPP^{+}16, WW17, WBC14, WZCHS10, Wos19, YP17, GS17, CPB19]. small-boat [CC17a]. Small-Island [DMB^{+}18]. Small-Scale [SALS17, HWO^{+}18, HIBG15, LCC^{+}18, RTM19a, SNA^{+}17, SC10, ARSHH16, ATB^{+}19, BRR18b, BBE19, BCMH^{+}18, BAA^{+}18, Ben17b, BM15, CMAA13, CRB^{+}19, CBG14, CF13, CBS^{+}16, Cro15, DRGP19, Dav15, Deb18, DYN^{+}19, DRH^{+}18, DK16, EA19, FB15, FDGM16, GLVLGG19, GSJ18, GMD15, GHM^{+}18, GTL^{+}18, GMM15, HWL^{+}19, HDM13, HAC19, IW19, JTK18, JHN^{+}19, KL14, Lan15, LDF^{+}18, LLN17, LH16b, MB10, MGA^{+}10, MJQMVÁ17, MMTC16, MN18, MMC^{+}14a, MMC^{+}14b, OC17, PRBX19, Par15, PD15, Pin17, PVPF19, PGPS10, Pom12, PC18b, PRPS17, PKM^{+}19, PCLE17, RROR12, RBTR^{+}19, RV16b, RPC^{+}17, RCMBG^{+}19, RD13, SDRML15, SC19, Sal16, SLS18, Sat18, SCC^{+}17, SGNM19, SSD17a, SL15a, SSRS14, Sun12, TRM19, TTS14, TLA16, THE^{+}18, TLA18, VPR^{+}16, VPP^{+}16, WW17, WBC14, WZCHS10, Wos19, YP17]. Small-scale [CPB19]. ’small-scaleness’ [GS17]. smaller [KDG^{+}15]. smartphones [JHN^{+}19]. Smith [BFTS^{+}16]. smolt [HHL^{+}18]. snapper [ASWM18, CSMB^{+}12, SAdC15]. snappers [NWW^{+}15, NWW^{+}17]. snapshot [NMJS17]. snorkelling [RPG^{+}15]. snow [Han16, MI15, ØR18, Dav15, MMDS12]. Soares [GPB^{+}18]. Social [BTS^{+}17, Bre15, Bro16c, CMMBV^{+}19, CCS16, FSS15, GPA18, HMGR14, HdV17, KGAB14, KPF17, MF19c, MLTBMB14, SDRML15, SvdBH^{+}18, WM18a, AG18, AL17, APD^{+}19, BdKvT19, BBN^{+}13, BJS18, Ben17a, Big15, Big16, BCE^{+}17, BC13, BSP^{+}15, Car15, CKdS14, CF17, CJW^{+}16, CE13, CNFJ10, CKHJ^{+}17, CK18, DRGCPS17, ECSE^{+}17, FSWFG13, FG18, GGK17, GR16, GBW16, Gat18, GA19, HRJH15, IOG18, JWBJ13, JG13, KVA18, KMF^{+}18, Kim10, KFB19, LSP^{+}17, LGFB^{+}19, LCC^{+}18, MKX17, MMP15, MWR^{+}17, MR10, MCH18, MBBDS19, MAHP16, MI15, MD15, MDM16, NMvP19, NHPD19, ÖSS^{+}11, Par15, PO10, PR13, RCD^{+}18, Sol14, SGGM14, ŠST^{+}18, SH10, TMG^{+}18, Vil12, VGG12, VGG15, VGBG15, VBMM17, WLX19, WCCO18, dlTCFB^{+}17, BD10, Bro10, CSCBR10, DCY^{+}19, FMN15, JQD13, MJQMVÁ17, PDS15, SSD17b]. Social [UAZ13, XJT18, YBD^{+}19]. social-ecological [AG18, Ben17a, Big16, BCE^{+}17, CNFJ10, GGK17, IOG18, LSP^{+}17, MI15, MDM16, PO10, TMG^{+}18, Vil12, dlTCFB^{+}17, YBD^{+}19]. societal [HNC14]. society [BC13, DTN10, GSW18, HdF^{+}19, RMRN^{+}16b, RDG^{+}13, SH15]. Socio [BDB19, CS14, CVP^{+}17, CA17, EJB19, KVMM10, RvdBG13, SG17, SMS^{+}18, SRGGVL14, CCC17, CSP^{+}14, GFA^{+}18, IH18, LLF^{+}16, NMvPNL15, PC18a, RDP^{+}18, RCUR13, RMH^{+}15, Ros18, WHP^{+}16, GAKP^{+}19]. socio-cultural [IH18, RCUR13]. Socio-ecological [BDB19, CCC17, LLF^{+}16, NMvPNL15, RDP^{+}18]. Socio-economic [CS14, CVP^{+}17, CA17, EJB19, KVMM10, RvdBG13, SG17, SMS^{+}18, SRGGVL14, CSP^{+}14, PC18a, RMH^{+}15, Ros18, WHP^{+}16, GAKP^{+}19]. socio-political [GFA^{+}18]. Sociocultural [Tir17]. Socioecological [PFMA^{+}19]. Socioeconomic [DYN^{+}19, CMW10, MJS16, SKS^{+}19, WSG18, AOAM13, SKS^{+}19]. Sociological [HK10]. SOFIA [PF12a, PZ17b, PZ19]. SOFIA’S [LJ17]. SoFISHticated [GBW16]. soft [CCRT16, CHL^{+}19, MOR15, MDNT19]. soft-shell [MOR15]. software [CVFP^{+}12]. Sole [CMVLGG14]. Solent [MPF12]. Solomon [WMVR18, AOA^{+}15, Bre13, Bro16b, CSP^{+}14, DGVS^{+}16b, PRT10, RLAT19]. solution [AG16, All14, ESM^{+}19, MNHN18]. Solutions [ACTG17, AOY16, HTD^{+}18, JHB^{+}18, LHJ^{+}15, LZ16, PLP14, SLS18, Val13, VS18, VKGN^{+}14b, vAS19]. solve [LV18, RM17]. Solving [JE11, CSBE14]. Somali [CGP^{+}18]. Some [KMTS16, Bra12, PHEM^{+}17, SLS^{+}13, ACA^{+}13, Bro16a, Gil11a, WM17, ZMZ16]. Something [TMR^{+}14]. Sometimes [vHHB14]. Song [GDG^{+}18]. sooner [MCMB^{+}18]. Sound [BHG11, RGWH18, San18, SLJ^{+}11, WLP10]. sounding [vS15]. Sounds [TMG15]. soundscape [HGH^{+}18]. source [Nij17, PG19, Rin18, dVM17, ZZ15b, ZZ17b]. sources [BDIS18, BV19, DMdJ^{+}15, PFCP^{+}14, VRT13]. Sourcing [CLDP17, SSWP10]. South [BS10, KJH^{+}10, LU13, MAR^{+}15, Nor15, PR13, RHP^{+}17, Wit15b, FPG^{+}17, HEJK10, HR17, JRS^{+}16, MdMGV^{+}18, PRBX19, RPdO^{+}19, UA13, BF15, Bro10, CPRP19, CL18, COS14, GFC^{+}15, GPB^{+}18, GNN^{+}17, Gre15, HBS^{+}19, IW19, KC13a, KLY14, Kim15, Kim19, LLS^{+}15, LM14, LM15, LSMW10, LLZ^{+}19, Lyo14, MVRG13, OS12, PDT^{+}14, PIH^{+}10, PGC16, Sco18, SPRH19, dOSL18, SC10, SSRS14, Sun12, TvN13, TJ18, VBSD17, Wit15a, Zha18, ZZ17c, dS17a, dCW15, vAS19]. south-east [UA13]. south-west [HR17]. Southeast [DHM^{+}17, GMP^{+}16, Pom12, PPCM16, PGP^{+}19, TH15, YSJU^{+}17]. southeastern [ATB^{+}18, DRMA18, FPG^{+}17, Kam18, TJ13, MMN18, STIC19, SD18]. southeastern/south [FPG^{+}17]. Southern [RG19, CVFP^{+}12, DBB^{+}14, Lan15, LEFC19, MGG^{+}19, NSB^{+}18a, SMBL13, Sea17a, SBMD13, SPRH19, SGP^{+}17, ZVP^{+}13, Bol17, CJB15, FMB^{+}13, JL11, LB18, MMS^{+}14, Nei15, OPMB12, PP10, PAD^{+}17, Pol12a, SRSN17, Tan17].

Southwest [RBW18, JHN^{+}19, PGL^{+}16, BMW17]. southwestern [LBHS16]. sovereignty [ABBD17, LLN17]. SOx [KS19]. sp [BSSS17]. Space [TMG^{+}18, AOY16, BW15, Car19, JHW13, Lan18, vdBAMJT19]. spaces [LLA^{+}19, ŠST^{+}18]. SPAF [QSICTdV19]. Spain [ASP16, CdL17, FVM14, GLVLGG19, dlFFVRC16, MJRdC16, dlÁPAVS19, PFMA^{+}19, RGGFA17, SRS14, Thi13a, dO13, ADGGVL12, AM15, CVFP^{+}12, FMMA^{+}15, GGSRVL17, MÁSSI13, MPMGGAHR18, RRMSSE16, RVdCGN11, SGNM19, Thi13b, VRMF11]. Spanish [AG12, CMVLGG14, CMFDCLV10, dHJTG17, GGVRVL16, GGADSR17, MPPSdVSOHR14, MSGM^{+}12, RRR17, RR18, VRMF12, dVM12]. Spatial [ADL^{+}18, BFGR14a, BRR15, CDR^{+}16, DSM^{+}13, De 11, DTM18, FOO^{+}15, FHL18, GCK^{+}11, HDG^{+}12, KBB^{+}15, Kid13, LDF^{+}18, MO14, MG16, OBC^{+}18, RROR12, TK14, vH19, AdSC11, BdKvT19, BBN^{+}13, BW15, BDF^{+}10, BSW19, BHS19, BB16, BFGR14b, BMGF^{+}17, CNJ^{+}10, CBN^{+}12, CB13, CMP^{+}15, CACL18, CSVHS14, Car13a, CC13, CL16a, CSBE14, CGM^{+}18, Cra19, DTJ10, FCdS^{+}15, FÓ12, FJ12, FMB^{+}13, FHM^{+}10, FWK16, GSR19, GSG^{+}15, GdV16, GFC^{+}12, GM14, HF16, HGG^{+}18, HD10a, HID12, HDJD15, HCJ15, HGBJ16, JMJ^{+}18, JKK13, JBKB15, JAO^{+}16, JJ19, JLQ16, KMC^{+}17, KGST14, KS18b, KRK13, KMR^{+}13, LCH^{+}13, LWJH11, LLX^{+}15, LBD^{+}18, MACM14, NVK^{+}17, NMvPNL15, NHPD19, NKBS17, NM18, OO19, OMF14, OLF19, Pap16, PWH^{+}16, PVVB17, PGL^{+}16, PFCP^{+}14, PGB^{+}17, PGB19, Por15, PV16, QJ13]. spatial [QSICTdV19, RRA^{+}10, RRMSSE16, SDF^{+}14b, SDF^{+}14a, SOCA15, SPO^{+}18, SJ17, Smi18a, SM18c, SBC^{+}16, SLS^{+}13, SMKBM16, TGT^{+}19, TvN13, Tam15, TKAO^{+}19, TEVN19, TGdCR11, VCdR^{+}18, WS14, WK19, Wes18, ZPT^{+}16, dGNC18, dVM12, DTMJH16]. Spatially [RFHEJK13, TMDSF19, CGM^{+}18, SBS^{+}13, SFC^{+}15b]. spatio [HAH^{+}16, HDTS14, SRA15, SMS^{+}18]. spatio-temporal [HAH^{+}16, HDTS14, SRA15, SMS^{+}18]. Spatiotemporal [CCD^{+}13]. spawning [HGB^{+}19]. speak [HSRH16]. special [De 17, De 13, DMW19, FKMV16, HMK19, MHF13, Nei15, VKGN^{+}14a, WFSV15, Cho10, De 17, TTM18a]. Specialised [MCH19]. specially [MSH13]. Species [BW19, DN12, MGH14, BF15, BAO^{+}12, Coc15, CP18, FSB14, FSV^{+}16, FDS15, Gib10, GMM^{+}13, GJA16, HA13, HSB^{+}16, ITNL^{+}14, Kas16, KS12d, KCJ15, LOD^{+}15, LCC14, MACM14, MCCC16, McL15, MMBK17, MUHH18, NMvPNL15, NBF^{+}18, OGM^{+}14, SKFL15, SBTF15, Sch15a, SDC13, SGW^{+}15, SS15a, SH16, WHB^{+}18, WWH^{+}17, WBH^{+}15, dMAH^{+}18, vH19]. Species-specific [BW19]. specific [BW19, CSVHS14, CP18, MBD13]. spectre [FBAK14]. speculations [SM18a]. Speed [SHM12]. Spencer [RHP^{+}17]. Spermonde [MFRvP16]. SPICOSA [Rei14]. spill [Ber19, BW18b, CMFG15, CSO17, CL16b, GSS^{+}18, JLY18, PQLH15, SS13]. spill-over [BW18b]. Spillover [CSMR18]. spills [GNW19]. spinner [HVB^{+}17, MJS16]. spiny [BRRG^{+}18, HKR19, Uhr16]. Spirit [Cho10, CL16b]. spoil [LBN15]. Sponsoring [Oya18]. spontaneously [GB15]. spoon [Cus17]. Sport [WBSW13, Wen11]. spread [MO14, Pin13]. spurs [DB19b]. Squalus [HKR19]. square [SH14b]. squid [COS14, RRAH13]. Sri [MJS16, DAJM15, GB15, GKG16, GGK17, GJA16, HDK^{+}15]. St [BAO^{+}12, dOSL18, Sch15b]. St. [Doy15, dOSL18, vS15]. Stability [IHS13, HdVB15, MÁSSI13, Val13]. Stabilizing [RBW18, SRK14]. stabiplage [DV17]. stacking [GE17]. Stade’ [UA13]. Stafford [AGWMP19]. stage [AMvT11, BBN^{+}13, LWS19, UEMRG^{+}18]. Stakeholder [BY19, BSM^{+}14, CHKG16, CRO12, EGF^{+}17, HD10b, SNP^{+}18, SF19, TMHJ12, AOY16, ARDP17, ATDP16, BM15, CS14, Car19, CBP^{+}18, Dix16, DFK13, FJ12, GFC^{+}12, JBV^{+}18, KSK^{+}19, KHvH^{+}14, Kru16, KCCO18, KH15, LJS^{+}18, LRMM17, LEFC19, LHB^{+}17, MM18, MPPB10, MLR^{+}16, MB18a, PPCM16, RvLG^{+}15, RLHM14, SBTF15, SSB^{+}18, SBP^{+}19, SG15b, SWM11, SFCI14, TMR^{+}14, VvDM10, WB18, WKB15, YY19, vLRvH^{+}14, BN19, HGS13, HJKR14, HAC19, How18, MPF12, ODRRM12, QSICTdV19]. stakeholders [BK16, CKC^{+}13, CA17, EMP^{+}14, GOO^{+}19, KSB15, KCM^{+}18, MWGD11, MPF12, MNLHRMP17, SBC^{+}15b, SMW13, VPP^{+}16, YWC^{+}16, dLS10, PRLCL12, PPT10, FVT19, HMGR14, LJ14, TR18]. stakes [He15]. stalked [SFCI14]. stamp [Gek10]. standard [BF15, CA16, HF12, LJT19]. Standards [KW14, BB19b, BB19c, HBK17, JML16, WH18]. standing [NMvP19]. start [GGK17]. start-up [GGK17]. State [CW15, CZX^{+}19, NACS17, PF12a, TSK18, Vra19, ASWM18, AMdB12, BPB19, Bra12, CR10, CMFDCLV10, Goo17, GCW^{+}18, Kar17, KAPF^{+}19, Lea19, LZ16, Nei15, PDMÖ19, RSMM19, RBB^{+}16, SSCC19, TÁSLC19, Weg11, YLZ^{+}19, CL19, GLTB^{+}15, MCHO11, OMF14, QSHI17, Sch17, PZ17b]. state-led [LZ16]. state-wide [SSCC19]. Stated [LL15]. States [Gla19, HS19, PDC19, Ant16, BBR^{+}18, Kee18, Par10, RWG^{+}19, RNS19, WF18, Asa12, Ásg16, CKVO16, CJW^{+}16, DSH14, DMB^{+}18, GLW17, GBF^{+}12, HPP^{+}17b, Iri18, KS12c, LM11, Oya18, PMN^{+}15, SOBS18, SH17, YCA^{+}12]. static [BRR15]. Station [HGH^{+}18]. statistic [PZ17a, YBB^{+}17]. statistical [HF12, VGA15]. statistics [Cha14, MK14, SARF16, SHM13]. Status [AZAMG13, FWC^{+}18, Gil15, MMC^{+}14b, RRRM^{+}15, UNS^{+}18, BJS^{+}16, BRR^{+}12, Cam15, CCK19, COS14, CCRT16, CPB19, DFSB15, DAB^{+}14, FDS15, GTL^{+}18, HWN11, HAC19, Lea19, LPP17, MMGM16, MGB^{+}18, Oan18, PR19, TPKL16, TBJC18, VSRD16, VvDM10, FCC^{+}12, NALO19]. statutory [WHB^{+}18]. stealth [WRP^{+}10]. steaming [Man17]. Steller [Man10]. step [LWvdBL16, ZZ16]. steps [EGF^{+}17, MGH14, NM12, Par15]. stereotypes [FVH18]. Stevens [DSR12, KSH^{+}15]. Stewardship [Cha17a, GSA10, LSK18, LHN18, MHM16, TZBBL18, GDD11, LRJ^{+}19, RB15, RWG13]. Still [ZPT^{+}16]. stingray [FSCJM19]. stochastic [VVS18]. Stock [PW15, ZH14, CCK19, CPB19, DRPSA17, DDP16, DES10, FP13, GCHO19, Kai19, NWW^{+}17, RW15, SV15, TFBS19]. stock-rebuilding [DRPSA17]. Stockholm [RM18]. stocking [San10, San11, Sev13]. stocks [BIPS13, BHAHW17, CDA^{+}13, CSH17, CA16, CRO12, FBAK14, HK10, MCAM13, OL10, OHQ^{+}16, OSS14, SKT^{+}10, VvDM10, WNB^{+}15]. stones [DL19a]. Stopping [Hof19]. storage [MSBB14, Mil14, SGMWE14, ZZ17c]. stores [KL12]. storm [WSGID17]. stormy [LJ16]. story [JRP17, McC17]. Strait [Lal17, HDH^{+}15, HBH^{+}19, RYD18, SSL16, vPLB^{+}13]. Straits [HIN^{+}17, Rah14]. stranding [LLZ^{+}19]. Strandings [WLV15]. strange [MK14]. Strategic [BV19, NKAP13, Gol13, GNW19, LWvdBL16, PDR11, SS19a, SNTL18]. Strategies [CMSM^{+}16, JRP17, Lin13, MBvZ15, BLR11, CL13, COS14, Cro15, DSM^{+}13, DRG^{+}16, DGVS^{+}16a, DGVS^{+}16b, GW11, HPW^{+}18, KARR^{+}19, KKWS15, KVMM10, LSTD12, LH16a, LCC14, LSC14, MKB17, MH13, MSG18, MI15, NGS16, PR19, PKM^{+}19, RRG^{+}12, RDVSC16, SB15, SSD17b, SSD^{+}16, SC10, TK11, TL16, VMR^{+}18, VCdR^{+}18, WCMHO18, Weg11, ABBD17, Far19, SSD17a]. Strategy [AGB^{+}18, ACG^{+}17, BJGMG18, BR13a, BDG^{+}14, BFGR14a, CETQ16, CBE^{+}19, KPR^{+}17, KTD^{+}15, LJB^{+}14, MKU19, PRO16, RDSB10, SGGM14, dVM12, vHHB14, vLvHvT12, vLRvH^{+}14, Bai16, BLG^{+}18, BFGR14b, BGL^{+}16, CL16a, DRG^{+}16, FGKMP14, LLO16, MM15b, ODRRM12, PMC^{+}15, QSICTdV19, SM18a, SLBC19, SH14c, Thi13a, Thi13b, ZPB^{+}13, LLO17, Pen18, STJ^{+}12, dVM14]. stream [LDS11, WS14, WRW^{+}17, WCS18]. streamlined [Kru16]. streams [PAC^{+}19]. street [Sev13]. strengthen [GW10, RFL^{+}18]. Strengthening [BCE^{+}17, Chu10, Nik19, MM18]. strengths [FGKMP14, NACS17, MMM19, RGB^{+}19, SSC^{+}15]. stress [MI15]. stresses [MMP18]. Strictly [Han13b]. strike [MPT13]. strikes [RGBT16]. Striking [Goo17, SS19a]. string [MÁSSI13]. striped [Mat14, MSG18]. Stronger [ABBD17, dGNC18]. structural [Bro10, PWB19, SGNM19, Seu16]. structure [CCL^{+}19, CP15, DPJ^{+}13, Hav10, HM18, KAPF^{+}19, Soo17, ZLL19]. structures [ECL^{+}15, KYS10, RvLvTH14, RMRN^{+}16a, ZM15]. Structuring [ERCMD11, SGGM14, WBGJ18, GdV16]. struggle [BJS18]. Struggling [AG18]. Stubborn [IHF15]. students [HPH^{+}18]. studies [Bai16, ECR^{+}16, GFA^{+}18, HGS12, JLQ16, MHK^{+}10, PSCY12, SM18c, SMW13, WRW^{+}17, WM17, WSGM16, Yos17]. Study [dVM10b, ASP16, AMLK^{+}12, ASBAM13, ALMS14, ADL^{+}18, BAA10, BGA^{+}15, Ben17b, BPMY18, Bol17, BB19b, BB19c, BM15, BDH^{+}18, BHG11, CVFP^{+}12, CPRP19, CSP^{+}14, Cro15, DFP^{+}16, DPH^{+}15, DR14, DCC19, Don17, DV17, FD11, FFSF18, FSS17, FWHS17, FFH18a, FLWG14, FDGM16, GGADSR17, GLTB^{+}15, GA14, Gla19, GCMA^{+}12, Gor19, GSA10, GCK^{+}11, Gus18a, HvD13, HHL^{+}18, HAH^{+}16, HWY^{+}18, HBA^{+}14, HHL14, HVB^{+}17, HdF^{+}19, HP12, HSR^{+}18, JA10, JCD17, JSY^{+}19, Kat16, KBG^{+}18, KEG12, KH10, KJH^{+}10, KFM17, KHP14, KV18, LLS^{+}15, LKL15, LJB^{+}14, LU13, LSC14, LPM^{+}14, LEFC19, LT19, LSAL19, MSBB14, MKX17, MOE11, MCCC16, MSS19, MBB18, MMBK17, MSG11, PRT10, PIH^{+}10, PSH18, PBSG12, PW16, Pol12b, PV16, RPdO^{+}19, RBW18, Rei19, RNS19, RBTR^{+}19, RNCGB15, RJPS11, RBS^{+}15, SOUN16, SLJ^{+}11, San18]. study [Sea17b, Sev13, SPMW12, SOÖ12, TFBS19, TMG15, TMDSF19, THK^{+}18, VLKN17, WL18, WP12, WLT^{+}19b, WM18b, WSGID17, WTA^{+}16, WPRA18, Wri15, WvLvK12, XDG^{+}16, YPS13, ZTAO13, ZZZ^{+}17, ZLSL17, dMAH^{+}18, dO13, PW15]. sturgeon [ES18, MGG^{+}19]. styrofoam [LHJ^{+}15]. sub [MSBB14, WTA^{+}16, NVK^{+}17, TCC^{+}19]. Sub-Antarctic [NVK^{+}17]. sub-lethal [WTA^{+}16]. Sub-Saharan [TCC^{+}19]. sub-seabed [MSBB14]. subjective [AYA^{+}19]. suboptimal [AFR^{+}19]. Subsidies [RRAH13, SCC14, SKT^{+}10, SDB14, CO16, CMS19, GAC^{+}19, He15, KKSC19, LBVN19, Mal16, Mar10, SCC^{+}17, SLL^{+}16, SES^{+}19, Tip16]. subsidy [Cha16, CMSM^{+}16, DRGCPS17]. subsistence [BRW^{+}13, SSG12, Wen11]. substances [CPRP19]. substandard [CW15]. substantial [San10]. substantially [WHB^{+}18]. substantiating [GWM^{+}10]. substitution [DSM^{+}13, JG13]. subtleties [BHM16]. subtropical [SNA^{+}17]. Success [Kim15, BJS^{+}16, Ben17b, DSR12, GH15, JFT^{+}18, JLCB16, MLR^{+}16, NFHJ16, VPR^{+}16]. Successful [WMSB15, KH15, LPC18, OTOL17, RL14, THPJ17, WMWH18, WSGM16]. successfully [CBP^{+}18]. Sudoeste [SFCI14]. sufficiently [ABCDMG17]. suggest [SS16]. Suggested [GDG^{+}18, HO12, SMQ^{+}16]. Suggestions [LRS^{+}13, Yat14, MRHP19]. suitability [HIN^{+}17, TMDSF19]. sulfide [CCC^{+}13]. sulfides [HBT^{+}10]. sulphur [Gil14]. Sultanate [ASBAM13]. Sulu [MR10]. Sumatra [SWB^{+}14]. summary [KBB^{+}15, KSH^{+}15]. Sunken [RD12]. super [NSB^{+}18a, Sin19]. super-whale [Sin19]. Supercluster [DS18]. supermarkets [SGLY13]. superprofit [BV19]. Suppl [WMVR18]. Supplement [KS18a]. supplies [EZBL19]. Supply [RG19, GKG16, HHS19, MM15b, PP19, RW15, SP19b, TRC^{+}17, TCC^{+}19]. Support [BJvH18, MMC^{+}14b, AT14, AFR^{+}19, BD14, BHS19, BRR^{+}12, Bur16, CH16, CD15, CHL^{+}19, CP18, DS17b, DFL15, EFK^{+}19, FJG14b, HWL^{+}19, HCJ15, IOG18, JWM^{+}16, Kai16, KNB^{+}18, KARR^{+}19, LPR^{+}17, MGH14, MMJM16, MSBR15, MDR^{+}16, MSG18, NKHL18, NHPD19, OPGL16, PFCP^{+}14, PGB^{+}17, PGB19, PBJ^{+}14, RFG^{+}14a, RWL14, RNS19, SSB^{+}18, SSD^{+}16, SvHdB^{+}15, SLS^{+}13, WM19, WBGJ18, vTRvLvH14, AGB^{+}18]. supported [BDSB16, BLR11, GFH^{+}17, LBVN19]. supporting [RB15, VKGN^{+}14b, GSB^{+}17]. supportive [TvN13]. supports [GC18, TRCR^{+}19]. supranational [KNY14]. surf [RR14]. surfers’ [Rei16]. Surrender [Deb18]. surrounding [Lyo14, SKL16]. surveillance [BJGMG18, Cha14, MSH13, RBS^{+}19]. survey [ARDP17, AYA^{+}19, Car13b, CC17a, CSS15, DDvP^{+}16, FC19a, GPA18, JS12, JS13, JFA17, NY13, PTN^{+}17, TKM^{+}18, ZZ17b]. Surveying [PC13b]. surveys [AAL^{+}18, BFL18, BLG^{+}18, DPV^{+}15]. survival [MGA^{+}10, ZT19]. susceptibility [TRM19]. Sussex [GKSJ11]. Sustainability [CBG14, GSC^{+}16, HAW19, MML^{+}19, McC17, NBC^{+}12, AARK15, Ant16, AFCMET18, CCC^{+}14, DSR12, DM15, DS17b, DMB^{+}18, EGHC17a, ECL^{+}15, FGPP18, GPH18, GBF^{+}10, GAFD^{+}18, Kel18, KDG^{+}15, KNY14, MM15a, MMC^{+}14b, OP15b, PZ17a, PR13, RBTR^{+}19, Sch12, SU11, SS15b, SS15c, SNTL18, SF13, SJ15, SKT^{+}10, SSSO17, TH15, VG15, VGCF18, YP17, YCY^{+}15, YBB^{+}17, Zag18, DRH^{+}18, Gus18b]. Sustainable [Bro10, CPB19, Fab16b, FSWFG13, MS19, RFL^{+}18, SSFL10, VKGN^{+}14b, WL18, ZQZ19, AMdlCMS^{+}18, ASP16, ATAM10, Alf17, ARGW14, BF15, Ben18, BAB^{+}14, BYK^{+}15b, BWD^{+}19, Bol17, Cha11, CT16, CCD^{+}13, CKBS14, CBP^{+}18, DP17, DPB17, DMW19, Dru10, ETEWZD11, FWHS17, GSC^{+}16, HGS13, HS19, IMA16, KCKA^{+}18, KR14, KKSC19, Liu13, LTM19, LKMN16, MDG^{+}14, Mar10, MCY^{+}19, Mes12, MPPSdVSOHR14, MV17, MSS18, MDR^{+}16, NNPV17, Nik19, NNP^{+}14, OY19, OCdS14, PC19, PHC14, PRS^{+}19, PDC19, RD12, RWG^{+}19, RUH^{+}14, RTM19a, SvZdJ15, SC19, SKFL15, SSCC19, SDF^{+}14b, SDF^{+}14a, SOCA15, Sco16, Sea17a, SDP^{+}16, SH17, SCV^{+}17, TBKH13, Thu10, TM13, TRBS13, Vir18, VKGN^{+}14a, WBA18, dGNC18, vHFJ^{+}19, vOHO^{+}17, Car14, EMD^{+}18, ETEW15, KHH13, KSH^{+}15, NM18, SI18, SCMS^{+}18, SALS17, Vir18]. sustainably [JGG^{+}16, Vir18]. Sustained [NBD15]. Sustaining [BLCD16, CF13, DG14, TLR^{+}12]. Svalbard [TN15]. SW [CVP^{+}17, GTL^{+}18]. swapping [HdVB15]. Sweden [BW18b, KHP14, LE12, Sev13]. Swedish [CSH17, HNVZ12, JBJ12, San11, SCS16, SG12, WPJ10, WBH^{+}13, ZH14]. Sweeping [PC11]. swim [CKBS14, FSS15, KH10]. Swimming [KHH13]. switching [Bur15]. Sydney [KH13]. Symbiotic [Kel18]. symbolic [VG15]. syndrome [CSH17]. Synergies [HCC^{+}19, DMW19]. synergy [KB15]. synthesis [CMP^{+}15, GGLP16, JLH14]. System [BB19a, Big16, Bod17, Bor10, CSVHS14, CCC17, CLS10, Cha11, CRB^{+}19, Cou11, FB17, GS16, Her18b, HCF^{+}16, HCF^{+}17, JTP18, Joh19, KSC^{+}19, KDG^{+}15, LSP^{+}17, LG19, MDG^{+}14, MB10, McC17, MD15, MDM16, NM12, Par15, Pin13, RGGFA17, SG15a, San11, SMR^{+}15, Smi18a, SCS16, TKAO^{+}19, Tor16, Vil12, WBH^{+}13, YCHM14, dlFRG17, DM15, Sco18, STB16]. Systematic [GSW18, ACP^{+}14, AMR14, BHM16, ESM^{+}19, FFH18b, GVSZ19, LLS^{+}15, MMGM16, SCK15, OBH15]. systematically [EZBL19]. systems [ACM^{+}12, Ben17a, Big15, BCE^{+}17, BE15, FSWFG13, Gil14, JCBM10, JGMS16, Kim10, KJP14, LR15, MF19a, MNM10, MOR15, MMA16, OCdS14, ÖGB^{+}10, PO10, Rei14, RSS14, TDI^{+}14, WHP^{+}16, BJvH18, GNIL14].

table [GFC^{+}12]. taboos [Sha17]. TAC [HJN16, KS12d]. tackle [Gil14, HPH^{+}18]. Tackling [PF12b]. TACs [LQC10, RRC^{+}17]. tactics [MBvZ15]. Tagbanua [Cap10]. tagging [LPU^{+}16]. Taiarapu [TFBS19]. Taiwan [CLS10, Cha11, CLLS13, CL16a, Che10, Che11, Che12a, Che12b, Che12c, CCC^{+}14, CQ14, CC17b, HC10, HO10b, Hua11, HO12, LH12, LWJH11, Liu13, LCC14, LHH15, LKC15, LC19, LWS19, LSC14, LYB17, Shi13, SCC10, TOL12, TO10, WC17, YTSO15]. Taiwanese [LH16a, WC17]. take [AFCMET18, BRRG^{+}18, KBF12, MRG15, NKHL18, TM13, VAW14, KBGF12]. taken [Hua11, MSGM^{+}12, dBMA13]. takes [Yia18]. Taking [CPB19, JD16, NMvP19, SL15a]. tale [BB19b, BB19c, Hun13]. talk [MC19]. Tam [HRC15]. Tamil [CA17]. Taming [AL18]. Tangled [Bor10, Grö18]. tango [KHvH^{+}14, Yia18]. tanker [DZLL15, JSY^{+}19]. tankers [WZF^{+}18]. Tanzania [GLJdlTC14, Kam18, KMD15, Kat16, KRM14, Sha17, TCN^{+}13]. Tao [LC19]. target [CdLH^{+}18, Gib10, MB15, SPRH19, LGD^{+}19, RFL^{+}18]. targeted [BW18b]. Targeting [GLSA10, FC19a]. targets [BW18a, CG19, HdF^{+}19, Lon12, SLJ^{+}11, SFC^{+}15a, SS11, SWG18, TML^{+}19, VMO14, MSdAM16]. tariff [She16]. Tasmania [NBPF^{+}12]. Tasmanian [CKFLO19, RYJ^{+}17]. tax [Gek10, IHF15]. taxa [CBP19, JS18]. taxonomy [MGJB^{+}17]. teaching [FMD^{+}18]. Teasing [She15]. Tech [ZLL19]. Technical [SLS18, SV13, CVP^{+}17, EFK^{+}19, LBD^{+}18, MFP18, QSMGVA18, VPC19]. technique [DV17]. techniques [OCW^{+}17]. Techno [Nym19, VVS18]. techno-economic [VVS18]. Techno-optimism [Nym19]. technological [HHL^{+}18, SMS^{+}18, VS10]. technologies [RPSS17]. technology [Gil11b, GMP^{+}16, HD16a, KEM^{+}18, LZL19, MN18, PBSG12, Sea17b, SH15, ZZZL18]. TEDs [DBD19]. telemedicine [NPAR19]. temperate [HMGR14, PTN^{+}17]. temporal [BRR15, EGD18, FWK16, GM14, HAH^{+}16, HDTS14, LBHS16, MACM14, SRA15, SMS^{+}18]. temporary [GH15]. Ten [OGM^{+}14, RBJ^{+}11, SAH14]. tense [Gue19]. tenuous [Sch10b]. tenure [Cou11, PC18b]. term [CH12, CC18, DDS^{+}11, DCY^{+}19, HDG^{+}12, HGH^{+}18, LBHS16, Par10, SS11, SFDB18]. terminal [KMvT15]. termination [Árn17]. terrestrial [KFBG13, KJWP14, TE16, BL18]. territorial [NKBS17, THK^{+}18, ABCDMG17, QSHI17]. territories [ACR18, BAA^{+}15b, Cha17b, DBB^{+}14, FLWG11, KVMM10, MVRG13, RBC^{+}13, MCY^{+}19]. Territory [SSCC19, MPPB10]. test [LGM16, LEG19, MSdAM16, RWL14]. Testing [CH16, GNIL14]. Texas [Mat14]. text [Pas14]. TFC [LPM^{+}14]. Thailand [BD14, CLS16, NC10]. Thau [AYR^{+}14]. their [BJS^{+}16, BYK^{+}15b, Bro11, CPFS^{+}13, DFP15, DTM18, FFSF17a, FÓ12, GLTB^{+}15, GWM^{+}10, Jet18, JKS13, KGAB14, KP17, KR14, KVMM10, LEG19, MF19b, OP15a, PWH^{+}16, PSCY12, RPdO^{+}19, RBB^{+}16, SMGL17, Tec19, VRMF11, VGBG15, WP17, ZPT^{+}16]. them [Bro11, HRLFMP15, MBSL13]. theoretical [BHS19, VRA^{+}12]. theory [AOY16, COG^{+}19, IH18, JWM^{+}16, KPR^{+}18, LGFB^{+}19, ODF^{+}16, QSICTdV19, THRS15]. theory-based [QSICTdV19]. There [AGWMP19, CdLH^{+}18, CHL^{+}19, KGS^{+}19, MM15b, vHK17]. Thermal [JSG^{+}18, FMD^{+}12]. these [Her18b]. thing [MWR^{+}17]. think [MJvP^{+}18, RBC^{+}13]. thinking [CPFS^{+}13, FVT19, Gil14, GRW^{+}16, Rog15]. third [CMFG15]. third-party [CMFG15]. Thirty [CC17a]. though [CN11]. threat [BB16, GBK^{+}15, GOdFC16, Han14a, Mar17, RHP^{+}17, SVA^{+}12, TCCV16, VGCF18]. threatened [BAO^{+}12, Sch15a, SGW^{+}15, SNA^{+}17]. threatening [KHH13]. threats [BOF^{+}15, EAG18, FLWG11, GFC^{+}15, GCM11, RGB^{+}19, SSC^{+}15, SSJ17, Wea16]. Three [Han16, BAMAH17, BC13, DN12, DMdJ^{+}15, EJB19, HGS13, JGMS16, KM15, PAC^{+}19, SSWP10, WMK15]. three-country [KM15]. three-dimensional [BC13]. Threshold [SDC13]. thresholds [TML^{+}19]. thrive [PR13]. throughout [HHS19]. Thunnus [BB19b, BB19c, LLH^{+}14]. thynnus [LLH^{+}14]. tidal [ESM^{+}19, GOHDB17, LDS11, SM17a, San18, TE16, WS14, WRW^{+}17, WWM^{+}19, Wri16]. tide [HPH^{+}18, LHN18, NNPV17]. tides [RVdCGN11]. ties [vdVD16]. tiger [SMD14]. tilefish [TMHJ12]. Time [TTS14, AFR^{+}19, BDHSJ19, EB16, GOM17, KC15, LACG18, LM11, MHL^{+}15, MI15, Shi16, SPS14, Tam15, TMG^{+}18, WFK^{+}12b, ZP18, vPJH^{+}17]. time-series [ZP18]. timely [BWRT17]. times [NBFO18]. Timing [Tam15]. Timor [MTP^{+}17, PRPS17, RDVSC16]. Timor-Leste [MTP^{+}17, RDVSC16]. Tipping [THA15, SNDB^{+}16]. Today [AOY16, TLR^{+}10]. todisaster [MH13]. together [ABBD17, XJT18, dLS10]. Tomakomai [MKX17]. tomorrow [AOY16]. Tonga [HPW^{+}18, KH10, KH12, KHH13, TTM18a]. tonnage [Gek10]. Too [SBSK19, HPP^{+}17b, TN15]. tool [ASWM18, CVFP^{+}12, Coc15, Dru10, EE17, EB16, FSMG10, JHN^{+}19, KKSC19, MFB^{+}14, MHL13, RvLG^{+}15, SO13, SMBL13, SSB^{+}18, SKS^{+}19, SHM12, Syb15, TTP12, ZMM^{+}14, SKS^{+}19]. toolbox [CBL18, GSB^{+}17]. tools [BHS19, CH16, DGC^{+}17, De 18, GCMA^{+}12, GRW^{+}16, GAB^{+}17, LLO16, MJS16, PGB^{+}17, PGB19, SLS^{+}13]. toothfish [GRLJ15]. top [ACA^{+}13, Che10, Hey17, Hey18, LRS^{+}13, SC13a, SBC15a, GBF^{+}10]. top-down [ACA^{+}13, Che10, LRS^{+}13]. Tore [CLF^{+}19]. Toro [SJW^{+}16, SSM15]. Torres [Lal17, vPLB^{+}13]. Tortoise’ [JR14]. total [Bor18, MKSS18, PDS15, SAdC15, WHB^{+}18]. tour [FSS15]. Tourism [DM15, ADH18, ALM17, BLCD16, Che11, CT16, CC17b, CKBS14, DJS17, DHM^{+}17, FSB14, FFH18a, GH18, KEG12, KWT17, LTM19, MMTC16, PDT^{+}14, RNS19, RCMBG^{+}19, SBW^{+}17, SP19a, TK13, VMR^{+}18, VBMM17, Won19, ZKR^{+}19, dGB10, RV16b]. Tourist [HTD^{+}18, BJO19, KH10, KWT17]. touristic [RPG^{+}15]. tourists [BJO19, CS19, GLTB^{+}15, CT16]. Tournament [HSRH16]. tower [VvDM10]. traceability [BO12, He18]. TRACES [BB19a]. Tracing [TCN^{+}13]. track [GAB^{+}17, JD16]. tracked [DBB^{+}14]. Tracking [And15, VGA15, CPKF19, WSG18]. Tracks [Pat15]. Tradable [CL15]. Trade [ARS16, BB19b, BB19c, Cam16, SPRH19, WLT^{+}19a, ABBD17, BF15, BBM^{+}17, BTS^{+}17, BTS16, BB19a, CYI13, CBS^{+}16, DDV^{+}13, EAA^{+}18, FD11, FMB^{+}18, FW19, FKWV19, GRLJ15, He15, KKSC19, KV18, LGC16, MDE14, NBM15, Nij15, Nij17, OO19, OC17, PRT10, Pin13, PF12b, PKR^{+}17, SCV^{+}15, SMGL17, ST17, She16, SCMS^{+}18, TCN^{+}13, TBB^{+}10, VKA^{+}19, WCMHO18, WM17, WNLS17, You16a, dMAH^{+}18, STB16]. trade-off [DDV^{+}13]. Trade-offs [SPRH19, EAA^{+}18, FMB^{+}18, OO19, Pin13, SCMS^{+}18, WCMHO18, WLT^{+}19a]. tradeoffs [Lee10, LCH^{+}13, OLF19]. traders [WBC14]. Trading [RR17, GE17, ITNL^{+}14, PGC19, Ped13].

Traditional [BAMAH17, CN11, CC17b, GLVLGG19, Pom16, RYRYM17]. traditionally [TTS14, VPL^{+}13]. traffic [CMFDCLV10, CDR^{+}16, HTD^{+}18, LKL15, ZMZ16]. trafficking [SH19]. Tragedy [HBL14, LKC15, BMW17, OO19]. trailing [FSCJM19]. Training [Gle14, FFSF17b, SB14]. trait [SSBL^{+}15]. trait-based [SSBL^{+}15]. trans [NAYSN18, SDHS19]. trans-Arctic [SDHS19]. trans-Atlantic [NAYSN18]. Transaction [SS15c]. transactions [PRBX19]. Transboundary [JAO^{+}16, MLF^{+}17, SSS^{+}17, ACG^{+}17, ADHV16, FOO^{+}15, FMS17, GCG^{+}10, GCD^{+}11, HCBM15, Pla19, TKN19]. transdisciplinarity [GdV16]. Transdisciplinary [PIH^{+}10]. transfer [HD16a, MN18, Muk15]. transferability [CS13]. Transferable [CG13, LPM^{+}14, YCA^{+}12, BBG^{+}16, Bod17, DRH^{+}18, ESM^{+}19, EGGT12, EGHC16, FGW15, MNHN18, Mox12, MCBM18, Sha10, Sol14, SCS16, Tur10, MM18]. transformation [AMvT11, Car14]. transformative [BMAG17, CMMBV^{+}19]. Transforming [GW15, vBvTM16]. transhipment [SSW^{+}18]. transit [MGB^{+}18]. transition [CC17b, FGW15, ÖGB^{+}10, PW16, RD12, TBJC18, KEF18]. Transitional [HL15b, PS11a]. transitioning [GNN^{+}17]. transitions [Car15, Man17, SMH17]. Translating [COG^{+}19, JS18]. translation [MS19]. transmission [NAYSN18]. transnational [Fen17, Gla19, Tyn18, vAS19, HGG^{+}18, VWG19]. transpacific [PMM^{+}17]. Transparency [Ard18, ARJ18, BWD^{+}19, CAP15, Cor14, HdVB15, MBCR16, OHB^{+}18]. transparent [HTA^{+}15a]. transport [Bol17, GNN^{+}17, Næv17, NPS^{+}19, NNPV17, NNP^{+}14, VGA15]. transportation [Sea17a]. transshipment [ECSE^{+}17]. trap [SGP^{+}17, KLY14]. traps [Uhr16]. trash [HPH^{+}18]. Trauma [AT14]. travel [CACL18]. Trawl [RWG13, AAL^{+}18, BSKH17, Cha16, Cho12, DMdJ^{+}15, DBD19, FPP^{+}15, FFKO15, FV10, GMM15, GOM17, HNVZ12, KBBR16, MC15, MMdJ^{+}14, PC18a, PMC^{+}15, SBMD13, SMR^{+}15, SKT^{+}10, VPC19, VKA^{+}19, WTD^{+}15, WSG18, ZH14]. trawler [MUHH18, Mul19]. trawlers [HID12, MMN18]. Trawling [TKAO^{+}19, DMdJ^{+}15, ECF^{+}18, FTG^{+}18, KB15, LBVN19, MBD13, MMDS12]. trawls [JL11, Pat15]. treaties [AAGMH17, HL15b, RM16]. treatment [DMC^{+}17, FSV^{+}16]. Treaty [Bes11, Sco18, HL15b, Lil16, MDNT19, TDMN19]. tree [SM18b]. trend [Gar12]. Trends [SPMW12, BSL19, CP13, FSM^{+}12, GLW17, GZW^{+}19a, GZW^{+}19b, HGH^{+}18, KH13, LB17, MACM14, MMN18, MMC^{+}14a, MMC^{+}14b, Nag16, Nym19, PDMÖ19, PC18a, PJS13, ZEP^{+}19]. trial [MA15]. Trialling [JHN^{+}19]. Triangle [FAA^{+}13, HBZH16, RO13, BTOM13, DRG^{+}16, RvLG^{+}15, TTJ18, VSRD16, vLvT10, CCTG^{+}13, CF17, FEF^{+}12, FE12, GMP^{+}16, PPR^{+}15]. tribally [GRYA^{+}15]. tribally-led [GRYA^{+}15]. tribunals [Goo17]. Tributyltin [GOdFC16]. Trieste [CS14]. trinational [PMM^{+}17]. Trindade [dOSL18]. triple [CdLH^{+}18]. triumvirate [Sch10b]. troll [RLH18]. Trondheimsfjord [TMR^{+}14]. trophic [ET19, HD16b, vOHO^{+}17]. trophy [SGW^{+}15, Sch15a].

Tropical [FV10, MBD13, Uhr16, ACM^{+}12, CBG14, FSWFG13, GvdB13, GFC^{+}15, KVA18, Lal17, Nij15, PSV^{+}14, PGPS10, SBC^{+}14, WHMG^{+}16, vPLB^{+}13]. Troubled [HMH18, GGLP16]. troubles [PS18, SP19b]. true [SvZdJ15]. Trust [GTM^{+}12, dVvT11, CvPH^{+}18, EBR^{+}18, GC17, GDD11, Han19, KMF^{+}18, MSTB19, Mat16, WS15b, dVM10b, vdVD16, CS19]. truth [AGWMP19, LOD^{+}15, SJ19b]. try [Vra19]. trying [BHAHW17]. tsunami [MGA^{+}10]. Tubbataha [DSWG13]. Tuna [Par10, XDG^{+}16, AMdB12, AHB18, Bar10, BAA^{+}15b, BB19b, BB19c, BHT^{+}18, BBvZ^{+}17, Cam16, CCS14, DMMG15, DP17, DPB17, Dru10, DK16, EYN^{+}15, FBAK14, dHJTG17, Gil11b, GOK14, Hav10, Hav13, HSB^{+}16, JTK18, KGMP18, KVH14, LLH^{+}14, LPU^{+}16, LHN18, MBvZ15, MBM14, MHM16, NIFA17, NJ18, PSV^{+}14, STM16, SDB14, SCOS17, SH14c, TOL12, TZBBL18, WM19, Web11, YBRBB16, Pol12a, SH12, dBMA13]. Tunisia [DV17, HAH^{+}16]. turbines [WS14]. TURFs [ABCDMG17, QSHI17, Yos17]. Turkey [EA19]. Turks [RNS19]. turmoil [LGH^{+}12a, LGH^{+}12b]. turnaround [LJ13]. Turning [ADH18, HPH^{+}18, NNPV17, HHG10, Iri18]. Tursiops [FSS17]. turtle [BKKN10, BRM^{+}17, DBD19, GCMA^{+}12, MBSL13, RBC^{+}13]. turtles [ATB^{+}18, BHG11, HGS13, MRCG11, SRGD18, TTJ18]. Tuvalu [Bro18]. Twelfth [Rei10]. twenty [BSD^{+}17]. two [BB19b, BB19c, Car11, Dup14, FJA^{+}14, GFA^{+}18, Gus18b, Hea13, HL15b, KR18, KBBR16, KH15, Man17, MSdAM16, Pin15, PA10, RGSF19, WBC14, WTA^{+}16, vHCT^{+}16, BGS13, HGS12]. type [FSCJM19, HSB^{+}16]. types [RAZ16]. typologies [KPR^{+}17]. typology [AVH10, BM15, CBS^{+}16, JKK13]. typus [BFTS^{+}16].

U.K. [Gau10, SS16]. U.S. [BK16, BPMY18, BGC^{+}17, BRM^{+}17, Bur15, CKdS14, Das16, DL19a, DFSB15, DCY^{+}19, HL15a, HGH^{+}18, HGB^{+}19, KS12a, KS12b, Kai15b, Kai16, Kai17a, Kai17b, Kai17c, KBG^{+}18, Knu17, LPC18, MC15, MCB^{+}19, PB13, SG15b, SM18c, SGMWE14, TL16, TWA^{+}15, TBB^{+}10, TMKK^{+}15, WLT^{+}19b]. ubiquitous [Far19]. UK [BFGR14a, Tec19, BFGR14b, De 16, DHR16, FLWG11, Gek10, GFA^{+}18, HHP17, HA14, HAA15, HHA17, HBC^{+}13, HDTS14, JBdL^{+}14, Kel18, MM17a, MMOL15, MPCB14, MBB18, MCY^{+}19, McN18, PGL^{+}16, RBC^{+}13, RU12, RMJ12, RLHM14, RFG^{+}14b, SFG15, Smi13, SGLY13, SPP18, SP19b, TÁB^{+}10, WML17, WPRA18, WCK^{+}17]. Un-gendering [GPH18]. unacceptable [BRR18a]. Unbalanced [MHF13, KVA18]. uncertain [MMOL15]. uncertainties [Kno11, LLH^{+}14, OCCvT^{+}18]. Uncertainty [WHB^{+}18, DMMG15, EG15, FBAK14, HHL^{+}18, JFA17, NRB^{+}18, San10, SFC^{+}15b, WMJF16, XDG^{+}16, BHA^{+}14]. UNCLOS [Has18, Mar17]. unconstrained [LSMW10]. unconventional [PGP18]. Uncovering [SKL16]. uncovers [KB15]. under-10m [GKSJ11]. undermine [HS19, PDC19, SH17, TCS^{+}18]. underpinning [DDS^{+}11, HRJH15]. underpins [HTA^{+}15b]. underrated [CC18]. underrepresent [CBP19]. understand [FVT19, MLHP15, TCN^{+}13]. Understanding [AMdlCMS^{+}18, CP10, MSZS^{+}11, NMvPNL15, Rei19, RGWH18, SCV^{+}17, AWJ19, FTF^{+}17, HMGR14, KPF17, MB10, Muk15, RvLG^{+}15, She15, VGBG15, dlTCFB^{+}17, Ben12, DJ15, MF10, RCUR13, SAM13]. understandings [BK16]. underwater [JL11, LU13, LZ16, Dro13]. undesirable [TJ13]. undulatus [PF12b]. uneconomic [SvZdJ15]. unequal [ODRRM12]. UNESCO [Dro13]. unfamiliar [ZCD^{+}15]. UNFCCC [HK11]. Unfinished [Che12c]. unfolding [VDB16]. Unguja [BTS^{+}17]. Ungwana [TKAO^{+}19]. unhelpful [Bur16]. uniform [GS16]. unilateral [Muk15]. unilaterally [SJP17]. Unintended [MG19, CCS14, CSCBR10, DJS17, GLR17, Sat18]. Union [Ros17, vLvHvT12, PL12, Ant16, CGDB13, DRGP19, De 11, GSR19, GGD16, GCW^{+}18, GAC^{+}19, Had18, KFPR10, LP16, MdMGV^{+}18, MM15b, PW16, RWG^{+}19, SSFL10, SRGGVL13, Tei18, Tyn17, Vil10, dVMdC^{+}19, Weg11, ZPB^{+}13]. unique [RL14, TO10]. unit [IT15]. United [GLW17, Gla19, GBF^{+}12, HS19, KS12c, MPF12, PDC19, SHS^{+}14, Asa12, Ásg16, CJW^{+}16, DSH14, DG14, FDP19, GDN19, HPP^{+}17b, Iri18, LM11, Mil14, PMN^{+}15, RFG^{+}13, RFL^{+}18, SH17, Smi17, YCA^{+}12]. Units [MHM16]. University [Rei10]. Unjust [DR14, SvZdJ15]. Unpacking [MV16]. Unravelling [BJvH18, MFAA16, Had18]. unrealistic [KFBG13]. unregulated [BO12, LT17, MBM14, OY19, PF12b, PRPS17, Ros17, SDP^{+}16, VAP^{+}10, You16a]. Unreported [LGH^{+}12a, LGH^{+}12b, BO12, GPM19, LT17, MBM14, OY19, PF12b, PNPD14, PPM17, PPM19, VAP^{+}10, You16a]. unsafe [NPS^{+}19]. unsustainable [DR14, SAH14]. Untangling [NMvP19]. unveils [BHM16]. unwanted [LKC15]. update [MAAB15, Vil10]. Updated [SES^{+}19, BCP19, SLL^{+}16]. upgrade [VRB14]. Upgrading [Bur16, HHS16, ARSHH16, Lim16]. upon [MG19, MLA16, Mor17]. urban [CSP^{+}14, SRGGVL14, WS15a]. urchin [ONF15]. urea [MR10]. Urgent [Cha15]. Uruguay [PCCR15, GMD15, GHM^{+}18, TJ13]. US$50 [Rah19]. USA [BB19b, BB19c, BSP19, BHG11, JBV^{+}18, PTDH17, PNPD14, SSWP10]. usage [OIS^{+}18]. Use [ABCDMG17, Coc15, MBTV14, MKSS18, QSHI17, AAC^{+}13, ARGW14, APD^{+}19, BBM^{+}17, BAA^{+}15b, BAA^{+}15a, BYK^{+}15b, BDP^{+}17, BJvH18, BHS19, CT16, DMMG14, DWCS18, DJ15, DT10, GCK^{+}11, HF16, HGS13, HFM^{+}13, HHA17, HBC^{+}13, JHN^{+}19, KHH13, Kim19, KRM14, KED^{+}18, KHP14, KVMM10, LLF^{+}16, LGM16, Lin13, MMM19, MTP^{+}15, MMJM15, MMJM16, MS14, MHM16, NKBS17, PTN^{+}17, PM10, PV16, RWL14, RU12, RUH^{+}14, San18, Sco16, Shi13, SWM11, SV15, TvN13, THK^{+}18, VRB14, Vir18, WRW^{+}17, WHMG^{+}16, vHFJ^{+}19, vdBAMJT19, JBPBA16]. use-use [BHS19]. use/user [LLA^{+}19]. used [Lon11, Pas14, RMJ12]. User [Thu10, CSS15, CHL^{+}19, DHR16, LLA^{+}19]. users [CST10, GW15, KRM14, MMJM15, MSB^{+}18, MPCB14, NHA15, VGBG15, PGB19]. uses [PTDH17, RFHEJK13, RFKEJ^{+}15]. Using [CA16, DJTK17, FMD^{+}18, GGSRVL17, GBK^{+}15, HIN^{+}17, HVB^{+}17, KFB19, LGM16, LR15, MHL13, MF10, RWK15, SLJ^{+}11, SBTF15, SPO^{+}18, Sol14, SMKBM16, TTM18b, WS14, WK17, WML^{+}16, AVC18, ARGW14, ATDP16, BAA10, BH14a, BC13, BA19, Cha14, CL16a, DPJ^{+}13, DDP16, DMJ^{+}17, GS17, GZW^{+}19b, GW11, HF16, HNVZ12, HBZH16, HTH13, JL11, KYS10, Kai16, LSP^{+}17, LH12, Liu13, MMP15, MSdAM16, Mor17, NBM15, NHPD19, OCG14, RBW18, RJPS11, Shi13, SCOS17, TRM19, TRC^{+}17, WNB^{+}15, WTA^{+}16, YYN^{+}15, YSJU^{+}17, ZCD^{+}15, dlFRG17, Cro15, GOO^{+}19, LGDB17, PL15, PA10, SGN^{+}17, WCCO18]. usual [Cla15]. utilisation [Bes10, SNTL18]. utility [MDM16, Rei16, VSCH16, VRB14, WSO^{+}14]. utilization [BAHCG14, CHDJ^{+}19, HO12]. utilizing [KLN17, KEM^{+}18].

V [Cho10, Gla19, KFM17]. Validating [AARK15]. validation [WHB^{+}18]. Valuation [PSCY12, ŠST^{+}18, BBP^{+}14, DV17, LKL15, MTP^{+}15, RCF^{+}12, TH17, YWC^{+}16]. Value [Lim16, RPC^{+}17, BvRH11, BD19, CPKF19, Cha17a, CP15, CF16, CLC^{+}16, DPB17, DDV^{+}13, EG15, Fab16a, FDP18, FGW17, FSB14, FFH18b, HHS16, HH10, IH18, JHW13, JLY18, KKG16, MM17a, MGJB^{+}17, MOH11, MDM16, NJ18, NAYSN18, PTML16, PCK^{+}19, PMA^{+}17, PRS10, PAA^{+}12, PCLE17, RRA^{+}10, RUD^{+}12, RV16b, SDP^{+}16, SGLY13, Son15, SBW^{+}17, SNTL18, SL16, TCN^{+}13, TTFC19, VGL^{+}17, WSC19]. value-chain [PCLE17]. value-destructive [BD19]. Values [JRS^{+}16, SCJ13, LJ17, MMP15, MD15, MDM16, OL10, RLM10, RLM12, SLH^{+}13, VGBG15]. Valuing [CdlPS^{+}14, HTH13, MAR^{+}15, MAHP16, OL10, PWH^{+}16, ZCD^{+}15]. Vanuatu [LBK^{+}13, RDVSC16]. vaquita [AFMCMP12]. variability [BAHD10, CLLS13, IHS13, ISHP14, KNA10, WRW^{+}17, ZFG12]. variation [SKFL15, SRA15]. varies [WHB^{+}18]. varieties [SCOS17]. various [WC17]. vector [DMC^{+}17]. vehicle [FSS15, HDK^{+}15, KR18]. vent [Joh19, TTFC19, VAHG^{+}18]. VentBase [CKC^{+}13]. vents [MV19]. Venue [Pri10]. Verde [DB19a]. verrucosa [PGL^{+}16]. versa [BVB18]. versions [Rog15]. versus [BDF^{+}10, CG19, CE13, EA19, HGS13, KJP14, Mox12, PRBX19]. vertically [HID12]. very [JD16]. Vessel [KSC^{+}19, ZM15, BW19, CSVHS14, CLS10, Cha11, CDR^{+}16, CJ16, DWCS18, GM14, HTD^{+}18, HSW17, Hut19, Kai16, KH11, KFM17, LKL15, PAC^{+}19, RRG^{+}12, RGBT16, RBS^{+}19, SKFL15, SVA^{+}12, SDHS19, TKAO^{+}19, VGA15, ZH14, BSSS17, Hav10, Hav13, Kai15b]. Vessels [HDH^{+}15, BSP19, CW15, CL13, DBG^{+}16, GCP19, JLY18, Lin13, TS14, VPC19, VRMF12]. vexillum [SHS^{+}14]. via [CMFG15, GKG16, OLF19]. viability [GFH^{+}17, LR15, SCC^{+}17]. viable [MH13]. vice [BVB18]. Vicentina [SFCI14]. victims [FW15]. Victoria [CN17]. Viet [HvD13]. Vietnam [AMvT11, HRC15, Ngo18, NJ18, PHC14, VPC19]. Vietnamese [BD10, DFL15, SD18]. View [Yat14, BPBM17, FDP19, PT10]. Viewpoint [AGWMP19, SJ19b, JD16]. views [BHA^{+}14, GCHO19, HMGR14, HJKR14, JBV^{+}18, RCMBG^{+}19]. Vigo [SRGGVL14]. village [CD15, GSJ18]. village-level [CD15]. villages [BAMAH17, Lan15]. violacea [FSCJM19]. violations [MPMK16]. Virginia [Rei10, OTOL17]. visibility [GCD^{+}15]. visible [ZTAO13]. vision [Pau18]. visions [PS11b]. Visitors [MDK^{+}16, RPG^{+}14]. visits [EWG^{+}18, WWH^{+}17]. visual [ZCD^{+}15]. visualisation [LBL^{+}17]. Visualising [Nag16]. visualizations [RWL14]. Vitória [dOSL18]. VLMPAs [JD16]. VMS [CSVHS14, MMdJ^{+}14, TKAO^{+}19]. voice [CCR16, SBC^{+}15b]. voices [AWJ19, MAJM^{+}18]. volitans [CAGH16]. Volume [BB19b, BD19]. voluntary [BSKH17, WGB^{+}16, CPB19, SALS17, Che10]. voyage [Cha16]. voyaging [LK15]. vs [TZBBL18, WP13, dVM12]. VTS [LKL15]. vulgaris [SGP^{+}17, DYN^{+}19, VRMF12]. vulnerabilities [BLC16, RRC^{+}17]. Vulnerability [DCHC17, RAZ16, AYR^{+}14, CJW^{+}16, CPBQ19, FSGB^{+}15, JWBJ13, JWM^{+}16, KBK19, LCR16, MLTBMB14, SG17, TMG^{+}18, TLA16, THE^{+}18, TLA18, TMDSF19, WMJF16]. vulnerable [ACP^{+}14, HA13, MCB^{+}19, MSGM^{+}12, dMAH^{+}18, JL11].

W [McL15]. Wadden [SGRV16]. Waitangi [Bes11]. Wakatobi [Cli13]. wake [WBH^{+}13]. Wales [BS10, MBB18, SHS^{+}14, TSW19]. Wall [SHM13]. wander [BSW19]. want [KR14]. War [BG16, SBM^{+}17, Zha16]. warming [REA^{+}14]. warnings [EOC15]. wars [LBB^{+}12]. was [OHH^{+}16]. Washington [FWK16]. waste [JHB^{+}18, PAC^{+}19, SOS12, WMWH18, dlFRG17]. wastes [WBA18]. wastewater [Bre14]. watching [AFGRF^{+}17, Bai12, CCC17, Che11, CCD^{+}13, CMSKP10, KH13, Lee10, SJW^{+}16]. Water [LCC14, APD^{+}19, BSKH17, CLS10, FSV^{+}16, GCK^{+}11, HWN11, LWS19, MAR^{+}15, Mal13, MFF^{+}18, MSG11, NWW^{+}15, NWW^{+}17, NYJW15, RRCS^{+}19, SFC^{+}15a, SDLT15, SLDHS18, TOL12, TML^{+}19, VH16, WCS18, ZT19, vH19, HWN11]. waterline [LGDB17]. waters [BAA10, BBN^{+}13, Bes19, Bro11, Cap10, CGP^{+}18, Cho12, Cro11, CJHR14, DJTK17, EFK^{+}19, FWK16, FGKMP14, FDS15, GRLM15, GGLP16, HL15a, HMH18, HTH13, JBPBA16, KP13, KLY14, MCHO11, MPMK16, MAAB15, NALO19, PGL^{+}16, PVS^{+}19, SFC^{+}15a, STG11, SS16, SSA16, Zha16, HGH^{+}18, JKS16, LMA^{+}15]. waterways [JRS^{+}16]. wave [LDS11, LGM16, MM19, Rei16, WSGID17, Wri16]. way [ACM^{+}12, CBN^{+}12, CDA^{+}13, CT16, GJ15, San15, Sch15a, TJ13, ZPB^{+}13]. ways [And15, Gus18b, LCC19]. weakly [WGB^{+}16]. Weaknesses [HBK17, FGKMP14, NACS17, RGB^{+}19, SSC^{+}15]. Wealth [SW10, VKGN^{+}14a, ZS18]. Wealth-based [SW10]. wealthy [YBD^{+}19]. weather [PTDH17]. web [Fol13, Grö18]. webs [KMTS16]. Weh [KH15]. welfare [BR13b, BR14, DHSC14, EZPRZ18, HDJD15, NSB18b, PLL17, PBJ^{+}14, RAF14]. welfare-based [RAF14]. well [AYA^{+}19, BV19, Car15, HHP17, HM18, MD16, Ngo18, PRS^{+}19, RTD^{+}17, WPRA18]. well-being [AYA^{+}19, Car15, HHP17, HM18, MD16, Ngo18, PRS^{+}19, WPRA18]. well-regulated [BV19]. wellbeing [BSW16, BC13, GC18, Kel18, RFG^{+}14a, TJ13]. Welsh [MB18a, TLB19]. were [LCC^{+}18]. West [Cro15, DL19a, KL12, KBG^{+}18, MC15, MRSH13, MCB^{+}19, HR17, HK18, LB13, PRBX19, RROR12, BSD^{+}17, BDvTvKT17, BLC16, DSH14, MBM14, NW10, RLH18, SWB^{+}14, TGBL11, VKA^{+}19, WSG18]. Western [LMA^{+}15, CdLR^{+}15, CdLH^{+}18, EYN^{+}15, GFC^{+}15, HTD^{+}18, Han12, PGC16, RCdLS13, SKFL15, Sev14, All14, CdLR^{+}15, GW15, Gil11a, GOM17, GJA16, HR17, Hav10, Hav13, HCP^{+}17, Mac14, Man10, Nor15, RCdLS13, RTM^{+}19b, SDB14, SCS14, WFK^{+}12a, YBRBB16]. wetland [GOHDB17]. wetland-birds [GOHDB17]. wetlands [KNB^{+}18, SBC^{+}14]. Whale [Bai12, CCC17, KH13, LTM19, PC13b, Asa12, AFGRF^{+}17, BFTS^{+}16, BJS^{+}16, BGC^{+}17, CN11, Che11, CCD^{+}13, CMSKP10, GJ15, HTD^{+}18, HD11a, HD11b, Lee10, MPT13, MMB^{+}19, Sin19, SJW^{+}16, TJ18, WTA^{+}16, ZVP^{+}13]. Whale-watching [CCC17, CCD^{+}13, Lee10, SJW^{+}16]. whale/dolphin [Che11]. whales [BWRT17, BSP19, DB19b, GCP19, KH10, KH12, KHH13, RRG^{+}12, RGBT16, SCKP19, SVA^{+}12]. Whaling [Bai12, Ili10a, Ili10b, VBHC^{+}17, Cla15, CF16, WSWM17, Hea13]. Whalsay [CG13]. wheelhouse [Yat14]. Where [DES10, PML10, KGS^{+}19, OBJ^{+}12, Cha17a, EWG^{+}18]. Whether [CYI13]. Which [BFGR14a, BFGR14b, LPS12]. whiffiagonis [MACM14]. while [RTM^{+}19b, vD16]. white [ADH18, LHH15]. Whitefish [LBB^{+}12, MM17a, MM14]. Whither [TMKK^{+}15]. Whiting [SCS14, Vil12]. Who [Don17, Kru16, LAH^{+}13, PDS15, PPOM16, WCCO18, BSW19, EWG^{+}18, GSW18, PdSA^{+}13, RLAT19, FPG^{+}17]. whom [KW18]. Whose [GBF^{+}10, De 10]. wide [MVRG13, SSCC19]. Wider [FMM13]. Wild [LOS11, MC19, MAHP16, PS11b]. wild-catch [MC19]. wildlife [ADH18, BLCD16, EMP^{+}14, Gue19, TBB^{+}10, WWH^{+}17, WML^{+}16, WTA^{+}16, BLCD16]. Will [CJ16, SCOS17, SG12, CPR14, GC17, MHK^{+}10, Pan14]. willing [CS19]. willingness [BBE^{+}16, CL18, DV17, LHN18, TCCV16, WHMG^{+}16, ZMZ18, vOHO^{+}17]. willingness-and [WHMG^{+}16]. Wind [RPSS17, BDF^{+}10, BN19, GSG^{+}15, HL15a, HHP17, HDJD15, HAA15, HHA17, KS12a, KS12b, KS12c, Man17, MMOL15, RDP^{+}18, RRMSSE16, SS19a, Sea17b, VLPC14, VHA^{+}19, VLKN17, WKB15, ZZZL18, vdBKB^{+}17]. windfarms [AMR14, HA14]. winter [LG19]. within [BJGMG18, BSAT18, CR10, Car19, CSS15, CDR^{+}16, FBP^{+}18, JA10, MPF12, MV16, MK15, MS14, PGP^{+}19, PKM^{+}19, RMRN^{+}16a, RGGFA17, RL19, SRGD18, VDB16, VRT13, VBHC^{+}17, YFR10, dVM10b]. without [FSS17, SM17b]. wolffish [DN12]. Wolves [Gue19]. Women [HZH^{+}13, ZTAO13, CCR16, GPH18, KJH^{+}10, MAJM^{+}18, RB15, ZZZ^{+}17]. words [HSRH16, Pas14]. Work [THRS15, Bac18, ERB14, NPS^{+}19, SFDB18, TLR^{+}10, VSCH16]. workable [WC17]. workers [ZTAO13]. workflows [KS12c]. Working [CAGH16, LZ16, Næv17, OSS14, XJT18, EGP^{+}18]. works [LBK^{+}13, TKM^{+}18]. workshop [CKC^{+}13, FMP^{+}18, KBB^{+}15, NWW^{+}15, NWW^{+}17, TDI^{+}14]. workshops [LHJ^{+}15, LLO16]. World [Cla15, HMS^{+}17, PZ17b, WS15a, EZBL19, FG18, GAP12, Jac15, MKR12, MMOL15, Pin17, PVPF19, PS14, Sch10a, SF13, WB19, dVM10a, Fen17, JSG^{+}18, LJ17]. Worldviews [Deb18]. worldwide [Cha17b, KP17]. worst [HBA^{+}14]. worst-case [HBA^{+}14]. worth [BHAHW17, DPV^{+}15, GWM^{+}10, SDC13]. would [HS19, PDC19, SH17, HD16b]. wrasse [PF12b]. wrong [HHG10, JD16]. WTO [KKSC19]. WTP [MDK^{+}16].

Xiamen [HCJ15].

Yantai [PQLH15]. Ye [PZ17a]. year [BSD^{+}17, CSF^{+}18, CCS16, FMD^{+}18, KJH^{+}10, VGA15]. years [CC17a, FH15, Her18b, HdV17, KR14, MMJM16, Val13]. yellow [GMD15, Kim19]. yellowfin [BB19b, BB19c]. yield [CdLR^{+}15, GSC^{+}16, HCP^{+}17, HC18, Mes12, NLP11, PL15, RCdLS13, SS11, SAH14, Hil10]. Yilan [LSC14]. York [BSP19, LU13]. Youth [GLW15]. Yucatan [PS11a, Ben17b]. Yukon [Jen15].

Zalophus [CJB15]. Zambian [TCC^{+}19]. Zamboanga [RNN^{+}18]. Zanzibar [BTS^{+}17, CNFJ10, GLJdlTC14, TCN^{+}13]. Zealand [Bes11, LLF^{+}16, YFR10, YFC13, YCHM14, BS10, Bes10, Bes12, Bod17, Bor10, CKBS14, DMKL18, Edd14, ECRL17, FHJ16, Her18b, JY19, MM18, McC17, RNCGB15, RNEBG19, SC11, SL15b, SL16, SM17b, SH19, TMG15, Tor16]. Zealand’s [LLA^{+}19]. Zeller [YBB^{+}17]. zero [VPP^{+}16, Kno11]. Zhejiang [ZLL19]. Ziegler [GZW^{+}19b]. Zone [BH14a, Coo11, De 16, GSDB14, LJL^{+}14, Rah19, RSS14, TBB12, BPB19, CDO16, DTN10, DTMJH16, KHP14, LSMW10, MFB^{+}14, MTP^{+}15, MMDS12, PCCR15, SWS12, TE16, WM18b, LPP17, LJ16, Mul19, RBS^{+}19, TN15, VVS18]. Zones [LLS^{+}15, AFCMET18, BRRG^{+}18, MJQMVÁ17, MCH19, TM13, VAW14, MPCB14, MLA16, SHM12]. Zoning [HF16, LSC14, BLCD16, DSM^{+}13, GAW^{+}13, LLS^{+}15, LSTD12, LLX^{+}15]. zooplankton [AAL^{+}18].

References

Alam:2017:SAB

[AA17] Shawkat Alam and Saif Uddin Ahammad. Shrimp aquaculture in Bangladesh: Domestic regulatory responses and compliance with MEAs. Marine Policy, 82(??):122–133, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300210.

Ahtiainen:2013:PPR

[AAC^{+}13] Heini Ahtiainen, Janne Artell, Mikołaj Czajkowski, Berit Hasler, Linus Hasselström, Kari Hyytiäinen, Jürgen Meyerhoff, James C. R. Smart, Tore Söderqvist, Katrin Zimmer, Julia Khaleeva, Olga Rastrigina, and Heidi Tuhkanen. Public preferences regarding use and condition of the Baltic Sea — an international comparison informing marine policy. Marine Policy, 42(??):20–30, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000201.

Al-Abdulrazzak:2017:OIG

[AAGMH17] Dalal Al-Abdulrazzak, Grantly R. Galland, Loren Mc-Clenachan, and John Hocevar. Opportunities for improving global marine conservation through multilateral treaties. Marine Policy, 86(??):247–252, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300118.

Aswani:2017:OSD

[AAL17] Shankar Aswani, Simon Albert, and Mark Love. One size does not fit all: Critical insights for effective communitybased resource management in Melanesia. Marine Policy, 81(??):381–391, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303979.

Aubert:2018:NMR

[AAL^{+}18] A. Aubert, E. Antajan, C. Lynam, S. Pitois, A. Pliru, S. Vaz, and D. Thibault. No more reason for ignoring gelatinous zooplankton in ecosystem assessment and marine management: Concrete cost-effective methodology during routine fishery trawl surveys. Marine Policy, 89 (??):100–108, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305018.

Adiga:2015:VRS

[AARK15] M. Suresha Adiga, P. S. Ananthan, V. Ramasubramanian, and H. V. Divya Kumari. Validating RAPFISH sustainability indicators: Focus on multi-disciplinary aspects of Indian marine fisheries. Marine Policy, 60(??):202–207, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002006.

Al-Abdulrazzak:2012:CLP

[AAT12] Dalal Al-Abdulrazzak and Stephen C. Trombulak. Classifying levels of protection in Marine Protected Areas. Marine Policy, 36(3):576–582, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001680.

Aanesen:2012:CEF

[AAvH12] Margrethe Aanesen, Claire W. Armstrong, and Luc van Hoof. The changing environment of fisheries policy in Europe. Marine Policy, 36(5):1172–1177, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000085.

Asche:2014:DFF

[ABB14] Frank Asche, Marianne Tranberg Bjørndal, and Trond Bjørndal. Development in fleet fishing capacity in rights based fisheries. Marine Policy, 44(??):166–171, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001723.

Allen:2017:STS

[ABBD17] Maggie Allen, Stoney Bird, Sara Breslow, and Nives Dolšak. Stronger together: Strategies to protect local sovereignty, ecosystems, and place-based communities from the global fossil fuel trade. Marine Policy, 80(??):168–176, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306704.

Aceves-Bueno:2017:TUR

[ABCDMG17] Eréndira Aceves-Bueno, Jorge Cornejo-Donoso, Steve J. Miller, and Steven D. Gaines. Are Territorial Use Rights in Fisheries (TURFs) sufficiently large? Marine Policy, 78(??):189–195, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307229.

Al-Belushi:2015:DMB

[ABSB15] Kawther I. A. Al-Belushi, Selina M. Stead, and J. Grant Burgess. The development of marine biotechnology in Oman: Potential for capacity building through open innovation. Marine Policy, 57(??):147–157, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000500.

Al-Belushi:2018:MOI

[ABSGB18] Kawther I. A. Al-Belushi, Selina M. Stead, Tim Gray, and J. Grant Burgess. Measurement of open innovation in the marine biotechnology sector in Oman. Marine Policy, 98(??):164–173, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306796.

Arceo:2013:MBT

[ACA^{+}13] Hazel O. Arceo, Bertrand Cazalet, Porfirio M. Aliño, Luisa Mangialajo, and Patrice Francour. Moving beyond a top-down fisheries management approach in the northwestern Mediterranean: Some lessons from the Philippines. Marine Policy, 39(??):29–42, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002047.

Authier:2017:CCM

[ACG^{+}17] Matthieu Authier, Florence Descroix Commanducci, Tilen Genov, Draško Holcer, Vincent Ridoux, Maÿlis Salivas, M. Begoña Santos, and Jérôme Spitz. Cetacean conservation in the Mediterranean and Black Seas: Fostering transboundary collaboration through the European Marine Strategy Framework Directive. Marine Policy, 82 (??):98–103, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307321.

Aswani:2012:WFE

[ACM^{+}12] Shankar Aswani, Patrick Christie, Nyawira A. Muthiga, Robin Mahon, Jurgenne H. Primavera, Lori A. Cramer, Edward B. Barbier, Elise F. Granek, Chris J. Kennedy, Eric Wolanski, and Sally Hacker. The way forward with ecosystem-based management in tropical contexts: Reconciling with existing management systems. Marine Policy, 36(1):1–10, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000509.

Ardron:2014:SAT

[ACP^{+}14] Jeff A. Ardron, Malcolm R. Clark, Andrew J. Penney, Thomas F. Hourigan, Ashley A. Rowden, Piers K. Dunstan, Les Watling, Timothy M. Shank, Di M. Tracey, Mathew R. Dunn, and Steven J. Parker. A systematic approach towards the identification and protection of vulnerable marine ecosystems. Marine Policy, 49(??): 146–154, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002765.

Asch:2018:FME

[ACR18] Rebecca G. Asch, William W. L. Cheung, and Gabriel Reygondeau. Future marine ecosystem drivers, biodiversity, and fisheries maximum catch potential in Pacific Island countries and territories under climate change. Marine Policy, 88(??):285–294, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301409.

Alder:2010:APM

[ACSK^{+}10] Jacqueline Alder, Sarika Cullis-Suzuki, Vasiliki Karpouzi, Kristin Kaschner, Suzanne Mondoux, Wilf Swartz, Pablo Trujillo, Reg Watson, and Daniel Pauly. Aggregate performance in managing marine ecosystems of 53 maritime countries. Marine Policy, 34(3):468–476, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001614.

Ahmed:2017:SBC

[ACTG17] Nesar Ahmed, William W. L. Cheung, Shirley Thompson, and Marion Glaser. Solutions to blue carbon emissions: Shrimp cultivation, mangrove deforestation and climate change in coastal Bangladesh. Marine Policy, 82(??):68–75, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302427.

Amigo-Dobano:2012:PFM

[ADGGVL12] Lucy Amigo-Dobaño, M. Dolores Garza-Gil, and Manuel Varela-Lafuente. The perceptions of fisheries management options by Spain’s Atlantic fishermen. Marine Policy, 36(5):1105–1111, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000322.

Apps:2018:TWE

[ADH18] Kirin Apps, Kay Dimmock, and Charlie Huveneers. Turning wildlife experiences into conservation action: Can white shark cage-dive tourism influence conservation behaviour? Marine Policy, 88(??):108–115, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304839.

Azmi:2016:DDB

[ADHV16] Kamal Azmi, Ruth Davis, Quentin Hanich, and Adam Vrahnos. Defining a disproportionate burden in transboundary fisheries: Lessons from international law. Marine Policy, 70(??):164–173, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302871.

Auge:2018:FMK

[ADL^{+}18] Amélie A. Augé, Maria P. Dias, Ben Lascelles, Alastair M. M. Baylis, Andy Black, P. Dee Boersma, Paulo Catry, Sarah Crofts, Filippo Galimberti, Jose Pedro Granadeiro, April Hedd, Katrin Ludynia, Juan F. Masello, William Montevecchi, Richard A. Phillips, Klemens Pütz, Petra Quillfeldt, Ginger A. Rebstock, Simona Sanvito, Iain J. Staniland, Andrew Stanworth, Dave Thompson, Megan Tierney, Philip N. Trathan, and John P. Croxall. Framework for mapping key areas for marine megafauna to inform Marine Spatial Planning: the Falkland Islands case study. Marine Policy, 92(??):61–72, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307790.

Agardy:2011:MGA

[AdSC11] Tundi Agardy, Giuseppe Notarbartolo di Sciara, and Patrick Christie. Mind the gap: Addressing the shortcomings of marine protected areas through large scale marine spatial planning. Marine Policy, 35(2):226–232, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001740.

Astorkiza:2018:EAP

[AdV18] Kepa Astorkiza and Ikerne del Valle. An economic analysis of private side of fishermen’s cofradías’ activity on the Cantabrian Sea. Marine Policy, 90(??):152–159, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307236.

Alam:2014:LRS

[AF14] Shawkat Alam and Abdullah Faruque. Legal regulation of the shipbreaking industry in Bangladesh: the international regulatory framework and domestic implementation challenges. Marine Policy, 47(??):46–56, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000384.

Ayer:2018:HSM

[AFCMET18] Austin Ayer, Stuart Fulton, Jacobo Alejandro Caamal-Madrigal, and Alejandro Espinoza-Tenorio. Halfway to sustainability: Management lessons from community-based, marine no-take zones in the Mexican Caribbean. Marine Policy, 93(??):22–30, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308989.

Alvarez-Fernandez:2017:MPM

[ÁFFSCF17] Inma Álvarez-Fernández, Nuria Fernández, Noela Sánchez-Carnero, and Juan Freire. The management performance of marine protected areas in the North-east Atlantic Ocean. Marine Policy, 76(??):159–168, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305085.

Avila-Foucat:2017:DPD

[AFGRF^{+}17] Veronique Sophie Avila-Foucat, Diane Gendron, Daniel Revollo-Fernandez, Ernesto Israel Popoca, and Alejandra Ramírez. Determinants of the potential demand for whale watching in Loreto Bay National Park. Marine Policy, 81(??):37–44, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304882.

Avila-Forcada:2012:CVM

[AFMCMP12] Sara Avila-Forcada, Adán L. Martínez-Cruz, and Carlos Muñoz-Piña. Conservation of vaquita marina in the Northern Gulf of California. Marine Policy, 36(3):613–622, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1100162X.

Afonso:2019:PSC

[AFR^{+}19] André S. Afonso, Leonardo L. Fidelis, Pedro L. Roque, Renato Galindo, Wanderson Dionisio, Leonardo B. Veras, and Fábio H. V. Hazin. Public support for conservation may decay with increasing residence time in suboptimal marine protected areas. Marine Policy, 108(??):

Article 103665, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300089.

Asche:2012:IFM

[AG12] Frank Asche and Jordi Guillen. The importance of fishing method, gear and origin: the Spanish hake market. Marine Policy, 36(2):365–369, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001205.

Ahmed:2016:CAM

[AG16] Nesar Ahmed and Marion Glaser. Coastal aquaculture, mangrove deforestation and blue carbon emissions: Is REDD+ a solution? Marine Policy, 66(??):58–66, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000166.

Aburto:2018:SSE

[AG18] Jaime A. Aburto and Carlos F. Gaymer. Struggling with social-ecological mismatches in marine management and conservation at Easter Island. Marine Policy, 92(??):21–29, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307649.

Agnew:2013:FCD

[AGB13] David J. Agnew, Nicolas L. Gutiérrez, and Doug S. Butterworth. Fish catch data: Less than what meets the eye. Marine Policy, 42(??):268–269, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000961.

Abramic:2018:ISR

[AGB^{+}18] Andrej Abramic, Daniel Gonzalez, Emanuele Bigagli, Anne Che-Bohnenstengel, and Paul Smits. INSPIRE: Support for and requirement of the Marine Strategy Framework Directive. Marine Policy, 92(??):86–100, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303123.

Agnew:2013:RFP

[AGSP^{+}13] D. J. Agnew, N. L. Gutiérrez, A. Stern-Pirlot, A. D. M. Smith, C. Zimmermann, and K. Sainsbury. Rebuttal to Froese and Proelss “Evaluation and legal assessment of certified seafood”. Marine Policy, 38(??):551–553, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001625.

Avery-Gomm:2019:TNC

[AGWMP19] Stephanie Avery-Gomm, Tony R. Walker, Mark L. Mallory, and Jennifer F. Provencher. There is nothing convenient about plastic pollution. Rejoinder to Stafford and Jones “Viewpoint — Ocean plastic pollution: a convenient but distracting truth?”. Marine Policy, 106(??):Article 103552, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301927.

Alexander:2019:HSM

[AH19] K. A. Alexander and M. Haward. The human side of marine ecosystem-based management (EBM): ‘Sectoral interplay’ as a challenge to implementing EBM. Marine Policy, 101(??):33–38, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306080.

Ayers:2018:BTC

[AHB18] Adam L. Ayers, Justin Hospital, and Christofer Boggs. Bigeye tuna catch limits lead to differential impacts for Hawai‘i longliners. Marine Policy, 94(??):93–105, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301313.

Appleby:2012:MCA

[AJ12] Tom Appleby and Peter J. S. Jones. The marine and coastal access act — a hornets’ nest? Marine Policy, 36(1):73–77, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000686.

Altamirano-Jimenez:2017:SOB

[AJ17] Isabel Altamirano-Jiménez. “The sea is our bread”: Interrupting green neoliberalism in Mexico. Marine Policy, 80(??):28–34, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300167.

Apgar-Kurtz:2015:FAL

[AK15] Breena Apgar-Kurtz. Factors affecting local permit ownership in Bristol Bay. Marine Policy, 56(??):71–77, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000482.

Almeida:2015:SMP

[AKV15] Cheila Almeida, Vanja Karadzic, and Sofia Vaz. The seafood market in Portugal: Driving forces and consequences. Marine Policy, 61(??):87–94, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002134.

Allo:2017:RSN

[AL17] Maria Alló and Maria L. Loureiro. The role of social norms on conservation programmes in shellfish fisheries. Marine Policy, 84(??):134–141, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306297.

Almklov:2018:TGI

[AL18] Petter Grytten Almklov and Gunnar M. Lamvik. Taming a globalized industry — Forces and counter forces influencing maritime safety. Marine Policy, 96(??):175–183, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301209.

Alfnes:2017:SOS

[Alf17] Frode Alfnes. Selling only sustainable seafood: Attitudes toward public regulation and retailer policies. Marine Policy, 78(??):74–79, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304122.

Allen:2014:CSP

[All14] Stewart Allen. Catch share programs in the Western Pacific — a solution in search of a problem? Marine Policy, 45(??):121–127, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002509.

Atmodjo:2017:FMC

[ALM17] Ery Atmodjo, Machiel Lamers, and Arthur Mol. Financing marine conservation tourism: Governing entrance fees in Raja Ampat, Indonesia. Marine Policy, 78(??):181–188, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300106.

Allan:2014:EIM

[ALMS14] G. J. Allan, P. Lecca, P. G. McGregor, and J. K. Swales. The economic impacts of marine energy developments: a case study from Scotland. Marine Policy, 43(??): 122–131, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001073.

Abecasis:2013:MCR

[ALSC13] Rita Costa Abecasis, Nancy Longnecker, Luisa Schmidt, and Julian Clifton. Marine conservation in remote small island settings: Factors influencing marine protected area establishment in the Azores. Marine Policy, 40(??):1–9, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000031.

Aranda:2015:AFP

[AM15] Martin Aranda and Arantza Murillas. Allocation of fishing possibilities, incentives and outcomes: Insights from Basque fishermen’s organisations in Spain. Marine Policy, 61(??):171–178, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002250.

Aranda:2012:MFC

[AMdB12] Martin Aranda, Hilario Murua, and Paul de Bruyn. Managing fishing capacity in tuna regional fisheries management organisations (RFMOs): Development and state of the art. Marine Policy, 36(5):985–992, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000103.

Acuna-Marrero:2018:UHA

[AMdlCMS^{+}18] David Acuña-Marrero, Raquel de la Cruz-Modino, Adam N. H. Smith, Pelayo Salinas de León, Matthew D. M. Pawley, and Marti J. Anderson. Understanding human attitudes towards sharks to promote sustainable coexistence. Marine Policy, 91(??):122–128, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303056.

Agnarsson:2016:CDQ

[AMG16] S. Agnarsson, T. Matthiasson, and F. Giry. Consolidation and distribution of quota holdings in the Icelandic fisheries. Marine Policy, 72(??):263–270, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302238.

Ainsworth:2012:EEB

[AMLK^{+}12] C. H. Ainsworth, H. Morzaria-Luna, I. C. Kaplan, P. S. Levin, E. A. Fulton, R. Cudney-Bueno, P. Turk-Boyer, J. Torre, G. D. Danemann, and T. Pfister. Effective ecosystem-based management must encourage regulatory compliance: a Gulf of California case study. Marine Policy, 36(6):1275–1283, November 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000607.

Ashley:2014:POW

[AMR14] M. C. Ashley, S. C. Mangi, and L. D. Rodwell. The potential of offshore windfarms to act as marine protected areas — a systematic review of current evidence. Marine Policy, 45(??):301–309, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001991.

Armitage:2011:EST

[AMvT11] Derek Armitage, Melissa Marschke, and Truong van Tuyen. Early-stage transformation of coastal marine governance in Vietnam? Marine Policy, 35(5):703–711, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000479.

Anderson:2015:TMF

[And15] Tarryn-Anne Anderson. Tracking the movement of fish: Skippers’ logbooks and contestations over ways of knowing the sea. Marine Policy, 60(??):318–324, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002619.

Anonymous:2010:EBAa

[Ano10a] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(1):ifc, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001274.

Anonymous:2010:EBAb

[Ano10b] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(2):ifc, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0900150X.

Anonymous:2010:EBAc

[Ano10c] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(3):ifc, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000308.

Anonymous:2010:EBAd

[Ano10d] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(4):ifc, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000758.

Anonymous:2010:EBAe

[Ano10e] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(5):ifc, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000874.

Anonymous:2010:EBAf

[Ano10f] Anonymous. Editorial Board/aims and scope. Marine Policy, 34(6):ifc, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001211.

Anonymous:2010:PJ

[Ano10g] Anonymous. Pages 1–196 (January 2010). Marine Policy, 34(1):??, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2010:PN

[Ano10h] Anonymous. Pages 1117-1388 (November 2010). Marine Policy, 34(6):??, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2010:PMa

[Ano10i] Anonymous. Pages 197-352 (March 2010). Marine Policy, 34(2):??, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2010:PMb

[Ano10j] Anonymous. Pages 353-738 (May 2010). Marine Policy, 34 (3):??, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2010:PS

[Ano10k] Anonymous. Pages 821-1116 (September 2010). Marine Policy, 34(5):??, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2011:EBAa

[Ano11a] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(1):ifc, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001570.

Anonymous:2011:EBAb

[Ano11b] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(2):ifc, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001867.

Anonymous:2011:EBAc

[Ano11c] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(3):ifc, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002186.

Anonymous:2011:EBAd

[Ano11d] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(4):ifc, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000339.

Anonymous:2011:EBAe

[Ano11e] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(5):ifc, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000571.

Anonymous:2011:EBAf

[Ano11f] Anonymous. Editorial Board/aims and scope. Marine Policy, 35(6):ifc, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000832.

Anonymous:2011:PJ

[Ano11g] Anonymous. Pages 1–94 (January 2011). Marine Policy, 35(1):??, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2011:PMb

[Ano11h] Anonymous. Pages 271-426 (May 2011). Marine Policy, 35 (3):??, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2011:PS

[Ano11i] Anonymous. Pages 559-736 (September 2011). Marine Policy, 35(5):??, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2011:PMa

[Ano11j] Anonymous. Pages 95-270 (March 2011). Marine Policy, 35(2):??, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2011:PN

[Ano11k] Anonymous. Publisher’s note. Marine Policy, 35(5): 559, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1100073X.

Anonymous:2012:EBAa

[Ano12a] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(1):ifc, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001266.

Anonymous:2012:EBAb

[Ano12b] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(2):ifc, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001746.

Anonymous:2012:EBAc

[Ano12c] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(3):ifc, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1100203X.

Anonymous:2012:EBAd

[Ano12d] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(4):ifc, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000279.

Anonymous:2012:EBAe

[Ano12e] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(5):ifc, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000772.

Anonymous:2012:EBAf

[Ano12f] Anonymous. Editorial Board/aims and scope. Marine Policy, 36(6):ifc, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1200108X.

Anonymous:2012:PJa

[Ano12g] Anonymous. Pages 1–306 (January 2012). Marine Policy, 36(1):??, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2012:PN

[Ano12h] Anonymous. Pages 1193-1314 (November 2012). Marine Policy, 36(6):??, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2012:PMa

[Ano12i] Anonymous. Pages 307–566 (March 2012). Marine Policy, 36(2):??, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2012:PMb

[Ano12j] Anonymous. Pages 567-844 (May 2012). Marine Policy, 36 (3):??, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2012:PJb

[Ano12k] Anonymous. Pages 845-946 (July 2012). Marine Policy, 36 (4):??, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2012:PS

[Ano12l] Anonymous. Pages 947-1192 (September 2012). Marine Policy, 36(5):??, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2013:EBAa

[Ano13a] Anonymous. Editorial Board/aims and scope. Marine Policy, 37(??):ifc, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001807.

Anonymous:2013:EBAb

[Ano13b] Anonymous. Editorial Board/aims and scope. Marine Policy, 38(??):ifc, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002278.

Anonymous:2013:EBAc

[Ano13c] Anonymous. Editorial Board/aims and scope. Marine Policy, 39(??):ifc, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000316.

Anonymous:2013:EBAd

[Ano13d] Anonymous. Editorial Board/aims and scope. Marine Policy, 40(??):ifc, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000584.

Anonymous:2013:EBAe

[Ano13e] Anonymous. Editorial Board/aims and scope. Marine Policy, 41(??):ifc, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000742.

Anonymous:2013:EBAf

[Ano13f] Anonymous. Editorial Board/aims and scope. Marine Policy, 42(??):ifc, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001115.

Anonymous:2013:PJ

[Ano13g] Anonymous. Pages 1–204 (July 2013). Marine Policy, 40 (??):1–204, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2013:PN

[Ano13h] Anonymous. Pages 1–336 (November 2013). Marine Policy, 42(??):1–336, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2013:PMb

[Ano13i] Anonymous. Pages 1–364 (May 2013). Marine Policy, 39 (??):1–364, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2013:PMa

[Ano13j] Anonymous. Pages 1–566 (March 2013). Marine Policy, 38(??):1–566, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:EBAa

[Ano14a] Anonymous. Editorial Board/aims and scope. Marine Policy, 43(??):ifc, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001814.

Anonymous:2014:EBAb

[Ano14b] Anonymous. Editorial Board/aims and scope. Marine Policy, 44(??):ifc, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002546.

Anonymous:2014:EBAc

[Ano14c] Anonymous. Editorial Board/aims and scope. Marine Policy, 45(??):ifc, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000153.

Anonymous:2014:EBAd

[Ano14d] Anonymous. Editorial Board/aims and scope. Marine Policy, 46(??):ifc, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1400058X.

Anonymous:2014:EBAe

[Ano14e] Anonymous. Editorial Board/aims and scope. Marine Policy, 47(??):ifc, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001018.

Anonymous:2014:EBAf

[Ano14f] Anonymous. Editorial Board/aims and scope. Marine Policy, 48(??):ifc, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001201.

Anonymous:2014:EBAg

[Ano14g] Anonymous. Editorial Board/aims and scope. Marine Policy, 49(??):ifc, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001511.

Anonymous:2014:EBAh

[Ano14h] Anonymous. Editorial Board/aims and scope. Marine Policy, 50(??):ifc, December 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002140.

Anonymous:2014:PJb

[Ano14i] Anonymous. Pages 1–162 (July 2014). Marine Policy, 47 (??):1–162, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PN

[Ano14j] Anonymous. Pages 1–194 (November 2014). Marine Policy, 49(??):1–194, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PMb

[Ano14k] Anonymous. Pages 1–196 (May 2014). Marine Policy, 46 (??):1–196, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PS

[Ano14l] Anonymous. Pages 1–212 (September 2014). Marine Policy, 48(??):1–212, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PMa

[Ano14m] Anonymous. Pages 1–368 (March 2014). Marine Policy, 45(??):1–368, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PJa

[Ano14n] Anonymous. Pages 1–386 (January 2014). Marine Policy, 43(??):1–386, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2014:PF

[Ano14o] Anonymous. Pages 1–478 (February 2014). Marine Policy, 44(??):1–478, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:EB

[Ano15a] Anonymous. Editorial Board. Marine Policy, 62(??): ifc, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003371.

Anonymous:2015:EBAa

[Ano15b] Anonymous. Editorial Board/aims and scope. Marine Policy, 51(??):ifc, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003066.

Anonymous:2015:EBAb

[Ano15c] Anonymous. Editorial Board/aims and scope. Marine Policy, 52(??):ifc, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003364.

Anonymous:2015:EBAc

[Ano15d] Anonymous. Editorial Board/aims and scope. Marine Policy, 53(??):ifc, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000081.

Anonymous:2015:EBAd

[Ano15e] Anonymous. Editorial Board/aims and scope. Marine Policy, 54(??):ifc, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000329.

Anonymous:2015:EBAe

[Ano15f] Anonymous. Editorial Board/aims and scope. Marine Policy, 55(??):ifc, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001086.

Anonymous:2015:EBAf

[Ano15g] Anonymous. Editorial Board/aims and scope. Marine Policy, 56(??):ifc, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001190.

Anonymous:2015:EBAg

[Ano15h] Anonymous. Editorial Board/aims and scope. Marine Policy, 57(??):ifc, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1500127X.

Anonymous:2015:EBAh

[Ano15i] Anonymous. Editorial Board/aims and scope. Marine Policy, 58(??):ifc, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001499.

Anonymous:2015:EBAi

[Ano15j] Anonymous. Editorial Board/aims and scope. Marine Policy, 59(??):ifc, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001839.

Anonymous:2015:EBAj

[Ano15k] Anonymous. Editorial Board/aims and scope. Marine Policy, 60(??):ifc, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002407.

Anonymous:2015:EBAk

[Ano15l] Anonymous. Editorial Board/aims and scope. Marine Policy, 61(??):ifc, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003152.

Anonymous:2015:PAb

[Ano15m] Anonymous. Pages 1–124 (August 2015). Marine Policy, 58(??):1–124, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PJb

[Ano15n] Anonymous. Pages 1–124 (June 2015). Marine Policy, 56 (??):1–124, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PS

[Ano15o] Anonymous. Pages 1–134 (September 2015). Marine Policy, 59(??):1–134, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PAa

[Ano15p] Anonymous. Pages 1–142 (April 2015). Marine Policy, 54 (??):1–142, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PF

[Ano15q] Anonymous. Pages 1–172 (February 2015). Marine Policy, 52(??):1–172, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PJc

[Ano15r] Anonymous. Pages 1–192 (July 2015). Marine Policy, 57 (??):1–192, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PM

[Ano15s] Anonymous. Pages 1–226 (March 2015). Marine Policy, 53(??):1–226, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PO

[Ano15t] Anonymous. Pages 1–352 (October 2015). Marine Policy, 60(??):1–352, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PD

[Ano15u] Anonymous. Pages 1–358 (December 2015). Marine Policy, 62(??):1–358, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PN

[Ano15v] Anonymous. Pages 1–430 (November 2015). Marine Policy, 61(??):1–430, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PJa

[Ano15w] Anonymous. Pages 1–592 (January 2015). Marine Policy, 51(??):1–592, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2015:PRR

[Ano15x] Anonymous. Pages R1-R2, 1-116 (May 2015). Marine Policy, 55(??):R1–, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Marine Policy, 63 (??):ifc, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003498.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Marine Policy, 64 (??):ifc, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003711.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Marine Policy, 65(??):ifc, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000300.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Marine Policy, 66(??):ifc, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000531.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Marine Policy, 67(??):ifc, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301002.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Marine Policy, 68(??):ifc, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301646.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. Marine Policy, 69(??):ifc, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302639.

Anonymous:2016:EBh

[Ano16h] Anonymous. Editorial Board. Marine Policy, 70(??):ifc, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303426.

Anonymous:2016:EBi

[Ano16i] Anonymous. Editorial Board. Marine Policy, 71(??): ifc, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304444.

Anonymous:2016:EBj

[Ano16j] Anonymous. Editorial Board. Marine Policy, 72 (??):ifc, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305140.

Anonymous:2016:EBk

[Ano16k] Anonymous. Editorial Board. Marine Policy, 73(??): ifc, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305929.

Anonymous:2016:EBl

[Ano16l] Anonymous. Editorial Board. Marine Policy, 74(??): ifc, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307400.

Anonymous:2016:PMa

[Ano16m] Anonymous. Pages 1–126 (March 2016). Marine Policy, 65(??):1–126, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PAa

[Ano16n] Anonymous. Pages 1–142 (April 2016). Marine Policy, 66 (??):1–142, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PF

[Ano16o] Anonymous. Pages 1–174 (February 2016). Marine Policy, 64(??):1–174, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PMb

[Ano16p] Anonymous. Pages 1–188 (May 2016). Marine Policy, 67 (??):1–188, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PJa

[Ano16q] Anonymous. Pages 1–212 (January 2016). Marine Policy, 63(??):1–212, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PAb

[Ano16r] Anonymous. Pages 1–226 (August 2016). Marine Policy, 70(??):1–226, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PJc

[Ano16s] Anonymous. Pages 1–236 (July 2016). Marine Policy, 69 (??):1–236, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PJb

[Ano16t] Anonymous. Pages 1–240 (June 2016). Marine Policy, 68 (??):1–240, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PN

[Ano16u] Anonymous. Pages 1–276 (November 2016). Marine Policy, 73(??):1–276, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PO

[Ano16v] Anonymous. Pages 1–284 (October 2016). Marine Policy, 72(??):1–284, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PS

[Ano16w] Anonymous. Pages 1–334 (September 2016). Marine Policy, 71(??):1–334, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2016:PD

[Ano16x] Anonymous. Pages 1–350 (December 2016). Marine Policy, 74(??):1–350, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Marine Policy, 75 (??):ifc, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307679.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Marine Policy, 76 (??):ifc, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16308314.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Marine Policy, 77(??):ifc, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300283.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. Marine Policy, 78(??):ifc, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301306.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Marine Policy, 79(??):ifc, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301501.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Marine Policy, 80(??):ifc, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730221X.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Marine Policy, 81(??):ifc, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302853.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. Marine Policy, 82(??):ifc, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303664.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. Marine Policy, 83(??): ifc, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17304566.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. Marine Policy, 84 (??):ifc, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730595X.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial Board. Marine Policy, 85(??): ifc, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730622X.

Anonymous:2017:EBl

[Ano17l] Anonymous. Editorial Board. Marine Policy, 86(??): ifc, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306851.

Anonymous:2017:PNb

[Ano17m] Anonymous. Pages 1–150 (November 2017). Marine Policy, 85(??):1–150, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PAa

[Ano17n] Anonymous. Pages 1–206 (April 2017). Marine Policy, 78 (??):1–206, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PF

[Ano17o] Anonymous. Pages 1–210 (February 2017). Marine Policy, 76(??):1–210, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PAb

[Ano17p] Anonymous. Pages 1–234 (August 2017). Marine Policy, 82(??):1–234, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PMa

[Ano17q] Anonymous. Pages 1–238 (March 2017). Marine Policy, 77(??):1–238, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PD

[Ano17r] Anonymous. Pages 1–258 (December 2017). Marine Policy, 86(??):1–258, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PS

[Ano17s] Anonymous. Pages 1–268 (September 2017). Marine Policy, 83(??):1–268, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PJa

[Ano17t] Anonymous. Pages 1–318 (January 2017). Marine Policy, 75(??):1–318, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PO

[Ano17u] Anonymous. Pages 1–332 (October 2017). Marine Policy, 84(??):1–332, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PJb

[Ano17v] Anonymous. Pages 1–418 (July 2017). Marine Policy, 81 (??):1–418, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PMb

[Ano17w] Anonymous. Pages 1–92 (May 2017). Marine Policy, 79 (??):1–92, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2017:PNa

[Ano17x] Anonymous. Publisher’s Note. Marine Policy, 75 (??):iv, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16308090.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Marine Policy, 87 (??):ifc, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308126.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Marine Policy, 88 (??):ii, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300071.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Marine Policy, 89 (??):ii, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830054X.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Marine Policy, 90(??):ii, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301052.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Marine Policy, 91(??):ii, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301702.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Marine Policy, 92(??):ii, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302148.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Marine Policy, 93(??):ii, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303476.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Marine Policy, 94(??):ii, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304263.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Marine Policy, 95(??): ii, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830558X.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Marine Policy, 96(??):ii, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306511.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Marine Policy, 97 (??):ii, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306912.

Anonymous:2018:EBl

[Ano18l] Anonymous. Editorial Board. Marine Policy, 98(??):ii, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307814.

Anonymous:2018:PJb

[Ano18m] Anonymous. Pages 1–148 (June 2018). Marine Policy, 92 (??):1–148, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PFb

[Ano18n] Anonymous. Pages 1–154 (February 2018). Marine Policy, 89(??):1–154, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PM

[Ano18o] Anonymous. Pages 1–166 (May 2018). Marine Policy, 91 (??):1–166, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PAa

[Ano18p] Anonymous. Pages 1–202 (April 2018). Marine Policy, 90 (??):1–202, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PAb

[Ano18q] Anonymous. Pages 1–282 (August 2018). Marine Policy, 94(??):1–282, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PO

[Ano18r] Anonymous. Pages 1–298 (October 2018). Marine Policy, 96(??):1–298, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PJc

[Ano18s] Anonymous. Pages 1–306 (July 2018). Marine Policy, 93 (??):1–306, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PN

[Ano18t] Anonymous. Pages 1–314 (November 2018). Marine Policy, 97(??):1–314, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PD

[Ano18u] Anonymous. Pages 1–324 (December 2018). Marine Policy, 98(??):1–324, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PFa

[Ano18v] Anonymous. Pages 1–364 (February 2018). Marine Policy, 88(??):1–364, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PJa

[Ano18w] Anonymous. Pages 1–378 (January 2018). Marine Policy, 87(??):1–378, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2018:PS

[Ano18x] Anonymous. Pages 1–402 (September 2018). Marine Policy, 95(??):1–402, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:A

[Ano19a] Anonymous. August 2019. Marine Policy, 106(??):??, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:D

[Ano19b] Anonymous. December 2019. Marine Policy, 110(??):??, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:EBa

[Ano19c] Anonymous. Editorial Board. Marine Policy, 99 (??):ii, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830887X.

Anonymous:2019:EBb

[Ano19d] Anonymous. Editorial Board. Marine Policy, 100 (??):ii, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18309898.

Anonymous:2019:EBc

[Ano19e] Anonymous. Editorial Board. Marine Policy, 101(??):ii, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19300612.

Anonymous:2019:EBd

[Ano19f] Anonymous. Editorial Board. Marine Policy, 102(??):ii, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301447.

Anonymous:2019:EBe

[Ano19g] Anonymous. Editorial Board. Marine Policy, 103(??):ii, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301861.

Anonymous:2019:EBf

[Ano19h] Anonymous. Editorial Board. Marine Policy, 104(??):ii, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302556.

Anonymous:2019:EBg

[Ano19i] Anonymous. Editorial Board. Marine Policy, 105(??):ii, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303574.

Anonymous:2019:EBh

[Ano19j] Anonymous. Editorial Board. Marine Policy, 106(??): Article 103578, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19304336.

Anonymous:2019:EBi

[Ano19k] Anonymous. Editorial Board. Marine Policy, 107(??): Article 103607, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19305020.

Anonymous:2019:EBj

[Ano19l] Anonymous. Editorial Board. Marine Policy, 108(??): Article 103684, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19306724.

Anonymous:2019:EBk

[Ano19m] Anonymous. Editorial Board. Marine Policy, 109(??): Article 103734, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19308115.

Anonymous:2019:EBl

[Ano19n] Anonymous. Editorial Board. Marine Policy, 110(??): Article 103751, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19308528.

Anonymous:2019:N

[Ano19o] Anonymous. November 2019. Marine Policy, 109(??):??, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:O

[Ano19p] Anonymous. October 2019. Marine Policy, 108(??):??, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PJc

[Ano19q] Anonymous. Pages 1–180 (July 2019). Marine Policy, 105 (??):1–180, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PMb

[Ano19r] Anonymous. Pages 1–192 (May 2019). Marine Policy, 103 (??):1–192, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PJb

[Ano19s] Anonymous. Pages 1–238 (June 2019). Marine Policy, 104 (??):1–238, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PMa

[Ano19t] Anonymous. Pages 1–286 (March 2019). Marine Policy, 101(??):1–286, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PF

[Ano19u] Anonymous. Pages 1–382 (February 2019). Marine Policy, 100(??):1–382, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PJa

[Ano19v] Anonymous. Pages 1–422 (January 2019). Marine Policy, 99(??):1–422, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PA

[Ano19w] Anonymous. Pages 1–68 (April 2019). Marine Policy, 102 (??):1–68, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2019:PN

[Ano19x] Anonymous. Publisher’s Note. Marine Policy, 106(??): Article 103579, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19304348.

Anonymous:2019:S

[Ano19y] Anonymous. September 2019. Marine Policy, 107(??):??, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Antonova:2016:RRF

[Ant16] Anna S. Antonova. The rhetoric of “responsible fishing”: Notions of human rights and sustainability in the European Union’s bilateral fishing agreements with developing states. Marine Policy, 70(??):77–84, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301749.

Albert:2015:RRP

[AOA^{+}15] Joelle A. Albert, Andrew D. Olds, Simon Albert, Annabelle Cruz-Trinidad, and Anne-Maree Schwarz. Reaping the reef: Provisioning services from coral reefs in Solomon Islands. Marine Policy, 62(??):244–251, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002742.

Ahmed:2013:ICC

[AOAM13] Nesar Ahmed, Anna Occhipinti-Ambrogi, and James F. Muir. The impact of climate change on prawn postlarvae fishing in coastal Bangladesh: Socioeconomic and ecological perspectives. Marine Policy, 39(??):224–233, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002060.

Abdullah:2016:TPT

[AOY16] Nazirah Mohamad Abdullah, Abdullah Hisam Omar, and Omar Yaakob. Today’s problem, tomorrow’s solutions: Lay theory explanations of marine space stakeholder management in the Malaysian context. Marine Policy, 73(??):162–171, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302883.

Ayles:2016:DIF

[APC16] Burton Ayles, Louie Porta, and Red McV Clarke. Development of an integrated fisheries co-management framework for new and emerging commercial fisheries in the Canadian Beaufort Sea. Marine Policy, 72(??):246–254, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302226.

Audouit:2019:CSR

[APD^{+}19] Christelle Audouit, Vanina Pasqualini, Rutger De Wit, Hervé Flanquart, Philippe Deboudt, and Caroline Rufin-Soler. Comparing social representation of water quality in coastal lagoons with normative use of ecological indicators. Marine Policy, 101(??):137–146, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307552.

Aksenov:2017:FNA

[APY^{+}17] Yevgeny Aksenov, Ekaterina E. Popova, Andrew Yool, A. J. George Nurser, Timothy D. Williams, Laurent Bertino, and Jon Bergh. On the future navigability of Arctic sea routes: High-resolution projections of the Arctic Ocean and sea ice. Marine Policy, 75(??):300–317, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000038.

Ardron:2018:TOI

[Ard18] Jeff A. Ardron. Transparency in the operations of the International Seabed Authority: an initial assessment.

Marine Policy, 95(??):324–331, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303281.

Alcaide:2017:EPS

[ARDP17] Juan Ignacio Alcaide, Emilio Rodríguez-Díaz, and Francisco Piniella. European policies on ship recycling: a stakeholder survey. Marine Policy, 81(??):262–272, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308545.

Ardron:2014:SUC

[ARGW14] Jeff A. Ardron, Rosemary Rayfuse, Kristina Gjerde, and Robin Warner. The sustainable use and conservation of biodiversity in ABNJ: What can be achieved using existing international agreements? Marine Policy, 49(??): 98–108, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000499.

Ardron:2018:ITG

[ARJ18] Jeff A. Ardron, Henry A. Ruhl, and Daniel O. B. Jones. Incorporating transparency into the governance of deepseabed mining in the Area beyond national jurisdiction. Marine Policy, 89(??):58–66, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304207.

Arnadottir:2017:ECJ

[Árn17] Snjólaug Árnadóttir. Ecological changes justifying termination or revision of EEZ and EFZ boundaries. Marine Policy, 84(??):287–292, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302440.

Asche:2016:TIS

[ARS16] Frank Asche, Cathy A. Roheim, and Martin D. Smith. Trade intervention: Not a silver bullet to address environmental externalities in global aquaculture. Marine Policy, 69(??):194–201, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001803.

Adhuri:2016:GMS

[ARSHH16] Dedi Supriadi Adhuri, Laksmi Rachmawati, Hirmen Sofyanto, and Natasha Hamilton-Hart. Green market for small people: Markets and opportunities for upgrading in small-scale fisheries in Indonesia. Marine Policy, 63 (??):198–205, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000706.

Asaro:2012:GAM

[Asa12] Michael J. Asaro. Geospatial analysis of management areas implemented for protection of the North Atlantic right whale along the northern Atlantic coast of the United States. Marine Policy, 36(4):915–921, July 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200005X.

Alter:2010:FCC

[ASB10] S. Elizabeth Alter, Mark P. Simmonds, and John R. Brandon. Forecasting the consequences of climatedriven shifts in human behavior on cetaceans. Marine Policy, 34(5):943–954, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000278.

Al-Subhi:2013:FCM

[ASBAM13] Khalid Khalifa Nasser Al-Subhi, Shekar Bose, and Hussein Samh Al-Masroori. Fishers’ compliance motivations: a case study of the Sultanate of Oman. Marine Policy, 37(??):141–148, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000590.

Asgeirsdottir:2016:SMB

[Ásg16] Áslaug Ásgeirsdóttir. Settling of the maritime boundaries of the United States: Cost of settlement and the benefits of legal certainty. Marine Policy, 73(??):187–195, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303268.

Aguado:2016:TSF

[ASP16] S. Hernández Aguado, I. Segado Segado, and Tony J. Pitcher. Towards sustainable fisheries: a multi-criteria participatory approach to assessing indicators of sustainable fishing communities: a case study from Cartagena (Spain). Marine Policy, 65(??):97–106, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003942.

Amorim:2018:GKI

[ASWM18] Patrícia Amorim, Pedro Sousa, Megan Westmeyer, and Gui M. Menezes. Generic Knowledge Indicator (GKI): a tool to evaluate the state of knowledge of fisheries applied to snapper and grouper. Marine Policy, 89(??): 40–49, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303068.

Abila:2014:TPT

[AT14] Sanley S. Abila and Lijun Tang. Trauma, post-trauma, and support in the shipping industry: the experience of Filipino seafarers after pirate attacks. Marine Policy, 46(??):132–136, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000244.

Ahmed:2010:PPF

[ATAM10] Nesar Ahmed, Max Troell, Edward H. Allison, and James F. Muir. Prawn postlarvae fishing in coastal Bangladesh: Challenges for sustainable livelihoods. Marine Policy, 34(2):218–227, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X09000979.

Awabdi:2018:ICA

[ATB^{+}18] Danielle Rodrigues Awabdi, Davi Castro Tavares, Ana Cristina Vigliar Bondioli, Camilah Antunes Zappes, and Ana Paula Madeira Di Beneditto. Influences of conservation action on attitudes and knowledge of fishermen towards sea turtles along the southeastern Brazil. Marine Policy, 95(??):57–68, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300897.

Alava:2019:MCB

[ATB^{+}19] Juan José Alava, Bradley Tatar, María José Barragán, Cristina Castro, Patricia Rosero, Judith Denkinger, Pedro J. Jiménez, Raúl Carvajal, and Jorge Samaniego. Mitigating cetacean bycatch in coastal Ecuador: Governance challenges for small-scale fisheries. Marine Policy, 110(??):Article 102769, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308508.

Aziz:2016:APR

[ATDP16] Ammar Abdul Aziz, Sebastian Thomas, Paul Dargusch, and Stuart Phinn. Assessing the potential of REDD+ in a production mangrove forest in Malaysia using stakeholder analysis and ecosystem services mapping. Marine Policy, 74(??):6–17, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303748.

Ancic:2018:DEP

[AVC18] Ivica Ančić, Nikola Vladimir, and Dae-Seung Cho. Determining environmental pollution from ships using Index of Energy Efficiency and Environmental Eligibility (I4E). Marine Policy, 95(??):1–7, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302008.

Angulo-Valdes:2010:NTB

[AVH10] Jorge A. Angulo-Valdés and Bruce G. Hatcher. A new typology of benefits derived from marine protected areas. Marine Policy, 34(3):635–644, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001882.

Avoyan:2017:PBS

[AvTT17] Emma Avoyan, Jan van Tatenhove, and Hilde Toonen. The performance of the Black Sea Commission as a collaborative governance regime. Marine Policy, 81(??):285–292, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300714.

Alexander:2013:ASF

[AWH13] Karen A. Alexander, Thomas A. Wilding, and Johanna Jacomina Heymans. Attitudes of Scottish fishers towards marine renewable energy. Marine Policy, 37(??):239–244, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000930.

Agnisola:2019:VMN

[AWJ19] Giulia Agnisola, Stephanie Weir, and Kate Johnson. The voices that matter: a narrative approach to understanding Scottish Fishers’ perspectives of Brexit. Marine Policy, 110(??):Article 103563, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308546.

Anna:2019:FHE

[AYA^{+}19] Zuzy Anna, Arief A. Yusuf, Armida S. Alisjahbana, Aisyah A. Ghina, and Rahma. Are fishermen happier? Evidence from a large-scale subjective wellbeing survey in a lower-middle-income country. Marine Policy, 106(??):Article 103559, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308960.

Ankamah-Yeboah:2018:MIC

[AYB18] Isaac Ankamah-Yeboah and Julia Bronnmann. Market integration in the crustaceans market: Evidence from Germany. Marine Policy, 87(??):72–77, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306036.

Agundez:2014:MEV

[AYR^{+}14] José A. Pérez Agúndez, Eden Yimam, Pascal Raux, Hélène Rey-Valette, and Sophie Girard. Modeling economic vulnerability: As applied to microbiological contamination on the Thau Lagoon shellfish farming industry. Marine Policy, 46(??):143–151, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000256.

Al-Zaidan:2013:SKF

[AZAMG13] Amani S. Y. Al-Zaidan, Salim Y. Al-Mohanna, and Preeta George. Status of Kuwait’s fishery resources: Assessment and perspective. Marine Policy, 38(??):1–7, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001017.

Bye:2019:NMA

[BA19] Rolf Johan Bye and Petter G. Almklov. Normalization of maritime accident data using AIS. Marine Policy, 109(??):Article 103675, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930154X.

Ban:2010:CIM

[BAA10] Natalie C. Ban, Hussein M. Alidina, and Jeff A. Ardron. Cumulative impact mapping: Advances, relevance and limitations to marine management and conservation, using Canada’s Pacific waters as a case study. Marine Policy, 34(5):876–886, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000114.

Bell:2015:OUN

[BAA^{+}15a] Johann D. Bell, Joelle Albert, Serge Andréfouët, Neil L. Andrew, Michel Blanc, Philip Bright, Deidre Brogan, Brooke Campbell, Hugh Govan, John Hampton, Quentin Hanich, Shelton Harley, Arthur Jorari, Marcus Lincoln Smith, Scott Pontifex, Michael K. Sharp, William Sokimi, and Arthur Webb. Optimising the use of nearshore fish aggregating devices for food security in the Pacific Islands. Marine Policy, 56(??):98–105, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000457.

Bell:2015:DUT

[BAA^{+}15b] Johann D. Bell, Valerie Allain, Edward H. Allison, Serge Andréfouët, Neil L. Andrew, Michael J. Batty, Michel Blanc, Jeffrey M. Dambacher, John Hampton, Quentin Hanich, Shelton Harley, Anne Lorrain, Michael McCoy, Nicholas McTurk, Simon Nicol, Graham Pilling, David Point, Michael K. Sharp, Paula Vivili, and Peter Williams. Diversifying the use of tuna to improve food security and public health in Pacific Island countries and territories. Marine Policy, 51(??):584–591, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400267X.

Bell:2018:OAO

[BAA^{+}18] Johann D. Bell, Joelle Albert, George Amos, Christopher Arthur, Michel Blanc, Don Bromhead, Scott F. Heron, Alistair J. Hobday, Andrew Hunt, David Itano, Philip A. S. James, Patrick Lehodey, Gang Liu, Simon Nicol, Jim Potemra, Gabriel Reygondeau, Jason Rubani, Joe Scutt Phillips, Inna Senina, and William Sokimi. Operationalising access to oceanic fisheries resources by smallscale fishers to improve food security in the Pacific Islands. Marine Policy, 88(??):315–322, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301161.

Blasiak:2014:PSO

[BAB^{+}14] Robert Blasiak, James L. Anderson, Peter Bridgewater, Ken Furuya, Benjamin S. Halpern, Hisashi Kurokura, Joji Morishita, Nobuyuki Yagi, and Akane Minohara. Paradigms of sustainable ocean management. Marine Policy, 48(??):206–211, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000931.

Backer:2018:RWP

[Bac18] Hermanni Backer. Regional work on prevention of pollution from ships in the Baltic Sea — a paradox or a global forerunner? Marine Policy, 98(??):255–263, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306857.

Broggiato:2014:FES

[BAHCG14] Arianna Broggiato, Sophie Arnaud-Haond, Claudio Chiarolla, and Thomas Greiber. Fair and equitable sharing of benefits from the utilization of marine genetic resources in areas beyond national jurisdiction: Bridging the gaps between science and policy. Marine Policy, 49(??):176–185, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000505.

Badjeck:2010:ICV

[BAHD10] Marie-Caroline Badjeck, Edward H. Allison, Ashley S. Halls, and Nicholas K. Dulvy. Impacts of climate variability and change on fishery-based livelihoods. Marine Policy, 34(3):375–383, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001237.

Bailey:2012:WWB

[Bai12] Jennifer L. Bailey. Whale watching, the Buenos Aires Group and the politics of the International Whaling Commission. Marine Policy, 36(2):489–494, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001473.

Bailey:2016:ACD

[Bai16] Jennifer Bailey. Adventures in cross-disciplinary studies: Grand strategy and fisheries management. Marine Policy, 63(??):18–27, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1500264X.

Bose:2017:TFE

[BAMAH17] Shekar Bose, Hussein Samh Al-Masroori, and Al-Mutasim Hammoud Ahmed Al-Habsi. Traditional fisheries enforcement program: a case of three coastal villages in the eastern part of Oman. Marine Policy, 78(??):61–67, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303578.

Boxall:2012:AEB

[BAO^{+}12] P. C. Boxall, W. L. Adamowicz, M. Olar, G. E. West, and G. Cantin. Analysis of the economic benefits associated with the recovery of threatened marine mammal species in the Canadian St. Lawrence Estuary. Marine Policy, 36(1):189–197, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000923.

Barclay:2010:ITI

[Bar10] Kate Barclay. Impacts of tuna industries on coastal communities in Pacific Island countries. Marine Policy, 34 (3):406–413, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001389.

Brochier:2018:COF

[BAT^{+}18] T. Brochier, P. Auger, D. Thiao, A. Bah, S. Ly, T. Nguyen-Huu, and P. Brehmer. Can overexploited fisheries recover by self-organization? Reallocation of fishing effort as an emergent form of governance. Marine Policy, 95(??):46–56, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302130.

Brownscombe:2019:BKA

[BAY^{+}19] Jacob W. Brownscombe, Aaron J. Adams, Nathan Young, Lucas P. Griffin, Peter E. Holder, John Hunt, Alejandro Acosta, Danielle Morley, Ross Boucek, Steven J. Cooke, and Andy J. Danylchuk. Bridging the knowledge-action gap: a case of research rapidly impacting recreational fisheries policy. Marine Policy, 104(??):210–215, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307978.

Boon:2016:ECT

[BB16] Pei Ya Boon and Maria Beger. The effect of contrasting threat mitigation objectives on spatial conservation priorities. Marine Policy, 68(??):23–29, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000634.

Biondo:2019:MTM

[BB19a] Monica V. Biondo and Rainer P. Burki. Monitoring the trade in marine ornamental fishes through the European Trade Control and Expert System TRACES: Challenges and possibilities. Marine Policy, 108(??):Article 103620, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300193.

Borland:2019:CTT

[BB19b] Meghan E. Borland and Megan Bailey. Corrigendum to “A tale of two standards: a case study of the Fair Trade USA certified Maluku handline yellowfin tuna (Thunnus albacares) fishery” [Marine Policy, Volume 100, 2019, Pages 353–360]. Marine Policy, 101(??):285, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301356.

Borland:2019:TTS

[BB19c] Meghan E. Borland and Megan Bailey. A tale of two standards: a case study of the Fair Trade USA certified Maluku handline yellowfin tuna (Thunnus albacares) fishery. Marine Policy, 100(??):353–360, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302586.

Bronnmann:2019:AAR

[BB19d] Julia Bronnmann and Thomas Bittmann. Asymmetric adjustment of retail cod and herring prices in Germany: a NARDL approach. Marine Policy, 106(??):Article 103513, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308352.

Belwal:2015:FOS

[BBA15] Rakesh Belwal, Shweta Belwal, and Omar Al Jabri. The fisheries of Oman: a situation analysis. Marine Policy, 61(??):237–248, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002183.

Bosch-Belmar:2017:JBP

[BBAP^{+}17] Mar Bosch-Belmar, Ernesto Azzurro, Kristian Pulis, Giacomo Milisenda, Verónica Fuentes, Ons Kéfi-Daly Yahia, Anton Micallef, Alan Deidun, and Stefano Piraino. Jellyfish blooms perception in Mediterranean finfish aquaculture. Marine Policy, 76(??):1–7, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305449.

Byron:2011:ISM

[BBCPC11] Carrie Byron, David Bengtson, Barry Costa-Pierce, and John Calanni. Integrating science into management: Ecological carrying capacity of bivalve shellfish aquaculture. Marine Policy, 35(3):363–370, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001843.

Brouwer:2016:PWP

[BBE^{+}16] R. Brouwer, S. Brouwer, M. A. Eleveld, M. Verbraak, A. J. Wagtendonk, and H. J. van der Woerd. Public willingness to pay for alternative management regimes of remote marine protected areas in the North Sea. Marine Policy, 68(??):195–204, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300823.

Barr:2019:FIP

[BBE19] Rhona Barr, Aaron Bruner, and Scott Edwards. Fisheries Improvement Projects and small-scale fisheries: the need for a modified approach. Marine Policy, 105(??):109–115, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308072.

Bertolotti:2016:ITQ

[BBG^{+}16] María I. Bertolotti, Fabiola Baltar, Patricia Gualdoni, Andrea Pagani, and Lautaro Rotta. Individual transferable quotas in Argentina: Policy and performance. Marine Policy, 71(??):132–137, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303116.

Barreto:2017:RUT

[BBM^{+}17] R. R. Barreto, H. Bornatowski, F. S. Motta, J. Santander-Neto, G. M. S. Vianna, and R. Lessa. Rethinking use and trade of pelagic sharks from Brazil. Marine Policy, 85(??):114–122, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300933.

Ban:2013:SSM

[BBN^{+}13] Natalie C. Ban, Karin M. Bodtker, David Nicolson, Carolyn K. Robb, Krista Royle, and Charlie Short. Setting the stage for marine spatial planning: Ecological and social data collation and analyses in Canada’s Pacific waters. Marine Policy, 39(??):11–20, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002163.

Borger:2014:IES

[BBP^{+}14] Tobias Börger, Nicola J. Beaumont, Linwood Pendleton, Kevin J. Boyle, Philip Cooper, Stephen Fletcher, Tim Haab, Michael Hanemann, Tara L. Hooper, S. Salman Hussain, Rosimeiry Portela, Mavra Stithou, Joanna Stockill, Tim Taylor, and Melanie C. Austen. Incorporating ecosystem services in marine planning: the role of valuation. Marine Policy, 46(??):161–170, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000311.

Banerjee:2018:EBI

[BBR^{+}18] Onil Banerjee, Kevin Boyle, Cassandra T. Rogers, Janice Cumberbatch, Barbara Kanninen, Michele Lemay, and Maja Schling. Estimating benefits of investing in resilience of coastal infrastructure in small island developing states: an application to Barbados. Marine Policy, 90(??):78–87, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300891.

Bjorndal:2016:FPP

[BBRT16] Trond Bjørndal, Ana Brasão, Jorge Ramos, and Amalie Tusvik. Fish processing in Portugal: an industry in expansion. Marine Policy, 72(??):94–106, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630313X.

Bush:2017:PPP

[BBvZ^{+}17] Simon R. Bush, Megan Bailey, Paul van Zwieten, Momo Kochen, Budy Wiryawan, Amanda Doddema, and Stephani C. Mangunsong. Private provision of public information in tuna fisheries. Marine Policy, 77(??):130–135, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303876.

Britton:2013:ASW

[BC13] Easkey Britton and Sarah Coulthard. Assessing the social wellbeing of Northern Ireland’s fishing society using a three-dimensional approach. Marine Policy, 37 (??):28–36, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000723.

Blythe:2017:SPH

[BCE^{+}17] Jessica Blythe, Philippa Cohen, Hampus Eriksson, Joshua Cinner, Delvene Boso, Anne-Maree Schwarz, and Neil Andrew. Strengthening post-hoc analysis of community-based fisheries management through the social-ecological systems framework. Marine Policy, 82(??):50–58, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300684.

Burgess:2018:FRP

[BCM^{+}18] Matthew G. Burgess, Michaela Clemence, Grant R. Mc-Dermott, Christopher Costello, and Steven D. Gaines. Five rules for pragmatic blue growth. Marine Policy, 87 (??):331–339, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307916.

Bell:2018:AMC

[BCMH^{+}18] Johann D. Bell, Andrés Cisneros-Montemayor, Quentin Hanich, Johanna E. Johnson, Patrick Lehodey, Bradley R. Moore, Morgan S. Pratchett, Gabriel Reygondeau, Inna Senina, John Virdin, and Colette C. C. Wabnitz. Adaptations to maintain the contributions of smallscale fisheries to food security in the Pacific Islands. Marine Policy, 88(??):303–314, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300015.

Berger:2019:UOM

[BCP19] Michael F. Berger, Vincent Caruso, and Emily Peterson. An updated orientation to marine conservation funding flows. Marine Policy, 107(??):Article 103497, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304196.

Boonstra:2010:HBL

[BD10] Wiebren J. Boonstra and Nguyen Bach Dang. A history of breaking laws — Social dynamics of non-compliance in Vietnamese marine fisheries. Marine Policy, 34(6):

1261–1267, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001016.

Bundy:2013:KCE

[BD13] Alida Bundy and Anthony Davis. Knowing in context: an exploration of the interface of marine harvesters’ local ecological knowledge with ecosystem approaches to management. Marine Policy, 38(??):277–286, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200142X.

Bennett:2014:WLP

[BD14] Nathan James Bennett and Philip Dearden. Why local people do not support conservation: Community perceptions of marine protected area livelihood impacts, governance and management in Thailand. Marine Policy, 44(??):107–116, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001711.

Bertheussen:2019:NCI

[BD19] Bernt Arne Bertheussen and Bent Magne Dreyer. Is the Norwegian cod industry locked into a value-destructive volume logic? Marine Policy, 103(??):113–120, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307243.

Blanchard:2019:SER

[BDB19] Catherine Blanchard, Carole Durussel, and Ben Boteler. Socio-ecological resilience and the law: Exploring the adaptive capacity of the BBNJ agreement. Marine Policy, 108(??):Article 103612, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303641.

Berkenhagen:2010:DBM

[BDF^{+}10] Jörg Berkenhagen, Ralf Döring, Heino O. Fock, Matthias H. F. Kloppmann, Søren A. Pedersen, and Torsten Schulze. Decision bias in marine spatial planning of offshore wind farms: Problems of singular versus cumulative assessments of economic impacts on fisheries. Marine Policy, 34(3):733–736, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001900.

Bertram:2014:CBA

[BDG^{+}14] Christine Bertram, Thomas Dworak, Stefan Görlitz, Eduard Interwies, and Katrin Rehdanz. Cost-benefit analysis in the context of the EU Marine Strategy Framework Directive: the case of Germany. Marine Policy, 43 (??):307–312, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001437.

Brooker:2018:SCS

[BDH^{+}18] Esther Brooker, Emilie Devenport, Charlotte Rachael Hopkins, Sebastian Hennige, J. Murray Roberts, and Calum Duncan. Scotland as a case study for how benefits of marine ecosystem services may contribute to the commercial fishing industry. Marine Policy, 93(??):271–283, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305619.

Blandon:2019:CFD

[BDHSJ19] Abigayil Blandon, Tim Daw, L. Jamila Haider, and Samantha Stone-Jovicich. Conceptualisations of fisheries development in Eastern Africa over time and between actors. Marine Policy, 107(??):Article 103512, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305074.

Barkin:2018:DSI

[BDIS18] J. Samuel Barkin, Elizabeth R. DeSombre, Atsushi Ishii, and Isao Sakaguchi. Domestic sources of international fisheries diplomacy: a framework for analysis. Marine Policy, 94(??):256–263, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305808.

Bakker:2019:RSC

[BdKvT19] Y. W. Bakker, J. de Koning, and J. van Tatenhove. Resilience and social capital: the engagement of fisheries communities in marine spatial planning. Marine Policy, 99(??):132–139, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301441.

Blasiak:2017:RNN

[BDP^{+}17] Robert Blasiak, Carole Durussel, Jeremy Pittman, Carole-Anne Sénit, Matilda Petersson, and Nobuyuki Yagi. The role of NGOs in negotiating the use of biodiversity in marine areas beyond national jurisdiction. Marine Policy, 81(??):1–8, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300787.

Bohorquez:2019:CGM

[BDP19] John J. Bohorquez, Anthony Dvarskas, and Ellen K. Pikitch. Categorizing global MPAs: a cluster analysis approach. Marine Policy, 108(??):Article 103663, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304439.

Bolton:2016:DDC

[BDSB16] Alexis E. Bolton, Bradford A. Dubik, Joshua S. Stoll, and Xavier Basurto. Describing the diversity of community supported fishery programs in North America. Marine Policy, 66(??):21–29, April 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000129.

Barnes-Dabban:2017:IER

[BDvTvKT17] Harry Barnes-Dabban, Jan P. M. van Tatenhove, Kris C. S. A. van Koppen, and Katrien J. A. M. Termeer. Institutionalizing environmental reform with sense-making:

West and Central Africa ports and the ‘green port’ phenomenon. Marine Policy, 86(??):111–120, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304736.

Boyes:2015:ECN

[BE15] Suzanne J. Boyes and Michael Elliott. The excessive complexity of national marine governance systems

— Has this decreased in England since the introduction of the Marine and Coastal Access Act 2009? Marine Policy, 51(??):57–65, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001936.

Benson:2012:SAU

[Ben12] Catherine S. Benson. Shifting accountabilities? Understanding the connections between national and provincial fisheries in Papua New Guinea. Marine Policy, 36 (4):859–866, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001904.

Benham:2017:APP

[Ben17a] Claudia F. Benham. Aligning public participation with local environmental knowledge in complex marine socialecological systems. Marine Policy, 82(??):16–24, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305322.

Bennett:2017:INS

[Ben17b] Abigail Bennett. The influence of neoliberalization on the success and failure of fishing cooperatives in contemporary small-scale fishing communities: a case study from Yucatán, Mexico. Marine Policy, 80(??):96–106, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303943.

Bennett:2018:NJI

[Ben18] Nathan J. Bennett. Navigating a just and inclusive path towards sustainable oceans. Marine Policy, 97(??): 139–146, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301465.

Bertram:2019:OSS

[Ber19] Douglas F. Bertram. Oil spill settlement funds directed to seabird conservation. Marine Policy, 108(??):Article 103622, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302490.

Bess:2010:MBB

[Bes10] Randall Bess. Maintaining a balance between resource utilisation and protection of the marine environment in New Zealand. Marine Policy, 34(3):690–698, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000023.

Bess:2011:NZT

[Bes11] Randall Bess. New Zealand’s Treaty of Waitangi and the doctrine of discovery: Implications for the foreshore and seabed. Marine Policy, 35(1):85–94, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001478.

Bess:2012:PMN

[Bes12] Randall Bess. Public management in New Zealand and its effect on institutional arrangements for managing fisheries. Marine Policy, 36(2):550–558, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000546.

Beseng:2019:CCW

[Bes19] Maurice Beseng. Cameroon’s choppy waters: the anatomy of fisheries crime in the maritime fisheries sector. Marine Policy, 108(??):Article 103669, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303264.

Barendse:2015:TSN

[BF15] Jaco Barendse and Junaid Francis. Towards a standard nomenclature for seafood species to promote more sustainable seafood trade in South Africa. Marine Policy, 53 (??):180–187, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003340.

Ban:2018:IPR

[BF18] Natalie C. Ban and Alejandro Frid. Indigenous peoples’ rights and marine protected areas. Marine Policy, 87 (??):180–185, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305547.

Burden:2019:BFM

[BF19] Merrick Burden and Rod Fujita. Better fisheries management can help reduce conflict, improve food security, and increase economic productivity in the face of climate change. Marine Policy, 108(??):Article 103610, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308959.

Brennan:2014:CEM

[BFGR14a] Jonathon Brennan, Clare Fitzsimmons, Tim Gray, and Laura Raggatt. Corrigendum to “EU Marine Strategy Framework Directive (MSFD) and Marine Spatial Planning (MSP): Which is the more dominant and practicable contributor to maritime policy in the UK?” [JMPO 43C (2013) 359–366]. Marine Policy, 44(??):478, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002200. See [BFGR14b].

Brennan:2014:EMS

[BFGR14b] Jonathon Brennan, Clare Fitzsimmons, Tim Gray, and Laura Raggatt. EU marine strategy framework directive (MSFD) and marine spatial planning (MSP): Which is the more dominant and practicable contributor to maritime policy in the UK? Marine Policy, 43(??):359–366, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1300153X. See corrigen-

dum [BFGR14a].

Barbeaux:2018:ELF

[BFL18] Steven J. Barbeaux, Lowell Fritz, and Elizabeth Logerwell. Exploring local fishery management through cooperative acoustic surveys in the Aleutian Islands. Marine Policy, 90(??):68–77, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307492.

Bailey:2016:CCI

[BFO^{+}16] Megan Bailey, Brett Favaro, Sarah P. Otto, Anthony Charles, Rodolphe Devillers, Anna Metaxas, Peter Tyedmers, Natalie C. Ban, Taylor Mason, Carie Hoover, Thomas J. Duck, Lucia Fanning, Chris Milley, Andrés M. Cisneros-Montemayor, Daniel Pauly, William W. L. Cheung, Sarika Cullis-Suzuki, Louise Teh, and U. Rashid Sumaila. Canada at a crossroad: the imperative for realigning ocean policy with ocean science. Marine Policy, 63(??):53–60, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002869.

Barbosa-Filho:2016:IBW

[BFTS^{+}16] Márcio Luiz Vargas Barbosa-Filho, Davi Castro Tavares, Salvatore Siciliano, Jailson Fulgencio de Moura, Eraldo Medeiros Costa-Neto, Fabio dos Santos Motta, and Christine Del Vechio Koike. Interactions between whale sharks, Rhincodon typus Smith, 1928 (Orectolobiformes, Rhincodontidae), and Brazilian fisheries: the need for effective conservation measures. Marine Policy, 73(??):210–215, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301385.

Blau:2015:AIN

[BG15] Jason Blau and Lee Green. Assessing the impact of a new approach to ocean management: Evidence to date from five ocean plans. Marine Policy, 56(??):1–8, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000299.

Bertelsen:2016:RCP

[BG16] Rasmus Gjedssø Bertelsen and Vincent Gallucci. The return of China, post-Cold War Russia, and the Arctic: Changes on land and at sea. Marine Policy, 72 (??):240–245, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302214.

Basurko:2015:FFM

[BGA^{+}15] Oihane C. Basurko, Gorka Gabiña, Marga Andrés, Anna Rubio, Ainhize Uriarte, and Iñigo Krug. Fishing for floating marine litter in SE Bay of Biscay: Review and feasibility study. Marine Policy, 61(??):103–112, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002110.

Borggaard:2017:MUA

[BGC^{+}17] Diane L. Borggaard, David M. Gouveia, Mary A. Colligan, Richard Merrick, Kathryn S. Swails, Michael J. Asaro, John Kenney, Glenn Salvador, and John Higgins. Managing U.S. Atlantic large whale entanglements: Four guiding principles. Marine Policy, 84(??):202–212, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630639X.

Bao:2019:CDC

[BGG19] Junlin Bao, Shu Gao, and Jianxiong Ge. Centralization and decentralization: Coastal management pattern changes since the late 19th century, Jiangsu Province, China. Marine Policy, 109(??):Article 103705, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300708.

Busch:2016:CSS

[BGL^{+}16] D. Shallin Busch, Roger Griffis, Jason Link, Karen Abrams, Jason Baker, Russell E. Brainard, Michael Ford, Jonathan A. Hare, Amber Himes-Cornell, Anne Hollowed, Nathan J. Mantua, Sam McClatchie, Michelle Mc-Clure, Mark W. Nelson, Kenric Osgood, Jay O. Peterson, Michael Rust, Vincent Saba, Michael F. Sigler, Seth Sykora-Bodie, Christopher Toole, Eric Thunberg, Robin S. Waples, and Richard Merrick. Climate science strategy of the US National Marine Fisheries Service. Marine Policy, 74(??):58–67, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305723.

Bown:2013:CMA

[BGS13] Natalie K. Bown, Tim S. Gray, and Selina M. Stead. Co-management and adaptive co-management: Two modes of governance in a Honduran marine protected area. Marine Policy, 39(??):128–134, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001947.

Bennett:2015:OG

[BGS15] Nathan James Bennett, Hugh Govan, and Terre Satterfield. Ocean grabbing. Marine Policy, 57(??):61–68, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000755.

Breen:2014:SEP

[BH14a] Benjamin Breen and Stephen Hynes. Shortcomings in the European principles of Integrated Coastal Zone Management (ICZM): Assessing the implications for locally orientated coastal management using Biome Portfolio Analysis (BPA). Marine Policy, 44(??):406–418, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002273.

Bucaram:2014:FIE

[BH14b] Santiago J. Bucaram and Alex Hearn. Factors that influence the entry–exit decision and intensity of participation of fishing fleet for the Galapagos lobster fishery. Marine Policy, 43(??):80–88, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001097.

Bhattacharya:2012:EIC

[Bha12] Syamantak Bhattacharya. The effectiveness of the ISM Code: a qualitative enquiry. Marine Policy, 36(2):528–535, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001497.

Blanchard:2014:HRU

[BHA^{+}14] Anne Blanchard, Kjellrun Hiis Hauge, Gisle Andersen, Jan Helge Fosså, Bjørn Einar Grøsvik, Nils Olav Handegard, Matthias Kaiser, Sonnich Meier, Erik Olsen, and Frode Vikebø. Harmful routines? Uncertainty in science and conflicting views on routine petroleum operations in Norway. Marine Policy, 43(??):313–320, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001425.

Ben-Hasan:2017:AMD

[BHAHW17] A. Ben-Hasan, M. Al-Husaini, and C. Walters. Adaptive management of declining fisheries: When is it worth trying to rebuild stocks through fishery regulation? Marine Policy, 85(??):107–113, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302804.

Byrd:2011:EFE

[BHG11] Barbie L. Byrd, Aleta A. Hohn, and Matthew H. Godfrey. Emerging fisheries, emerging fishery interactions with sea turtles: a case study of the large-mesh gillnet fishery for flounder in Pamlico Sound, North Carolina, USA. Marine Policy, 35(3):271–285, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001715.

Blasiak:2017:PDI

[BHI^{+}17] Robert Blasiak, Julia Hsiang-Wen Huang, Hiroe Ishihara, Ingrid Kelling, Sopha Lieng, Hannah Lindoff, Alastair Macfarlane, Akane Minohara, Yasuyuki Miyakoshi, Herman Wisse, and Nobuyuki Yagi. Promoting diversity and inclusiveness in seafood certification and ecolabelling: Prospects for Asia. Marine Policy, 85(??):42–47, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730502X.

Brechon:2016:SAA

[BHM16] Amanda L. Bréchon, Robert Hanner, and Stefano Mariani. A systematic analysis across North Atlantic countries unveils subtleties in cod product labelling. Marine Policy, 69(??):124–133, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301865.

Bonnevie:2019:AUU

[BHS19] Ida Maria Bonnevie, Henning Sten Hansen, and Lise Schrøder. Assessing use-use interactions at sea: a theoretical framework for spatial decision support tools facilitating co-location in maritime spatial planning. Marine Policy, 106(??):Article 103533, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306729.

Bucaram:2018:AFE

[BHT^{+}18] Santiago J. Bucaram, Alex Hearn, Ana M. Trujillo, Willington Rentería, Rodrigo H. Bustamante, Guillermo Morán, Gunther Reck, and José L. García. Assessing fishing effects inside and outside an MPA: the impact of the Galapagos Marine Reserve on the Industrial pelagic tuna fisheries during the first decade of operation. Marine Policy, 87(??):212–225, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303949.

Bigagli:2015:ELF

[Big15] Emanuele Bigagli. The EU legal framework for the management of marine complex social–ecological systems. Marine Policy, 54(??):44–51, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003261.

Bigagli:2016:ILF

[Big16] Emanuele Bigagli. The international legal framework for the management of the global oceans social-ecological system. Marine Policy, 68(??):155–164, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300860.

Bailey:2013:MBC

[BIPS13] Megan Bailey, Gakushi Ishimura, Richard Paisley, and U. Rashid Sumaila. Moving beyond catch in allocation approaches for internationally shared fish stocks. Marine Policy, 40(??):124–136, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002552.

Bedford:2018:PPC

[BJGMG18] Jacob Bedford, David Johns, Simon Greenstreet, and Abigail McQuatters-Gollop. Plankton as prevailing conditions: a surveillance role for plankton indicators within the Marine Strategy Framework Directive. Marine Policy, 89(??):109–115, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306711.

Bonnichsen:2019:FMT

[BJO19] O. Bonnichsen, C. L. Jensen, and S. B. Olsen. Fishing for more tourists — an empirical investigation of tourist anglers’ preferences for angling site quality. Marine Policy, 106(??):Article 103532, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307091.

Bejder:2016:ECS

[BJS^{+}16] Michelle Bejder, David W. Johnston, Joshua Smith, Ari Friedlaender, and Lars Bejder. Embracing conservation success of recovering humpback whale populations: Evaluating the case for downlisting their conservation status in Australia. Marine Policy, 66(??):137–141, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001402.

Bavinck:2018:FSS

[BJS18] Maarten Bavinck, Svein Jentoft, and Joeri Scholtens. Fisheries as social struggle: a reinvigorated social science research agenda. Marine Policy, 94(??):46–52, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17309028.

Bolman:2018:UMU

[BJvH18] Bas Bolman, Robbert G. Jak, and Luc van Hoof. Unravelling the myth — The use of Decisions Support Systems in marine management. Marine Policy, 87(??):241–249, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730725X.

Biedron:2016:TSU

[BK16] Ingrid S. Biedron and Barbara A. Knuth. Toward shared understandings of ecosystem-based fisheries management among fishery management councils and stakeholders in the U.S. Mid-Atlantic and New England regions. Marine Policy, 70(??):40–48, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301774.

Breen:2016:IPQ

[BKH16] Benjamin Breen, Hugh Kelley, and Stephen Hynes. The impact of precautionary quota constraints on the composition of multispecies harvest portfolios. Marine Policy, 69(??):13–23, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002900.

Bennett:2018:CIC

[BKHA^{+}18] Nathan J. Bennett, Maery Kaplan-Hallam, Gerry Augustine, Natalie Ban, Dyhia Belhabib, Irene Brueckner-Irwin, Anthony Charles, John Couture, Sondra Eger, Lucia Fanning, Paul Foley, Anne Marie Goodfellow, Larry Greba, Edward Gregr, Don Hall, Sarah Harper, Brandon Maloney, Jim McIsaac, Wanli Ou, Evelyn Pinkerton, Darren Porter, Richard Sparrow, Robert Stephenson, Allison Stocks, U. Rashid Sumaila, Tasha Sutcliffe, and Megan Bailey. Coastal and Indigenous community access to marine resources and the ocean: a policy imperative for Canada. Marine Policy, 87(??):186–193, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306413.

Bartram:2010:STB

[BKKN10] Paul K. Bartram, J. John Kaneko, and Katrina Kucey-Nakamura. Sea turtle bycatch to fish catch ratios for differentiating Hawaii longline-caught seafood products. Marine Policy, 34(1):145–149, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000888.

Barrio:2018:EMO

[BL18] Melina Barrio and Maria L. Loureiro. Evaluating management options for a Marine and Terrestrial National Park: Heterogeneous preferences in choice experiments. Marine Policy, 95(??):85–94, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306644.

Blasiak:2015:BER

[Bla15] Robert Blasiak. Balloon effects reshaping global fisheries. Marine Policy, 57(??):18–20, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000627.

Blanchard:2017:FHS

[Bla17] Catherine Blanchard. Fragmentation in high seas fisheries: Preliminary reflections on a global oceans governance approach. Marine Policy, 84(??):327–332, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730249X.

Belhabib:2016:OWA

[BLC16] Dyhia Belhabib, Vicky W. Y. Lam, and William W. L. Cheung. Overview of West African fisheries under climate change: Impacts, vulnerabilities and adaptive responses of the artisanal and industrial sectors. Marine Policy, 71(??):15–28, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302901.

Bentz:2016:SMW

[BLCD16] Julia Bentz, Fernando Lopes, Helena Calado, and Philip Dearden. Sustaining marine wildlife tourism through linking Limits of Acceptable Change and zoning in the Wildlife Tourism Model. Marine Policy, 68(??):100–107, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000683.

Baudrier:2018:OFF

[BLG^{+}18] Jérôme Baudrier, Alain Lefebvre, François Galgani, Claire Saraux, and Mathieu Doray. Optimising French fisheries surveys for marine strategy framework directive integrated ecosystem monitoring. Marine Policy, 94(??):10–19, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301283.

Brinson:2011:DMS

[BLR11] Ayeisha Brinson, Min-Yang Lee, and Barbara Rountree. Direct marketing strategies: the rise of community supported fishery programs. Marine Policy, 35(4):542–548, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000157.

Bjorndal:2015:EAP

[BLR15] Trond Bjørndal, Alena Lappo, and Jorge Ramos. An economic analysis of the Portuguese fisheries sector 1960– 2011. Marine Policy, 51(??):21–30, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001687.

Berque:2013:CBM

[BM13] Joannès Berque and Osamu Matsuda. Coastal biodiversity management in Japanese satoumi. Marine Policy, 39(??):191–200, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002126.

Brewer:2015:TFT

[BM15] T. D. Brewer and K. Moon. Towards a functional typology of small-scale fisheries co-management informed by stakeholder perceptions: a coral reef case study. Marine Policy, 51(??):48–56, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001948.

Brewer:2017:ATL

[BMAG17] Jennifer F. Brewer, Kyle Molton, Robin Alden, and Carla Guenther. Accountability, transformative learning, and alternate futures for New England groundfish catch shares. Marine Policy, 80(??):113–122, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630570X.

Ban:2014:BIS

[BMD^{+}14] Natalie C. Ban, Sara M. Maxwell, Daniel C. Dunn, Alistair J. Hobday, Nicholas J. Bax, Jeff Ardron, Kristina M. Gjerde, Edward T. Game, Rodolphe Devillers, David M. Kaplan, Piers K. Dunstan, Patrick N. Halpin, and Robert L. Pressey. Better integration of sectoral planning and management approaches for the interlinked ecology of the open oceans. Marine Policy, 49(??):127–136, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002832.

Bastari:2016:LMP

[BMF^{+}16] Azzurra Bastari, Fiorenza Micheli, Francesco Ferretti, Antonio Pusceddu, and Carlo Cerrano. Large marine protected areas (LMPAs) in the Mediterranean Sea: the opportunity of the Adriatic Sea. Marine Policy, 68 (??):165–177, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630094X.

Buhl-Mortensen:2017:MEB

[BMGF^{+}17] Lene Buhl-Mortensen, Ibon Galparsoro, Tomás Vega Fernández, Kate Johnson, Giovanni D’Anna, Fabio Badalamenti, Germana Garofalo, Julia Carlström, Joanna Piwowarczyk, Marijn Rabaut, Jan Vanaverbeke, Cor Schipper, Jan van Dalfsen, Vassiliki Vassilopoulou, Yiannis Issaris, Luc van Hoof, Ellen Pecceu, Kris Hostens, Marie Louise Pace, Leyla Knittweis, Vanessa Stelzenmüller, Valentina Todorova, and Valentina Doncheva. Maritime ecosystem-based management in practice: Lessons learned from the application of a generic spatial planning framework in Europe. Marine Policy, 75(??):174–186, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000373.

Butler:2011:PCS

[BMGH11] James R. A. Butler, Stuart J. Middlemas, Isla M. Graham, and Robert N. Harris. Perceptions and costs of seal impacts on Atlantic salmon fisheries in the Moray Firth, Scotland: Implications for the adaptive co-management of sealfishery conflict. Marine Policy, 35(3):317–323, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000179X.

Bartlett:2010:PIP

[BMPV10] C. Y. Bartlett, T. Maltali, G. Petro, and P. Valentine. Policy implications of protected area discourse in the Pacific islands. Marine Policy, 34(1):99–104, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000724.

Barnett:2017:ECN

[BMW17] Allain J. Barnett, Robin A. Messenger, and Melanie G. Wiber. Enacting and contesting neoliberalism in fisheries: the tragedy of commodifying lobster access rights in Southwest Nova Scotia. Marine Policy, 80(??):60–68, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300124.

Broadbent:2019:RPC

[BN19] Ian D. Broadbent and Caroline L. B. Nixon. Refusal of planning consent for the Docking Shoal offshore wind farm: Stakeholder perspectives and lessons learned. Marine Policy, 110(??):Article 103529, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309345.

Borit:2012:EFR

[BO12] Melania Borit and Petter Olsen. Evaluation framework for regulatory requirements related to data recording and traceability designed to prevent illegal, unreported and unregulated fishing. Marine Policy, 36(1):96–102, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000716.

Bodwitch:2017:CNZ

[Bod17] Hekia Bodwitch. Challenges for New Zealand’s individual transferable quota system: Processor consolidation, fisher exclusion, & Māori quota rights. Marine Policy, 80(??):88–95, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307060.

Boonstra:2015:WMG

[BOF^{+}15] W. J. Boonstra, K. M. Ottosen, A. S. A. Ferreira, A. Richter, L. A. Rogers, M. W. Pedersen, A. Kokkalis, H. Bardarson, S. Bonanomi, W. Butler, F. K. Diekert, N. Fouzai, M. Holma, R. E. Holt, K.Ø. Kvile, E. Malanski, J. I. Macdonald, E. Nieminen, G. Romagnoni, M. Snickars, B. Weigel, P. Woods, J. Yletyinen, and J. D Whittington. What are the major global threats and impacts in marine environments? Investigating the contours of a shared perception among marine scientists from the bottom-up. Marine Policy, 60(??):197–201, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001669.

Bohman:2018:LRA

[Boh18] Brita Bohman. Lessons from the regulatory approaches to combat eutrophication in the Baltic Sea region. Marine Policy, 98(??):227–236, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830678X.

Bohman:2019:RCA

[Boh19] Brita Bohman. Regulatory control of adaptive fisheries: Reflections on the implementation of the landing obligation in the EU common fisheries policy. Marine Policy, 110(??):Article 103557, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930329X.

Bola:2017:PSS

[Bol17] Amelia Bola. Potential for sustainable sea transport: a case study of the Southern Lomaiviti, Fiji Islands. Marine Policy, 75(??):260–270, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000464.

Borch:2010:TLN

[Bor10] Trude Borch. Tangled lines in New Zealand’s quota management system: the process of including recreational fisheries. Marine Policy, 34(3):655–662, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001912.

Borges:2018:STA

[Bor18] Lisa Borges. Setting of total allowable catches in the 2013 EU common fisheries policy reform: possible impacts. Marine Policy, 91(??):97–103, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303044.

Barr:2013:OMP

[BP13] Lissa M. Barr and Hugh P. Possingham. Are outcomes matching policy commitments in Australian marine conservation planning? Marine Policy, 42(??):39–48, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000213.

Batista:2019:IDL

[BPB19] Celene Milanés Batista, Cristina I. Pereira, and Camilo M. Botero. Improving a decree law about coastal zone management in a small island developing state: the case of Cuba. Marine Policy, 101(??):93–107, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301830.

Burdon:2017:MRB

[BPBM17] Daryl Burdon, Tavis Potts, Carol Barbone, and Lucas Mander. The matrix revisited: a bird’s-eye view of marine ecosystem service provision. Marine Policy, 77(??):78–89, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306650.

Blasiak:2016:LRE

[BPF^{+}16] Robert Blasiak, Erich Pacheco, Ken Furuya, Christopher D. Golden, Ahmed Riyaz Jauharee, Yoji Natori, Hiroaki Saito, Hussein Sinan, Takehiro Tanaka, Nobuyuki Yagi, and Evonne Yiu. Local and regional experiences with assessing and fostering ocean health. Marine Policy, 71(??):54–59, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302937.

Blomberg:2018:EUE

[BPMY18] Brittany N. Blomberg, Jennifer Beseres Pollack, Paul A. Montagna, and David W. Yoskowitz. Evaluating the U.S. Estuary Restoration Act to inform restoration policy implementation: a case study focusing on oyster reef projects. Marine Policy, 91(??):161–166, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304402.

Banzon:2019:PCB

[BPYC19] Estelle Marina H. Banzon, Maria Isabel O. Padilla, Jonathan Patrick Yan, and Ronald Allan L. Cruz. Proposed categories of bycatch based on an assessment of data from the Anilao Fish Port, Batangas, Philippines. Marine Policy, 100(??):1–7, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305293.

Bertram:2013:EEE

[BR13a] Christine Bertram and Katrin Rehdanz. On the environmental effectiveness of the EU Marine Strategy Framework Directive. Marine Policy, 38(??):25–40, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001042.

Butterworth:2013:RAW

[BR13b] Andrew Butterworth and Mary Richardson. A review of animal welfare implications of the Canadian commercial seal hunt. Marine Policy, 38(??):457–469, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001662.

Butterworth:2014:RAW

[BR14] Andrew Butterworth and Mary Richardson. A review of animal welfare implications of the Canadian commercial seal hunt — a response to critique of paper MP13 172. Marine Policy, 43(??):379–381, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001528.

Bergseth:2018:DCC

[BR18] Brock J. Bergseth and Matthew Roscher. Discerning the culture of compliance through recreational fisher’s perceptions of poaching. Marine Policy, 89(??):132–141, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306772.

Branch:2012:FSF

[Bra12] Trevor A. Branch. FAO’s state of fisheries and aquaculture: Correcting some misrepresentations by Pauly and Froese. Marine Policy, 36(5):1191–1192, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000437.

Brewer:2013:DDA

[Bre13] T. D. Brewer. Dominant discourses, among fishers and middlemen, of the factors affecting coral reef fish distributions in Solomon Islands. Marine Policy, 37(??):245–253, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000942.

Bremer:2014:NRR

[Bre14] Scott Bremer. ‘No right to rubbish’: Mobilising postnormal science for planning Gisborne’s wastewater outfall. Marine Policy, 46(??):22–30, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000037.

Breslow:2015:ANS

[Bre15] Sara Jo Breslow. Accounting for neoliberalism: “social drivers” in environmental management. Marine Policy, 61 (??):420–429, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003194.

Bretos:2017:FLE

[BRM^{+}17] Fernando Bretos, Julia Azanza Ricardo, Félix Moncada, S. Hoyt Peckham, Jorge A. Angulo Valdés, Antonio Diego, and Katie R. Thompson. Fisheries learning exchanges and sea turtle conservation: an effort between Mexico, Cuba and the U.S. to engage Cuban coastal communities in non-consumptive alternative behaviors. Marine Policy, 77(??):227–230, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303062.

Brooks:2010:SDS

[Bro10] Kate Brooks. Sustainable development: Social outcomes of structural adjustments in a South Australian fishery. Marine Policy, 34(3):671–678, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001948.

Brosnan:2011:DAG

[Bro11] Ian G. Brosnan. The diminishing age gap between polar cruisers and their ships: a new reason to codify the IMO Guidelines for ships operating in polar waters and make them mandatory? Marine Policy, 35(2):261–265, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001685.

Brown:2015:RCN

[Bro15] Alistair M. Brown. A reporting compliance of national Fisheries Authority of Papua New Guinea. Marine Policy, 51(??):196–201, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002267.

Bromley:2016:RBF

[Bro16a] Daniel W. Bromley. Rights-based fisheries and contested claims of ownership: Some necessary clarifications. Marine Policy, 72(??):231–236, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302457.

Brown:2016:RCF

[Bro16b] Alistair Brown. Reporting challenges facing the Solomon Islands Ministry of Fisheries & Marine Resources. Marine Policy, 73(??):181–186, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302597.

Brown:2016:SEE

[Bro16c] Christopher James Brown. Social, economic and environmental effects of closing commercial fisheries to enhance recreational fishing. Marine Policy, 73(??):204– 209, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302469.

Brown:2018:FRN

[Bro18] Alistair Brown. The financial reporting of the National Fisheries Corporation of Tuvalu: the case for alternative indigenous reporting mechanisms. Marine Policy, 88 (??):93–100, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306133.

Breen:2012:EAR

[BRR^{+}12] P. Breen, L. A. Robinson, S. I. Rogers, A. M. Knights, G. Piet, T. Churilova, P. Margonski, N. Papadopoulou, E. Akoglu, A. Eriksson, Z. Finenko, V. Fleming-Lehtinen, B. Galil, F. Goodsir, M. Goren, O. Kryvenko, J. M. Leppanen, V. Markantonatou, S. Moncheva, T. Oguz, L. Paltriguera, K. Stefanova, F. Timofte, and F. Thomsen. An environmental assessment of risk in achieving good environmental status to support regional prioritisation of management in Europe. Marine Policy, 36(5):1033–1043, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000097.

Brown:2015:STA

[BRR15] Susie L. Brown, David Reid, and Emer Rogan. Spatial and temporal assessment of potential risk to cetaceans from static fishing gears. Marine Policy, 51(??):267–280, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002425.

Ballesteros:2018:AUP

[BRR18a] Hugo M. Ballesteros and Gonzalo Rodríguez-Rodríguez. “Acceptable’’ and “unacceptable” poachers: Lessons in managing poaching from the Galician shellfish sector. Marine Policy, 87(??):104–110, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304529.

Ballesteros:2018:HMC

[BRR18b] Hugo M. Ballesteros and Gonzalo Rodríguez-Rodríguez. How much in the clan are you? The community as an explanatory factor of the acceptance of poaching in smallscale fisheries. Marine Policy, 97(??):188–196, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301088.

Buglass:2018:EEC

[BRRG^{+}18] Salomé Buglass, Harry Reyes, Jorge Ramirez-González, Tyler D. Eddy, Pelayo Salinas de León, and José Marin Jarrin. Evaluating the effectiveness of coastal no-take zones of the Galapagos Marine Reserve for the red spiny lobster, Panulirus penicillatus. Marine Policy, 88(??): 204–212, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306048.

Busilacchi:2013:RSF

[BRW^{+}13] Sara Busilacchi, Garry R. Russ, Ashley J. Williams, Stephen G. Sutton, and Gavin A. Begg. The role of subsistence fishing in the hybrid economy of an indigenous community. Marine Policy, 37(??):183–191, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000863.

Banks:2010:IMR

[BS10] Simon A. Banks and Greg A. Skilleter. Implementing marine reserve networks: a comparison of approaches in New South Wales (Australia) and New Zealand. Marine Policy, 34(2):197–207, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000955.

Bradley:2018:AEI

[BS18] Melanie Bradley and Alison Swaddling. Addressing environmental impact assessment challenges in Pacific island countries for effective management of deep sea minerals activities. Marine Policy, 95(??):356–362, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303700.

Bourrel:2018:BCC

[BSAT18] Marie Bourrel, Alison Swaddling, Vira Atalifo, and Akuila Tawake. Building in-country capacity and expertise to ensure good governance of the deep sea minerals industry within the Pacific region. Marine Policy, 95(??):372–379, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301240.

Ba:2017:PED

[BSD^{+}17] Aliou Ba, Jörn Schmidt, Moustpha Dème, Kira Lancker, Christian Chaboud, Philippe Cury, Djiga Thiao, Malick Diouf, and Patrice Brehmer. Profitability and economic drivers of small pelagic fisheries in West Africa: a twenty year perspective. Marine Policy, 76(??):152–158, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305462.

Bell:2019:RFS

[BSH^{+}19] Johann D. Bell, Michael K. Sharp, Elizabeth Havice, Michael Batty, Karen E. Charlton, Joanna Russell, Wayne Adams, Kamal Azmi, Alessandro Romeo, Colette C. C. Wabnitz, Neil L. Andrew, Leonard Rodwell, Steve Gu’urau, and Robert Gillett. Realising the food security benefits of canned fish for Pacific Island countries. Marine Policy, 100(??):183–191, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303403.

Bethoney:2017:BBM

[BSKH17] N. David Bethoney, Bradley P. Schondelmeier, Jeff Kneebone, and William S. Hoffman. Bridges to best management: Effects of a voluntary bycatch avoidance program in a mid-water trawl fishery. Marine Policy, 83(??):172–178, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301914.

Belhabib:2019:FAE

[BSL19] Dyhia Belhabib, U. Rashid Sumaila, and Philippe Le Billon. The fisheries of Africa: Exploitation, policy, and maritime security trends. Marine Policy, 101(??):80–92, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18305773.

Butler:2014:SPE

[BSM^{+}14] J. R. A. Butler, T. Skewes, D. Mitchell, M. Pontio, and T. Hills. Stakeholder perceptions of ecosystem service declines in Milne Bay, Papua New Guinea: Is human population a more critical driver than climate change? Marine Policy, 46(??):1–13, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002947.

Borujeni:2015:EAB

[BSMS15] Pezhman Alaei Borujeni, Stein Ivar Steinshamn, Sayed Abolghasem Mortazavi, and Hasan Salehi. Enhance and advance: the benefits of recruitment enhancement in the case of the Iranian Kutum Fishery. Marine Policy, 61 (??):23–32, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001918.

Brooks:2015:SAS

[BSP^{+}15] Kate Brooks, Jacki Schirmer, Sean Pascoe, Lianos Triantafillos, Eddie Jebreen, Toni Cannard, and Cathy M. Dichmont. Selecting and assessing social objectives for Australian fisheries management. Marine Policy, 53(??): 111–122, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003248.

Brown:2019:PEB

[BSP19] Danielle M. Brown, Paul L. Sieswerda, and E. C. M. Parsons. Potential encounters between humpback whales (Megaptera novaeangliae) and vessels in the New York Bight apex, USA. Marine Policy, 106(??):Article 103527, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309758.

Bellquist:2017:IRI

[BSSS17] Lyall Bellquist, Brice Semmens, Stephen Stohs, and Alayna Siddall. Impacts of recently implemented recreational fisheries regulations on the Commercial Passenger Fishing Vessel fishery for Paralabrax sp. in California. Marine Policy, 86(??):134–143, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308120.

Biedenweg:2016:HFI

[BSW16] Kelly Biedenweg, Kari Stiles, and Katharine Wellman. A holistic framework for identifying human wellbeing indicators for marine policy. Marine Policy, 64(??):31–37, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003218.

Boerder:2019:AWW

[BSW19] Kristina Boerder, Laurenne Schiller, and Boris Worm. Not all who wander are lost: Improving spatial protection for large pelagic fishes. Marine Policy, 105(??):80–90, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301866.

Bourrel:2018:CHM

[BTC18] Marie Bourrel, Torsten Thiele, and Duncan Currie. The common of heritage of mankind as a means to assess and advance equity in deep sea mining. Marine Policy, 95(??):311–316, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304729.

Bennett:2017:ACC

[BTO^{+}17] Nathan J. Bennett, Lydia Teh, Yoshitaka Ota, Patrick Christie, Adam Ayers, Jon C. Day, Phil Franks, David Gill, Rebecca L. Gruby, John N. Kittinger, J. Zachary Koehn, Nai‘a. Lewis, John Parks, Marjo Vierros, Tara S. Whitty, Aulani Wilhelm, Kim Wright, Jaime A. Aburto, Elena M. Finkbeiner, Carlos F. Gaymer, Hugh Govan, Noella Gray, Rebecca M. Jarvis, Maery Kaplan-Hallam, and Terre Satterfield. An appeal for a code of conduct for marine conservation. Marine Policy, 81(??):411–418, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300672.

Bush:2013:DTM

[BTOM13] Simon R. Bush, Hilde Toonen, Peter Oosterveer, and Arthur P. J. Mol. The ‘devils triangle’ of MSC certification: Balancing credibility, accessibility and continuous improvement. Marine Policy, 37(??):288–293, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000991.

Bellmann:2016:GTF

[BTS16] Christophe Bellmann, Alice Tipping, and U. Rashid Sumaila. Global trade in fish and fishery products: an overview. Marine Policy, 69(??):181–188, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003899.

Barrowclift:2017:SET

[BTS^{+}17] E. Barrowclift, A. J. Temple, S. Stead, N. S. Jiddawi, and P. Berggren. Social, economic and trade characteristics of the elasmobranch fishery on Unguja Island, Zanzibar, East Africa. Marine Policy, 83(??):128–136, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301276.

Bueger:2015:WMS

[Bue15] Christian Bueger. What is maritime security? Marine Policy, 53(??):159–164, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003327.

Burroughs:2015:FSE

[Bur15] Richard Burroughs. Feedback switching and the evolution of U.S. coastal management. Marine Policy, 53(??):94–100, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14003224.

Burmester:2016:UUD

[Bur16] Brent Burmester. Upgrading or unhelpful? Defiant corporate support for a marine protected area. Marine Policy, 63(??):206–212, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000688.

Bertheussen:2019:SSS

[BV19] Bernt Arne Bertheussen and Terje Vassdal. Strategic sources of superprofit in a well-regulated fishery. Marine Policy, 106(??):Article 103551, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309369.

Boonstra:2018:SMC

[BVB18] Wiebren J. Boonstra, Matilda Valman, and Emma Björkvik. A sea of many colours — How relevant is Blue Growth for capture fisheries in the Global North, and vice versa? Marine Policy, 87(??):340–349, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305857.

Bradley:2015:PPA

[BvPS15] Michael Bradley, Ingrid van Putten, and Marcus Sheaves. The pace and progress of adaptation: Marine climate change preparedness in Australia’s coastal communities. Marine Policy, 53(??):13–20, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002966.

Barry:2011:IRV

[BvRH11] Luke Barry, Tom M. van Rensburg, and Stephen Hynes. Improving the recreational value of Ireland’s coastal resources: a contingent behavioural application. Marine Policy, 35(6):764–771, December 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000108.

Becker-Weinberg:2015:PLR

[BW15] Vasco Becker-Weinberg. Portugal’s legal regime on marine spatial planning and management of the national maritime space. Marine Policy, 61(??):46–53, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001736.

Blasiak:2018:AFA

[BW18a] Robert Blasiak and Colette C. C. Wabnitz. Aligning fisheries aid with international development targets and goals. Marine Policy, 88(??):86–92, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306310.

Blomquist:2018:SPI

[BW18b] Johan Blomquist and Staffan Waldo. Scrapping programmes and ITQs: Labour market outcomes and spill-over effects on non-targeted fisheries in Sweden. Marine Policy, 88(??):41–47, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304219.

Braccini:2019:SSV

[BW19] J. Matias Braccini and Daniela Waltrick. Species-specific at-vessel mortality of sharks and rays captured by demersal longlines. Marine Policy, 99(??):94–98, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301337.

Blasiak:2019:TGT

[BWD^{+}19] Robert Blasiak, Colette C. C. Wabnitz, Tim Daw, Michael Berger, Abigayil Blandon, Gonçalo Carneiro, Beatrice Crona, Mary Frances Davidson, Solène Guggisberg, Jeremy Hills, Felix Mallin, Edmund McManus, Karim ould Chih, Jeremy Pittman, Xose Santos, Lena Westlund, Hanna Wetterstrand, and Kai Wiegler. Towards greater transparency and coherence in funding for sustainable marine fisheries and healthy oceans. Marine Policy, 107(??):Article 103508, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930274X.

Brillant:2017:TOP

[BWRT17] Sean W. Brillant, Tonya Wimmer, Robert W. Rangeley, and Christopher T. Taggart. A timely opportunity to protect North Atlantic right whales in Canada. Marine Policy, 81(??):160–166, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307989.

Blasiak:2016:SIA

[BY16] Robert Blasiak and Nobuyuki Yagi. Shaping an international agreement on marine biodiversity beyond areas of national jurisdiction: Lessons from high seas fisheries. Marine Policy, 71(??):210–216, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300021.

Buchan:2019:SDP

[BY19] P. M. Buchan and K. L. Yates. Stakeholder dynamics, perceptions and representation in a regional coastal partnership. Marine Policy, 101(??):125–136, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308254.

Blasiak:2015:IHS

[BYK15a] Robert Blasiak, Nobuyuki Yagi, and Hisashi Kurokura. Impacts of hegemony and shifts in dominance on marine capture fisheries. Marine Policy, 52(??):52–58, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002930.

Blasiak:2015:MES

[BYK^{+}15b] Robert Blasiak, Nobuyuki Yagi, Hisashi Kurokura, Kaoru Ichikawa, Kazumi Wakita, and Aimee Mori. Marine ecosystem services: Perceptions of indispensability and pathways to engaging citizens in their sustainable use. Marine Policy, 61(??):155–163, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002249.

Carruthers:2016:USD

[CA16] Thomas R. Carruthers and David J. Agnew. Using simulation to determine standard requirements for recovery rates of fish stocks. Marine Policy, 73(??):146–153, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300173.

Colwell:2017:SEI

[CA17] Julia M. Novak Colwell and Mark Axelrod. Socioeconomic impacts of a closed fishing season on resourcedependent stakeholders in Tamil Nadu, India: Differences in income and expenditure effects by occupational group. Marine Policy, 77(??):182–190, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305231.

Camp:2018:ATD

[CACL18] Edward V. Camp, Robert N. M. Ahrens, Chelsey Crandall, and Kai Lorenzen. Angler travel distances: Implications for spatial approaches to marine recreational fisheries governance. Marine Policy, 87(??):263–274, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308223.

Chapman:2016:WAL

[CAGH16] Jennifer K. Chapman, Lucy G. Anderson, Charlotte L. A. Gough, and Alasdair R. Harris. Working up an appetite for lionfish: a market-based approach to manage the invasion of Pterois volitans in Belize. Marine Policy, 73(??):256–262, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304857.

Campbell:2015:SAF

[Cam15] Michael Campbell. The status of artisanal fishing in El Salvador. Marine Policy, 52(??):33–37, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400270X.

Campling:2016:TPG

[Cam16] Liam Campling. Trade politics and the global production of canned tuna. Marine Policy, 69(??):220–228, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000506.

Capistrano:2010:RAW

[Cap10] Robert Charles G. Capistrano. Reclaiming the ancestral waters of indigenous peoples in the Philippines: the Tagbanua experience with fishing rights and indigenous rights. Marine Policy, 34(3):453–460, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900147X.

Clark:2015:EBE

[CAP15] Nichola A. Clark, Jeff A. Ardron, and Linwood H. Pendleton. Evaluating the basic elements of transparency of regional fisheries management organizations. Marine Policy, 57(??):158–166, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000524.

Carneiro:2011:MMH

[Car11] Gonçalo Carneiro. Marine management for human development: a review of two decades of scholarly evidence. Marine Policy, 35(3):351–362, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001831.

Carneiro:2013:EMS

[Car13a] Gonçalo Carneiro. Evaluation of marine spatial planning. Marine Policy, 37(??):214–229, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000917.

Carothers:2013:SUH

[Car13b] Courtney Carothers. A survey of US halibut IFQ holders: Market participation, attitudes, and impacts. Marine Policy, 38(??):515–522, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001741.

Carter:2014:TSF

[Car14] Caitríona Carter. The transformation of Scottish fisheries: Sustainable interdependence from ‘net to plate’. Marine Policy, 44(??):131–138, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001681.

Carothers:2015:FPS

[Car15]  Courtney Carothers.  Fisheries privatization, social

transitions, and well-being in Kodiak, Alaska. Marine Policy, 61(??):313–322, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003200.

Carr:2019:SSC

[Car19] Liam M. Carr. Seeking stakeholder consensus within Ireland’s conflicted salmon aquaculture space. Marine Policy, 99(??):201–212, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300678.

Calado:2013:PMS

[CB13] H. Calado and J. Bentz. The Portuguese maritime spatial plan. Marine Policy, 42(??):325–333, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000900.

Cavallo:2019:IAI

[CBE^{+}19] Marianna Cavallo, Ángel Borja, Michael Elliott, Victor Quintino, and Julia Touza. Impediments to achieving integrated marine management across borders: the case of the EU Marine Strategy Framework Directive. Marine Policy, 103(??):68–73, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306146.

Cinner:2012:IDC

[CBF^{+}12] J. E. Cinner, Xavier Basurto, Pedro Fidelman, John Kuange, Rachael Lahari, and Ahmad Mukminin. Institutional designs of customary fisheries management arrangements in Indonesia, Papua New Guinea, and Mexico. Marine Policy, 36(1):278–285, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001072.

Cisse:2014:STS

[CBG14] Abdoul Ahad Cissé, Fabian Blanchard, and Olivier Guyader. Sustainability of tropical small-scale fisheries: Integrated assessment in French Guiana. Marine Policy, 44(??):397–405, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002285.

Christie:2017:WPM

[CBG^{+}17] Patrick Christie, Nathan J. Bennett, Noella J. Gray, T. ‘Aulani Wilhelm, Nai‘a Lewis, John Parks, Natalie C. Ban, Rebecca L. Gruby, Lindsay Gordon, Jon Day, Sue Taei, and Alan M. Friedlander. Why people matter in ocean governance: Incorporating human dimensions into large-scale marine protected areas. Marine Policy, 84(??):273–284, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300532.

Chan:2018:ERF

[CBL18] Maggie N. Chan, Anne H. Beaudreau, and Philip A. Loring. Evaluating the recreational fishery management toolbox: Charter captains’ perceptions of harvest controls, limited access, and quota leasing in the guided halibut fishing sector in Alaska. Marine Policy, 91(??):129–135, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730739X.

Calado:2012:NIM

[CBN^{+}12] Helena Calado, Julia Bentz, Kiat Ng, Anna Zivian, Nicole Schaefer, Colin Pringle, David Johnson, and Mike Phillips. NGO involvement in marine spatial planning: a way forward? Marine Policy, 36(2):382–388, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001242.

Copello:2016:EPI

[CBP^{+}16] Sofía Copello, Gabriela S. Blanco, Juan Pablo Seco Pon, Flavio Quintana, and Marco Favero. Exporting the problem: Issues with fishing closures in seabird conservation. Marine Policy, 74(??):120–127, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300793.

Creed:2018:MTS

[CBP^{+}18] R. Creed, B. Baily, J. Potts, M. Bray, and R. Austin. Moving towards sustainable coasts: a critical evaluation of a stakeholder engagement group in successfully delivering the mechanism of adaptive management. Marine Policy, 90(??):184–193, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303871.

Cashion:2019:OCD

[CBP19] Madeline S. Cashion, Nicolas Bailly, and Daniel Pauly. Official catch data underrepresent shark and ray taxa caught in Mediterranean and Black Sea fisheries. Marine Policy, 105(??):1–9, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303099.

Crona:2016:TTI

[CBS^{+}16] Beatrice I. Crona, Xavier Basurto, Dale Squires, Stefan Gelcich, Tim M. Daw, Ahmed Khan, Elizabeth Havice, Victoria Chomo, Max Troell, Eny A. Buchary, and Edward H. Allison. Towards a typology of interactions between small-scale fisheries and global seafood trade. Marine Policy, 65(??):1–10, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003607.

Castrejon:2013:IFC

[CC13] Mauricio Castrejón and Anthony Charles. Improving fisheries co-management through ecosystem-based spatial management: the Galapagos Marine Reserve.

Marine Policy, 38(??):235–245, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001388.

Chambers:2017:TYA

[CC17a] Catherine Chambers and Courtney Carothers. Thirty years after privatization: a survey of Icelandic small-boat fishermen. Marine Policy, 80(??):69–80, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300136.

Chen:2017:TBT

[CC17b] Chung-Ling Chen and Ya-Chiao Chang. A transition beyond traditional fisheries: Taiwan’s experience with developing fishing tourism. Marine Policy, 79(??):84–91, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301045.

Costa:2018:BAO

[CC18] Sónia Costa and Rui Caldeira. Bibliometric analysis of ocean literacy: an underrated term in the scientific literature. Marine Policy, 87(??):149–157, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305109.

Collins:2013:PEI

[CCC^{+}13] Patrick Colman Collins, Peter Croot, Jens Carlsson, Ana Colaço, Anthony Grehan, Kiseong Hyeong, Robert Kennedy, Christian Mohn, Samantha Smith, Hiroyuki Yamamoto, and Ashley Rowden. A primer for the Environmental Impact Assessment of mining at seafloor massive sulfide deposits. Marine Policy, 42(??):198–209, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000298.

Chen:2014:DCM

[CCC^{+}14] Jyun-Long Chen, Jun-Yu Chen, Ching-Ta Chuang, Hsueh-Jung Lu, Hsiang-Hsi Liu, and Yu-Shang Lin. Developing a co-management financing mechanism to enhance the financial sustainability of marine protected areas in Taiwan. Marine Policy, 48(??):126–133, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000943.

Chalcobsky:2017:WWP

[CCC17] Bárbara Ailén Chalcobsky, Enrique Alberto Crespo, and Mariano Alberto Coscarella. Whale-watching in Patagonia: What regulation scheme should be implemented when the socio-ecological system is changing? Marine Policy, 75(??):165–173, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304596.

Chion:2013:SMP

[CCD^{+}13] Clément Chion, Guy Cantin, Suzan Dionne, Benoit Dubeau, Philippe Lamontagne, Jacques-André Landry, Danielle Marceau, Cristiane C. A. Martins, Nadia Ménard, Robert Michaud, Lael Parrott, and Samuel Turgeon. Spatiotemporal modelling for policy analysis: Application to sustainable management of whale-watching activities. Marine Policy, 38(??):151–162, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001297.

Chrysafi:2019:EEK

[CCK19] Anna Chrysafi, Jason M. Cope, and Anna Kuparinen. Eliciting expert knowledge to inform stock status for datalimited stock assessments. Marine Policy, 101(??):167–176, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307352.

Chen:2019:SMD

[CCL^{+}19] Chung-Ling Chen, Ya-Chieh Chuang, Tuey-Chih Lee, Chien-Ho Liu, and Cheng-Lang Yang. Not out of sight but out of mind: Developing a multi-criteria evaluation structure for green fishery harbors. Marine Policy, 100(??):324–331, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302392.

Calhoun:2016:AVW

[CCR16] Sarah Calhoun, Flaxen Conway, and Suzanne Russell. Acknowledging the voice of women: implications for fisheries management and policy. Marine Policy, 74(??):292–299, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301993.

Collins:2016:ESD

[CCRT16] Patrick C. Collins, Jens Carlsson, Petrina Rowcroft, and Brian Tibbles. Ecosystem status of the deep Black Sea, soft sediment, benthic community. Marine Policy, 73(??):216–223, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304833.

Chan:2014:FRT

[CCS14] Valerie Chan, Raymond Clarke, and Dale Squires. Full retention in tuna fisheries: Benefits, costs and unintended consequences. Marine Policy, 45(??):213–221, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002418.

Clay:2016:SBR

[CCS16] Patricia M. Clay, Lisa L. Colburn, and Tarsila Seara. Social bonds and recovery: an analysis of Hurricane Sandy in the first year after landfall. Marine Policy, 74(??):334–340, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302500.

Cabral:2013:CSI

[CCTG^{+}13] Reniel Cabral, Annabelle Cruz-Trinidad, Rollan Geronimo, Lydia Napitupulu, Paul Lokani, Delvene Boso, Christine Marie Casal, Nurulhuda Ahmad Fatan, and Porfirio Aliño. Crisis sentinel indicators: Averting a potential meltdown in the Coral Triangle. Marine Policy, 39(??):241–247, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002114.

Chaigneau:2015:IVL

[CD15] Tomas Chaigneau and Tim M. Daw. Individual and villagelevel effects on community support for Marine Protected Areas (MPAs) in the Philippines. Marine Policy, 51 (??):499–506, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002231.

Cardinale:2013:REF

[CDA^{+}13] M. Cardinale, H. Dörner, A. Abella, J. L. Andersen, J. Casey, R. Döring, E. Kirkegaard, A. Motova, J. Anderson, E. J. Simmonds, and C. Stransky. Rebuilding EU fish stocks and fisheries, a process under way? Marine Policy, 39(??):43–52, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200200X.

Corral:2017:PAF

[CdL17] Serafin Corral and David Romero Manrique de Lara. Participatory artisanal fisheries management in islands: Application to the Canary Islands (Spain). Marine Policy, 81(??):45–52, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306601.

Caputi:2018:EBF

[CdLH^{+}18] Nick Caputi, Simon de Lestang, Jason How, Fabian Trinnie, and Warrick (Rick) Fletcher. Ecosystem-based fisheries management (or ‘triple bottom line’) assessments of the western rock lobster resource: Is there an optimal target for fishing? Marine Policy, 94(??):264–274, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305766.

Christensen:2014:VSP

[CdlPS^{+}14] Villy Christensen, Santiago de la Puente, Juan Carlos Sueiro, Jeroen Steenbeek, and Patricia Majluf. Valuing seafood: the Peruvian fisheries sector. Marine Policy, 44(??):302–311, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002194.

Caputi:2015:MEY

[CdLR^{+}15] Nick Caputi, Simon de Lestang, Chris Reid, Alex Hesp, and Jason How. Maximum economic yield of the western rock lobster fishery of Western Australia after moving from effort to quota control. Marine Policy, 51(??):452–464, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002681.

Chircop:2016:OLR

[CDO16] Aldo Chircop, David Dzidzornu, and Chidi Oguamanam. Ocean law reform: a multi-level comparative law analysis of Nigerian maritime zone legislation. Marine Policy, 67(??):60–75, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000178.

Coomber:2016:DVT

[CDR^{+}16] Frazer Guy Coomber, Massimo D’Incà, Massimiliano Rosso, Paola Tepsich, Giuseppe Notarbartolo di Sciara, and Aurélie Moulins. Description of the vessel traffic within the north Pelagos Sanctuary: Inputs for Marine Spatial Planning and management implications within an existing international Marine Protected Area. Marine Policy, 69(??):102–113, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301853.

Crilly:2013:SVL

[CE13] R. Crilly and A. Esteban. Small versus large-scale, multi-fleet fisheries: the case for economic, social and environmental access criteria in European fisheries. Marine Policy, 37(??):20–27, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000875.

Clarke:2017:GLG

[CE17] Maurice Clarke and Afra Egan. Good luck or good governance? The recovery of Celtic Sea herring. Marine Policy, 78(??):163–170, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305796.

Clifton:2012:MCP

[CEB^{+}12] Julian Clifton, Michelle Etienne, David K. A. Barnes, Richard S. K. Barnes, David J. Suggett, and David J. Smith. Marine conservation policy in Seychelles: Current constraints and prospects for improvement. Marine Policy, 36(3):823–831, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001941.

Cavallo:2016:ARC

[CETQ16] Marianna Cavallo, Michael Elliott, Julia Touza, and Victor Quintino. The ability of regional coordination and policy integration to produce coherent marine management: Implementing the Marine Strategy Framework Directive in the North-East Atlantic. Marine Policy, 68(??):108–116, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000658.

Cavallo:2017:BII

[CETQ17] Marianna Cavallo, Michael Elliott, Julia Touza, and Victor Quintino. Benefits and impediments for the integrated and coordinated management of European seas. Marine Policy, 86(??):206–213, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300982.

Cohen:2013:SSS

[CF13] Philippa J. Cohen and Simon J. Foale. Sustaining small-scale fisheries with periodically harvested marine reserves. Marine Policy, 37(??):278–287, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200098X.

Cote:2016:SVS

[CF16] Isabelle M. Côté and Corinna Favaro. The scientific value of scientific whaling. Marine Policy, 74(??):88–90, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303074.

Clifton:2017:EIP

[CF17] Julian Clifton and Simon Foale. Extracting ideology from policy: Analysing the social construction of conservation priorities in the Coral Triangle region. Marine Policy, 82(??):189–196, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730163X.

Cramer:2018:GFP

[CFC^{+}18] Lori A. Cramer, Courtney Flathers, Deanna Caracciolo, Suzanne M. Russell, and Flaxen Conway. Graying of the fleet: Perceived impacts on coastal resilience and local policy. Marine Policy, 96(??):27–35, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308631.

Cardwell:2013:TQE

[CG13] Emma Cardwell and Robert Gear. Transferable quotas, efficiency and crew ownership in Whalsay, Shetland. Marine Policy, 40(??):160–166, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000225.

Cabral:2018:HIC

[CG18] Reniel B. Cabral and Rollan C. Geronimo. How important are coral reefs to food security in the Philippines? Diving deeper than national aggregates and averages. Marine Policy, 91(??):136–141, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308357.

Campbell:2019:AEV

[CG19] Lisa M. Campbell and Noella J. Gray. Area expansion versus effective and equitable management in international marine protected areas goals and targets. Marine Policy, 100(??):192–199, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302331.

Condie:2014:ISF

[CGC14] H. M. Condie, A. Grant, and T. L. Catchpole. Incentivising selective fishing under a policy to ban discards; lessons from European and global fisheries. Marine Policy, 45(??):287–292, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300198X.

Cheilari:2013:EFP

[CGDB13] Anna Cheilari, Jordi Guillen, Dimitrios Damalas, and Thomas Barbas. Effects of the fuel price crisis on the energy efficiency and the economic performance of the European Union fishing fleets. Marine Policy, 40(??):18–24, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002473.

Chang:2014:MEG

[CGF14] Yen-Chiang Chang, Warwick Gullett, and David L. Fluharty. Marine environmental governance networks and approaches: Conference report. Marine Policy, 46 (??):192–196, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000323.

Coccoli:2018:CAR

[CGM^{+}18] Caroline Coccoli, Ibon Galparsoro, Arantza Murillas, Kemal Pınarbaşı, and Jose A. Fernandes. Conflict analysis and reallocation opportunities in the framework of marine spatial planning: a novel, spatially explicit Bayesian belief network approach for artisanal fishing and aquaculture. Marine Policy, 94(??):119–131, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301635.

Cashion:2018:FSW

[CGP^{+}18] Tim Cashion, Sarah M. Glaser, Lo Persson, Paige M. Roberts, and Dirk Zeller. Fisheries in Somali waters: Reconstruction of domestic and foreign catches for 1950– 2015. Marine Policy, 87(??):275–283, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305717.

Chu:2017:IEF

[CGS^{+}17] Jingjie Chu, Taryn M. Garlock, Patrick Sayon, Frank Asche, and James L. Anderson. Impact evaluation of a fisheries development project. Marine Policy, 85(??): 141–149, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730564X.

Caveen:2013:MPW

[CGSP13] Alex J. Caveen, Tim S. Gray, Selina M. Stead, and Nicholas V. C. Polunin. MPA policy: What lies behind the science? Marine Policy, 37(??):3–10, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000668.

Cardinale:2012:SMS

[CH12] Massimiliano Cardinale and Joakim Hjelm. Size matters: Short term loss and long term gain in a size-selective fishery. Marine Policy, 36(4):903–906, July 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000024.

Carr:2016:TFD

[CH16] Liam M. Carr and William D. Heyman. Testing fisherdeveloped alternatives to fishery management tools for community support and regulatory effectiveness. Marine Policy, 67(??):40–53, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000403.

Chang:2011:AVM

[Cha11] Shui-Kai Chang. Application of a vessel monitoring system to advance sustainable fisheries management — Benefits received in Taiwan. Marine Policy, 35(2):116–121, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001491.

Chang:2014:CLL

[Cha14] Shui-Kai Chang. Constructing logbook-like statistics for coastal fisheries using coastal surveillance radar and fish market data. Marine Policy, 43(??):338–346, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001450.

Chang:2015:PCB

[Cha15] Shui-Kai Chang. Precious corals become more precious in the northwestern Pacific: Urgent need for integrated policy. Marine Policy, 52(??):103–107, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002826.

Chang:2016:SEE

[Cha16] Shui-Kai Chang. From subsidy evaluation to effort estimation: Advancing the function of voyage data recorders for offshore trawl fishery management. Marine Policy, 74 (??):99–107, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303232.

Chandler:2017:SOP

[Cha17a] John Chandler. Stewardship of offshore petroleum: Where is the value? Marine Policy, 81(??):64–70, July 2017.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300222.

Charles:2017:MSB

[Cha17b] Kevin Charles. Marine science and blue growth: Assessing the marine academic production of 123 cities and territories worldwide. Marine Policy, 84(??):119–129, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630762X.

Chang:2019:SCE

[Cha19] Yen-Chiang Chang. The Sino–Canadian Exchange on the Arctic: Conference report. Marine Policy, 99(??): 76–79, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306250.

Carrasquilla-Henao:2019:MFR

[CHBRJ19] Mauricio Carrasquilla-Henao, Natalie Ban, Mario Rueda, and Francis Juanes. The mangrove-fishery relationship: a local ecological knowledge perspective. Marine Policy, 108(??):Article 103656, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302374.

Collins:2019:IIE

[CHDJ^{+}19] Jane Eva Collins, Harriet Harden-Davies, Marcel Jaspars, Torsten Thiele, Thomas Vanagt, and Isabelle Huys. Inclusive innovation: Enhancing global participation in and benefit sharing linked to the utilization of marine genetic resources from areas beyond national jurisdiction. Marine Policy, 109(??):Article 103696, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301393.

Chen:2010:FIP

[Che10] Chung-Ling Chen. Factors influencing participation of ‘topdown but voluntary’ fishery management — Empirical evidence from Taiwan. Marine Policy, 34(1):150–155, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0900089X.

Chen:2011:CWM

[Che11] Chung-Ling Chen. From catching to watching: Moving towards quality assurance of whale/dolphin watching tourism in Taiwan. Marine Policy, 35(1):10–17, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000134X.

Chen:2012:MPC

[Che12a] Chih-Shin Chen. Management of the precious coral fishery in Taiwan: Progress and perspectives. Marine Policy, 36(3):623–629, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001667.

Chen:2012:TRI

[Che12b] Chung-Ling Chen. Taiwan’s response to international fisheries management after 2005 as influenced by ICCAT and fishers’ perception. Marine Policy, 36(2):350–357, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001151.

Chen:2012:UBT

[Che12c] Chung-Ling Chen. Unfinished business: Taiwan’s experience with rights-based coastal fisheries management. Marine Policy, 36(5):955–962, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000164.

Cavanagh:2016:SPE

[CHKG16] Rachel D. Cavanagh, Simeon L. Hill, Cheryl A. Knowland, and Susie M. Grant. Stakeholder perspectives on ecosystem-based management of the Antarctic krill fishery. Marine Policy, 68(??):205–211, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300872.

Cosquer:2019:RUK

[CHL^{+}19] Alix Cosquer, Michael Hughes, Nicolas Le Corre, Aude Saint-Pierre, Ingrid Peuziat, Thierry Michot, and Nicolas Bernard. Recreation user knowledge, support and engagement in French MPAs: Are there reverse sideeffects of the French soft regulation and management approach? Marine Policy, 104(??):108–117, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830160X.

Cho:2010:LCF

[Cho10] Dong-Oh Cho. Limitations of 1992 CLC/FC and enactment of the Special Law on M/V Hebei Spirit incident in Korea. Marine Policy, 34(3):447–452, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001468.

Cho:2011:RDF

[Cho11] Dong-Oh Cho. Removing derelict fishing gear from the deep seabed of the East Sea. Marine Policy, 35(5): 610–614, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000236.

Cho:2012:EIB

[Cho12] Dong-Oh Cho. Eliminating illegal bottom trawl fishing in the coastal waters of Korea. Marine Policy, 36(2):321–326, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1100114X.

Clark:2017:DNE

[CHSW17] Robert Clark, John Humphreys, Jean-Luc Solandt, and Catherine Weller. Dialectics of nature: the emergence of policy on the management of commercial fisheries in English European Marine Sites. Marine Policy, 78(??):11–17, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306819.

Chung:2010:SRG

[Chu10] Suh-Yong Chung. Strengthening regional governance to protect the marine environment in Northeast Asia: From a fragmented to an integrated approach. Marine Policy, 34(3):549–556, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001742.

Cvitanovic:2015:OBK

[CHvKM15] C. Cvitanovic, A. J. Hobday, L. van Kerkhoff, and N. A. Marshall. Overcoming barriers to knowledge exchange for adaptive resource management; the perspectives of Australian marine scientists. Marine Policy, 52 (??):38–44, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002887.

Curtis:2016:WCM

[CJ16] Hazel Curtis and Edward Jones. Will I clear my feet? Perspectives on a fishing vessel scrapping scheme in Scotland. Marine Policy, 71(??):94–105, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303013.

Christiansen:2017:DNG

[CJ17] Erling A. N. Christiansen and Stig-Erik Jakobsen. Diversity in narratives to green the Norwegian salmon farming industry. Marine Policy, 75(??):156–164, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304365.

Cook:2015:APC

[CJB15] Taylor C. Cook, Kira James, and Maddalena Bearzi. Angler perceptions of California sea lion (Zalophus californianus) depredation and marine policy in Southern California. Marine Policy, 51(??):573–583, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400253X.

Cronin:2014:FSI

[CJHR14] M. Cronin, M. Jessopp, J. Houle, and D. Reid. Fisheryseal interactions in Irish waters: Current perspectives and future research priorities. Marine Policy, 44(??): 120–130, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001693.

Colburn:2016:ICC

[CJW^{+}16] Lisa L. Colburn, Michael Jepson, Changhua Weng, Tarsila Seara, Jeremy Weiss, and Jonathan A. Hare. Indicators of climate change and social vulnerability in fishing dependent communities along the Eastern and Gulf Coasts of the United States. Marine Policy, 74(??):323–333, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302123.

Cummins:2011:ERI

[CK11] Valerie Anne Cummins and Kathrin Kopke. Editorial: Reflections from Ireland. Marine Policy, 35(6):737–738, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1100008X.

Curtin:2018:MMS

[CK18] Richard Curtin and Michael Keatinge. A methodology to measure the social impact of the EU quota setting procedure. Marine Policy, 95(??):248–255, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305481.

Cowling:2014:ESS

[CKBS14] M. Cowling, R. Kirkwood, L. J. Boren, and C. Scarpaci. The effects of seal-swim activities on the New Zealand fur seal (Arctophoca australis forsteri) in the Bay of Plenty, New Zealand, and recommendations for a sustainable tourism industry. Marine Policy, 45(??):39–44, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002522.

Collins:2013:VDC

[CKC^{+}13] Patrick Colman Collins, Bob Kennedy, Jon Copley, Rachel Boschen, Nicholas Fleming, James Forde, Se-Jong Ju, Dhugal Lindsay, Leigh Marsh, Verity Nye, Adrian Patterson, Hirome Watanabe, Hiroyuki Yamamoto, Jens Carlsson, and Andrew David Thaler. VentBase: Developing a consensus among stakeholders in the deep-sea regarding environmental impact assessment for deep-sea mining–A workshop report. Marine Policy, 42(??):334–336, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000705.

Clay:2014:MSE

[CKdS14] Patricia M. Clay, Andrew Kitts, and Patricia Pinto da Silva. Measuring the social and economic performance of catch share programs: Definition of metrics and application to the U.S. Northeast Region groundfish fishery. Marine Policy, 44(??):27–36, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001632.

Cullen-Knox:2019:PSM

[CKFLO19] C. Cullen-Knox, A. Fleming, L. Lester, and E. Ogier. Publicised scrutiny and mediatised environmental conflict: the case of Tasmanian salmon aquaculture. Marine Policy, 100(??):307–315, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304913.

Cullen-Knox:2017:SLO

[CKHJ^{+}17] Coco Cullen-Knox, Marcus Haward, Julia Jabour, Emily Ogier, and Sean R. Tracey. The social licence to operate and its role in marine governance: Insights from Australia. Marine Policy, 79(??):70–77, May 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630642X.

Chapsos:2019:ILF

[CKN19] Ioannis Chapsos, Juliette Koning, and Math Noortmann. Involving local fishing communities in policy making: Addressing Illegal fishing in Indonesia. Marine Policy, 109(??):Article 103708, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300399.

Carpenter:2016:LBI

[CKVO16] Griffin Carpenter, Richard Kleinjans, Sebastian Villasante, and Bethan C. O’Leary. Landing the blame: the influence of EU Member States on quota setting. Marine Policy, 64(??):9–15, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003206.

Campbell:2013:CMA

[CKY^{+}13] Stuart J. Campbell, Tasrif Kartawijaya, Irfan Yulianto, Rian Prasetia, and Julian Clifton. Co-management approaches and incentives improve management effectiveness in the Karimunjawa National Park, Indonesia. Marine Policy, 41(??):72–79, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002631.

Chen:2013:FGD

[CL13] Chung-Ling Chen and Ta-Kang Liu. Fill the gap: Developing management strategies to control garbage pollution from fishing vessels. Marine Policy, 40(??):34–40, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000092.

Call:2015:TPP

[CL15] Isabel L. Call and Daniel K. Lew. Tradable permit programs: What are the lessons for the new Alaska halibut catch sharing plan? Marine Policy, 52(??):125–137, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002760.

Chang:2016:IMS

[CL16a] Yi Chang and Bo-Han Lin. Improving marine spatial planning by using an incremental amendment strategy: the case of Anping, Taiwan. Marine Policy, 68(??):30–38, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000488.

Chung:2016:COS

[CL16b] Suh-Yong Chung and Goueun Lee. Combating oil spill accidents in Northeast Asia: a case of the NOWPAP and Hebei Spirit oil spill. Marine Policy, 72(??):14–20, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303724.

Choi:2018:WPR

[CL18] Eun Chul Choi and Joo Suk Lee. The willingness to pay for removing the microplastics in the ocean

— The case of Seoul metropolitan area, South Korea. Marine Policy, 93(??):93–100, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304803.

Chang:2019:DSO

[CL19] Yen-Chiang Chang and Xiuhua Li. The disappearance of the State Oceanic Administration in China? — Current developments. Marine Policy, 107(??):Article 103588, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300144.

Clapham:2015:JWF

[Cla15] Phillip J. Clapham. Japan’s whaling following the International Court of Justice ruling: Brave new world — or business as usual? Marine Policy, 51(??):238–241, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002279.

Cristini:2016:CVM

[CLC^{+}16] Luisa Cristini, Richard S. Lampitt, Vanessa Cardin, Eric Delory, Peter Haugan, Nick O’Neill, George Petihakis, and Henry A. Ruhl. Cost and value of multidisciplinary fixed-point ocean observatories. Marine Policy, 71(??):138–146, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300380.

Caveen:2017:RAS

[CLDP17] A. J. Caveen, W. Lart, H. Duggan, and T. Pickerell. The Risk Assessment for Sourcing Seafood (RASS): Empowering businesses to buy responsibly. Marine Policy, 75 (??):1–10, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302767.

Campos:2019:PFS

[CLF^{+}19] A. Campos, P. Lopes, P. Fonseca, I. Figueiredo, V. Henriques, N. Gouveia, J. Delgado, L. Gouveia, A. Amorim, G. Araujo, T. Drago, and A. dos Santos. Portuguese fisheries in seamounts of Madeira–Tore (NE Atlantic). Marine Policy, 99(??):50–57, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301386.

Clifton:2013:RCT

[Cli13] Julian Clifton. Refocusing conservation through a cultural lens: Improving governance in the Wakatobi National Park, Indonesia. Marine Policy, 41(??):80–86, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002564.

Chen:2017:DSR

[CLLL17] Jihong Chen, Kevin X. Li, Xiang Liu, and Haobo Li. The development of ship registration policy in China: Response to flags of convenience. Marine Policy, 83(??): 22–28, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16308193.

Chang:2013:AFM

[CLLS13] Yi Chang, Ming-An Lee, Kuo-Tien Lee, and Kwang-Tsao Shao. Adaptation of fisheries and mariculture management to extreme oceanic environmental changes and climate variability in Taiwan. Marine Policy, 38(??):476–482, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001698.

Chang:2010:DWF

[CLS10] Shui-Kai Chang, Kun-Yu Liu, and Yann-Huei Song. Distant water fisheries development and vessel monitoring system implementation in Taiwan — History and driving forces. Marine Policy, 34(3):541–548, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001730.

Chantavanich:2016:USF

[CLS16] Supang Chantavanich, Samarn Laodumrongchai, and Christina Stringer. Under the shadow: Forced labour among sea fishers in Thailand. Marine Policy, 68(??):1–7, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003851.

Cabral:2016:ESA

[CLV^{+}16] P. Cabral, H. Levrel, F. Viard, K. Frangoudes, S. Girard, and P. Scemama. Ecosystem services assessment and compensation costs for installing seaweed farms. Marine Policy, 71(??):157–165, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303220.

Cooke:2012:MRM

[CLW^{+}12] Justin Cooke, Russell Leaper, Paul Wade, David Lavigne, and Barbara Taylor. Management rules for marine mammal populations: a response to Lonergan. Marine Policy, 36(2):389–392, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001114.

Chapsos:2017:MSI

[CM17] Ioannis Chapsos and James A. Malcolm. Maritime security in Indonesia: Towards a comprehensive agenda? Marine Policy, 76(??):178–184, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305668.

Cabral:2015:DMP

[CMA15] Reniel B. Cabral, Samuel S. Mamauag, and Porfirio M. Aliño. Designing a marine protected areas network in a data-limited situation. Marine Policy, 59(??):64– 76, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000950.

Castello:2013:AHS

[CMAA13] Leandro Castello, David G. McGrath, Caroline C. Arantes, and Oriana T. Almeida. Accounting for heterogeneity in small-scale fisheries management: the Amazon case. Marine Policy, 38(??):557–565, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001765.

Curtis:2015:MCM

[CMB^{+}15] K. Alexandra Curtis, Jeffrey E. Moore, Charlotte Boyd, Peter W. Dillingham, Rebecca L. Lewison, Barbara L. Taylor, and Kelsey C. James. Managing catch of marine megafauna: Guidelines for setting limit reference points. Marine Policy, 61(??):249–263, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002031.

Clark:2013:PBC

[CMC^{+}13] Douglas A. Clark, Chanda Meek, John Cheechoo, Susan Clark, A. Lee Foote, David Lee, and Geoff York. Polar bears and CITES: a rejoinder to Parsons and Cornick. Marine Policy, 38(??):365–368, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001534.

Cisneros-Montemayor:2013:EII

[CMCMHP13] Andrés M. Cisneros-Montemayor, Miguel A. Cisneros-Mata, Sarah Harper, and Daniel Pauly. Extent and implications of IUU catch in Mexico’s marine fisheries. Marine Policy, 39(??):283–288, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002448.

Castillo-Manzano:2010:FTP

[CMFDCLV10] José I. Castillo-Manzano, David Florido-Del-Corral, and Lourdes Lopez-Valpuesta. Fisheries traffic: the poor relation of port devolution. Lessons from Spanish state ports of general interest. Marine Policy, 34(1):189–192, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000517.

Caballero-Miguez:2015:IAA

[CMFG15] Gonzalo Caballero-Miguez and Raquel Fernández-González. Institutional analysis, allocation of liabilities and thirdparty enforcement via courts: the case of the Prestige oil spill. Marine Policy, 55(??):90–101, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000056.

Cardenosa:2019:PGG

[CMH19] Diego Cardeñosa, Wessley Merten, and John Hyde. Prioritizing global genetic capacity building assistance to implement CITES shark and ray listings. Marine Policy, 106(??):Article 103544, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305748.

Cisneros-Montemayor:2019:SEB

[CMMBV^{+}19] Andrés M. Cisneros-Montemayor, Marcia Moreno-Báez, Michelle Voyer, Edward H. Allison, William W. L. Cheung, Margot Hessing-Lewis, Muhammed A. Oyinlola, Gerald G. Singh, Wilf Swartz, and Yoshitaka Ota. Social equity and benefits as the nexus of a transformative Blue Economy: a sectoral review of implications. Marine Policy, 109(??):Article 103702, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305093.

Caldow:2015:BAF

[CMP^{+}15] Chris Caldow, Mark E. Monaco, Simon J. Pittman, Matthew S. Kendall, Theresa L. Goedeke, Charles Menza, Brian P. Kinlan, and Bryan M. Costa. Biogeographic assessments: a framework for information synthesis in marine spatial planning. Marine Policy, 51(??):423–432, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14001973.

Cisneros-Montemayor:2019:BMH

[CMS19] Andrés M. Cisneros-Montemayor and U. Rashid Sumaila. Busting myths that hinder an agreement to end harmful fisheries subsidies. Marine Policy, 109(??):Article 103699, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19306803.

Cisneros-Montemayor:2010:GPW

[CMSKP10] A. M. Cisneros-Montemayor, U. R. Sumaila, K. Kaschner, and D. Pauly. The global potential for whale watching. Marine Policy, 34(6):1273–1278, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000103X.

Cisneros-Montemayor:2016:SRF

[CMSM^{+}16] Andrés M. Cisneros-Montemayor, Enrique Sanjurjo, Gordon R. Munro, Victor Hernández-Trejo, and U. Rashid Sumaila. Strategies and rationale for fishery subsidy reform. Marine Policy, 69(??):229–236, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002857.

Caballero-Miguez:2014:ICF

[CMVLGG14] Gonzalo Caballero-Miguez, Manuel M. Varela-Lafuente, and María Dolores Garza-Gil. Institutional change, fishing rights and governance mechanisms: the dynamics of the Spanish 300 fleet on the Grand Sole fishing grounds. Marine Policy, 44(??):465–472, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002406.

Cinner:2010:DLS

[CMW10] J. E. Cinner, T. R. McClanahan, and A. Wamukota. Differences in livelihoods, socioeconomic characteristics, and knowledge about the sea between fishers and non-fishers living near and far from marine parks on the Kenyan coast. Marine Policy, 34(1):22–28, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000542.

Carter:2011:EEC

[CN11] Brian T. G. Carter and Erik A. Nielsen. Exploring ecological changes in Cook Inlet beluga whale habitat though traditional and local ecological knowledge of contributing factors for population decline. Marine Policy, 35(3):299–308, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001776.

Cepic:2017:JNC

[CN17]  Dražen Cepić and Fiona Nunan.  Justifying non-

compliance: the morality of illegalities in small scale fisheries of Lake Victoria, East Africa. Marine Policy, 86 (??):104–110, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17304104.

Crona:2010:MCS

[CNFJ10] Beatrice Crona, Magnus Nyström, Carl Folke, and Narriman Jiddawi. Middlemen, a critical social-ecological link in coastal communities of Kenya and Zanzibar. Marine Policy, 34(4):761–771, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000242.

Calado:2010:MSP

[CNJ^{+}10] H. Calado, K. Ng, D. Johnson, L. Sousa, M. Phillips, and F. Alves. Marine spatial planning: Lessons learned from the Portuguese debate. Marine Policy, 34(6):1341– 1349, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001132.

Chou:2016:ORD

[CO16] Yi Chou and Ching-Hsiewn Ou. The opportunity to regulate domestic fishery subsidies through international agreements. Marine Policy, 63(??):118–125, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500278X.

Cochrane:2015:UMC

[Coc15] Kevern Cochrane. Use and misuse of CITES as a management tool for commercially-exploited aquatic species. Marine Policy, 59(??):16–31, September 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000974.

Chung:2019:TRB

[COG^{+}19] Anne Chung, Thomas Oliver, Jamison Gove, Kelvin Gorospe, Darla White, Kristine Davidson, and William Walsh. Translating resilience-based management theory to practice for coral bleaching recovery in Hawai‘ i. Marine Policy, 99(??):58–68, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303841.

Cooper:2011:PIC

[Coo11] J. A. G. Cooper. Progress in Integrated Coastal Zone Management (ICZM) in Northern Ireland. Marine Policy, 35 (6):794–799, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000030.

Corten:2014:EMF

[Cor14] Ad Corten. EU–Mauritania fisheries partnership in need of more transparency. Marine Policy, 49(??):1–11, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14001122.

Cochrane:2014:ACS

[COS14] K. L. Cochrane, B. Oliver, and W. Sauer. An assessment of the current status of the chokka squid fishery in South Africa and an evaluation of alternative allocation strategies. Marine Policy, 43(??):149–163, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001206.

Coulthard:2011:MTJ

[Cou11] Sarah Coulthard. More than just access to fish: the pros and cons of fisher participation in a customary marine tenure (Padu) system under pressure. Marine Policy, 35(3):405–412, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002113.

Curtin:2010:UME

[CP10] Richard Curtin and Raúl Prellezo. Understanding marine ecosystem based management: a literature review. Marine Policy, 34(5):821–830, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000047.

Campbell:2013:MGA

[CP13] Brooke Campbell and Daniel Pauly. Mariculture: a global analysis of production trends since 1950. Marine Policy, 39(??):94–100, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002072.

Coglan:2015:CCM

[CP15] Louisa Coglan and Sean Pascoe. Corporate-cooperative management of fisheries: a potential alternative governance structure for low value small fisheries? Marine Policy, 57(??):27–35, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1500072X.

Curtin:2018:IIM

[CP18] Polly Curtin and Sarah Papworth. Increased information and marketing to specific individuals could shift conservation support to less popular species. Marine Policy, 88(??):101–107, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305572.

Courtney:2019:TSS

[CPB19] Catherine A. Courtney, Robert Pomeroy, and Stephen H. Brooks. Taking stock of the status of implementation of the Voluntary Guidelines for Securing Sustainable Smallscale Fisheries: a country-level assessment framework. Marine Policy, 100(??):361–370, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302240.

Comte:2019:CAG

[CPBQ19] Adrien Comte, Linwood H. Pendleton, Denis Bailly, and Emmanuelle Quillérou. Conceptual advances on global scale assessments of vulnerability: Informing investments for coastal populations at risk of climate change. Marine Policy, 99(??):391–399, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301623.

Chuenpagdee:2013:MPA

[CPFS^{+}13] Ratana Chuenpagdee, Jose J. Pascual-Fernández, Emese Szeliánszky, Juan Luis Alegret, Julia Fraga, and Svein Jentoft. Marine protected areas: Re-thinking their inception. Marine Policy, 39(??):234–240, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002151.

Carruthers:2019:OBF

[CPKF19] Erin H. Carruthers, Courtenay E. Parlee, Robert Keenan, and Paul Foley. Onshore benefits from fishing: Tracking value from the northern shrimp fishery to communities in Newfoundland and Labrador. Marine Policy, 103(??):130–137, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306122.

Carbonetti:2014:OLP

[CPR14] Benjamin Carbonetti, Robert Pomeroy, and David L. Richards. Overcoming the lack of political will in small scale fisheries. Marine Policy, 44(??):295–301, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002170.

Cho:2019:DMP

[CPRP19] Hyungmin Cho, Gibyung Park, Jewan Ryu, and Heekyung Park. Decision-making process for places of refuge in hazardous and noxious substances incident: Case study of South Korea. Marine Policy, 108(??):Article 103643, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307474.

Chen:2014:LBJ

[CQ14] Chung-Ling Chen and Guo-Hao Qiu. The long and bumpy journey: Taiwan’s aquaculture development and management. Marine Policy, 48(??):152–161, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000980.

Carollo:2010:EBM

[CR10] Cristina Carollo and Dave J. Reed. Ecosystem-based management institutional design: Balance between federal, state, and local governments within the Gulf of Mexico Alliance. Marine Policy, 34(1):178–181, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000931.

Crona:2011:OLA

[CR11] B. Crona and S. Rosendo. Outside the law? Analyzing policy gaps in addressing fishers’ migration in East Africa. Marine Policy, 35(3):379–388, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002010.

Craig:2019:FAM

[Cra19] Robin Kundis Craig. Fostering adaptive marine aquaculture through procedural innovation in marine spatial planning. Marine Policy, 110(??):Article 103555, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303197.

Chuenpagdee:2019:GIS

[CRB^{+}19] Ratana Chuenpagdee, Delphine Rocklin, David Bishop, Matthew Hynes, Randal Greene, Miguel R. Lorenzi, and Rodolphe Devillers. The global information system on small-scale fisheries (ISSF): a crowdsourced knowledge platform. Marine Policy, 101(??):158–166, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300155.

Celliers:2013:PIC

[CRCD13] L. Celliers, S. Rosendo, I. Coetzee, and G. Daniels. Pathways of integrated coastal management from national policy to local implementation: Enabling climate change adaptation. Marine Policy, 39(??):72–86, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002035.

Cronin:2011:CSI

[Cro11] M. A. Cronin. The conservation of seals in Irish waters: How research informs policy. Marine Policy, 35 (6):748–755, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000078.

Coers:2012:SPM

[CRO12] Aukje Coers, Jesper Raakjær, and Christian Olesen. Stakeholder participation in the management of North East Atlantic pelagic fish stocks: the future role of the Pelagic Regional Advisory Council in a reformed CFP. Marine Policy, 36(3):689–695, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001679.

Cross:2015:WFU

[Cro15] Helen Cross. Why fish? Using entry-strategies to inform governance of the small-scale sector: a case-study in the Bijagós Archipelago (West Africa). Marine Policy, 51 (??):128–135, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1400181X.

Catchpole:2017:CLO

[CRSM^{+}17] Thomas L. Catchpole, Ana Ribeiro-Santos, Stephen C. Mangi, Christopher Hedley, and Tim S. Gray. The challenges of the landing obligation in EU fisheries. Marine Policy, 82(??):76–86, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305875.

Criddle:2013:DDC

[CS13] Keith R. Criddle and James Strong. Dysfunction by design: Consequences of limitations on transferability of catch shares in the Alaska pollock fishery. Marine Policy, 40(??):91–99, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000158.

Canu:2014:SEA

[CS14] Donata Melaku Canu and Cosimo Solidoro. Socio-economic analysis and stakeholder involvement: Mussel-farming in the Gulf of Trieste. Marine Policy, 43(??):55–62, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001024.

Chen:2017:SAJ

[CS17] Gang Chen and Desai Shan. Seafarers’ access to jurisdictions over labour matters. Marine Policy, 77(??):1–8, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305425.

Casey:2019:PNP

[CS19] James F. Casey and Peter W. Schuhmann. PACT or no PACT are tourists willing to contribute to the Protected Areas Conservation Trust in order to enhance marine resource conservation in Belize? Marine Policy, 101(??):8–14, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300617.

Christie:2014:CLA

[CSBE14] N. Christie, K. Smyth, R. Barnes, and M. Elliott. Colocation of activities and designations: a means of solving or creating problems in marine spatial planning? Marine Policy, 43(??):254–261, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001280.

Cinti:2010:UCF

[CSCBR10] A. Cinti, W. Shaw, R. Cudney-Bueno, and M. Rojo. The unintended consequences of formal fisheries policies: Social disparities and resource overuse in a major fishing community in the Gulf of California, Mexico. Marine Policy, 34(2):328–339, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001109.

Carr:2013:ISF

[CSF^{+}13] Lindsey A. Carr, Adrian C. Stier, Katharina Fietz, Ignacio Montero, Austin J. Gallagher, and John F. Bruno. Illegal shark fishing in the Galápagos Marine Reserve. Marine Policy, 39(??):317–321, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002461.

Chiba:2018:HFA

[CSF^{+}18] Sanae Chiba, Hideaki Saito, Ruth Fletcher, Takayuki Yogi, Makino Kayo, Shin Miyagi, Moritaka Ogido, and Katsunori Fujikura. Human footprint in the abyss: 30 year records of deep-sea plastic debris. Marine Policy, 96(??):204–212, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17305195.

Chagaris:2019:MCO

[CSF^{+}19] David Chagaris, Skyler Sagarese, Nick Farmer, Behzad Mahmoudi, Kim de Mutsert, Steven VanderKooy, William F. Patterson, Morgan Kilgour, Amy Schueller, Robert Ahrens, and Matthew Lauretta. Management challenges are opportunities for fisheries ecosystem models in the Gulf of Mexico. Marine Policy, 101(??):1–7, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303026.

Cardinale:2017:ERS

[CSH17] Massimiliano Cardinale, Anders Svenson, and Joakim Hjelm. The “easy restriction” syndrome drive local fish stocks to extinction: the case of the management of Swedish coastal populations. Marine Policy, 83(??): 179–183, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300660.

Cannon:2018:FIP

[CSK^{+}18] James Cannon, Pedro Sousa, Isidora Katara, Pedro Veiga, Braddock Spear, Douglas Beveridge, and Tracy Van Holt. Fishery improvement projects: Performance over the past decade. Marine Policy, 97(??):179–187, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300393.

Cullis-Suzuki:2012:RSD

[CSMB^{+}12] Sarika Cullis-Suzuki, Murdoch McAllister, Pamela Baker, Tom Carruthers, and T. J. Tate. Red snapper discards in the Gulf of Mexico: fishermen’s perceptions following the implementation of Individual Fishing Quotas. Marine Policy, 36(3):583–591, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001539.

Cuervo-Sanchez:2018:SMP

[CSMR18] Rafael Cuervo-Sánchez, Jorge H. Maldonado, and Mario Rueda. Spillover from marine protected areas on the Pacific coast in Colombia: a bioeconomic modelling approach for shrimp fisheries. Marine Policy, 88(??):182–188, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307509.

Caballero:2017:ECJ

[CSO17] Gonzalo Caballero and David Soto-Oñate. Environmental crime and judicial rectification of the Prestige oil spill: the polluter pays. Marine Policy, 84(??):213–219, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300878.

Cullis-Suzuki:2010:FHS

[CSP10] Sarika Cullis-Suzuki and Daniel Pauly. Failing the high seas: a global evaluation of regional fisheries management organizations. Marine Policy, 34(5):1036–1042, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000540.

Cleasby:2014:SEC

[CSP^{+}14] Nathan Cleasby, Anne-Maree Schwarz, Michael Phillips, Chris Paul, Jharendu Pant, Janet Oeta, Tim Pickering, Alex Meloty, Michael Laumani, and Max Kori. The socioeconomic context for improving food security through land based aquaculture in Solomon Islands: a peri-urban case study. Marine Policy, 45(??):89–97, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002741.

Colbert-Sangree:2015:CBF

[CSS15] Nathanial Colbert-Sangree and Jordan F. Suter. Community based fishery management within the Menai Bay conservation area: a survey of the resource user. Marine Policy, 60(??):171–177, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001761.

Cinti:2010:IUI

[CST10] A. Cinti, W. Shaw, and J. Torre. Insights from the users to improve fisheries performance: Fishers’ knowledge and attitudes on fisheries policies in Bahía de Kino, Gulf of California, Mexico. Marine Policy, 34(6):1322–1334, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001119.

Carpenter:2018:PPM

[CSvH^{+}18] Angela Carpenter, Rebecca Shellock, Robin von Haartman, Steve Fletcher, and Gillian Glegg. Public perceptions of management priorities for the English Channel region. Marine Policy, 97(??):294–304, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305013.

Campbell:2014:MFM

[CSVHS14] Maria S. Campbell, Kilian M. Stehfest, Stephen C. Votier, and Jason M. Hall-Spencer. Mapping fisheries for marine spatial planning: Gear-specific vessel monitoring system (VMS), marine conservation and offshore renewable energy. Marine Policy, 45(??):293–300, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002121.

Chen:2016:MPC

[CT16] Chung-Ling Chen and Ning Teng. Management priorities and carrying capacity at a high-use beach from tourists’ perspectives: a way towards sustainable beach tourism.

Marine Policy, 74(??):213–219, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303712.

Cusumano:2017:ESS

[Cus17] Eugenio Cusumano. Emptying the sea with a spoon? Non-governmental providers of migrants search and rescue in the Mediterranean. Marine Policy, 75(??):91–98, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304821.

Cabrera:2012:STM

[CVFP^{+}12] R. Cabrera, C. Valero-Franco, F. Piniella, C. Díaz, M. C. Gómez-Cama, M. C. Soriguer, and J. A. Hernando. A software tool for monitoring legal minimum length of landings: Case study of a fishery in southern Spain. Marine Policy, 36(4):895–902, July 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001989.

Carvalho:2017:SEO

[CVP^{+}17] André N. Carvalho, Paulo Vasconcelos, David Piló, Fábio Pereira, and Miguel B. Gaspar. Socio-economic, operational and technical characterisation of the harvesting of gooseneck barnacle (Pollicipes pollicipes) in SW Portugal: Insights towards fishery co-management. Marine Policy, 78(??):34–44, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630776X.

Cvitanovic:2018:BTA

[CvPH^{+}18] C. Cvitanovic, E. I. van Putten, A. J. Hobday, M. Mackay, R. Kelly, J. McDonald, K. Waples, and P. Barnes. Building trust among marine protected area managers and community members through scientific research: Insights from the Ningaloo Marine Park, Australia. Marine Policy, 93(??):195–206, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301325.

Cariou:2015:ISV

[CW15] Pierre Cariou and Francois-Charles Wolff. Identifying substandard vessels through Port State Control inspections: a new methodology for Concentrated Inspection Campaigns. Marine Policy, 60(??):27–39, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001463.

Cox:2016:CLG

[CWP16] Michael Cox, Margaret Wilson, and Tyler Pavlowich. The challenges of local governance: Gear-based fragmentation in the Dominican fishery of Buen Hombre. Marine Policy, 63(??):109–117, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002808.

Chang:2013:WET

[CYI13] Shui-Kai Chang, Ya-Ching Yang, and Nozomu Iwasaki. Whether to employ trade controls or fisheries management to conserve precious corals (Coralliidae) in the Northern Pacific Ocean. Marine Policy, 39(??):144–153, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001960.

Chen:2019:IKF

[CZX^{+}19] Jihong Chen, Shuanghong Zhang, Lang Xu, Zheng Wan, Yijie Fei, and Tianxiao Zheng. Identification of key factors of ship detention under Port State Control. Marine Policy, 102(??):21–27, April 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307279.

Dale:2011:MMC

[DA11] Aaron Dale and Derek Armitage. Marine mammal comanagement in Canada’s Arctic: Knowledge co-production for learning and adaptive capacity. Marine Policy, 35 (4):440–449, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1000206X.

Danielsen:2018:FPF

[DA18] Rannvá Danielsen and Sveinn Agnarsson. Fisheries policy in the Faroe Islands: Managing for failure? Marine Policy, 94(??):204–214, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302756.

Dunn:2014:CBD

[DAB^{+}14] Daniel C. Dunn, Jeff Ardron, Nicholas Bax, Patricio Bernal, Jesse Cleary, Ian Cresswell, Ben Donnelly, Piers Dunstan, Kristina Gjerde, David Johnson, Kristin Kaschner, Ben Lascelles, Jake Rice, Henning von Nordheim, Louisa Wood, and Patrick N. Halpin. The Convention on Biological Diversity’s Ecologically or Biologically Significant Areas: Origins, development, and current status. Marine Policy, 49(??):137–145, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002856.

deAndrade:2017:OMP

[dAdOS17] Aline Batista de Andrade and Marcelo de Oliveira Soares. Offshore marine protected areas: Divergent perceptions of divers and artisanal fishers. Marine Policy, 76(??): 107–113, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301786.

DAmico:2016:NPL

[DAGG16] P. D’Amico, A. Armani, D. Gianfaldoni, and A. Guidi. New provisions for the labelling of fishery and aquaculture products: Difficulties in the implementation of Regulation (EU) n. 1379/2013. Marine Policy, 71(??):147–156, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630135X.

Deepananda:2015:IKB

[DAJM15] K. H. M. Ashoka Deepananda, Upali S. Amarasinghe, and Udith K. Jayasinghe-Mudalige. Indigenous knowledge in the beach seine fisheries in Sri Lanka: an indispensable factor in community-based fisheries management. Marine Policy, 57(??):69–77, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000779.

Das:2016:FAF

[Das16] Chhandita Das. Fisheries annual fixed cost data collection and estimation methodology: an application in the Northeast, U.S. Marine Policy, 71(??):184–193, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303190.

Davidson:2010:CBL

[Dav10] Arthur Davidson. The cost-benefit ledger of quota leasing. Marine Policy, 34(5):1115–1116, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000014X.

Davis:2012:POR

[Dav12] Mary E. Davis. Perceptions of occupational risk by US commercial fishermen. Marine Policy, 36(1):28–33, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000558.

Davis:2015:ASC

[Dav15] Reade Davis. ‘All in’: Snow crab, capitalization, and the future of small-scale fisheries in Newfoundland. Marine Policy, 61(??):323–330, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000901.

Dancette:2019:AAP

[DB19a] Raphaëlle Dancette and Jean-Claude Brêthes. An analysis of actors’ perceptions of Maio Island’s (Cape Verde) marine governance. Marine Policy, 104(??):177–197, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305347.

Davies:2019:MHC

[DB19b] Kimberley T. A. Davies and Sean W. Brillant. Mass human-caused mortality spurs federal action to protect endangered North Atlantic right whales in Canada. Marine Policy, 104(??):157–162, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306675.

Delord:2014:AIS

[DBB^{+}14] Karine Delord, Christophe Barbraud, Charles-André Bost, Bernard Deceuninck, Thierry Lefebvre, Rose Lutz, Thierry Micol, Richard A. Phillips, Phil N. Trathan, and Henri Weimerskirch. Areas of importance for seabirds tracked from French southern territories, and recommendations for conservation. Marine Policy, 48(??):1–13, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000694.

Duarte:2019:CAT

[DBD19] Dérien L. V. Duarte, Matt K. Broadhurst, and Luiz F. C. Dumont. Challenges in adopting turtle excluder devices (TEDs) in Brazilian penaeid-trawl fisheries. Marine Policy, 99(??):374–381, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305025.

Dolman:2016:TEA

[DBE^{+}16] Sarah Dolman, Sarah Baulch, Peter G. H. Evans, Fiona Read, and Fabian Ritter. Towards an EU Action Plan on Cetacean Bycatch. Marine Policy, 72(??):67–75, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300306.

Davis:2016:AIA

[DBG^{+}16] Andrew R. Davis, Allison Broad, Warwick Gullett, James Reveley, Chantel Steele, and Clive Schofield. Anchors away? The impacts of anchor scour by ocean-going vessels and potential response options. Marine Policy, 73 (??):1–7, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302615.

deBruyn:2013:PAF

[dBMA13] Paul de Bruyn, Hilario Murua, and Martín Aranda. The Precautionary approach to fisheries management: How this is taken into account by Tuna regional fisheries management organisations (RFMOs). Marine Policy, 38 (??):397–406, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001583.

Djoumessi:2019:FII

[DCC19] Armand Djoumessi, Shu-Ling Chen, and Stephen Cahoon. Factors influencing innovation in maritime clusters: an empirical study from Australia. Marine Policy, 108(??):Article 103558, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309539.

Davies:2019:GMA

[DCF^{+}19] Ian P. Davies, Valerie Carranza, Halley E. Froehlich, Rebecca R. Gentry, Peter Kareiva, and Benjamin S. Halpern. Governance of marine aquaculture: Pitfalls, potential, and pathways forward. Marine Policy, 104(??):29–36, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302537.

Ding:2017:VIC

[DCHC17] Qi Ding, Xinjun Chen, Ray Hilborn, and Yong Chen. Vulnerability to impacts of climate change on marine fisheries and food security. Marine Policy, 83(??):55–61, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300040.

DaRocha:2012:CFP

[DCV12] José-María Da Rocha, Santiago Cerviño, and Sebastian Villasante. The Common Fisheries Policy: an enforcement problem. Marine Policy, 36(6):1309–1314, November 2012.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000425.

deConing:2015:TNC

[dCW15] Eve de Coning and Emma Witbooi. Towards a new’fisheries crime’ paradigm: South Africa as an illustrative example. Marine Policy, 60(??):208–215, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500189X.

Dubik:2019:GFF

[DCY^{+}19] Bradford A. Dubik, Elizabeth C. Clark, Talia Young, Sarah Bess Jones Zigler, Mikaela M. Provost, Malin L. Pinsky, and Kevin St. Martin. Governing fisheries in the face of change: Social responses to long-term geographic shifts in a U.S. fishery. Marine Policy, 99(??):243–251, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305463.

Day:2013:EGL

[DD13] Jon C. Day and Kirstin Dobbs. Effective governance of a large and complex cross-jurisdictional marine protected area: Australia’s Great Barrier Reef. Marine Policy, 41(??):14–24, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002618.

Dudley:2017:DMP

[DDL^{+}17] Nigel Dudley, Jon Day, Dan Laffoley, Marc Hockings, and Sue Stolton. Defining marine protected areas: a response to Horta e Costa et al. Marine Policy, 77(??):191–192, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307011.

DaRos:2019:DSN

[DDM^{+}19] Zaira Da Ros, Antonio Dell’Anno, Telmo Morato, Andrew K. Sweetman, Marina Carreiro-Silva, Chris J. Smith, Nadia Papadopoulou, Cinzia Corinaldesi, Silvia Bianchelli, Cristina Gambi, Roberto Cimino, Paul Snelgrove, Cindy Lee Van Dover, and Roberto Danovaro. The deep sea: the new frontier for ecological restoration. Marine Policy, 108(??):Article 103642, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309199.

Dichmont:2016:HMA

[DDP16] Catherine M. Dichmont, Roy A. Deng, and André E. Punt. How many of Australia’s stock assessments can be conducted using stock assessment packages? Marine Policy, 74(??):279–287, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305048.

Dobbs:2011:DLT

[DDS^{+}11] Kirstin Dobbs, Jon Day, Hilary Skeat, John Baldwin, Fergus Molloy, Laurence McCook, Margaret Johnson, Bruce Elliot, Andrew Skeat, Karen Vohland, David Wachenfeld, and Richard Kenchington. Developing a long-term outlook for the Great Barrier Reef, Australia: a framework for adaptive management reporting underpinning an ecosystem-based management approach. Marine Policy, 35(2):233–240, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001752.

Dowling:2013:LHV

[DDV^{+}13] N. A. Dowling, C. M. Dichmont, W. Venables, A. D. M. Smith, D. C. Smith, D. Power, and D. Galeano. From low- to high-value fisheries: Is it possible to quantify the trade-off between management cost, risk and catch? Marine Policy, 40(??):41–52, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002503.

Dutra:2016:HIC

[DDvP^{+}16] Leo X. C. Dutra, Catherine M. Dichmont, Ingrid E. van Putten, Olivier Thébaud, Roy A. Deng, Ricardo Pascual, Randall Owens, Eddie Jebreen, Carolyn Thompson, M. St. J. Warne, R. Quinn, J. Bennett, M. Read, D. Wachenfeld, Catherine Collier, Michelle Waycott, J. Davies, A. Garland, M. Dunning, and J. Playford. How important is the coast? A survey of coastal objectives in an Australian regional city. Marine Policy, 71(??):229–241, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303104.

DeSanto:2010:WSP

[De 10] Elizabeth M. De Santo. ‘Whose science?’ Precaution and power-play in European marine environmental decisionmaking. Marine Policy, 34(3):414–420, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001390.

DeSanto:2011:EJI

[De 11] Elizabeth M. De Santo. Environmental justice implications of Maritime Spatial Planning in the European Union. Marine Policy, 35(1):34–38, January 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001399.

DeSanto:2013:DMS

[De 13] Elizabeth M. De Santo. The Darwin Mounds special area of conservation: Implications for offshore marine governance. Marine Policy, 41(??):25–32, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300016X.

DeSanto:2016:APP

[De 16] Elizabeth M. De Santo. Assessing public “participation” in environmental decision-making: Lessons learned from the UK Marine Conservation Zone (MCZ) site selection process. Marine Policy, 64(??):91–101, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500322X.

DeLucia:2017:AEB

[De 17] Vito De Lucia. The Arctic environment and the BBNJ negotiations. Special rules for special circumstances? Marine Policy, 86(??):234–240, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304086.

DeSanto:2018:ICA

[De 18] Elizabeth M. De Santo. Implementation challenges of area-based management tools (ABMTs) for biodiversity beyond national jurisdiction (BBNJ). Marine Policy, 97 (??):34–43, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303166.

Deb:2018:SNW

[Deb18] Apurba Krishna Deb. ‘Surrender to nature’: Worldviews and rituals of the small-scale coastal fishers of Bangladesh. Marine Policy, 92(??):1–12, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306826.

Deligiannis:2017:SPI

[Del17] Panos Deligiannis. Ship performance indicator. Marine Policy, 75(??):204–209, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X1630029X.

Diekert:2010:WCC

[DES10] Florian K. Diekert, Anne Maria Eikeset, and Nils Chr. Stenseth. Where could catch shares prevent stock collapse? Marine Policy, 34(3):710–712, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001419.

Duggan:2013:IFS

[DFK13] Deirdre E. Duggan, Keith D. Farnsworth, and Sarah B. M. Kraak. Identifying functional stakeholder clusters to maximise communication for the ecosystem approach to fisheries management. Marine Policy, 42(??):56–67, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000407.

Duy:2015:GSP

[DFL15] Nguyen Ngoc Duy, Ola Flaaten, and Le Kim Long. Government support and profitability effects — Vietnamese offshore fisheries. Marine Policy, 61(??):77–86, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002146.

Dalton:2015:CMM

[DFP15] Tracey Dalton, Graham Forrester, and Richard Pollnac. Are Caribbean MPAs making progress toward their goals and objectives? Marine Policy, 54(??):69–76, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003467.

DAnna:2016:GAE

[DFP^{+}16] Giovanni D’Anna, Tomás Vega Fernández, Carlo Pipitone, Germana Garofalo, and Fabio Badalamenti. Governance analysis in the Egadi Islands Marine Protected Area: a Mediterranean case study. Marine Policy, 71(??): 301–309, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003796.

DellApa:2015:SMC

[DFSB15] Andrea Dell’Apa, Adam Fullerton, Franklin Schwing, and Margaret M. Brady. The status of marine and coastal ecosystem-based management among the network of U.S. federal programs. Marine Policy, 60(??):249–258, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002122.

Duran:2015:CMC

[DFV15] Roi Durán, Begoña A. Farizo, and María Xosé Vázquez. Conservation of maritime cultural heritage: a discrete choice experiment in a European Atlantic Region. Marine Policy, 51(??):356–365, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002565.

Druel:2014:SML

[DG14] Elisabeth Druel and Kristina M. Gjerde. Sustaining marine life beyond boundaries: Options for an implementing agreement for marine biodiversity beyond national jurisdiction under the United Nations Convention on the Law of the Sea. Marine Policy, 49(??):90–97, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002820.

deGroot:2010:PDT

[dGB10] Jiska de Groot and Simon R. Bush. The potential for dive tourism led entrepreneurial marine protected areas in Curacao. Marine Policy, 34(5):1051–1059, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000564.

Ding:2014:BCC

[DGC14] Juan Ding, Xueqian Ge, and Ryan Casey. “Blue competition” in China: Current situation and challenges. Marine Policy, 44(??):351–359, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300225X.

Darling:2017:RDA

[DGC^{+}17] John A. Darling, Bella S. Galil, Gary R. Carvalho, Marc Rius, Frédérique Viard, and Stefano Piraino. Recommendations for developing and applying genetic tools to assess and manage biological invasions in marine ecosystems. Marine Policy, 85(??):54–64, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303421.

deGrunt:2018:TSI

[dGNC18] L. S. de Grunt, K. Ng, and H. Calado. Towards sustainable implementation of maritime spatial planning in Europe: a peek into the potential of the Regional Sea Conventions playing a stronger role. Marine Policy, 95(??):102–110, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300460.

Dey:2016:EICa

[DGVS^{+}16a] Madan Mohan Dey, Kamal Gosh, Rowena Valmonte-Santos, Mark W. Rosegrant, and Oai Li Chen. Economic impact of climate change and climate change adaptation strategies for fisheries sector in Fiji. Marine Policy, 67 (??):164–170, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003930.

Dey:2016:EICb

[DGVS^{+}16b] Madan Mohan Dey, Kamal Gosh, Rowena Valmonte-Santos, Mark W. Rosegrant, and Oai Li Chen. Economic impact of climate change and climate change adaptation strategies for fisheries sector in Solomon Islands: Implication for food security. Marine Policy, 67(??):171–178, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000099.

Dahlstrom:2011:RIR

[DHC11] Alisha Dahlstrom, Chad L. Hewitt, and Marnie L. Campbell. A review of international, regional and national biosecurity risk assessment frameworks. Marine Policy, 35(2):208–217, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001697.

Daigle:2016:CCP

[DHFL^{+}16] Rémi M. Daigle, Wolfgang Haider, Sergio Fernández-Lozada, Kimberly Irwin, Philippe Archambault, and Isabelle M. Côté. From coast to coast: Public perception of ocean-derived benefits in Canada. Marine Policy, 74 (??):77–84, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303323.

Garcia-del-Hoyo:2017:DAS

[dHJTG17] Juan José García del Hoyo, Ramón Jiménez-Toribio, and Patrice Guillotreau. A demand analysis of the Spanish canned tuna market. Marine Policy, 86(??):127–133, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302129.

DeBrauwer:2017:ECM

[DHM^{+}17] Maarten De Brauwer, Euan S. Harvey, Jennifer L. McIlwain, Jean-Paul A. Hobbs, Jamaluddin Jompa, and Michael Burton. The economic contribution of the muck dive industry to tourism in Southeast Asia. Marine Policy, 83(??):92–99, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300581.

Dubois:2016:RSF

[DHR16] Mark Dubois, Maria Hadjimichael, and Jesper Raakjær. The rise of the scientific fisherman: Mobilising knowledge and negotiating user rights in the Devon inshore brown crab fishery, UK. Marine Policy, 65(??):48–55, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003838.

Daoust:2014:RAW

[DHSC14] Pierre-Yves Daoust, Mike Hammill, Garry Stenson, and Charles Caraguel. A review of animal welfare implications of the Canadian commercial seal hunt: a critique. Marine Policy, 43(??):367–371, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001541.

Dixon:2016:MEO

[Dix16] Zachary Parke Dixon. Material expertise: an ontological approach to stakeholder participation in marine policy. Marine Policy, 72(??):107–114, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304146.

Deighan:2015:FRU

[DJ15] L. K. Deighan and L. D. Jenkins. Fishing for recognition: Understanding the use of NGO guidelines in fishery improvement projects. Marine Policy, 51(??):476–485, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002711.

DeSanto:2011:FCS

[DJM11] E. M. De Santo, P. J. S. Jones, and A. M. M. Miller. Fortress conservation at sea: a commentary on the Chagos marine protected area. Marine Policy, 35(2):258–260, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1000165X.

Diz:2018:MMB

[DJR^{+}18] Daniela Diz, David Johnson, Michael Riddell, Sian Rees, Jessica Battle, Kristina Gjerde, Sebastian Hennige, and J. Murray Roberts. Mainstreaming marine biodiversity into the SDGs: the role of other effective area-based conservation measures (SDG 14.5). Marine Policy, 93(??):251–261, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302178.

Dawson:2017:UCR

[DJS17] Jackie Dawson, Margaret Johnston, and Emma Stewart. The unintended consequences of regulatory complexity: the case of cruise tourism in Arctic Canada. Marine Policy, 76(??):71–78, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304080.

Dalton:2017:UNE

[DJTK17] Tracey Dalton, Di Jin, Robert Thompson, and Allison Katzanek. Using normative evaluations to plan for and manage shellfish aquaculture development in Rhode Island coastal waters. Marine Policy, 83(??):194–203, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303657.

DeMers:2012:OFF

[DK12] Amanda DeMers and Viktoria Kahui. An overview of Fiji’s fisheries development. Marine Policy, 36(1):174–179, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1100090X.

Dentoni:2015:CMP

[DK15] Domenico Dentoni and Laurens Klerkx. Co-managing public research in Australian fisheries through convergence– divergence processes. Marine Policy, 60(??):259–271, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1500202X.

Duggan:2016:SSB

[DK16] Deirdre E. Duggan and Momo Kochen. Small in scale but big in potential: Opportunities and challenges for fisheries certification of Indonesian small-scale tuna fisheries. Marine Policy, 67(??):30–39, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000130.

Dawson:2019:MEB

[DL19a] Claire Dawson and Phillip S. Levin. Moving the ecosystembased fisheries management mountain begins by shifting small stones: a critical analysis of EBFM on the U.S. West Coast. Marine Policy, 100(??):58–65, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305517.

Donlan:2019:ECS

[DL19b] C. Josh Donlan and Gloria M. Luque. Exploring the causes of seafood fraud: a meta-analysis on mislabeling and price. Marine Policy, 100(??):258–264, February 2019.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304731.

Pineiro-Antelo:2019:OGS

[dlÁPAVS19] María de los Ángeles Piñeiro-Antelo, Francisco Ramón Durán Villa, and Xosé M. Santos. ODA in Galicia (Spain). The importance of the fisheries sector and the cultural priority. Marine Policy, 107(??):Article 103455, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304111.

Garcia-de-la-Fuente:2016:MAI

[dlFFVRC16] Laura García de-la Fuente, Esteban Fernández-Vázquez, and Carmen Ramos-Carvajal. A methodology for analyzing the impact of the artisanal fishing fleets on regional economies: an application for the case of Asturias (Spain). Marine Policy, 74(??):165–176, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305735.

delaFuente:2017:SCE

[dlFRG17] Santiago Suárez de la Fuente, David Roberge, and Alistair R. Greig. Safety and CO_{2} emissions: Implications of using organic fluids in a ship’s waste heat recovery system. Marine Policy, 75(??):191–203, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000580.

Durden:2018:EIA

[DLM^{+}18] Jennifer M. Durden, Laura E. Lallier, Kevin Murphy, Aline Jaeckel, Kristina Gjerde, and Daniel O. B. Jones. Environmental Impact Assessment process for deep-sea mining in ‘the Area’. Marine Policy, 87(??):194–202, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305316.

deReynier:2010:BSS

[dLS10] Yvonne L. deReynier, Phillip S. Levin, and Noriko L. Shoji. Bringing stakeholders, scientists, and managers together through an integrated ecosystem assessment process. Marine Policy, 34(3):534–540, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001729.

delaTorre-Castro:2017:GAB

[dlTCFB^{+}17] Maricela de la Torre-Castro, Sara Fröcklin, Sanna Börjesson, Janine Okupnik, and Narriman S. Jiddawi. Gender analysis for better coastal management — Increasing our understanding of social-ecological seascapes. Marine Policy, 83(??):62–74, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308259.

delaTorre-Castro:2010:FIA

[dlTCL10] Maricela de la Torre-Castro and Lars Lindström. Fishing institutions: Addressing regulative, normative and cultural-cognitive elements to enhance fisheries management. Marine Policy, 34(1):77–84, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000670.

Dimmock:2015:SDT

[DM15] Kay Dimmock and Ghazali Musa. Scuba Diving Tourism System: a framework for collaborative management and sustainability. Marine Policy, 54(??):52–58, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003455.

deMitcheson:2018:CMM

[dMAH^{+}18] Yvonne Sadovy de Mitcheson, Astrid Alexandra Andersson, Alex Hofford, Calton S. W. Law, Loby C. Y. Hau, and Daniel Pauly. Out of control means off the menu: the case for ceasing consumption of luxury products from highly vulnerable species when international trade cannot be adequately controlled; shark fin as a case study. Marine Policy, 98(??):115–120, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303038.

Dunstan:2018:HCC

[DMB^{+}18] Piers K. Dunstan, Bradley R. Moore, Johann D. Bell, Neil J. Holbrook, Eric C. J. Oliver, James Risbey, Scott D. Foster, Quentin Hanich, Alistair J. Hobday, and Nathan J. Bennett. How can climate predictions improve sustainability of coastal fisheries in Pacific Small-Island Developing States? Marine Policy, 88(??):295–302, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301963.

Davidson:2017:PPB

[DMC^{+}17] Ian C. Davidson, Mark S. Minton, Katharine J. Carney, A. Whitman Miller, and Gregory M. Ruiz. Pioneering patterns of ballast treatment in the emerging era of marine vector management. Marine Policy, 78(??):158–162, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307370.

Demestre:2015:NFS

[DMdJ^{+}15] Montserrat Demestre, Alba Muntadas, Silvia de Juan, Chryssi Mitilineou, Paolo Sartor, Julio Mas, Stefanos Kavadas, and Javier Martín. The need for finescale assessment of trawl fishing effort to inform on an ecosystem approach to fisheries: Exploring three data sources in Mediterranean trawling grounds. Marine Policy, 62(??):134–143, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002638.

Durden:2017:PFR

[DMJ^{+}17] Jennifer M. Durden, Kevin Murphy, Aline Jaeckel, Cindy L. Van Dover, Sabine Christiansen, Kristina Gjerde, Aleyda Ortega, and Daniel O. B. Jones. A procedural framework for robust environmental management of deep-sea mining projects using a conceptual model. Marine Policy, 84(??):193–201, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300465.

Davies:2018:EMP

[DMKL18] Kathryn Davies, Ani A. Murchie, Vince Kerr, and Carolyn Lundquist. The evolution of marine protected area planning in Aotearoa New Zealand: Reflections on participation and process. Marine Policy, 93(??):113–127, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305067.

Davies:2014:PPF

[DMMG14] Tim K. Davies, Chris C. Mees, and E. J. Milner-Gulland. The past, present and future use of drifting fish aggregating devices (FADs) in the Indian Ocean. Marine Policy, 45(??):163–170, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002972.

Davies:2015:SGU

[DMMG15] Tim K. Davies, Chris C. Mees, and E. J. Milner-Gulland. Second-guessing uncertainty: Scenario planning for management of the Indian Ocean tuna purse seine fishery. Marine Policy, 62(??):169–177, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002705.

Diz:2019:MPS

[DMW19] Daniela Diz, Elisa Morgera, and Meriwether Wilson. Marine policy special issue: SDG synergies for sustainable fisheries and poverty alleviation. Marine Policy, 110(??):Article 102860, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305493.

Dawe:2012:SRC

[DN12] Jennifer L. Dawe and Barbara Neis. Species at risk in Canada: Lessons learned from the listing of three species of wolffish. Marine Policy, 36(2):405–413, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001126.

deOliveira:2013:FAM

[dO13] Lucia Perez de Oliveira. Fishers as advocates of marine protected areas: a case study from Galicia (NW Spain). Marine Policy, 41(??):95–102, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002655.

Doloreux:2017:WMC

[Dol17] David Doloreux. What is a maritime cluster? Marine Policy, 83(??):215–220, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303470.

Donkersloot:2016:CCA

[Don16] Rachel Donkersloot. Considering community allocations: Power and the politics of enclosure in the Gulf of Alaska. Marine Policy, 74(??):300–308, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302007.

Donda:2017:WBF

[Don17] Steve Donda. Who benefits from fisheries co-management? A case study in Lake Chiuta, Malawi. Marine Policy, 80(??):147–153, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306273.

Soares:2018:TLR

[dOSL18] Marcelo de Oliveira Soares and Caroline Costa Lucas. Towards large and remote protected areas in the South Atlantic Ocean: St. Peter and St. Paul’s Archipelago and the Vitória–Trindade Seamount Chain. Marine Policy, 93(??):101–103, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301350.

Doyon:2015:LGM

[Doy15] Sabrina Doyon. Losing ground: the marginalization of the St. Lawrence eel Fisheries in Québec, Canada. Marine Policy, 61(??):331–338, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003121.

Digal:2017:FAA

[DP17] Larry N. Digal and Shemaiah Gail P. Placencia. Factors affecting the adoption of sustainable tuna fishing practices: the case of municipal fishers in Maasim, Sarangani Province, Region 12, Philippines. Marine Policy, 77 (??):30–36, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306431.

Digal:2017:MAI

[DPB17] Larry N. Digal, Shemaiah Gail P. Placencia, and Carol Q. Balgos. Market assessment on the incentives and disincentives for the adoption of sustainable practices along the tuna value chain in Region 12, Philippines. Marine Policy, 86(??):39–46, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301197.

Diogo:2015:HED

[DPH^{+}15] Hugo Diogo, João G. Pereira, Ruth M. Higgins, Ângela Canha, and Dália Reis. History, effort distribution and landings in an artisanal bottom longline fishery: an empirical study from the North Atlantic Ocean. Marine Policy, 51(??):75–85, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001961.

Dichmont:2013:CFG

[DPJ^{+}13] Catherine Mary Dichmont, Sean Pascoe, Edward Jebreen, Rachel Pears, Kate Brooks, and Pascal Perez. Choosing a fishery’s governance structure using data poor methods. Marine Policy, 37(??):123–131, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000358.

Diogo:2016:CMI

[DPS16] Hugo Diogo, João Gil Pereira, and Mara Schmiing. Catch me if you can: Non-compliance of limpet protection in the Azores. Marine Policy, 63(??):92–99, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002912.

Dennis:2015:CBF

[DPV^{+}15] Darren Dennis, Éva Plagányi, Ingrid Van Putten, Trevor Hutton, and Sean Pascoe. Cost benefit of fisheryindependent surveys: Are they worth the money? Marine Policy, 58(??):108–115, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000986.

Davies:2010:BCE

[DR10] R. W. D. Davies and R. Rangeley. Banking on cod: Exploring economic incentives for recovering Grand Banks and North Sea cod fisheries. Marine Policy, 34(1):92–98, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000712.

Ditty:2014:UUC

[DR14] J. M. Ditty and C. E. Rezende. Unjust and unsustainable: a case study of the Açu port industrial complex. Marine Policy, 45(??):82–88, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002777.

Draper:2015:NCR

[Dra15] Kelsey Draper. Networks of capital: Reframing knowledge in the Namibian hake fishery. Marine Policy, 60 (??):293–299, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002309.

Dey:2016:AEI

[DRG^{+}16] Madan Mohan Dey, Mark W. Rosegrant, Kamal Gosh, Oai Li Chen, and Rowena Valmonte-Santos. Analysis of the economic impact of climate change and climate change adaptation strategies for fisheries sector in Pacific coral triangle countries: Model, estimation strategy, and baseline results. Marine Policy, 67(??):156–163, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003814.

Da-Rocha:2017:SCF

[DRGCPS17] José-María Da-Rocha, Javier García-Cutrín, Raúl Prellezo, and Jaume Sempere. The social cost of fishery subsidy reforms. Marine Policy, 83(??):236–242, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302154.

Da-Rocha:2019:BGA

[DRGP19] José-María Da-Rocha, Jordi Guillen, and Raúl Prellezo. (Blue) Growth accounting in small-scale European Union fleets. Marine Policy, 100(??):200–206, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305232.

Dinesen:2018:ITQ

[DRH^{+}18] Grete E. Dinesen, Iben Wiene Rathje, Mathilde Højrup, Francois Bastardie, Finn Larsen, Thomas Kirk Sørensen, Erik Hoffmann, and Ole R. Eigaard. Individual transferable quotas, does one size fit all? Sustainability analysis of an alternative model for quota allocation in a smallscale coastal fishery. Marine Policy, 88(??):23–31, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730146X.

DiasAlves:2018:EAF

[DRMA18] Lázaro Dias Alves, Eduardo Manuel Rosa Bulhões, Ana Paula Madeira Di Beneditto, and Camilah Antunes Zappes. Ethnoclimatology of Artisanal fishermen: Interference in coastal fishing in southeastern Brazil. Marine Policy, 95(??):69–76, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308527.

Dromgoole:2013:RPM

[Dro13] Sarah Dromgoole. Reflections on the position of the major maritime powers with respect to the UNESCO Convention on the Protection of the Underwater Cultural Heritage 2001. Marine Policy, 38(??):116–123, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200125X.

Da-Rocha:2017:DEE

[DRPSA17] José-María Da-Rocha, Raúl Prellezo, Jaume Sempere, and Luis Taboada Antelo. A dynamic economic equilibrium model for the economic assessment of the fishery stockrebuilding policies. Marine Policy, 81(??):185–195, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630820X.

Druon:2010:HMA

[Dru10] Jean-Noël Druon. Habitat mapping of the Atlantic bluefin tuna derived from satellite data: Its potential as a tool for the sustainable management of pelagic fisheries. Marine Policy, 34(2):293–297, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001055.

Dolman:2010:TBE

[DS10] Sarah Dolman and Mark Simmonds. Towards best environmental practice for cetacean conservation in developing Scotland’s marine renewable energy. Marine Policy, 34(5):1021–1027, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000527.

daSilva:2017:BRA

[dS17a] Alexandre Pereira da Silva. Brazil’s recent agenda on the sea and the South Atlantic contemporary scenario. Marine Policy, 85(??):25–32, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307564.

Drymon:2017:API

[DS17b] J. Marcus Drymon and Steven B. Scyphers. Attitudes and perceptions influence recreational angler support for shark conservation and fisheries sustainability. Marine Policy, 81(??):153–159, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307941.

Doloreux:2018:MMC

[DS18] David Doloreux and Richard Shearmur. Moving maritime clusters to the next level: Canada’s Ocean Supercluster initiative. Marine Policy, 98(??):33–36, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303506.

Doloreux:2016:QCM

[DSF16] David Doloreux, Richard Shearmur, and Diego Figueiredo. Québec’ coastal maritime cluster: Its impact on regional economic development, 2001–2011. Marine Policy, 71(??): 201–209, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301713.

DeAlessi:2014:LRI

[DSH14] Michael De Alessi, Joseph M. Sullivan, and Ray Hilborn. The legal, regulatory, and institutional evolution of fishing cooperatives in Alaska and the West Coast of the United States. Marine Policy, 43(??):217–225, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001322.

Dutra:2019:GMF

[DSH^{+}19] Leo X. C. Dutra, Ilva Sporne, Marcus Haward, Shankar Aswani, Kevern L. Cochrane, Stewart Frusher, Maria A. Gasalla, Sônia M. F. Gianesella, Tanith Grant, Alistair J. Hobday, Sarah Jennings, Éva Plagányi, Gretta Pecl, Shyam S. Salim, Warwick Sauer, Manuela B. Taboada, and Ingrid E. van Putten. Governance mapping: a framework for assessing the adaptive capacity of marine resource governance to environmental change. Marine Policy, 106(??):Article 103392, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301398.

Dale:2019:CMI

[DSK^{+}19] Pat Dale, Ilva Sporne, Jon Knight, Marcus Sheaves, Leila Eslami-Andergoli, and Patrick Dwyer. A conceptual model to improve links between science, policy and practice in coastal management. Marine Policy, 103(??):42–49, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830383X.

DeFreitas:2013:SSS

[DSM^{+}13] Débora M. De Freitas, Stephen G. Sutton, James M. Moloney, Elodie J. I. Lédée, and Renae C. Tobin. Spatial substitution strategies of recreational fishers in response to zoning changes in the Great Barrier Reef Marine Park. Marine Policy, 40(??):145–153, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000134.

Dailianis:2018:HAR

[DSP^{+}18] Thanos Dailianis, Christopher J. Smith, Nadia Papadopoulou, Vasilis Gerovasileiou, Katerina Sevastou, Trine Bekkby, Meri Bilan, David Billett, Christoffer Boström, Marina Carreiro-Silva, Roberto Danovaro, Simonetta Fraschetti, Karine Gagnon, Cristina Gambi, Anthony Grehan, Silvija Kipson, Jonne Kotta, Chris J. McOwen, Telmo Morato, Henn Ojaveer, Christopher K. Pham, and Rachael Scrimgeour. Human activities and resultant pressures on key European marine habitats: an analysis of mapped resources. Marine Policy, 98 (??):1–10, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302884.

DellApa:2012:MSA

[DSR12] Andrea Dell’Apa, Lisa Schiavinato, and Roger A. Rulifson. The Magnuson–Stevens act (1976) and its reauthorizations: Failure or success for the implementation of fishery sustainability and management in the US? Marine Policy, 36(3):673–680, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001825.

Dygico:2013:AME

[DSWG13] Marivel Dygico, Angelique Songco, Alan T. White, and Stuart J. Green. Achieving MPA effectiveness through application of responsive governance incentives in the Tubbataha reefs. Marine Policy, 41(??):87–94, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002722.

Driscoll:2010:FUG

[DT10] John Driscoll and Peter Tyedmers. Fuel use and greenhouse gas emission implications of fisheries management: the case of the New England Atlantic herring fishery. Marine Policy, 34(3):353–359, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001122.

Dalton:2010:MHD

[DTJ10] Tracey Dalton, Robert Thompson, and Di Jin. Mapping human dimensions in marine spatial planning and management: an example from Narragansett Bay, Rhode Island. Marine Policy, 34(2):309–319, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001080.

Dominguez-Tejo:2018:PDG

[DTM18] Elianny Domínguez-Tejo and Graciela Metternicht. Poorlydesigned goals and objectives in resource management plans: Assessing their impact for an Ecosystem-Based Approach to Marine Spatial Planning. Marine Policy, 88 (??):122–131, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302361.

Dominguez-Tejo:2016:MSP

[DTMJH16] Elianny Domínguez-Tejo, Graciela Metternicht, Emma Johnston, and Luke Hedge. Marine Spatial Planning advancing the Ecosystem-Based Approach to coastal zone management: a review. Marine Policy, 72(??):115–130, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300161.

Diedrich:2010:BSS

[DTN10] A. Diedrich, J. Tintoré, and F. Navinés. Balancing science and society through establishing indicators for integrated coastal zone management in the Balearic Islands. Marine Policy, 34(4):772–781, July 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000187.

Dupont:2014:RBM

[Dup14] Diane P. Dupont. Rights-based management in Canada: Lessons from two coasts and a centre. Marine Policy, 44 (??):60–64, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001577.

Dribek:2017:CVA

[DV17] Abderraouf Dribek and Louinord Voltaire. Contingent valuation analysis of willingness to pay for beach erosion control through the stabiplage technique: a study in Djerba (Tunisia). Marine Policy, 86(??):17–23, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300404.

Dvarskas:2019:EEA

[Dva19] Anthony Dvarskas. Experimental ecosystem accounting for coastal and marine areas: a pilot application of the SEEA-EEA in Long Island coastal bays. Marine Policy, 100(??):141–151, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303233.

deVos:2016:NMF

[dVDA^{+}16] B. I. de Vos, R. Döring, M. Aranda, F. C. Buisman, K. Frangoudes, L. Goti, C. Macher, C. D. Maravelias, A. Murillas-Maza, O. van der Valk, and P. Vasilakopoulos. New modes of fisheries governance: Implementation of the landing obligation in four European countries. Marine Policy, 64(??):1–8, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003243.

deVivero:2010:OGC

[dVM10a] Juan L. Suárez de Vivero and Juan C. Rodríguez Mateos. Ocean governance in a competitive world. The BRIC countries as emerging maritime powers — building new geopolitical scenarios. Marine Policy, 34(5):967–978, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000448.

deVos:2010:CTR

[dVM10b] Birgit I. de Vos and Arthur P. J. Mol. Changing trust relations within the Dutch fishing industry: the case of National Study Groups. Marine Policy, 34(5):887– 895, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000126.

deVivero:2012:SAM

[dVM12] Juan Luis Suárez de Vivero and Juan Carlos Rodríguez Mateos. The Spanish approach to marine spatial planning. Marine Strategy Framework Directive vs. EU Integrated Maritime Policy. Marine Policy, 36(1):18–27, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000522.

Vivero:2014:CMS

[dVM14] Juan L. Suárez de Vivero and Juan C. Rodríguez Mateos. Changing maritime scenarios. The geopolitical dimension of the EU Atlantic Strategy. Marine Policy, 48 (??):59–72, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000888.

Vivero:2017:FGR

[dVM17] Juan L. Suárez de Vivero and Juan C. Rodríguez Mateos. Forecasting geopolitical risks: Oceans as source of instability. Marine Policy, 75(??):19–28, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306571.

Vivero:2019:FSM

[dVMdC^{+}19] Juan L. Suárez de Vivero, Juan C. Rodríguez Mateos, David Florido del Corral, María José Barragán, Helena Calado, Marian Kjellevold, and Ewelina Janowska Miasik. Food security and maritime security: a new challenge for the European Union’s ocean policy. Marine Policy, 108(??):Article 103640, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830962X.

deVos:2011:TRB

[dVvT11] Birgit I. de Vos and Jan P. M. van Tatenhove. Trust relationships between fishers and government: New challenges for the co-management arrangements in the Dutch flatfish industry. Marine Policy, 35(2):218–225, March 2011.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001703.

Davis:2013:IPA

[DW13] Brendal Davis and Boris Worm. The International Plan of Action for Sharks: How does national implementation measure up? Marine Policy, 38(??):312–320, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001467.

Davies:2018:DSM

[DWCS18] Peter Davies, Chris Williams, Griffin Carpenter, and Bryce D. Stewart. Does size matter? Assessing the use of vessel length to manage fisheries in England. Marine Policy, 97(??):202–210, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830068X.

Diedhiou:2019:SDO

[DYN^{+}19] Idrissa Diedhiou, Zhengyong Yang, Mansor Ndour, Moustapha Dème, Massal Fall, Modou Thiaw, Ndiaga Thiam, and Sheng Li. Socioeconomic dimension of the octopus “Octopus vulgaris” in the context of fisheries management of both small-scale and industrial fisheries in Senegal. Marine Policy, 106(??):Article 103517, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306171.

DuBois:2012:CSB

[DZ12] Carolyn DuBois and Christos Zografos. Conflicts at sea between artisanal and industrial fishers: Inter-sectoral interactions and dispute resolution in Senegal. Marine Policy, 36(6):1211–1220, November 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000516.

Dong:2015:AIC

[DZLL15] Bingying Dong, Ling Zhu, Kevin Li, and Meifeng Luo. Acceptance of the international compensation regime for tanker oil pollution — And its implications for China. Marine Policy, 61(??):179–186, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002201.

Ertor-Akyazi:2019:FVI

[EA19] Pinar Ertör-Akyazi. Formal versus informal institutions: Extraction and earnings in framed field experiments with small-scale fishing communities in Turkey. Marine Policy, 109(??):Article 103673, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301691.

Epstein:2018:HDE

[EAA^{+}18] Graham Epstein, Evan Andrews, Derek Armitage, Paul Foley, Jeremy Pittman, and Rebecca Brushett. Human dimensions of ecosystem-based management: Lessons in managing trade-offs from the Northern Shrimp Fishery in Northern Peninsula, Newfoundland. Marine Policy, 97 (??):10–17, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308667.

Eagling:2015:IRO

[EAE15] Lawrence E. Eagling, Elizabeth C. Ashton, and Josh Eagle. The incentives of a resource owner: Evidence from a private oyster fishery. Marine Policy, 58(??):28–35, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500086X.

Easman:2018:APA

[EAG18] Emily S. Easman, Kirsten E. Abernethy, and Brendan J. Godley. Assessing public awareness of marine environmental threats and conservation efforts. Marine Policy, 87(??):234–240, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303615.

Eliasen:2016:FSR

[EB16] Søren Qvist Eliasen and Nikolaj Bichel. Fishers sharing real-time information about “bad” fishing locations. A tool for quota optimisation under a regime of landing obligations. Marine Policy, 64(??):16–23, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003267.

Ebel:2018:PPC

[EBR^{+}18] Sarah A. Ebel, Christine M. Beitl, Jocelyn Runnebaum, Robin Alden, and Teresa R. Johnson. The power of participation: Challenges and opportunities for facilitating trust in cooperative fisheries research in the Maine lobster fishery. Marine Policy, 90(??):47–54, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730492X.

Evans:2018:ERN

[EBS18] Karen Evans, Nicholas J. Bax, and David C. Smith. Enhancing the robustness of a national assessment of the marine environment. Marine Policy, 98(??):133–145, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302070.

Eichert:2018:ERF

[ECF^{+}18] Moritz Eichert, Aida Campos, Paulo Fonseca, Pedro Lopes, Lino Marques, and Margarida Castro. Effects of reallocating fishing effort from trawling to creels in a Norway lobster fishery. Marine Policy, 93(??):142–149, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307662.

Eriksson:2015:GSS

[ECL^{+}15] Hampus Eriksson, Chantal Conand, Alessandro Lovatelli, Nyawira A. Muthiga, and Steven W. Purcell. Governance structures and sustainability in Indian Ocean sea cucumber fisheries. Marine Policy, 56(??):16–22, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000305.

Evariste:2016:CEC

[ECR^{+}16] Emmanuelle Evariste, Pascal Claquin, Jean-Paul Robin, Arnaud Auber, Abigail McQuatters-Gollop, and Jean-Claude Dauvin. The Channel ecosystem, a cross-roads of anthropogenic pressures and scientific studies: Lessons learned from the European INTERREG IV projects (2009– 2015). Marine Policy, 63(??):158–165, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002961.

Ellis:2017:EMF

[ECRL17] J. I. Ellis, M. R. Clark, H. L. Rouse, and G. Lamarche. Environmental management frameworks for offshore mining: the New Zealand approach. Marine Policy, 84(??): 178–192, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306972.

Ewell:2017:PES

[ECSE^{+}17] Christopher Ewell, Sarika Cullis-Suzuki, Mikaela Ediger, John Hocevar, Dana Miller, and Jennifer Jacquet. Potential ecological and social benefits of a moratorium on transshipment on the high seas. Marine Policy, 81(??):293–300, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300623.

Eddy:2014:OHF

[Edd14] Tyler D. Eddy. One hundred-fold difference between perceived and actual levels of marine protection in New Zealand. Marine Policy, 46(??):61–67, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000062.

Edwards:2017:LBT

[EE17] Danielle N. Edwards and Daniel G. Edwards. Licence Banks as a tool to mitigate corporate control of fisheries: a British Columbia groundfish example. Marine Policy, 80(??):141–146, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307278.

Eliasen:2019:LOC

[EFK^{+}19] Søren Qvist Eliasen, Jordan Feekings, Ludvig Krag, Tiago Veiga-Malta, Lars O. Mortensen, and Clara Ulrich. The landing obligation calls for a more flexible technical gear regulation in EU waters — Greater industry involvement could support development of gear modifications. Marine Policy, 99(??):173–180, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302926.

Estevez:2015:PMC

[EG15] Rodrigo A. Estévez and Stefan Gelcich. Participative multi-criteria decision analysis in marine management and conservation: Research progress and the challenge of integrating value judgments and uncertainty. Marine Policy, 61(??):1–7, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001815.

Eriksen:2018:MPB

[EGD18] Torkild Eriksen, Harm Greidanus, and Conor Delaney. Metrics and provider-based results for completeness and temporal resolution of satellite-based AIS services. Marine Policy, 93(??):80–92, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305535.

Evans:2017:SPM

[EGF^{+}17] Ally J. Evans, Brian Garrod, Louise B. Firth, Stephen J. Hawkins, Elisabeth S. Morris-Webb, Harry Goudge, and Pippa J. Moore. Stakeholder priorities for multi-functional coastal defence developments and steps to effective implementation. Marine Policy, 75(??):143–155, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303645.

Emery:2012:ICR

[EGGT12] Timothy J. Emery, Bridget S. Green, Caleb Gardner, and John Tisdell. Are input controls required in individual transferable quota fisheries to address ecosystem based fisheries management objectives? Marine Policy, 36(1):122–131, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000777.

Emery:2016:RGI

[EGHC16] Timothy J. Emery, Caleb Gardner, Klaas Hartmann, and Ian Cartwright. The role of government and industry in resolving assignment problems in fisheries with individual transferable quotas. Marine Policy, 73(??):46–52, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302019.

Emery:2017:BSG

[EGHC17a] Timothy J. Emery, Caleb Gardner, Klaas Hartmann, and Ian Cartwright. Beyond sustainability: is government obliged to increase economic benefit from fisheries in the face of industry resistance? Marine Policy, 76 (??):48–54, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305450.

Emery:2017:IEF

[EGHC17b] Timothy J. Emery, Caleb Gardner, Klaas Hartmann, and Ian Cartwright. Incorporating economics into fisheries management frameworks in Australia. Marine Policy, 77 (??):136–143, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305000.

Elliott:2018:IBH

[EGP^{+}18] Sophie A. M. Elliott, Laurent Guérin, Roland Pesch, Petra Schmitt, Bryony Meakins, Cristina Vina-Herbon, Jose M. González-Irusta, Ana de la Torriente, and Alberto Serrano. Integrating benthic habitat indicators: Working towards an ecosystem approach. Marine Policy, 90(??):88–94, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306371.

Eliasen:2015:DIR

[EHR15] Søren Q. Eliasen, Troels J. Hegland, and Jesper Raakjær. Decentralising: the implementation of regionalisation and co-management under the post-2013 Common Fisheries Policy. Marine Policy, 62(??):224–232, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002730.

Eriksson:2019:SEI

[EJB19] Björn Eriksson, Frank Johansson, and Malgorzata Blicharska. Socio-economic impacts of marine conservation efforts in three Indonesian fishing communities. Marine Policy, 103(??):59–67, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304810.

Englender:2014:CCC

[EKSZ14] Dorota Englender, Jenny Kirschey, Aleke Stöfen, and Andreas Zink. Cooperation and compliance control in areas beyond national jurisdiction. Marine Policy, 49(??): 186–194, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002819.

Eero:2012:WEB

[EKV12] Margit Eero, Friedrich W. Köster, and Morten Vinther. Why is the Eastern Baltic cod recovering? Marine Policy, 36(1):235–240, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000996.

Escobar:2019:DBF

[ELAC19] Rosana Escobar, Andrea Luna-Acosta, and Susana Caballero. DNA barcoding, fisheries and communities: What do we have? Science and local knowledge to improve resource management in partnership with communities in the Colombian Caribbean. Marine Policy, 99(??): 407–413, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301891.

Eliasen:2014:CAA

[Eli14] Søren Qvist Eliasen. Cod avoidance by area regulations in Kattegat — experiences for the implementation of a discard ban in the EU. Marine Policy, 45(??):108–113, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002790.

Elofsson:2010:CEB

[Elo10] Katarina Elofsson. Cost-effectiveness of the Baltic Sea Action Plan. Marine Policy, 34(5):1043–1050, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000552.

Eikeset:2018:WBG

[EMD^{+}18] Anne Maria Eikeset, Anna B. Mazzarella, Brynhildur Davídsdóttir, Dane H. Klinger, Simon A. Levin, Elena Rovenskaya, and Nils Chr. Stenseth. What is blue growth? The semantics of “Sustainable Development” of marine environments. Marine Policy, 87(??):177–179, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306905.

Engel:2014:PAS

[EMP^{+}14] Mônica Tais Engel, Silvio Marchini, Ana Carolina Pont, Rodrigo Machado, and Larissa Rosa de Oliveira. Perceptions and attitudes of stakeholders towards the wildlife refuge of Ilha dos Lobos, a marine protected area in Brazil. Marine Policy, 45(??):45–51, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002674.

Engstrom:2018:CBS

[Eng18] Viljam Engström. Complexities of the Baltic Sea regulatory framework. Marine Policy, 98(??):191–200, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306754.

Emery:2019:CLR

[ENWL19] Timothy J. Emery, Rocio Noriega, Ashley J. Williams, and James Larcombe. Changes in logbook reporting by commercial fishers following the implementation of electronic monitoring in Australian Commonwealth fisheries. Marine Policy, 104(??):135–145, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307218.

Ertor:2015:PLE

[EOC15] Irmak Ertör and Miquel Ortega-Cerdà. Political lessons from early warnings: Marine finfish aquaculture conflicts in Europe. Marine Policy, 51(??):202–210, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001924.

Edwards:2019:RIC

[EP19] Danielle N. Edwards and Evelyn Pinkerton. Rise of the investor class in the British Columbia Pacific halibut fishery. Marine Policy, 109(??):Article 103676, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302696.

Ek:2014:RBS

[ERB14] Åsa Ek, Marcus Runefors, and Jonas Borell. Relationships between safety culture aspects — a work process to enable interpretation. Marine Policy, 44(??):179–186, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001784.

Espinosa-Romero:2011:SDM

[ERCMD11] Maria J. Espinosa-Romero, Kai M. A. Chan, Timothy McDaniels, and Denise M. Dalmer. Structuring decision-making for ecosystem-based management. Marine Policy, 35(5):575–583, September 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000200.

Eikeset:2013:BEA

[ERD^{+}13] Anne Maria Eikeset, Andries P. Richter, Dorothy J. Dankel, Erin S. Dunlop, Mikko Heino, Ulf Dieckmann, and Nils Chr. Stenseth. A bio-economic analysis of harvest control rules for the Northeast Arctic cod fishery. Marine Policy, 39(??):172–181, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002199.

Ermolin:2018:ASC

[ES18] Ilya Ermolin and Linas Svolkinas. Assessment of the sturgeon catches and seal bycatches in an IUU fishery in the Caspian Sea. Marine Policy, 87(??):284–290, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303329.

Edwards:2013:INR

[ESGC13] P. E. T. Edwards, A. E. Sutton-Grier, and G. E. Coyle. Investing in nature: Restoring coastal habitat blue infrastructure and green job creation. Marine Policy, 38(??):65–71, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001182.

Elliott:2019:SRT

[ESM^{+}19] Kathryn Elliott, Helen C. M. Smith, Fran Moore, Adriaan H. van der Weijde, and Iraklis Lazakis. A systematic review of transferable solution options for the environmental impacts of tidal lagoons. Marine Policy, 99 (??):190–200, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830294X.

Ellis:2019:CFS

[ET19] John Ellis and Rachel Tiller. Conceptualizing future scenarios of integrated multi-trophic aquaculture (IMTA) in the Norwegian salmon industry. Marine Policy, 104 (??):198–209, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300162.

Espinoza-Tenorio:2015:AIA

[ETEW15] Alejandro Espinoza-Tenorio, Ileana Espejel, and Matthias Wolff. From adoption to implementation? An academic perspective on Sustainable Fisheries Management in a developing country. Marine Policy, 62(??):252–260, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002523.

Espinoza-Tenorio:2011:CFI

[ETEWZD11] Alejandro Espinoza-Tenorio, Ileana Espejel, Matthias Wolff, and José Alberto Zepeda-Domínguez. Contextual factors influencing sustainable fisheries in Mexico. Marine Policy, 35(3):343–350, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000182X.

Elliott:2018:RVM

[EWG^{+}18] Lewis R. Elliott, Mathew P. White, James Grellier, Siân E. Rees, Ruth D. Waters, and Lora E. Fleming. Recreational visits to marine and coastal environments in England: Where, what, who, why, and when? Marine Policy, 97(??):305–314, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301520.

Evans:2015:OFM

[EYN^{+}15] K. Evans, J. W. Young, S. Nicol, D. Kolody, V. Allain, J. Bell, J. N. Brown, A. Ganachaud, A. J. Hobday, B. Hunt, J. Innes, A. Sen Gupta, E. van Sebille, R. Kloser, T. Patterson, and A. Singh. Optimising fisheries management in relation to tuna catches in the western central Pacific Ocean: a review of research priorities and opportunities. Marine Policy, 59(??):94–104, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001335.

Edwards:2019:MMM

[EZBL19] Peter Edwards, Wenbo Zhang, Ben Belton, and David C. Little. Misunderstandings, myths and mantras in aquaculture: Its contribution to world food supplies has been systematically over reported. Marine Policy, 106(??): Article 103547, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300788.

Exarchopoulos:2018:SWC

[EZPRZ18] Georgios Exarchopoulos, Pengfei Zhang, Nicola Pryce-Roberts, and Minghua Zhao. Seafarers’ welfare: a critical review of the related legal issues under the Maritime Labour Convention 2006. Marine Policy, 93(??):62–70, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308382.

Foale:2013:FSC

[FAA^{+}13] Simon Foale, Dedi Adhuri, Porfiro Aliño, Edward H. Allison, Neil Andrew, Philippa Cohen, Louisa Evans, Michael Fabinyi, Pedro Fidelman, Christopher Gregory, Natasha Stacey, John Tanzer, and Nireka Weeratunge. Food security and the Coral Triangle Initiative. Marine Policy, 38(??):174–183, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001315.

Fabinyi:2016:PCL

[Fab16a] Michael Fabinyi. Producing for Chinese luxury seafood value chains: Different outcomes for producers in the Philippines and North America. Marine Policy, 63(??): 184–190, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000731.

Fabinyi:2016:SSC

[Fab16b] Michael Fabinyi. Sustainable seafood consumption in China. Marine Policy, 74(??):85–87, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303402.

Fraser:2017:PPP

[FAPP17] Kerrie A. Fraser, Vanessa M. Adams, Robert L. Pressey, and John M. Pandolfi. Purpose, policy, and practice: Intent and reality for on-ground management and outcomes of the Great Barrier Reef Marine Park. Marine Policy, 81(??):301–311, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300763.

Farady:2019:MNU

[Far19] Susan E. Farady. Microplastics as a new, ubiquitous pollutant: Strategies to anticipate management and advise seafood consumers. Marine Policy, 104(??):103–107, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830633X.

Finkbeiner:2015:RDC

[FB15] Elena M. Finkbeiner and Xavier Basurto. Re-defining comanagement to facilitate small-scale fisheries reform: an illustration from northwest Mexico. Marine Policy, 51 (??):433–441, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002723.

Frangoudes:2017:FOM

[FB17] Katia Frangoudes and Manuel Bellanger. Fishers’ opinions on marketization of property rights and the quota system in France. Marine Policy, 80(??):107–112, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730012X.

Fromentin:2014:SUM

[FBAK14] Jean-Marc Fromentin, Sylvain Bonhommeau, Haritz Arrizabalaga, and Laurence T. Kell. The spectre of uncertainty in management of exploited fish stocks: the illustrative case of Atlantic bluefin tuna. Marine Policy, 47(??):8–14, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1400030X.

Fraser-Baxter:2018:EMR

[FBM18] Sam Fraser-Baxter and Fabien Medvecky. Evaluating the media’s reporting of public and political responses to human–shark interactions in N.S.W., Australia. Marine Policy, 97(??):109–118, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300423.

Famuditi:2018:AMC

[FBP^{+}18] Taye Famuditi, Malcolm Bray, Jonathan Potts, Brian Baily, and Robert Inkpen. Adaptive management and community reaction: the activities of Coastal Action Groups (CAGs) within the shoreline management process in England. Marine Policy, 97(??):270–277, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304445.

Figus:2019:CSR

[FC19a] Elizabeth Figus and Keith R. Criddle. Comparing self-reported incidental catch among fishermen targeting Pacific halibut and a fishery independent survey. Marine Policy, 100(??):371–381, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730667X.

Ford:2019:PIC

[FC19b] J. Ford and D. Clark. Preparing for the impacts of climate change along Canada’s Arctic coast: the importance of search and rescue. Marine Policy, 108(??): Article 103662, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300260.

Fenberg:2012:SEM

[FCC^{+}12] Phillip B. Fenberg, Jennifer E. Caselle, Joachim Claudet, Michaela Clemence, Steven D. Gaines, Jose Antonio García-Charton, Emanuel J. Gonçalves, Kirsten Grorud-Colvert, Paolo Guidetti, Stuart R. Jenkins, Peter J. S. Jones, Sarah E. Lester, Rob McAllen, Even Moland, Serge Planes, and Thomas K. Sørensen. The science of European marine reserves: Status, efficacy, and future needs. Marine Policy, 36(5):1012–1021, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000383.

Ferreira:2015:CTM

[FCdS^{+}15] Maria Adelaide Ferreira, Helena Calado, Carlos Pereira da Silva, António Domingos Abreu, Francisco Andrade, Catarina Fonseca, Emanuel J. Gonçalves, José Guerreiro, Francisco Noronha, Margarida Pereira, Carlos Pinto Lopes, Marta Chantal Ribeiro, Yorgos Stratoudakis, and Lia Vasconcelos. Contributions towards maritime spatial planning (MSP) in Portugal — Conference report. Marine Policy, 59(??):61–63, September 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000998.

Fabinyi:2011:PPL

[FD11] Michael Fabinyi and Dante Dalabajan. Policy and practice in the live reef fish for food trade: a case study from Palawan, Philippines. Marine Policy, 35(3):371–378, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1000196X.

Freed:2016:ESS

[FDGM16] Sarah Freed, Veronica Dujon, Elise F. Granek, and Jaffar Mouhhidine. Enhancing small-scale fisheries management through community engagement and multi-community partnerships: Comoros case study. Marine Policy, 63 (??):81–91, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002882.

Fabinyi:2018:MBF

[FDP18] Michael Fabinyi, Wolfram H. Dressler, and Michael D. Pido. Moving beyond financial value in seafood commodity chains. Marine Policy, 94(??):89–92, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308904.

Forse:2019:TMU

[FDP19] Andy Forse, Benjamin Drakeford, and Jonathan Potts. Towards managing the United Kingdom’s fisheries: a Brexit view from Scotland. Marine Policy, 109(??):Article 103709, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19300855.

Froese:2015:OIG

[FDS15] Rainer Froese, Nazli Demirel, and Arlene Sampang. An overall indicator for the good environmental status of marine waters based on commercially exploited species. Marine Policy, 51(??):230–237, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001869.

Fidelman:2012:MSI

[FE12] Pedro Fidelman and Julia A. Ekstrom. Mapping seascapes of international environmental arrangements in the Coral Triangle. Marine Policy, 36(5):993–1004, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000139.

Fidelman:2012:GLS

[FEF^{+}12] Pedro Fidelman, Louisa Evans, Michael Fabinyi, Simon Foale, Josh Cinner, and Franciska Rosen. Governing largescale marine commons: Contextual challenges in the Coral Triangle. Marine Policy, 36(1):42–53, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000662.

Fenton:2017:RPC

[Fen17] Paul Fenton. The role of port cities and transnational municipal networks in efforts to reduce greenhouse gas emissions on land and at sea from shipping

— an assessment of the World Ports Climate Initiative. Marine Policy, 75(??):271–277, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003826.

Folkersen:2018:DSM

[FFH18a] Maja Vinde Folkersen, Christopher M. Fleming, and Syezlin Hasan. Deep sea mining’s future effects on Fiji’s tourism industry: a contingent behaviour study. Marine Policy, 96(??):81–89, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300836.

Folkersen:2018:EVD

[FFH18b] Maja Vinde Folkersen, Christopher M. Fleming, and Syezlin Hasan. The economic value of the deep sea: a systematic review and meta-analysis. Marine Policy, 94 (??):71–80, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300940.

Fissel:2015:DPE

[FFKO15] Benjamin E. Fissel, Ronald G. Felthoven, Stephen Kasperski, and Christopher O’Donnell. Decomposing productivity and efficiency changes in the Alaska head and gut factory trawl fleet. Marine Policy, 62(??):337–346, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001773.

Fan:2017:CCC

[FFSF17a] Lidong Fan, Jiangang Fei, Ulf Schriever, and Si Fan. The communicative competence of Chinese seafarers and their employability in the international maritime labour market. Marine Policy, 83(??):137–145, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308168.

Fan:2017:CES

[FFSF17b] Lidong Fan, Jiangang Fei, Ulf Schriever, and Si Fan. A critical examination of the seafaring English education and training in China. Marine Policy, 86(??):56–63, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303937.

Fan:2018:ESC

[FFSF18] Lidong Fan, Jiangang Fei, Ulf Schriever, and Si Fan. An empirical study on the communicative competence of Chinese seafarers. Marine Policy, 87(??):65–71, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302671.

Filer:2018:HCN

[FG18] Colin Filer and Jennifer Gabriel. How could Nautilus Minerals get a social licence to operate the world’s first deep sea mine? Marine Policy, 95(??):394–400, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306625.

Freire-Gibb:2014:GSW

[FGKMP14] Lucio Carlos Freire-Gibb, Rebecca Koss, Piotr Margonski, and Nadia Papadopoulou. Governance strengths and weaknesses to implement the marine strategy framework directive in European waters. Marine Policy, 44(??): 172–178, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001796.

Ferrer:2018:ILB

[FGPP18] Alice Joan G. Ferrer, Len Garces, Maripaz Perez, and Robert Pomeroy. Inter-local bay alliances in Northern Mindanao, Philippines: Experiences and lessons learned in establishment and organisational sustainability. Marine Policy, 94(??):81–88, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301167.

Friedman:2018:MBF

[FGR18] K. Friedman, S. M. Garcia, and J. Rice. Mainstreaming biodiversity in fisheries. Marine Policy, 95(??):209– 220, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305444.

Fare:2015:PCF

[FGW15] Rolf Färe, Shawna Grosskopf, and John Walden. Productivity change and fleet restructuring after transition to individual transferable quota management. Marine Policy, 62(??):318–325, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001578.

Fare:2017:MCV

[FGW17] Rolf Färe, Shawna Grosskopf, and John B. Walden. Measuring capital value in a commercial fishery: a distance function approach. Marine Policy, 81(??):109–115, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300064.

Fritz:2015:EIM

[FH15] Jan-Stefan Fritz and John Hanus. The European Integrated Maritime Policy: the next five years. Marine Policy, 53(??):1–4, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002978.

Floysand:2016:DRP

[FHJ16] Arnt Fløysand, Kristine Håland, and Stig-Erik Jakobsen. Discourses, risk perceptions and the “green” profile of the New Zealand salmon farming industry. Marine Policy, 74(??):230–235, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306285.

Flannery:2018:ENP

[FHL18] Wesley Flannery, Noel Healy, and Marcos Luna. Exclusion and non-participation in Marine Spatial Planning. Marine Policy, 88(??):32–40, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730324X.

Foley:2010:GEP

[FHM^{+}10] Melissa M. Foley, Benjamin S. Halpern, Fiorenza Micheli, Matthew H. Armsby, Margaret R. Caldwell, Caitlin M. Crain, Erin Prahler, Nicole Rohr, Deborah Sivas, Michael W. Beck, Mark H. Carr, Larry B. Crowder, J. Emmett Duffy, Sally D. Hacker, Karen L. McLeod, Stephen R. Palumbi, Charles H. Peterson, Helen M. Regan, Mary H. Ruckelshaus, Paul A. Sandifer, and Robert S. Steneck. Guiding ecological principles for marine spatial planning. Marine Policy, 34(5):955–966, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000436.

Frost:2018:RMP

[FHN^{+}18] Robyn Frost, Paul Hibberd, Masio Nidung, Emily Artack, and Marie Bourrel. Redrawing the map of the Pacific. Marine Policy, 95(??):302–310, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630330X.

Fleming:2012:CAG

[FJ12] D. M. Fleming and P. J. S. Jones. Challenges to achieving greater and fairer stakeholder involvement in marine spatial planning as illustrated by the Lyme Bay scallop dredging closure. Marine Policy, 36(2):370–377, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001217.

Freestone:2014:CEI

[FJA^{+}14] David Freestone, David Johnson, Jeff Ardron, Kate Killerlain Morrison, and Sebastian Unger. Can existing institutions protect biodiversity in areas beyond national jurisdiction? Experiences from two on-going processes. Marine Policy, 49(??):167–175, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300290X.

Fletcher:2014:EEM

[FJG^{+}14a] Stephen Fletcher, Rebecca Jefferson, Gillian Glegg, Lynda Rodwell, and Wendy Dodds. England’s evolving marine and coastal governance framework. Marine Policy, 45 (??):261–268, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002042.

Friedrich:2014:PPS

[FJG14b] Laura A. Friedrich, Rebecca Jefferson, and Gillian Glegg. Public perceptions of sharks: Gathering support for shark conservation. Marine Policy, 47(??):1–7, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000360.

Fernandez:2016:ISI

[FKMV16] Linda Fernandez, Brooks Kaiser, Sue Moore, and Niels Vestergaard. Introduction to special issue: Arctic marine resource governance. Marine Policy, 72(??):237–239, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302329.

Foster:2019:GST

[FKWV19] Sarah J. Foster, Ting-Chun Kuo, Anita Kar Yan Wan, and Amanda C. J. Vincent. Global seahorse trade defies export bans under CITES action and national legislation. Marine Policy, 103(??):33–41, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302197.

Flaaten:2010:FRC

[Fla10] Ola Flaaten. Fisheries rent creation and distribution — the imaginary case of Codland. Marine Policy, 34(6):

1268–1272, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001028.

Flaaten:2013:IQC

[Fla13] Ola Flaaten. Institutional quality and catch performance of fishing nations. Marine Policy, 38(??):267–276, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001418.

Francis:2011:PPF

[FLH11] Tessa B. Francis, Phillip S. Levin, and Chris J. Harvey. The perils and promise of futures analysis in marine ecosystem-based management. Marine Policy, 35(5): 675–681, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000431.

French:2019:CAS

[FLT^{+}19] R. P. French, J. M. Lyle, W. M. Twardek, S. J. Cooke, and J. M. Semmens. A characterization of Australian shortfin mako shark anglers. Marine Policy, 110(??):Article 103550, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305992.

Forster:2011:MBC

[FLWG11] J. Forster, I. R. Lake, A. R. Watkinson, and J. A. Gill. Marine biodiversity in the Caribbean UK overseas territories: Perceived threats and constraints to environmental management. Marine Policy, 35(5):647–657, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000297.

Forster:2014:MDL

[FLWG14] J. Forster, I. R. Lake, A. R. Watkinson, and J. A. Gill. Marine dependent livelihoods and resilience to environmental change: a case study of Anguilla. Marine Policy, 45(??):204–212, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1300242X.

Fletcher:2013:EPM

[FMB^{+}13] Stephen Fletcher, Emma McKinley, Kenneth C. Buchan, Ness Smith, and Karen McHugh. Effective practice in marine spatial planning: a participatory evaluation of experience in Southern England. Marine Policy, 39(??):341–348, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001789.

Finkbeiner:2018:ETO

[FMB^{+}18] Elena M. Finkbeiner, Fiorenza Micheli, Nathan J. Bennett, Adam L. Ayers, Elodie Le Cornu, and Angee N. Doerr. Exploring trade-offs in climate change response in the context of Pacific Island fisheries. Marine Policy, 88(??):359–364, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301872.

Fujita:2012:ROT

[FMD^{+}12] Rod Fujita, Alexander C. Markham, Julio E. Diaz Diaz, Julia Rosa Martinez Garcia, Courtney Scarborough, Patrick Greenfield, Peter Black, and Stacy E. Aguilera. Revisiting ocean thermal energy conversion. Marine Policy, 36(2):463–465, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000972.

Fujita:2013:MRO

[FMD^{+}13] R. Fujita, J. H. Moxley, H. DeBey, T. Van Leuvan, A. Leumer, K. Honey, S. Aguilera, and M. Foley. Managing for a resilient ocean. Marine Policy, 38(??):538–544, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001236.

Fauville:2018:UCI

[FMD^{+}18] Géraldine Fauville, Patricia McHugh, Christine Domegan, Åsa Mäkitalo, Lene Friis Møller, Martha Papathanassiou, Carla Alvarez Chicote, Susana Lincoln, Vanessa Batista, Evy Copejans, Fiona Crouch, and Susan Gotensparre. Using collective intelligence to identify barriers to teaching 12–19 year olds about the ocean in Europe. Marine Policy, 91(??):85–96, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307443.

Fernandez-Macho:2016:IAM

[FMGV16] Javier Fernández-Macho, Pilar González, and Jorge Virto. An index to assess maritime importance in the European Atlantic economy. Marine Policy, 64(??):72–81, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003346.

Frisch:2015:GPO

[FMKT15] L. C. Frisch, J. T. Mathis, N. P. Kettle, and S. F. Trainor. Gauging perceptions of ocean acidification in Alaska. Marine Policy, 53(??):101–110, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003236.

Fanning:2013:ALM

[FMM13] Lucia Fanning, Robin Mahon, and Patrick McConney. Applying the large marine ecosystem (LME) governance framework in the Wider Caribbean Region. Marine Policy, 42(??):99–110, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000511.

Fernandez-Macho:2015:MME

[FMMA^{+}15] Javier Fernández-Macho, Arantza Murillas, Alberto Ansuategi, Marta Escapa, Carmen Gallastegui, Pilar González, Raul Prellezo, and Jorge Virto. Measuring the maritime economy: Spain in the European Atlantic Arc. Marine Policy, 60(??):49–61, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001438.

Foley:2015:GEC

[FMN15] Paul Foley, Charles Mather, and Barbara Neis. Governing enclosure for coastal communities: Social embeddedness in a Canadian shrimp fishery. Marine Policy, 61(??): 390–400, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003030.

Forsman:2018:FHW

[FMP^{+}18] Z. H. Forsman, P. Maurin, M. Parry, A. Chung, C. Sartor, M. A. Hixon, K. Hughes, K. Rodgers, I. S. S. Knapp, D. A. Gulko, E. C. Franklin, L. Del Rio Torres, N. T. Chan, C. S. Wolke, R. D. Gates, and R. J. Toonen. The first Hawai’i workshop for coral restoration & nurseries. Marine Policy, 96(??):133–135, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301143.

Flores:2017:MTE

[FMS17] Cristina Pallero Flores, Juan Manuel Barragán Muñoz, and Marinez Eymael García Scherer. Management of transboundary estuaries in Latin America and the Caribbean. Marine Policy, 76(??):63–70, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305656.

Flannery:2012:RMS

[FÓ12] Wesley Flannery and Micheál Ó Cinnéide. A roadmap for marine spatial planning: a critical examination of the European Commission’s guiding principles based on their application in the Clyde MSP Pilot Project. Marine Policy, 36(1):265–271, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001059.

Fock:2011:NEC

[Foc11] Heino O. Fock. Natura 2000 and the European Common Fisheries Policy. Marine Policy, 35(2):181–188, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001648.

Falaleeva:2011:TCA

[FOG^{+}11] Maria Falaleeva, Cathal O’Mahony, Stefan Gray, Margaret Desmond, Jeremy Gault, and Valerie Cummins. Towards climate adaptation and coastal governance in Ireland: Integrated architecture for effective management? Marine Policy, 35(6):784–793, December 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000066.

Foley:2013:MIF

[Fol13] C. M. R. Foley. Management implications of fishing up, down, or through the marine food web. Marine Policy, 37 (??):176–182, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000851.

Flannery:2015:ECT

[FOO^{+}15] Wesley Flannery, Anne Marie O’Hagan, Cathal O’Mahony, Heather Ritchie, and Sarah Twomey. Evaluating conditions for transboundary Marine Spatial Planning: Challenges and opportunities on the island of Ireland. Marine Policy, 51(??):86–95, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400195X.

Froese:2012:ELA

[FP12] Rainer Froese and Alexander Proelss. Evaluation and legal assessment of certified seafood. Marine Policy, 36(6): 1284–1289, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000619.

Froese:2013:SOI

[FP13] Rainer Froese and Alexander Proelss. Is a stock overfished if it is depleted by overfishing? A response to the rebuttal of Agnew et al. to Froese and Proelss “Evaluation and legal assessment of certified seafood”.

Marine Policy, 38(??):548–550, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001613.

Fouqueray:2019:EAC

[FP19] Mondane Fouqueray and Elissaios Papyrakis. An empirical analysis of the cross-national determinants of marine protected areas. Marine Policy, 99(??):87–93, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307571.

Fiedler:2017:PLF

[FPG^{+}17] Fernando N. Fiedler, Dagoberto Port, Bruno B. Giffoni, Gilberto Sales, and Fabiane Fisch. Pelagic longline fisheries in southeastern/south Brazil. Who cares about the law? Marine Policy, 77(??):56–64, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308211.

Fernandes:2015:DCI

[FPP^{+}15] Ana Cláudia Fernandes, Nélida Pérez, Nuno Prista, Juan Santos, and Manuela Azevedo. Discards composition from Iberian trawl fleets. Marine Policy, 53(??):33–44, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002747.

Froese:2013:RRC

[FQ13] Rainer Froese and Martin Quaas. Rio+20 and the reform of the Common Fisheries Policy in Europe. Marine Policy, 39(??):53–55, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002059.

Fauconnet:2016:FSI

[FR16] Laurence Fauconnet and Marie-Joëlle Rochet. Fishing selectivity as an instrument to reach management objectives in an ecosystem approach to fisheries. Marine Policy, 64(??):46–54, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003231.

Fritz:2010:TNF

[Fri10] Jan-Stefan Fritz. Towards a ‘new form of governance’ in science-policy relations in the European Maritime Policy. Marine Policy, 34(1):1–6, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000505.

Farr:2014:NCT

[FSB14] Marina Farr, Natalie Stoeckl, and Rabiul Alam Beg. The non-consumptive (tourism) ‘value’ of marine species in the Northern section of the Great Barrier Reef. Marine Policy, 43(??):89–103, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001061.

Francois:2019:EHT

[FSCJM19] Poisson François, Catteau Sidonie, Chiera Caroline, and Groul Jean-Marc. The effect of hook type and trailing gear on hook shedding and fate of pelagic stingray (Pteroplatytrygon violacea): New insights to develop effective mitigation approaches. Marine Policy, 107(??):Article 103594, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307437.

Ferrol-Schulte:2015:CLV

[FSGB^{+}15] Daniella Ferrol-Schulte, Philipp Gorris, Wasistini Baitoningsih, Dedi S. Adhuri, and Sebastian C. A. Ferse. Coastal livelihood vulnerability to marine resource degradation: a review of the Indonesian national coastal and marine policy framework. Marine Policy, 52(??):163–171, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002590.

Fox:2012:EGP

[FSM^{+}12] Helen E. Fox, Carrie S. Soltanoff, Michael B. Mascia, Kelly M. Haisfield, Alfonso V. Lombana, Christopher R. Pyke, and Louisa Wood. Explaining global patterns and trends in marine protected area (MPA) development. Marine Policy, 36(5):1131–1138, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000140.

Fletcher:2010:EBF

[FSMG10] W. J. Fletcher, J. Shaw, S. J. Metcalf, and D. J. Gaughan. An Ecosystem Based Fisheries Management framework: the efficient, regional-level planning tool for management agencies. Marine Policy, 34(6):1226–1238, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000849.

Farr:2014:RFB

[FSS14] Marina Farr, Natalie Stoeckl, and Stephen Sutton. Recreational fishing and boating: Are the determinants the same? Marine Policy, 47(??):126–137, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000529.

Filby:2015:SSV

[FSS15] Nicole E. Filby, Karen A. Stockin, and Carol Scarpaci. Social science as a vehicle to improve dolphin-swim tour operation compliance? Marine Policy, 51(??):40–47, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14001845.

Filby:2017:CMP

[FSS17] Nicole E. Filby, Karen A. Stockin, and Carol Scarpaci. Can Marine Protected Areas be developed effectively without baseline data? A case study for Burrunan dolphins (Tursiops australis). Marine Policy, 77(??):152–163, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300100.

Fernandes:2016:CBE

[FSV^{+}16] Jose A. Fernandes, Lionel Santos, Thomas Vance, Tim Fileman, David Smith, John D. D. Bishop, Frédérique Viard, Ana M. Queirós, Gorka Merino, Erik Buisman, and Melanie C. Austen. Costs and benefits to European shipping of ballast-water and hull-fouling treatment: Impacts of native and non-indigenous species. Marine Policy, 64(??):148–155, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003590.

Ferrol-Schulte:2013:SLA

[FSWFG13] Daniella Ferrol-Schulte, Matthias Wolff, Sebastian Ferse, and Marion Glaser. Sustainable Livelihoods Approach in tropical coastal and marine social–ecological systems: a review. Marine Policy, 42(??):253–258, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000833.

Forster:2017:ECU

[FTF^{+}17] Johanna Forster, Rachel A. Turner, Clare Fitzsimmons, Angelie M. Peterson, Robin Mahon, and Selina M. Stead. Evidence of a common understanding of proximate and distal drivers of reef health. Marine Policy, 84(??): 263–272, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630731X.

Ferra:2018:MCB

[FTG^{+}18] Carmen Ferrà, Anna Nora Tassetti, Fabio Grati, Giulio Pellini, Piero Polidori, Giuseppe Scarcella, and Gianna Fabi. Mapping change in bottom trawling activity in the Mediterranean Sea through AIS data. Marine Policy, 94(??):275–281, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301859.

Foster:2010:TST

[FV10] S. J. Foster and A. C. J. Vincent. Tropical shrimp trawl fisheries: Fishers’ knowledge of and attitudes about a doomed fishery. Marine Policy, 34(3):437–446, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001456.

Freitag:2018:BST

[FVH18] Amy Freitag, Bruce Vogt, and Troy Hartley. Breaking stereotypes through network analysis of the Chesapeake oyster community. Marine Policy, 90(??):146–151, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17304360.

Fernandez-Vidal:2014:FFA

[FVM14] Duarte Fernández-Vidal and Ramón Muiño. Fact or fiction? Assessing governance and co-management of Marine Reserves of Fishing Interest in Cedeira and Lira (NW Spain). Marine Policy, 47(??):15–22, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000281.

Foley:2011:RFI

[FvRA11] Naomi S. Foley, Tom M. van Rensburg, and Claire W. Armstrong. The rise and fall of the Irish orange roughy fishery: an economic analysis. Marine Policy, 35(6):756–763, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000042.

Freitag:2019:YTW

[FVT19] Amy Freitag, James Vasslides, and Howard Townsend. Are you thinking what I’m thinking? A conceptual modeling approach to understand stakeholders’ assessments of the fate of Chesapeake oysters. Marine Policy, 99 (??):99–110, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302719.

Faure:2015:CVE

[FW15] Michael Faure and Hui Wang. Compensating victims of a European Deepwater Horizon accident: OPOL revisited. Marine Policy, 62(??):25–36, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002365.

Ford:2019:SLD

[FW19] Jessica H. Ford and Chris Wilcox. Shedding light on the dark side of maritime trade — a new approach for identifying countries as flags of convenience. Marine Policy, 99(??):298–303, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303257.

Froese:2018:SRE

[FWC^{+}18] Rainer Froese, Henning Winker, Gianpaolo Coro, Nazli Demirel, Athanassios C. Tsikliras, Donna Dimarchopoulou, Giuseppe Scarcella, Martin Quaas, and Nele Matz-Lück. Status and rebuilding of European fisheries. Marine Policy, 93(??):159–170, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307364.

Fleming:2017:SDG

[FWHS17] Aysha Fleming, Russell M. Wise, Heidi Hansen, and Linda Sams. The sustainable development goals: a case study. Marine Policy, 86(??):94–103, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304414.

Freeman:2016:APS

[FWK16] Mikaela C. Freeman, Libby Whiting, and Ryan P. Kelly. Assessing potential spatial and temporal conflicts in Washington’s marine waters. Marine Policy, 70(??):137–144, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302512.

Girard:2014:EOM

[GA14] Sophie Girard and José A. Pérez Agúndez. The effects of the oyster mortality crisis on the economics of the shellfish farming sector: Preliminary review and prospects from a case study in Marennes–Oleron Bay (France). Marine Policy, 48(??):142–151, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000967.

Goti-Aralucea:2019:ASE

[GA19] Leyre Goti-Aralucea. Assessing the social and economic impact of small scale fisheries management measures in a marine protected area with limited data. Marine Policy, 101(??):246–256, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308430.

Gullestad:2017:TEB

[GAB^{+}17] Peter Gullestad, Anne Marie Abotnes, Gunnstein Bakke, Mette Skern-Mauritzen, Kjell Nedreaas, and Guldborg Søvik. Towards ecosystem-based fisheries management in Norway — Practical tools for keeping track of relevant issues and prioritising management efforts. Marine Policy, 77(??):104–110, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305383.

Guillen:2019:ASE

[GAC^{+}19] Jordi Guillen, Frank Asche, Natacha Carvalho, José M. Fernández Polanco, Ignacio Llorente, Rasmus Nielsen, Max Nielsen, and Sebastian Villasante. Aquaculture subsidies in the European Union: Evolution, impact and future potential for growth. Marine Policy, 104(??):19–28, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309400.

Gerhartz-Abraham:2016:ICC

[GAFAV16] Adrian Gerhartz-Abraham, Lucia M. Fanning, and Jorge Angulo-Valdes. ICZM in Cuba: Challenges and opportunities in a changing economic context. Marine Policy, 73(??):69–76, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301968.

Goti-Aralucea:2018:OSO

[GAFD^{+}18] Leyre Goti-Aralucea, Mike Fitzpatrick, Ralf Döring, David Reid, John Mumford, and Anna Rindorf. “Overarching sustainability objectives overcome incompatible directions in the Common Fisheries Policy”. Marine Policy, 91(??):49– 57, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306681.

Glaus:2019:FPF

[GAKP^{+}19] Kerstin B. J. Glaus, Irene Adrian-Kalchhauser, Susanna Piovano, Sharon A. Appleyard, Juerg M. Brunnschweiler, and Ciro Rico. Fishing for profit or food? Socioeconomic drivers and fishers’ attitudes towards sharks in Fiji. Marine Policy, 100(??):249–257, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303270.

Gammell:2011:MSH

[Gam11] Alistair Gammell. More a safe haven than fortress. Marine Policy, 35(5):728, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X11000133.

Gregr:2012:RCE

[GAP12] Edward J. Gregr, Andrea L. Ahrens, and R. Ian Perry. Reconciling classifications of ecologically and biologically significant areas in the world’s oceans. Marine Policy, 36(3):716–726, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1100159X.

Garibaldi:2012:FGC

[Gar12] Luca Garibaldi. The FAO global capture production database: a six-decade effort to catch the trend. Marine Policy, 36(3):760–768, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X11001928.

Gattringer:2018:RCP

[Gat18] Clemens W. Gattringer. A revisited conceptualization of plastic pollution accumulation in marine environments: Insights from a social ecological economics perspective. Marine Policy, 96(??):221–226, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730516X.

Gauci:2010:UKM

[Gau10] Gotthard Mark Gauci. The U.K. Marine and Coastal Access Bill — a missed opportunity to enhance protection from marine environmental pollution? Marine Policy, 34(3):498–505, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001675.

Grantham:2013:CZA

[GAW^{+}13] Hedley S. Grantham, Vera N. Agostini, Joanne Wilson, Sangeeta Mangubhai, Nur Hidayat, Andreas Muljadi, Muhajir, Chris Rotinsulu, Meity Mongdong, Michael W. Beck, and Hugh P. Possingham. A comparison of zoning analyses to inform the planning of a marine protected area network in Raja Ampat, Indonesia. Marine Policy, 38(??):184–194, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001339.

Galappaththi:2015:CCM

[GB15] Eranga K. Galappaththi and Fikret Berkes. Can comanagement emerge spontaneously? Collaborative management in Sri Lankan shrimp aquaculture. Marine Policy, 60(??):1–8, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001426.

Gullestad:2015:DBP

[GBBB15] P. Gullestad, G. Blom, G. Bakke, and B. Bogstad. The “Discard Ban Package”: Experiences in efforts to improve the exploitation patterns in Norwegian fisheries. Marine Policy, 54(??):1–9, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002589.

Gonzalez-Bernat:2017:LOB

[GBC17] Maria J. Gonzalez-Bernat and Julian Clifton. “Living with our backs to the sea”: a critical analysis of marine and coastal governance in Guatemala. Marine Policy, 81(??):9– 20, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307163.

Glaser:2010:WST

[GBF^{+}10] Marion Glaser, Wasistini Baitoningsih, Sebastian C. A. Ferse, Muhammad Neil, and Rio Deswandi. Whose sustainability? Top–down participation and emergent rules in marine protected area management in Indonesia. Marine Policy, 34(6):1215–1225, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000837.

Grimm:2012:ACS

[GBF^{+}12] Dietmar Grimm, Ivan Barkhorn, David Festa, Kate Bonzon, Judd Boomhower, Valerie Hovland, and Jason Blau. Assessing catch shares’ effects evidence from Federal United States and associated British Columbian fisheries. Marine Policy, 36(3):644–657, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001643.

Giakoumi:2015:UTM

[GBK^{+}15] Sylvaine Giakoumi, Christopher J. Brown, Stelios Katsanevakis, Megan I. Saunders, and Hugh P. Possingham. Using threat maps for cost-effective prioritization of actions to conserve coastal habitats. Marine Policy, 61 (??):95–102, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002055.

Gansbauer:2016:SPS

[GBW16] Anja Gänsbauer, Ulrike Bechtold, and Harald Wilfing. SoFISHticated policy — social perspectives on the fish conflict in the Northeast Atlantic. Marine Policy, 66 (??):93–103, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000191.

Gerwing:2017:ETG

[GC17] Travis G. Gerwing and Kieran Cox. Erosion of trust in government consultation will impede the creation of environmental policy. Marine Policy, 83(??):126–127, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302774.

Gillam:2018:FBS

[GC18] Cintia Gillam and Anthony Charles. Fishers in a Brazilian Shantytown: Relational wellbeing supports recovery from environmental disaster. Marine Policy, 89(??): 77–84, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16308533.

Grorud-Colvert:2019:HPI

[GCCSS^{+}19] Kirsten Grorud-Colvert, Vanessa Constant, Jenna Sullivan-Stack, Katherine Dziedzic, Sara L. Hamilton, Zachary Randell, Heather Fulton-Bennett, Zechariah D. Meunier, Silke Bachhuber, Alissa J. Rickborn, Barbara Spiecker, and Jane Lubchenco. High-profile international commitments for ocean protection: Empty promises or meaningful progress? Marine Policy, 105(??):52–66, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307371.

Guerreiro:2011:RIE

[GCD^{+}11] José Guerreiro, Aldo Chircop, David Dzidzornu, Catarina Grilo, Raquel Ribeiro, Rudy van der Elst, and Ana Viras. The role of international environmental instruments in enhancing transboundary marine protected areas: an approach in East Africa. Marine Policy, 35(2):95–104, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001193.

Griffin:2015:IVC

[GCD^{+}15] Robert Griffin, Nicolas Chaumont, Douglas Denu, Anne Guerry, Choong-Ki Kim, and Mary Ruckelshaus. Incorporating the visibility of coastal energy infrastructure into multi-criteria siting decisions. Marine Policy, 62(??): 218–223, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002754.

Guerreiro:2010:ETN

[GCG^{+}10] José Guerreiro, Aldo Chircop, Catarina Grilo, Ana Viras, Raquel Ribeiro, and Rudy van der Elst. Establishing a transboundary network of marine protected areas: Diplomatic and management options for the east African context. Marine Policy, 34(5):896–910, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000151.

Grilli:2019:AVS

[GCHO19] Gianluca Grilli, John Curtis, Stephen Hynes, and Paul O’Reilly. Anglers’ views on stock conservation: Sea bass angling in Ireland. Marine Policy, 99(??):34–41, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308321.

Gray:2011:SCM

[GCK^{+}11] Darcy L. Gray, Rosaline R. Canessa, C. Peter Keller, Philip Dearden, and Rick B. Rollins. Spatial characterization of marine recreational boating: Exploring the use of an on-the-water questionnaire for a case study in the Pacific Northwest. Marine Policy, 35(3):286–298, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001739.

Grech:2011:BSA

[GCM11] A. Grech, R. Coles, and H. Marsh. A broad-scale assessment of the risk to coastal seagrasses from cumulative threats. Marine Policy, 35(5):560–567, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000534.

Gonzalez-Carman:2012:LIT

[GCMA^{+}12] Victoria González-Carman, Natalia Machain, Diego Albareda, Hermes Mianzan, and Claudio Campagna. Legal and institutional tools to mitigate marine turtle bycatch: Argentina as a case study. Marine Policy, 36(6): 1265–1274, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000589.

Garcia-Cegarra:2019:CRA

[GCP19] Ana M. García-Cegarra and Aldo S. Pacheco. Collision risk areas between fin and humpback whales with large cargo vessels in Mejillones Bay (23^{o}S), northern Chile. Marine Policy, 103(??):182–186, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300861.

Graziano:2018:PSC

[GCW^{+}18] Armando Graziano, Pierre Cariou, François-Charles Wolff, Maximo Q. Mejia, and Jens-Uwe Schröder-Hinrichs. Port state control inspections in the European Union: Do inspector’s number and background matter? Marine Policy, 88(??):230–241, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305626.

Gilmour:2011:BIQ

[GDD11] Patrick W. Gilmour, Peter D. Dwyer, and Robert W. Day. Beyond individual quotas: the role of trust and cooperation in promoting stewardship of five Australian abalone fisheries. Marine Policy, 35(5):692–702, September 2011.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000455.

Gilmour:2013:EAF

[GDD13] Patrick W. Gilmour, Peter D. Dwyer, and Robert W. Day. Enhancing the agency of fishers: a conceptual model of self-management in Australian abalone fisheries. Marine Policy, 37(??):165–175, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200084X.

Gulas:2017:DAO

[GDD^{+}17] Sarah Gulas, Mitchell Downton, Kareina D’Souza, Kelsey Hayden, and Tony R. Walker. Declining Arctic Ocean oil and gas developments: Opportunities to improve governance and environmental pollution control. Marine Policy, 75(??):53–61, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630416X.

Gulas:2019:CDA

[GDD^{+}19] Sarah Gulas, Mitchell Downton, Kareina D’Souza, Kelsey Hayden, and Tony R. Walker. Corrigendum to “Declining Arctic Ocean oil and gas developments: Opportunities to improve governance and environmental pollution control” [Mar. Pol. 75 (2017) 53–61]. Marine Policy, 105(??):177–179, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302064.

Gourlie:2018:PNS

[GDG^{+}18] Don Gourlie, Ruth Davis, Hugh Govan, James Marshman, and Quentin Hanich. Performing “A New Song”: Suggested Considerations for Drafting Effective Coastal Fisheries Legislation Under Climate Change. Marine Policy, 88(??):342–349, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301288.

Gavrilov:2019:AUN

[GDN19] Viatcheslav Gavrilov, Roman Dremliuga, and Rustambek Nurimbetov. Article 234 of the 1982 United Nations Convention on the law of the sea and reduction of ice cover in the Arctic Ocean. Marine Policy, 106(??):Article 103518, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308297.

Gissi:2016:EMS

[GdV16] Elena Gissi and Juan Luis Suarez de Vivero. Exploring marine spatial planning education: Challenges in structuring transdisciplinarity. Marine Policy, 74(??):43–57, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301129.

Gren:2017:CSN

[GE17] Ing-Marie Gren and Katarina Elofsson. Credit stacking in nutrient trading markets for the Baltic Sea. Marine Policy, 79(??):1–7, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307965.

Gekara:2010:SNU

[Gek10] Victor Oyaro Gekara. The stamp of neoliberalism on the UK tonnage tax and the implications for British seafaring. Marine Policy, 34(3):487–494, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001651.

Germond:2015:GDM

[Ger15] Basil Germond. The geopolitical dimension of maritime security. Marine Policy, 54(??):137–142, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003509.

Graziano:2018:ESP

[GFA^{+}18] Marcello Graziano, Clive J. Fox, Karen Alexander, Cristina Pita, J. J. Heymans, Margaret Crumlish, Adam Hughes, Joly Ghanawi, and Lorenzo Cannella. Environmental and socio-political shocks to the seafood sector: What does this mean for resilience? Lessons from two UK case studies, 1945–2016. Marine Policy, 87(??):301–313, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304864.

Gopnik:2012:CTE

[GFC^{+}12] Morgan Gopnik, Clare Fieseler, Laura Cantral, Kate Mc-Clellan, Linwood Pendleton, and Larry Crowder. Coming to the table: Early stakeholder engagement in marine spatial planning. Marine Policy, 36(5):1139–1149, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1200019X.

Garla:2015:PAE

[GFC^{+}15] Ricardo C. Garla, Renato H. A. Freitas, Janaina F. Calado, Gustavo B. C. Paterno, and Adriana R. Carvalho. Public awareness of the economic potential and threats to sharks of a tropical oceanic archipelago in the western South Atlantic. Marine Policy, 60(??):128–133, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001712.

Godwin:2017:TEV

[GFH^{+}17] Sean C. Godwin, Fiona T. Francis, Brett R. Howard, Luis Malpica-Cruz, and Allison L. Witter. Towards the economic viability of local seafood programs: Key features for the financial performance of community supported fisheries. Marine Policy, 81(??):375–380, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630505X.

Garcia-Florez:2014:NSA

[GFMG^{+}14] Lucía García-Flórez, Jesús Morales, Miguel B. Gaspar, David Castilla, Estanis Mugerza, Patrick Berthou, Laura García de la Fuente, Manuela Oliveira, Oscar Moreno, Juan José García del Hoyo, Luis Arregi, Celine Vignot, Rosa Chapela, and Arantza Murillas. A novel and simple approach to define artisanal fisheries in Europe. Marine Policy, 44(??):152–159, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001759.

Garza-Gil:2017:IGE

[GGADSR17] M. Dolores Garza-Gil, Lucy Amigo-Dobaño, and Juan C. Surís-Regueiro. Institutions and governance in the European Common Fisheries Policy: an empirical study of Spanish fishers’ attitudes toward greater participation. Marine Policy, 79(??):33–39, May 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303505.

Germond:2016:OGM

[GGD16] Basil Germond and Celine Germond-Duret. Ocean governance and maritime security in a placeful environment: the case of the European Union. Marine Policy, 66 (??):124–131, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000154.

Galappaththi:2017:CSM

[GGK17] Iroshani M. Galappaththi, Eranga K. Galappaththi, and Sarath S. Kodithuwakku. Can start-up motives influence social-ecological resilience in community-based entrepreneurship setting? Case of coastal shrimp farmers in Sri Lanka. Marine Policy, 86(??):156–163, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303226.

Greaker:2017:PRR

[GGL17] Mads Greaker, Kristine Grimsrud, and Lars Lindholt. The potential resource rent from Norwegian fisheries. Marine Policy, 84(??):156–166, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301719.

Goldsmith:2016:BTW

[GGLP16] Kaitlin Goldsmith, Elise Granek, Amy Lubitow, and Michael Papenfus. Bridge over troubled waters: a synthesis session to connect scientific and decision making sectors. Marine Policy, 70(??):30–39, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301877.

Garza-Gil:2017:UIO

[GGSRVL17] Ma Dolores Garza-Gil, Juan C. Surís-Regueiro, and Manuel M. Varela-Lafuente. Using input–output methods to assess the effects of fishing and aquaculture on a regional economy: the case of Galicia, Spain. Marine Policy, 85(??):48–53, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302592.

Garza-Gil:2016:MAE

[GGVRVL16] M. Dolores Garza-Gil, Ma. Xosé Vázquez-Rodríguez, and Manuel M. Varela-Lafuente. Marine aquaculture and environment quality as perceived by Spanish consumers. The case of shellfish demand. Marine Policy, 74(??):1–5, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630210X.

Gnanalingam:2015:FTF

[GH15] Gaya Gnanalingam and Chris Hepburn. Flexibility in temporary fisheries closure legislation is required to maximise success. Marine Policy, 61(??):39–45, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002018.

Gallagher:2018:ECS

[GH18] Austin J. Gallagher and Charlie P. M. Huveneers. Emerging challenges to shark-diving tourism. Marine Policy, 96(??):9–12, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302264.

Gianelli:2018:OEA

[GHM^{+}18] I. Gianelli, S. Horta, G. Martínez, A. de la Rosa, and O. Defeo. Operationalizing an ecosystem approach to small-scale fisheries in developing countries: the case of Uruguay. Marine Policy, 95(??):180–188, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300101.

Grealis:2017:EIA

[GHO^{+}17] Eoin Grealis, Stephen Hynes, Cathal O’Donoghue, Amaya Vega, Suzanne Van Osch, and Cian Twomey. The economic impact of aquaculture expansion: an input-output approach. Marine Policy, 81(??):29–36, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300453.

Gormley:2015:AMI

[GHP^{+}15] Kate S. G. Gormley, Angela D. Hull, Joanne S. Porter, Michael C. Bell, and William G. Sanderson. Adaptive management, international co-operation and planning for marine conservation hotspots in a changing climate. Marine Policy, 53(??):54–66, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003182.

Gibbs:2010:WIT

[Gib10] Mark T. Gibbs. Why ITQs on target species are inefficient at achieving ecosystem based fisheries management outcomes. Marine Policy, 34(3):708–709, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001407.

Gillett:2011:RPS

[Gil11a] Robert Gillett. Replacing purse seining with pole-and-line fishing in the central and Western Pacific: Some aspects of the baitfish requirements. Marine Policy, 35(2):148–154, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001533.

Gilman:2011:BGB

[Gil11b] Eric L. Gilman. Bycatch governance and best practice mitigation technology in global tuna fisheries. Marine Policy, 35(5):590–609, September 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000224.

Gilbert:2014:RST

[Gil14] Paul Gilbert. From reductionism to systems thinking: How the shipping sector can address sulphur regulation and tackle climate change. Marine Policy, 43(??):376–378, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001516.

Gilman:2015:SIM

[Gil15] Eric Gilman. Status of international monitoring and management of abandoned, lost and discarded fishing gear and ghost fishing. Marine Policy, 60(??):225–239, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1500175X.

Geijer:2015:NAM

[GJ15] Christina K. A. Geijer and Peter J. S. Jones. A network approach to migratory whale conservation: Are MPAs the way forward or do all roads lead to the IMO? Marine Policy, 51(??):1–12, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001663.

Gunawardena:2016:PSC

[GJA16] Nuwan D. P. Gunawardena, Tuanthong Jutagate, and Upali S. Amarasinghe. Patterns of species composition of beach seine fisheries off North-Western coast of Sri Lanka, fishers’ perceptions and implications for comanagement. Marine Policy, 72(??):131–138, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302561.

Galappaththi:2016:CEM

[GKG16] Eranga K. Galappaththi, Sarath S. Kodithuwakku, and Iroshani M. Galappaththi. Can environment management integrate into supply chain management? Information sharing via shrimp aquaculture cooperatives in northwestern Sri Lanka. Marine Policy, 68(??):187–194, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300963.

Gray:2011:QDD

[GKSJ11] Tim Gray, R. C. Korda, Selina Stead, and Estelle Jones. Quota discarding and distributive justice: the case of the under-10m fishing fleet in Sussex, England. Marine Policy, 35(2):122–129, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001508.

Glazewski:2019:LPC

[Gla19] Jan Glazewski. Legal and practical challenges around restitution, secrecy and asset recovery in transnational fisheries crime: a case study of United States v Bengis, 2013. Marine Policy, 105(??):151–157, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309515.

Glegg:2014:TMP

[Gle14] Gillian Glegg. Training for marine planners: Present and future needs. Marine Policy, 43(??):13–20, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000870.

Gustavsson:2014:PDJ

[GLJdlTC14] Madeleine Gustavsson, Lars Lindström, Narriman S. Jiddawi, and Maricela de la Torre-Castro. Procedural and distributive justice in a community-based managed Marine Protected Area in Zanzibar, Tanzania. Marine Policy, 46(??):91–100, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000074.

Gardner:2017:IUO

[GLR17] Charlie J. Gardner, Julia E. Latham, and Steve Rocliffe. Intended and unintended outcomes in fisheries learning exchanges: Lessons from Mexico and Madagascar. Marine Policy, 77(??):219–226, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302354.

Garnacho:2010:TER

[GLSA10] Eva Garnacho, Robin J. Law, Ronny Schallier, and Joan Albaiges. Targeting European R&D for accidental marine pollution. Marine Policy, 34(5):1068–1075, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000588.

Ghilardi-Lopes:2015:PCT

[GLTB^{+}15] Natalia Pirani Ghilardi-Lopes, Alexander Turra, Marcos S. Buckeridge, Amanda Cristina Silva, Flávio Augusto de Souza Berchez, and Valéria Marques de Oliveira. On the perceptions and conceptions of tourists with regard to global environmental changes and their consequences for coastal and marine environments: a case study of the northern São Paulo State coast, Brazil. Marine Policy, 57(??):85–92, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000548.

Garcia-Lorenzo:2019:APS

[GLVLGG19] Iria García-Lorenzo, Manuel María Varela-Lafuente, and M. Dolores Garza-Gil. Adaptative processes in smallscale traditional fishermen’s organisations. The case of Cofradías in Galicia (NW Spain). Marine Policy, 99 (??):382–390, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304548.

Guest:2015:YSO

[GLW15] Haley Guest, Heike K. Lotze, and Douglas Wallace. Youth and the sea: Ocean literacy in Nova Scotia, Canada. Marine Policy, 58(??):98–107, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000895.

Georgianna:2017:CTN

[GLW17] Daniel Georgianna, Min-Yang Lee, and John Walden. Contrasting trends in the Northeast United States groundfish and scallop processing industries. Marine Policy, 85(??):100–106, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730266X.

Guillen:2014:ITS

[GM14] Jordi Guillen and Francesc Maynou. Importance of temporal and spatial factors in the ex-vessel price formation for red shrimp and management implications. Marine Policy, 47(??):66–70, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000426.

Germond:2019:CCM

[GM19] Basil Germond and Antonios D. Mazaris. Climate change and maritime security. Marine Policy, 99(??):262–266, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306365.

Gascuel:2012:TII

[GMD^{+}12] D. Gascuel, G. Merino, R. Döring, J. N. Druon, L. Goti, S. Guénette, C. Macher, K. Soma, M. Travers-Trolet, and S. Mackinson. Towards the implementation of an integrated ecosystem fleet-based management of European fisheries. Marine Policy, 36(5):1022–1032, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000152.

Gianelli:2015:EAS

[GMD15] Ignacio Gianelli, Gastón Martínez, and Omar Defeo. An ecosystem approach to small-scale co-managed fisheries: the yellow clam fishery in Uruguay. Marine Policy, 62(??):196–202, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002766.

Guillen:2013:EMM

[GMM^{+}13] Jordi Guillen, Claire Macher, Mathieu Merzéréaud, Michel Bertignac, Spyros Fifas, and Olivier Guyader. Estimating MSY and MEY in multi-species and multi-fleet fisheries, consequences and limits: an application to the Bay of Biscay mixed fishery. Marine Policy, 40(??):64–74, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002709.

Guanais:2015:DFA

[GMM15] José Hugo Gondim Guanais, Rodrigo Pereira Medeiros, and Patrick A. McConney. Designing a framework for addressing bycatch problems in Brazilian small-scale trawl fisheries. Marine Policy, 51(??):111–118, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400178X.

Gormley:2014:CME

[GMP^{+}14] Kate S. G. Gormley, Lauren H. McWhinnie, Joanne S. Porter, Angela D. Hull, Teresa F. Fernandes, and William G. Sanderson. Can management effort be predicted for marine protected areas? New considerations for network design. Marine Policy, 47(??):138–146, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000372.

Gorospe:2016:MST

[GMP^{+}16] Kelvin D. Gorospe, William Michaels, Robert Pomeroy, Christopher Elvidge, Patrick Lynch, Supin Wongbusarakum, and Russell E. Brainard. The mobilization of science and technology fisheries innovations towards an ecosystem approach to fisheries management in the Coral Triangle and Southeast Asia. Marine Policy, 74(??):143–152, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302093.

Galland:2018:ICS

[GNHM18] Grantly R. Galland, Amanda E. M. Nickson, Rachel Hopkins, and Shana K. Miller. On the importance of clarity in scientific advice for fisheries management. Marine Policy, 87(??):250–254, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304906.

Gari:2014:TAS

[GNIL14] Sirak R. Gari, Alice Newton, John Icely, and Christopher D. Lowe. Testing the application of the Systems Approach Framework (SAF) for the management of eutrophication in the Ria Formosa. Marine Policy, 43 (??):40–45, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000936.

Goundar:2017:KCM

[GNN^{+}17] Avnita Goundar, Alison Newell, Peter Nuttall, Isabelle Rojon, and Jale Samuwai. King Canute muses in the South Seas: Why aren’t Pacific Islands transitioning to low carbon sea transport futures? Marine Policy, 81(??):80–90, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304900.

Grubesic:2019:SPA

[GNW19] Tony H. Grubesic, Jake R. Nelson, and Ran Wei. A strategic planning approach for protecting environmentally sensitive coastlines from oil spills: Allocating response resources on a limited budget. Marine Policy, 108(??):Article 103549, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830784X.

Grimon:2016:TIG

[GOdFC16] René Oscar Rodríguez Grimón, María Fernanda Arroyo Osorio, Débora M. de Freitas, and Ítalo Braga Castro. Tributyltin impacts in Galapagos Islands and Ecuadorian shore: Marine protected areas under threat. Marine Policy, 69(??):24–31, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630015X.

Garcia-Oliva:2017:EIT

[GOHDB17] Miriam Garcia-Oliva, Tara Hooper, Slobodan Djordjević, and Michael Belmont. Exploring the implications of tidal farms deployment for wetland-birds habitats in a highly protected estuary. Marine Policy, 81(??):359–367, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300647.

Gilman:2014:ERA

[GOK14] Eric Gilman, Matthew Owens, and Thomas Kraft. Ecological risk assessment of the Marshall Islands longline tuna fishery. Marine Policy, 44(??):239–255, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001966.

Goldstein:2013:CFE

[Gol13] Lyle J. Goldstein. Chinese fisheries enforcement: Environmental and strategic implications. Marine Policy, 40(??):187–193, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002485.

Guijarro:2017:IEE

[GOM17] Beatriz Guijarro, Francesc Ordines, and Enric Massutí. Improving the ecological efficiency of the bottom trawl fishery in the Western Mediterranean: It’s about time! Marine Policy, 83(??):204–214, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304092.

Goodman:2017:SRB

[Goo17] Camille Goodman. Striking the right balance? Applying the jurisprudence of international tribunals to coastal state innovations in international fisheries governance. Marine Policy, 84(??):293–299, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730180X.

Gray:2019:CFU

[GOO^{+}19] S. R. J. Gray, C. O’Mahony, B. O’Dwyer, S. A. Gray, and J. Gault. Caught by the fuzz: Using FCM to prevent coastal adaptation stakeholders from fleeing the scene. Marine Policy, 109(??):Article 103688, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303811.

Gorris:2019:MGB

[Gor19] P. Gorris. Mind the gap between aspiration and practice in co-managing marine protected areas: a case study from Negros Occidental, Philippines. Marine Policy, 105(??):12– 19, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304664.

Gill:2019:MMD

[GOTS19] David A. Gill, Hazel A. Oxenford, Rachel A. Turner, and Peter W. Schuhmann. Making the most of data-poor fisheries: Low cost mapping of small island fisheries to inform policy. Marine Policy, 101(??):198–207, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302312.

Gonson:2018:SCC

[GPA18] Charles Gonson, Dominique Pelletier, and Frederique Alban. Social carrying capacity assessment from questionnaire and counts survey: Insights for recreational settings management in coastal areas. Marine Policy, 98 (??):146–157, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303536.

Giglio:2018:LRM

[GPB^{+}18] Vinicius J. Giglio, Hudson T. Pinheiro, Mariana G. Bender, Roberta M. Bonaldo, Letícia V. Costa-Lotufo, Carlos E. L. Ferreira, Sergio R. Floeter, Andrea Freire, João L. Gasparini, Jean-Christophe Joyeux, João Paulo Krajewski, Alberto Lindner, Guilherme O. Longo, Tito M. C. Lotufo, Rafael Loyola, Osmar J. Luiz, Raphael M. Macieira, Rafael A. Magris, Thayná J. Mello, Juan P. Quimbayo, Luiz A. Rocha, Bárbara Segal, João B. Teixeira, Daniele A. Vila-Nova, Ciro C. Vilar, Carla Zilberberg, and Ronaldo B. Francini-Filho. Large and remote marine protected areas in the South Atlantic Ocean are flawed and raise concerns: Comments on Soares and Lucas (2018). Marine Policy, 96(??):13–17, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303373.

Gissi:2018:GOW

[GPH18] E. Gissi, M. E. Portman, and A. K. Hornidge. Ungendering the ocean: Why women matter in ocean governance for sustainability. Marine Policy, 94(??):215–219, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730787X.

Ganapathiraju:2019:EIU

[GPM19] Pramod Ganapathiraju, Tony J. Pitcher, and Gopikrishna Mantha. Estimates of illegal and unreported seafood imports to Japan. Marine Policy, 108(??):Article 103439, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830945X.

Gussow:2010:OIF

[GPO^{+}10] Kerstin Güssow, Alexander Proelss, Andreas Oschlies, Katrin Rehdanz, and Wilfried Rickels. Ocean iron fertilization: Why further research is needed. Marine Policy, 34(5):911–918, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000163.

Gall:2016:ESA

[GR16]  Sarah C. Gall and Lynda D. Rodwell.  Evaluat-

ing the social acceptability of Marine Protected Areas. Marine Policy, 65(??):30–38, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500367X.

Grafton:2010:ACC

[Gra10] R. Quentin Grafton. Adaptation to climate change in marine capture fisheries. Marine Policy, 34(3):606–615, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001845.

Green:2015:FSP

[Gre15] Lesley Green. Fisheries science, Parliament and fishers’ knowledge in South Africa: an attempt at scholarly diplomacy. Marine Policy, 60(??):345–352, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002048.

Gritsenko:2017:RGE

[Gri17] Daria Gritsenko. Regulating GHG Emissions from shipping: Local, global, or polycentric approach? Marine Policy, 84(??):130–133, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X17301215.

Grilly:2015:PFG

[GRLJ15] Emily Grilly, Keith Reid, Sarah Lenel, and Julia Jabour. The price of fish: a global trade analysis of Patagonian (Dissostichus eleginoides) and Antarctic toothfish (Dissostichus mawsoni). Marine Policy, 60(??):186–196, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001657.

Garcia-Rivera:2015:QQA

[GRLM15] Santiago García-Rivera, Jose Luis Sánchez Lizaso, and Jose María Bellido Millán. A quantitative and qualitative assessment of the discard ban in European Mediterranean waters. Marine Policy, 53(??):149–158, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003303.

Gelcich:2018:AIM

[GRMAC18] Stefan Gelcich, Francisca Reyes-Mendy, Rodrigo Arriagada, and Bárbara Castillo. Assessing the implementation of marine ecosystem based management into national policies: Insights from agenda setting and policy responses. Marine Policy, 92(??):40–47, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304037.

Gronholm:2018:TWB

[Grö18] Sam Grönholm. A tangled web: Baltic Sea Region governance through networks. Marine Policy, 98(??): 201–210, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306766.

Goodspeed:2016:ADT

[GRW^{+}16] Robert Goodspeed, Catherine Riseng, Kevin Wehrly, Wenting Yin, Lacey Mason, and Benjamin Schoenfeldt. Applying design thinking methods to ecosystem management tools: Creating the Great Lakes Aquatic Habitat Explorer. Marine Policy, 69(??):134–145, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301890.

Gadamus:2015:BIE

[GRYA^{+}15] Lily Gadamus, Julie Raymond-Yakoubian, Roy Ashenfelter, Austin Ahmasuk, Vera Metcalf, and George Noongwook. Building an indigenous evidence-base for tribally-led habitat conservation policies. Marine Policy, 62(??):116–124, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002596.

Gunnlaugsson:2016:IFI

[GS16] Stefan B. Gunnlaugsson and Hordur Saevaldsson. The Icelandic fishing industry: Its development and financial performance under a uniform individual quota system. Marine Policy, 71(??):73–81, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301725.

Gibson:2017:DDS

[GS17] Darah Gibson and U. Rashid Sumaila. Determining the degree of ’small-scaleness’ using fisheries in British Columbia as an example. Marine Policy, 86(??):121–126, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17304815.

Goyert:2010:PPM

[GSA10] Wendy Goyert, Raphael Sagarin, and John Annala. The promise and pitfalls of Marine Stewardship Council certification: Maine lobster as a case study. Marine Policy, 34(5):1103–1109, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000062X.

Garmendia:2017:AMP

[GSB^{+}17] M. Garmendia, D. Sauzade, N. Beaumont, B. Boteler, M. Pascual, T. Boudine, M. Breil, E. Furlan, A. Kontogianni, I. Krüger, J. Le Tellier, E. Gileva, D. March, G. Roeleveld, P. Ronco, A. Shivarov, M. Skourtos, and A. Markandya. The Adaptive Marine Policy (AMP) toolbox: Supporting policy-makers developing adaptive policies in the Mediterranean and Black Sea. Marine Policy, 84(??):99–109, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304505.

Guillen:2016:SNL

[GSC^{+}16] Jordi Guillen, Angel Calvo Santos, Griffin Carpenter, Natacha Carvalho, John Casey, Jordi Lleonart, Francesc Maynou, Gorka Merino, and Anton Paulrud. Sustainability now or later? Estimating the benefits of pathways to maximum sustainable yield for EU Northeast Atlantic fisheries. Marine Policy, 72(??):40–47, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630149X.

Guimaraes:2014:EBM

[GSDB14] Maria Helena Guimarães, Carlos Sousa, Tomaz Dentinho, and Tomasz Boski. Economic base model for the Guadiana estuary, Portugal an application for Integrated Coastal Zone Management. Marine Policy, 43(??):63–70, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000869.

Gimpel:2015:GMF

[GSG^{+}15] Antje Gimpel, Vanessa Stelzenmüller, Britta Grote, Bela H. Buck, Jens Floeter, Ismael Núñez-Riboni, Bernadette Pogoda, and Axel Temming. A GIS modelling framework to evaluate marine spatial planning scenarios: Co-location of offshore wind farms and aquaculture in the German EEZ. Marine Policy, 55(??):102–115, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000238.

Gehrig:2018:OIR

[GSJ18] Stefan Gehrig, Achim Schlüter, and Narriman S. Jiddawi. Overlapping identities: the role of village and occupational group for small-scale fishers’ perceptions on environment and governance. Marine Policy, 96(??):100–110, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308837.

Garcia:2019:RMS

[GSR19] Pablo Quero García, Javier García Sanabria, and Juan Adolfo Chica Ruiz. The role of maritime spatial planning on the advance of blue energy in the European Union. Marine Policy, 99(??):123–131, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830304X.

Greening:2018:LIN

[GSS^{+}18] Holly Greening, Roberta Swann, Kerry St. Pé, Susan Testroet-Bergeron, Ray Allen, Mark Alderson, Jennifer Hecker, and Sarah P. Bernhardt. Local implementation of a national program: the National Estuary Program response following the Deepwater Horizon oil spill in the Gulf of Mexico. Marine Policy, 87(??):60–64, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304724.

Gerovasileiou:2019:HME

[GSS^{+}19] Vasilis Gerovasileiou, Christopher J. Smith, Katerina Sevastou, Nadia Papadopoulou, Thanos Dailianis, Trine Bekkby, Dario Fiorentino, Chris J. McOwen, Teresa Amaro, Elizabeth Grace Tunka Bengil, Meri Bilan, Christoffer Boström, Marina Carreiro-Silva, Emma Cebrian, Carlo Cerrano, Roberto Danovaro, Simonetta Fraschetti, Karine Gagnon, Cristina Gambi, Anthony Grehan, Bernat Hereu, Silvija Kipson, Jonne Kotta, Cristina Linares, Telmo Morato, Henn Ojaveer, Helen Orav-Kotta, Antonio Sarà, and Rachael Scrimgeour. Habitat mapping in the European Seas — is it fit for purpose in the marine restoration agenda? Marine Policy, 106(??):Article 103521, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307061.

Grimvall:2018:SEA

[GSW18] Anders Grimvall, Eva-Lotta Sundblad, and Are Wallin. Systematic exploration of actors in society who influence the input of nutrients into the sea. Marine Policy, 96 (??):65–71, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307266.

Giglio:2018:HCP

[GTL^{+}18] Vinicius J. Giglio, Maria L. F. Ternes, Osmar J. Luiz, Cleverson Zapelini, and Matheus O. Freitas. Human consumption and popular knowledge on the conservation status of groupers and sharks caught by smallscale fisheries on Abrolhos Bank, SW Atlantic. Marine Policy, 89(??):142–146, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303627.

Glenn:2012:TFS

[GTM^{+}12] Helen Glenn, Diana Tingley, Sonia Sánchez Maroño, Dennis Holm, Laurence Kell, Gurpreet Padda, Ingi Runar Edvardsson, Johann Asmundsson, Alexis Conides, Kostas Kapiris, Mintewab Bezabih, Premachandra Wattage, and Sakari Kuikka. Trust in the fisheries scientific community. Marine Policy, 36(1):54–72, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000674.

Guerra:2018:MOR

[Gue18] Flávia Guerra. Mapping offshore renewable energy governance. Marine Policy, 89(??):21–33, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304323.

Guerra:2019:WSM

[Gue19] Ana Sofia Guerra. Wolves of the sea: Managing human-wildlife conflict in an increasingly tense ocean. Marine Policy, 99(??):369–373, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303208.

Guggisberg:2019:FCM

[Gug19] Solène Guggisberg. Funding coastal and marine fisheries projects under the climate change regime. Marine Policy, 107(??):Article 103352, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304962.

Gustafsson:2018:BSR

[Gus18a] Åsa Gustafsson. The Baltic Sea Region Border Control Cooperation (BSRBCC) and border management in the Baltic Sea region: a case study. Marine Policy, 98(??): 309–316, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18307000.

Gustavsson:2018:ECS

[Gus18b] Madeleine Gustavsson. Examining the ‘cultural sustainability’ of two different ways of governing fishing practices. Marine Policy, 97(??):262–269, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301593.

Gagern:2013:CRF

[GvdB13] Antonius Gagern and Jeroen van den Bergh. A critical review of fishing agreements with tropical developing countries. Marine Policy, 38(??):375–386, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001558.

Gambelli:2019:EPO

[GVSZ19] Danilo Gambelli, Daniela Vairo, Francesco Solfanelli, and Raffaele Zanoli. Economic performance of organic aquaculture: a systematic review. Marine Policy, 108(??): Article 103542, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306468.

Godin:2010:KLH

[GW10] Aurelie Cosandey Godin and Boris Worm. Keeping the lead: How to strengthen shark conservation and management policies in Canada. Marine Policy, 34(5):995– 1001, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000497.

Griffin:2011:DPE

[GW11] Wade L. Griffin and Richard T. Woodward. Determining policy-efficient management strategies in fisheries using data envelopment analysis (DEA). Marine Policy, 35(4):496–507, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002162.

Gibbs:2015:TSH

[GW15] Leah Gibbs and Andrew Warren. Transforming shark hazard policy: Learning from ocean-users and shark encounter in Western Australia. Marine Policy, 58(??):116–124, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000962.

Glenn:2010:MPA

[GWM^{+}10] H. Glenn, P. Wattage, S. Mardle, T. Van Rensburg, A. Grehan, and N. Foley. Marine protected areas — substantiating their worth. Marine Policy, 34(3):421–430, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001420.

Gilbert:2017:RME

[GWWH17] P. Gilbert, P. Wilson, C. Walsh, and P. Hodgson. The role of material efficiency to reduce CO _{2} emissions during ship manufacture: a life cycle approach. Marine Policy, 75(??):227–237, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630077X.

Greer:2019:GTC

[GZW^{+}19a] Krista Greer, Dirk Zeller, Jessika Woroniak, Angie Coulter, Maeve Winchester, M. L. Deng Palomares, and Daniel Pauly. Global trends in carbon dioxide (CO_{2}) emissions from fuel combustion in marine fisheries from 1950 to 2016. Marine Policy, 107(??):Article 103382, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730893X.

Greer:2019:RZA

[GZW^{+}19b] Krista Greer, Dirk Zeller, Jessika Woroniak, Angie Coulter, Maeve Winchester, M. L. Deng Palomares, and Daniel Pauly. Reply to Ziegler et al. “Adding perspectives to: Global trends in carbon dioxide (CO

2) emissions from fuel combustion in marine fisheries from 1950–2016” and addressing concerns of using fishing effort to predict carbon dioxide emissions. Marine Policy, 107(??):Article 103491, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302040.

Heupel:2013:ELS

[HA13] Eric Heupel and Peter J. Auster. Eco-labeling seafood: Addressing impacts to vulnerable seafloor species, communities, habitats and ecosystems in data-poor regions. Marine Policy, 38(??):8–15, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001029.

Hooper:2014:CLO

[HA14] Tara Hooper and Melanie Austen. The co-location of offshore windfarms and decapod fisheries in the UK: Constraints and opportunities. Marine Policy, 43(??): 295–300, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001371.

Hooper:2015:PFD

[HAA15] Tara Hooper, Matthew Ashley, and Melanie Austen. Perceptions of fishers and developers on the co-location of offshore wind farms and decapod fisheries in the UK. Marine Policy, 61(??):16–22, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001967.

Henriquez-Antipa:2019:SMP

[HAC19] Luis A. Henríquez-Antipa and F. Cárcamo. Stakeholder’s multidimensional perceptions on policy implementation gaps regarding the current status of Chilean small-scale seaweed aquaculture. Marine Policy, 103(??):138–147, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301969.

Hadjimichael:2018:CBD

[Had18] Maria Hadjimichael. A call for a blue degrowth: Unravelling the European Union’s fisheries and maritime policies. Marine Policy, 94(??):158–164, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305687.

Humber:2015:EEE

[HAGB15] Frances Humber, Mialy Andriamahefazafy, Brendan John Godley, and Annette Cameron Broderick. Endangered, essential and exploited: How extant laws are not enough to protect marine megafauna in Madagascar. Marine Policy, 60(??):70–83, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001396.

Halouani:2016:STE

[HAH^{+}16] Ghassen Halouani, Khaled Abdou, Tarek Hattab, Mohamed Salah Romdhane, Frida Ben Rais Lasram, and François Le Loc’h. A spatio-temporal ecosystem model to simulate fishing management plans: a case of study in the Gulf of Gabes (Tunisia). Marine Policy, 69(??):62–72, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300677.

Hilborn:2019:PAM

[HAK^{+}19] Ray Hilborn, Christopher M. Anderson, Gordon H. Kruse, André E. Punt, Michael Sissenwine, Chris Oliver, James N. Ianelli, Robert J. Trumble, David J. Agnew, and Nicole Baker. Pramod et al. methods to estimate IUU are not credible. Marine Policy, 108(??):Article 103632, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303318.

Hannesson:2011:RBF

[Han11] Rögnvaldur Hannesson. Rights based fishing on the high seas: Is it possible? Marine Policy, 35(5):667–674, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000315.

Hanich:2012:DBC

[Han12] Quentin Hanich. Distributing the bigeye conservation burden in the western and central Pacific fisheries. Marine Policy, 36(2):327–332, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001230.

Hannesson:2013:NEI

[Han13a] Rögnvaldur Hannesson. Norway’s experience with ITQs. Marine Policy, 37(??):264–269, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000966.

Hannesson:2013:SBE

[Han13b] Rögnvaldur Hannesson. Strictly for the birds?: On ecosystem services of forage fish. Marine Policy, 38(??):109–115, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001248.

Hannesson:2014:DTM

[Han14a] Rögnvaldur Hannesson. Does threat of mutually assured destruction produce quasi-cooperation in the mackerel fishery? Marine Policy, 44(??):342–350, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002236.

Hannesson:2014:NEI

[Han14b] Rögnvaldur Hannesson. Norway’s experience with ITQs: a rejoinder. Marine Policy, 44(??):473–474, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300167X.

Hansen:2016:TMC

[Han16] Harald Sakarias Brøvig Hansen. Three major challenges in managing non-native sedentary Barents Sea snow crab (Chionoecetes opilio). Marine Policy, 71(??): 38–43, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302950.

Hanich:2019:ENP

[Han19] Quentin Hanich. Editorial note on publication of Malin et al.: In oceans we trust: conservation, philanthropy, and the political economy of the Phoenix Islands Protected Area. Marine Policy, 107(??):Article 103523, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302775.

Hartley:2010:FMG

[Har10] Troy W. Hartley. Fishery management as a governance network: Examples from the Gulf of Maine and the potential for communication network analysis research in fisheries. Marine Policy, 34(5):1060–1067, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000576.

Harris:2014:FSH

[Har14] Peter Harris. Fortress, safe haven or home? The Chagos MPA in political context. Marine Policy, 46(??):19–21, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000049.

Hart:2015:CSI

[Har15] Anthony M. Hart. Commercial scale invertebrate fisheries enhancement in Australia: Experiences, challenges and opportunities. Marine Policy, 62(??):82–93, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002602.

Harvey:2019:PPP

[Har19] Nick Harvey. Protecting private properties from the sea: Australian policies and practice. Marine Policy, 107(??):Article 103566, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830753X.

Hassanali:2018:AIA

[Has18] Kahlil Hassanali. Approaching the implementing agreement to UNCLOS on biodiversity in ABNJ: Exploring favorable outcomes for CARICOM. Marine Policy, 98 (??):92–96, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304743.

Havice:2010:STA

[Hav10] Elizabeth Havice. The structure of tuna access agreements in the Western and Central Pacific Ocean: Lessons for Vessel Day Scheme planning. Marine Policy, 34(5): 979–987, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000461.

Havice:2013:RBM

[Hav13] Elizabeth Havice. Rights-based management in the Western and Central Pacific Ocean tuna fishery: Economic and environmental change under the Vessel Day Scheme. Marine Policy, 42(??):259–267, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000717.

Hossain:2019:SIC

[HAW19] Tahazzud Hossain, Michelle Adams, and Tony R. Walker. Sustainability initiatives in Canadian ports. Marine Policy, 106(??):Article 103519, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307413.

Hospital:2014:CSM

[HB14] Justin Hospital and Courtney Beavers. Catch shares and the main Hawaiian Islands bottomfish fishery: Linking fishery conditions and fisher perceptions. Marine Policy, 44 (??):9–17, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001607.

Holzer:2019:DHO

[HB19] Jorge Holzer and Kristen Byler. Deepwater Horizon Oceanic Fish Restoration Project: an application of market mechanisms for conservation. Marine Policy, 108(??):Article 103538, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308273.

Hauge:2014:IRA

[HBA^{+}14] Kjellrun Hiis Hauge, Anne Blanchard, Gisle Andersen, Ragnhild Boland, Bjørn Einar Grøsvik, Daniel Howell, Sonnich Meier, Erik Olsen, and Frode Vikebø. Inadequate risk assessments — a study on worst-case scenarios related to petroleum exploitation in the Lofoten area. Marine Policy, 44(??):82–89, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001504.

Hughes:2013:CPU

[HBC^{+}13] Adam D Hughes, Kenneth D Black, Iona Campbell, Johanna J Heymans, Kyla K Orr, Michele S. Stanley, and Maeve S. Kelly. Comments on ‘Prospects for the use of macroalgae for fuel in Ireland and UK: an overview of marine management issues’. Marine Policy, 38(??):554–556, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001686.

Halldorsson:2012:FRC

[HBG12] Jón E. Halldórsson, Björn Björnsson, and Stefán B. Gunnlaugsson. Feasibility of ranching coastal cod (Gadus morhua) compared with on-growing, full-cycle farming and fishing. Marine Policy, 36(1):11–17, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000510.

Huntington:2019:RAA

[HBH^{+}19] Henry P. Huntington, Sarah Bobbe, Andrew Hartsig, Erika J. Knight, Alexey Knizhnikov, Alexander Moiseev, Olga Romanenko, Melanie A. Smith, and Benjamin K. Sullender. The role of areas to be avoided in the governance of shipping in the greater Bering Strait region. Marine Policy, 110(??):Article 103564, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307942.

Haugen:2017:WEF

[HBK17] Arne Sveinson Haugen, Scott Bremer, and Matthias Kaiser. Weaknesses in the ethical framework of aquaculture related standards. Marine Policy, 75(??):11–18, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306698.

Hadjimichael:2014:TFP

[HBL14] Maria Hadjimichael, Adriana Bruggeman, and Manfred A. Lange. Tragedy of the few? A political ecology perspective of the right to the sea: the Cyprus marine aquaculture sector. Marine Policy, 49(??):12–19, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001146.

Hopkins:2016:SMP

[HBP16] Charlotte Rachael Hopkins, David Mark Bailey, and Tavis Potts. Scotland’s Marine Protected Area network: Reviewing progress towards achieving commitments for marine conservation. Marine Policy, 71(??):44–53, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303001.

Hills:2019:DBR

[HBS^{+}19] Jeremy Hills, Shirleen Bala, Anama Solofa, Piers Dunstan, Mibu Fischer, and Donna Hayes. The disjuncture between regional ocean priorities and development assistance in the South Pacific. Marine Policy, 107(??):Article 103420, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304305.

Hoagland:2010:DSM

[HBT^{+}10] Porter Hoagland, Stace Beaulieu, Maurice A. Tivey, Roderick G. Eggert, Christopher German, Lyle Glowka, and Jian Lin. Deep-sea mining of seafloor massive sulfides. Marine Policy, 34(3):728–732, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001870.

Hossain:2016:MHH

[HBZH16] Mohammad Shawkat Hossain, Japar Sidik Bujang, Muta Harah Zakaria, and Mazlan Hashim. Marine and human habitat mapping for the Coral Triangle Initiative region of Sabah using Landsat and Google Earth imagery. Marine Policy, 72(??):176–191, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300884.

Huang:2010:FCM

[HC10] Hsiang-Wen Huang and Ching-Ta Chuang. Fishing capacity management in Taiwan: Experiences and prospects. Marine Policy, 34(1):70–76, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000669.

Hilborn:2018:PBG

[HC18] Ray Hilborn and Chris Costello. The potential for blue growth in marine fish yield, profit and abundance of fish in the ocean. Marine Policy, 87(??):350–355, January 2018.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730074X.

Harvey:2019:CRA

[HC19] Nick Harvey and Beverley Clarke. 21st Century reform in Australian coastal policy and legislation. Marine Policy, 103(??):27–32, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18308327.

Hanich:2015:RFE

[HCBM15] Quentin Hanich, Brooke Campbell, Megan Bailey, and Erik Molenaar. Research into fisheries equity and fairness — addressing conservation burden concerns in transboundary fisheries. Marine Policy, 51(??):302–304, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002449.

Holden:2019:SCC

[HCC^{+}19] Jessica J. Holden, Brenna Collicutt, Garth Covernton, Kieran D. Cox, Darienne Lancaster, Sarah E. Dudas, Natalie C. Ban, and Aerin L. Jacob. Synergies on the coast: Challenges facing shellfish aquaculture development on the central and north coast of British Columbia. Marine Policy, 101(??):108–117, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302379.

HortaeCosta:2016:RBC

[HCF^{+}16] Bárbara Horta e Costa, Joachim Claudet, Gustavo Franco, Karim Erzini, Anthony Caro, and Emanuel J. Gonçalves. A regulation-based classification system for Marine Protected Areas (MPAs). Marine Policy, 72(??):192–198, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300197.

HortaeCosta:2017:RBC

[HCF^{+}17] Barbara Horta e Costa, Joachim Claudet, Gustavo Franco, Karim Erzini, Anthony Caro, and Emanuel J.

Gonçalves. A regulation-based classification system for marine protected areas: a response to Dudley et al. [9]. Marine Policy, 77(??):193–195, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307485.

Huang:2015:REE

[HCJ15] Wei Huang, James J. Corbett, and Di Jin. Regional economic and environmental analysis as a decision support for marine spatial planning in Xiamen. Marine Policy, 51(??):555–562, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002395.

Hesp:2017:IFE

[HCP^{+}17] Sybrand Hesp, Nick Caputi, Jim Penn, Mervi Kangas, Errol Sporer, Brett Hogan, and James Clement. Improving fleet efficiency to maximise economic yield in a Western Australian prawn fishery. Marine Policy, 86(??):82–93, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307230.

Hermansen:2010:CSS

[HD10a] Øystein Hermansen and Bent Dreyer. Challenging spatial and seasonal distribution of fish landings — The experiences from rural community quotas in Norway. Marine Policy, 34(3):567–574, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001766.

Human:2010:SCD

[HD10b] Brett A. Human and Amanda Davies. Stakeholder consultation during the planning phase of scientific programs. Marine Policy, 34(3):645–654, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001894.

Hinch:2011:EFC

[HD11a] Patricia R. Hinch and Elizabeth M. De Santo. Erratum to “Factors to consider in evaluating the management and conservation effectiveness of a whale sanctuary to protect and conserve the North Atlantic right whale (Eubalaena glacialis)”. [Mar. Policy 35(2) (2010) 163–180]. Marine Policy, 35(4):557, July 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000145.

Hinch:2011:FCE

[HD11b] Patricia R. Hinch and Elizabeth M. De Santo. Factors to consider in evaluating the management and conservation effectiveness of a whale sanctuary to protect and conserve the North Atlantic right whale (Eubalaena glacialis). Marine Policy, 35(2):163–180, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001557.

Harden-Davies:2016:MST

[HD16a] Harriet Harden-Davies. Marine science and technology transfer: Can the Intergovernmental Oceanographic Commission advance governance of biodiversity beyond national jurisdiction? Marine Policy, 74(??):260–267, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303955.

Hermida:2016:HTL

[HD16b] Margarida Hermida and João Delgado. High trophic level and low diversity: Would Madeira benefit from fishing down? Marine Policy, 73(??):130–137, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302056.

HortaeCosta:2019:COC

[HdF^{+}19] Bárbara Horta e Costa, Jorge Manuel dos Santos Gonçalves, Gustavo Franco, Karim Erzini, Ricardo Furtado, Célia Mateus, Emília Cadeireiro, and Emanuel João Gonçalves. Categorizing ocean conservation targets to avoid a potential false sense of protection to society: Portugal as a case-study. Marine Policy, 108(??):Article 103553, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830616X.

Halpern:2012:NTP

[HDG^{+}12] Benjamin S. Halpern, Jordan Diamond, Steve Gaines, Stefan Gelcich, Mary Gleason, Simon Jennings, Sarah Lester, Amber Mace, Laurence McCook, Karen McLeod, Nicholas Napoli, Kit Rawson, Jake Rice, Andrew Rosenberg, Mary Ruckelshaus, Bettina Saier, Paul Sandifer, Astrid Scholz, and Anna Zivian. Near-term priorities for the science, policy and practice of Coastal and Marine Spatial Planning (CMSP). Marine Policy, 36(1):198–205, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000935.

Huntington:2015:VRR

[HDH^{+}15] Henry P. Huntington, Raychelle Daniel, Andrew Hartsig, Kevin Harun, Marilyn Heiman, Rosa Meehan, George Noongwook, Leslie Pearson, Melissa Prior-Parks, Martin Robards, and George Stetson. Vessels, risks, and rules: Planning for safe shipping in Bering Strait. Marine Policy, 51(??):119–127, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002012.

Hoagland:2015:AAS

[HDJD15] P. Hoagland, T. M. Dalton, D. Jin, and J. B. Dwyer. An approach for analyzing the spatial welfare and distributional effects of ocean wind power siting: the Rhode Island/Massachusetts area of mutual interest. Marine Policy, 58(??):51–59, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000925.

Harkes:2015:SAV

[HDK^{+}15] I. H. T. Harkes, A. Drengstig, M. P. Kumara, J. M. P. K. Jayasinghe, and M. Huxham. Shrimp aquaculture as a vehicle for Climate Compatible Development in Sri Lanka. The case of Puttalam Lagoon. Marine Policy, 61(??):273–283, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002225.

Haward:2013:CCS

[HDL^{+}13] Marcus Haward, Julie Davidson, Michael Lockwood, Marc Hockings, Lorne Kriwoken, and Robyn Allchin. Climate change, scenarios and marine biodiversity conservation. Marine Policy, 38(??):438–446, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001649.

Hauzer:2013:ECB

[HDM13] Melissa Hauzer, Philip Dearden, and Grant Murray. The effectiveness of community-based governance of small-scale fisheries, Ngazidja island, Comoros. Marine Policy, 38 (??):346–354, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001510.

Hull:2014:ICS

[HDTS14] Steve Hull, Ian Dickie, Rob Tinch, and Justine Saunders. Issues and challenges in spatio-temporal application of an ecosystem services framework to UK seas. Marine Policy, 45(??):359–367, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300208X.

Hoefnagel:2017:SEC

[HdV17] Ellen Hoefnagel and Birgit de Vos. Social and economic consequences of 40 years of Dutch quota management. Marine Policy, 80(??):81–87, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305838.

Hoefnagel:2013:MIG

[HdVB13] Ellen Hoefnagel, Birgit de Vos, and Erik Buisman. Marine informational governance, a conceptual framework. Marine Policy, 42(??):150–156, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000493.

Hoefnagel:2015:QSR

[HdVB15] Ellen Hoefnagel, Birgit de Vos, and Erik Buisman. Quota swapping, relative stability, and transparency. Marine Policy, 57(??):111–119, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000615.

He:2015:CPP

[He15] Juan He. Chinese public policy on fisheries subsidies: Reconciling trade, environmental and food security stakes. Marine Policy, 56(??):106–116, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000020.

He:2018:COL

[He18] Juan He. From country-of-origin labelling (COOL) to seafood import monitoring program (SIMP): How far can seafood traceability rules go? Marine Policy, 96(??): 163–174, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300745.

Heazle:2013:SYC

[Hea13] Michael Heazle. “See you in court!”: Whaling as a two level game in Australian politics and foreign policy. Marine Policy, 38(??):330–336, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001492.

Heen:2014:CAR

[Hee14] Knut Heen. Comment to the article by R. Hannesson “Norway’s experience with ITQs”. Marine Policy, 44 (??):475–477, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001668.

Hadjimichael:2010:DBF

[HEJK10] Maria Hadjimichael, Gareth Edwards-Jones, and Michel J. Kaiser. Distribution of the burden of fisheries regulations in Europe: the north/south divide. Marine Policy, 34(4):795–802, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000229.

Hermanson:2018:ESM

[Her18a] Ann-Sofie Hermanson. Energy security in a multi-level governance perspective. Marine Policy, 98(??):301–308, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306997.

Hersoug:2018:AAY

[Her18b] Bjørn Hersoug. “After all these years” — New Zealand’s quota management system at the crossroads. Marine Policy, 92(??):101–110, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307698.

Haas:2016:CCP

[HES16] Andrea R. Haas, Danielle N. Edwards, and U. Rashid Sumaila. Corporate concentration and processor control: Insights from the salmon and herring fisheries in British Columbia. Marine Policy, 68(??):83–90, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000713.

Heylen:2017:RPM

[Hey17] Brigitte Heylen. RETRACTED: Protecting marine top predators through adaptive and dynamic management. Marine Policy, 86(??):190–197, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306522.

Heylen:2018:RNP

[Hey18] Brigitte Heylen. Retraction notice to “Protecting marine top predators through adaptive and dynamic management” [Marine Policy 86 (2017) 190–197]. Marine Policy, 92(??):148, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302100.

Hynes:2012:DSS

[HF12] Stephen Hynes and Niall Farrelly. Defining standard statistical coastal regions for Ireland. Marine Policy, 36 (2):393–404, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001199.

Habtemariam:2016:ZMU

[HF16] Bereket Tesfamariam Habtemariam and Qinhua Fang. Zoning for a multiple-use marine protected area using spatial multi-criteria analysis: the case of the Sheik Seid Marine National Park in Eritrea. Marine Policy, 63(??): 135–143, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1500295X.

Hosch:2011:FCC

[HFF11] Gilles Hosch, Gianluca Ferraro, and Pierre Failler. The 1995 FAO Code of Conduct for Responsible Fisheries: Adopting, implementing or scoring results? Marine Policy, 35(2):189–200, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001661.

Hawkins:2013:DRR

[HFM^{+}13] S. J. Hawkins, L. B. Firth, M. McHugh, E. S. Poloczanska, R. J. H. Herbert, M. T. Burrows, M. A. Kendall, P. J. Moore, R. C. Thompson, S. R. Jenkins, D. W. Sims, M. J. Genner, and N. Mieszkowska. Data rescue and re-use: Recycling old information to address new policy concerns. Marine Policy, 42(??):91–98, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000420.

Hatchard:2014:RAC

[HG14] Jenny L. Hatchard and Tim S. Gray. From RACs to Advisory Councils: Lessons from North Sea discourse for the 2014 reform of the European Common Fisheries Policy. Marine Policy, 47(??):87–93, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000530.

Heyman:2019:CMA

[HGB^{+}19] William D. Heyman, Arnaud Grüss, Christopher R. Biggs, Shin’ichi Kobara, Nicholas A. Farmer, Mandy Karnauskas, Sue Lowerre-Barbieri, and Brad Erisman. Cooperative monitoring, assessment, and management of fish spawning aggregations and associated fisheries in the U.S. Gulf of Mexico. Marine Policy, 109(??):Article 103689, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302179.

Hynes:2016:DCM

[HGBJ16] Stephen Hynes, Hans Gerritsen, Benjamin Breen, and Mark Johnson. Discrete choice modelling of fisheries with nuanced spatial information. Marine Policy, 72 (??):156–165, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300070.

Herndon:2010:CIC

[HGDB10] Andrew Herndon, Vincent F. Gallucci, Douglas DeMaster, and William Burke. The case for an International Commission for the Conservation and Management of Sharks (IC-CMS). Marine Policy, 34(6):1239–1248, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000850.

Hassler:2018:CAA

[HGG^{+}18] Björn Hassler, Kira Gee, Michael Gilek, Anne Luttmann, Andrea Morf, Fred Saunders, Igne Stalmokaite, Helena Strand, and Jacek Zaucha. Collective action and agency in Baltic Sea marine spatial planning: Transnational policy coordination in the promotion of regional coherence. Marine Policy, 92(??):138–147, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308229.

Haver:2018:MLT

[HGH^{+}18] Samara M. Haver, Jason Gedamke, Leila T. Hatch, Robert P. Dziak, Sofie Van Parijs, Megan F. McKenna, Jay Barlow, Catherine Berchok, Eva DiDonato, Brad Hanson, Joseph Haxel, Marla Holt, Danielle Lipski, Haru Matsumoto, Christian Meinig, David K. Mellinger, Sue E. Moore, Erin M. Oleson, Melissa S. Soldevilla, and Holger Klinck. Monitoring long-term soundscape trends in U.S. Waters: the NOAA/NPS Ocean Noise Reference Station Network. Marine Policy, 90(??):6–13, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306309.

Hastings:2012:SBC

[HGS12] Jesse G. Hastings, Rebecca L. Gruby, and Leila S. Sievanen. Science-based coastal management in Fiji: Two case studies from the NGO sector. Marine Policy, 36(4):907–914, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000036.

Hart:2013:CVN

[HGS13] Katharine A. Hart, Tim Gray, and Selina M. Stead. Consumptive versus non-consumptive use of sea turtles? Stakeholder perceptions about sustainable use in three communities near Cahuita National Park, Costa Rica. Marine Policy, 42(??):236–244, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000845.

Hannesson:2010:VPS

[HH10] Rögnvaldur Hannesson and Samuel F. Herrick. The value of Pacific sardine as forage fish. Marine Policy, 34(5): 935–942, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000254.

Hooper:2017:RUO

[HHA17] Tara Hooper, Caroline Hattam, and Melanie Austen. Recreational use of offshore wind farms: Experiences and opinions of sea anglers in the UK. Marine Policy, 78(??):55–60, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307618.

Hind:2010:CBC

[HHG10] E. J. Hind, M. C. Hiponia, and T. S. Gray. From community-based to centralised national management — a wrong turning for the governance of the marine protected area in Apo Island, Philippines? Marine Policy, 34(1):54–62, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0900061X.

Heen:2014:CGF

[HHL14] Eirik Eriksen Heen, Knut Heen, and PingSun Leung. Conflicting goals in fisheries management — a study of the Norwegian cod fisheries. Marine Policy, 49(??):73–80, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000864.

Haapasaari:2015:PAM

[HHL^{+}15] Päivi Haapasaari, Inari Helle, Annukka Lehikoinen, Jouni Lappalainen, and Sakari Kuikka. A proactive approach for maritime safety policy making for the Gulf of Finland: Seeking best practices. Marine Policy, 60(??):107–118, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001621.

Hagspiel:2018:ROU

[HHL^{+}18] Verena Hagspiel, Jørgen Hannevik, Maria Lavrutich, Magnus Naustdal, and Henrik Struksnæs. Real options under technological uncertainty: a case study of investment in a post-smolt facility in Norway. Marine Policy, 88(??):158–166, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304098.

Hattam:2017:WBF

[HHP17] Caroline Hattam, Tara Hooper, and Eleni Papathanasopoulou. A well-being framework for impact evaluation: the case of the UK offshore wind industry. Marine Policy, 78(??):122–131, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303621.

Hamilton-Hart:2016:UEF

[HHS16] Natasha Hamilton-Hart and Christina Stringer. Upgrading and exploitation in the fishing industry: Contributions of value chain analysis. Marine Policy, 63(??):166–171, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1500069X.

Holmes:2019:EBF

[HHS19] Glen Holmes, Quentin Hanich, and Mark Soboil. Economic benefits of FAD set limits throughout the supply chain. Marine Policy, 103(??):1–8, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308194.

Haque:2015:SSF

[HIBG15] C. Emdad Haque, C. Julián Idrobo, Fikret Berkes, and Dale Giesbrecht. Small-scale fishers’ adaptations to change: the role of formal and informal credit in Paraty, Brazil. Marine Policy, 51(??):401–407, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002644.

Hermansen:2012:CSS

[HID12] Øystein Hermansen, John R. Isaksen, and Bent Dreyer. Challenging spatial and seasonal distribution of fish landings — Experiences from vertically integrated trawlers and delivery obligations in Norway. Marine Policy, 36 (1):206–213, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000947.

Hilborn:2010:PGY

[Hil10] Ray Hilborn. Pretty Good Yield and exploited fishes. Marine Policy, 34(1):193–196, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000682.

Hashim:2017:UFK

[HIN^{+}17] Mazlan Hashim, Shuhei Ito, Shinya Numata, Tetsuro Hosaka, Mohamad S. Hossain, Syarifuddin Misbari, Nurul N. Yahya, and Samsudin Ahmad. Using fisher knowledge, mapping population, habitat suitability and risk for the conservation of dugongs in Johor Straits of Malaysia. Marine Policy, 78(??):18–25, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304602.

Hendriksen:2014:FOS

[HJKR14] Astrid Hendriksen, Charlène Jouanneau, Rebecca Koss, and Jesper Raakjaer. Fishing for opinions: Stakeholder views on MSFD implementation in European Seas. Marine Policy, 50(??):353–363, December 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000815.

Hedeholm:2016:LAC

[HJN16] Rasmus Berg Hedeholm, Rikke Becker Jacobsen, and Einar Eg Nielsen. Learning from ‘apparent consensus’ in TAC disputes: Exploring knowledge overlaps in LEK and genetic categorization of Atlantic cod. Marine Policy, 69(??):114–120, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301920.

Haapasaari:2010:FEK

[HK10] Päivi Haapasaari and Timo P. Karjalainen. Formalizing expert knowledge to compare alternative management plans:

Sociological perspective to the future management of Baltic salmon stocks. Marine Policy, 34(3):477–486, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001626.

Heitmann:2011:ACD

[HK11] Nadine Heitmann and Setareh Khalilian. Accounting for carbon dioxide emissions from international shipping: Burden sharing under different UNFCCC allocation options and regime scenarios. Marine Policy, 35(5): 682–691, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000443.

Hsueh:2018:ICS

[HK18] Lily Hsueh and Stephen Kasperski. The impact of catch shares on multiregional fishery participation and effort: the case of west coast harvesters in the Alaska fisheries. Marine Policy, 95(??):123–132, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308473.

Harrould-Kolieb:2019:GFI

[HKHG19] Ellycia R. Harrould-Kolieb and Ove Hoegh-Guldberg. A governing framework for international ocean acidification policy. Marine Policy, 102(??):10–20, April 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305323.

Howard:2019:OCS

[HKR19] Sunkita Howard, Viktoria Kahui, and Jenny Rock. The opportunity cost of spiny dogfish (Squalus acanthias) bycatch in a ling (Genypterus blacodes) longline fishery. Marine Policy, 105(??):38–43, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307127.

Holmes:2010:GEM

[HL10] John Holmes and John Lock. Generating the evidence for marine fisheries policy and management. Marine Policy, 34(1):29–35, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000554.

Hall:2015:DWL

[HL15a] Damon M. Hall and Eli D. Lazarus. Deep waters: Lessons from community meetings about offshore wind resource development in the U.S. Marine Policy, 57(??):9–17, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000536.

Hong:2015:TMC

[HL15b] Gi Hoon Hong and Young Joo Lee. Transitional measures to combine two global ocean dumping treaties into a single treaty. Marine Policy, 55(??):47–56, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000184.

Helson:2010:PRP

[HLC^{+}10] Jeremy Helson, Stefan Leslie, George Clement, Richard Wells, and Ray Wood. Private rights, public benefits: Industry-driven seabed protection. Marine Policy, 34 (3):557–566, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001754.

Hunter:2018:MTA

[HLL^{+}18] Chelsea E. Hunter, Matthew Lauer, Arielle Levine, Sally Holbrook, and Andrew Rassweiler. Maneuvering towards adaptive co-management in a coral reef fishery. Marine Policy, 98(??):77–84, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304585.

Halpern:2013:EPC

[HLM^{+}13] Benjamin S. Halpern, Catherine Longo, Karen L. McLeod, Roger Cooke, Baruch Fischhoff, Jameal F. Samhouri, and Courtney Scarborough. Elicited preferences for components of ocean health in the California Current. Marine Policy, 42(??):68–73, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000286.

Holma:2019:CEB

[HLRO19] Maija Holma, Marko Lindroos, Atso Romakkaniemi, and Soile Oinonen. Comparing economic and biological management objectives in the commercial Baltic salmon fisheries. Marine Policy, 100(??):207–214, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308746.

Hori:2018:SHW

[HM18] Juri Hori and Mitsutaku Makino. The structure of human well-being related to ecosystem services in coastal areas: a comparison among the six North Pacific countries. Marine Policy, 95(??):221–226, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305419.

Hattam:2014:SIT

[HMGR14] C. E. Hattam, S. C. Mangi, S. C. Gall, and L. D. Rodwell. Social impacts of a temperate fisheries closure: understanding stakeholders’ views. Marine Policy, 45(??):269–278, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002029.

Holley:2018:TWR

[HMH18] Jason R. Holley, Katherine A. McComas, and Matthew P. Hare. Troubled waters: Risk perception and the case of oyster restoration in the closed waters of the Hudson–Raritan Estuary. Marine Policy, 91(??):104–112, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302385.

Hersoug:2019:GEA

[HMK19] Bjørn Hersoug, Eirik Mikkelsen, and Kine Mari Karlsen. “Great expectations” — Allocating licenses with special requirements in Norwegian salmon farming. Marine Policy, 100(??):152–162, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304366.

Hedge:2017:IMF

[HMS^{+}17] P. Hedge, F. Molloy, H. Sweatman, K. R. Hayes, J. M. Dambacher, J. Chandler, N. Bax, M. Gooch, K. Anthony, and B. Elliot. An integrated monitoring framework for the Great Barrier Reef World Heritage Area. Marine Policy, 77(??):90–96, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306121.

Hynes:2014:ISA

[HNC14] Stephen Hynes, Danny Norton, and Rebecca Corless. Investigating societal attitudes towards the marine environment of Ireland. Marine Policy, 47(??):57–65, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000359.

Hornborg:2012:IEA

[HNVZ12] Sara Hornborg, Per Nilsson, Daniel Valentinsson, and Friederike Ziegler. Integrated environmental assessment of fisheries management: Swedish, Nephrops trawl fisheries evaluated using a life cycle approach. Marine Policy, 36(6):1193–1201, November 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000346.

Ho:2010:IAS

[Ho10a] Joshua Ho. The implications of Arctic sea ice decline on shipping. Marine Policy, 34(3):713–715, May 2010.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001717.

Huang:2010:PCF

[HO10b] Ming-Ho Huang and Ching-Hsiewn Ou. Precious coral fisheries management in Taiwan — Past, present & future. Marine Policy, 34(5):1002–1009, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000503.

Huang:2012:DMD

[HO12] Ming-Ho Huang and Ching-Hsiewn Ou. A discussion of management disputes arising from the multiple utilization of flying fish resources in Taiwan and suggested countermeasures. Marine Policy, 36(2):512–519, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001485.

Hobbs:2014:GOA

[Hob14] Jean-Paul Adrian Hobbs. A glaring omission in Australia’s marine conservation planning. Marine Policy, 44 (??):149–151, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001747.

Hofman:2019:SOL

[Hof19] Robert J. Hofman. Stopping overexploitation of living resources on the high seas. Marine Policy, 103(??):91–100, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302938.

Holahan:2012:IIG

[Hol12] Robert Holahan. Investigating interest group representation on the Pacific Fisheries Management Council. Marine Policy, 36(3):782–789, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001898.

Horwood:2010:MEM

[Hor10] Joe Horwood. Marine ecosystem management: Fish abundance and size under exploitation. Marine Policy, 34(6): 1203–1206, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000734.

Houghton:2014:INP

[Hou14] Katherine Houghton. Identifying new pathways for ocean governance: the role of legal principles in areas beyond national jurisdiction. Marine Policy, 49(??):118–126, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14001298.

Howard:2018:BGS

[How18] Brian Clark Howard. Blue growth: Stakeholder perspectives. Marine Policy, 87(??):375–377, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306930.

House:2012:ISE

[HP12] Christopher House and Michael R. Phillips. Integrating the science education nexus into coastal governance: a Mediterranean and Black Sea case study. Marine Policy, 36(2):495–501, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001229.

Hastings:2013:MLP

[HP13] Emily Hastings and Tavis Potts. Marine litter: Progress in developing an integrated policy approach in Scotland. Marine Policy, 42(??):49–55, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000419.

Heenan:2015:CIE

[HPB^{+}15] Adel Heenan, Robert Pomeroy, Johann Bell, Philip L. Munday, William Cheung, Cheryl Logan, Russell Brainard, Affendi Yang Amri, Porfirio Aliño, Nygiel Armada, Laura David, Rebecca Rivera-Guieb, Stuart Green, Jamaluddin Jompa, Teresa Leonardo, Samuel Mamauag, Britt Parker, Janna Shackeroff, and Zulfigar Yasin. A climate-informed, ecosystem approach to fisheries management. Marine Policy, 57(??):182–192, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000676.

Hartley:2018:TTT

[HPH^{+}18] Bonny L. Hartley, Sabine Pahl, Matthew Holland, Iro Alampei, Joana M. Veiga, and Richard C. Thompson. Turning the tide on trash: Empowering European educators and school students to tackle marine litter. Marine Policy, 96(??):227–234, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305122.

Heck:2017:MPS

[HPP^{+}17a] Nadine Heck, Adina Paytan, Donald C. Potts, Brent Haddad, and Karen Lykkebo Petersen. Management priorities for seawater desalination plants in a marine protected area: a multi-criteria analysis. Marine Policy, 86(??):64–71, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17304761.

Helvey:2017:CUS

[HPP^{+}17b] Mark Helvey, Caroline Pomeroy, Naresh C. Pradhan, Dale Squires, and Stephen Stohs. Can the United States have its fish and eat it too? Marine Policy, 75(??):62–67, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305255.

Heck:2016:CRL

[HPPH16] Nadine Heck, Adina Paytan, Donald C. Potts, and Brent Haddad. Coastal residents’ literacy about seawater desalination and its impacts on marine ecosystems in California. Marine Policy, 68(??):178–186, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300859.

Hill:2018:IGI

[HPW^{+}18] Nicholas J. Hill, Thomas Peatman, Corey B. Wakefield, Stephen J. Newman, Tuikolongahau Halafihi, Jeff Kinch, Charles T. T. Edwards, Simon J. Nicol, and Ashley J. Williams. Improving guidelines for implementing harvest strategies in capacity-limited fisheries — Lessons from Tonga’s deepwater line fishery. Marine Policy, 98(??):85–91, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300654.

Houghton:2014:IAG

[HR14] Katherine Houghton and Julien Rochette. Introduction: Advancing governance of areas beyond national jurisdiction. Marine Policy, 49(??):81–84, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001304.

Harring:2016:GER

[HR16] Niklas Harring and Björn Rönnerstrand. Government effectiveness, regulatory compliance and public preference for marine policy instruments. An experimental approach. Marine Policy, 71(??):106–110, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303128.

Hastings:2017:DPN

[HR17] Kirrily Hastings and Karina L. Ryan. Differences in perception of a newly created Marine Park in south-west Western Australia by boat-based recreational fishers and the broader community. Marine Policy, 77(??):65–77, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303359.

Hyder:2015:MMC

[HRA^{+}15] Kieran Hyder, Axel G. Rossberg, J. Icarus Allen, Melanie C. Austen, Rosa M. Barciela, Hayley J. Bannister, Paul G. Blackwell, Julia L. Blanchard, Michael T. Burrows, Emma Defriez, Tarquin Dorrington, Karen P. Edwards, Bernardo Garcia-Carreras, Michael R. Heath, Deborah J. Hembury, Johanna J. Heymans, Jason Holt, Jennifer E. Houle, Simon Jennings, Steve Mackinson, Stephen J. Malcolm, Ruaraidh McPike, Laurence Mee, David K. Mills, Caron Montgomery, Dean Pearson, John K. Pinnegar, Marilena Pollicino, Ekaterina E. Popova, Louise Rae, Stuart I. Rogers, Douglas Speirs, Michael A. Spence, Robert Thorpe, R. Kerry Turner, Johan van der Molen, Andrew Yool, and David M. Paterson. Making modelling count — increasing the contribution of shelf-seas community and ecosystem models to policy development and management. Marine Policy, 61(??):291–302, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500216X.

Ho:2015:PSF

[HRC15] Nga Thi Thanh Ho, Helen Ross, and Jeffrey Coutts. Power sharing in fisheries co-management in Tam Giang Lagoon, Vietnam. Marine Policy, 53(??):171–179, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003339.

Holm:2015:CSC

[HRJH15] Petter Holm, Jesper Raakjær, Rikke Becker Jacobsen, and Edgar Henriksen. Contesting the social contracts underpinning fisheries — Lessons from Norway, Iceland and Greenland. Marine Policy, 55(??):64–72, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000226.

Herrera-Racionero:2015:UTF

[HRLFMP15] Paloma Herrera-Racionero, Emmánuel Lizcano-Fernández, and Lluís Miret-Pastor. “Us” and “them”. fishermen from Gandía and the loss of institutional legitimacy. Marine Policy, 54(??):130–136, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003558.

Hueter:2019:RRU

[HS19] Robert E. Hueter and David S. Shiffman. Rebuttal to “Response to ‘A United States shark fin ban would undermine sustainable shark fisheries’ I. F. Porcher et al., Marine Policy 104 (2019) 85–89”. Marine Policy, 110(??):Article 103601, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302118.

Huang:2016:IHT

[HSB^{+}16] Hsiang-Wen Huang, Yonat Swimmer, Keith Bigelow, Alexis Gutierrez, and Daniel G. Foster. Influence of hook type on catch of commercial and bycatch species in an Atlantic tuna fishery. Marine Policy, 65(??):68–75, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003863.

Hentati-Sundberg:2014:CFM

[HSH14] J. Hentati-Sundberg and J. Hjelm. Can fisheries management be quantified? Marine Policy, 48(??):18–20, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000712.

Hynes:2018:PAT

[HSR^{+}18] Stephen Hynes, Kathrine Skoland, Elisa Ravagnan, Brita Gjerstad, and Anne Vatland Krøvel. Public attitudes toward aquaculture: an Irish and Norwegian comparative study. Marine Policy, 96(??):1–8, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308503.

Heard:2016:ASL

[HSRH16] Matthew Heard, Stephen Sutton, Paul Rogers, and Charlie Huveneers. Actions speak louder than words: Tournament angling as an avenue to promote best practice for pelagic shark fishing. Marine Policy, 64(??):168–173, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003632.

Holland:2017:CVB

[HSW17] Daniel S. Holland, Erin Steiner, and Amanda Warlick. Can vessel buybacks pay off: an evaluation of an industry funded fishing vessel buyback. Marine Policy, 82(??):8–15, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305711.

Holland:2015:UCS

[HTA^{+}15a] Daniel S. Holland, Eric Thunberg, Juan Agar, Scott Crosson, Chad Demarest, Stephen Kasperski, Larry Perruso, Erin Steiner, Jessica Stephen, Andy Strelcheck, and Mike Travis. US catch share markets: a review of data availability and impediments to transparent markets. Marine Policy, 57(??):103–110, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000767.

Hyder:2015:CCS

[HTA^{+}15b] Kieran Hyder, Bryony Townhill, Lucy G. Anderson, Jane Delany, and John K. Pinnegar. Can citizen science contribute to the evidence-base that underpins marine policy? Marine Policy, 59(??):112–120, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001049.

Halliday:2018:TVT

[HTD^{+}18] William D. Halliday, Pierre-Louis Têtu, Jackie Dawson, Stephen J. Insley, and R. Casey Hilliard. Tourist vessel traffic in important whale areas in the western Canadian Arctic: Risks and possible management solutions. Marine Policy, 97(??):72–81, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305359.

Holmberg:2015:PGR

[HTF^{+}15] Robert J. Holmberg, Michael F. Tlusty, Elizabeth Futoma, Les Kaufman, James A. Morris, and Andrew L. Rhyne. The 800-pound grouper in the room: Asymptotic body size and invasiveness of marine aquarium fishes. Marine Policy, 53(??):7–12, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002863.

Hynes:2013:VIC

[HTH13] Stephen Hynes, Dugald Tinch, and Nick Hanley. Valuing improvements to coastal waters using choice experiments: an application to revisions of the EU Bathing Waters Directive. Marine Policy, 40(??):137–144, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000067.

Hanich:2010:CAP

[HTT10] Quentin Hanich, Feleti Teo, and Martin Tsamenyi. A collective approach to Pacific islands fisheries management: Moving beyond regional agreements. Marine Policy, 34 (1):85–91, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000700.

Huang:2011:BHS

[Hua11] Hsiang-Wen Huang. Bycatch of high sea longline fisheries and measures taken by Taiwan: Actions and challenges. Marine Policy, 35(5):712–720, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000480.

Huang:2014:MAI

[Hua14] Jie Huang. Maritime archaeology and identification of historic shipwrecks: a legal perspective. Marine Policy, 44(??):256–264, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002145.

Holen:2018:OSAa

[HUHA18a] Siri Mariane Holen, Ingrid Bouwer Utne, Ingunn Marie Holmen, and Halvard Aasjord. Occupational safety in aquaculture — Part 1: Injuries in Norway. Marine Policy, 96(??):184–192, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303906.

Holen:2018:OSAb

[HUHA18b] Siri Mariane Holen, Ingrid Bouwer Utne, Ingunn Marie Holmen, and Halvard Aasjord. Occupational safety in aquaculture — Part 2: Fatalities in Norway 1982– 2015. Marine Policy, 96(??):193–199, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303918.

Hunt:2013:BOC

[Hun13] Colin Hunt. Benefits and opportunity costs of Australia’s Coral Sea marine protected area: a precautionary tale. Marine Policy, 39(??):352–360, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002539.

Hutniczak:2019:CBR

[Hut19] Barbara Hutniczak. Coordination between RFMOs on mutual recognition of IUU vessel lists. Marine Policy, 107(??):Article 103596, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930137X.

Heenehan:2017:UAP

[HVB^{+}17] Heather L. Heenehan, Sofie M. Van Parijs, Lars Bejder, Julian A. Tyne, and David W. Johnston. Using acoustics to prioritize management decisions to protect coastal dolphins: a case study using Hawaiian spinner dolphins. Marine Policy, 75(??):84–90, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306078.

Ha:2013:FLN

[HvD13] Tran Thi Phung Ha and Han van Dijk. Fishery livelihoods and (non-)compliance with fishery regulations — a case study in Ca Mau Province, Mekong Delta, Viet Nam. Marine Policy, 38(??):417–427, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001601.

Hornborg:2019:EBF

[HvPN^{+}19] Sara Hornborg, Ingrid van Putten, Camilla Novaglio, Elizabeth A. Fulton, Julia L. Blanchard, Éva Plagányi, Cathy Bulman, and Keith Sainsbury. Ecosystembased fisheries management requires broader performance indicators for the human dimension. Marine Policy, 108(??):Article 103639, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309126.

Holland:2010:FFP

[HW10] Daniel S. Holland and Joshua Wiersma. Free form property rights for fisheries: the decentralized design of rights-based management through groundfish “sectors” in New England. Marine Policy, 34(5):1076–1081, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000059X.

Halim:2019:DFD

[HWL^{+}19] Abdul Halim, Budy Wiryawan, Neil R. Loneragan, Adrian Hordyk, M. Fedi A. Sondita, Alan T. White, Sonny Koeshendrajana, Toni Ruchimat, Robert S. Pomeroy, and Christiana Yuni. Developing a functional definition of small-scale fisheries in support of marine capture fisheries management in Indonesia. Marine Policy, 100(??):238–248, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303063.

Hartnett:2011:IEW

[HWN11] Michael Hartnett, James G. Wilson, and Stephen Nash. Irish estuaries: Water quality status and monitoring implications under the water framework directive. Marine Policy, 35(6):810–818, December 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1100011X.

Hanich:2018:SSF

[HWO^{+}18] Quentin Hanich, Colette C. C. Wabnitz, Yoshitaka Ota, Moses Amos, Connie Donato-Hunt, and Andrew Hunt. Small-scale fisheries under climate change in the Pacific Islands region. Marine Policy, 88(??):279–284, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307340.

Han:2018:SEM

[HWY^{+}18] Yu Han, Fangyi Wei, Guanqiong Ye, Shengyun Yang, Panpan Ma, and Wenjia Hu. A study on evaluation the marine carrying capacity in Guangxi Province, China. Marine Policy, 91(??):66–74, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305651.

Harper:2013:WFC

[HZH^{+}13] Sarah Harper, Dirk Zeller, Melissa Hauzer, Daniel Pauly, and Ussif Rashid Sumaila. Women and fisheries: Contribution to food security and local economies. Marine Policy, 39(??):56–63, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002175.

Ignatius:2018:JTA

[IH18] Suvi Ignatius and Päivi Haapasaari. Justification theory for the analysis of the socio-cultural value of fish and fisheries: the case of Baltic salmon. Marine Policy, 88(??):167–173, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300027.

Isaksen:2015:SFT

[IHF15] John R. Isaksen, Øystein Hermansen, and Ola Flaaten. Stubborn fuel tax concessions: the case of fisheries in Norway. Marine Policy, 52(??):85–92, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002905.

Ishimura:2013:SCM

[IHS13] Gakushi Ishimura, Samuel Herrick, and Ussif Rashid Sumaila. Stability of cooperative management of the Pacific sardine fishery under climate variability. Marine Policy, 39(??):333–340, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002497.

Iliff:2010:CIF

[Ili10a] Mike Iliff. Contemporary initiatives on the future of the International Whaling Commission. Marine Policy, 34(3):461–467, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001481.

Iliff:2010:HIF

[Ili10b] Mike Iliff. The Hogarth initiative on the future of the International Whaling Commission. Marine Policy, 34 (3):360–366, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001213.

Islam:2016:EIS

[IMA16] Md Monirul Islam, Essam Yassin Mohammed, and Liaquat Ali. Economic incentives for sustainable hilsa fishing in Bangladesh: an analysis of the legal and institutional framework. Marine Policy, 68(??):8–22, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1600049X.

Ingram:2018:RCS

[IOG18] Rebecca J. Ingram, Kirsten L. L. Oleson, and Jamison M. Gove. Revealing complex social-ecological interactions through participatory modeling to support ecosystembased management in Hawai‘ i. Marine Policy, 94(??): 180–188, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308941.

Irish:2018:IEH

[Iri18] Onni Irish. Identifying ecological hotspots in the United States and Norway: Turning ecosystem-based management into practice? Marine Policy, 98(??):65–76, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302458.

Islam:2014:LBA

[ISHP14] Md. Monirul Islam, Susannah Sallu, Klaus Hubacek, and Jouni Paavola. Limits and barriers to adaptation to climate variability and change in Bangladeshi coastal fishing communities. Marine Policy, 43(??):208–216, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001334.

Islam:2017:ECC

[ISM^{+}17] Mohammad Mahmudul Islam, Md Mostafa Shamsuzzaman, Mohammad Mojibul Hoque Mozumder, Xu Xiangmin, Yu Ming, and Md. Abu Sayed Jewel. Exploitation and conservation of coastal and marine fisheries in Bangladesh: Do the fishery laws matter? Marine Policy, 76(??):143–151, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630553X.

Iftekhar:2012:CSC

[IT12] M. S. Iftekhar and J. G. Tisdell. Comparison of simultaneous and combinatorial auction designs in fisheries quota market. Marine Policy, 36(2):446–453, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001424.

Iftekhar:2015:BPM

[IT15] Md Sayed Iftekhar and John G. Tisdell. Bidding and performance in multiple unit combinatorial fishery quota auctions: Role of information feedbacks. Marine Policy, 62(??):233–243, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002651.

IJsseldijk:2018:CBC

[ItDD^{+}18] Lonneke L. IJsseldijk, Mariel T. I. ten Doeschate, Nicholas J. Davison, Andrea Gröne, and Andrew C. Brownlow. Crossing boundaries for cetacean conservation: Setting research priorities to guide management of harbour porpoises. Marine Policy, 95(??):77–84, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301799.

Innes:2014:EPT

[ITNL^{+}14] James Innes, Olivier Thébaud, Ana Norman-López, L. Richard Little, and John Kung. Evidence of package trading in a mature multi-species ITQ market. Marine Policy, 46(??):68–71, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002960.

Isaacs:2019:FCH

[IW19] Moenieba Isaacs and Emma Witbooi. Fisheries crime, human rights and small-scale fisheries in South Africa: a case of bigger fish to fry. Marine Policy, 105(??):158–168, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309503.

Jakobsen:2010:IFC

[JA10] Stig-Erik Jakobsen and Bernt Aarset. Institutions as facilities for change?: a study of the coherence between political regulations and innovations within the pelagic fisheries sector in Norway. Marine Policy, 34(5):928–934, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000199.

Jacques:2015:WFG

[Jac15] Peter J. Jacques. Are world fisheries a global panarchy? Marine Policy, 53(??):165–170, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400325X.

Jacobsen:2019:PDF

[Jac19] Jóannes Jacobsen. Path dependence in Faroese fisheries (mis)management. Marine Policy, 108(??):Article 103615, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302878.

Jaeckel:2016:DSM

[Jae16] Aline Jaeckel. Deep seabed mining and adaptive management: the procedural challenges for the International Seabed Authority. Marine Policy, 70(??):205–211, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300756.

Jaeckel:2016:SBC

[JAG16] Aline Jaeckel, Jeff A. Ardron, and Kristina M. Gjerde. Sharing benefits of the common heritage of mankind — Is the deep seabed mining regime ready? Marine Policy, 70(??):198–204, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300744.

Jay:2016:TDM

[JAO^{+}16] Stephen Jay, Fátima L. Alves, Cathal O’Mahony, Maria Gomez, Aoibheann Rooney, Margarida Almodovar, Kira Gee, Juan Luis Suárez de Vivero, Jorge M. S. Gonçalves, Maria da Luz Fernandes, Olvido Tello, Sarah Twomey, Inmaculada Prado, Catarina Fonseca, Luis Bentes, Guida Henriques, and Aldino Campos. Transboundary dimensions of marine spatial planning: Fostering interjurisdictional relations and governance. Marine Policy, 65(??):85–96, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003954.

Johansen:2019:NSI

[JBBV^{+}19] Ulf Johansen, Heidi Bull-Berg, Lars H. Vik, Arne M. Stokka, Roger Richardsen, and Ulf Winther. The Norwegian seafood industry — Importance for the national economy. Marine Policy, 110(??):Article 103561, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830914X.

Jefferson:2014:PPU

[JBdL^{+}14] R. L. Jefferson, I. Bailey, D. d’A. Laffoley, J. P. Richards, and M. J. Attrill. Public perceptions of the UK marine environment. Marine Policy, 43(??):327–337, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001462.

Jagers:2012:WCA

[JBJ12] Sverker C. Jagers, Daniel Berlin, and Svein Jentoft. Why comply? Attitudes towards harvest regulations among Swedish fishers. Marine Policy, 36(5):969–976, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000115.

Ji:2015:MME

[JBK15] Xichen Ji, Jan Brinkhuis, and Sabine Knapp. A method to measure enforcement effort in shipping with incomplete information. Marine Policy, 60(??):162–170, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001748.

Jarvis:2015:CSP

[JBKB15] Rebecca M. Jarvis, Barbara Bollard Breen, Christian U. Krägeloh, and D. Rex Billington. Citizen science and the power of public participation in marine spatial planning. Marine Policy, 57(??):21–26, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000603.

Jacquet:2016:RUA

[JBPBA16] Jennifer Jacquet, Eli Blood-Patterson, Cassandra Brooks, and David Ainley. ‘Rational use’ in Antarctic waters. Marine Policy, 63(??):28–34, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002821.

Jackman:2018:CCS

[JBV^{+}18] Jennifer Jackman, Lauren Bettencourt, Jerry Vaske, Michele Sweeney, Katharine Bloom, Allen Rutberg, and Brandi Brook. Conflict and consensus in stakeholder views of seal management on Nantucket Island, MA, USA. Marine Policy, 95(??):166–173, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306450.

Jentoft:2010:PRA

[JCBM10] Svein Jentoft, Ratana Chuenpagdee, Alida Bundy, and Robin Mahon. Pyramids and roses: Alternative images for the governance of fisheries systems. Marine Policy, 34(6):1315–1321, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001107.

Jiang:2017:ECC

[JCD17] Daokui Jiang, Zhuo Chen, and Guilin Dai. Evaluation of the carrying capacity of marine industrial parks: a case study in China. Marine Policy, 77(??):111–119, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305243.

Jones:2016:VRR

[JD16] P. J. S. Jones and E. M. De Santo. Viewpoint — Is the race for remote, very large marine protected areas (VLMPAs) taking us down the wrong track? Marine Policy, 73(??):231–234, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630481X.

Jones:2019:EEM

[JDM^{+}19] Daniel O. B. Jones, Jennifer M. Durden, Kevin Murphy, Kristina M. Gjerde, Aleksandra Gebicka, Ana Colaço, Telmo Morato, Daphne Cuvelier, and David S. M. Billett. Existing environmental management approaches relevant to deep-sea mining. Marine Policy, 103(??):172–181, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303956.

Jones:2013:IEF

[JDQV13] P. J. S. Jones, E. M. De Santo, W. Qiu, and O. Vestergaard. Introduction: an empirical framework for deconstructing the realities of governing marine protected areas. Marine Policy, 41(??):1–4, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002667.

Johnsen:2011:SCF

[JE11] Jahn Petter Johnsen and Søren Eliasen. Solving complex fisheries management problems: What the EU can learn from the Nordic experiences of reduction of discards. Marine Policy, 35(2):130–139, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000151X.

Jenkins:2015:INP

[Jen15] David Jenkins. Impacts of neoliberal policies on nonmarket fishing economies on the Yukon River, Alaska. Marine Policy, 61(??):356–365, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003315.

Jetoo:2018:MLG

[Jet18] Savitri Jetoo. Multi-level governance innovations of the Baltic Sea and the North American Great Lakes: New actors and their roles in building adaptive capacity for eutrophication governance. Marine Policy, 98(??):237– 245, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306845.

Jensen:2017:FRS

[JFA17] Frank Jensen, Hans Frost, and Jens Abildtrup. Fisheries regulation: a survey of the literature on uncertainty, compliance behavior and asymmetric information. Marine Policy, 81(??):167–178, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305292.

Johnson:2018:CCL

[JFK18] David Johnson, Maria Adelaide Ferreira, and Ellen Kenchington. Climate change is likely to severely limit the effectiveness of deep-sea ABMTs in the North Atlantic. Marine Policy, 87(??):111–122, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304268.

Johnson:2018:REP

[JFT^{+}18] David E. Johnson, Christopher Barrio Froján, Phillip J. Turner, Philip Weaver, Vikki Gunn, Daniel C. Dunn, Patrick Halpin, Nicholas J. Bax, and Piers K. Dunstan. Reviewing the EBSA process: Improving on success. Marine Policy, 88(??):75–85, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730711X.

Jenkins:2013:FGS

[JG13] Lekelia D. Jenkins and Karen Garrison. Fishing gear substitution to reduce bycatch and habitat impacts: an example of social–ecological research to inform policy. Marine Policy, 38(??):293–303, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001443.

Jaeckel:2017:CCH

[JGA17] Aline Jaeckel, Kristina M. Gjerde, and Jeff A. Ardron. Conserving the common heritage of humankind — Options for the deep-seabed mining regime. Marine Policy, 78(??):150–157, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306716.

Jaffry:2016:EBM

[JGG^{+}16] S. Jaffry, H. Glenn, Y. Ghulam, T. Willis, and C. Delanbanque. Are expectations being met? Consumer preferences and rewards for sustainably certified fisheries. Marine Policy, 73(??):77–91, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300719.

Jones:2016:CAT

[JGMS16] Estelle Victoria Jones, Tim Gray, Donald Macintosh, and Selina Stead. A comparative analysis of three marine governance systems for implementing the Convention on Biological Diversity (CBD). Marine Policy, 66(??):30–38, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1600021X.

Jambeck:2018:CES

[JHB^{+}18] Jenna Jambeck, Britta Denise Hardesty, Amy L. Brooks, Tessa Friend, Kristian Teleki, Joan Fabres, Yannick Beaudoin, Abou Bamba, Julius Francis, Anthony J. Ribbink, Tatjana Baleta, Hindrik Bouwman, Jonathan Knox, and Chris Wilcox. Challenges and emerging solutions to the land-based plastic waste issue in Africa. Marine Policy, 96(??):256–263, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305286.

Jeffers:2019:TUS

[JHN^{+}19] Victoria F. Jeffers, Frances Humber, Thierry Nohasiarivelo, Radonirina Botosoamananto, and Lucy G. Anderson. Trialling the use of smartphones as a tool to address gaps in small-scale fisheries catch data in southwest Madagascar. Marine Policy, 99(??):267–274, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303634.

Jin:2013:EAE

[JHW13] Di Jin, Porter Hoagland, and Brooke Wikgren. An empirical analysis of the economic value of ocean space associated with commercial fishing. Marine Policy, 42(??): 74–84, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000237.

Jay:2019:TFH

[JJ19] Stephen Jay and Hannah Jones. Towards a framework for higher education for marine spatial planning. Marine Policy, 99(??):230–238, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304561.

Janssen:2013:STS

[JKK13] Holger Janßen, Sue Kidd, and Torunn Kvinge. A spatial typology for the sea: a contribution from the Baltic. Marine Policy, 42(??):190–197, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000699.

Johnson:2013:MRC

[JKS13] Kate Johnson, Sandy Kerr, and Jonathan Side. Marine renewables and coastal communities — Experiences from the offshore oil industry in the 1970s and their relevance to marine renewables in the 2010s. Marine Policy, 38(??):491–499, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001716.

Johnson:2016:PFO

[JKS16] Kate R. Johnson, Sandy A. Kerr, and Jonathan C. Side. The Pentland Firth and Orkney Waters and Scotland — Planning Europe’s Atlantic gateway. Marine Policy, 71(??):285–292, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003693.

Jones:2011:DVM

[JL11] Christopher D. Jones and Susanne J. Lockhart. Detecting Vulnerable Marine Ecosystems in the Southern Ocean using research trawls and underwater imagery. Marine Policy, 35(5):732–736, September 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000285.

Junior:2016:MWM

[JLCB16] José Gilmar C. Oliveira Júnior, Richard J. Ladle, Ricardo Correia, and Vandick S. Batista. Measuring what matters — Identifying indicators of success for Brazilian marine protected areas. Marine Policy, 74(??):91–98, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300434.

Jacobsen:2014:GSE

[JLH14] Kelsey I. Jacobsen, Sarah E. Lester, and Benjamin S. Halpern. A global synthesis of the economic multiplier effects of marine sectors. Marine Policy, 44(??): 273–278, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002169.

Jones:2016:MSP

[JLQ16] Peter J. S. Jones, L. M. Lieberknecht, and W. Qiu. Marine spatial planning in reality: Introduction to case studies and discussion of findings. Marine Policy, 71(??): 256–264, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302147.

Jaiteh:2017:ESF

[JLW17] Vanessa F. Jaiteh, Neil R. Loneragan, and Carol Warren. The end of shark finning? Impacts of declining catches and fin demand on coastal community livelihoods. Marine Policy, 82(??):224–233, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301720.

Jin:2018:PVB

[JLY18] Se-Jun Jin, Seul-Ye Lim, and Seung-Hoon Yoo. The public value of building large oil spill response vessels in Korea. Marine Policy, 88(??):242–247, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305559.

James:2018:ADI

[JMJ^{+}18] Mark James, Tania Mendo, Esther L. Jones, Kyla Orr, Ali McKnight, and John Thompson. AIS data to inform small scale fisheries management and marine spatial planning. Marine Policy, 91(??):113–121, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305845.

Jakobsen:2016:ESD

[JML16] Ingvild Ulrikke Jakobsen and Nele Matz-Lück. Environmental standards for deep seabed mining: Introduction. Marine Policy, 70(??):174, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630183X.

Jamero:2019:SAA

[JOE^{+}19] Ma. Laurice Jamero, Motoharu Onuki, Miguel Esteban, Christopher Chadwick, Nicholson Tan, Ven Paolo Valenzuela, Richard Crichton, and John Erick Avelino. In-situ adaptation against climate change can enable relocation of impoverished small islands. Marine Policy, 108(??): Article 103614, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303324.

Johnson:2019:PLC

[Joh19] David Edward Johnson. Protecting the lost city hydrothermal vent system: All is not lost, or is it? Marine Policy, 107(??):Article 103593, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303252.

Jones:2013:GAG

[Jon13] Peter J. S. Jones. A governance analysis of the Galápagos Marine Reserve. Marine Policy, 41(??):65–71, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002606.

Jones:2013:GMP

[JQD13] P. J. S. Jones, W. Qiu, and E. M. De Santo. Governing marine protected areas: Social–ecological resilience through institutional diversity. Marine Policy, 41(??):5–13, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002679.

Jouanneau:2014:HTL

[JR14] Charlène Jouanneau and Jesper Raakjær. ‘The Hare and the Tortoise’: Lessons from Baltic Sea and Mediterranean Sea governance. Marine Policy, 50(??):331–338, December 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000785.

Jones:2017:SAC

[JRP17] Russ Jones, Catherine Rigg, and Evelyn Pinkerton. Strategies for assertion of conservation and local management rights: a Haida Gwaii herring story. Marine Policy, 80 (??):154–167, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306005.

Jones:2016:VTW

[JRS^{+}16] Natalie A. Jones, Helen Ross, Sylvie Shaw, Katherine Witt, Breanna Pinner, and David Rissik. Values towards waterways in south east Queensland: Why people care. Marine Policy, 71(??):121–131, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300550.

Jardine:2012:CSP

[JS12] Sunny L. Jardine and James N. Sanchirico. Catch share programs in developing countries: a survey of the literature. Marine Policy, 36(6):1242–1254, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000541.

Jardine:2013:CCS

[JS13] Sunny L. Jardine and James N. Sanchirico. Corrigendum to “Catch share programs in developing countries: a survey of the literature” [J. Mar. Policy 2012; 36 (6), 1242–1254]. Marine Policy, 42(??):166, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300002X.

Jenkins:2018:ACB

[JS18] Tom L. Jenkins and Jamie R. Stevens. Assessing connectivity between MPAs: Selecting taxa and translating genetic data to inform policy. Marine Policy, 94(??):165–173, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307984.

Jimenez:2019:EFI

[JSD19] Jesús M. Jiménez and Lina M. Saavedra-Díaz. Evaluating formal and informal rules as a basis for implementing coastal marine artisanal fisheries management in Colombia. Marine Policy, 101(??):225–236, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302276.

Johnson:2018:PPP

[JSG^{+}18] David Edward Johnson, Erick Ross Salazar, Anthony Gallagher, Aldous Rees, Carolina Sheridan Rodriguez, Sergio Cambronero Solano, Gustavo Rojas Ortega, and Christopher Barrio Froján. Preventing plastics pervading an oceanic oasis: Building the case for the Costa Rica Thermal Dome to become a World Heritage site in ABNJ. Marine Policy, 96(??):235–242, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300150.

Jin:2019:MPP

[JSLL19] Mengjie Jin, Wenming Shi, K. C. Lin, and Kevin X. Li. Marine piracy prediction and prevention: Policy implications. Marine Policy, 108(??):Article 103528, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830705X.

Jin:2019:OTR

[JSY^{+}19] Mengjie Jin, Wenming Shi, Kum Fai Yuen, Yi Xiao, and Kevin X. Li. Oil tanker risks on the marine environment: an empirical study and policy implications. Marine Policy, 108(??):Article 103655, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303008.

Jenkins:2017:SFL

[JTBP17] Lekelia D. Jenkins, Katie R. Thompson, Luis Bourillon, and S. Hoyt Peckham. The scope of fisheries learning exchanges for conservation. Marine Policy, 77(??):196–204, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302913.

James:2018:IIT

[JTK18] Philip A. S. James, Alex Tidd, and Lotokufaki Paka Kaitu. The impact of industrial tuna fishing on small-scale fishers and economies in the Pacific. Marine Policy, 95(??):189–198, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308825.

Jacob:2018:MBO

[JTP18] Céline Jacob, Sébastien Thorin, and Sylvain Pioch. Marine biodiversity offsetting: an analysis of the emergence of an environmental governance system in California. Marine Policy, 93(??):128–141, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306383.

Jacob:2010:EFD

[JWBJ10] Steve Jacob, Priscilla Weeks, Benjamin G. Blount, and Michael Jepson. Exploring fishing dependence in Gulf Coast communities. Marine Policy, 34(6):1307– 1314, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001090.

Jacob:2013:DES

[JWBJ13] Steve Jacob, Priscilla Weeks, Ben Blount, and Michael Jepson. Development and evaluation of social indicators of vulnerability and resiliency for fishing communities in the Gulf of Mexico. Marine Policy, 37(??):86–95, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000759.

Johnson:2016:ARV

[JWM^{+}16] Johanna E. Johnson, David J. Welch, Jeffrey A. Maynard, Johann D. Bell, Gretta Pecl, Julie Robins, and Thor Saunders. Assessing and reducing vulnerability to climate change: Moving from theory to practical decisionsupport. Marine Policy, 74(??):220–229, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304626.

Jarvis:2019:KRP

[JY19] Rebecca M. Jarvis and Tim Young. Key research priorities for the future of marine science in New Zealand. Marine Policy, 106(??):Article 103539, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309059.

Kaplan:2015:LEC

[KAF^{+}15] Katherine A. Kaplan, Gabby N. Ahmadia, Helen Fox, Louise Glew, Emily F. Pomeranz, and Patrick Sullivan. Linking ecological condition to enforcement of marine protected area regulations in the greater Caribbean region. Marine Policy, 62(??):186–195, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002699.

Kaiser:2015:HCA

[Kai15a] Mark J. Kaiser. Hurricane clean-up activity in the Gulf of Mexico, 2004–2013. Marine Policy, 51(??):512–526, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002346.

Kaiser:2015:OSV

[Kai15b] Mark J. Kaiser. Offshore Service Vessel activity forecast and regulatory modeling in the U.S. Gulf of Mexico, 2012–2017. Marine Policy, 57(??):132–146, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000718.

Kaiser:2016:SVA

[Kai16] Mark J. Kaiser. Service vessel activity in the U.S. Gulf of Mexico in support of the oil and gas industry using AIS data, 2009–2010. Marine Policy, 63(??):61–80, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002833.

Kaiser:2017:FPD

[Kai17a] Mark J. Kaiser. FERC pipeline decommissioning cost in the U.S. Gulf of Mexico, 1995–2015. Marine Policy, 82 (??):167–180, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306327.

Kaiser:2017:OPC

[Kai17b] Mark J. Kaiser. Offshore pipeline construction cost in the U.S. Gulf of Mexico. Marine Policy, 82(??):147–166, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305851.

Kaiser:2017:UGM

[Kai17c]  Mark J. Kaiser.  U.S. Gulf of Mexico deepwater

pipeline construction — a review of lessons learned. Marine Policy, 86(??):214–233, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304335.

Kaifu:2019:CAJ

[Kai19] Kenzo Kaifu. Challenges in assessments of Japanese eel stock. Marine Policy, 102(??):1–4, April 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308169.

Kamat:2018:DDM

[Kam18] Vinay R. Kamat. Dispossession and disenchantment: the micropolitics of marine conservation in southeastern Tanzania. Marine Policy, 88(??):261–268, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303482.

Koen-Alonso:2019:NAF

[KAPF^{+}19] Mariano Koen-Alonso, Pierre Pepin, Michael J. Fogarty, Andrew Kenny, and Ellen Kenchington. The Northwest Atlantic Fisheries Organization Roadmap for the development and implementation of an Ecosystem Approach to Fisheries: structure, state of development, and challenges. Marine Policy, 100(??):342–352, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305694.

Karnad:2017:NCL

[Kar17] Divya Karnad. Navigating customary law and state fishing legislation to create effective fisheries governance in India. Marine Policy, 86(??):241–246, December 2017.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304396.

Katila:2019:DQS

[KARR^{+}19] Jenny Katila, Katariina Ala-Rämi, Sari Repka, Esteban Rendon, and Janne Törrönen. Defining and quantifying the sea-based economy to support regional blue growth strategies — Case Gulf of Bothnia. Marine Policy, 100(??):215–225, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301659.

Kasperski:2016:OMH

[Kas16] Stephen Kasperski. Optimal multispecies harvesting in the presence of a nuisance species. Marine Policy, 64 (??):55–63, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003280.

Katikiro:2016:IAL

[Kat16] Robert E. Katikiro. Improving alternative livelihood interventions in marine protected areas: a case study in Tanzania. Marine Policy, 70(??):22–29, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302160.

Kaye:2015:APR

[Kay15] Stuart Kaye. Australian practice in respect of the continental shelf beyond 200 nautical miles. Marine Policy, 51 (??):339–346, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002498.

Krumholz:2015:FCG

[KB15] Jason S. Krumholz and Michael L. Brennan. Fishing for common ground: Investigations of the impact of trawling on ancient shipwreck sites uncovers a potential for management synergy. Marine Policy, 61(??):127–133, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002109.

Kaiser:2015:SIA

[KBB^{+}15] Brooks A. Kaiser, Sergey Bakanev, Rasmus Gjedsø Bertelsen, Marcus Carson, Arne Eide, Linda Fernandez, Patrick Halpin, Sergei Izmalkov, Line A. Kyhn, Henrik Österblom, Maarten Punt, Lars Ravn-Jonsen, James Sanchirico, Konstantin Sokolov, Jan H. Sundet, Gudrun G. Thorarinsdóttir, and Niels Vestergaard. Spatial issues in Arctic marine resource governance workshop summary and comment. Marine Policy, 58(??):1–5, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000822.

Kossmann:2016:AIC

[KBB16] Christina M. Kossmann, Jelle H. Behagel, and Megan Bailey. Action and inertia in collaborative governance. Marine Policy, 72(??):21–30, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630080X.

Kuriyama:2016:CSL

[KBBR16] Peter T. Kuriyama, Trevor A. Branch, Marlene A. Bellman, and Kate Rutherford. Catch shares have not led to catch-quota balancing in two North American multispecies trawl fisheries. Marine Policy, 71(??):60–70, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302925.

Kraak:2013:LFM

[KBC^{+}13] Sarah B. M. Kraak, Nick Bailey, Massimiliano Cardinale, Chris Darby, José A. A. De Oliveira, Margit Eero, Norman Graham, Steven Holmes, Tore Jakobsen, Alexander Kempf, Eskild Kirkegaard, John Powell, Robert D. Scott, E. John Simmonds, Clara Ulrich, Willy Vanhee, and Morten Vinther. Lessons for fisheries management from the EU cod recovery plan. Marine Policy, 37(??):200–213, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000905.

Kearney:2012:ANT

[KBF12] R. Kearney, C. D. Buxton, and G. Farebrother. Australia’s no-take marine protected areas: Appropriate conservation or inappropriate management of fishing? Marine Policy, 36(5):1064–1071, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000413.

Kauer:2018:RBT

[KBG^{+}18] Kate Kauer, Lyall Bellquist, Mary Gleason, Aliya Rubinstein, Joe Sullivan, Dwayne Oberhoff, Lisa Damrosch, Michelle Norvell, and Michael Bell. Reducing bycatch through a risk pool: a case study of the U.S. West Coast groundfish fishery. Marine Policy, 96(??):90–99, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302069.

Kearney:2012:QIP

[KBGF12] R. Kearney, C. D. Buxton, P. Goodsell, and G. Farebrother. Questionable interpretation of the Precautionary Principle in Australia’s implementation of ‘no-take’ marine protected areas. Marine Policy, 36(3):592–597, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001692.

Kim:2019:AVK

[KBK19] Bong-Tae Kim, Christopher L. Brown, and Do-Hoon Kim. Assessment on the vulnerability of Korean aquaculture to climate change. Marine Policy, 99(??):111–122, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303105.

Kampf:2013:HRE

[KC13a] Jochen Kämpf and Beverley Clarke. How robust is the environmental impact assessment process in South Australia? Behind the scenes of the Adelaide seawater desalination project. Marine Policy, 38(??):500–506, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001728.

Ku:2013:CPB

[KC13b] Kuo-Cheng Ku and Tsen-Chien Chen. A conceptual process-based reference model for collaboratively managing recreational scuba diving in Kenting National Park. Marine Policy, 39(??):1–10, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002096.

Koslow:2015:POO

[KC15] J. Anthony Koslow and Jennifer Couture. Pacific Ocean observation programs: Gaps in ecological time series. Marine Policy, 51(??):408–414, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400236X.

Krupa:2018:RPR

[KCCO18] Meagan Boltwood Krupa, Molly McCarthy Cunfer, S. Jeanette Clark, and Emily O’Dean. Resurrecting the public record: Assessing stakeholder participation in Alaska’s fisheries. Marine Policy, 96(??):36–43, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301817.

Kraska:2015:BLM

[KCJ15] James Kraska, Guillermo Ortuño Crespo, and David W. Johnston. Bio-logging of marine migratory species in the law of the sea. Marine Policy, 51(??):394–400, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002322.

Kerr:2015:ROS

[KCJW15] Sandy Kerr, John Colton, Kate Johnson, and Glen Wright. Rights and ownership in sea country: implications of marine renewable energy for indigenous and local communities. Marine Policy, 52(??):108–115, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002942.

Kenny:2018:DSF

[KCKA^{+}18] Andrew J. Kenny, Neil Campbell, Mariano Koen-Alonso, Pierre Pepin, and Daniela Diz. Delivering sustainable fisheries through adoption of a risk-based framework as part of an ecosystem approach to fisheries management. Marine Policy, 93(??):232–240, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306856.

Kienker:2018:BHA

[KCM^{+}18] S. E. Kienker, R. A. Coleman, R. L. Morris, P. Steinberg, B. Bollard, R. Jarvis, K. A. Alexander, and E. M. A. Strain. Bringing harbours alive: Assessing the importance of ecoengineered coastal infrastructure for different stakeholders and cities. Marine Policy, 94(??):238–246, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300666.

Kuznetsov:2015:TSM

[KDG^{+}15] Andrei Kuznetsov, John Dinwoodie, David Gibbs, Mark Sansom, and Harriet Knowles. Towards a sustainability management system for smaller ports. Marine Policy, 54(??):59–68, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003534.

Knowles:2015:EMC

[KDS^{+}15] John E. Knowles, Emma Doyle, Steven R. Schill, Lynnette M. Roth, Amy Milam, and George T. Raber. Establishing a marine conservation baseline for the insular Caribbean. Marine Policy, 60(??):84–97, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001384.

Kittinger:2010:CFC

[KDW10] John N. Kittinger, Kristin N. Duin, and Bruce A. Wilcox. Commercial fishing, conservation and compatibility in the Northwestern Hawaiian Islands. Marine Policy, 34 (2):208–217, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000967.

Klinger:2018:MBG

[KED^{+}18] Dane H. Klinger, Anne Maria Eikeset, Brynhildur Davídsdóttir, Anna-Marie Winter, and James R. Watson. The mechanics of blue growth: Management of oceanic natural resource use with multiple, interacting sectors. Marine Policy, 87(??):356–362, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305869.

Keessen:2018:WSC

[Kee18] Andrea Keessen. What states can do to adapt to climate change in the Baltic Sea. Marine Policy, 98(??): 295–300, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306985.

Kelly:2018:CCM

[KEF18] Christina Kelly, Geraint Ellis, and Wesley Flannery. Conceptualising change in marine governance: Learning from Transition Management. Marine Policy, 95(??): 24–35, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307480.

Kelly:2012:RPM

[KEG12] Claire Kelly, Stephen Essex, and Gillian Glegg. Reflective practice for marine planning: a case study of marine nature-based tourism partnerships. Marine Policy, 36(3):769–781, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001916.

Kelly:2018:NSS

[Kel18] Catherine Kelly. ‘I Need the Sea and the Sea Needs Me’: Symbiotic coastal policy narratives for human wellbeing and sustainability in the UK. Marine Policy, 97(??): 223–231, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301738.

King:2018:MBD

[KEM^{+}18] Andrew S. King, Nicholas G. Elliott, Catriona K. Macleod, Mark A. James, and Jeffrey M. Dambacher. Making better decisions: Utilizing qualitative signed digraphs modeling to enhance aquaculture production technology selection. Marine Policy, 91(??):22–33, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308217.

Keppel:2014:IEK

[Kep14] Gunnar Keppel. The importance of expert knowledge in conservation planning — Comment to an article by C. J. Klein et al. Marine Policy, 48(??):202–203, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000839.

Kluger:2019:UMA

[KFB19] Lotta C. Kluger, Ramón Filgueira, and Carrie J. Byron. Using media analysis to scope priorities in social carrying capacity assessments: a global perspective. Marine Policy, 99(??):252–261, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303981.

Kearney:2013:HTM

[KFBG13] R. Kearney, G. Farebrother, C. D. Buxton, and P. Goodsell. How terrestrial management concepts have led to unrealistic expectations of marine protected areas. Marine Policy, 38(??):304–311, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001455.

Konar:2017:DBP

[KFM17] B. Konar, L. Frisch, and S. B. Moran. Development of best practices for scientific research vessel operations in a changing Arctic: a case study for R/V Sikuliaq. Marine Policy, 86(??):182–189, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305043.

Khalilian:2010:DFC

[KFPR10] Setareh Khalilian, Rainer Froese, Alexander Proelss, and Till Requate. Designed for failure: a critique of the Common Fisheries Policy of the European Union. Marine Policy, 34(6):1178–1182, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000709.

Karadzic:2014:SLF

[KGAB14] Vanja Karadzic, John Grin, Paula Antunes, and Marija Banovic. Social learning in fish producers’ organizations: How fishers perceive their membership experience and what they learn from it. Marine Policy, 44(??):427–437, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002327.

Khan:2018:IFM

[KGMP18] Alexander M. A. Khan, Tim S. Gray, Aileen C. Mill, and Nicholas V. C. Polunin. Impact of a fishing moratorium on a tuna pole-and-line fishery in eastern Indonesia. Marine Policy, 94(??):143–149, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300799.

Krovel:2019:EAT

[KGS^{+}19] A. V. Krøvel, B. Gjerstad, K. Skoland, K. M. Lindland, S. Hynes, and E. Ravagnan. Exploring attitudes toward aquaculture in Norway — Is there a difference between the Norwegian general public and local communities where the industry is established? Marine Policy, 108(??):Article 103648, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309606.

Kelly:2014:REM

[KGST14] Christina Kelly, Lorraine Gray, Rachel Shucksmith, and Jacqueline F. Tweddle. Review and evaluation of marine spatial planning in the Shetland Islands. Marine Policy, 46(??):152–160, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000293.

Kessler:2010:ATI

[KH10] Megan Kessler and Robert Harcourt. Aligning tourist, industry and government expectations: a case study from the swim with whales industry in Tonga. Marine Policy, 34(6):1350–1356, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001144.

Knudsen:2011:ILI

[KH11] Olav F. Knudsen and Björn Hassler. IMO legislation and its implementation: Accident risk, vessel deficiencies and national administrative practices. Marine Policy, 35 (2):201–207, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001673.

Kessler:2012:MIC

[KH12] Megan Kessler and Robert Harcourt. Management implications for the changing interactions between people and whales in Ha’apai, Tonga. Marine Policy, 36(2):440–445, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001394.

Kessler:2013:WWR

[KH13] Megan Kessler and Robert Harcourt. Whale watching regulation compliance trends and the implications for management off Sydney, Australia. Marine Policy, 42(??):

14–19, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000250.

Kusumawati:2015:KFS

[KH15] Ika Kusumawati and Hsiang-Wen Huang. Key factors for successful management of marine protected areas: a comparison of stakeholders’ perception of two MPAs in Weh Island, Sabang, Aceh, Indonesia. Marine Policy, 51(??):465–475, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002620.

Karlsen:2012:ELS

[KHD12] Kine Mari Karlsen, Øystein Hermansen, and Bent M. Dreyer. Eco-labeling of seafood: Does it affect the harvesting patterns of Norwegian fishermen? Marine Policy, 36(5):1123–1130, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000474.

Kessler:2013:SWT

[KHH13] Megan Kessler, Robert Harcourt, and Gillian Heller. Swimming with whales in Tonga: Sustainable use or threatening process? Marine Policy, 39(??):314–316, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002540.

Kratzer:2014:UOC

[KHP14] Susanne Kratzer, E. Therese Harvey, and Petra Philipson. The use of ocean color remote sensing in integrated coastal zone management — a case study from Himmerfjärden, Sweden. Marine Policy, 43(??):29–39, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001036.

Katarzyte:2017:PMS

[KHT17] Marija Kataržytė, Sven Hille, and Regina Terlecka. Promoting marine science: International science camp as a platform. Marine Policy, 84(??):76–81, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301410.

Kraan:2014:HDT

[KHvH^{+}14] Marloes Kraan, Astrid Hendriksen, Luc van Hoof, Judith van Leeuwen, and Charlène Jouanneau. How to dance? The tango of stakeholder involvement in marine governance research. Marine Policy, 50(??):347–352, December 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001407.

Kidd:2013:RIA

[Kid13] Sue Kidd. Rising to the integration ambitions of Marine Spatial Planning: Reflections from the Irish Sea. Marine Policy, 39(??):273–282, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002369.

Kim:2010:FDN

[Kim10] Suam Kim. Fisheries development in northeastern Asia in conjunction with changes in climate and social systems. Marine Policy, 34(4):803–809, July 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000485.

Kim:2015:SHN

[Kim15] Hyun Jung Kim. Success in heading north?: South Korea’s master plan for Arctic policy. Marine Policy, 61 (??):264–272, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002213.

Kim:2017:SDS

[Kim17] Rakhyun E. Kim. Should deep seabed mining be allowed? Marine Policy, 82(??):134–137, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308119.

Kim:2019:SKU

[Kim19] Hyun Jung Kim. South Korea’s use of force against Chinese illegal fishing in the course of law enforcement in the Yellow Sea. Marine Policy, 99(??):148–156, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302501.

Kirkfeldt:2019:OCW

[Kir19] Trine Skovgaard Kirkfeldt. An ocean of concepts: Why choosing between ecosystem-based management, ecosystem-based approach and ecosystem approach makes a difference. Marine Policy, 106(??):Article 103541, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309655.

Ko:2010:FYP

[KJH^{+}10] Jae-Young Ko, Glenn A. Jones, Moon-Soo Heo, Young-Su Kang, and Sang-Hyuck Kang. A fifty-year production and economic assessment of common property-based management of marine living common resources: a case study for the women divers communities in Jeju, South Korea. Marine Policy, 34(3):624–634, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001869.

Klein:2014:SCP

[KJP14] Carissa J. Klein, Stacy D. Jupiter, and Hugh P. Possingham. Setting conservation priorities in Fiji: Decision science versus additive scoring systems. Marine Policy, 48(??):204–205, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000803.

Kerr:2014:PEI

[KJS14] S. Kerr, K. Johnson, and J. C. Side. Planning at the edge: Integrating across the land sea divide. Marine Policy, 47(??):118–125, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000396.

Klein:2014:EIC

[KJWP14] Carissa J. Klein, Stacy D. Jupiter, Matthew Watts, and Hugh P. Possingham. Evaluating the influence of candidate terrestrial protected areas on coral reef condition in Fiji. Marine Policy, 44(??):360–365, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002261.

Knutsson:2016:EFM

[KKG16] Ögmundur Knútsson, Dadi Már Kristófersson, and Helgi Gestsson. The effects of fisheries management on the Icelandic demersal fish value chain. Marine Policy, 63 (??):172–179, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000640.

Kumar:2019:EFS

[KKSC19] Radika Kumar, Ronald Ravinesh Kumar, Peter Josef Stauvermann, and Jadhav Chakradhar. The effectiveness of fisheries subsidies as a trade policy tool to achieving sustainable development goals at the WTO. Marine Policy, 100(??):132–140, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304822.

King:2015:DKF

[KKWS15] Tanya King, Sue Kilpatrick, Karen Willis, and Christopher Speldewinde. “A different kettle of fish”: Mental health strategies for Australian fishers, and farmers. Marine Policy, 60(??):134–140, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001724.

Kaplan:2012:KCS

[KL12] Isaac C. Kaplan and Jerry Leonard. From krill to convenience stores: Forecasting the economic and ecological effects of fisheries management on the US West Coast. Marine Policy, 36(5):947–954, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000127.

Karper:2014:PCE

[KL14] Marjoleine A. M. Karper and Priscila F. M. Lopes. Punishment and compliance: Exploring scenarios to improve the legitimacy of small-scale fisheries management rules on the Brazilian coast. Marine Policy, 44(??):457–464, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002376.

Katsanevakis:2015:MCC

[KLC^{+}15] Stelios Katsanevakis, Noam Levin, Marta Coll, Sylvaine Giakoumi, Daniel Shkedi, Peter Mackelworth, Ran Levy, Adonis Velegrakis, Drosos Koutsoubas, Hrvoje Caric, Eran Brokovich, Bayram Öztürk, and Salit Kark. Marine conservation challenges in an era of economic crisis and geopolitical instability: the case of the Mediterranean Sea. Marine Policy, 51(??):31–39, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001870.

Kleih:2013:FSS

[KLM^{+}13] Ulrich Kleih, John Linton, Ana Marr, Murdoch Mactaggart, Diego Naziri, and John E. Orchard. Financial services for small and medium-scale aquaculture and fisheries producers. Marine Policy, 37(??):106–114, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200067X.

Kyriazi:2015:BNG

[KLMD15] Zacharoula Kyriazi, Raul Lejano, Frank Maes, and Steven Degraer. Bargaining a net gain compensation agreement between a marine renewable energy developer and a marine protected area manager. Marine Policy, 60(??):40–48, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001645.

Kotrikla:2017:AAP

[KLN17] Anna Maria Kotrikla, Theodoros Lilas, and Nikitas Nikitakos. Abatement of air pollution at an Aegean island port utilizing shore side electricity and renewable energy. Marine Policy, 75(??):238–248, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000397.

Kim:2014:EDF

[KLY14] Sang-Goo Kim, Won-IL Lee, and Moon Yuseok. The estimation of derelict fishing gear in the coastal waters of South Korea: Trap and gill-net fisheries. Marine Policy, 46(??):119–122, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000086.

Kildow:2010:IEC

[KM10] J. T. Kildow and A. McIlgorm. The importance of estimating the contribution of the oceans to national economies. Marine Policy, 34(3):367–374, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001225.

Kosenius:2015:EBC

[KM15] Anna-Kaisa Kosenius and Ollikainen Markku. Ecosystem benefits from coastal habitats — a three-country choice experiment. Marine Policy, 58(??):15–27, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000810.

Kafas:2017:SPM

[KMC^{+}17] Andronikos Kafas, Anne McLay, Marianna Chimienti, Beth E. Scott, Ian Davies, and Matthew Gubbins. ScotMap: Participatory mapping of inshore fishing activity to inform marine spatial planning in Scotland. Marine Policy, 79(??):8–18, May 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305097.

Katikiro:2015:CFL

[KMD15] Robert E. Katikiro, Edison D. Macusi, and K. H. M. Ashoka Deepananda. Challenges facing local communities in Tanzania in realising locally-managed marine areas. Marine Policy, 51(??):220–229, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002206.

Kamiyama:2018:DES

[KMF^{+}18] Ryutaro Kamiyama, Tsutom Miyata, Alice Joan G. Ferrer, Hisashi Kurokura, and Satoshi Ishikawa. Differences in the effects of social network, trust, and co-operation on fishery co-management. Marine Policy, 87(??):314–320, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730386X.

Kempf:2016:MCM

[KML^{+}16] Alexander Kempf, John Mumford, Polina Levontin, Adrian Leach, Ayoe Hoff, Katell G. Hamon, Heleen Bartelings, Morten Vinther, Moritz Stäbler, Jan Jaap Poos, Sophie Smout, Hans Frost, Sander van den Burg, Clara Ulrich, and Anna Rindorf. The MSY concept in a multiobjective fisheries environment — Lessons from the North Sea. Marine Policy, 69(??):146–158, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301841.

Kyriazi:2013:INC

[KMR^{+}13] Zacharoula Kyriazi, Frank Maes, Marijn Rabaut, Magda Vincx, and Steven Degraer. The integration of nature conservation into the marine spatial planning process. Marine Policy, 38(??):133–139, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001273.

Kopp:2016:SEI

[KMTS16] Dorothée Kopp, Robert Marianne, Chouvelon Tiphaine, and Méhault Sonia. Some expected impacts of the Common Fishery Policy on marine food webs. Marine Policy, 66(??):8–14, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000075.

Korbee:2015:ECC

[KMvT15] Dorien Korbee, Arthur P. J. Mol, and Jan P. M. van Tatenhove. Ecological considerations in constructing marine infrastructure: the Falmouth cruise terminal development, Jamaica. Marine Policy, 56(??):23–32, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000287.

Knott:2017:PFO

[KN17] Christine Knott and Barbara Neis. Privatization, financialization and ocean grabbing in New Brunswick herring fisheries and salmon aquaculture. Marine Policy, 80(??):10–18, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630690X.

Kalikoski:2010:BAC

[KNA10] Daniela C. Kalikoski, Pedro Quevedo Neto, and Tiago Almudi. Building adaptive capacity to climate variability: the case of artisanal fisheries in the estuary of the Patos Lagoon, Brazil. Marine Policy, 34(4):742–751, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000045X.

Knapp:2011:LPO

[Kna11] Gunnar Knapp. Local permit ownership in Alaska salmon fisheries. Marine Policy, 35(5):658–666, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000303.

Karstens:2018:FWN

[KNB^{+}18] Svenja Karstens, Cristina Nazzari, Carmen Bâlon, Malgorzata Bielecka, Žilvinas Grigaitis, Johanna Schumacher, Nardine Stybel, and Arturas Razinkovas-Baziukas. Floating wetlands for nutrient removal in eutrophicated coastal lagoons: Decision support for site selection and permit process. Marine Policy, 97(??):51–60, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305384.

Knol:2010:SAI

[Kno10] Maaike Knol. Scientific advice in integrated ocean management: the process towards the Barents Sea plan. Marine Policy, 34(2):252–260, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001018.

Knol:2011:UPZ

[Kno11] Maaike Knol. The uncertainties of precaution: Zero discharges in the Barents Sea. Marine Policy, 35(3):399–404, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10002058.

Knutson:2017:ECN

[Knu17] P. Knutson. Escaping the corporate net: Pragmatics of small boat direct marketing in the U.S. salmon fishing industry of the Northeastern Pacific. Marine Policy, 80 (??):123–129, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301221.

Kvalvik:2014:NSF

[KNY14] Ingrid Kvalvik, Bjoerg H. Noestvold, and James A. Young. National or supranational fisheries sustainability certification schemes? A critical analysis of Norwegian and Icelandic responses. Marine Policy, 46(??):137–142, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1400027X.

Kopke:2011:PKC

[KO11] Kathrin Kopke and Cathal O’Mahony. Preparedness of key coastal and marine sectors in Ireland to adapt to climate change. Marine Policy, 35(6):800–809, December 2011.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000091.

Kalfagianni:2013:FMW

[KP13] Agni Kalfagianni and Philipp Pattberg. Fishing in muddy waters: Exploring the conditions for effective governance of fisheries and aquaculture. Marine Policy, 38(??):124–132, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001261.

Kastrisios:2017:NCC

[KP17] Christos Kastrisios and Maria Pilikou. Nautical cartography competences and their effect to the realisation of a worldwide Electronic Navigational Charts database, the performance of ECDIS and the fulfilment of IMO chart carriage requirements. Marine Policy, 75(??):29–37, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304250.

Kao:2012:AAS

[KPF12] Shih-Ming Kao, Nathaniel S. Pearre, and Jeremy Firestone. Adoption of the Arctic search and rescue agreement: a shift of the Arctic regime toward a hard law basis? Marine Policy, 36(3):832–838, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001977.

Kelly:2017:SLM

[KPF17] Rachel Kelly, Gretta T. Pecl, and Aysha Fleming. Social licence in the marine sector: a review of understanding and application. Marine Policy, 81(??):21–28, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306558.

Koliousis:2017:SPF

[KPR^{+}17] Ioannis G. Koliousis, Stratos Papadimitriou, Elena Riza, Peter J. Stavroulakis, and Vangelis Tsioumas. Strategy, policy, and the formulation of maritime cluster typologies. Marine Policy, 86(??):31–38, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304190.

Koliousis:2018:STM

[KPR^{+}18] Ioannis G. Koliousis, Stratos Papadimitriou, Elena Riza, Peter J. Stavroulakis, and Vangelis Tsioumas. Scarcity theory and maritime clusters: From paradox to modelling. Marine Policy, 93(??):40–46, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308862.

Kincaid:2014:WFW

[KR14] Kate Barley Kincaid and George A. Rose. Why fishers want a closed area in their fishing grounds: Exploring perceptions and attitudes to sustainable fisheries and conservation 10 years post closure in Labrador, Canada. Marine Policy, 46(??):84–90, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000098.

Koivurova:2018:CVI

[KR18] Timo Koivurova and Allan Rosas. The CBSS as a vehicle for institutionalised governance in the Baltic Sea Area, in comparison with its two sister organisations in the north. Marine Policy, 98(??):211–219, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306778.

Koehn:2013:PPS

[KRK13] J. Zachary Koehn, Daniel R. Reineman, and John N. Kittinger. Progress and promise in spatial human dimensions research for ecosystem-based ocean planning. Marine Policy, 42(??):31–38, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000249.

Kincaid:2014:FPM

[KRM14] Kate Barley Kincaid, George Rose, and Humphrey Mahudi. Fishers’ perception of a multiple-use marine protected area: Why communities and gear users differ at Mafia Island, Tanzania. Marine Policy, 43(??):226–235, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001310.

Krupa:2016:WWK

[Kru16] Meagan B. Krupa. Who’s who in the Kenai River Fishery SES: a streamlined method for stakeholder identification and investment analysis. Marine Policy, 71(??): 194–200, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303384.

King:2010:RNL

[KS10] Dennis M. King and Jon G. Sutinen. Rational noncompliance and the liquidation of Northeast groundfish resources. Marine Policy, 34(1):7–21, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000529.

Kaiser:2012:MDC

[KS12a] Mark J. Kaiser and Brian Snyder. Modeling the decommissioning cost of offshore wind development on the U.S. Outer Continental Shelf. Marine Policy, 36(1):153–164, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000807.

Kaiser:2012:OWC

[KS12b] Mark J. Kaiser and Brian Snyder. Offshore wind capital cost estimation in the U.S. Outer Continental Shelf

— a reference class approach. Marine Policy, 36(5): 1112–1122, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000073.

Kaiser:2012:OWD

[KS12c] Mark J. Kaiser and Brian Snyder. Offshore wind decommissioning regulations and workflows in the Outer Continental Shelf United States. Marine Policy, 36(1):113–121, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000765.

Klaer:2012:DPC

[KS12d] Neil L. Klaer and David C. Smith. Determining primary and companion species in a multi-species fishery: Implications for TAC setting. Marine Policy, 36(3):606–612, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001540.

Kaiser:2013:FOD

[KS13] Mark J. Kaiser and Brian Snyder. The five offshore drilling rig markets. Marine Policy, 39(??):201–214, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002187.

Kaplan:2016:CBC

[KS16] Maxwell B. Kaplan and Susan Solomon. A coming boom in commercial shipping? The potential for rapid growth of noise from commercial ships by 2030. Marine Policy, 73(??):119–121, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303335.

Kaplan:2018:RRM

[KS18a] Maxwell B. Kaplan and Susan Solomon. Response to rejoinder to: M. B. Kaplan, S. Solomon, A coming boom in commercial shipping? The potential for rapid growth of noise from commercial ships by 2030, Mar. Policy 73 (Supplement C) (2016) 119–121. Marine Policy, 93(??):54, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830201X.

Kern:2018:EAM

[KS18b] Kristine Kern and Sara Söderström. The ecosystem approach to management in the Baltic Sea Region: Analyzing regional environmental governance from a spatial perspective. Marine Policy, 98(??):271–277, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306948.

Kim:2019:RSE

[KS19] A-Rom Kim and Young-Joon Seo. The reduction of SOx emissions in the shipping industry: the case of Korean companies. Marine Policy, 100(??):98–106, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307188.

Kahmann:2015:NJH

[KSB15] Birte Kahmann, Klara Helene Stumpf, and Stefan Baumgärtner. Notions of justice held by stakeholders of the Newfoundland fishery. Marine Policy, 62(??):37–50, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002316.

Kumawat:2015:CBN

[KSC^{+}15] Tarachand Kumawat, Latha Shenoy, Sushant K. Chakraborty, Vinay D. Deshmukh, and Sadashiv G. Raje. Compliance of bag net fishery of Maharashtra coast, India with Article 7 of the FAO Code of Conduct for Responsible Fisheries. Marine Policy, 56(??):9–15, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000263.

Kroetz:2019:EPA

[KSC^{+}19] Kailin Kroetz, James N. Sanchirico, Elsa Galarza Contreras, David Corderi Novoa, Nestor Collado, and Elaine W. Swiedler. Examination of the Peruvian Anchovy Individual Vessel Quota (IVQ) system. Marine Policy, 101(??):15–24, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308539.

Kuriyama:2015:IFL

[KSH^{+}15] Peter T. Kuriyama, Margaret C. Siple, Emma E. Hodgson, Elizabeth M. Phillips, Merrick Burden, David Fluharty, André E. Punt, Timothy E. Essington, John Henderschedt, and David A. Armstrong. Issues at the fore in the land of Magnuson and Stevens: a summary of the 14th Bevan Series on Sustainable Fisheries. Marine Policy, 54(??):118–121, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003492.

Kleitou:2019:ILM

[KSK^{+}19] Periklis Kleitou, Ioannis Savva, Demetris Kletou, Jason M. Hall-Spencer, Charalampos Antoniou, Yiannis Christodoulides, Niki Chartosia, Louis Hadjioannou, Andreas C. Dimitriou, Carlos Jimenez, Antonis Petrou, Spyros Sfenthourakis, and Sian Rees. Invasive lionfish in the Mediterranean: Low public awareness yet high stakeholder concerns. Marine Policy, 104(??):66–74, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307085.

Kim:2017:MRE

[KSS17] Do-Hoon Kim, Chang K. Seung, and Young-Il Seo. Multiregional economic impacts of recreational fisheries: Analysis of small sea ranch in Gyeong-Nam Province, Korea. Marine Policy, 84(??):90–98, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301938.

Keen:2018:TDB

[KSWS18] Meg R. Keen, Anne-Maree Schwarz, and Lysa Wini-Simeon. Towards defining the Blue Economy: Practical lessons from Pacific Ocean governance. Marine Policy, 88(??):333–341, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308235.

Klinger:2013:MBF

[KTA^{+}13] Dane H. Klinger, Mary Turnipseed, James L. Anderson, Frank Asche, Larry B. Crowder, Atle G. Guttormsen, Benjamin S. Halpern, Mary I. O’Connor, Raphael Sagarin, Kimberly A. Selkoe, Geoffrey G. Shester, Martin D. Smith, and Peter Tyedmers. Moving beyond the fished or farmed dichotomy. Marine Policy, 38(??):369–374, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001546.

Kontogianni:2015:MEJ

[KTD^{+}15] A. Kontogianni, C. Tourkolias, D. Damigos, M. Skourtos, and B. Zanou. Modeling expert judgment to assess costeffectiveness of EU Marine Strategy Framework Directive programs of measures. Marine Policy, 62(??):203–212, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002535.

Kuo:2018:ACI

[KV18] Ting-Chun Kuo and Amanda Vincent. Assessing the changes in international trade of marine fishes under CITES regulations — a case study of seahorses. Marine Policy, 88(??):48–57, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303767.

Kabbadj:2018:STI

[KVA18] Lina Kabbadj, Simon Van Wynsberge, and Serge Andréfouët. Scaling tropical island conservation planning to the regional level can lead to unbalanced ecological representation and poor social equity among islands. Marine Policy, 93(??):31–39, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730708X.

Kirby:2014:AEL

[KVH14] David Seán Kirby, Candice Visser, and Quentin Hanich. Assessment of eco-labelling schemes for Pacific tuna fisheries. Marine Policy, 43(??):132–142, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001085.

Kronen:2010:SED

[KVMM10] Mecki Kronen, Aliti Vunisea, Franck Magron, and Brian McArdle. Socio-economic drivers and indicators for artisanal coastal fisheries in Pacific island countries and territories and their use for fisheries management strategies. Marine Policy, 34(6):1135–1143, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000655.

Kirby:2014:SEM

[KW14] David Seán Kirby and Peter Ward. Standards for the effective management of fisheries bycatch. Marine Policy, 44(??):419–426, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002339.

Koslow:2016:ISD

[KW16] J. Anthony Koslow and Melaina Wright. Ichthyoplankton sampling design to monitor marine fish populations and communities. Marine Policy, 68(??):55–64, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000646.

Kopsel:2018:CLW

[KW18] Vera Köpsel and Cormac Walsh. “Coastal landscapes for whom? Adaptation challenges and landscape management in Cornwall”. Marine Policy, 97(??):278–286, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304032.

Krelling:2017:DPR

[KWT17] Allan Paul Krelling, Allan Thomas Williams, and Alexander Turra. Differences in perception and reaction of tourist groups to beach marine debris that can influence a loss of tourism revenue in coastal areas.

Marine Policy, 85(??):87–99, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301689.

Kaiser:2010:EFU

[KYS10] Mark J. Kaiser, Yunke Yu, and Brian Snyder. Economic feasibility of using offshore oil and gas structures in the Gulf of Mexico for platform-based aquaculture. Marine Policy, 34(3):699–707, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000035.

Langholz:2014:RPE

[LA14] Jeff A. Langholz and Adina Abeles. Rethinking postgraduate education for marine conservation. Marine Policy, 43 (??):372–375, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001401.

LeMeur:2018:DSM

[LACG18] Pierre-Yves Le Meur, Nicholas Arndt, Patrice Christmann, and Vincent Geronimi. Deep-sea mining prospects in French Polynesia: Governance and the politics of time. Marine Policy, 95(??):380–387, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304717.

Laffoley:2014:FOC

[Laf14] Dan Laffoley. ‘Future Ocean’ — Communicating and conserving the blue heart of the planet. Marine Policy, 45 (??):259–260, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002133.

LeManach:2013:WGW

[LAH^{+}13] Frédéric Le Manach, Mialy Andriamahefazafy, Sarah Harper, Alasdair Harris, Gilles Hosch, Glenn-Marie Lange, Dirk Zeller, and Ussif Rashid Sumaila. Who gets what? Developing a more equitable framework for EU fishing agreements. Marine Policy, 38(??):257–266, March 2013.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001406.

Lalancette:2017:CNI

[Lal17] Annie Lalancette. Creeping in? Neoliberalism, indigenous realities and tropical rock lobster (kaiar) management in Torres Strait, Australia. Marine Policy, 80(??):47–59, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300379.

Langdon:2015:FHN

[Lan15] Steve J. Langdon. Foregone harvests and neoliberal policies: Creating opportunities for rural, small-scale, community-based fisheries in southern Alaskan coastal villages. Marine Policy, 61(??):347–355, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000561.

Langlet:2018:SSD

[Lan18] David Langlet. Scale, space and delimitation in marine legal governance — Perspectives from the Baltic Sea. Marine Policy, 98(??):278–285, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830695X.

Leleu:2012:FPI

[LAP^{+}12] Kevin Leleu, Frédérique Alban, Dominique Pelletier, Eric Charbonnel, Yves Letourneur, and Charles F. Boudouresque. Fishers’ perceptions as indicators of the performance of Marine Protected Areas (MPAs). Marine Policy, 36(2):414–422, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X11001047.

Laxe:2010:DCF

[Lax10] Fernando González Laxe. Dysfunctions in common fishing regulations. Marine Policy, 34(1):182–188, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000943.

Levy:2013:MIC

[LB13] Jessica S. Levy and Natalie C. Ban. A method for incorporating climate change modelling into marine conservation planning: an Indo-west Pacific example. Marine Policy, 38(??):16–24, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001030.

Leitao:2017:DBP

[LB17] Francisco Leitão and Vânia Baptista. The discard ban policy, economic trends and opportunities for the Portuguese fisheries sector. Marine Policy, 75(??):75–83, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304274.

Liu:2018:CCP

[LB18] Nengye Liu and Cassandra M. Brooks. China’s changing position towards marine protected areas in the Southern Ocean: Implications for future Antarctic governance. Marine Policy, 94(??):189–195, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308977.

Little:2012:WWP

[LBB^{+}12] David C. Little, Simon R. Bush, Ben Belton, Nguyen Thanh Phuong, James A. Young, and Francis J. Murray. Whitefish wars: Pangasius, politics and consumer confusion in Europe. Marine Policy, 36(3):738–745, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001564.

Lyons:2016:CCF

[LBC^{+}16] Courtney Lyons, Benjamin Blount, Courtney Carothers, Meredith Marchioni, Reade Davis, and Philip Loring. Considering communities in fisheries management.

Marine Policy, 74(??):288–291, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302809.

Lucchetti:2018:SPP

[LBD^{+}18] Alessandro Lucchetti, Andrea Belardinelli, Lorenzo D’Andrea, Bojan Marčeta, Michela Martinelli, Tommaso Russo, Vanja Čikeš Keč, Barbara Zorica, and Massimo Virgili. Small pelagic purse seines in the Adriatic Sea: a spatial analysis and technical overview in relation to Mediterranean Regulation provisions. Marine Policy, 98(??): 104–114, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301210.

Lima:2016:FKI

[LBHS16] Evelyn G. Lima, Alpina Begossi, Gustavo Hallwass, and Renato A. M. Silvano. Fishers’ knowledge indicates short-term temporal changes in the amount and composition of catches in the southwestern Atlantic. Marine Policy, 71(??):111–120, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630286X.

Leopold:2013:CBM

[LBK^{+}13] Marc Léopold, Jennifer Beckensteiner, Jeremie Kaltavara, Jason Raubani, and Sophie Caillon. Community-based management of near-shore fisheries in Vanuatu: What works? Marine Policy, 42(??):167–176, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000560.

Levontin:2017:RVF

[LBL^{+}17] Polina Levontin, Paul Baranowski, Adrian W. Leach, Alexandria Bailey, John D. Mumford, Antoni Quetglas, and Laurence T. Kell. On the role of visualisation in fisheries management. Marine Policy, 78(??):114–121, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630793X.

Lankester:2015:MRR

[LBN15] Ally J. Lankester, Erin Bohensky, and Maxine Newlands. Media representations of risk: the reporting of dredge spoil disposal in the Great Barrier Reef Marine Park at Abbot Point. Marine Policy, 60(??):149–161, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001682.

Lomboy:2019:BHE

[LBP^{+}19] Cristopher G. Lomboy, Flora Belinario, Robert Pomeroy, Joey Pedrajas, Raquel Sanchez Tirona, Steve Box, Paolo Roberto Domondon, and Kristine Balbido-Ramirez. Building household economic resilience to secure a future for near shore fishers in the Philippines. Marine Policy, 99(??):334–342, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18305050.

LeManach:2019:PSS

[LBVN19] Frédéric Le Manach, Laetitia Bisiaux, Sebastian Villasante, and Claire Nouvian. Public subsidies have supported the development of electric trawling in Europe. Marine Policy, 104(??):225–231, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308819.

Liu:2019:CLK

[LC19] Tuz-Ming Liu and Shui-Kai Chang. Changes in local knowledge and its impacts on ecological resources management: the case of flyingfish culture of the Tao in Taiwan. Marine Policy, 103(??):74–83, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306456.

Liu:2014:MSP

[LCC14] Ta-Kang Liu, Chi-Hsin Chang, and Ming-Li Chou. Management strategies to prevent the introduction of nonindigenous aquatic species in response to the Ballast Water Convention in Taiwan. Marine Policy, 44(??):

187–195, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001772.

Lloret:2018:SSC

[LCC^{+}18] Josep Lloret, Ian G. Cowx, Henrique Cabral, Margarida Castro, Toni Font, Jorge M. S. Gonçalves, Ana Gordoa, Ellen Hoefnagel, Sanja Matić-Skoko, Eirik Mikkelsen, Beatriz Morales-Nin, Dimitrios K. Moutopoulos, Marta Muñoz, Miguel Neves dos Santos, Pedro Pintassilgo, Cristina Pita, Konstantinos I. Stergiou, Vahdet Ünal, Pedro Veiga, and Karim Erzini. Small-scale coastal fisheries in European Seas are not what they were: Ecological, social and economic changes. Marine Policy, 98(??):176–186, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302482.

Lyons:2019:ACD

[LCC19] Courtney Lyons, Courtney Carothers, and Jesse Coleman. Alaska’s community development quota program: a complex institution affecting rural communities in disparate ways. Marine Policy, 108(??):Article 103560, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18308789.

Lester:2013:ETA

[LCH^{+}13] Sarah E. Lester, Christopher Costello, Benjamin S. Halpern, Steven D. Gaines, Crow White, and John A. Barth. Evaluating tradeoffs among ecosystem services to inform marine spatial planning. Marine Policy, 38 (??):80–89, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001200.

Lyons:2016:MMC

[LCR16] Courtney Lyons, Courtney Carothers, and Katherine Reedy. Means, meanings, and contexts: a framework for integrating detailed ethnographic data into assessments of fishing community vulnerability. Marine Policy, 74(??):341–350, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301944.

Long:2015:KPM

[LCS15] Rachel D. Long, Anthony Charles, and Robert L. Stephenson. Key principles of marine ecosystem-based management. Marine Policy, 57(??):53–60, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500024X.

LeCornu:2018:SMS

[LDF^{+}18] Elodie Le Cornu, Angee N. Doerr, Elena M. Finkbeiner, Don Gourlie, and Larry B. Crowder. Spatial management in small-scale fisheries: a potential approach for climate change adaptation in Pacific Islands. Marine Policy, 88(??):350–358, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301690.

Langton:2011:SCT

[LDS11] R. Langton, I. M. Davies, and B. E. Scott. Seabird conservation and tidal stream and wave power generation: Information needs for predicting and managing potential impacts. Marine Policy, 35(5):623–630, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000261.

Lescrauwaet:2010:FPH

[LDVM10] Ann-Katrien Lescrauwaet, Heidi Debergh, Magda Vincx, and Jan Mees. Fishing in the past: Historical data on sea fisheries landings in Belgium. Marine Policy, 34(6): 1279–1289, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001041.

Lidskog:2012:SBS

[LE12] Rolf Lidskog and Ingemar Elander. Sweden and the Baltic Sea pipeline: Between ecology and economy. Marine Policy, 36(2):333–338, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001084.

Leary:2019:ADW

[Lea19] David Leary. Agreeing to disagree on what we have or have not agreed on: the current state of play of the BBNJ negotiations on the status of marine genetic resources in areas beyond national jurisdiction. Marine Policy, 99(??):21–29, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305499.

Lee:2010:ETG

[Lee10] Min-Yang Lee. Economic tradeoffs in the Gulf of Maine ecosystem: Herring and whale-watching. Marine Policy, 34(1):156–162, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000906.

Lucrezi:2019:ESE

[LEFC19] Serena Lucrezi, Minoo Hassani Esfehani, Eliana Ferretti, and Carlo Cerrano. The effects of stakeholder education and capacity building in marine protected areas: a case study from southern Mozambique. Marine Policy, 108(??):Article 103645, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309710.

Lucrezi:2019:TCM

[LEG19] Serena Lucrezi, Suria Ellis, and Enrico Gennari. A test of causative and moderator effects in human perceptions of sharks, their control and framing. Marine Policy, 109(??):Article 103687, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302398.

Lee:2010:SEM

[LEKB10] Dae-In Lee, Ki-Hyuk Eom, Gui-Young Kim, and Gun-Wook Baeck. Scoping the effective marine environmental assessment of dredging and ocean disposal of coastal sediments in Korea. Marine Policy, 34(5):1082–1092, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000606.

Levin:2018:BEF

[LEM^{+}18] Phillip S. Levin, Timothy E. Essington, Kristin N. Marshall, Laura E. Koehn, Lee G. Anderson, Alida Bundy, Courtney Carothers, Felicia Coleman, Leah R. Gerber, Jonathan H. Grabowski, Edward Houde, Olaf P. Jensen, Christian Möllmann, Kenneth Rose, James N. Sanchirico, and Anthony D. M. Smith. Building effective fishery ecosystem plans. Marine Policy, 92(??):48–57, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306954.

Lewis:2016:MPC

[Lew16] Justin S. Lewis. Maritime piracy confrontations across the globe: Can crew action shape the outcomes? Marine Policy, 64(??):116–122, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003358.

Li:2017:MPA

[LF17] Yunzhou Li and David L. Fluharty. Marine protected area networks in China: Challenges and prospects. Marine Policy, 85(??):8–16, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301987.

Little:2015:BMC

[LFGM15] Cedric Little, Christian Felzensztein, Eli Gimmon, and Pablo Muñoz. The business management of the Chilean salmon farming industry. Marine Policy, 54(??):108–117, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14003571.

Lensu:2019:BMD

[LG19] Mikko Lensu and Floris Goerlandt. Big maritime data for the Baltic Sea with a focus on the winter navigation system. Marine Policy, 104(??):53–65, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305335.

Leroy:2016:ERT

[LGC16] Antonia Leroy, Florence Galletti, and Christian Chaboud. The EU restrictive trade measures against IUU fishing. Marine Policy, 64(??):82–90, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002973.

Lemieux:2019:HRA

[LGD^{+}19] Christopher J. Lemieux, Paul A. Gray, Rodolphe Devillers, Pamela A. Wright, Philip Dearden, Elizabeth A. Halpenny, Mark Groulx, Thomas J. Beechey, and Karen Beazley. How the race to achieve Aichi Target 11 could jeopardize the effective conservation of biodiversity in Canada and beyond. Marine Policy, 99(??):312–323, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304706.

Loerzel:2017:SDA

[LGDB17] Jarrod L. Loerzel, Theresa L. Goedeke, Maria K. Dillard, and Greg Brown. SCUBA divers above the waterline: Using participatory mapping of coral reef conditions to inform reef management. Marine Policy, 76(??):79–89, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300501.

Lemee:2019:WMD

[LGFB^{+}19] C. Lemée, M. Guillard, G. Fleury-Bahi, N. Krien, C. Chadenas, E. Chauveau, M. Desse, M. Coquet, M. Lamarre, and O. Navarro. What meaning do individuals give to coastal risks? Contribution of the social representation theory. Marine Policy, 108(??):Article 103629, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18305360.

LeManach:2012:EUF

[LGH^{+}12a] Frédéric Le Manach, Charlotte Gough, Alasdair Harris, Frances Humber, Sarah Harper, and Dirk Zeller. Erratum to “Unreported fishing, hungry people and political turmoil: the recipe for a food security crisis in Madagascar” [Mar. Policy 36 (1) (2012) 218– 225]. Marine Policy, 36(2):564, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001436.

LeManach:2012:UFH

[LGH^{+}12b] Frédéric Le Manach, Charlotte Gough, Alasdair Harris, Frances Humber, Sarah Harper, and Dirk Zeller. Unreported fishing, hungry people and political turmoil: the recipe for a food security crisis in Madagascar? Marine Policy, 36(1):218–225, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000960.

Lowe:2014:HEI

[LGM14] Christopher D. Lowe, Alison J. Gilbert, and Laurence D. Mee. Human–environment interaction in the Baltic Sea. Marine Policy, 43(??):46–54, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000821.

Lees:2016:UKD

[LGM16] Kirsty J. Lees, Andrew J. Guerin, and Elizabeth A. Masden. Using kernel density estimation to explore habitat use by seabirds at a marine renewable wave energy test facility. Marine Policy, 63(??):35–44, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002845.

Lin:2012:CBE

[LH12] Cherng-Yuan Lin and Tsan-Huang Huang. Cost–benefit evaluation of using biodiesel as an alternative fuel for fishing boats in Taiwan. Marine Policy, 36(1):103–107, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000741.

Liao:2016:CSF

[LH16a] Chun-Pei Liao and Hsiang-Wen Huang. The cooperation strategies of fisheries between Taiwanese purse seiners and Pacific Island Countries. Marine Policy, 66(??):67–74, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000233.

Lozano:2016:PRC

[LH16b] Alejandro J. García Lozano and Joel T. Heinen. Property relations and the co-management of small-scale fisheries in Costa Rica: Lessons from marine areas for responsible fishing in the Gulf of Nicoya. Marine Policy, 73(??):196–203, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305279.

Lynham:2017:CSP

[LHB^{+}17] J. Lynham, B. S. Halpern, T. Blenckner, T. Essington, J. Estes, M. Hunsicker, C. Kappel, A. K. Salomon, C. Scarborough, K. A. Selkoe, and A. Stier. Costly stakeholder participation creates inertia in marine ecosystems. Marine Policy, 76(??):122–129, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307199.

Liu:2015:SCE

[LHH15] Ta-Kang Liu, Hsiao-Yin Huang, and Shao-Liang Hsu. Saving the critically endangered Chinese white dolphin in Taiwan: Debate regarding the designation of an MPA. Marine Policy, 61(??):113–120, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001888.

Lee:2015:FSS

[LHJ^{+}15] Jongmyoung Lee, Sunwook Hong, Yong Chang Jang, Mi Jeong Lee, Daeseok Kang, and Won Joon Shim. Finding solutions for the styrofoam buoy debris problem through participatory workshops. Marine Policy, 51 (??):182–189, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002243.

Lim:2018:MSC

[LHN18] Kar H. Lim, Wuyang Hu, and Rodolfo M. Nayga. Is Marine Stewardship Council’s ecolabel a rising tide for all? Consumers’ willingness to pay for origin-differentiated ecolabeled canned tuna. Marine Policy, 96(??):18–26, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306991.

Lily:2016:RDS

[Lil16] Hannah Lily. A regional deep-sea minerals treaty for the Pacific Islands? Marine Policy, 70(??):220–226, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300896.

Lim:2016:VCU

[Lim16] Guanie Lim. Value chain upgrading: Evidence from the Singaporean aquaculture industry. Marine Policy, 63 (??):191–197, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000652.

Lin:2013:SPB

[Lin13] Cherng-Yuan Lin. Strategies for promoting biodiesel use in marine vessels. Marine Policy, 40(??):84–90, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000109.

Liu:2013:MOC

[Liu13] Wen-Hong Liu. Managing the offshore and coastal fisheries in Taiwan to achieve sustainable development using policy indicators. Marine Policy, 39(??):162–171, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002205.

Livernois:2010:EEC

[Liv10] John Livernois. The economics of ending Canada’s commercial harp seal hunt. Marine Policy, 34(1):42–53, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000578.

Linke:2013:CTS

[LJ13] Sebastian Linke and Svein Jentoft. A communicative turnaround: Shifting the burden of proof in European fisheries governance. Marine Policy, 38(??):337–345, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001509.

Linke:2014:EPD

[LJ14] Sebastian Linke and Svein Jentoft. Exploring the phronetic dimension of stakeholders’ knowledge in EU fisheries governance. Marine Policy, 47(??):153–161, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002352.

Lieberknecht:2016:SSD

[LJ16] Louise M. Lieberknecht and Peter J. S. Jones. From stormy seas to the doldrums: the challenges of navigating towards an ecologically coherent marine protected area network through England’s Marine Conservation Zone process. Marine Policy, 71(??):275–284, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302172.

Lobo:2017:SCD

[LJ17] Rafaella Lobo and Peter J. Jacques. SOFIA’S choices: Discourses, values, and norms of the World Ocean Regime. Marine Policy, 78(??):26–33, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302111.

Levrel:2014:MCM

[LJB^{+}14] Harold Levrel, Céline Jacob, Denis Bailly, Mahe Charles, Olivier Guyader, Schéhérazade Aoubid, Adeline Bas, Alexia Cujus, Marjolaine Frésard, Sophie Girard, Julien Hay, Yann Laurans, Jérôme Paillet, José A. Pérez Agúndez, and Rémi Mongruel. The maintenance costs of marine natural capital: a case study from the initial assessment of the Marine Strategy Framework Directive in France. Marine Policy, 49(??):37–47, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001079.

Lodge:2014:SMI

[LJL^{+}14] Michael Lodge, David Johnson, Gwenaëlle Le Gurun, Markus Wengler, Phil Weaver, and Vikki Gunn. Seabed mining: International Seabed Authority environmental management plan for the Clarion–Clipperton Zone. A partnership approach. Marine Policy, 49(??):66–72, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14001171.

Lembo:2018:ASE

[LJS^{+}18] Giuseppe Lembo, Alfred Jokumsen, Maria Teresa Spedicato, Maria Teresa Facchini, and Isabella Bitetto. Assessing stakeholder’s experience and sensitivity on key issues for the economic growth of organic aquaculture production. Marine Policy, 87(??):84–93, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304116.

Luthman:2019:GSF

[LJT19] O. Luthman, M. Jonell, and M. Troell. Governing the salmon farming industry: Comparison between national regulations and the ASC salmon standard. Marine Policy, 106(??):Article 103534, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830842X.

Lee:2015:BVN

[LK15] Taedong Lee and Hyun Jung Kim. Barriers of voyaging on the Northern Sea Route: a perspective from shipping Companies. Marine Policy, 62(??):264–270, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002572.

Liu:2015:TUC

[LKC15] Ta-Kang Liu, Jui-Chuang Kao, and Ping Chen. Tragedy of the unwanted commons: Governing the marine debris in Taiwan’s oyster farming. Marine Policy, 53(??):123–130, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14003285.

Lee:2015:EEV

[LKL15] Gunwoo Lee, Soo-Yeob Kim, and Min-Kyu Lee. Economic evaluation of vessel traffic service (VTS): a contingent valuation study. Marine Policy, 61(??):149–154, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002304.

Lucchetti:2016:TSF

[LKMN16] A. Lucchetti, S. E. A. Kholeif, H. H. Mahmoud, and E. Notti. Towards sustainable fisheries management in emerging markets: an overview of properties, gaps and opportunities in Egypt. Marine Policy, 72(??):1–10, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303244.

Lew:2015:SPS

[LL15] Daniel K. Lew and Douglas M. Larson. Stated preferences for size and bag limits of Alaska charter boat anglers. Marine Policy, 61(??):66–76, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002080.

LeHeron:2019:DCP

[LLA^{+}19] Erena Le Heron, June Logie, Will Allen, Richard Le Heron, Paula Blackett, Kate Davies, Alison Greenaway, Bruce Glavovic, and Daniel Hikuroa. Diversity, contestation, participation in Aotearoa New Zealand’s multi-use/user marine spaces. Marine Policy, 106(??):Article 103536, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309680.

LeHeron:2016:NSS

[LLF^{+}16] Richard Le Heron, Nick Lewis, Karen Fisher, Simon Thrush, Carolyn Lundquist, Judi Hewitt, and Joanne Ellis. Non-sectarian scenario experiments in socioecological knowledge building for multi-use marine environments: Insights from New Zealand’s Marine Futures project. Marine Policy, 67(??):10–21, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1600035X.

Leach:2014:IPU

[LLH^{+}14] Adrian W. Leach, Polina Levontin, Johnson Holt, Laurence T. Kell, and John D. Mumford. Identification and prioritization of uncertainties for management of Eastern Atlantic bluefin tuna (Thunnus thynnus). Marine Policy, 48(??):84–92, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000827.

Levkoe:2017:FFE

[LLN17] Charles Z. Levkoe, Kristen Lowitt, and Connie Nelson. “Fish as food”: Exploring a food sovereignty approach to small-scale fisheries. Marine Policy, 85(??):65–70, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303317.

Loizidou:2016:MSF

[LLO16] Xenia I. Loizidou, Michael I. Loizides, and Demetra L. Orthodoxou. Marine strategy framework directive: Defining joint monitoring opportunities for the Eastern Mediterranean and the Black Sea, through dedicated decision making workshops and innovative policy tools. Marine Policy, 67(??):76–82, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000518.

Loizidou:2017:MSF

[LLO17] Xenia I. Loizidou, Michael I. Loizides, and Demetra L. Orthodoxou. Marine Strategy Framework Directive: Innovative and participatory decision-making method for the identification of common measures in the Mediterranean. Marine Policy, 84(??):82–89, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300131.

Lauer:2015:LSA

[LLS^{+}15] Peter Lauer, Lambertus López, Emmanuelle Sloan, Sean Sloan, and Mehdi Doroudi. Learning from the systematic approach to aquaculture zoning in South Australia: a case study of aquaculture (Zones — Lower Eyre Peninsula) Policy 2013. Marine Policy, 59(??):77–84, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001013.

Lu:2015:CMS

[LLX^{+}15] Wen-Hai Lu, Jie Liu, Xian-Quan Xiang, Wei-Ling Song, and Alistair McIlgorm. A comparison of marine spatial planning approaches in China: Marine functional zoning and the marine ecological red line. Marine Policy, 62(??):94–101, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002559.

Liu:2019:OCS

[LLZ^{+}19] Mingming Liu, Mingli Lin, Peijun Zhang, Tianfei Xue, and Songhai Li. An overview of cetacean stranding around Hainan Island in the South China Sea, 1978– 2016: Implications for research, conservation and management. Marine Policy, 101(??):147–153, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300198.

Lovecraft:2011:HDM

[LM11] Amy Lauren Lovecraft and Chanda L. Meek. The human dimensions of marine mammal management in a time of rapid change: comparing policies in Canada, Finland and the United States. Marine Policy, 35(4):427–429, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10002095.

Lee:2014:NCA

[LM14] Sang-Go Lee and Amaj Rahimi Midani. National comprehensive approaches for rebuilding fisheries in South Korea. Marine Policy, 45(??):156–162, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002935.

Lee:2015:FSG

[LM15]  Sang-Go Lee and Amaj Rahimi Midani.  Fishery

self-governance in fishing communities of South Korea. Marine Policy, 53(??):27–32, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003029.

Lallier:2016:EIA

[LM16] Laura E. Lallier and Frank Maes. Environmental impact assessment procedure for deep seabed mining in the area: Independent expert review and public participation. Marine Policy, 70(??):212–219, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300768.

Leroy:2018:IDM

[LM18] Antonia Leroy and Michel Morin. Innovation in the decision-making process of the RFMOs. Marine Policy, 97 (??):156–162, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308369.

LeFloch:2015:RMF

[LMA^{+}15] Pascal Le Floc’h, Arantza Murillas, Martin Aranda, Fabienne Daurès, Mike Fitzpatrick, Olivier Guyader, Aaron Hatcher, Claire Macher, and Paul Marchal. The regional management of fisheries in European Western Waters. Marine Policy, 51(??):375–384, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002553.

Levin:2016:DSH

[LMG^{+}16] Lisa A. Levin, Kathryn Mengerink, Kristina M. Gjerde, Ashley A. Rowden, Cindy Lee Van Dover, Malcolm R. Clark, Eva Ramirez-Llodra, Bronwen Currie, Craig R. Smith, Kirk N. Sato, Natalya Gallo, Andrew K. Sweetman, Hannah Lily, Claire W. Armstrong, and Joseph Brider. Defining “serious harm” to the marine environment in the context of deep-seabed mining. Marine Policy, 74(??):245–259, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630495X.

Leisher:2012:MBC

[LMH^{+}12] Craig Leisher, Sangeeta Mangubhai, Sebastiaan Hess, Hesti Widodo, Tri Soekirman, Salomina Tjoe, Stevanus Wawiyai, S. Neil Larsen, Lukas Rumetna, A. Halim, and M. Sanjayan. Measuring the benefits and costs of community education and outreach in marine protected areas. Marine Policy, 36(5):1005–1011, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000395.

Lescroel:2016:SOT

[LMP^{+}16] Amélie Lescroël, Raphaël Mathevet, Clara Péron, Matthieu Authier, Pascal Provost, Akinori Takahashi, and David Grémillet. Seeing the ocean through the eyes of seabirds: a new path for marine conservation? Marine Policy, 68(??):212–220, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000671.

Lehtiniemi:2015:DTM

[LOD^{+}15] Maiju Lehtiniemi, Henn Ojaveer, Matej David, Bella Galil, Stephan Gollasch, Cynthia McKenzie, Dan Minchin, Anna Occhipinti-Ambrogi, Sergej Olenin, and Judith Pederson. Dose of truth — Monitoring marine non-indigenous species to serve legislative requirements. Marine Policy, 54(??):26–35, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003522.

Lonergan:2011:PBR

[Lon11] Mike Lonergan. Potential biological removal and other currently used management rules for marine mammal populations: a comparison. Marine Policy, 35(5):584– 589, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000212.

Lonergan:2012:TMR

[Lon12] Mike Lonergan. The targets of management rules for marine mammal populations require justification: a reply to Cooke et al. (2012). Marine Policy, 36(5):1188– 1190, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000176.

Liu:2011:WFS

[LOS11] Yajie Liu, Jon Olaf Olaussen, and Anders Skonhoft. Wild and farmed salmon in Norway — a review. Marine Policy, 35(3):413–418, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002125.

Leite:2016:RPF

[LP16] Laura Leite and Cristina Pita. Review of participatory fisheries management arrangements in the European Union. Marine Policy, 74(??):268–278, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304742.

Lange:2018:GCM

[LPC18] Marcus Lange, Glenn Page, and Valerie Cummins. Governance challenges of marine renewable energy developments in the U.S. — Creating the enabling conditions for successful project development. Marine Policy, 90(??):37–46, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307029.

Lucchetti:2014:TFC

[LPM^{+}14] Alessandro Lucchetti, Corrado Piccinetti, Uriano Meconi, Cristina Frittelloni, Mara Marchesan, Silvia Palladino, and Massimo Virgili. Transferable Fishing Concessions (TFC): a pilot study on the applicability in the Mediterranean Sea. Marine Policy, 44(??):438–447, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002340.

Lee:2017:CLS

[LPP17] Seokwoo Lee, Young Kil Park, and Hansan Park. The complex legal status of the Current Fishing Pattern Zone in the East China Sea. Marine Policy, 81(??):219–228, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300409.

Lillebo:2017:HCM

[LPR^{+}17] A. I. Lillebø, C. Pita, J. Garcia Rodrigues, S. Ramos, and S. Villasante. How can marine ecosystem services support the Blue Growth agenda? Marine Policy, 81(??):132–142, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16308107.

Levrel:2012:CMM

[LPS12] Harold Levrel, Sylvain Pioch, and Richard Spieler. Compensatory mitigation in marine ecosystems: Which indicators for assessing the “no net loss” goal of ecosystem services and ecological functions? Marine Policy, 36(6): 1202–1210, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000486.

Leroy:2016:IBA

[LPU^{+}16] Bruno Leroy, Thomas Peatman, Thomas Usu, Sylvain Caillot, Brad Moore, Ashley Williams, and Simon Nicol. Interactions between artisanal and industrial tuna fisheries: Insights from a decade of tagging experiments. Marine Policy, 65(??):11–19, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003644.

Leal:2010:WFA

[LQC10] Carlos P. Leal, Renato A. Quiñones, and Carlos Chávez. What factors affect the decision making process when setting TACs?: the case of Chilean fisheries. Marine Policy, 34(6):1183–1195, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000710.

Levine:2015:UCP

[LR15] Arielle Levine and Laurie Richmond. Using commonpool resource design principles to assess the viability of community-based fisheries co-management systems in American Samoa and Hawai’i. Marine Policy, 62(??): 9–17, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002389.

Lee:2019:DIG

[LRJ^{+}19] Lynn Chi Lee, Mike Reid, Russ Jones, Janet Winbourne, Murray Rutherford, and Anne Katherine Salomon. Drawing on indigenous governance and stewardship to build resilient coastal fisheries: People and abalone along Canada’s northwest coast. Marine Policy, 109(??):Article 103701, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301204.

Lester:2017:ESP

[LRMM17] Sarah E. Lester, Elizabeth O. Ruff, Kevin Mayall, and Jennifer McHenry. Exploring stakeholder perceptions of marine management in Bermuda. Marine Policy, 84(??): 235–243, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302683.

Lopes:2013:SFT

[LRS^{+}13] P. F. M. Lopes, E. M. Rosa, S. Salyvonchyk, V. Nora, and A. Begossi. Suggestions for fixing top-down coastal fisheries management through participatory approaches. Marine Policy, 40(??):100–110, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X13000043.

Lundmark:2019:MEK

[LSAL19] Carina Lundmark, Annica Sandström, Klas Andersson, and Linda Laikre. Monitoring the effects of knowledge communication on conservation managers’ perception of genetic biodiversity — a case study from the Baltic Sea. Marine Policy, 99(??):223–229, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304998.

Lu:2014:ZSM

[LSC14] Shiau-Yun Lu, Cheng-Han Shen, and Wen-Yan Chiau. Zoning strategies for marine protected areas in Taiwan: Case study of Gueishan Island in Yilan County, Taiwan. Marine Policy, 48(??):21–29, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000724.

Lajus:2018:IMS

[LSK18] D. Lajus, D. Stogova, and E. C. H. Keskitalo. The implementation of Marine Stewardship Council (MSC) certification in Russia: Achievements and considerations. Marine Policy, 90(??):105–114, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307728.

Linnane:2010:IUE

[LSMW10] Adrian Linnane, Sean Sloan, Richard McGarvey, and Tim Ward. Impacts of unconstrained effort: Lessons from a rock lobster (Jasus edwardsii) fishery decline in the northern zone management region of South Australia. Marine Policy, 34(5):844–850, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000072.

Leenhardt:2017:ESE

[LSP^{+}17] Pierre Leenhardt, Vanessa Stelzenmüller, Nicolas Pascal, Wolfgang Nikolaus Probst, Annie Aubanel, Tamatoa Bambridge, Mahé Charles, Eric Clua, François Féral, Bran Quinquis, Bernard Salvat, and Joachim Claudet. Exploring social-ecological dynamics of a coral reef resource system using participatory modeling and empirical data. Marine Policy, 78(??):90–97, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306108.

Ledee:2012:RAS

[LSTD12] Elodie J. I. Lédée, Stephen G. Sutton, Renae C. Tobin, and Debora M. De Freitas. Responses and adaptation strategies of commercial and charter fishers to zoning changes in the Great Barrier Reef Marine Park. Marine Policy, 36(1):226–234, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000984.

Larkin:2017:SCC

[LSW17] Alice Larkin, Tristan Smith, and Paul Wrobel. Shipping in changing climates. Marine Policy, 75(??):188–190, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303207.

Lindley:2017:OCI

[LT17] Jade Lindley and Erika J. Techera. Overcoming complexity in illegal, unregulated and unreported fishing to achieve effective regulatory pluralism. Marine Policy, 81(??):71–79, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307035.

Lukic:2019:FDC

[LT19] Anamarija Lukić and Željka Tonković. Fisheries-dependent community — Obstacle or opportunity for further development? A case study of Kali in Croatia. Marine Policy, 100(??):116–121, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303713.

Lowe:2019:LLI

[LTM19] Judi Lowe, Johann Friedrich C. Tejada, and Mark G. Meekan. Linking livelihoods to improved biodiversity conservation through sustainable integrated coastal management and community based dive tourism : Oslob Whale Sharks. Marine Policy, 108(??):Article 103630, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307589.

LoBue:2013:POU

[LU13] Carl LoBue and Jay Udelhoven. Private ownership of underwater lands in Great South Bay, New York: a case study in degradation, restoration and protection. Marine Policy, 41(??):103–109, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200262X.

Lopes:2018:PPS

[LV18] Priscila F. M. Lopes and Sebastián Villasante. Paying the price to solve fisheries conflicts in Brazil’s Marine Protected Areas. Marine Policy, 93(??):1–8, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306577.

Laso:2018:RLM

[LVRM^{+}18] Jara Laso, Ian Vázquez-Rowe, María Margallo, Ángel Irabien, and Rubén Aldaco. Revisiting the LCA+DEA method in fishing fleets. How should we be measuring efficiency? Marine Policy, 91(??):34–40, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304700.

Liu:2011:RLG

[LWJH11] Wen-Hong Liu, Chin-Cheng Wu, Hao-Tang Jhan, and Ching-Hsien Ho. The role of local Government in marine spatial planning and management in Taiwan. Marine Policy, 35(2):105–115, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001466.

Liu:2019:IBW

[LWS19] Ta-Kang Liu, Yu-Cheng Wang, and Ping-Hsuan Su. Implementing the ballast water management convention: Taiwan’s experience and challenges in the early stage. Marine Policy, 109(??):Article 103706, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305663.

Li:2016:HCS

[LWvdBL16] Ruiqian Li, Johan Woltjer, Margo van den Brink, and Yongfu Li. How coastal strategic planning reflects interrelationships between ecosystem services: a four-step method. Marine Policy, 70(??):114–127, August 2016.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300264.

Liu:2017:RDC

[LXWX17] Baiqiong Liu, Min Xu, Jing Wang, and Sumei Xie. Regional disparities in China’s marine economy. Marine Policy, 82(??):1–7, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300945.

Lee:2014:RCF

[LY14] Min-Kyu Lee and Seung-Hoon Yoo. The role of the capture fisheries and aquaculture sectors in the Korean national economy: an input–output analysis. Marine Policy, 44(??):448–456, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300239X.

Lu:2017:FCE

[LYB17] Yu-Heng Lu, Nobuyuki Yagi, and Robert Blasiak. Factors contributing to effective management in the Sakuraebi (Sergia lucens) fishery of Donggang, Taiwan. Marine Policy, 86(??):72–81, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300957.

Lynham:2014:HCS

[Lyn14] John Lynham. How have catch shares been allocated? Marine Policy, 44(??):42–48, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001619.

Lyons:2014:PSI

[Lyo14] Youna Lyons. Prospects for satellite imagery of insular features and surrounding marine habitats in the South China Sea. Marine Policy, 45(??):146–155, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002911.

Lu:2016:CSL

[LZ16] Bingbin Lu and Shichao Zhou. China’s state-led working model on protection of underwater cultural heritage: Practice, challenges, and possible solutions. Marine Policy, 65(??):39–47, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003668.

Lin:2019:MCS

[LZL19] Xianghong Lin, Li Zheng, and Weiwen Li. Measurement of the contributions of science and technology to the marine fisheries industry in the coastal regions of China. Marine Policy, 108(??):Article 103647, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308248.

Msomphora:2015:CQM

[MA15] Mbachi Ruth Msomphora and Margrethe Aanesen. Is the catch quota management (CQM) mechanism attractive to fishers? A preliminary analysis of the Danish 2011 CQM trial project. Marine Policy, 58(??):78–87, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000937.

McKinley:2018:CCM

[MA18] Emma McKinley and Tim G. Acott. Coastal communities: the missing link in marine policy? Marine Policy, 97(??):220–222, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305001.

Murray:2015:AMC

[MAAB15] Cathryn Clarke Murray, Selina Agbayani, Hussein M. Alidina, and Natalie C. Ban. Advancing marine cumulative effects mapping: an update in Canada’s Pacific waters. Marine Policy, 58(??):71–77, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000858.

Macinko:2014:LCS

[Mac14] Seth Macinko. Lipstick and catch shares in the Western Pacific: Beyond evangelism in fisheries policy? Marine Policy, 44(??):37–41, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001589.

Macdonald:2014:FKI

[MACM14] P. Macdonald, C. H. Angus, I. R. Cleasby, and C. T. Marshall. Fishers’ knowledge as an indicator of spatial and temporal trends in abundance of commercial fish species: Megrim (Lepidorhombus whiffiagonis) in the northern North Sea. Marine Policy, 45(??):228–239, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002431.

Morton:2016:VWS

[MAHP16] J. Morton, E. Ariza, M. Halliday, and C. Pita. Valuing the wild salmon fisheries of Scotland: the social and political dimensions of management. Marine Policy, 73 (??):35–45, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304559.

Mcleod:2018:RVP

[MAJM^{+}18] Elizabeth Mcleod, Seema Arora-Jonsson, Yuta J. Masuda, Mae Bruton-Adams, Carol O. Emaurois, Berna Gorong, C. J. Hudlow, Robyn James, Heather Kuhlken, Barbara Masike-Liri, Emeliana Musrasrik-Carl, Agnes Otzelberger, Kathryn Relang, Bertha M. Reyuw, Betty Sigrah, Christina Stinnett, Julita Tellei, and Laura Whitford. Raising the voices of Pacific Island women to inform climate adaptation policies. Marine Policy, 93(??):178–185, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300344.

Mallory:2013:CDW

[Mal13] Tabitha Grace Mallory. China’s distant water fishing industry: Evolving policies and implications. Marine Policy, 38(??):99–108, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001224.

Mallory:2016:FSC

[Mal16] Tabitha Grace Mallory. Fisheries subsidies in China: Quantitative and qualitative assessment of policy coherence and effectiveness. Marine Policy, 68(??):74–82, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000415.

Mallin:2018:SLR

[Mal18] Marc-Andrej Felix Mallin. From sea-level rise to seabed grabbing: the political economy of climate change in Kiribati. Marine Policy, 97(??):244–252, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302434.

Mangel:2010:SIE

[Man10] Marc Mangel. Scientific inference and experiment in Ecosystem Based Fishery Management, with application to Steller sea lions in the Bering Sea and Western Gulf of Alaska. Marine Policy, 34(5):836–843, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000060.

Mander:2017:SSN

[Man17] Sarah Mander. Slow steaming and a new dawn for wind propulsion: a multi-level analysis of two low carbon shipping transitions. Marine Policy, 75(??):210–216, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301300.

Markus:2010:TSF

[Mar10] Till Markus. Towards sustainable fisheries subsidies: Entering a new round of reform under the Common Fisheries Policy. Marine Policy, 34(6):1117–1124, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000631.

MacDonald:2015:VCW

[MAR^{+}15] Darla Hatton MacDonald, Ali Ardeshiri, John M. Rose, Bayden D. Russell, and Sean D. Connell. Valuing coastal water quality: Adelaide, South Australia metropolitan area. Marine Policy, 52(??):116–124, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002954.

Marciniak:2017:NIA

[Mar17] Konrad Jan Marciniak. New implementing agreement under UNCLOS: a threat or an opportunity for fisheries governance? Marine Policy, 84(??):320–326, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302580.

Mata-Alvarez-Santullano:2013:FEC

[MÁSSI13] Francisco Mata-Álvarez-Santullano and Antonio Souto-Iglesias. Fishing effort control policies and ship stability: Analysis of a string of accidents in Spain in the period 2004–2007. Marine Policy, 40(??):10–17, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002680.

Matlock:2014:PAN

[Mat14] Gary C. Matlock. The precautionary approach to non-native fisheries — The case of striped bass in Texas. Marine Policy, 47(??):94–98, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000475.

Matera:2016:LDI

[Mat16] Jaime Matera. Livelihood diversification and institutional (dis-)trust: Artisanal fishing communities under resource management programs in Providencia and Santa Catalina, Colombia. Marine Policy, 67(??):22–29, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000282.

Mazaris:2017:MMP

[Maz17] Antonios D. Mazaris. Manifestation of maritime piracy as an additional challenge for global conservation. Marine Policy, 77(??):171–175, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X16306145.

Marin:2010:NAU

[MB10] Andrés Marín and Fikret Berkes. Network approach for understanding small-scale fisheries governance: the case of the Chilean coastal co-management system. Marine Policy, 34(5):851–858, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000084.

Meinesz:2015:DMP

[MB15] Alexandre Meinesz and Aurélie Blanfuné. 1983–2013: Development of marine protected areas along the French Mediterranean coasts and perspectives for achievement of the Aichi target. Marine Policy, 54(??):10–16, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002875.

McKinley:2018:WLN

[MB18a] Emma McKinley and Rhoda C. Ballinger. Welsh legislation in a new era: a stakeholder perspective for coastal management. Marine Policy, 97(??):253–261, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304640.

Mullins:2018:FRD

[MB18b] Peter Mullins and Lee Burns. The fiscal regime for deep sea mining in the Pacific region. Marine Policy, 95(??): 337–345, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304808.

McKinley:2018:SES

[MBB18] Emma McKinley, Rhoda C. Ballinger, and Nicola J. Beaumont. Saltmarshes, ecosystem services, and an evolving policy landscape: a case study of Wales, UK. Marine Policy, 91(??):1–10, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303743.

Morrissey:2019:EDI

[MBBDS19] Karyn Morrissey, Shakeel B. Burthoo-Barah, Mukesh Dawoonauth, and Pasquale Lucio Scandizzo. Exploring the distributional impact of investment in the port sector on households in Mauritius: a social accounting matrix approach. Marine Policy, 99(??):324–333, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302914.

Madureira:2016:EPN

[MBCR16] Pedro Madureira, Harald Brekke, Georgy Cherkashov, and Marzia Rovere. Exploration of polymetallic nodules in the Area: Reporting practices, data management and transparency. Marine Policy, 70(??):101–107, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301270.

Macfadyen:2013:TST

[MBD13] Graeme Macfadyen, Richard Banks, and Robin Davies. Tropical shrimp trawling: Developing a management blueprint and adapting and implementing it in specific countries and fisheries. Marine Policy, 40(??):25–33, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000079.

Miller:2014:PEE

[MBM14] Alice M. M. Miller, Simon R. Bush, and Arthur P. J. Mol. Power Europe: EU and the illegal, unreported and unregulated tuna fisheries regulation in the West and Central Pacific Ocean. Marine Policy, 45(??):138–145, March 2014.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002923.

Madrigal-Ballestero:2013:WMT

[MBSL13] Róger Madrigal-Ballestero, Achim Schlüter, and Maria Claudia Lopez. What makes them follow the rules? Empirical evidence from turtle egg harvesters in Costa Rica. Marine Policy, 37(??):270–277, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000978.

Marino:2019:LAS

[MBT^{+}19] Mónica Mariño, Annette Breckwoldt, Mirta Teichberg, Alfred Kase, and Hauke Reuter. Livelihood aspects of seaweed farming in Rote Island, Indonesia. Marine Policy, 107(??):Article 103600, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300545.

Mellado:2014:ULK

[MBTV14] Tiscar Mellado, Timothée Brochier, Julien Timor, and Javier Vitancurt. Use of local knowledge in marine protected area management. Marine Policy, 44(??): 390–396, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002297.

Macusi:2015:STT

[MBvZ15] Edison D. Macusi, Ricardo P. Babaran, and P. A. M. van Zwieten. Strategies and tactics of tuna fishers in the payao (anchored FAD) fishery from general Santos City, Philippines. Marine Policy, 62(??):63–73, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002511.

Miller:2014:FLO

[MC14] Peter I. Miller and Stelios Christodoulou. Frequent locations of oceanic fronts as an indicator of pelagic diversity: Application to marine protected areas and renewables. Marine Policy, 45(??):318–329, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002066.

Mamula:2015:MPE

[MC15] Aaron Mamula and Trevor Collier. Multifactor productivity, environmental change, and regulatory impacts in the U.S. West Coast groundfish trawl fishery, 1994–2013. Marine Policy, 62(??):326–336, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500161X.

Moon:2016:DRB

[MC16] Ruby Moon and Flaxen Conway. Does the relationship between fishermen and enforcers impact regulatory compliance? Marine Policy, 74(??):316–322, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301981.

Mazur:2019:LTF

[MC19] Nicole A. Mazur and Allan L. Curtis. Let’s talk fish: Examining public judgements of Australia’s wild-catch commercial fishing industry. Marine Policy, 99(??):14–20, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308242.

Mulazzani:2013:MGM

[MCAM13] Luca Mulazzani, Richard Curtin, Marga Andrés, and Giulio Malorgio. Multilevel governance and management of shared stocks with integrated markets: the European anchovy case. Marine Policy, 38(??):407–416, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001595.

Moore:2019:IFC

[MCB^{+}19] Stephanie K. Moore, Michael R. Cline, Kathryn Blair, Terrie Klinger, Anna Varney, and Karma Norman. An index of fisheries closures due to harmful algal blooms and a framework for identifying vulnerable fishing communities on the U.S. West Coast. Marine Policy, 110(??):Article 103543, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307073.

Mulazzani:2018:ITE

[MCBM18] Luca Mulazzani, Luca Camanzi, Alfonso Bonezzi, and Giulio Malorgio. Individual transferable effort quotas for Italian fisheries? A preliminary analysis. Marine Policy, 91(??):14–21, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306565.

McCormack:2017:SNZ

[McC17] Fiona McCormack. Sustainability in New Zealand’s quota management system: a convenient story. Marine Policy, 80(??):35–46, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303785.

Malpica-Cruz:2016:MMI

[MCCC16] Luis Malpica-Cruz, Laís C. T. Chaves, and Isabelle M. Côté. Managing marine invasive species through public participation: Lionfish derbies as a case study. Marine Policy, 74(??):158–164, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304304.

McNeill:2018:AMP

[MCH18] Asha McNeill, Julian Clifton, and Euan S. Harvey. Attitudes to a marine protected area are associated with perceived social impacts. Marine Policy, 94(??):106–118, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1730831X.

McNeill:2019:SRF

[MCH19] Asha McNeill, Julian Clifton, and Euan S. Harvey. Specialised recreational fishers reject sanctuary zones and favour fisheries management. Marine Policy, 107(??):

Article 103592, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19301216.

Mellett:2011:AEB

[MCHO11] M. Mellett, C. Curtin, T. Hennessey, and A. M. O’Hagan. Attainment of ecosystem based governance in European waters — a State property rights regime approach for Ireland. Marine Policy, 35(6):739–747, December 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000121.

McLaughlin:2015:BLM

[McL15] Richard J. McLaughlin. Bio-logging as marine scientific research under the law of the sea: a commentary responding to James Kraska, Guillermo Ortuño Crespo, David W. Johnston, bio-logging of marine migratory species in the law of the sea, Marine Policy 51 (2015) 394–

400. Marine Policy, 60(??):178–181, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001694.

Mangin:2018:CMD

[MCMB^{+}18] Tracey Mangin, Miguel Ángel Cisneros-Mata, Jennifer Bone, Christopher Costello, Steven D. Gaines, Gavin McDonald, Laura Rodriguez, C. Kent Strauss, and Pedro Zapata. The cost of management delay: the case for reforming Mexican fisheries sooner rather than later. Marine Policy, 88(??):1–10, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730475X.

McNeil:2018:ARB

[McN18] Ewan M. McNeil. Antifouling: Regulation of biocides in the UK before and after Brexit. Marine Policy, 92(??):58–60, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308308.

McManus:2019:CSU

[MCY^{+}19] Edmund McManus, Martin Collins, Oliver Yates, Matthew Sanders, Bryony Townhill, Stephen Mangi, and Emmanouil Tyllianakis. Commonwealth SIDS and UK Overseas Territories sustainable fisheries programmes: an overview of projects and benefits of official development assistance funding. Marine Policy, 107(??):Article 103437, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301149.

Murray:2015:SSS

[MD15] Grant Murray and Linda D’Anna. Seeing shellfish from the seashore: the importance of values and place in perceptions of aquaculture and marine social–ecological system interactions. Marine Policy, 62(??):125–133, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002560.

Mahajan:2016:PES

[MD16] Shauna L. Mahajan and Tim Daw. Perceptions of ecosystem services and benefits to human well-being from community-based marine protected areas in Kenya. Marine Policy, 74(??):108–119, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300276.

Manez:2014:FLF

[MDE14] Kathleen Schwerdtner Máñez, Lorel Dandava, and Werner Ekau. Fishing the last frontier: the introduction of the marine aquarium trade and its impact on local fishing communities in Papua New Guinea. Marine Policy, 44(??):279–286, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002157.

Malone:2014:GOO

[MDG^{+}14] Thomas C. Malone, Paul M. DiGiacomo, Emanuel Gonçalves, Anthony H. Knap, Liana Talaue-McManus, and Stephen de Mora. A global ocean observing system framework for sustainable development. Marine Policy, 43 (??):262–272, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001346.

Marzetti:2016:VAI

[MDK^{+}16] S. Marzetti, M. Disegna, E. Koutrakis, A. Sapounidis, V. Marin, S. Martino, S. Roussel, H. Rey-Valette, and C. Paoli. Visitors’ awareness of ICZM and WTP for beach preservation in four European Mediterranean regions. Marine Policy, 63(??):100–108, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002894.

Murray:2016:MWW

[MDM16] Grant Murray, Linda D’Anna, and Patricia MacDonald. Measuring what we value: the utility of mixed methods approaches for incorporating values into marine socialecological system management. Marine Policy, 73(??): 61–68, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301294.

Maynou:2018:FPE

[MdMGV^{+}18] Francesc Maynou, Maria del Mar Gil, Sergio Vitale, Giovan Battista Giusto, Antigoni Foutsi, Mafalda Rangel, Rita Rainha, Karim Erzini, Jorge M. S. Gonçalves, Luis Bentes, Claudio Viva, Paolo Sartor, Francesco De Carlo, Ilaria Rossetti, Maria Christou, Konstantinos Stergiou, Christos D. Maravelias, and Dimitrios Damalas. Fishers’ perceptions of the European Union discards ban: perspective from south European fisheries. Marine Policy, 89(??):147–153, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301008.

Mendenhall:2019:STH

[MDNT19] Elizabeth Mendenhall, Elizabeth De Santo, Elizabeth Nyman, and Rachel Tiller. A soft treaty, hard to reach: the second inter-governmental conference for biodiversity beyond national jurisdiction. Marine Policy, 108(??):

Article 103664, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19303094.

Moore:2016:IEF

[MDR^{+}16] Cordelia Moore, Jeffrey C. Drazen, Ben T. Radford, Christopher Kelley, and Stephen J. Newman. Improving essential fish habitat designation to support sustainable ecosystem-based fisheries management. Marine Policy, 69(??):32–41, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300616.

Meek:2011:PUP

[Mee11] Chanda L. Meek. Putting the US polar bear debate into context: the disconnect between old policy and new problems. Marine Policy, 35(4):430–439, July 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002034.

Mendenhall:2018:OPR

[Men18] Elizabeth Mendenhall. Oceans of plastic: a research agenda to propel policy development. Marine Policy, 96 (??):291–298, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302902.

Mendenhall:2019:ILS

[Men19] Elizabeth Mendenhall. Interpreting the law of the sea ‘regime of islands’: an opportunity for productive US leadership. Marine Policy, 99(??):213–215, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300769.

Mesnil:2012:HEM

[Mes12] Benoit Mesnil. The hesitant emergence of maximum sustainable yield (MSY) in fisheries policies in Europe. Marine Policy, 36(2):473–480, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001412.

McCrimmon:2010:UMU

[MF10] Donald McCrimmon and Lucia Fanning. Using Memoranda of Understanding to facilitate marine management in Canada. Marine Policy, 34(6):1335–1340, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001120.

McKinley:2012:IME

[MF12] Emma McKinley and Stephen Fletcher. Improving marine environmental health through marine citizenship: a call for debate. Marine Policy, 36(3):839–843, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001813.

Mahon:2019:ROGa

[MF19a] Robin Mahon and Lucia Fanning. Regional ocean governance: Integrating and coordinating mechanisms for polycentric systems. Marine Policy, 107(??):Article 103589, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302763.

Mahon:2019:ROGb

[MF19b] Robin Mahon and Lucia Fanning. Regional ocean governance: Polycentric arrangements and their role in global ocean governance. Marine Policy, 107(??):Article 103590, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302751.

Mather:2019:SLA

[MF19c] Charles Mather and Lucia Fanning. Social licence and aquaculture: Towards a research agenda. Marine Policy, 99(??):275–282, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304603.

Moura:2016:UDM

[MFAA16] Teresa Moura, António Fernandes, Ricardo Alpoim, and Manuela Azevedo. Unravelling the dynamics of a multi-gear fleet — Inputs for fisheries assessment and management under the Common Fisheries Policy. Marine Policy, 72(??):219–230, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630121X.

Maccarrone:2014:IBS

[MFB^{+}14] Vincenzo Maccarrone, Francesco Filiciotto, Gaspare Buffa, Salvatore Mazzola, and Giuseppa Buscaino. The ICZM Balanced Scorecard: a tool for putting integrated coastal zone management into action. Marine Policy, 44(??):321–334, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002212.

Moura:2018:MDW

[MFF^{+}18] Teresa Moura, António Fernandes, Ivone Figueiredo, Ricardo Alpoim, and Manuela Azevedo. Management of deep-water sharks’ by-catch in the Portuguese anglerfish fishery: from EU regulations to practice. Marine Policy, 90(??):55–67, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17301458.

Mahon:2010:GCL

[MFMP10] Robin Mahon, Lucia Fanning, Patrick McConney, and Richard Pollnac. Governance characteristics of large marine ecosystems. Marine Policy, 34(5):919–927, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000175.

Madau:2018:TES

[MFP18] Fabio A. Madau, Roberto Furesi, and Pietro Pulina. The technical efficiency in Sardinian fisheries cooperatives. Marine Policy, 95(??):111–116, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306516.

Minarro:2016:RPC

[MFRvP16] Sara Miñarro, Gabriela Navarrete Forero, Hauke Reuter, and Ingrid E. van Putten. The role of patron-client relations on the fishing behaviour of artisanal fishermen in the Spermonde Archipelago (Indonesia). Marine Policy, 69(??):73–83, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300574.

McGhee:2019:WDB

[MFT19] Clara McGhee, Lynne Falconer, and Trevor Telfer. What does ‘beyond compliance’ look like for the Scottish salmon aquaculture industry? Marine Policy, 109(??):Article 103668, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830798X.

Morris:2014:MRS

[MG14] Corey J. Morris and John M. Green. MPA regulations should incorporate adaptive management — The case of Gilbert Bay Labrador Atlantic cod (Gadus morhua). Marine Policy, 49(??):20–28, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000979.

Murray:2016:SMM

[MG16] Alexander G. Murray and Matt Gubbins. Spatial management measures for disease mitigation as practiced in Scottish aquaculture. Marine Policy, 70(??):93–100, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301476.

Mallol:2019:UCA

[MG19] Sandra Mallol and Raquel Goñi. Unintended changes of artisanal fisheries métiers upon implementation of an MPA. Marine Policy, 101(??):237–245, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306583.

Marin:2010:TCD

[MGA^{+}10] Andrés Marín, Stefan Gelcich, Gonzalo Araya, Gonzalo Olea, Miguel Espíndola, and Juan C. Castilla. The 2010 tsunami in Chile: Devastation and survival of coastal small-scale fishing communities. Marine Policy, 34(6):1381–1384, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001168.

Milakovic:2018:CSF

[MGB^{+}18] Aleksandar-Saša Milaković, Bjørn Gunnarsson, Sergey Balmasov, Sungwon Hong, Kitae Kim, Peter Schütz, and Sören Ehlers. Current status and future operational models for transit shipping along the Northern Sea Route. Marine Policy, 94(??):53–60, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308576.

Mirrasooli:2019:FAI

[MGG^{+}19] Elham Mirrasooli, Rasoul Ghorbani, Saeid Gorgin, Seyed Mostafa Aghilinejhad, and Ali Jalali. Factors associated with illegal fishing and fisher attitudes toward sturgeon conservation in the southern Caspian Sea. Marine Policy, 100(??):107–115, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305736.

Mileski:2018:CDD

[MGGS18] Joan Mileski, Amir Gharehgozli, Lawrence Ghoram, and Ryan Swaney. Cooperation in developing a disaster prevention and response plan for Arctic shipping. Marine Policy, 92(??):131–137, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308692.

Marshall:2014:SDM

[MGH14] C. E. Marshall, G. A. Glegg, and K. L. Howell. Species distribution modelling to support marine conservation planning: the next steps. Marine Policy, 45(??):330–332, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002005.

McQuatters-Gollop:2017:MMC

[MGJB^{+}17] Abigail McQuatters-Gollop, David G. Johns, Eileen Bresnan, Jennifer Skinner, Isabelle Rombouts, Rowena Stern, Anaïs Aubert, Marie Johansen, Jacob Bedford, and Antony Knights. From microscope to management: the critical value of plankton taxonomy to marine policy and biodiversity conservation. Marine Policy, 83(??):1–10, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307874.

Mileski:2013:FMA

[MH13] Joan P. Mileski and Jared Honeycutt. Flexibility in maritime assets and pooling strategies: a viable response todisaster. Marine Policy, 40(??):111–116, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000146.

McDonald:2017:IBA

[MHA^{+}17] Gavin McDonald, Bill Harford, Alejandro Arrivillaga, Elizabeth A. Babcock, Ramon Carcamo, James Foley, Rod Fujita, Todd Gedamke, Janet Gibson, Kendra Karr, Julie Robinson, and Jono Wilson. An indicator-based adaptive management framework and its development for data-limited fisheries in Belize. Marine Policy, 76 (??):28–37, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307527.

Macfadyen:2019:MCP

[MHC19] Graeme Macfadyen, Tim Huntington, and Rod Cappell. Maximising the chances for positive impacts of fisheries and aquaculture Official Development Assistance. Marine Policy, 107(??):Article 103218, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830280X.

Mackelworth:2013:UGC

[MHF13] P. Mackelworth, D. Holcer, and C. M. Fortuna. Unbalanced governance: the Cres–Lošinj special marine reserve, a missed conservation opportunity. Marine Policy, 41(??): 126–133, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002588.

McIlgorm:2010:HWC

[MHK^{+}10] Alistair McIlgorm, Susan Hanna, Gunnar Knapp, Pascal Le Floc’h., Frank Millerd, and Minling Pan. How will climate change alter fishery governance — insights from seven international case studies. Marine Policy, 34 (1):170–177, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0900092X.

Mackelworth:2013:UCT

[MHL13] Peter Mackelworth, Draško Holcer, and Bojan Lazar. Using conservation as a tool to resolve conflict: Establishing the Piran–Savudrija international Marine Peace Park. Marine Policy, 39(??):112–119, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001996.

Maxwell:2015:DOM

[MHL^{+}15] Sara M. Maxwell, Elliott L. Hazen, Rebecca L. Lewison, Daniel C. Dunn, Helen Bailey, Steven J. Bograd, Dana K. Briscoe, Sabrina Fossette, Alistair J. Hobday, Meredith Bennett, Scott Benson, Margaret R. Caldwell, Daniel P. Costa, Heidi Dewar, Tomo Eguchi, Lucie Hazen, Suzanne Kohin, Tim Sippel, and Larry B. Crowder. Dynamic ocean management: Defining and conceptualizing real-time management of the ocean. Marine Policy, 58(??):42–50, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000639.

Moreno:2016:FFC

[MHM16] Guillermo Moreno, Miguel Herrera, and Julio Morón. To FAD or not to FAD: a challenge to the Marine Stewardship Council and its Conformity Assessment Bodies on the use of Units of Assessment and Units of Certification for industrial purse seine tuna fisheries. Marine Policy, 73(??):100–107, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303360.

Murray:2015:ATS

[MI15] Grant D. Murray and Danny Ings. Adaptation in a time of stress: a social-ecological perspective on changing fishing strategies in the Canadian snow crab fishery. Marine Policy, 60(??):280–286, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002158.

Milligan:2014:POC

[Mil14] Ben Milligan. Planning for offshore CO _{2} storage: Law and policy in the United Kingdom. Marine Policy, 48 (??):162–171, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001080.

Misund:2019:NOP

[Mis19] Andreas Ugelvik Misund. From a natural occurring parasitic organism to a management object: Historical perceptions and discourses related to salmon lice in Norway. Marine Policy, 99(??):400–406, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301453.

Maya-Jariego:2017:ELC

[MJQMVÁ17] Isidro Maya-Jariego, José F. Querevalú-Miñán, Lourdes G. Varela, and Javier Ávila. Escape the lion cage: Social networks by catch zones of small-scale fisheries in the oil settlement of Lobitos (Peru). Marine Policy, 81(??):340–349, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300489.

Maya-Jariego:2016:RBP

[MJRdC16] Isidro Maya-Jariego, Daniel Holgado Ramos, and David Florido del Corral. Relations between professional groups in the Atlantic and Mediterranean fishing enclaves of Andalusia (Spain): a personal networks approach with clustered graphs. Marine Policy, 72(??):48–58, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303736.

Mattingley:2016:EAB

[MJS16] Lauren Mattingley, Sevvandi Jayakody, and Beth E. Scott. Exploring the applicability of biological and socioeconomic tools in developing EAFM plans for data absent areas: Spinner dolphin EAFM for Kalpitiya, Sri Lanka. Marine Policy, 68(??):136–145, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300641.

Mackay:2018:WPC

[MJvP^{+}18] M. Mackay, S. Jennings, E. I. van Putten, H. Sibly, and S. Yamazaki. When push comes to shove in recreational fishing compliance, think ‘nudge’. Marine Policy, 95(??):256–266, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300939.

Moutopoulos:2014:FSD

[MK14] Dimitrios K. Moutopoulos and Constantin Koutsikopoulos. Fishing strange data in national fisheries statistics of Greece. Marine Policy, 48(??):114–122, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400089X.

Masud:2015:CAT

[MK15] Muhammad Mehedi Masud and Fatimah Binti Kari. Community attitudes towards environmental conservation behaviour: an empirical investigation within MPAs, Malaysia. Marine Policy, 52(??):138–144, February 2015.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002772.

Macusi:2017:IEF

[MKB17] Edison D. Macusi, Robert E. Katikiro, and Ricardo P. Babaran. The influence of economic factors in the change of fishing strategies of anchored FAD fishers in the face of declining catch, General Santos City, Philippines. Marine Policy, 78(??):98–106, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306649.

Mason:2012:ECA

[MKMS12] Janet Mason, Rosemary Kosaka, Aaron Mamula, and Cameron Speir. Effort changes around a marine reserve: the case of the California Rockfish Conservation Area. Marine Policy, 36(5):1054–1063, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000450.

Marinesque:2012:GIM

[MKR12] Sophie Marinesque, David M. Kaplan, and Lynda D. Rodwell. Global implementation of marine protected areas: Is the developing world being left behind? Marine Policy, 36(3):727–737, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001606.

Militz:2018:UTA

[MKSS18] Thane A. Militz, Jeff Kinch, David S. Schoeman, and Paul C. Southgate. Use of total allowable catch to regulate a selective marine aquarium fishery. Marine Policy, 90(??):160–167, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17301185.

Myrberg:2019:POS

[MKU19] Kai Myrberg, Samuli Korpinen, and Laura Uusitalo. Physical oceanography sets the scene for the Marine Strategy Framework Directive implementation in the Baltic Sea.

Marine Policy, 107(??):Article 103591, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830191X.

Mabon:2017:CSI

[MKX17] Leslie Mabon, Jun Kita, and Ziqiu Xue. Challenges for social impact assessment in coastal regions: a case study of the Tomakomai CCS Demonstration Project. Marine Policy, 83(??):243–251, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303330.

Mitchell:2019:OID

[ML19] Rebecca J. Mitchell and Reidar P. Lystad. Occupational injury and disease in the Australian aquaculture industry. Marine Policy, 99(??):216–222, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305864.

Moore:2016:ASE

[MLA16] Francesca Moore, Jessica Lamond, and Thomas Appleby. Assessing the significance of the economic impact of Marine Conservation Zones in the Irish Sea upon the fisheries sector and regional economy in Northern Ireland. Marine Policy, 74(??):136–142, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303980.

Matz-Luck:2014:IOP

[MLF14] Nele Matz-Lück and Johannes Fuchs. The impact of OSPAR on protected area management beyond national jurisdiction: Effective regional cooperation or a network of paper parks? Marine Policy, 49(??):155–166, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002844.

Maldonado:2017:TFM

[MLF^{+}17] Adriana Guzmán Maldonado, Priscila Fabiana Macedo Lopes, Carlos Alberto Rodríguez Fernández, Carlos Andrés Lasso Alcala, and Ussif Rashid Sumalia. Transboundary fisheries management in the Amazon: Assessing current policies for the management of the ornamental silver arawana (Osteoglossum bicirrhosum). Marine Policy, 76(??):192–199, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305772.

Mulvaney:2015:CNB

[MLHP15] Kate K. Mulvaney, Seungyoon Lee, Tomas O. Höök, and Linda S. Prokopy. Casting a net to better understand fisheries management: an affiliation network analysis of the Great Lakes Fishery Commission. Marine Policy, 57(??):120–131, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000573.

McDonald:2016:CSP

[MLR^{+}16] Sara L. McDonald, Rebecca L. Lewison, Stephen E. Roady, Randall A. Kramer, Deborah Rigling-Gallagher, and Andrew J. Read. Comparing stakeholder perceptions with empirical outcomes from negotiated rulemaking policies: Is participant satisfaction a proxy for policy success? Marine Policy, 73(??):224–230, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303591.

Morzaria-Luna:2014:SIV

[MLTBMB14] Hem Nalini Morzaria-Luna, Peggy Turk-Boyer, and Marcia Moreno-Baez. Social indicators of vulnerability for fishing communities in the Northern Gulf of California, Mexico: Implications for climate change. Marine Policy, 45(??):182–193, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002388.

Meek:2011:AGH

[MLV^{+}11] Chanda L. Meek, Amy Lauren Lovecraft, Riku Varjopuro, Martha Dowsley, and Aaron T. Dale. Adaptive governance and the human dimensions of marine mammal management: Implications for policy in a changing North. Marine Policy, 35(4):466–476, July 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002083.

Mulazzani:2014:ECD

[MM14] Luca Mulazzani and Giulio Malorgio. The external and commercial dimensions of the EU fisheries policy: an institutional approach applied to the whitefish case. Marine Policy, 46(??):123–131, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000232.

Madin:2015:ICF

[MM15a] Elizabeth M. P. Madin and Peter I. Macreadie. Incorporating carbon footprints into seafood sustainability certification and eco-labels. Marine Policy, 57(??):178–181, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000585.

Mulazzani:2015:TCE

[MM15b] Luca Mulazzani and Giulio Malorgio. Is there coherence in the European Union’s strategy to guarantee the supply of fish products from abroad? Marine Policy, 52 (??):1–10, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002802.

Mardle:2017:ICE

[MM17a] Simon Mardle and Sebastien Metz. Impacts of current EU regulation on the UK whitefish value chain. Marine Policy, 84(??):52–59, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300052.

Mulazzani:2017:BGE

[MM17b] Luca Mulazzani and Giulio Malorgio. Blue growth and ecosystem services. Marine Policy, 85(??):17–24, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301021.

Mackinson:2018:EEA

[MM18] S. Mackinson and D. A. J. Middleton. Evolving the ecosystem approach in European fisheries: Transferable lessons from New Zealand’s experience in strengthening stakeholder involvement. Marine Policy, 90(??):194–202, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308400.

McVeigh:2019:SWF

[MM19] Joanne McVeigh and Malcolm MacLachlan. A silver wave? Filipino shipmates’ experience of merchant seafaring. Marine Policy, 99(??):283–297, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306584.

Murillas-Maza:2016:PIF

[MMA16] Arantza Murillas-Maza and Marga Andrés. Providing incentives for fishermen through rights-based comanagement systems. An impact-assessment on Basque fisheries. Marine Policy, 70(??):128–136, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300604.

Myers:2019:RFP

[MMB^{+}19] Hannah J. Myers, Michael J. Moore, Mark F. Baumgartner, Sean W. Brillant, Steven K. Katona, Amy R. Knowlton, Lyne Morissette, Heather M. Pettis, Geoff Shester, and Timothy B. Werner. Ropeless fishing to prevent large whale entanglements: Ropeless Consortium report. Marine Policy, 107(??):Article 103587, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301575.

Morfin:2017:NNS

[MMBK17] Marie Morfin, Sonia Méhault, Hugues P. Benoît, and Dorothée Kopp. Narrowing down the number of species requiring detailed study as candidates for the EU Common Fisheries Policy discard ban. Marine Policy, 77(??):23–29, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302895.

Muallil:2014:CTP

[MMC^{+}14a] Richard N. Muallil, Samuel S. Mamauag, Jeffrey T. Cababaro, Hazel O. Arceo, and Porfirio M. Aliño. Catch trends in Philippine small-scale fisheries over the last five decades: the fishers’ perspectives. Marine Policy, 47 (??):110–117, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000463.

Muallil:2014:STC

[MMC^{+}14b] Richard N. Muallil, Samuel S. Mamauag, Reniel B. Cabral, Emerlinda O. Celeste-Dizon, and Porfirio M. Aliño. Status, trends and challenges in the sustainability of smallscale fisheries in the Philippines: Insights from FISHDA (Fishing Industries’ Support in Handling Decisions Application) model. Marine Policy, 44(??):212–221, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001929.

Martin:2014:PNM

[MMdJ^{+}14] P. Martín, A. Muntadas, S. de Juan, P. Sánchez, and M. Demestre. Performance of a northwestern Mediterranean bottom trawl fleet: How the integration of landings and VMS data can contribute to the implementation of ecosystem-based fisheries management. Marine Policy, 43(??):112–121, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001231.

Mullowney:2012:IBT

[MMDS12] Darrell R. J. Mullowney, Corey J. Morris, Earl G. Dawe, and Katherine R. Skanes. Impacts of a bottom trawling exclusion zone on Snow Crab abundance and fish harvester behavior in the Labrador Sea, Canada. Marine Policy, 36(3):567–575, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001552.

Martin:2016:SQR

[MMGM16] Carol L. Martin, Salim Momtaz, Troy Gaston, and Natalie A. Moltschaniwskyj. A systematic quantitative review of coastal and marine cultural ecosystem services: Current status and future research. Marine Policy, 74 (??):25–32, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300902.

Martin:2015:AES

[MMJM15] Carol L. Martin, Salim Momtaz, Alan Jordan, and Natalie A. Moltschaniwskyj. An assessment of the effectiveness of in-situ signage in multiple-use marine protected areas in providing information to different recreational users. Marine Policy, 56(??):78–85, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000512.

Martin:2016:ERF

[MMJM16] Carol L. Martin, Salim Momtaz, Alan Jordan, and Natalie A. Moltschaniwskyj. Exploring recreational fishers’ perceptions, attitudes, and support towards a multipleuse marine protected area six years after implementation. Marine Policy, 73(??):138–145, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304961.

Malvarosa:2019:SIA

[MML^{+}19] Loretta Malvarosa, Arantza Murillas, Sigrid Lehuta, J. Rasmus Nielsen, Claire Macher, Leyre Goti, Arina Motova, Ralf Doering, Gunnar Haraldson, Paolo Accadia, Katell G. Hamon, Francois Bastardie, Christos D. Maravelias, Simon Mardle, and Thomas Thøgersen. Sustainability Impact Assessment (SIA) in fisheries: Implementation in EU fishing regions. Marine Policy, 101(??):63–79, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308564.

Macedo:2019:MUM

[MMM19] Heitor Schulz Macedo, Rodrigo Pereira Medeiros, and Patrick McConney. Are multiple-use marine protected areas meeting fishers’ proposals? Strengths and constraints in fisheries’ management in Brazil. Marine Policy, 99 (??):351–358, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302367.

Martins:2018:ETI

[MMN18] R. R. M. Martins and C. Monteiro-Neto. Economic trends of industrial double-rig bottom trawlers in Southeastern Brazil. Marine Policy, 90(??):125–136, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306589.

Masden:2015:RED

[MMOL15] Elizabeth A. Masden, Aly McCluskie, Ellie Owen, and Rowena H. W. Langston. Renewable energy developments in an uncertain world: the case of offshore wind and birds in the UK. Marine Policy, 51(??):169–172, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400222X.

MacDonald:2015:CSV

[MMP15] Patricia A. MacDonald, Grant Murray, and Michele Patterson. Considering social values in the seafood sector using the Q-method. Marine Policy, 52(??):68–76, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002917.

Malakar:2018:PMS

[MMP18] Krishna Malakar, Trupti Mishra, and Anand Patwardhan. Perceptions of multi-stresses impacting livelihoods of marine fishermen. Marine Policy, 97(??):18–26, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303531.

McCay:2014:CCC

[MMPD^{+}14] Bonnie J. McCay, Fiorenza Micheli, Germán Ponce-Díaz, Grant Murray, Geoff Shester, Saudiel Ramirez-Sanchez, and Wendy Weisman. Cooperatives, concessions, and co-management on the Pacific coast of Mexico. Marine Policy, 44(??):49–59, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001553.

Morandeau:2014:WDF

[MMS^{+}14] G. Morandeau, C. Macher, F. Sanchez, N. Bru, L. Fauconnet, and N. Caill-Milly. Why do fishermen discard? Distribution and quantification of the causes of discards in the Southern Bay of Biscay passive gear fisheries. Marine Policy, 48(??):30–38, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000761.

McDonald:2016:DFO

[MMTC16] Gavin McDonald, Tracey Mangin, Lennon R. Thomas, and Christopher Costello. Designing and financing optimal enforcement for small-scale fisheries and dive tourism industries. Marine Policy, 67(??):105–117, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000476.

Morgera:2018:LSS

[MN18] Elisa Morgera and Mara Ntona. Linking small-scale fisheries to international obligations on marine technology transfer. Marine Policy, 93(??):295–306, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303238.

Morales-Nin:2015:SEO

[MNA15] Beatriz Morales-Nin and Joan Albaigés. SEAS-ERA: an overarching effort to coordinate marine research policies across Europe. Marine Policy, 61(??):33–38, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1500192X.

Murray:2018:DCO

[MNG^{+}18] Fiona Murray, Katherine Needham, Kate Gormley, Sally Rouse, Joop W. P. Coolen, David Billett, Jennifer Dannheim, Silvana N. R. Birchenough, Kieran Hyder, Richard Heard, Joseph S. Ferris, Jan M. Holstein, Lea-Anne Henry, Oonagh McMeel, Jan-Bart Calewaert, and J. Murray Roberts. Data challenges and opportunities for environmental management of North Sea oil and gas decommissioning in an era of blue growth. Marine Policy, 97(??):130–138, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302355.

Merayo:2018:ITQ

[MNHN18] Eugenia Merayo, Rasmus Nielsen, Ayoe Hoff, and Max Nielsen. Are individual transferable quotas an adequate solution to overfishing and overcapacity? Evidence from Danish fisheries. Marine Policy, 87(??):167–176, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302014.

Martinez-Novo:2017:ASR

[MNLHRMP17] Rodrigo Martínez-Novo, Emmánuel Lizcano, Paloma Herrera-Racionero, and Lluís Miret-Pastor. Aquaculture stakeholders role in fisheries co-management. Marine Policy, 76(??):130–135, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303219.

Mansouri:2010:PMF

[MNM10] Mo Mansouri, Roshanak Nilchiani, and Ali Mostashari. A policy making framework for resilient port infrastructure systems. Marine Policy, 34(6):1125–1134, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000643.

Morrissey:2012:IME

[MO12] Karyn Morrissey and Cathal O’Donoghue. The Irish marine economy and regional development. Marine Policy, 36(2):358–364, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001138.

Morrissey:2013:RMS

[MO13] Karyn Morrissey and Cathal O’Donoghue. The role of the marine sector in the Irish national economy: an input– output analysis. Marine Policy, 37(??):230–238, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000929.

Merrie:2014:IAP

[MO14] Andrew Merrie and Per Olsson. An innovation and agency perspective on the emergence and spread of Marine Spatial Planning. Marine Policy, 44(??):366–374, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002315.

MacKenzie:2011:HEP

[MOE11] Brian R. MacKenzie, Henn Ojaveer, and Margit Eero. Historical ecology provides new insights for ecosystem management: eastern Baltic cod case study. Marine Policy, 35(2):266–270, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001727.

Morrissey:2011:QVM

[MOH11] Karyn Morrissey, Cathal O’Donoghue, and Stephen Hynes. Quantifying the value of multi-sectoral marine commercial activity in Ireland. Marine Policy, 35(5):721–727, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000492.

McClenachan:2015:ACC

[MOR15] Loren McClenachan, Grace O’Connor, and Travis Reynolds. Adaptive capacity of co-management systems in the face of environmental change: the soft-shell clam fishery and invasive green crabs in Maine. Marine Policy, 52(??): 26–32, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002851.

Morgan:2017:ICU

[Mor17] Richard Morgan. An investigation of constraints upon fisheries diversification using the Analytic Hierarchy Process (AHP). Marine Policy, 86(??):24–30, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300969.

Moxnes:2012:ITQ

[Mox12] Erling Moxnes. Individual transferable quotas versus auctioned seasonal quotas: an experimental investigation. Marine Policy, 36(2):339–349, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001187.

Martino:2012:CBM

[MP12] Simone Martino and Lindsay M. Parson. A comparison between manganese nodules and cobalt crust economics in a scenario of mutual exclusivity. Marine Policy, 36 (3):790–800, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1100193X.

McAuliffe:2014:EAP

[MPCB14] Sarah McAuliffe, Jonathan Potts, Rosaline Canessa, and Brian Baily. Establishing attitudes and perceptions of recreational boat users based in the River Hamble Estuary, UK, towards Marine Conservation Zones. Marine Policy, 45(??):98–107, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002510.

Maguire:2012:RSM

[MPF12] Bernadine Maguire, Jonathan Potts, and Stephen Fletcher. The role of stakeholders in the marine planning process — Stakeholder analysis within the Solent, United Kingdom. Marine Policy, 36(1):246–257, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001011.

Miret-Pastor:2018:AFD

[MPMGGAHR18] Lluís Miret-Pastor, Agustín Molina-García, César García-Aranda, and Paloma Herrera-Racionero. Analysis of the fisheries diversification funds in Spain during the period 2007–2014. Marine Policy, 93(??):150–158, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300757.

Moutopoulos:2016:QIC

[MPMK16] Dimitrios K. Moutopoulos, George Prodromitis, Irene Mantzouni, and Constantin Koutsikopoulos. Quantifying the implementation of Common Fisheries Policy: Patterns of fisheries violations and penalties imposed in Greek waters. Marine Policy, 70(??):65–76, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301336.

Mason:2010:CTE

[MPPB10] Claire Mason, Gillian Paxton, Joanna Parr, and Naomi Boughen. Charting the territory: Exploring stakeholder reactions to the prospect of seafloor exploration and mining in Australia. Marine Policy, 34(6):1374–1380, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001181.

Miret-Pastor:2014:EAS

[MPPSdVSOHR14] Lluís Miret-Pastor, Ángel Peiró-Signes, Maria del Val Segarra-Oña, and Paloma Herrera-Racionero. Empirical analysis of sustainable fisheries and the relation to economic performance enhancement: the case of the Spanish fishing industry. Marine Policy, 46(??):105–110, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000116.

Mullen:2013:DPC

[MPT13] Kaitlyn A. Mullen, Michael L. Peterson, and Sean K. Todd. Has designating and protecting critical habitat had an impact on endangered North Atlantic right whale ship strike mortality? Marine Policy, 42(??):293–304, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000973.

Mayo-Ramsay:2010:ELS

[MR10] Julia Mayo-Ramsay. Environmental, legal and social implications of ocean urea fertilization: Sulu Sea example. Marine Policy, 34(5):831–835, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000059.

Morton:2016:RPS

[MR16] Alexandra Morton and Richard Routledge. Risk and precaution: Salmon farming. Marine Policy, 74(??): 205–212, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304390.

Malick:2017:CCI

[MRC17] Michael J. Malick, Murray B. Rutherford, and Sean P. Cox. Confronting challenges to integrating Pacific salmon into ecosystem-based management policies. Marine Policy, 85(??):123–132, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301902.

McClellan:2011:CCM

[MRCG11] Catherine M. McClellan, Andrew J. Read, Wendy M. Cluse, and Matthew H. Godfrey. Conservation in a complex management environment: the by-catch of sea turtles in North Carolina’s commercial fisheries. Marine Policy, 35(2):241–248, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001764.

McDonald:2015:PPC

[MRG15] Sara L. McDonald and Deborah Rigling-Gallagher. Participant perceptions of consensus-based, marine mammal take reduction planning. Marine Policy, 61(??): 216–226, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002237.

Matear:2019:CBB

[MRHP19] Liam Matear, James R. Robbins, Michelle Hale, and Jonathan Potts. Cetacean biodiversity in the Bay of Biscay: Suggestions for environmental protection derived from citizen science data. Marine Policy, 109(??):Article 103672, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302313.

Metcalfe:2013:MCS

[MRSH13] Kristian Metcalfe, Thomas Roberts, Robert J. Smith, and Stuart R. Harrop. Marine conservation science and governance in North–West Europe: Conservation planning and international law and policy. Marine Policy, 39(??):289–295, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002436.

McCain:2016:HAJ

[MRSL16] J. Scott P. McCain, Robert W. Rangeley, David C. Schneider, and Heike K. Lotze. Historical abundance of juvenile commercial fish in coastal habitats: Implications for fish habitat management in Canada. Marine Policy, 73(??):235–243, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302305.

Miller:2014:FUB

[MS14] Dana D. Miller and U. Rashid Sumaila. Flag use behavior and IUU activity within the international fishing fleet: Refining definitions and identifying areas of concern. Marine Policy, 44(??):204–211, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001930.

Milewski:2019:SAC

[MS19] Inka Milewski and Ruth E. Smith. Sustainable aquaculture in Canada: Lost in translation. Marine Policy, 107(??):Article 103571, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301332.

Masud:2018:CRF

[MSA^{+}18] Muhammad Mehedi Masud, Ahmad S. Sackor, A. S. A. Ferdous Alam, Abul Quasem Al-Amin, and Ahmad Bashawir Abdul Ghani. Community responses to flood risk management — an empirical investigation of the Marine Protected Areas (MPAs) in Malaysia. Marine Policy, 97(??):119–126, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300319.

McCagh:2015:KSM

[MSB15] Christine McCagh, Joanne Sneddon, and Dominque Blache. Killing sharks: the media’s role in public and political response to fatal human–shark interactions. Marine Policy, 62(??):271–278, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002675.

Mavruk:2018:GFN

[MSB^{+}18] Sinan Mavruk, İsmet Saygu, Fethi Bengil, Vahit Alan, and Ernesto Azzurro. Grouper fishery in the Northeastern Mediterranean: an assessment based on interviews on resource users. Marine Policy, 87(??):141–148, January 2018.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304517.

Mabon:2014:PSS

[MSBB14] Leslie Mabon, Simon Shackley, and Nathan Bower-Bir. Perceptions of sub-seabed carbon dioxide storage in Scotland and implications for policy: a qualitative study. Marine Policy, 45(??):9–15, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002662.

McComas:2015:CAM

[MSBR15] Katherine A. McComas, Jonathon P. Schuldt, Colleen A. Burge, and Sungjong Roh. Communicating about marine disease: the effects of message frames on policy support. Marine Policy, 57(??):45–52, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000470.

Mouro:2018:PPD

[MSC18] Carla Mouro, Tânia Santos, and Paula Castro. Pastpresent discontinuity in ecological change and marine governance: an integrated narrative approach to artisanal fishing. Marine Policy, 97(??):163–169, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307534.

Marques:2016:APA

[MSdAM16] Simone Marques, Andrea Q. Steiner, and Marcelo de Almeida Medeiros. Assessing the performance of the Aichi Biodiversity Targets in Brazil: a test using two regional-scale indices related to coastal and marine ecosystem conservation. Marine Policy, 67(??):130–138, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000439.

Munoz:2011:OCW

[MSG11] P. Durán Muñoz and M. Sayago-Gil. An overview of cold-water coral protection on the high seas: the Hatton bank (NE Atlantic) — a case study. Marine Policy, 35(5):615–622, September 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000248.

Murphy:2018:POK

[MSG18] Robert Murphy, Steven Scyphers, and Jonathan Grabowski. Perceptions outweigh knowledge in predicting support for management strategies in the recreational Striped Bass (Morone saxatilis) fishery. Marine Policy, 97(??): 44–50, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308795.

Munoz:2012:ATF

[MSGM^{+}12] P. Durán Muñoz, M. Sayago-Gil, F. J. Murillo, J. L. Del Río, L. J. López-Abellán, M. Sacau, and R. Sarralde. Actions taken by fishing nations towards identification and protection of vulnerable marine ecosystems in the high seas: the Spanish case (Atlantic Ocean). Marine Policy, 36(2):536–543, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001503.

Miller:2013:MCS

[MSH13] Denzil G. M. Miller, Natasha M. Slicer, and Quentin Hanich. Monitoring, control and surveillance of protected areas and specially managed areas in the marine domain. Marine Policy, 39(??):64–71, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002023.

Msomphora:2016:CRD

[Mso16] Mbachi Ruth Msomphora. Conflict resolution and the delegation of authority in fisheries management: the case of Outer Hebrides Inshore Fisheries Group in Scotland.

Marine Policy, 73(??):263–275, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301233.

Mohammed:2018:FRS

[MSS18] Essam Yassin Mohammed, Dave Steinbach, and Paul Steele. Fiscal reforms for sustainable marine fisheries governance: Delivering the SDGs and ensuring no one is left behind. Marine Policy, 93(??):262–270, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301574.

Manejar:2019:LMB

[MSS19] Arvie Joy A. Manejar, Liza May H. Sandoy, and Rodelio F. Subade. Linking marine biodiversity conservation and poverty alleviation: a case study in selected rural communities of Sagay Marine Reserve, Negros Occidental. Marine Policy, 104(??):12–18, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830592X.

Mallin:2019:OWT

[MSTB19] Marc-Andrej Felix Mallin, Dennis C. Stolz, Benjamin S. Thompson, and Mads Barbesgaard. In oceans we trust: Conservation, philanthropy, and the political economy of the Phoenix Islands Protected Area. Marine Policy, 107(??):Article 103421, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304214.

McCrea-Strub:2011:UCE

[MSZS^{+}11] Ashley McCrea-Strub, Dirk Zeller, Ussif Rashid Sumaila, Jay Nelson, Andrew Balmford, and Daniel Pauly. Understanding the cost of establishing marine protected areas. Marine Policy, 35(1):1–9, January 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001338.

Martin:2015:GMA

[MTF^{+}15] C. S. Martin, M. J. Tolley, E. Farmer, C. J. Mcowen, J. L. Geffert, J. P. W. Scharlemann, H. L. Thomas, J. H. van Bochove, D. Stanwell-Smith, J. M. Hutton, B. Lascelles, J. D. Pilgrim, J. M. M. Ekstrom, and D. P. Tittensor. A global map to aid the identification and screening of critical habitat for marine industries. Marine Policy, 53(??):45–53, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002991.

Marre:2015:UES

[MTP^{+}15] Jean-Baptiste Marre, Olivier Thebaud, Sean Pascoe, Sarah Jennings, Jean Boncoeur, and Louisa Coglan. The use of ecosystem services valuation in Australian coastal zone management. Marine Policy, 56(??):117–124, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000469.

Mills:2017:LDD

[MTP^{+}17] David J. Mills, Alexander Tilley, Mario Pereira, Denis Hellebrandt, Avelino Pereira Fernandes, and Philippa J. Cohen. Livelihood diversity and dynamism in Timor-Leste; insights for coastal resource governance and livelihood development. Marine Policy, 82(??):206–215, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302695.

Mortensen:2018:ICS

[MUHH18] Lars O. Mortensen, Clara Ulrich, Jan Hansen, and Rasmus Hald. Identifying choke species challenges for an individual demersal trawler in the North Sea, lessons from conversations and data analysis. Marine Policy, 87(??):1–11, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303809.

Mukherjee:2015:EAU

[Muk15] Zinnia Mukherjee. An economic approach to understanding the international transfer of bycatch from unilateral bycatch reduction policies. Marine Policy, 51(??):

190–195, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001985.

Mullie:2019:ARI

[Mul19] Wim C. Mullié. Apparent reduction of illegal trawler fishing effort in Ghana’s Inshore Exclusive Zone 2012– 2018 as revealed by publicly available AIS data. Marine Policy, 108(??):Article 103623, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300867.

Marschke:2016:SSU

[MV16] Melissa Marschke and Peter Vandergeest. Slavery scandals: Unpacking labour challenges and policy responses within the off-shore fisheries sector. Marine Policy, 68(??):39–46, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000622.

Mizuta:2017:SCI

[MV17] Darien Danielle Mizuta and Eirini Ioanna Vlachopoulou. Satoumi concept illustrated by sustainable bottom-up initiatives of Japanese Fisheries Cooperative Associations. Marine Policy, 78(??):143–149, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304791.

Menini:2019:APH

[MV19] Elisabetta Menini and Cindy Lee Van Dover. An atlas of protected hydrothermal vents. Marine Policy, 108(??): Article 103654, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18309394.

Macedo:2013:GWC

[MVRG13] Heitor Schulz Macedo, Melissa Vivacqua, Helio Castro Lima Rodrigues, and Leopoldo Cavaleri Gerhardinger. Governing wide coastal-marine protected territories: a governance analysis of the Baleia Franca Environmental Protection Area in South Brazil. Marine Policy, 41(??):

118–125, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000171.

Mackinson:2011:ESF

[MWGD11] S. Mackinson, D. C. Wilson, P. Galiay, and B. Deas. Engaging stakeholders in fisheries and marine research. Marine Policy, 35(1):18–24, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X10001375.

Martin:2017:DRT

[MWR^{+}17] Victoria Y. Martin, Betty Weiler, Arianne Reis, Kay Dimmock, and Pascal Scherrer. ‘Doing the right thing’: How social science can help foster pro-environmental behaviour change in marine protected areas. Marine Policy, 81 (??):236–246, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300751.

Muawanah:2018:RNL

[MYA^{+}18] Umi Muawanah, Gellwynn Yusuf, Luky Adrianto, Jimy Kalther, Robert Pomeroy, Habibi Abdullah, and Toni Ruchimat. Review of national laws and regulation in Indonesia in relation to an ecosystem approach to fisheries management. Marine Policy, 91(??):150–160, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305468.

Nielsen:2018:MPF

[NAB^{+}18] Max Nielsen, Frank Asche, Ole Bergesen, Johan Blomquist, Edgar Henriksen, Ayoe Hoff, Rasmus Nielsen, Jónas R. Vidarsson, and Staffan Waldo. The myth of the poor fisher: Evidence from the Nordic countries. Marine Policy, 93(??):186–194, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306693.

Nobre:2017:GCE

[NACS17] Danieli Marinho Nobre, Daniela Trigueirinho Alarcon, Ana Cinti, and Alexandre Schiavetti. Governance of the Cassurubá Extractive Reserve, Bahia State, Brazil: an analysis of strengths and weaknesses to inform policy. Marine Policy, 77(??):44–55, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308077.

Naevestad:2017:SCW

[Næv17] Tor-Olav Nævestad. Safety culture, working conditions and personal injuries in Norwegian maritime transport. Marine Policy, 84(??):251–262, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301136.

Nagasaka:2016:VHT

[Nag16] Naohiko Nagasaka. Visualising historical trends in global maritime boundary delimitations since the 1940s. Marine Policy, 71(??):29–37, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300689.

Nakhshina:2016:CCP

[Nak16] Maria Nakhshina. Constraints on community participation in salmon fisheries management in Northwest Russia. Marine Policy, 74(??):309–315, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630238X.

Nieminen:2019:EBA

[NALO19] Emmi Nieminen, Heini Ahtiainen, Carl-Johan Lagerkvist, and Soile Oinonen. The economic benefits of achieving Good Environmental Status in the Finnish marine waters of the Baltic Sea. Marine Policy, 99(??):181–189, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308102.

Nielsen:2018:PTT

[NAYSN18] Max Nielsen, Isaac Ankamah-Yeboah, Lisa Staahl, and Rasmus Nielsen. Price transmission in the trans-Atlantic northern shrimp value chain. Marine Policy, 93(??):71–79, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300435.

Norse:2012:SDS

[NBC^{+}12] Elliott A. Norse, Sandra Brooke, William W. L. Cheung, Malcolm R. Clark, Ivar Ekeland, Rainer Froese, Kristina M. Gjerde, Richard L. Haedrich, Selina S. Heppell, Telmo Morato, Lance E. Morgan, Daniel Pauly, Rashid Sumaila, and Reg Watson. Sustainability of deep-sea fisheries. Marine Policy, 36(2):307–320, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001102.

Nilssen:2015:SCA

[NBD15] Jon Nilssen, Bernt Arne Bertheussen, and Bent Dreyer. Sustained competitive advantage based on high quality input. Marine Policy, 52(??):145–154, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002735.

Newman:2018:RAP

[NBF^{+}18] Stephen J. Newman, Joshua I. Brown, David V. Fairclough, Brent S. Wise, Lynda M. Bellchambers, Brett W. Molony, Rodney C. J. Lenanton, Gary Jackson, Kim A. Smith, Daniel J. Gaughan, Warrick (Rick) J. Fletcher, Rory B. McAuley, and Corey B. Wakefield. A risk assessment and prioritisation approach to the selection of indicator species for the assessment of multi-species, multi-gear, multi-sector fishery resources. Marine Policy, 88(??):11–22, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17305596.

Nursey-Bray:2018:DCM

[NBFO18] Melissa Nursey-Bray, Pedro Fidelman, and Mensah Owusu. Does co-management facilitate adaptive capacity in times of environmental change? Insights from fisheries in Australia. Marine Policy, 96(??):72–80, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301490.

Natale:2015:ADI

[NBM15] Fabrizio Natale, Alessandra Borrello, and Arina Motova. Analysis of the determinants of international seafood trade using a gravity model. Marine Policy, 60(??):98–106, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1500158X.

Nursey-Bray:2018:ACP

[NBMB^{+}18] Melissa Nursey-Bray, Anders Magnusson, Nathan Bicknell, Meagan Magnusson, Julian Morison, and Andrew Sullivan. Adapting to change: Prioritising management for the future of the Marine Scalefish Fishery. Marine Policy, 95(??):153–165, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308771.

Nursey-Bray:2012:CCC

[NBPF^{+}12] Melissa Nursey-Bray, G. T. Pecl, S. Frusher, C. Gardner, M. Haward, A. J. Hobday, S. Jennings, A. E. Punt, H. Revill, and I. van Putten. Communicating climate change: Climate change risk perceptions and rock lobster fishers, Tasmania. Marine Policy, 36(3):753–759, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001655.

Nasuchon:2010:CIF

[NC10] Nopparat Nasuchon and Anthony Charles. Community involvement in fisheries management: Experiences in the Gulf of Thailand countries. Marine Policy, 34 (1):163–169, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000918.

Natale:2013:IFD

[NCH^{+}13] Fabrizio Natale, Natacha Carvalho, Michael Harrop, Jordi Guillen, and Katia Frangoudes. Identifying fisheries dependent communities in EU coastal areas. Marine Policy, 42(??):245–252, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000948.

Nicoll:2017:CAI

[ND17] Rob Nicoll and Jon C. Day. Correct application of the IUCN protected area management categories to the CCAMLR Convention Area. Marine Policy, 77(??):9–12, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304614.

Neis:2015:ISI

[Nei15] Barbara Neis. Introduction to the special issue marine policy science, state and fishers knowledge in the Benguela ecosystem in Southern Africa. Marine Policy, 60(??): 287–292, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002280.

Neilson:2018:CIM

[Nei18] Alasdair Neilson. Considering the importance of metaphors for marine conservation. Marine Policy, 97(??):239– 243, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301556.

Nadarajah:2017:GAG

[NF17] Suthamathy Nadarajah and Ola Flaaten. Global aquaculture growth and institutional quality. Marine Policy, 84 (??):142–151, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302142.

Nilsson:2016:CMA

[NFHJ16] Jessica A. Nilsson, Elizabeth A. Fulton, Marcus Haward, and Craig Johnson. Consensus management in Antarctica’s high seas — Past success and current challenges. Marine Policy, 73(??):172–180, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304389.

Ngoc:2018:IEH

[Ngo18] Quach Thi Khanh Ngoc. Impacts on the ecosystem and human well-being of the marine protected area in Cu Lao Cham, Vietnam. Marine Policy, 90(??):174–183, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302701.

Neto:2016:RSB

[NGS16] Joaquim R. Santos Neto, Vinicius J. Giglio, and Alexandre Schiavetti. A reply to Santos and Brannstrom: Livelihood strategies in a marine extractive reserve: Implications for conservation interventions. Marine Policy, 71 (??):71–72, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300240.

Nielsen:2015:RBM

[NHA15] Kåre Nolde Nielsen, Petter Holm, and Michaela Aschan. Results based management in fisheries: Delegating responsibility to resource users. Marine Policy, 51(??): 442–451, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002693.

Natale:2013:IBA

[NHFV13] Fabrizio Natale, Johann Hofherr, Gianluca Fiore, and Jarno Virtanen. Interactions between aquaculture and fisheries. Marine Policy, 38(??):205–213, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001352.

Noble:2019:LSE

[NHPD19] Mae M. Noble, David Harasti, Jamie Pittock, and Bruce Doran. Linking the social to the ecological using GIS methods in marine spatial planning and management to support resilience: a review. Marine Policy, 108(??):Article 103657, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306158.

Nakatsuka:2017:LRP

[NIFA17] Shuya Nakatsuka, Yukimasa Ishida, Hiromu Fukuda, and Tetsuya Akita. A limit reference point to prevent recruitment overfishing of Pacific bluefin tuna. Marine Policy, 78(??):107–113, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304973.

Nijman:2015:CLE

[Nij15] Vincent Nijman. CITES-listings, EU eel trade bans and the increase of export of tropical eels out of Indonesia. Marine Policy, 58(??):36–41, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000883.

Nijman:2017:NAS

[Nij17] Vincent Nijman. North Africa as a source for European eel following the 2010 EU CITES eel trade ban. Marine Policy, 85(??):133–137, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307953.

Nikcevic:2019:SRL

[Nik19] Jelena Nikčević. Strengthening the role of local government to ensure sustainable development of the cruise sector: the case of Kotor. Marine Policy, 109(??):Article 103693, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1930332X.

Nguyen:2018:BIA

[NJ18] Kim Anh Thi Nguyen and Curtis M. Jolly. Balancing interests of actors in the ocean tuna value chain of Khanh Hoa province, Vietnam. Marine Policy, 98(??):11–22, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308916.

Noble:2013:SEA

[NKAP13] Bram Noble, Skye Ketilson, Alec Aitken, and Greg Poelzer. Strategic environmental assessment opportunities and risks for Arctic offshore energy planning and development. Marine Policy, 39(??):296–302, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002527.

Nomura:2017:CAF

[NKBS17] Keiko J. Nomura, David M. Kaplan, Jennifer Beckensteiner, and Andrew M. Scheld. Comparative analysis of factors influencing spatial distributions of marine protected areas and territorial use rights for fisheries in Japan. Marine Policy, 82(??):59–67, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630759X.

Nordlund:2013:CIC

[NKCR13] Lina M. Nordlund, Ulrike Kloiber, Eleanor Carter, and Sibylle Riedmiller. Chumbe Island Coral Park — governance analysis. Marine Policy, 41(??):110–117, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1200259X.

Navarro:2018:RFS

[NKHL18] Matthew Navarro, Marit E. Kragt, Atakelty Hailu, and Tim J. Langlois. Recreational fishers’ support for notake marine reserves is high and increases with reserve age. Marine Policy, 96(??):44–52, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301611.

Norman-Lopez:2011:NEE

[NLP11] Ana Norman-López and Sean Pascoe. Net economic effects of achieving maximum economic yield in fisheries. Marine Policy, 35(4):489–495, July 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002149.

Noye:2012:FST

[NM12] Jeremy Noye and Kwame Mfodwo. First steps towards a quota allocation system in the Indian Ocean. Marine Policy, 36(4):882–894, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11002016.

Ntona:2018:CSO

[NM18] Mara Ntona and Elisa Morgera. Connecting SDG 14 with the other Sustainable Development Goals through marine spatial planning. Marine Policy, 93(??):214–222, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300520.

Neilson:2019:RMR

[NM19] Alison Laurie Neilson and Rita São Marcos. Reframing marine resource management with relational ontologies and hybrid entanglements: Fishing for empathy between Azorean fishers and scientists. Marine Policy, 105(??):30– 37, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301970.

Niner:2017:GSM

[NMJS17] Holly J. Niner, Ben Milligan, Peter J. S. Jones, and Craig A. Styan. A global snapshot of marine biodiversity offsetting policy. Marine Policy, 81(??):368–374, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306091.

Naranjo-Madrigal:2019:LBR

[NMvP19] Helven Naranjo-Madrigal and Ingrid van Putten. The link between risk taking, fish catches, and social standing: Untangling a complex cultural landscape. Marine Policy, 100(??):173–182, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300721.

Naranjo-Madrigal:2015:USE

[NMvPNL15] Helven Naranjo-Madrigal, Ingrid van Putten, and Ana Norman-López. Understanding socio-ecological drivers of spatial allocation choice in a multi-species artisanal fishery: a Bayesian network modeling approach. Marine Policy, 62(??):102–115, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002547.

Nuttall:2014:RSS

[NNP^{+}14] Peter Nuttall, Alison Newell, Biman Prasad, Joeli Veitayaki, and Elisabeth Holland. A review of sustainable sea-transport for Oceania: Providing context for renewable energy shipping for the Pacific. Marine Policy, 43(??):283–287, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001358.

Newell:2017:TTN

[NNPV17] Alison Newell, Peter Nuttall, Biman Prasad, and Joeli Veitayaki. Turning the Tide: the need for sustainable sea transport in the Pacific. Marine Policy, 75(??): 249–259, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000142.

Nolan:2019:PAI

[Nol19] Callum Nolan. Power and access issues in Ghana’s coastal fisheries: a political ecology of a closing commodity frontier. Marine Policy, 108(??):Article 103621, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300879.

Norton:2015:MMR

[Nor15] Marieke Norton. The militarisation of marine resource conservation and law enforcement in the Western Cape, South Africa. Marine Policy, 60(??):338–344, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002401.

Nittari:2019:RMA

[NPAR19] Giulio Nittari, Isabel Pirillo, Francesco Amenta, and Giovanna Ricci. The right to medical assistance for seafarers. Ethical and practical consequences of the introduction of telemedicine to improve healthcare on board ships. Marine Policy, 106(??):Article 103525, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306498.

Naevestad:2019:SCM

[NPS^{+}19] Tor-Olav Nævestad, Ross O. Phillips, Kristine V. Størkersen, Alexandra Laiou, and George Yannis. Safety culture in maritime transport in Norway and Greece: Exploring national, sectorial and organizational influences on unsafe behaviours and work accidents. Marine Policy, 99 (??):1–13, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301106.

Niiranen:2018:GCC

[NRB^{+}18] S. Niiranen, A. Richter, T. Blenckner, L. C. Stige, M. Valman, and A.-M. Eikeset. Global connectivity and crossscale interactions create uncertainty for Blue Growth of Arctic fisheries. Marine Policy, 87(??):321–330, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306917.

Nielsen:2014:GGF

[NRN14] Max Nielsen, Lars Ravensbeck, and Rasmus Nielsen. Green growth in fisheries. Marine Policy, 46(??):43–52, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000050.

Nilsen:2018:PPC

[NS18] Marie Nilsen and Kristine Vedal Størkersen. Permitted to be powerful? A comparison of the possibilities to regulate safety in the Norwegian petroleum and maritime industries. Marine Policy, 92(??):30–39, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307753.

Nahuelhual:2018:SFB

[NSB^{+}18a] L. Nahuelhual, G. Saavedra, G. Blanco, E. Wesselink, G. Campos, and X. Vergara. On super fishers and black capture: Images of illegal fishing in artisanal fisheries of southern Chile. Marine Policy, 95(??):36–45, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301404.

Nunny:2018:RSK

[NSB18b] Laetitia Nunny, Mark P. Simmonds, and Andrew Butterworth. A review of seal killing practice in Europe: Implications for animal welfare. Marine Policy, 98(??): 121–132, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307303.

Nahuelhual:2017:MES

[NVK^{+}17] L. Nahuelhual, X. Vergara, A. Kusch, G. Campos, and D. Droguett. Mapping ecosystem services for marine spatial planning: Recreation opportunities in Sub-Antarctic Chile. Marine Policy, 81(??):211–218, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308296.

Njock:2010:MRM

[NW10] Jean-Calvin Njock and Lena Westlund. Migration, resource management and global change: Experiences from fishing communities in West and Central Africa. Marine Policy, 34(4):752–760, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000217.

Nikitine:2018:DFE

[NWD18] J. Nikitine, A. M. W. Wilson, and T. P. Dawson. Developing a framework for the efficient design and management of large scale marine protected areas. Marine Policy, 94(??):196–203, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830109X.

Newman:2015:IWM

[NWW^{+}15] Stephen J. Newman, Corey B. Wakefield, Ashley J. Williams, Joseph M. O’Malley, Simon J. Nicol, Edward E. DeMartini, Tuikolongahau Halafihi, Jeremie Kaltavara, Robert L. Humphreys, Brett M. Taylor, Allen H. Andrews, and Ryan S. Nichols. International workshop on methodological evolution to improve estimates of life history parameters and fisheries management of data-poor deep-water snappers and groupers. Marine Policy, 60(??):182–185, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001797.

Newman:2017:IWA

[NWW^{+}17] Stephen J. Newman, Corey B. Wakefield, Ashley J. Williams, Joseph M. O’Malley, Brett M. Taylor, Simon J. Nicol, Ryan S. Nichols, S. Alex Hesp, Norm G. Hall, Nicholas Hill, Joyce J. L. Ong, Allen H. Andrews, Claire M. Wellington, Euan S. Harvey, Peter Mous, Zack S. Oyafuso, Cassandra Pardee, Michael Bunce, Joseph D. DiBattista, and Bradley R. Moore. International workshop on advancing methods to overcome challenges associated with life history and stock assessments of data-poor deep-water snappers and groupers. Marine Policy, 79(??):78–83, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630700X.

Neff:2013:SBP

[NY13] Christopher L. Neff and Jean Y. H. Yang. Shark bites and public attitudes: Policy implications from the first before and after shark bite survey. Marine Policy, 38(??):545–547, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1200156X.

Nichols:2015:FAA

[NYJW15] Rachel Nichols, Satoshi Yamazaki, Sarah Jennings, and Reg A. Watson. Fishing access agreements and harvesting decisions of host and distant water fishing nations. Marine Policy, 54(??):77–85, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400356X.

Nyman:2019:TOO

[Nym19] Elizabeth Nyman. Techno-optimism and ocean governance: New trends in maritime monitoring. Marine Policy, 99 (??):30–33, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303750.

Oanta:2018:IOD

[Oan18] Gabriela A. Oanta. International organizations and deep-sea fisheries: Current status and future prospects. Marine Policy, 87(??):51–59, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302816.

Oppel:2018:SSM

[OBC^{+}18] Steffen Oppel, Mark Bolton, Ana P. B. Carneiro, Maria P. Dias, Jonathan A. Green, Juan F. Masello, Richard A. Phillips, Ellie Owen, Petra Quillfeldt, Annalea Beard, Sophie Bertrand, Jez Blackburn, P. Dee Boersma, Alder Borges, Annette C. Broderick, Paulo Catry, Ian Cleasby, Elizabeth Clingham, Jeroen Creuwels, Sarah Crofts, Richard J. Cuthbert, Hanneke Dallmeijer, Delia Davies, Rachel Davies, Ben J. Dilley, Herculano Andrade Dinis, Justine Dossa, Michael J. Dunn, Marcio A. Efe, Annette L. Fayet, Leila Figueiredo, Adelcides Pereira Frederico, Carina Gjerdrum, Brendan J. Godley, José Pedro Granadeiro, Tim Guilford, Keith C. Hamer, Carolina Hazin, April Hedd, Leeann Henry, Marcos Hernández-Montero, Jefferson Hinke, Nobuo Kokubun, Eliza Leat, Laura Mc-Farlane Tranquilla, Benjamin Metzger, Teresa Militão, Gilson Montrond, Wim Mullié, Oliver Padget, Elizabeth J. Pearmain, Ingrid L. Pollet, Klemens Pütz, Flavio Quintana, Norman Ratcliffe, Robert A. Ronconi, Peter G. Ryan, Sarah Saldanha, Akiko Shoji, Jolene Sim, Cleo Small, Louise Soanes, Akinori Takahashi, Phil Trathan, Wayne Trivelpiece, Jan Veen, Ewan Wakefield, Nicola Weber, Sam Weber, Laura Zango, Francis Daunt, Motohiro Ito, Michael P. Harris, Mark A. Newell, Sarah Wanless, Jacob González-Solís, and John Croxall. Spatial scales of marine conservation management for breeding seabirds. Marine Policy, 98(??):37–46, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302422.

OLeary:2015:BPS

[OBH15] Bethan C. O’Leary, Helen R. Bayliss, and Neal R. Haddaway. Beyond PRISMA: Systematic reviews to inform marine science and policy. Marine Policy, 62(??):261–263, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002778.

OLeary:2012:FNM

[OBJ^{+}12] B. C. O’Leary, R. L. Brown, D. E. Johnson, H. von Nordheim, J. Ardron, T. Packeiser, and C. M. Roberts. The first network of marine protected areas (MPAs) in the high seas: the process, the challenges and where next. Marine Policy, 36(3):598–605, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001837.

Oliver:2015:GPB

[OBNH15] Shelby Oliver, Matias Braccini, Stephen J. Newman, and Euan S. Harvey. Global patterns in the bycatch of sharks and rays. Marine Policy, 54(??):86–97, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003546.

ONeill:2017:ANS

[OC17] Elizabeth Drury O’Neill and Beatrice Crona. Assistance networks in seafood trade — a means to assess benefit distribution in small-scale fisheries. Marine Policy, 78 (??):196–205, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306443.

Ounanian:2018:GME

[OCCvT^{+}18] Kristen Ounanian, Eira Carballo-Cárdenas, Jan P. M. van Tatenhove, Alyne Delaney, K. Nadia Papadopoulou, and Christopher J. Smith. Governing marine ecosystem restoration: the role of discourses and uncertainties. Marine Policy, 96(??):136–144, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302033.

Olson:2014:PSS

[OCdS14] Julia Olson, Patricia M. Clay, and Patricia Pinto da Silva. Putting the seafood in sustainable food systems. Marine Policy, 43(??):104–111, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300105X.

Oliveira:2014:EPQ

[OCG14] M. M. Oliveira, A. S. Camanho, and M. B. Gaspar. Enhancing the performance of quota managed fisheries using seasonality information: the case of the Portuguese artisanal dredge fleet. Marine Policy, 45(??):114–120, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002698.

Oliveira:2015:PIR

[OCG15] M. M. Oliveira, A. S. Camanho, and M. B. Gaspar. The phycotoxins’ impact on the revenue of the Portuguese artisanal dredge fleet. Marine Policy, 52(??):45–51, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1400284X.

Oliveira:2017:FBL

[OCW^{+}17] Manuela M. Oliveira, Ana S. Camanho, John B. Walden, Vera L. Miguéis, Nuno B. Ferreira, and Miguel B. Gaspar. Forecasting bivalve landings with multiple regression and data mining techniques: the case of the Portuguese Artisanal Dredge Fleet. Marine Policy, 84(??):110–118, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303159.

Ogier:2016:FMA

[ODF^{+}16] Emily M. Ogier, Julie Davidson, Pedro Fidelman, Marcus Haward, Alistair J. Hobday, Neil J. Holbrook, Eriko Hoshino, and Gretta T. Pecl. Fisheries management approaches as platforms for climate change adaptation: Comparing theory and practice in Australian fisheries. Marine Policy, 71(??):82–93, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302962.

Ovando:2013:CIC

[ODL^{+}13] Daniel A. Ovando, Robert T. Deacon, Sarah E. Lester, Christopher Costello, Tonya Van Leuvan, Karlynn McIlwain, C. Kent Strauss, Michael Arbuckle, Rod Fujita, Stefan Gelcich, and Hirotsugu Uchida. Conservation incentives and collective choices in cooperative fisheries. Marine Policy, 37(??):132–140, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000565.

Ounanian:2012:UFS

[ODRRM12] K. Ounanian, A. Delaney, J. Raakjær, and P. Ramirez-Monsalve. On unequal footing: Stakeholder perspectives on the marine strategy framework directive as a mechanism of the ecosystem-based approach to marine management. Marine Policy, 36(3):658–666, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001588.

Osterblom:2010:MEA

[ÖGB^{+}10] H. Österblom, A. Gårdmark, L. Bergström, B. Müller-Karulis, C. Folke, M. Lindegren, M. Casini, P. Olsson, R. Diekmann, T. Blenckner, C. Humborg, and C. Möllmann. Making the ecosystem approach operational — Can regime shifts in ecological- and governance systems facilitate the transition? Marine Policy, 34(6):

1290–1299, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001053.

Ojaveer:2014:TRA

[OGM^{+}14] Henn Ojaveer, Bella S. Galil, Dan Minchin, Sergej Olenin, Ana Amorim, João Canning-Clode, Paula Chainho, Gordon H. Copp, Stephan Gollasch, Anders Jelmert, Maiju Lehtiniemi, Cynthia McKenzie, Josip Mikuš, Laurence Miossec, Anna Occhipinti-Ambrogi, Marijana Pećarević, Judith Pederson, Gemma Quilez-Badia, Jeroen W. M. Wijsman, and Argyro Zenetos. Ten recommendations for advancing the assessment and management of non-indigenous species in marine ecosystems. Marine Policy, 44(??):160–165, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001735.

Ovesen:2018:MDS

[OHB^{+}18] Vidar Ovesen, Ron Hackett, Lee Burns, Peter Mullins, and Scott Roger. Managing deep sea mining revenues for the public good- ensuring transparency and distribution equity. Marine Policy, 95(??):332–336, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301264.

Olsen:2016:HIO

[OHH^{+}16] Erik Olsen, Silje Holen, Alf Håkon Hoel, Lene Buhl-Mortensen, and Ingolf Røttingen. How Integrated Ocean governance in the Barents Sea was created by a drive for increased oil production. Marine Policy, 71(??): 293–300, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15003681.

Opitz:2016:AMC

[OHQ^{+}16] Silvia Opitz, Julia Hoffmann, Martin Quaas, Nele Matz-Lück, Crispina Binohlan, and Rainer Froese. Assessment of MSC-certified fish stocks in the Northeast Atlantic. Marine Policy, 71(??):10–14, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300082.

Oozeki:2018:REI

[OIS^{+}18] Yoshioki Oozeki, Denzou Inagake, Tsutomu Saito, Makoto Okazaki, Ippei Fusejima, Makoto Hotai, Tomowo Watanabe, Hiroya Sugisaki, and Masanori Miyahara. Reliable estimation of IUU fishing catch amounts in the northwestern Pacific adjacent to the Japanese EEZ: Potential for usage of satellite remote sensing images. Marine Policy, 88(??):64–74, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303998.

Ojea:2010:VRO

[OL10] Elena Ojea and Maria L. Loureiro. Valuing the recovery of overexploited fish stocks in the context of existence and option values. Marine Policy, 34(3):514–521, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001699.

OHagan:2011:ELP

[OL11] A. M. O’Hagan and A. W. Lewis. The existing law and policy framework for ocean energy development in Ireland. Marine Policy, 35(6):772–783, December 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000054.

Olaussen:2018:EPR

[Ola18] Jon Olaf Olaussen. Environmental problems and regulation in the aquaculture industry. Insights from Norway. Marine Policy, 98(??):158–163, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300794.

Oyafuso:2019:EBT

[OLF19] Zack S. Oyafuso, PingSun Leung, and Erik C. Franklin. Evaluating bioeconomic tradeoffs of fishing reserves via spatial optimization. Marine Policy, 100(??):163– 172, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301945.

Olson:2011:PNE

[Ols11] Julia Olson. Producing nature and enacting difference in ecosystem-based fisheries management: an example from the Northeastern US. Marine Policy, 35(4):528–535, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000170.

OHara:2019:FIS

[OM19] Jeffrey K. O’Hara and Loren McClenachan. Factors influencing ‘Sea to School’ purchases of local seafood products. Marine Policy, 100(??):76–82, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830486X.

Olsen:2014:SRI

[OMF14] Stephen B. Olsen, Jennifer H. McCann, and Grover Fugate. The State of Rhode Island’s pioneering marine spatial plan. Marine Policy, 45(??):26–38, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002455.

Ourens:2015:MBB

[ONF15] Rosana Ouréns, Inés Naya, and Juan Freire. Mismatch between biological, exploitation, and governance scales and ineffective management of sea urchin (Paracentrotus lividus) fisheries in Galicia. Marine Policy, 51(??): 13–20, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001894.

Olsen:2017:MFA

[OO17a] Marit Schei Olsen and Tonje Cecilie Osmundsen. Media framing of aquaculture. Marine Policy, 76(??): 19–27, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304286.

Osmundsen:2017:ICA

[OO17b] Tonje C. Osmundsen and Marit Schei Olsen. The imperishable controversy over aquaculture. Marine Policy, 76 (??):136–142, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304353.

OHiggins:2019:RTC

[OO19] Tim O’Higgins and Anne Marie O’Hagan. A return to the tragedy of the commons? Brexit tradeoffs and spatial analysis, an Irish perspective. Marine Policy, 106(??):Article 103524, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300181.

Ottersen:2011:NPI

[OOvdM^{+}11] Geir Ottersen, Erik Olsen, Gro I. van der Meeren, Are Dommasnes, and Harald Loeng. The Norwegian plan for integrated ecosystem-based management of the marine environment in the Norwegian Sea. Marine Policy, 35(3):389–398, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002046.

OBryhim:2015:IKA

[OP15a] Jason R. O’Bryhim and E. C. M. Parsons. Increased knowledge about sharks increases public concern about their conservation. Marine Policy, 56(??):43–47, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000421.

ORyan:2015:PIS

[OP15b] Raúl O’Ryan and Mauricio Pereira. Participatory indicators of sustainability for the salmon industry: the case of Chile. Marine Policy, 51(??):322–330, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002437.

OBryhim:2016:ESS

[OPGL16] Jason R. O’Bryhim, E. C. M. Parsons, Michael P. Gilmore, and Stacey L. Lance. Evaluating support for shark conservation among artisanal fishing communities in Costa Rica. Marine Policy, 71(??):1–9, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630032X.

Okes:2012:EPC

[OPMB12] Nicola C. Okes, Samantha Petersen, Liziwe McDaid, and Janine Basson. Enabling people to create change: Capacity building for Ecosystem Approach to Fisheries (EAF) implementation in Southern Africa. Marine Policy, 36 (1):286–296, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001096.

ORegan:2015:HPM

[O’R15] Sacha M. O’Regan. Harvesters’ perspectives on the management of British Columbia’s giant red sea cucumber fishery. Marine Policy, 51(??):103–110, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001997.

Osthagen:2018:CHD

[ØR18] Andreas Østhagen and Andreas Raspotnik. Crab! How a dispute over snow crab became a diplomatic headache between Norway and the EU. Marine Policy, 98(??): 58–64, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302550.

Ohms:2019:FPA

[OR19] Verena R. Ohms and Jesper Raakjær. The future of the Pelagic Advisory Council: Repositioning the organization in the face of BREXIT. Marine Policy, 106(??): Article 103535, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18309217.

Owen:2012:DSC

[OS12] Nick A. Owen and Clive H. Schofield. Disputed South China Sea hydrocarbons in perspective. Marine Policy, 36(3):809–822, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001953.

Osterblom:2011:ISE

[ÖSS^{+}11] H. Österblom, M. Sissenwine, D. Symes, M. Kadin, T. Daw, and C. Folke. Incentives, social–ecological feedbacks and European fisheries. Marine Policy, 35(5):568–574, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000194.

Oremus:2014:RRU

[OSS14] Kimberly Lai Oremus, Lisa Suatoni, and Brad Sewell. The requirement to rebuild US fish stocks: Is it working? Marine Policy, 47(??):71–75, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000451.

Oreska:2017:BSA

[OTOL17] Matthew P. J. Oreska, Barry Truitt, Robert J. Orth, and Mark W. Luckenbach. The bay scallop (Argopecten irradians) industry collapse in Virginia and its implications for the successful management of scallop-seagrass habitats. Marine Policy, 75(??):116–124, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303529.

Okafor-Yarwood:2015:GBS

[OY15] Ifesinachi Okafor-Yarwood. The Guinea-Bissau–Senegal maritime boundary dispute. Marine Policy, 61(??):284– 290, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002274.

Okafor-Yarwood:2019:IUU

[OY19] Ifesinachi Okafor-Yarwood. Illegal, unreported and unregulated fishing, and the complexities of the sustainable development goals (SDGs) for countries in the Gulf of Guinea. Marine Policy, 99(??):414–422, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303445.

Oyarce:2018:SSA

[Oya18] Ximena Hinrichs Oyarce. Sponsoring States in the Area: Obligations, liability and the role of developing States. Marine Policy, 95(??):317–323, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303293.

Pitcher:2010:RCT

[PA10] Tony J. Pitcher and Cameron H. Ainsworth. Resilience to change in two coastal communities: Using the maximum dexterity fleet. Marine Policy, 34(4):810–814, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000230.

Prellezo:2012:REB

[PAA^{+}12] Raúl Prellezo, Paolo Accadia, Jesper L. Andersen, Bo S. Andersen, Erik Buisman, Alyson Little, J. Rasmus Nielsen, Jan Jaap Poos, Jeff Powell, and Christine Röckmann. A review of EU bio-economic models for fisheries: the value of a diversity of models. Marine Policy, 36(2):423–431, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001382.

Pittman:2015:GFC

[PAA^{+}15] Jeremy Pittman, Derek Armitage, Steven Alexander, Donovan Campbell, and Marium Alleyne. Governance fit for climate change in a Caribbean coastal-marine context. Marine Policy, 51(??):486–498, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002255.

Parks:2019:QMT

[PAC^{+}19] Melissa Parks, Austin Ahmasuk, Barry Compagnoni, Andrew Norris, and Roger Rufe. Quantifying and mitigating three major vessel waste streams in the northern Bering Sea. Marine Policy, 106(??):Article 103530, August 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308315.

Pita:2017:RMP

[PAD^{+}17] Pablo Pita, Iñaqui Artetxe, Hugo Diogo, Pedro Gomes, Ana Gordoa, Kieran Hyder, João Pereira, Cristina Pita, Mafalda Rangel, João Garcia-Rodrigues, Oscar Sagué, Pedro Veiga, José Vingada, and Sebastián Villasante. Research and management priorities for Atlantic marine recreational fisheries in Southern Europe. Marine Policy, 86(??):1–8, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303846.

Pan:2014:ECM

[Pan14] Minling Pan. Economic characteristics and management challenges of the Hawaii pelagic longline fisheries: Will a catch share program help? Marine Policy, 44(??): 18–26, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001620.

Papageorgiou:2016:MSP

[Pap16] Marilena Papageorgiou. Marine spatial planning and the Greek experience. Marine Policy, 74(??):18–24, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301506.

Parris:2010:TDT

[Par10] Hannah Parris. Tuna dreams and tuna realities: Defining the term “maximising economic returns from the tuna fisheries” in six Pacific Island states. Marine Policy, 34 (1):105–113, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000736.

Partelow:2015:KSO

[Par15] Stefan Partelow. Key steps for operationalizing social– ecological system framework research in small-scale fisheries: a heuristic conceptual approach. Marine Policy, 51 (??):507–511, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002383.

Piniella:2017:PSR

[PARD17] Francisco Piniella, Juan Ignacio Alcaide, and Emilio Rodríguez-Díaz. The Panama Ship Registry: 1917– 2017. Marine Policy, 77(??):13–22, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306996.

Pastoors:2014:EGN

[Pas14] M. A. Pastoors. Exponential growth in the number of words used for the European Common Fisheries Policy (CFP): Does better management require more text? Marine Policy, 46(??):101–104, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000104.

Paterson:2015:TTL

[Pat15] Barbara Paterson. Tracks, trawls and lines — Knowledge practices of skippers in the Namibian hake fisheries. Marine Policy, 60(??):309–317, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001912.

Pauly:2018:VMF

[Pau18] Daniel Pauly. A vision for marine fisheries in a global blue economy. Marine Policy, 87(??):371–374, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306929.

Patrick:2013:EEI

[PB13] Wesley S. Patrick and Lee R. Benaka. Estimating the economic impacts of bycatch in U.S. commercial fisheries. Marine Policy, 38(??):470–475, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001674.

Potts:2014:DMP

[PBJ^{+}14] Tavis Potts, Daryl Burdon, Emma Jackson, Jonathan Atkins, Justine Saunders, Emily Hastings, and Olivia Langmead. Do marine protected areas deliver flows of ecosystem services to support human welfare? Marine Policy, 44(??):139–148, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001656.

Piovano:2012:EBR

[PBSG12] Susanna Piovano, Giovanni Basciano, Yonat Swimmer, and Cristina Giacoma. Evaluation of a bycatch reduction technology by fishermen: a case study from Sicily. Marine Policy, 36(1):272–277, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001060.

Pettipas:2016:CPF

[PBW16] Shauna Pettipas, Meagan Bernier, and Tony R. Walker. A Canadian policy framework to mitigate plastic marine pollution. Marine Policy, 68(??):117–122, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300665.

Pettipas:2017:CCP

[PBW17] Shauna Pettipas, Meagan Bernier, and Tony R. Walker. Corrigendum to “A Canadian policy framework to mitigate plastic marine pollution” [Marine Policy 68 (2016) 117–122]. Marine Policy, 75(??):187, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302779.

Parsons:2011:SSD

[PC11] E. C. M. Parsons and L. A. Cornick. Sweeping scientific data under a polar bear skin rug: the IUCN and the proposed listing of polar bears under CITES Appendix I. Marine Policy, 35(5):729–731, September 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1100025X.

Parsons:2013:PPP

[PC13a] E. C. M. Parsons and L. A. Cornick. Politics, people and polar bears: a rebuttal of Clark et al. (2013). Marine Policy, 42(??):178–179, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000468.

Peterson:2013:WIA

[PC13b] Megan J. Peterson and Courtney Carothers. Whale interactions with Alaskan sablefish and Pacific halibut fisheries: Surveying fishermen perception, changing fishing practices and mitigation. Marine Policy, 42(??):315–324, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001048.

Pipitone:2018:RTP

[PC18a] Vito Pipitone and Francesco Colloca. Recent trends in the productivity of the Italian trawl fishery: the importance of the socio-economic context and overexploitation. Marine Policy, 87(??):135–140, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730252X.

Pomeroy:2018:PCM

[PC18b] Robert Pomeroy and Catherine A. Courtney. The Philippines context for marine tenure and small-scale fisheries. Marine Policy, 95(??):283–293, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830263X.

Patchell:2019:RIF

[PC19] Jerry Patchell and Christopher Cheng. Resilience of an inshore fishing population in Hong Kong: Paradox and potential for sustainable fishery policy. Marine Policy, 99 (??):157–169, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306747.

Pinto:2015:CEM

[PCC15] Hugo Pinto, Ana Rita Cruz, and Colin Combe. Cooperation and the emergence of maritime clusters in the Atlantic: Analysis and implications of innovation and human capital for blue growth. Marine Policy, 57(??):167–177, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000780.

Perez-Cayeiro:2015:EPI

[PCCR15] María Luisa Pérez-Cayeiro and Juan Adolfo Chica-Ruiz. Evaluation of a programme of integrated coastal zone management: the Ecoplata Programme (Uruguay). Marine Policy, 51(??):527–535, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002413.

Petza:2019:OFA

[PCK^{+}19] Dimitra Petza, Christos Chalkias, Nikoleta Koukourouvli, Marta Coll, Vassiliki Vassilopoulou, Paraskevi K. Karachle, Vasiliki Markantonatou, Athanasios C. Tsikliras, and Stelios Katsanevakis. An operational framework to assess the value of fisheries restricted areas for marine conservation. Marine Policy, 102(??):28–39, April 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307103.

Purcell:2017:DER

[PCLE17] Steven W. Purcell, Beatrice I. Crona, Watisoni Lalavanua, and Hampus Eriksson. Distribution of economic returns in small-scale fisheries for international markets: a valuechain analysis. Marine Policy, 86(??):9–16, December 2017.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304244.

Pan:2014:GEI

[PCS14] Minling Pan, Keith Criddle, and Craig Severance. Guest Editors’ introduction: Catch shares and Pacific Islands Region Fisheries. Marine Policy, 44(??):1–2, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001644.

Pinkerton:2015:NPE

[PD15] Evelyn Pinkerton and Reade Davis. Neoliberalism and the politics of enclosure in North American small-scale fisheries. Marine Policy, 61(??):303–312, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000743.

Pascoe:2017:DMM

[PD17] Sean Pascoe and Catherine M. Dichmont. Does membership matter? Individual influences in natural resource management decision making. Marine Policy, 83(??): 48–54, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17301550.

Pascoe:2013:MOQ

[PDB^{+}13] Sean Pascoe, Catherine Mary Dichmont, Kate Brooks, Rachel Pears, and Edward Jebreen. Management objectives of Queensland fisheries: Putting the horse before the cart. Marine Policy, 37(??):115–122, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000334.

Parker:2010:CAF

[PDC10] Joan Parker, David Doulman, and Jean Collins. Citation analysis for the 1995 FAO Code of Conduct for Responsible Fisheries. Marine Policy, 34(1):139–144, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000876.

Porcher:2019:RUS

[PDC19] Ila France Porcher, Brian W. Darvell, and Gilles Cuny. Response to “A United States shark fin ban would undermine sustainable shark fisheries” D. S. Shiffman & R. E. Hueter, Marine Policy 85 (2017) 138–140. Marine Policy, 104(??):85–89, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301179.

Petersson:2019:PTN

[PDMÖ19] Matilda Tove Petersson, Lisa Maria Dellmuth, Andrew Merrie, and Henrik Österblom. Patterns and trends in nonstate actor participation in regional fisheries management organizations. Marine Policy, 104(??):146–156, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308722.

Parsons:2011:NOS

[PDR11] James Parsons, John Dinwoodie, and Michael Roe. Northern opportunities: a strategic review of Canada’s Arctic icebreaking services. Marine Policy, 35(4):549–556, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000182.

Pares:2015:WSS

[PDS15] Claudio Parés, Jorge Dresdner, and Hugo Salgado. Who should set the total allowable catch? Social preferences and legitimacy in fisheries management institutions. Marine Policy, 54(??):36–43, April 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003480.

Portman:2013:WHL

[PdSA^{+}13] Michelle E. Portman, Giuseppe Notarbartolo di Sciara, Tundi Agardy, Stelios Katsanevakis, Hugh P. Possingham, and Giuseppe Di-Carlo. He who hesitates is lost: Why conservation in the Mediterranean Sea is necessary and possible now. Marine Policy, 42(??):270–279, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000729.

Pascoe:2014:EPI

[PDT^{+}14] Sean Pascoe, Amar Doshi, Olivier Thébaud, Colette R. Thomas, Heidi Z. Schuttenberg, Scott F. Heron, Naneng Setiasih, James C. H. Tan, James True, Kristy Wallmo, Christy Loper, and Emma Calgaro. Estimating the potential impact of entry fees for marine parks on dive tourism in South East Asia. Marine Policy, 47(??):147–152, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000554.

Pinkerton:2010:IMF

[PE10] Evelyn Pinkerton and Danielle N. Edwards. Ignoring market failure in quota leasing? Marine Policy, 34(5): 1110–1114, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10000138.

Pedroza:2013:MIT

[Ped13] Carmen Pedroza. Middlemen, informal trading and its linkages with IUU fishing activities in the port of Progreso, Mexico. Marine Policy, 39(??):135–143, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002102.

Penca:2018:EPS

[Pen18] Jerneja Penca. European Plastics Strategy: What promise for global marine litter? Marine Policy, 97(??):197– 201, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18301519.

Pauly:2012:CFS

[PF12a] Daniel Pauly and Rainer Froese. Comments on FAO’s State of Fisheries and Aquaculture, or ‘SOFIA 2010’. Marine Policy, 36(3):746–752, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X11001722.

Poh:2012:TIU

[PF12b] Tun-Min Poh and Lucia M. Fanning. Tackling illegal, unregulated, and unreported trade towards Humphead wrasse (Cheilinus undulatus) recovery in Sabah, Malaysia. Marine Policy, 36(3):696–702, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001618.

Pilar-Fonseca:2014:IFD

[PFCP^{+}14] Tereza Pilar-Fonseca, Aida Campos, João Pereira, Ana Moreno, Silvia Lourenço, and Manuel Afonso-Dias. Integration of fishery-dependent data sources in support of octopus spatial management. Marine Policy, 45(??):69–75, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002753.

Porobic:2018:IEB

[PFF^{+}18] Javier Porobic, Elizabeth A. Fulton, Stewart Frusher, Carolina Parada, Marcus Haward, Billy Ernst, and Diana Stram. Implementing Ecosystem-based Fisheries Management: Lessons from Chile’s experience. Marine Policy, 97(??):82–90, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307510.

Pauly:2016:BHI

[PFH16] Daniel Pauly, Rainer Froese, and Sidney J. Holt. Balanced harvesting: the institutional incompatibilities. Marine Policy, 69(??):121–123, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301543.

Pita:2019:SCD

[PFMA^{+}19] Pablo Pita, Diana Fernández-Márquez, Manel Antelo, Gonzalo Macho, and Sebastián Villasante. Socioecological changes in data-poor S-fisheries: a hidden shellfisheries crisis in Galicia (NW Spain). Marine Policy, 101(??):208–224, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18301581.

Pomeroy:2017:ALP

[PFP17] Robert Pomeroy, Alice Joan Ferrer, and Joey Pedrajas. An analysis of livelihood projects and programs for fishing communities in the Philippines. Marine Policy, 81 (??):250–255, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300362.

Pedroza-Gutierrez:2019:GDL

[PG19] Carmen Pedroza-Gutiérrez. The gender division of labor in fish processing in Lake Chapala. A source of bargaining power. Marine Policy, 107(??):Article 103597, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307140.

Pinarbasi:2017:DST

[PGB^{+}17] Kemal Pınarbaşı, Ibon Galparsoro, Ángel Borja, Vanessa Stelzenmüller, Charles N. Ehler, and Antje Gimpel. Decision support tools in marine spatial planning: Present applications, gaps and future perspectives. Marine Policy, 83(??):83–91, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17301100.

Pinarbasi:2019:EUP

[PGB19] Kemal Pınarbaşı, Ibon Galparsoro, and Ángel Borja. End users’ perspective on decision support tools in marine spatial planning. Marine Policy, 108(??):Article 103658, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301101.

Pires:2016:ERS

[PGC16] Natalia M. Pires, Ricardo C. Garla, and Adriana R. Carvalho. The economic role of sharks in a major ecotourism archipelago in the western South Atlantic. Marine Policy, 72(??):31–39, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303839.

Pascoe:2019:EFE

[PGC19] Sean Pascoe, Nicholas Giles, and Louisa Coglan. Extracting fishery economic performance information from quota trading data. Marine Policy, 102(??):61–67, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307772.

Pikesley:2016:PSF

[PGL^{+}16] Stephen K. Pikesley, Brendan J. Godley, Holly Latham, Peter B. Richardson, Laura M. Robson, Jean-Luc Solandt, Colin Trundle, Chris Wood, and Matthew J. Witt. Pink sea fans (Eunicella verrucosa) as indicators of the spatial efficacy of Marine Protected Areas in southwest UK coastal waters. Marine Policy, 64(??):38–45, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002948.

Pinello:2018:UAE

[PGP18] Dario Pinello, Jennifer Gee, and Konstantinos Polymeros. An unconventional approach to estimating crew remuneration in fisheries. Marine Policy, 87(??):226–233, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302191.

Pomeroy:2019:RSW

[PGP^{+}19] Robert S. Pomeroy, Len R. Garces, Michael D. Pido, John E. Parks, and Geronimo Silvestre. The role of scale within an Ecosystem Approach to fisheries management: Policy and practice in Southeast Asian seas. Marine Policy, 106(??):Article 103531, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309047.

Pomeroy:2010:EBF

[PGPS10] Robert Pomeroy, Len Garces, Michael Pido, and Geronimo Silvestre. Ecosystem-based fisheries management in small-scale tropical marine fisheries: Emerging models of governance arrangements in the Philippines. Marine Policy, 34(2):298–308, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001079.

Peckett:2014:AQD

[PGR14] Frances J. Peckett, Gillian A. Glegg, and Lynda D. Rodwell. Assessing the quality of data required to identify effective marine protected areas. Marine Policy, 45 (??):333–341, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002108.

Pan:2016:PAF

[PH16] Min Pan and Henry P. Huntington. A precautionary approach to fisheries in the Central Arctic Ocean: Policy, science, and China. Marine Policy, 63(??):153–157, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002997.

Pomeroy:2015:PPF

[PHAC15] Caroline Pomeroy, Madeleine Hall-Arber, and Flaxen Conway. Power and perspective: Fisheries and the ocean commons beset by demands of development. Marine Policy, 61(??):339–346, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003170.

Pham:2014:FBB

[PHC14] Thi Duy Thanh Pham, Hsiang-Wen Huang, and Ching-Ta Chuang. Finding a balance between economic performance and capacity efficiency for sustainable fisheries: Case of the Da Nang gillnet fishery, Vietnam. Marine Policy, 44(??):287–294, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002182.

Plet-Hansen:2017:REM

[PHEM^{+}17] Kristian S. Plet-Hansen, Søren Q. Eliasen, Lars O. Mortensen, Heidrikur Bergsson, Hans J. Olesen, and Clara Ulrich. Remote electronic monitoring and the landing obligation — some insights into fishers’ and fishery inspectors’ opinions. Marine Policy, 76(??):98–106, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306030.

Peng:2019:DDL

[PHLM19] Daomin Peng, Xiaomei Hou, Yang Li, and Yongtong Mu. The difference in development level of marine shellfish industry in 10 major producing countries. Marine Policy, 106(??):Article 103516, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308807.

Pecceu:2016:GAM

[PHM16] Ellen Pecceu, Kris Hostens, and Frank Maes. Governance analysis of MPAs in the Belgian part of the North Sea. Marine Policy, 71(??):265–274, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003875.

Poulsen:2018:DER

[PHS18] René Taudal Poulsen, Roberto Rivas Hermann, and Carla K. Smink. Do eco-rating schemes improve the environmental performance of ships? Marine Policy, 87 (??):94–103, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17304475.

Paterson:2010:TCO

[PIH^{+}10] B. Paterson, M. Isaacs, M. Hara, A. Jarre, and C. L. Moloney. Transdisciplinary co-operation for an ecosystem approach to fisheries: a case study from the South African sardine fishery. Marine Policy, 34(4):782–794, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000205.

Pinkerton:2013:AIE

[Pin13] Evelyn Pinkerton. Alternatives to ITQs in equity– efficiency–effectiveness trade-offs: How the lay-up system spread effort in the BC halibut fishery. Marine Policy, 42(??):5–13, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000195.

Pinkerton:2015:RME

[Pin15] Evelyn Pinkerton. The role of moral economy in two British Columbia fisheries: Confronting neoliberal policies. Marine Policy, 61(??):410–419, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000913.

Pinkerton:2017:HRD

[Pin17] Evelyn Pinkerton. Hegemony and resistance: Disturbing patterns and hopeful signs in the impact of neoliberal policies on small-scale fisheries around the world. Marine Policy, 80(??):1–9, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307345.

Porter:2013:ECT

[PJS13] Read D. Porter, Zachary Jylkka, and Greta Swanson. Enforcement and compliance trends under IFQ management in the Gulf of Mexico commercial reef fish fishery. Marine Policy, 38(??):45–53, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001066.

Petersen:2016:NSG

[PKA^{+}16] S. Petersen, A. Krätschell, N. Augustin, J. Jamieson, J. R. Hein, and M. D. Hannington. News from the seabed — Geological characteristics and resource potential of deepsea mineral resources. Marine Policy, 70(??):175–187, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300732.

Prosperi:2019:ASS

[PKM^{+}19] Paolo Prosperi, James Kirwan, Damian Maye, Fabio Bartolini, Daniele Vergamini, and Gianluca Brunori. Adaptation strategies of small-scale fisheries within changing market and regulatory conditions in the EU. Marine Policy, 100(??):316–323, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300496.

Pentz:2018:CRF

[PKOF18] Brian Pentz, Nicole Klenk, Sharron Ogle, and Jonathan A. D. Fisher. Can regional fisheries management organizations (RFMOs) manage resources effectively during climate change? Marine Policy, 92(??):13–20, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305560.

Perepelkin:2010:IMM

[PKPdP10] Mihail Perepelkin, Sabine Knapp, German Perepelkin, and Michiel de Pooter. An improved methodology to measure flag performance for the shipping industry. Marine Policy, 34(3):395–405, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001377.

Prakash:2017:MAT

[PKR^{+}17] Sanjeevi Prakash, Thipramalai Thangappan Ajith Kumar, Rajeev Raghavan, Andrew Rhyne, Michael F. Tlusty, and Thanumalaya Subramoniam. Marine aquarium trade in India: Challenges and opportunities for conservation and policy. Marine Policy, 77(??):120–129, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305073.

Perez-Labajos:2012:AIB

[PL12] Carlos A. Perez-Labajos. Analysis of inequalities between the fishing capacities of the fleets of the European union. Marine Policy, 36(3):630–635, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001709.

Patrick:2015:HPS

[PL15] Wesley S. Patrick and Jason S. Link. Hidden in plain sight: Using optimum yield as a policy framework to operationalize ecosystem-based fisheries management. Marine Policy, 62(??):74–81, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500233X.

Platjouw:2019:DTL

[Pla19] Froukje Maria Platjouw. Dimensions of transboundary legal coherence needed to foster ecosystem-based governance in the Arctic. Marine Policy, 110(??):Article 103666, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303185.

Papastavrou:2017:WMD

[PLL17] Vassili Papastavrou, Russell Leaper, and David Lavigne. Why management decisions involving marine mammals should include animal welfare. Marine Policy, 79(??):19– 24, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306960.

Purcell:2014:CSM

[PLP14] S. W. Purcell, A. Lovatelli, and K. Pakoa. Constraints and solutions for managing Pacific Island sea cucumber fisheries with an ecosystem approach. Marine Policy, 45 (??):240–250, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002479.

Powers:2010:CFU

[PM10] Joseph E. Powers and Melissa H. Monk. Current and future use of indicators for ecosystem based fisheries management. Marine Policy, 34(3):723–727, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001821.

Pierce:2014:PPC

[PM14] Brett Pierce and Pallab Mozumder. Perceptions and preferences of commercial fishers for dedicated access privilege framework in a multispecies fishery. Marine Policy, 45(??):52–59, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002480.

Porras:2017:PPP

[PMA^{+}17] Ina Porras, Essam Y. Mohammed, Liaquat Ali, Md. Shahajat Ali, and Md. Belayet Hossain. Power, profits and payments for ecosystem services in Hilsa fisheries in Bangladesh: a value chain analysis. Marine Policy, 84 (??):60–68, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300994.

Pranovi:2015:PTF

[PMC^{+}15] Fabio Pranovi, Marco Anelli Monti, Alberto Caccin, Daniele Brigolin, and Matteo Zucchetta. Permanent trawl fishery closures in the Mediterranean Sea: an effective management strategy? Marine Policy, 60(??):272–279, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002043.

Peron:2010:WDF

[PML10] Guillaume Péron, Jean François Mittaine, and Bertrand Le Gallic. Where do fishmeal and fish oil products come from? An analysis of the conversion ratios in the global fishmeal industry. Marine Policy, 34(4):815–820, July 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1000028X.

Peckham:2017:CFC

[PMM^{+}17] S. Hoyt Peckham, David Maldonado, Yoshimasa Matsuzawa, Kama Dean, and Irene Kinan Kelly. Connecting fishers to conserve a transpacific ambassador: a trinational fisheries learning exchange. Marine Policy, 77(??):231–237, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304298.

Pelc:2015:FAB

[PMN^{+}15] Robin A. Pelc, Lisa M. Max, Wendy Norden, Santi Roberts, Rachel Silverstein, and Sam R. Wilding. Further action on bycatch could boost United States fisheries performance. Marine Policy, 56(??):56–60, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000275.

Pramod:2014:EIU

[PNPD14] Ganapathiraju Pramod, Katrina Nakamura, Tony J. Pitcher, and Leslie Delagran. Estimates of illegal and unreported fish in seafood imports to the USA. Marine Policy, 48(??):102–113, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000918.

Pepin-Neff:2018:RFI

[PNW18] Christopher L. Pepin-Neff and Thomas Wynter. Reducing fear to influence policy preferences: an experiment with sharks and beach safety policy options. Marine Policy, 88(??):222–229, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302397.

Perry:2010:ICG

[PO10] R. Ian Perry and Rosemary E. Ommer. Introduction: Coping with global change in marine social-ecological systems. Marine Policy, 34(4):739–741, July 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000266.

Pater:2014:DMH

[PO14] Christopher Pater and Ian Oxley. Developing marine historic environment management policy: the English Heritage experience. Marine Policy, 45(??):342–348, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002078.

Polacheck:2012:AIF

[Pol12a] Tom Polacheck. Assessment of IUU fishing for Southern Bluefin Tuna. Marine Policy, 36(5):1150–1165, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1200036X.

Polacheck:2012:PIS

[Pol12b] Tom Polacheck. Politics and independent scientific advice in RFMO processes: a case study of crossing boundaries. Marine Policy, 36(1):132–141, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000789.

Pomeroy:2012:MOS

[Pom12] Robert S. Pomeroy. Managing overcapacity in small-scale fisheries in Southeast Asia. Marine Policy, 36(2):520–527, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001527.

Pomeroy:2016:RFT

[Pom16] Robert Pomeroy. A research framework for traditional fisheries: Revisited. Marine Policy, 70(??):153–163, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302949.

Porter:2010:FOE

[Por10] Read D. Porter. Fisheries observers as enforcement assets: Lessons from the North Pacific. Marine Policy, 34(3):583–589, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0900178X.

Portman:2015:MSP

[Por15] M. E. Portman. Marine spatial planning in the Middle East: Crossing the policy-planning divide. Marine Policy, 61(??):8–15, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001906.

Porcelli:2017:CBC

[Por17] Apollonya Maria Porcelli. Comparing bonding capital in New England groundfish and scallop fisheries: Differing effects of privatization. Marine Policy, 84(??):244–250, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306200.

Paterson:2010:EIS

[PP10] Barbara Paterson and Samantha L. Petersen. EAF implementation in Southern Africa: Lessons learnt. Marine Policy, 34(2):276–292, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001043.

Pramod:2019:DSI

[PP19] Ganapathiraju Pramod and Tony J. Pitcher. In defence of seafood import analysis: Credulity bamboozled by supply chain laundering. Marine Policy, 108(??): Article 103651, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19305226.

Pomeroy:2016:IMF

[PPCM16] Robert Pomeroy, John Parks, Kitty Courtney, and Nives Mattich. Improving marine fisheries management in Southeast Asia: Results of a regional fisheries stakeholder analysis. Marine Policy, 65(??):20–29, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003656.

Pramod:2017:REI

[PPM17] Ganapathiraju Pramod, Tony J. Pitcher, and Gopikrishna Mantha. RETRACTED: Estimates of illegal and unreported seafood imports to Japan. Marine Policy, 84 (??):42–51, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300817.

Pramod:2019:RNE

[PPM19] Ganapathiraju Pramod, Tony J. Pitcher, and Gopikrishna Mantha. Retraction notice to “Estimates of illegal and unreported seafood imports to Japan”. Marine Policy, 107(??):Article 103574, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304245.

Pomeroy:2016:DIF

[PPML16] Robert Pomeroy, John Parks, Karina Lorenz Mrakovcich, and Christopher LaMonica. Drivers and impacts of fisheries scarcity, competition, and conflict on maritime security. Marine Policy, 67(??):94–104, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000105.

Potts:2016:WCE

[PPOM16] Tavis Potts, Cristina Pita, Tim O’Higgins, and Laurence Mee. Who cares? European attitudes towards marine and coastal environments. Marine Policy, 72(??):59–66, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303670.

Pomeroy:2015:MTE

[PPR^{+}15] Robert Pomeroy, Kevin Hiew Wai Phang, K. Ramdass, Jasmin Mohd Saad, Paul Lokani, Grizelda Mayo-Anda, Edward Lorenzo, Gidor Manero, Zhazha Maguad, Michael D. Pido, and Gilliam Goby. Moving towards an ecosystem approach to fisheries management in the Coral Triangle region. Marine Policy, 51(??):211–219, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002292.

Pita:2010:SPF

[PPT10] Cristina Pita, Graham J. Pierce, and Ioannis Theodossiou. Stakeholders’ participation in the fisheries management decision-making process: Fishers’ perceptions of participation. Marine Policy, 34(5):1093–1102, September 2010.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000618.

Pan:2015:EED

[PQLH15] Guangchen Pan, Shengyao Qiu, Xin Liu, and Xiaoke Hu. Estimating the economic damages from the Penglai 19-3 oil spill to the Yantai fisheries in the Bohai Sea of northeast China. Marine Policy, 62(??):18–24, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002262.

Pierce:2013:OTR

[PR13] Janine Pierce and Guy Robinson. Oysters thrive in the right environment: the social sustainability of oyster farming in the Eyre Peninsula, South Australia. Marine Policy, 37(??):77–85, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000693.

Prasad:2019:FDE

[PR19] Ravita D. Prasad and Atul Raturi. Fuel demand and emissions for maritime sector in Fiji: Current status and low-carbon strategies. Marine Policy, 102(??):40–50, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303397.

Parappurathu:2019:FVI

[PRBX19] Shinoj Parappurathu, C. Ramachandran, K. K. Baiju, and Antony Kurisunkal Xavier. Formal versus informal: Insights into the credit transactions of small-scale fishers along the south west coast of India. Marine Policy, 103(??):101–112, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1830647X.

Parappurathu:2017:WAF

[PRG^{+}17] Shinoj Parappurathu, C. Ramachandran, A. Gopalakrishnan, Dilip Kumar, M. K. Poddar, Manas Choudhury, R. Geetha, K. Mohammed Koya, R. Narayana Kumar, K. P. Salini, and P. V. Sunil. What ails fisheries insurance in India? An assessment of issues, challenges and future potential. Marine Policy, 86(??):144–155, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305304.

Princen:2010:VSP

[Pri10] Sebastiaan Princen. Venue shifts and policy change in EU fisheries policy. Marine Policy, 34(1):36–41, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000566.

Perez-Ramirez:2012:MCA

[PRLCL12] Mónica Pérez-Ramírez, Salvador Lluch-Cota, and Mario Lasta. MSC certification in Argentina: Stakeholders’ perceptions and lessons learned. Marine Policy, 36(5): 1182–1187, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000553.

Probst:2016:PRD

[PRO16] Wolfgang Nikolaus Probst, Andrea Rau, and Daniel Oesterwind. A proposal for restructuring Descriptor 3 of the Marine Strategy Framework Directive (MSFD). Marine Policy, 74(??):128–135, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305905.

Perez-Ramirez:2012:PIF

[PRPLBLC12] Mónica Pérez-Ramírez, Bruce Phillips, Daniel Lluch-Belda, and Salvador Lluch-Cota. Perspectives for implementing fisheries certification in developing countries. Marine Policy, 36(1):297–302, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001163.

Prescott:2017:MSL

[PRPS17] James Prescott, James Riwu, Andhika P. Prasetyo, and Natasha Stacey. The money side of livelihoods: Economics of an unregulated small-scale Indonesian sea cucumber fishery in the Timor Sea. Marine Policy, 82(??):197–205, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301793.

Prayaga:2010:VRF

[PRS10] Prabha Prayaga, John Rolfe, and Natalie Stoeckl. The value of recreational fishing in the Great Barrier Reef, Australia: a pooled revealed preference and contingent behaviour model. Marine Policy, 34(2):244–251, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001006.

Pittman:2019:MPC

[PRS^{+}19] S. J. Pittman, L. D. Rodwell, R. J. Shellock, M. Williams, M. J. Attrill, J. Bedford, K. Curry, S. Fletcher, S. C. Gall, J. Lowther, A. McQuatters-Gollop, K. L. Moseley, and S. E. Rees. Marine parks for coastal cities: a concept for enhanced community well-being, prosperity and sustainable city living. Marine Policy, 103(??):160–171, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307486.

Parsons:2010:TLI

[PRT10] E. C. M. Parsons, N. A. Rose, and T. M. Telecky. The trade in live Indo-Pacific bottlenose dolphins from Solomon Islands — a CITES decision implementation case study. Marine Policy, 34(3):384–388, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001249.

Phillipson:2010:RIF

[PS10] Jeremy Phillipson and David Symes. Recontextualising inshore fisheries: the changing face of British inshore fisheries management. Marine Policy, 34(6):1207–1214, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000825.

Pedroza:2011:RFS

[PS11a] Carmen Pedroza and Silvia Salas. Responses of the fishing sector to transitional constraints: From reactive to proactive change, Yucatan fisheries in Mexico. Marine Policy, 35(1):39–49, January 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001405.

Pinkerton:2011:CCC

[PS11b] Evelyn Pinkerton and Jennifer Silver. Cadastralizing or coordinating the clam commons: Can competing community and government visions of wild and farmed fisheries be reconciled? Marine Policy, 35(1):63–72, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001429.

Pontecorvo:2012:ELD

[PS12] Giulio Pontecorvo and William E. Schrank. The expansion, limit and decline of the global marine fish catch. Marine Policy, 36(5):1178–1181, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000498.

Pontecorvo:2014:CDW

[PS14] Giulio Pontecorvo and William E. Schrank. The continued decline in the world catch of marine fish. Marine Policy, 44(??):117–119, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300170X.

Phillipson:2018:STB

[PS18] Jeremy Phillipson and David Symes. ’A sea of troubles’: Brexit and the fisheries question. Marine Policy, 90(??):168–173, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307376.

Platjouw:2019:RRL

[PS19] Froukje Maria Platjouw and Niko Soininen. Reconciling the rule of law with adaptive regulation of marine ecosystems — Challenges and opportunities for the Arctic and beyond. Marine Policy, 110(??):Article 103726, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307948.

Plantier-Santos:2012:GME

[PSCY12] Carlota Plantier-Santos, Cristina Carollo, and David W. Yoskowitz. Gulf of Mexico Ecosystem Service Valuation Database (GecoServ): Gathering ecosystem services valuation studies to promote their inclusion in the decisionmaking process. Marine Policy, 36(1):214–217, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000959.

Pihlajamaki:2018:FSS

[PSH18] Mia Pihlajamäki, Simo Sarkki, and Päivi Haapasaari. Food security and safety in fisheries governance — a case study on Baltic herring. Marine Policy, 97(??):211–219, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306528.

Piniella:2010:AFA

[PSP10] Francisco Piniella, Milagrosa Casimiro Soriguer, and Francisco Pastoriza. Artisanal fishing in Andalusia (and III): “The Day After… ”. Marine Policy, 34(1):120–132, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000852.

Poisson:2014:CRD

[PSV^{+}14] François Poisson, Bernard Séret, Anne-Lise Vernet, Michel Goujon, and Laurent Dagorn. Collaborative research: Development of a manual on elasmobranch handling and release best practices in tropical tuna purse-seine fisheries. Marine Policy, 44(??):312–320, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002224.

Progoulaki:2010:HRM

[PT10] Maria Progoulaki and Ioannis Theotokas. Human resource management and competitive advantage: an application of resource-based view in the shipping industry. Marine Policy, 34(3):575–582, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001778.

Patrolia:2017:IWR

[PTDH17] Emily Patrolia, Robert Thompson, Tracey Dalton, and Porter Hoagland. The influence of weather on the recreational uses of coastal lagoons in Rhode Island, USA. Marine Policy, 83(??):252–258, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305802.

Pendleton:2016:VGM

[PTML16] Linwood H. Pendleton, Olivier Thébaud, Rémi C. Mongruel, and Harold Levrel. Has the value of global marine and coastal ecosystem services changed? Marine Policy, 64(??):156–158, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003620.

Paxton:2017:SSN

[PTN^{+}17] Avery B. Paxton, J. Christopher Taylor, Douglas P. Nowacek, Julian Dale, Elijah Cole, Christine M. Voss, and Charles H. Peterson. Seismic survey noise disrupted fish use of a temperate reef. Marine Policy, 78(??):68–73, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307382.

Pita:2013:PSI

[PTP13] Cristina Pita, Ioannis Theodossiou, and Graham J. Pierce. The perceptions of Scottish inshore fishers about marine protected areas. Marine Policy, 37(??):254–263, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000954.

Prestrelo:2016:IMU

[PV16]  Luana Prestrelo and Marcelo Vianna.  Identifying

multiple-use conflicts prior to marine spatial planning: a case study of a multi-legislative estuary in Brazil. Marine Policy, 67(??):83–93, May 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000452.

Pita:2018:MSD

[PVA^{+}18] Pablo Pita, Sebastián Villasante, Robert Arlinghaus, Pedro Gomes, Harry V. Strehlow, Pedro Veiga, José Vingada, and Kieran Hyder. A matter of scales: Does the management of marine recreational fisheries follow the ecosystem approach to fisheries in Europe? Marine Policy, 97(??):61–71, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304512.

Pokki:2018:CEA

[PVK18] H. Pokki, J. Virtanen, and S. Karvinen. Comparison of economic analysis with financial analysis of fisheries: Application of the perpetual inventory method to the Finnish fishing fleet. Marine Policy, 95(??):239–247, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307078.

Pita:2019:MSS

[PVPF19] Cristina Pita, Sebastián Villasante, and José J. Pascual-Fernández. Managing small-scale fisheries under data poor scenarios: lessons from around the world. Marine Policy, 101(??):154–157, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19301113.

Popova:2019:ECB

[PVS^{+}19] Ekaterina Popova, David Vousden, Warwick H. H. Sauer, Essam Y. Mohammed, Valerie Allain, Nicola Downey-Breedt, Ruth Fletcher, Kristina M. Gjerde, Patrick N. Halpin, Stephen Kelly, David Obura, Gretta Pecl, Michael Roberts, Dionysios E. Raitsos, Alex Rogers, Melita Samoilys, Ussif Rashid Sumaila, Sean Tracey, and Andrew Yool. Ecological connectivity between the areas beyond national jurisdiction and coastal waters: Safeguarding interests of coastal communities in developing countries. Marine Policy, 104(??):90–102, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300764.

Pennino:2017:DMS

[PVVB17] Maria Grazia Pennino, Raul Vilela, Julio Valeiras, and Jose M. Bellido. Discard management: a spatial multicriteria approach. Marine Policy, 77(??):144–151, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307333.

Pan:2015:MPS

[PW15] Minling Pan and John Walden. Measuring Productivity in a Shared Stock Fishery: a Case Study of the Hawaii Longline Fishery. Marine Policy, 62(??):302–308, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002195.

Piwowarczyk:2016:DLG

[PW16] Joanna Piwowarczyk and Borys Wróbel. Determinants of legitimate governance of marine Natura 2000 sites in a post-transition European Union country: a case study of Puck Bay, Poland. Marine Policy, 71(??):310–317, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000269.

Pittman:2019:GAS

[PWB19] Jeremy Pittman, Colette C. C. Wabnitz, and Robert Blasiak. A global assessment of structural change in development funding for fisheries. Marine Policy, 109(??):Article 103644, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305165.

Papathanasopoulou:2016:VHB

[PWH^{+}16] Eleni Papathanasopoulou, Mathew P. White, Caroline Hattam, Aisling Lannin, Andrea Harvey, and Anne Spencer. Valuing the health benefits of physical activities in the marine environment and their importance for marine spatial planning. Marine Policy, 63(??):144–152, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002936.

Purcell:2018:CMP

[PWN18] Steven W. Purcell, David H. Williamson, and Poasi Ngaluafe. Chinese market prices of beche-de-mer: Implications for fisheries and aquaculture. Marine Policy, 91(??):58–65, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308035.

Penney:2017:MSG

[PWR17] Ryan Penney, Geoff Wilson, and Lynda Rodwell. Managing sino-ghanaian fishery relations: a political ecology approach. Marine Policy, 79(??):46–53, May 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308132.

Pauly:2017:BCD

[PZ17a] Daniel Pauly and Dirk Zeller. The best catch data that can possibly be? Rejoinder to Ye et al. “FAO’s statistic data and sustainability of fisheries and aquaculture”. Marine Policy, 81(??):406–410, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301380.

Pauly:2017:CFS

[PZ17b] Daniel Pauly and Dirk Zeller. Comments on FAO’s State of World Fisheries and Aquaculture (SOFIA 2016). Marine Policy, 77(??):176–181, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305516.

Pauly:2019:AFC

[PZ19] Daniel Pauly and Dirk Zeller. Agreeing with FAO: Comments on SOFIA 2018. Marine Policy, 100(??): 332–333, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18305669.

Qiu:2017:QAM

[QG17] Wenxian Qiu and Warwick Gullett. Quantitative analysis for maritime delimitation: Reassessing the Bay of Bengal delimitation between Bangladesh and Myanmar. Marine Policy, 78(??):45–54, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301245.

Qiu:2013:SCR

[Qiu13] Wanfei Qiu. The Sanya Coral Reef National Marine Nature Reserve, China: a governance analysis. Marine Policy, 41(??):50–56, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002710.

Qiu:2013:EPL

[QJ13] Wanfei Qiu and Peter J. S. Jones. The emerging policy landscape for marine spatial planning in Europe. Marine Policy, 39(??):182–190, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002084.

Quynh:2017:TUR

[QSHI17] Chi Nguyen Thi Quynh, Steven Schilizzi, Atakelty Hailu, and Sayed Iftekhar. Territorial Use Rights for Fisheries (TURFs): State of the art and the road ahead. Marine Policy, 75(??):41–52, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303086.

Quesada-Silva:2019:SPA

[QSICTdV19] Michele Quesada-Silva, Alejandro Iglesias-Campos, Alexander Turra, and Juan L. Suárez de Vivero. Stakeholder Participation Assessment Framework (SPAF): a theory-based strategy to plan and evaluate marine spatial planning participatory processes. Marine Policy, 108(??):Article 103619, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302258.

Quijano:2018:FCT

[QSMGVA18] Daniel Quijano, Silvia Salas, Carmen Monroy-García, and Iván Velázquez-Abunader. Factors contributing to technical efficiency in a mixed fishery: Implications in buyback programs. Marine Policy, 94(??):61–70, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303305.

Qi:2018:DMR

[QZH^{+}18] Zhixin Qi, Chunchang Zhang, Junsong Han, Yali Gao, and Deqi Xiong. A decision model responding to the refuge request from a ship in need of assistance. Marine Policy, 95(??):294–300, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307820.

Reithe:2014:MPA

[RAF14] Siv Reithe, Claire W. Armstrong, and Ola Flaaten. Marine protected areas in a welfare-based perspective. Marine Policy, 49(??):29–36, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001134.

Rodrigues:2018:HCR

[RAG18] Amanda Ricci Rodrigues, Patrízia Raggi Abdallah, and Maria A. Gasalla. Harvesting costs and revenues: Implication of the performance of open-access industrial fishing fleets off Rio Grande, Brazil. Marine Policy, 93 (??):104–112, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307625.

Rahman:2014:MSM

[Rah14] Nazira Abdul Rahman. Multilateralism in the Straits of Malacca and Singapore. Marine Policy, 44(??):232– 238, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001978.

Rahman:2019:BRC

[Rah19] Chris Rahman. Book review: China’s Maritime Gray Zone Operations, Andrew S. Erickson, Ryan D. Martinson (Eds.). Naval Institute Press, Annapolis, MD (2019), 324, US$50, -13: 978-1-59114-693-3. Marine Policy, 110(??):Article 103586, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304270.

Rayfuse:2019:RLR

[Ray19] Rosemary Rayfuse. The role of law in the regulation of fishing activities in the Central Arctic Ocean. Marine Policy, 110(??):Article 103562, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303665.

Rasheed:2016:VDT

[RAZ16] A. Rifaee Rasheed, Ameer Abdulla, and Nabeeh I. Zakariyya. Vulnerability of different types of fishers to potential implementation of a management plan in a Marine Protected Area (MPA) in the Maldives. Marine Policy, 74(??):195–204, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301622.

Rochette:2012:IPR

[RB12] Julien Rochette and Raphaël Billé. ICZM Protocols to Regional Seas Conventions: What? Why? How? Marine Policy, 36(5):977–984, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000218.

Ram-Bidesi:2015:RRW

[RB15] Vina Ram-Bidesi. Recognizing the role of women in supporting marine stewardship in the Pacific Islands. Marine Policy, 59(??):1–8, September 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001025.

Rogers:2016:SLP

[RBB^{+}16] Kerrylee Rogers, Paul I. Boon, Simon Branigan, Norman C. Duke, Colin D. Field, James A. Fitzsimons, Hugh Kirkman, Jock R. Mackenzie, and Neil Saintilan. The state of legislation and policy protecting Australia’s mangrove and salt marsh and their ecosystem services. Marine Policy, 72(??):139–155, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304067.

Richardson:2013:LTC

[RBC^{+}13] Peter B. Richardson, Annette C. Broderick, Michael S. Coyne, Shannon Gore, James C. Gumbs, Arlington Pickering, Susan Ranger, Matthew J. Witt, and Brendan J. Godley. Leatherback turtle conservation in the Caribbean UK overseas territories: Act local, think global? Marine Policy, 38(??):483–490, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001704.

Ross:2011:TGP

[RBJ^{+}11] Peter S. Ross, Jay Barlow, Thomas A. Jefferson, Brendan E. Hickie, Tatiana Lee, Christina MacFarquhar, E. Christien Parsons, Kimberly N. Riehl, Naomi A. Rose, Elisabeth Slooten, Chia-Yang Tsai, John Y. Wang, Andrew J. Wright, and Shih Chu Yang. Ten guiding principles for the delineation of priority habitat for endangered small cetaceans. Marine Policy, 35(4):483–488, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002022.

Rochette:2015:ROG

[RBM^{+}15] Julien Rochette, Raphaël Billé, Erik J. Molenaar, Petra Drankier, and Lucien Chabason. Regional oceans governance mechanisms: a review. Marine Policy, 60 (??):9–19, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001451.

Rossetto:2015:MCD

[RBS^{+}15] Marisa Rossetto, Isabella Bitetto, Maria Teresa Spedicato, Giuseppe Lembo, Monica Gambino, Paolo Accadia, and Paco Melià. Multi-criteria decision-making for fisheries management: a case study of Mediterranean demersal fisheries. Marine Policy, 53(??):83–93, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400298X.

Rowlands:2019:SSF

[RBS^{+}19] Gwilym Rowlands, Judith Brown, Bradley Soule, Pablo Trueba Boluda, and Alex D. Rogers. Satellite surveillance of fishing vessel activity in the Ascension Island Exclusive Economic Zone and Marine Protected Area. Marine Policy, 101(??):39–50, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303002.

Robotham:2019:CSS

[RBTR^{+}19] Hugo Robotham, Eduardo Bustos, Francisco Ther-Rios, Marcela Avila, Matías Robotham, Carlos Hidalgo, and Jorge Muñoz. Contribution to the study of sustainability of small-scale artisanal fisheries in Chile. Marine Policy, 106(??):Article 103514, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303774.

Rehn:2018:SRU

[RBW18] A. Carson Rehn, Allain J. Barnett, and Melanie G. Wiber. Stabilizing risk using public participatory GIS: a case study on mitigating marine debris in the Bay of Fundy, Southwest New Brunswick, Canada. Marine Policy, 96(??):264–269, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305110.

Robb:2011:CFC

[RBWL11] Carolyn K. Robb, Karin M. Bodtker, Kim Wright, and Jennifer Lash. Commercial fisheries closures in marine protected areas on Canada’s Pacific coast: the exception, not the rule. Marine Policy, 35(3):309–316, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001788.

Ringer:2018:GCP

[RCD^{+}18] Danielle Ringer, Courtney Carothers, Rachel Donkersloot, Jesse Coleman, and Paula Cullenberg. For generations to come? The privatization paradigm and shifting social baselines in Kodiak, Alaska’s commercial fisheries. Marine Policy, 98(??):97–103, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300411.

Reid:2013:AEM

[RCdLS13] Chris Reid, Nick Caputi, Simon de Lestang, and Peter Stephenson. Assessing the effects of moving to maximum economic yield effort level in the western rock lobster fishery of Western Australia. Marine Policy, 39(??):303–313, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002394.

Raheem:2012:ANM

[RCF^{+}12] N. Raheem, S. Colt, E. Fleishman, J. Talberth, P. Swedeen, K. J. Boyle, M. Rudd, R. D. Lopez, D. Crocker, D. Bohan, T. O’Higgins, C. Willer, and R. M. Boumans. Application of non-market valuation to California’s coastal policy decisions. Marine Policy, 36(5):1166–1171, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000061.

Rosendo:2018:DMS

[RCM18] S. Rosendo, L. Celliers, and M. Mechisso. Doing more with the same: a reality-check on the ability of local government to implement Integrated Coastal Management for climate change adaptation. Marine Policy, 87(??):29–39, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17306504.

Rubio-Cisneros:2019:PFD

[RCMBG^{+}19] Nadia T. Rubio-Cisneros, Marcia Moreno-Báez, Jeffrey Glover, Dominique Rissolo, Andrea Sáenz-Arroyo, Christöpher Götz, Silvia Salas, Anthony Andrews, Gustavo Marín, Sara Morales-Ojeda, Francisca Antele, and Jorge Herrera-Silveira. Poor fisheries data, many fishers, and increasing tourism development: Interdisciplinary views on past and current small-scale fisheries exploitation on Holbox Island. Marine Policy, 100(??):8–20, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303779.

Reed:2013:BFC

[RCUR13] Matt Reed, Paul Courtney, Julie Urquhart, and Natalie Ross. Beyond fish as commodities: Understanding the socio-cultural role of inshore fisheries in England. Marine Policy, 37(??):62–68, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200070X.

Rangeley:2012:RSB

[RD12] Robert W. Rangeley and Robin W. D. Davies. Raising the “sunken billions”: Financing the transition to sustainable fisheries. Marine Policy, 36(5):1044–1046, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000371.

Ruddle:2013:HRN

[RD13] Kenneth Ruddle and Anthony Davis. Human rights and neo-liberalism in small-scale fisheries: Conjoined priorities and processes. Marine Policy, 39(??):87–93, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002138.

Rigby:2017:IAP

[RDBB17] Benjamin Rigby, Reade Davis, Dean Bavington, and Christopher Baird. Industrial aquaculture and the politics of resignation. Marine Policy, 80(??):19–27, June 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306777.

Ruckelshaus:2013:SOB

[RDG^{+}13] M. Ruckelshaus, S. C. Doney, H. M. Galindo, J. P. Barry, F. Chan, J. E. Duffy, C. A. English, S. D. Gaines, J. M. Grebmeier, A. B. Hollowed, N. Knowlton, J. Polovina, N. N. Rabalais, W. J. Sydeman, and L. D. Talley. Securing ocean benefits for society in the face of climate change. Marine Policy, 40(??):154–159, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000183.

Raoux:2018:ACS

[RDP^{+}18] Aurore Raoux, Jeffrey M. Dambacher, Jean-Philippe Pezy, Camille Mazé, Jean-Claude Dauvin, and Nathalie Niquil. Assessing cumulative socio-ecological impacts of offshore wind farm development in the Bay of Seine (English Channel). Marine Policy, 89(??):11–20, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730444X.

Ratz:2010:CRE

[RDSB10] Hans-Joachim Rätz, Hendrik Dörner, Robert Scott, and Thomas Barbas. Complementary roles of European and national institutions under the Common Fisheries Policy and the Marine Strategy Framework Directive. Marine Policy, 34(5):1028–1035, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000539.

Rosegrant:2016:EIC

[RDVSC16] Mark W. Rosegrant, Madan Mohan Dey, Rowena Valmonte-Santos, and Oai Li Chen. Economic impacts of climate change and climate change adaptation strategies in Vanuatu and Timor-Leste. Marine Policy, 67(??):179–188, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003802.

Reeves:2014:DEC

[REA^{+}14] Randall R. Reeves, Peter J. Ewins, Selina Agbayani, Mads Peter Heide-Jørgensen, Kit M. Kovacs, Christian Lydersen, Robert Suydam, Wendy Elliott, Gert Polet, Yvette van Dijk, and Rosanne Blijleven. Distribution of endemic cetaceans in relation to hydrocarbon development and commercial shipping in a warming Arctic. Marine Policy, 44(??):375–389, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002303.

Reid:2010:CRT

[Rei10] Chris Reid. Conference report: Twelfth Conference of the North Atlantic Fisheries History Association, Old Dominion University, Norfolk, Virginia, 19–22 August 2009. Marine Policy, 34(3):737–738, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001432.

Reis:2014:ISA

[Rei14] Jeanette Reis. Introduction to systems approaches in coastal management — The legacy of the SPICOSA project. Marine Policy, 43(??):1–2, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000882.

Reineman:2016:USW

[Rei16] Daniel R. Reineman. The utility of surfers’ wave knowledge for coastal management. Marine Policy, 67(??):139–147, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000361.

Reidy:2019:UBR

[Rei19] Rhonda D. Reidy. Understanding the barriers to reconciling marine mammal-fishery conflicts: a case study in British Columbia. Marine Policy, 108(??):Article 103635, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18307255.

Renwick:2018:RCE

[Ren18] Alan Renwick. Regulatory challenges to economic growth in aquaculture: the case of licensing in the Irish oyster industry. Marine Policy, 88(??):151–157, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730310X.

Rees:2013:PQS

[RFG^{+}13] Siân Rees, Stephen Fletcher, Gillian Glegg, Charlotte Marshall, Lynda Rodwell, Rebecca Jefferson, Maria Campbell, Olivia Langmead, Matthew Ashley, Helen Bloomfield, Daniel Brutto, Andrew Colenutt, Alessandra Conversi, Bob Earll, Caroline Hattam, Simon Ingram, Emma McKinley, Laurence Mee, Jenny Oates, Francis Peckett, Jim Portus, Martin Reed, Stuart Rogers, Justine Saunders, Kylie Scales, and Russel Wynn. Priority questions to shape the marine and coastal policy research agenda in the United Kingdom. Marine Policy, 38(??):531–537, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001777.

Rees:2014:SBL

[RFG^{+}14a] Siân E. Rees, Stephen Fletcher, Sarah C. Gall, Laura A. Friedrich, Emma L. Jackson, and Lynda D. Rodwell. Securing the benefits: Linking ecology with marine planning policy to examine the potential of a network of Marine Protected Areas to support human wellbeing. Marine Policy, 44(??):335–341, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002248.

Rodwell:2014:MCP

[RFG^{+}14b] L. D. Rodwell, S. Fletcher, G. A. Glegg, M. Campbell, S. E Rees, M. Ashley, E. A. Linley, M. Frost, B. Earll, R. B. Wynn, L. Mee, P. Almada-Villela, D. Lear, P. Stanger, A. Colenutt, F. Davenport, N. J. Barker Bradshaw, and R. Covey. Marine and coastal policy in the UK: Challenges and opportunities in a new era. Marine Policy, 45(??):251–258, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300211X.

Ruiz-Frau:2013:SEE

[RFHEJK13] A. Ruiz-Frau, H. Hinz, G. Edwards-Jones, and M. J. Kaiser. Spatially explicit economic assessment of cultural ecosystem services: Non-extractive recreational uses of the coastal environment related to marine biodiversity. Marine Policy, 38(??):90–98, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001212.

Ruiz-Frau:2015:BEN

[RFKEJ^{+}15] Ana Ruiz-Frau, Michel J. Kaiser, Gareth Edwards-Jones, Carissa J. Klein, Daniel Segan, and Hugh P. Possingham. Balancing extractive and non-extractive uses in marine conservation plans. Marine Policy, 52(??):11–18, February 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002796.

Rees:2018:DQE

[RFL^{+}18] Siân E. Rees, Nicola L. Foster, Olivia Langmead, Simon Pittman, and David E. Johnson. Defining the qualitative elements of Aichi Biodiversity Target 11 with regard to the marine and coastal environment in order to strengthen global efforts for marine biodiversity conservation outlined in the United Nations Sustainable Development Goal 14. Marine Policy, 93(??):241–250, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306613.

Rizzari:2019:SRB

[RG19] Justin R. Rizzari and Caleb Gardner. Supply risk of bait in Australia’s Southern Rock Lobster Fishery. Marine Policy, 108(??):Article 103659, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302623.

Ronchi:2019:FLA

[RGB^{+}19] Francesca Ronchi, Francois Galgani, Flavia Binda, Milica Mandić, Monika Peterlin, Pero Tutman, Aikaterini Anastasopoulou, and Tomaso Fortibuoni. Fishing for Litter in the Adriatic-Ionian macroregion (Mediterranean Sea): Strengths, weaknesses, opportunities and threats. Marine Policy, 100(??):226–237, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301301.

Reimer:2016:MVS

[RGBT16] Julie Reimer, Caroline Gravel, Moira W. Brown, and Christopher T. Taggart. Mitigating vessel strikes: the problem of the peripatetic whales and the peripatetic fleet. Marine Policy, 68(??):91–99, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000695.

Ruff:2019:GMC

[RGC^{+}19] Elizabeth O. Ruff, Rebecca R. Gentry, Tyler Clavelle, Lennon R. Thomas, and Sarah E. Lester. Governance and mariculture in the Caribbean. Marine Policy, 107(??):Article 103565, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300107.

Rahimi:2016:FDI

[RGG^{+}16] Shireen Rahimi, Steven D. Gaines, Stefan Gelcich, Robert Deacon, and Dan Ovando. Factors driving the implementation of fishery reforms. Marine Policy, 71(??): 222–228, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301452.

Rivera:2017:HMC

[RGGFA17] Antonella Rivera, Stefan Gelcich, Lucía García-Flórez, and José Luis Acuña. Heterogeneous management and conservation perceptions within the gooseneck barnacle comanagement system in Asturias (N. Spain). Marine Policy, 81(??):229–235, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16305553.

Rohe:2019:LPP

[RGSF19] Janne R. Rohe, Hugh Govan, Achim Schlüter, and Sebastian C. A. Ferse. A legal pluralism perspective on coastal fisheries governance in two Pacific Island countries. Marine Policy, 100(??):90–97, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307601.

Richardson:2018:UCG

[RGWH18] Kelsey Richardson, Riki Gunn, Chris Wilcox, and Britta Denise Hardesty. Understanding causes of gear loss provides a sound basis for fisheries management. Marine Policy, 96(??):278–284, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307406.

Rumson:2017:CRA

[RHB17] Alexander G. Rumson, Stephen H. Hallett, and Timothy R. Brewer. Coastal risk adaptation: the potential role of accessible geospatial Big Data. Marine Policy, 83(??):100–110, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301756.

Robbins:2017:ATA

[RHP^{+}17] William D. Robbins, Charlie Huveneers, Guido J. Parra, Luciana Möller, and Bronwyn M. Gillanders. Anthropogenic threat assessment of marine-associated fauna in Spencer Gulf, South Australia. Marine Policy, 81 (??):392–400, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16306352.

Reiblich:2019:BCS

[RHW^{+}19] J. Reiblich, E. Hartge, L. M. Wedding, S. Killian, and G. M. Verutes. Bridging climate science, law, and policy to advance coastal adaptation planning. Marine Policy, 104(??):125–134, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306249.

Rahim:2016:RGS

[RIK16] Mia Mahmudur Rahim, Md. Tarikul Islam, and Sanjaya Kuruppu. Regulating global shipping corporations’ accountability for reducing greenhouse gas emissions in the seas. Marine Policy, 69(??):159–170, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301907.

Ringbom:2018:RSS

[Rin18] Henrik Ringbom. Regulation of ship-source pollution in the Baltic Sea. Marine Policy, 98(??):246–254, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306742.

Roberts:2013:NEK

[RJ13] Thomas Roberts and Peter J. S. Jones. North East Kent European marine site: Overcoming barriers to conservation through community engagement. Marine Policy, 41 (??):33–40, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002576.

Ringbom:2018:CAC

[RJ18a] Henrik Ringbom and Marko Joas. Concluding article: the changing regulatory landscape of the Baltic Sea — an analysis. Marine Policy, 98(??):317–324, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307012.

Ringbom:2018:IML

[RJ18b] Henrik Ringbom and Marko Joas. Introduction: Multilevel regulation in the Baltic Sea region. Marine Policy, 98 (??):187–190, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306833.

Rose:2011:SBS

[RJPS11] N. A. Rose, D. Janiger, E. C. M. Parsons, and M. Stachowitsch. Shifting baselines in scientific publications: a case study using cetacean research. Marine Policy, 35(4):477–482, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002009.

Rudd:2013:SPG

[RL13] Murray A. Rudd and Ricky N. Lawton. Scientists’ prioritization of global coastal research questions. Marine Policy, 39(??):101–111, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1200190X.

Rossiter:2014:WMS

[RL14] Jaime Speed Rossiter and Arielle Levine. What makes a “successful” marine protected area? The unique context of Hawai’i’s fish replenishment areas. Marine Policy, 44(??):196–203, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001760.

Rollins:2019:CFH

[RL19] Emily Rollins and Sabrina Lovell. Charter fishing in Hawaii: a multi-region analysis of the economic linkages and contributions within and outside Hawaii. Marine Policy, 100(??):277–287, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303610.

Rabbitt:2019:WCW

[RLAT19] Sheridan Rabbitt, Ian Lilley, Simon Albert, and Ian R. Tibbetts. What’s the catch in who fishes? Fisherwomen’s contributions to fisheries and food security in Marovo Lagoon, Solomon Islands. Marine Policy, 108(??):Article 103667, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308005.

Richerson:2018:PEI

[RLH18] Kate Richerson, Jerry Leonard, and Daniel S. Holland. Predicting the economic impacts of the 2017 West Coast salmon troll ocean fishery closure. Marine Policy, 95(??):142–152, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306668.

Rodwell:2014:FCM

[RLHM14] Lynda D. Rodwell, Jason Lowther, Charlotte Hunter, and Stephen C. Mangi. Fisheries co-management in a new era of marine policy in the UK: a preliminary assessment of stakeholder perceptions. Marine Policy, 45(??):279–286, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002054.

Ratsimbazafy:2019:EDM

[RLKH19] Hajaniaina Ratsimbazafy, Thierry Lavitra, Marc Kochzius, and Jean Hugé. Emergence and diversity of marine protected areas in Madagascar. Marine Policy, 105 (??):91–108, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304937.

Richerson:2010:AIE

[RLM10] Kate Richerson, Phillip S. Levin, and Marc Mangel. Accounting for indirect effects and non-commensurate values in ecosystem based fishery management (EBFM). Marine Policy, 34(1):114–119, January 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000748.

Richerson:2012:EAI

[RLM12] Kate Richerson, Phillip S. Levin, and Marc Mangel. Erratum to “Accounting for indirect effects and non-commensurate values in ecosystem based fishery management (EBFM)” [Mar. Policy 34 (2010) 114– 119]. Marine Policy, 36(2):565, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001448.

Rahman:2016:PCR

[RM16] S. M. Mizanur Rahman and Audrey L. Mayer. Policy compliance recommendations for international shipbreaking treaties for Bangladesh. Marine Policy, 73(??): 122–129, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304560.

Raubenheimer:2017:MPM

[RM17] Karen Raubenheimer and Alistair McIlgorm. Is the Montreal Protocol a model that can help solve the global marine plastic debris problem? Marine Policy, 81(??):322–329, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307096.

Raubenheimer:2018:CBS

[RM18] Karen Raubenheimer and Alistair McIlgorm. Can the Basel and Stockholm Conventions provide a global framework to reduce the impact of marine plastic litter? Marine Policy, 96(??):285–290, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305146.

Rindorf:2017:MBM

[RMB^{+}17] Anna Rindorf, John Mumford, Paul Baranowski, Lotte Worsøe Clausen, Dorleta García, Niels T. Hintzen, Alexander Kempf, Adrian Leach, Polina Levontin, Pamela Mace, Steven Mackinson, Christos Maravelias, Raúl Prellezo, Antoni Quetglas, George Tserpes, Rüdiger Voss, and David Reid. Moving beyond the MSY concept to reflect multidimensional fisheries management objectives. Marine Policy, 85(??):33–41, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307801.

Rees:2015:SEE

[RMH^{+}15] Siân E. Rees, Stephen C. Mangi, Caroline Hattam, Sarah C. Gall, Lynda D. Rodwell, Frankie J. Peckett, and Martin J. Attrill. The socio-economic effects of a Marine Protected Area on the ecosystem service of leisure and recreation. Marine Policy, 62(??):144–152, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002626.

Roberts:2012:SCL

[RMJ12] Stephen E. Roberts, Peter B. Marlow, and Bogdan Jaremin. Shipping casualties and loss of life in UK merchant shipping, UK second register and foreign flags used by UK shipping companies. Marine Policy, 36(3):703–712, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001849.

Roberts:2017:ACR

[RMR^{+}17] Annette Roberts, Max Munday, Neil Roche, Adam Brown, Mike Armstrong, Jodie Hargreaves, Sarah Pilgrim-Morrison, Kevin Williamson, and Kieran Hyder. Assessing the contribution of recreational sea angling to the English economy. Marine Policy, 83(??):146–152, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305474.

Ramirez-Monsalve:2016:ICP

[RMRN^{+}16a] P. Ramírez-Monsalve, J. Raakjær, K. N. Nielsen, U. Laksá, R. Danielsen, D. Degnbol, M. Ballesteros, and P. Degnbol. Institutional challenges for policy-making and fisheries advice to move to a full EAFM approach within the current governance structures for marine policies. Marine Policy, 69(??):1–12, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300483.

Ramirez-Monsalve:2016:EAF

[RMRN^{+}16b] P. Ramírez-Monsalve, J. Raakjær, K. N. Nielsen, J. L. Santiago, M. Ballesteros, U. Laksá, and P. Degnbol. Ecosystem Approach to Fisheries Management (EAFM) in the EU — Current science–policy–society interfaces and emerging requirements. Marine Policy, 66(??):83–92, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000221.

Rojas-Nazar:2015:MRE

[RNCGB15] U. A. Rojas-Nazar, R. Cullen, J. P. A. Gardner, and J. J. Bell. Marine reserve establishment and on-going management costs: a case study from New Zealand. Marine Policy, 60(??):216–224, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001943.

Rojas-Nazar:2019:LFM

[RNEBG19] Ursula A. Rojas-Nazar, Tyler D. Eddy, James J. Bell, and Jonathan P. A. Gardner. Lobster fishery and marine reserve interactions in central New Zealand. Marine Policy, 105(??):67–79, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307206.

Rola:2018:ICF

[RNN^{+}18] Agnes C. Rola, Teresita A. Narvaez, Maria Rio A. Naguit, Dulce D. Elazegui, Bing Baltazar C. Brillo, Merlyne M. Paunlagui, Hadji C. Jalotjot, and Catherine P. Cervantes. Impact of the closed fishing season policy for sardines in Zamboanga Peninsula, Philippines. Marine Policy, 87(??):40–50, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730338X.

Robinson:2019:CPL

[RNS19] Danielle Robinson, Steven P. Newman, and Selina M. Stead. Community perceptions link environmental decline to reduced support for tourism development in small island states: a case study in the Turks and Caicos Islands. Marine Policy, 108(??):Article 103671, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300149.

Rosen:2013:IEG

[RO13] Franciska Rosen and Per Olsson. Institutional entrepreneurs, global networks, and the emergence of international institutions for ecosystem-based management: the Coral Triangle Initiative. Marine Policy, 38(??):195–204, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001340.

Reilly:2015:APF

[ROD15] Kieran Reilly, Anne Marie O’Hagan, and Gordon Dalton. Attitudes and perceptions of fishermen on the island of Ireland towards the development of marine renewable energy projects. Marine Policy, 58(??):88–97, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000834.

Rogerson:2015:BHT

[Rog15] Jennifer J. M. Rogerson. Being heard: Thinking through different versions of rationality, epistemological policing and dissonances in marine conservation. Marine Policy, 60(??):325–330, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002474.

Ross:2013:ECF

[Ros13] Natalie Ross. Exploring concepts of fisheries ‘dependency’ and ‘community’ in Scotland. Marine Policy, 37 (??):55–61, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000644.

Rosello:2017:CUF

[Ros17] Mercedes Rosello. Cooperation and unregulated fishing: interactions between customary international law, and the European Union IUU fishing regulation. Marine Policy, 84(??):306–312, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302531.

Rosales:2018:SMS

[Ros18] Rina Maria P. Rosales. SEAT: Measuring socio-economic benefits of marine protected areas. Marine Policy, 92 (??):120–130, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300459.

Roy:2016:LCA

[Roy16] Anjan Kumer Dev Roy. Local community attitudes towards mangrove forest conservation: Lessons from Bangladesh. Marine Policy, 74(??):186–194, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301282.

Ricci:2017:ECM

[RPA17] Giovanna Ricci, Isabel Pirillo, and Francesco Amenta. Ethical challenges to medical assistance at sea. Marine Policy, 81(??):247–249, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302324.

Rosales:2017:VCA

[RPC^{+}17] Rina Maria Rosales, Robert Pomeroy, Ina Judith Calabio, Mabel Batong, Kimakarla Cedo, Nestor Escara, Vivien Facunla, Anecita Gulayan, Manuel Narvadez, Madzni Sarahadil, and Maria Aimee Sobrevega. Value chain analysis and small-scale fisheries management. Marine Policy, 83(??):11–21, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300908.

Rangel:2019:DFA

[RPdO^{+}19] Mafalda Rangel, Cristina Pita, Manuela Maria de Oliveira, Maria Helena Guimarães, Rita Rainha, Carlos Sonderblohm, Pedro Monteiro, Frederico Oliveira, Marta Ballesteros, Jorge M. S. Gonçalves, Graham J. Pierce, and Karim Erzini. Do fisher associations really represent their members’ needs and opinions? The case study of the octopus fishery in the Algarve (south Portugal). Marine Policy, 101(??):276–284, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308928.

Rangel:2014:DSG

[RPG^{+}14] M. O. Rangel, C. B. Pita, J. M. S. Gonçalves, F. Oliveira, C. Costa, and K. Erzini. Developing self-guided scuba dive routes in the Algarve (Portugal) and analysing visitors’ perceptions. Marine Policy, 45(??):194–203, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002364.

Rangel:2015:ETS

[RPG^{+}15] M. O. Rangel, C. B. Pita, J. M. S. Gonçalves, F. Oliveira, C. Costa, and K. Erzini. Eco-touristic snorkelling routes at Marinha beach (Algarve): Environmental education and human impacts. Marine Policy, 60(??):62–69, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001591.

Roberts:2013:CLL

[RPM13] Stephen E. Roberts, Stephen J. Pettit, and Peter B. Marlow. Casualties and loss of life in bulk carriers from 1980 to 2010. Marine Policy, 42(??):223–235, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000547.

Rehmatulla:2017:WTO

[RPSS17] Nishatabbas Rehmatulla, Sophia Parker, Tristan Smith, and Victoria Stulgis. Wind technologies: Opportunities and barriers to a low carbon shipping industry. Marine Policy, 75(??):217–226, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003917.

Rendle:2014:ASR

[RR14] Emma J. Rendle and Lynda D. Rodwell. Artificial surf reefs: a preliminary assessment of the potential to enhance a coastal economy. Marine Policy, 45(??):349–358, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002017.

Ramesh:2017:TCM

[RR17] Madhuri Ramesh and Nitin D. Rai. Trading on conservation: a marine protected area as an ecological fix. Marine Policy, 82(??):25–31, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301227.

Rodriguez-Rodriguez:2018:RAP

[RR18] David Rodríguez-Rodríguez. Rapid assessment of protection and ecological effectiveness of the Spanish Fishing Reserve Network. Marine Policy, 90(??):29–36, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308345.

Rees:2010:VMB

[RRA^{+}10] Siân E. Rees, Lynda D. Rodwell, Martin J. Attrill, Melanie C. Austen, and Steven C. Mangi. The value of marine biodiversity to the leisure and recreation industry and its application to marine spatial planning. Marine Policy, 34(5):868–875, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000102.

Ramirez-Rodriguez:2013:SJS

[RRAH13] Mauricio Ramírez-Rodríguez and Luis César Almendárez-Hernández. Subsidies in the jumbo squid fishery in the Gulf of California, Mexico. Marine Policy, 40(??):117–123, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000122.

Ruiz:2017:BEV

[RRC^{+}17] Javier Ruiz, Margarita María Rincón, David Castilla, Fernando Ramos, and Juan José García del Hoyo. Biological and economic vulnerabilities of fixed TACs in small pelagics: an analysis of the European anchovy (Engraulis encrasicolus) in the Gulf of Cádiz. Marine Policy, 78 (??):171–180, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16303852.

Rosenhaim:2019:SBW

[RRCS^{+}19] Ingrid Linck Rosenhaim, Kathrin Riemann-Campe, Hiroshi Sumata, Cornelia Koeberle, Ralf Brauner, Andreas Herber, and Rüdiger Gerdes. Simulated ballast water accumulation along Arctic shipping routes. Marine Policy, 103(??):9–18, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304421.

Reeves:2012:IAI

[RRG^{+}12] Randall Reeves, Cheryl Rosa, J. Craig George, Gay Sheffield, and Michael Moore. Implications of Arctic industrial growth and strategies to mitigate future vessel and fishing gear impacts on bowhead whales. Marine Policy, 36(2):454–462, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001400.

Rodriguez-Rodriguez:2016:ABG

[RRMSSE16] D. Rodríguez-Rodríguez, D. Abdul Malak, T. Soukissian, and A. Sánchez-Espinosa. Achieving Blue Growth through maritime spatial planning: Offshore wind energy optimization and biodiversity conservation in Spain. Marine Policy, 73(??):8–14, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303517.

Ramirez-Rodriguez:2012:SMS

[RROR12] Mauricio Ramírez-Rodríguez and Miguel Ángel Ojeda-Ruíz. Spatial management of small-scale fisheries on the west coast of Baja California Sur, Mexico. Marine Policy, 36(1):108–112, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000753.

Rodriguez-Rodriguez:2017:MDM

[RRR17] Gonzalo Rodríguez-Rodríguez and Roberto Bande Ramudo. Market driven management of climate change impacts in the Spanish mussel sector. Marine Policy, 83 (??):230–235, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302105.

Rodriguez-Rodriguez:2015:SMP

[RRRM^{+}15] D. Rodríguez-Rodríguez, S. Rees, G. Mannaerts, M. Sciberras, C. Pirie, G. Black, C. Aulert, E. V. Sheehan, S. Carrier, and M. J. Attrill. Status of the marine protected area network across the English Channel (La Manche): Cross-country similarities and differences in MPA designation, management and monitoring. Marine Policy, 51(??):536–546, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002541.

Rodriguez-Rodriguez:2016:MPA

[RRRM^{+}16] David Rodríguez-Rodríguez, J. Rodríguez, D. Abdul Malak, A. Nastasi, and P. Hernández. Marine protected areas and fisheries restricted areas in the Mediterranean: assessing “actual” marine biodiversity protection coverage at multiple scales. Marine Policy, 64(??):24–30, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003255.

Rush:2017:CPC

[RS17] S. Rush and J. L. Solandt. Challenges on providing conservation advice for a growing network of English Marine Protected Areas. Marine Policy, 83(??):75–82, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302822.

Rochwulaningsih:2019:MPB

[RSMM19] Yety Rochwulaningsih, Singgih Tri Sulistiyono, Noor Naelil Masruroh, and Nazala Noor Maulany. Marine policy basis of Indonesia as a maritime state: the importance of integrated economy. Marine Policy, 108(??):Article 103602, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301415.

Reis:2014:RSA

[RSS14] J. Reis, T. Stojanovic, and H. Smith. Relevance of systems approaches for implementing Integrated Coastal Zone Management principles in Europe. Marine Policy, 43 (??):3–12, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000894.

Rehmatulla:2017:RBE

[RST17] Nishatabbas Rehmatulla, Tristan Smith, and Laurienne Tibbles. The relationship between EU’s public procurement policies and energy efficiency of ferries in the EU. Marine Policy, 75(??):278–289, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003887.

Ramirez:2017:HWE

[RTD^{+}17] I. Ramirez, M. Tarzia, M. P. Dias, I. J. Burfield, J. A. Ramos, S. Garthe, and V. H. Paiva. How well is the EU protecting its seabirds? Progress in implementing the Birds Directive at sea. Marine Policy, 81(??):179–184, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306510.

Roux:2019:SSF

[RTM19a] Marie-Julie Roux, Ross F. Tallman, and Zoya A. Martin. Small-scale fisheries in Canada’s Arctic: Combining science and fishers knowledge towards sustainable management. Marine Policy, 101(??):177–186, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307631.

Ryan:2019:QSD

[RTM^{+}19b] Karina L. Ryan, Stephen M. Taylor, Rory McAuley, Gary Jackson, and Brett W. Molony. Quantifying shark depredation events while commercial, charter and recreational fishing in Western Australia. Marine Policy, 109(??):Article 103674, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300831.

Roberts:2012:PUM

[RU12] Thomas Roberts and Paul Upham. Prospects for the use of macro-algae for fuel in Ireland and the UK: an overview of marine management issues. Marine Policy, 36(5):1047–1053, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000449.

Rockmann:2012:AVP

[RUD^{+}12] Christine Röckmann, Clara Ulrich, Marion Dreyer, Ewen Bell, Edward Borodzicz, Päivi Haapasaari, Kjellrun Hiis Hauge, Daniel Howell, Samu Mäntyniemi, David Miller, George Tserpes, and Martin Pastoors. The added value of participatory modelling in fisheries management — what has been learnt? Marine Policy, 36(5):1072–1085, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000462.

Rochette:2014:RAC

[RUH^{+}14] Julien Rochette, Sebastian Unger, Dorothée Herr, David Johnson, Takehiro Nakamura, Tim Packeiser, Alexander Proelss, Martin Visbeck, Andrew Wright, and Daniel Cebrian. The regional approach to the conservation and sustainable use of marine biodiversity in areas beyond national jurisdiction. Marine Policy, 49(??):109–117, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000438.

Rivero:2016:WRP

[RV16a] Susana Rivero and Sebastian Villasante. What are the research priorities for marine ecosystem services? Marine Policy, 66(??):104–113, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X16000270.

Rodrigues:2016:DSV

[RV16b] João Garcia Rodrigues and Sebastián Villasante. Disentangling seafood value chains: Tourism and the local market driving small-scale fisheries. Marine Policy, 74 (??):33–42, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300446.

Rodrigues:2013:SEI

[RvdBG13] Luís C. Rodrigues, Jeroen C. J. M. van den Bergh, and Andrea Ghermandi. Socio-economic impacts of ocean acidification in the Mediterranean Sea. Marine Policy, 38 (??):447–456, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001650.

Rodriguez:2011:RTA

[RVdCGN11] Gonzalo Rodríguez Rodríguez, Sebastián Villasante, and María do Carme García-Negro. Are red tides affecting economically the commercialization of the Galician (NW Spain) mussel farming? Marine Policy, 35(2):252–257, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1000148X.

Rockmann:2015:ITT

[RvLG^{+}15] Christine Röckmann, Judith van Leeuwen, David Goldsborough, Marloes Kraan, and Gerjan Piet. The interaction triangle as a tool for understanding stakeholder interactions in marine ecosystem based management. Marine Policy, 52(??):155–162, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002814.

Raakjaer:2014:EBM

[RvLvTH14] Jesper Raakjaer, Judith van Leeuwen, Jan van Tatenhove, and Maria Hadjimichael. Ecosystem-based marine management in European regional seas calls for nested governance structures and coordination — a policy brief. Marine Policy, 50(??):373–381, December 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000797.

Raakjaer:2014:MGE

[RvT14]  Jesper Raakjær and Jan van Tatenhove.  Marine

governance of European Seas: Introduction. Marine Policy, 50(??):323–324, December 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000748.

Rudders:2015:OPE

[RW15] David B. Rudders and John M. Ward. Own-price elasticity of open access supply as a long-run measure of fish stock abundance. Marine Policy, 53(??):215–226, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002656.

Rousseau:2019:DGA

[RWBF19] Yannick Rousseau, Reg A. Watson, Julia L. Blanchard, and Elizabeth A. Fulton. Defining global artisanal fisheries. Marine Policy, 108(??):Article 103634, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309412.

Rieser:2013:TFC

[RWG13] Alison Rieser, Les Watling, and John Guinotte. Trawl fisheries, catch shares and the protection of benthic marine ecosystems: Has ownership generated incentives for seafloor stewardship? Marine Policy, 40(??):75–83, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12002692.

Rickels:2019:DEU

[RWG^{+}19] Wilfried Rickels, Christian Weigand, Patricia Grasse, Jörn Schmidt, and Rüdiger Voss. Does the European Union achieve comprehensive blue growth? Progress of EU coastal states in the Baltic and North Sea, and the Atlantic Ocean against sustainable development goal 14. Marine Policy, 106(??):Article 103515, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306614.

Ren:2018:RCM

[RWJ18] Wenhan Ren, Qi Wang, and Jianyue Ji. Research on China’s marine economic growth pattern: an empirical analysis of China’s eleven coastal regions. Marine Policy, 87(??):158–166, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17304426.

Read:2015:UCD

[RWK15] Andrew D. Read, Ronald J. West, and Brendan P. Kelaher. Using compliance data to improve marine protected area management. Marine Policy, 60(??):119–127, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001670.

Rehr:2014:TUC

[RWL14] Amanda P. Rehr, Gregory D. Williams, and Phillip S. Levin. A test of the use of computer generated visualizations in support of ecosystem-based management. Marine Policy, 46(??):14–18, May 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002959.

Robinson:2014:LDM

[RWW^{+}14] James P. W. Robinson, Easton R. White, Logan D. Wiwchar, Danielle C. Claar, Justin P. Suraci, and Julia K. Baum. The limitations of diversity metrics in directing global marine conservation. Marine Policy, 48(??):123–125, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000840.

Raymond-Yakoubian:2018:IAO

[RYD18] Julie Raymond-Yakoubian and Raychelle Daniel. An Indigenous approach to ocean planning and policy in the Bering Strait region of Alaska. Marine Policy, 97(??): 101–108, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300113.

Rust:2017:ECE

[RYJ^{+}17] Steven Rust, Satoshi Yamazaki, Sarah Jennings, Timothy Emery, and Caleb Gardner. Excess capacity and efficiency in the quota managed Tasmanian Rock Lobster Fishery. Marine Policy, 76(??):55–62, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304663.

Raymond-Yakoubian:2017:ITK

[RYRYM17] Julie Raymond-Yakoubian, Brenden Raymond-Yakoubian, and Catherine Moncrieff. The incorporation of traditional knowledge into Alaska federal fisheries management. Marine Policy, 78(??):132–142, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307825.

Standal:2018:HRN

[SA18] Dag Standal and Frank Asche. Hesitant reforms: the Norwegian approach towards ITQ’s. Marine Policy, 88 (??):58–63, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303561.

Solis:2015:ITF

[SAdC15] Daniel Solís, Juan J. Agar, and Julio del Corral. IFQs and total factor productivity changes: the case of the Gulf of Mexico red snapper fishery. Marine Policy, 62(??):347–357, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001608.

Soanes:2013:ICF

[SAGG13] L. M. Soanes, P. W. Atkinson, R. D. Gauvain, and J. A. Green. Individual consistency in the foraging behaviour of Northern Gannets: Implications for interactions with offshore renewable energy developments. Marine Policy, 38(??):507–514, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200173X.

Steadman:2014:MUY

[SAH14] Daniel Steadman, Thomas Appleby, and Julie Hawkins. Minimising unsustainable yield: Ten failing European fisheries. Marine Policy, 48(??):192–201, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001109.

Sala:2016:RES

[Sal16] Antonello Sala. Review of the EU small-scale driftnet fisheries. Marine Policy, 74(??):236–244, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302706.

Singleton:2017:CRF

[SALS17] R. L. Singleton, E. H. Allison, P. Le Billon, and U. R. Sumaila. Conservation and the right to fish: International conservation NGOs and the implementation of the Voluntary Guidelines for securing Sustainable Small-Scale Fisheries. Marine Policy, 84(??):22–32, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302762.

Sigfusson:2013:EII

[SAM13] Thor Sigfusson, Ragnar Arnason, and Karyn Morrissey. The economic importance of the Icelandic fisheries cluster — Understanding the role of fisheries in a small economy. Marine Policy, 39(??):154–161, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200214X.

Sandstrom:2010:ISU

[San10] Annica Sandström. Institutional and substantial uncertainty — Explaining the lack of adaptability in fish stocking policy. Marine Policy, 34(6):1357–1365, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001156.

Sandstrom:2011:NCP

[San11] Annica Sandström. Navigating a complex policy system — Explaining local divergences in Swedish fish stocking policy. Marine Policy, 35(3):419–425, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002137.

Santos:2015:FWL

[San15] Anna N. Santos. Fisheries as a way of life: Gendered livelihoods, identities and perspectives of artisanal fisheries in eastern Brazil. Marine Policy, 62(??):279–288, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002584.

Sangiuliano:2018:APE

[San18] Stephen J. Sangiuliano. Analysing the potentials and effects of multi-use between tidal energy development and environmental protection and monitoring: a case study of the inner sound of the Pentland Firth. Marine Policy, 96 (??):120–132, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300368.

Sangiuliano:2019:APE

[San19] Stephen J. Sangiuliano. Analysing the prevalence of ecosystem services in the objectives and policies of Scotland’s National Marine Plan. Marine Policy, 104(??):37–52, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18302720.

Saenz-Arroyo:2016:LEK

[SARF16] Andrea Sáenz-Arroyo and Daniel Revollo-Fernández. Local ecological knowledge concurs with fishing statistics: an example from the abalone fishery in Baja California, Mexico. Marine Policy, 71(??):217–221, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302585.

Satizabal:2018:UCR

[Sat18] Paula Satizábal. The unintended consequences of ‘responsible fishing’ for small-scale fisheries: Lessons from the Pacific coast of Colombia. Marine Policy, 89(??):

50–57, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303822.

Schernewski:2014:IBT

[SB14] Gerald Schernewski and Steffen Bock. Internet based training and education for coastal management in Germany: a critical evaluation. Marine Policy, 43(??):21–28, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13000912.

Santos:2015:LSM

[SB15] Anna N. Santos and Christian Brannstrom. Livelihood strategies in a marine extractive reserve: Implications for conservation interventions. Marine Policy, 59(??): 44–52, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15001347.

Sutcliffe:2018:RSE

[SB18] Sarah R. Sutcliffe and Michele L. Barnes. The role of shark ecotourism in conservation behaviour: Evidence from Hawaii. Marine Policy, 97(??):27–33, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301180.

Sheaves:2014:RRA

[SBC^{+}14] Marcus Sheaves, Justin Brookes, Rob Coles, Marnie Freckelton, Paul Groves, Ross Johnston, and Pia Winberg. Repair and revitalisation of Australia’s tropical estuaries and coastal wetlands: Opportunities and constraints for the reinstatement of lost function and productivity. Marine Policy, 47(??):23–38, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000402.

Sandstrom:2015:NGT

[SBC15a] Annica Sandström, Örjan Bodin, and Beatrice Crona. Network Governance from the top — The case of ecosystembased coastal and marine management. Marine Policy, 55(??):57–63, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000202.

Santiago:2015:ERR

[SBC^{+}15b] Jose L. Santiago, Marta A. Ballesteros, Rosa Chapela, Cristina Silva, Kåre N. Nielsen, Mafalda Rangel, Karim Erzini, Laura Wise, Aida Campos, Maria F. Borges, Antonello Sala, Massimo Virgili, Jónas R. Vidarsson, Alan Baudron, and Paul G. Fernandes. Is Europe ready for a results-based approach to fisheries management? The voice of stakeholders. Marine Policy, 56(??):86–97, June 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500041X.

Soanes:2016:IFA

[SBC^{+}16] L. M. Soanes, J. A. Bright, D. Carter, M. P. Dias, T. Fleming, K. Gumbs, G. Hughes, F. Mukhida, and J. A. Green. Important foraging areas of seabirds from Anguilla, Caribbean: Implications for marine spatial planning. Marine Policy, 70(??):85–92, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301919.

Seto:2017:WFF

[SBM^{+}17] Katherine Seto, Dyhia Belhabib, Josephus Mamie, Duncan Copeland, Jan Michael Vakily, Heiko Seilert, Andrew Baio, Sarah Harper, Dirk Zeller, Kyrstn Zylich, and Daniel Pauly. War, fish, and foreign fleets: the marine fisheries catches of Sierra Leone 1950–2015. Marine Policy, 83(??):153–163, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308284.

Silva:2013:REI

[SBMD13] Catarina N. S. Silva, Matt K. Broadhurst, Rodrigo P. Medeiros, and José H. Dias. Resolving environmental issues in the southern Brazilian artisanal penaeidtrawl fishery through adaptive co-management. Marine Policy, 42(??):133–141, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000456.

Seeteram:2019:REI

[SBP^{+}19] Nadia Seeteram, Mahadev Bhat, Brett Pierce, Kevin Cavasos, and David Die. Reconciling economic impacts and stakeholder perception: a management challenge in Florida Gulf Coast fisheries. Marine Policy, 108(??): Article 103628, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18305530.

Stelzenmuller:2013:MES

[SBS^{+}13] Vanessa Stelzenmüller, Patricia Breen, Tammy Stamford, Frank Thomsen, Fabio Badalamenti, Ángel Borja, Lene Buhl-Mortensen, Julia Carlstöm, Giovanni D’Anna, Norbert Dankers, Steven Degraer, Mike Dujin, Fabio Fiorentino, Ibon Galparsoro, Sylvaine Giakoumi, Michele Gristina, Kate Johnson, Peter J. S. Jones, Stelios Katsanevakis, Leyla Knittweis, Zacharoula Kyriazi, Carlo Pipitone, Joanna Piwowarczyk, Marijn Rabaut, Thomas K. Sørensen, Jan van Dalfsen, Vassiliki Vassilopoulou, Tomás Vega Fernández, Magda Vincx, Sandra Vöge, Anke Weber, Nicklas Wijkmark, Robbert Jak, Wanfei Qiu, and Remment ter Hofstede. Monitoring and evaluation of spatially managed areas: a generic framework for implementation of ecosystem based marine management and its application. Marine Policy, 37(??):149–164, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000735.

Soininen:2019:TIF

[SBSK19] Niko Soininen, Antti Belinskij, Jukka Similä, and Raine Kortet. Too important to fail? Evaluating legal adaptive capacity for increasing coastal and marine aquaculture production in EU–Finland. Marine Policy, 110(??):Article 103498, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308820.

Sumaila:2016:FFO

[SBT16] U. Rashid Sumaila, Christophe Bellmann, and Alice Tipping. Fishing for the future: an overview of challenges and opportunities. Marine Policy, 69(??):173–180, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000087.

Sawchuk:2015:USE

[SBTF15] Jennifer Heibult Sawchuk, Anne H. Beaudreau, Daniel Tonnes, and David Fluharty. Using stakeholder engagement to inform endangered species management and improve conservation. Marine Policy, 54(??):98–107, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14003510.

Spalding:2017:MGV

[SBW^{+}17] Mark Spalding, Lauretta Burke, Spencer A. Wood, Joscelyne Ashpole, James Hutchison, and Philine zu Ermgassen. Mapping the global value and distribution of coral reef tourism. Marine Policy, 82(??):104–113, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300635.

Sowman:2010:SSF

[SC10] Merle Sowman and Paula Cardoso. Small-scale fisheries and food security strategies in countries in the Benguela Current Large Marine Ecosystem (BCLME) region: Angola, Namibia and South Africa. Marine Policy, 34(6):1163–1170, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000680.

Stewart:2011:QCN

[SC11] James Stewart and Peter Callagher. Quota concentration in the New Zealand fishery: Annual catch entitlement and the small fisher. Marine Policy, 35(5):631–646, September 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000273.

Saarman:2013:CML

[SC13a] Emily T. Saarman and Mark H. Carr. The California Marine Life Protection Act: a balance of top down and bottom up governance in MPA planning. Marine Policy, 41(??):41–49, September 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000110.

Stewart:2013:IRS

[SC13b] James Stewart and Peter Callagher. Industry response to the 2003 set net restrictions for protection of Maui’s dolphin. Marine Policy, 42(??):210–222, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300064X.

Sloan:2016:APR

[SC16] James Sloan and Kevin Chand. An analysis of property rights in the Fijian qoliqoli. Marine Policy, 72 (??):76–81, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300847.

Said:2019:ASD

[SC19] Alicia Said and Ratana Chuenpagdee. Aligning the sustainable development goals to the small-scale fisheries guidelines: a case for EU fisheries governance. Marine Policy, 107(??):Article 103599, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307553.

Shih:2010:MSF

[SCC10] Yi-Che Shih, Chiu L. Chou, and Wen-Yan Chiau. Maritime safety for fishing boat operations and avoidable hijacking in Taiwan. Marine Policy, 34(2):349–351, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001067.

Squires:2014:SPG

[SCC14] Dale Squires, Raymond Clarke, and Valerie Chan. Subsidies, public goods, and external benefits in fisheries. Marine Policy, 45(??):222–227, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002443.

Schuhbauer:2017:HSA

[SCC^{+}17] Anna Schuhbauer, Ratana Chuenpagdee, William W. L. Cheung, Krista Greer, and U. Rashid Sumaila. How subsidies affect the economic viability of small-scale fisheries. Marine Policy, 82(??):114–121, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730177X.

Sala:2016:FBE

[SCD^{+}16] Enric Sala, Christopher Costello, Jaime De Bourbon Parme, Marco Fiorese, Geoff Heal, Kieran Kelleher, Russell Moffitt, Lance Morgan, Jayne Plunkett, Kristin D. Rechberger, Andrew A. Rosenberg, and Rashid Sumaila. Fish banks: an economic model to scale marine conservation. Marine Policy, 73(??):154–161, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300586.

Santeramo:2017:CPO

[SCD^{+}17] F. G. Santeramo, D. Carlucci, B. De Devitiis, G. Nardone, and R. Viscecchia. On consumption patterns in oyster markets: the role of attitudes. Marine Policy, 79(??):54–61, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300076.

Schumann:2010:APP

[Sch10a] Sarah Schumann. Application of participatory principles to investigation of the natural world: an example from Chile. Marine Policy, 34(6):1196–1202, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000722.

Schumann:2010:TTR

[Sch10b] Sarah Schumann. A tenuous triumvirate: the role of independent biologists in Chile’s co-management regime for shellfish. Marine Policy, 34(1):133–138, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000864.

Schreiber:2012:ELI

[Sch12] Milena Arias Schreiber. The evolution of legal instruments and the sustainability of the Peruvian anchovy fishery. Marine Policy, 36(1):78–89, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000698.

Schratwieser:2015:RSA

[Sch15a] Jason Schratwieser. A rejoinder to Shiffman et al., Trophy fishing for species threatened with extinction: a way forward building on a history of conservation. Marine Policy, 53(??):5–6, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003005.

Schultz:2015:DOR

[Sch15b] Oliver J. Schultz. Defiance and obedience: Regulatory compliance among artisanal fishers in St Helena Bay. Marine Policy, 60(??):331–337, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002450.

Schatz:2017:CFA

[Sch17] Valentin J. Schatz. The contribution of fisheries access agreements to flag State responsibility. Marine Policy, 84 (??):313–319, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730218X.

Song:2013:VIP

[SCJ13] Andrew M. Song, Ratana Chuenpagdee, and Svein Jentoft. Values, images, and principles: What they represent and how they may improve fisheries governance. Marine Policy, 40(??):167–175, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000274.

Sierra-Correa:2015:EBA

[SCK15] Paula Cristina Sierra-Correa and Jaime Ricardo Cantera Kintz. Ecosystem-based adaptation for improving coastal planning for sea-level rise: a systematic review for mangrove coasts. Marine Policy, 51(??):385–393, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002462.

Sebe:2019:DMF

[SCKP19] Maxime Sèbe, Christos, A. Kontovas, and Linwood Pendleton. A decision-making framework to reduce the risk of collisions between ships and whales. Marine Policy, 109(??):Article 103697, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304567.

Salgado:2015:IMA

[SCMS15] Hugo Salgado, Carlos A. Chávez, Montserrat Miller, and John K. Stranlund. ITQ markets with administrative costs: an application to the industrial common sardine and anchovy fishery in Chile. Marine Policy, 62(??):178–185, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002614.

Singh:2018:RAC

[SCMS^{+}18] Gerald G. Singh, Andrés M. Cisneros-Montemayor, Wilf Swartz, William Cheung, J. Adam Guy, Tiff-Annie Kenny, Chris J. McOwen, Rebecca Asch, Jan Laurens Geffert, Colette C. C. Wabnitz, Rashid Sumaila, Quentin Hanich, and Yoshitaka Ota. A rapid assessment of cobenefits and trade-offs among Sustainable Development Goals. Marine Policy, 93(??):223–231, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302026.

Scovazzi:2016:NBI

[Sco16] Tullio Scovazzi. The negotiations for a binding instrument on the conservation and sustainable use of marine biological diversity beyond national jurisdiction. Marine Policy, 70(??):188–191, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301488.

Scott:2018:WLD

[Sco18] Shirley V. Scott. What lessons does the Antarctic Treaty System offer for the future of peaceful relations in the South China Sea? Marine Policy, 87(??):295–300, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305882.

Sun:2017:WAC

[SCOS17] Chin-Hwa Jenny Sun, Fu-Sung Chiang, Matthew Owens, and Dale Squires. Will American consumers pay more for eco-friendly labeled canned tuna? Estimating US consumer demand for canned tuna varieties using scanner data. Marine Policy, 79(??):62–69, May 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304894.

Silva:2019:RCM

[SCR^{+}19] Priscila Silva, Henrique Cabral, Mafalda Rangel, João Pereira, and Cristina Pita. Ready for co-management? Portuguese artisanal octopus fishers’ preferences for management and knowledge about the resource. Marine Policy, 101(??):268–275, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308783.

Sylvia:2014:FCP

[SCS14] Gil Sylvia, Chris Cusack, and Josh Swanson. Fishery cooperatives and the Pacific Whiting Conservation Cooperative: Lessons and application to non-industrial fisheries in the Western Pacific. Marine Policy, 44(??):65–71, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001590.

Stage:2016:ESI

[SCS16] Jesper Stage, Anna Christiernsson, and Patrik Söderholm. The economics of the Swedish individual transferable quota system: Experiences and policy implications. Marine Policy, 66(??):15–20, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1600004X.

Sampaio:2015:PPA

[SCV^{+}15] Flavia D. F. Sampaio, A. Freire Carolina, M. Sampaio Tony Vinícius, R. S. Vitule Jean, and F. Fávaro Luís. The precautionary principle and its approach to risk analysis and quarantine related to the trade of marine ornamental fishes in Brazil. Marine Policy, 51(??):163–168, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400205X.

Steenbergen:2017:UIP

[SCV^{+}17] Dirk J. Steenbergen, Julian Clifton, Leontine E. Visser, Natasha Stacey, and Andrew McWilliam. Understanding influences in policy landscapes for sustainable coastal livelihoods. Marine Policy, 82(??):181–188, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302166.

Salomon:2013:CCS

[SD13] Markus Salomon and Miriam Dross. Challenges in crosssectoral marine protection in Europe. Marine Policy, 42 (??):142–149, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000559.

Schewe:2018:NFO

[SD18] Rebecca L. Schewe and Cassie Dutton. NOAA fishery observers and Vietnamese American fishers in the Southeastern US. Marine Policy, 96(??):145–151, October 2018.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300277.

Sumaila:2014:STF

[SDB14] U. Rashid Sumaila, Andrew Dyck, and Adam Baske. Subsidies to tuna fisheries in the Western Central Pacific Ocean. Marine Policy, 43(??):288–294, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001383.

Schultz:2013:WES

[SDC13] Jessica A. Schultz, Emily S. Darling, and Isabelle M. Côté. What is an endangered species worth? Threshold costs for protecting imperilled fishes in Canada. Marine Policy, 42(??):125–132, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000389.

Santos:2014:HSSb

[SDF^{+}14a] Catarina Frazão Santos, Tiago Domingos, Maria Adelaide Ferreira, Michael Orbach, and Francisco Andrade. How sustainable is sustainable marine spatial planning? Part I — Linking the concepts. Marine Policy, 49(??): 59–65, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14001158.

Santos:2014:HSSa

[SDF^{+}14b] Catarina Frazão Santos, Tiago Domingos, Maria Adelaide Ferreira, Michael Orbach, and Francisco Andrade. How sustainable is sustainable marine spatial planning? Part II — The Portuguese experience. Marine Policy, 49(??):48–58, November 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400116X.

Stevenson:2019:ETA

[SDHS19] Todd C. Stevenson, Jeremy Davies, Henry P. Huntington, and Whit Sheard. An examination of trans-Arctic vessel routing in the Central Arctic Ocean. Marine Policy, 100(??):83–89, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307334.

Scriven:2015:BWM

[SDLT15] Danielle R. Scriven, Claudio DiBacco, Andrea Locke, and Thomas W. Therriault. Ballast water management in Canada: a historical perspective and implications for the future. Marine Policy, 59(??):121–133, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001475.

Shephard:2016:AAV

[SDP^{+}16] Grace Elizabeth Shephard, Kari Dalen, Regina Peldszus, Sara Aparício, Larissa Beumer, Roger Birkeland, Nikolaos Gkikas, Melina Kourantidou, Panagiotis Ktenas, Peter Wilhelm Linde, Francesco Marazzi, Ruth Beatriz Mezzalira Pincinato, Jaziar Radianti, Bernhard Schartmüller, Eike Ingrid Stübner, Alix Varnajot, Magali Vullierme, and Irina Zhilina. Assessing the added value of the recent declaration on unregulated fishing for sustainable governance of the central Arctic Ocean. Marine Policy, 66(??):50–57, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1600018X.

Saavedra-Diaz:2015:SPC

[SDRML15] Lina M. Saavedra-Díaz, Andrew A. Rosenberg, and Berta Martín-López. Social perceptions of Colombian smallscale marine fisheries conflicts: Insights for management. Marine Policy, 56(??):61–70, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003273.

Searcy:2017:BIC

[Sea17a] Taylor Searcy. Bridging islands and calming seas: a material flow management approach to sustainable sea transportation for Fiji’s lower southern Lau Islands.

Marine Policy, 83(??):221–229, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307497.

Searcy:2017:HWC

[Sea17b] Taylor Searcy. Harnessing the wind: a case study of applying Flettner rotor technology to achieve fuel and cost savings for Fiji’s domestic shipping industry. Marine Policy, 86(??):164–172, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303780.

Sen:2010:DFD

[Sen10] Sevaly Sen. Developing a framework for displaced fishing effort programs in marine protected areas. Marine Policy, 34(6):1171–1177, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000692.

Sumaila:2019:UEA

[SES^{+}19] U. Rashid Sumaila, Naazia Ebrahim, Anna Schuhbauer, Daniel Skerritt, Yang Li, Hong Sik Kim, Tabitha Grace Mallory, Vicky W. L. Lam, and Daniel Pauly. Updated estimates and analysis of global fisheries subsidies. Marine Policy, 109(??):Article 103695, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303677.

Seung:2016:ICE

[Seu16] Chang K. Seung. Identifying channels of economic impacts: an inter-regional structural path analysis for Alaska fisheries. Marine Policy, 66(??):39–49, April 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000208.

Seva:2013:CCS

[Sev13] Mikael Sevä. A comparative case study of fish stocking between Sweden and Finland: Explaining differences in decision making at the street level. Marine Policy, 38 (??):287–292, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001431.

Severance:2014:SCC

[Sev14] Craig Severance. Sharing the catch or catching the shares: Catch shares for the western Pacific region? Marine Policy, 44(??):3–8, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001565.

Stojanovic:2013:DWO

[SF13] T. A. Stojanovic and C. J. Q. Farmer. The development of world oceans & coasts and concepts of sustainability. Marine Policy, 42(??):157–165, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000481.

Sun:2016:CED

[SF16] Chin-Hwa Jenny Sun and Leah Fine. A cost-effective discards-proportional at-sea monitoring allocation scheme for the groundfish fishery in New England. Marine Policy, 66(??):75–82, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000063.

Steeves:2019:SPC

[SF19] L. Steeves and R. Filgueira. Stakeholder perceptions of climate change in the context of bivalve aquaculture. Marine Policy, 103(??):121–129, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301118.

Schernewski:2015:IEM

[SFC^{+}15a] Gerald Schernewski, Rene Friedland, Marina Carstens, Ulrike Hirt, Wera Leujak, Günther Nausch, Thomas Neumann, Thorkild Petenati, Sigrid Sagert, Norbert Wasmund, and Mario von Weber. Implementation of European marine policy: New water quality targets for German Baltic waters. Marine Policy, 51(??):305–321, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002358.

Stelzenmuller:2015:AUA

[SFC^{+}15b] Vanessa Stelzenmüller, Tomás Vega Fernández, Katherine Cronin, Christine Röckmann, Maria Pantazi, Jan Vanaverbeke, Tammy Stamford, Kris Hostens, Ellen Pecceu, Steven Degraer, Lene Buhl-Mortensen, Julia Carlström, Ibon Galparsoro, Kate Johnson, Joanna Piwowarczyk, Vassiliki Vassilopoulou, Robbert Jak, Marie Louise Pace, and Luc van Hoof. Assessing uncertainty associated with the monitoring and evaluation of spatially managed areas. Marine Policy, 51(??):151–162, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002036.

Stewart:2014:ESP

[SFCI14] Akilah T. M. Stewart, Bruno D. D. Fragoso, Rodrigo Clímaco, and John D. Icely. Evaluation of stakeholder perspectives on the management of the stalked barnacles (Pollicipes pollicipes) resource in the Parque Natural do Sudoeste Alentejano e Costa Vicentina, Portugal. Marine Policy, 43(??):71–79, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000924.

Seyle:2018:LTI

[SFDB18] D. Conor Seyle, Karina G. Fernandez, Alexander Dimitrevich, and Chirag Bahri. The long-term impact of maritime piracy on seafarers’ behavioral health and work decisions. Marine Policy, 87(??):23–28, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303925.

Scarff:2015:NMM

[SFG15] Gavin Scarff, Clare Fitzsimmons, and Tim Gray. The new mode of marine planning in the UK: Aspirations and challenges. Marine Policy, 51(??):96–102, January 2015.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002000.

Svedang:2012:WRI

[SG12] H. Svedäng and L. Gipperth. Will regionalisation improve fisheries management in the EU? An analysis of the Swedish eel management plan reflects difficulties. Marine Policy, 36(3):801–808, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001965.

Saevaldsson:2015:IPS

[SG15a] Hordur Saevaldsson and Stefan B. Gunnlaugsson. The Icelandic pelagic sector and its development under an ITQ management system. Marine Policy, 61(??):207– 215, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002353.

Siddiki:2015:SAU

[SG15b] Saba Siddiki and Shilpi Goel. A stakeholder analysis of U.S. marine aquaculture partnerships. Marine Policy, 57(??):93–102, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1500055X.

Senapati:2017:SEV

[SG17] Sibananda Senapati and Vijaya Gupta. Socio-economic vulnerability due to climate change: Deriving indicators for fishing communities in Mumbai. Marine Policy, 76 (??):90–97, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307515.

Sundblad:2014:SSD

[SGGM14] Eva-Lotta Sundblad, Anders Grimvall, Lena Gipperth, and Andrea Morf. Structuring social data for the Marine Strategy Framework Directive. Marine Policy, 45(??):1–8, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002467.

Sogn-Grundvaag:2013:VLC

[SGLY13] Geir Sogn-Grundvåg, Thomas A. Larsen, and James A. Young. The value of line-caught and other attributes: an exploration of price premiums for chilled fish in UK supermarkets. Marine Policy, 38(??):41–44, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001054.

Siddons:2018:IRC

[SGMG18] Beth L. Siddons, Gillian Glegg, and Abigail McQuatters-Gollop. Inter-regional coherence: Can Northeast Atlantic pelagic habitat indicators be applied to the Arctic? Marine Policy, 96(??):53–64, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308643.

Sutton-Grier:2014:IES

[SGMWE14] Ariana E. Sutton-Grier, Amber K. Moore, Peter C. Wiley, and Peter E. T. Edwards. Incorporating ecosystem services into the implementation of existing U.S. natural resource management regulations: Operationalizing carbon sequestration and storage. Marine Policy, 43(??):246–253, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13001292.

Smith:2017:HDA

[SGN^{+}17] Carter S. Smith, Rachel K. Gittman, Isabelle P. Neylan, Steven B. Scyphers, Joseph P. Morton, F. Joel Fodrie, Jonathan H. Grabowski, and Charles H. Peterson. Hurricane damage along natural and hardened estuarine shorelines: Using homeowner experiences to promote naturebased coastal protection. Marine Policy, 81(??):350–358, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300477.

Semitiel-Garcia:2019:FPS

[SGNM19] María Semitiel-García and Pedro Noguera-Méndez. Fishers’ participation in small-scale fisheries. A structural analysis of the Cabo de Palos-Islas Hormigas MPA, Spain. Marine Policy, 101(??):257–267, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307613.

Sonderblohm:2017:PAM

[SGP^{+}17] Carlos P. Sonderblohm, Maria H. Guimarães, Cristina Pita, Mafalda Rangel, João Pereira, Jorge M. S. Gonçalves, and Karim Erzini. Participatory assessment of management measures for octopus vulgaris pot and trap fishery from southern Portugal. Marine Policy, 75(??):133–142, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16307187.

Slob:2016:GWS

[SGRV16] Adriaan F. L. Slob, T. R. A. Geerdink, C. Röckmann, and S. Vöge. Governance of the Wadden Sea. Marine Policy, 71(??):325–333, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302391.

Shiffman:2015:LCA

[SGW^{+}15] D. S. Shiffman, A. J. Gallagher, J. Wester, C. C. Macdonald, A. D. Thaler, S. J. Cooke, and N. Hammerschlag. A letter of clarification from the authors of “trophy fishing for species threatened with extinction”. Marine Policy, 53(??):213–214, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003017.

Sogn-Grundvaag:2019:LBF

[SGZI19] Geir Sogn-Grundvåg, Dengjun Zhang, and Audun Iversen. Large buyers at a fish auction: the case of the Norwegian pelagic auction. Marine Policy, 104(??):232–238, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300988.

Symes:2010:FPR

[SH10] David Symes and Ellen Hoefnagel. Fisheries policy, research and the social sciences in Europe: Challenges for the 21st century. Marine Policy, 34(2):268–275, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001031.

Sumaila:2012:MBT

[SH12] Ussif Rashid Sumaila and Ling Huang. Managing Bluefin Tuna in the Mediterranean Sea. Marine Policy, 36(2): 502–511, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1100145X.

Shen:2014:OMF

[SH14a] Gongming Shen and Mikko Heino. An overview of marine fisheries management in China. Marine Policy, 44 (??):265–272, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002091.

Standal:2014:BSO

[SH14b] Dag Standal and Bjørn Hersoug. Back to square one? Fisheries allocation under pressure. Marine Policy, 43 (??):236–245, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001309.

Sunoko:2014:ITF

[SH14c] Rahmadi Sunoko and Hsiang-Wen Huang. Indonesia tuna fisheries development and future strategy. Marine Policy, 43(??):174–183, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001255.

Standal:2015:STB

[SH15] Dag Standal and Bjørn Hersoug. Shaping technology, building society; the industrialization of the Norwegian cod fisheries. Marine Policy, 51(??):66–74, January 2015.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001730.

Sundet:2016:NMI

[SH16] Jan H. Sundet and Alf Håkon Hoel. The Norwegian management of an introduced species: the Arctic red king crab fishery. Marine Policy, 72(??):278–284, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302330.

Shiffman:2017:USS

[SH17] D. S. Shiffman and R. E. Hueter. A United States shark fin ban would undermine sustainable shark fisheries. Marine Policy, 85(??):138–140, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304384.

Stringer:2019:HTF

[SH19] Christina Stringer and Thomas Harré. Human trafficking as a fisheries crime? An application of the concept to the New Zealand context. Marine Policy, 105(??):169–176, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309473.

Shanks:2010:ITF

[Sha10] Steven Shanks. Introducing a transferable fishing day management regime for Pacific Island countries. Marine Policy, 34(5):988–994, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000473.

Shalli:2017:RLT

[Sha17] Mwanahija Salehe Shalli. The role of local taboos in the management of marine fisheries resources in Tanzania. Marine Policy, 85(??):71–78, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300805.

Sheppard:2014:RAP

[She14] Charles Sheppard. Response to the article by Peter Harris: Fortress, safe haven or home? The Chagos MPA in political context. Marine Policy, 47(??):85–86, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000542.

Shelmerdine:2015:TDH

[She15] Richard L. Shelmerdine. Teasing out the detail: How our understanding of marine AIS data can better inform industries, developments, and planning. Marine Policy, 54(??):17–25, April 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003479.

Shepotylo:2016:ENT

[She16] Oleksandr Shepotylo. Effect of non-tariff measures on extensive and intensive margins of exports in seafood trade. Marine Policy, 68(??):47–54, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1600066X.

Shiau:2013:SUM

[Shi13] Tzay-An Shiau. Sea use management using a hybrid operational model: Taiwan’s experience. Marine Policy, 39(??):265–272, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002412.

Shi:2016:RGG

[Shi16] Yubing Shi. Reducing greenhouse gas emissions from international shipping: Is it time to consider market-based measures? Marine Policy, 64(??):123–134, February 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003541.

Schuyler:2018:EIR

[SHL^{+}18] Qamar Schuyler, Britta Denise Hardesty, T. J. Lawson, Kimberley Opie, and Chris Wilcox. Economic incentives reduce plastic inputs to the ocean. Marine Policy, 96 (??):250–255, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305377.

Steckenreuter:2012:SRZ

[SHM12] André Steckenreuter, Robert Harcourt, and Luciana Möller. Are Speed Restriction Zones an effective management tool for minimising impacts of boats on dolphins in an Australian marine park? Marine Policy, 36 (1):258–264, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001023.

Song:2013:BGW

[SHM13] Wei Ling Song, Guang Shun He, and Alistair McIlgorm. From behind the Great Wall: the development of statistics on the marine economy in China. Marine Policy, 39(??):120–127, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001959.

Sambrook:2014:CCC

[SHS^{+}14] Katie Sambrook, Rohan H. F. Holt, Rowland Sharp, Kate Griffith, Ronan C. Roche, Rebekah G. Newstead, Gabrielle Wyn, and Stuart R. Jenkins. Capacity, capability and cross-border challenges associated with marine eradication programmes in Europe: the attempted eradication of an invasive non-native ascidian, Didemnum vexillum in Wales, United Kingdom. Marine Policy, 48(??):51–58, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000906.

Shamsuzzaman:2018:ALF

[SI18] Md. Mostafa Shamsuzzaman and Mohammad Mahmudul Islam. Analysing the legal framework of marine living resources management in Bangladesh: Towards achieving Sustainable Development Goal 14. Marine Policy, 87 (??):255–262, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302749.

Singleton:2019:ESW

[Sin19] Benedict E. Singleton. The evolution of the super-whale. Complexity and simplicity in environmental communication. Marine Policy, 99(??):170–172, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730845X.

Suuronen:2010:IRS

[SJ10] Petri Suuronen and Pekka Jounela. Introducing runsize driven fisheries management for the coastal fishery of Atlantic salmon: Preseason forecasts for policy makers. Marine Policy, 34(3):679–689, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900195X.

Sjostedt:2014:DER

[SJ14] Martin Sjöstedt and Sverker C. Jagers. Democracy and the environment revisited: the case of African fisheries. Marine Policy, 43(??):143–148, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001218.

Stoll:2015:UBS

[SJ15] Joshua S. Stoll and Teresa R. Johnson. Under the banner of sustainability: the politics and prose of an emerging US federal seafood certification. Marine Policy, 51 (??):415–422, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002607.

Smith:2017:MSP

[SJ17] Glen Smith and Svein Jentoft. Marine spatial planning in Scotland. Levelling the playing field? Marine Policy, 84(??):33–41, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730091X.

Smith:2019:MPF

[SJ19a] Hance D. Smith and Azmath Jaleel. Marine policy: the first four decades. Marine Policy, 108(??):Article 103652, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305822.

Stafford:2019:VOP

[SJ19b] Richard Stafford and Peter J. S. Jones. Viewpoint — ocean plastic pollution: a convenient but distracting truth? Marine Policy, 103(??):187–191, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830681X.

Stafford:2019:WSS

[SJ19c] Richard Stafford and Peter J. S. Jones. We should not separate out environmental issues, but the current approach to plastic pollution can be a distraction from meaningful action. A response to Avery-Gomm et al. Marine Policy, 107(??):Article 103585, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304178.

Senko:2017:LIB

[SJP17] Jesse Senko, Lekelia D. Jenkins, and S. Hoyt Peckham. At loggerheads over international bycatch: Initial effects of a unilaterally imposed bycatch reduction policy. Marine Policy, 76(??):200–209, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630327X.

Suuronen:2010:FRM

[SJT10] Petri Suuronen, Pekka Jounela, and Vesa Tschernij. Fishermen responses on marine protected areas in the Baltic cod fishery. Marine Policy, 34(2):237–243, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000992.

Sitar:2016:BOB

[SJW^{+}16] A. Sitar, May-Collado L. J, A. J. Wright, E. Peters-Burton, L. Rockwood, and E. C. M. Parsons. Boat operators in Bocas del Toro, Panama display low levels of compliance with national whale-watching regulations. Marine Policy, 68(??):221–228, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300951.

Shamshak:2015:CCL

[SK15] Gina Louise Shamshak and Jonathan R. King. From cannery to culinary luxury: the evolution of the global geoduck market. Marine Policy, 55(??):81–89, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000251.

Seung:2018:DCI

[SK18] Chang K. Seung and Do-Hoon Kim. Developing confidence intervals for economic impacts: a multi-regional analysis of a recreational fishery in Korea. Marine Policy, 94(??):20–27, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300125.

Samy-Kamal:2015:DVF

[SKFL15] Mohamed Samy-Kamal, Aitor Forcada, and José Luis Sánchez Lizaso. Daily variation of fishing effort and ex-vessel prices in a western Mediterranean multi-species fishery: Implications for sustainable management. Marine Policy, 61(??):187–195, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002341.

Sorensen:2016:UGM

[SKL16] Thomas Kirk Sørensen and Lotte Kindt-Larsen. Uncovering governance mechanisms surrounding harbour porpoise conservation in the Danish Skagerrak Sea. Marine Policy, 71(??):318–324, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000245.

Smith:2019:FSO

[SKS^{+}19] Sarah L. Smith, Rachel Karasik, Aristoteles Stavrinaky, Hirotsugu Uchida, and Merrick Burden. Fishery Socioeconomic Outcomes Tool: a rapid assessment tool for evaluating socioeconomic performance of fisheries management. Marine Policy, 105(??):20–29, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305815.

Sumaila:2010:SHS

[SKT^{+}10] Ussif Rashid Sumaila, Ahmed Khan, Louise Teh, Reg Watson, Peter Tyedmers, and Daniel Pauly. Subsidies to high seas bottom trawl fleets and the sustainability of deepsea demersal fish stocks. Marine Policy, 34(3):495–497, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001663.

Silva:2015:EFD

[SL15a] Monalisa R. O. Silva and Priscila F. M. Lopes. Each fisherman is different: Taking the environmental perception of small-scale fishermen into account to manage marine protected areas. Marine Policy, 51(??):347–355, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002528.

Stewart:2015:ENZ

[SL15b] James Stewart and Jonathan Leaver. Efficiency of the New Zealand annual catch entitlement market. Marine Policy, 55(??):11–22, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000160.

Stewart:2016:DVA

[SL16] James Stewart and Jonathan Leaver. Deemed value arbitrage in the New Zealand annual catch entitlement market. Marine Policy, 66(??):1–7, April 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000051.

Stoll:2019:EAD

[SLBC19] Joshua S. Stoll, Heather M. Leslie, Melissa L. Britsch, and Caitlin M. Cleaver. Evaluating aquaculture as a diversification strategy for Maine’s commercial fishing sector in the face of change. Marine Policy, 107(??):Article 103583, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18309229.

Stuer-Lauridsen:2018:SRA

[SLDHS18] Frank Stuer-Lauridsen, Guillaume Drillet, Flemming Thorbjørn Hansen, and Justine Saunders. Same Risk Area: an area-based approach for the management of bio-invasion risks from ships’ ballast water. Marine Policy, 97(??): 147–155, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308424.

Sala:2018:ARP

[SLGC^{+}18] Enric Sala, Jane Lubchenco, Kirsten Grorud-Colvert, Catherine Novelli, Callum Roberts, and U. Rashid Sumaila. Assessing real progress towards effective ocean protection. Marine Policy, 91(??):11–13, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307686.

Sanchirico:2013:CVM

[SLH^{+}13] James N. Sanchirico, Daniel K. Lew, Alan C. Haynie, David M. Kling, and David F. Layton. Conservation values in marine ecosystem-based management. Marine Policy, 38(??):523–530, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001753.

Samhouri:2011:UES

[SLJ^{+}11] Jameal F. Samhouri, Phillip S. Levin, C. Andrew James, Jessi Kershner, and Greg Williams. Using existing scientific capacity to set targets for ecosystem-based management: a Puget Sound case study. Marine Policy, 35(4):508–518, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10002150.

Sumaila:2016:GFS

[SLL^{+}16] U. Rashid Sumaila, Vicky Lam, Frédéric Le Manach, Wilf Swartz, and Daniel Pauly. Global fisheries subsidies: an updated estimate. Marine Policy, 69(??):189–193, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000026.

Stelzenmuller:2013:PTS

[SLS^{+}13] Vanessa Stelzenmüller, Janette Lee, Andy South, Jo Foden, and Stuart I. Rogers. Practical tools to support marine spatial planning: a review and some prototype tools. Marine Policy, 38(??):214–227, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001364.

Sala:2018:TSE

[SLS18] Antonello Sala, Alessandro Lucchetti, and Paolo Sartor. Technical solutions for European small-scale driftnets. Marine Policy, 94(??):247–255, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301854.

Sangiuliano:2017:SNS

[SM17a] Stephen Sangiuliano and Stanley Mastrantonis. From Scotland to New Scotland: Constructing a sectoral marine plan for tidal energy for Nova Scotia. Marine Policy, 84(??):1–11, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17300970.

Stewart:2017:FQA

[SM17b] James Stewart and Shane Moriarity. Fishing without quota: the ACE fisher in New Zealand’s inshore fishery. Marine Policy, 86(??):198–205, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304530.

Simmons:2018:SMS

[SM18a] Peter Simmons and Mehmet Ibrahim Mehmet. Shark management strategy policy considerations: Community preferences, reasoning and speculations. Marine Policy, 96 (??):111–119, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302082.

Smith:2018:DTC

[SM18b] Jenny Smith and Anna Metaxas. A decision tree that can address connectivity in the design of Marine Protected Area Networks (MPAn). Marine Policy, 88(??): 269–278, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307558.

Smythe:2018:LLM

[SM18c] Tiffany C. Smythe and Jennifer McCann. Lessons learned in marine governance: Case studies of marine spatial planning practice in the U.S. Marine Policy, 94(??):227–237, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308990.

Sanchez:2013:RFA

[SMBL13] F. Sanchez, G. Morandeau, N. Bru, and M. Lissardy. A restricted fishing area as a tool for fisheries management: Example of the Capbreton canyon, southern Bay of Biscay. Marine Policy, 42(??):180–189, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000523.

Salomon:2014:MPT

[SMD14] Markus Salomon, Till Markus, and Miriam Dross. Masterstroke or paper tiger — The reform of the EU’s Common Fisheries Policy. Marine Policy, 47(??):76–84, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000347.

Sharmina:2017:GES

[SMGL17] Maria Sharmina, Christophe McGlade, Paul Gilbert, and Alice Larkin. Global energy scenarios and their implications for future shipped trade. Marine Policy, 84 (??):12–21, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17302713.

Steenbergen:2017:ERL

[SMH17] Dirk J. Steenbergen, Cliff Marlessy, and Elisabeth Holle. Effects of rapid livelihood transitions: Examining local co-developed change following a seaweed farming boom. Marine Policy, 82(??):216–223, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301744.

Smith:2013:RDM

[Smi13] Hance D. Smith. The regional development and management of fisheries: the UK case. Marine Policy, 37 (??):11–19, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001327.

Smith:2017:AAA

[Smi17] Leland Holbrook Smith. To accede or not to accede: an analysis of the current US position related to the United Nations law of the sea. Marine Policy, 83(??):184–193, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730283X.

Smith:2018:GGR

[Smi18a] Glen Smith. Good governance and the role of the public in Scotland’s marine spatial planning system. Marine Policy, 94(??):1–9, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306966.

Smith:2018:IDS

[Smi18b] Hance D. Smith. Introduction to Deep Sea Mining Activities In The Pacific Region. Marine Policy, 95(??):

301, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18305542.

Strickland-Munro:2016:MSP

[SMKBM16] Jennifer Strickland-Munro, Halina Kobryn, Greg Brown, and Susan A. Moore. Marine spatial planning for the future: Using Public Participation GIS (PPGIS) to inform the human dimension for large marine parks. Marine Policy, 73(??):15–26, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303931.

Stratoudakis:2016:EAF

[SMQ^{+}16] Yorgos Stratoudakis, Catarina Sofia Mateus, Bernardo Ruivo Quintella, Carlos Antunes, and Pedro Raposo de Almeida. Exploited anadromous fish in Portugal: Suggested direction for conservation and management. Marine Policy, 73 (??):92–99, November 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302998.

Silva:2015:DRF

[SMR^{+}15] Cristina Silva, Hugo Mendes, Mafalda Rangel, Laura Wise, Karim Erzini, Maria de Fátima Borges, Marta Ballesteros, Jose Luis Santiago, Aida Campos, Jonas Vidarsson, and Kåre N. Nielsen. Development of a responsive fisheries management system for the Portuguese crustacean bottom trawl fishery: Lessons learnt. Marine Policy, 52(??):19–25, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002759.

Stephenson:2018:SET

[SMS^{+}18] Fabrice Stephenson, Aileen C. Mill, Catherine L. Scott, Gavin B. Stewart, Matthew J. Grainger, Nicholas V. C. Polunin, and Clare Fitzsimmons. Socio-economic, technological and environmental drivers of spatio-temporal changes in fishing pressure. Marine Policy, 88(??): 189–203, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305778.

Soomai:2013:CEI

[SMW13] Suzuette S. Soomai, Bertrum H. MacDonald, and Peter G. Wells. Communicating environmental information to the stakeholders in coastal and marine policy-making: Case studies from Nova Scotia and the Gulf of Maine/Bay of Fundy region. Marine Policy, 40(??):176–186, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000262.

Silvano:2017:GPF

[SNA^{+}17] Renato A. M. Silvano, Vinícius Nora, Tainá B. Andreoli, Priscila F. M. Lopes, and Alpina Begossi. The ‘ghost of past fishing’: Small-scale fisheries and conservation of threatened groupers in subtropical islands. Marine Policy, 75(??):125–132, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306261.

Serrao-Neumann:2016:MGA

[SNDB^{+}16] Silvia Serrao-Neumann, Julie L. Davidson, Claudia L. Baldwin, Aysin Dedekorkut-Howes, Joanna C. Ellison, Neil J. Holbrook, Michael Howes, Christine Jacobson, and Edward A. Morgan. Marine governance to avoid tipping points: Can we adapt the adaptability envelope? Marine Policy, 65(??):56–67, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003760.

Sharaan:2017:QBA

[SNIN17] Mahmoud Sharaan, Abdelazim Negm, Moheb Iskander, and Kazuo Nadaoka. Questionnaire-based assessment of Mediterranean fishing ports, Nile Delta, Egypt. Marine Policy, 81(??):98–108, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305498.

Soma:2018:SPF

[SNP^{+}18] K. Soma, J. R. Nielsen, N. Papadopoulou, H. Polet, M. Zengin, C. J. Smith, O. R. Eigaard, A. Sala, S. Bonanomi, S. W. K. van den Burg, G. J. Piet, El Buisman, and A. Gümüş. Stakeholder perceptions in fisheries management — Sectors with benthic impacts. Marine Policy, 92(??):73–85, June 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307066.

Stevens:2018:RAP

[SNTL18] Julien R. Stevens, Richard W. Newton, Michael Tlusty, and David C. Little. The rise of aquaculture by-products: Increasing food production, value, and sustainability through strategic utilisation. Marine Policy, 90(??):115–124, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17305328.

Sampaio:2013:BEL

[SO13] Flavia Duarte Ferraz Sampaio and Antonio Ostrensky. Brazilian environmental legislation as tool to conserve marine ornamental fish. Marine Policy, 42(??):280– 285, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000857.

Salpin:2018:MSR

[SOBS18] Charlotte Salpin, Vita Onwuasoanya, Marie Bourrel, and Alison Swaddling. Marine scientific research in Pacific Small Island Developing States. Marine Policy, 95(??): 363–371, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16304730.

Santos:2015:CIS

[SOCA15] Catarina Frazão Santos, Michael Orbach, Helena Calado, and Francisco Andrade. Challenges in implementing sustainable marine spatial planning: the new Portuguese legal framework case. Marine Policy, 61(??):196–206, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002298.

Soliman:2014:UIT

[Sol14] Adam Soliman. Using individual transferable quotas (ITQs) to achieve social policy objectives: a proposed intervention. Marine Policy, 45(??):76–81, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002807.

Song:2015:HDF

[Son15] Andrew M. Song. Human dignity: a fundamental guiding value for a human rights approach to fisheries? Marine Policy, 61(??):164–170, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002328.

Stange:2012:MOC

[SOÖ12] Kari Stange, Per Olsson, and Henrik Österblom. Managing organizational change in an international scientific network: a study of ICES reform processes. Marine Policy, 36(3):681–688, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001631.

Soomai:2017:SPI

[Soo17] Suzuette S. Soomai. The science-policy interface in fisheries management: Insights about the influence of organizational structure and culture on information pathways. Marine Policy, 81(??):53–63, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308363.

Stockhausen:2012:DMW

[SOS12] B. Stockhausen, R. A. Officer, and R. Scott. Discard mitigation — what we can learn from waste minimization practices in other natural resources? Marine Policy, 36(1):90–95, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000704.

Sakurai:2016:FAR

[SOUN16] Ryo Sakurai, Takahiro Ota, Takuro Uehara, and Ken’ichi Nakagami. Factors affecting residents’ behavioral intentions for coastal conservation: Case study at Shizugawa Bay, Miyagi, Japan. Marine Policy, 67(??):1–9, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16000117.

Spalding:2019:GPM

[SP19a] Mark Spalding and Cara L. Parrett. Global patterns in mangrove recreation and tourism. Marine Policy, 110(??):Article 103540, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306602.

Symes:2019:STB

[SP19b] David Symes and Jeremy Phillipson. ’A sea of troubles’ (2): Brexit and the UK seafood supply chain. Marine Policy, 102(??):5–9, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304573.

Schakner:2019:CAP

[SPB19] Zachary Schakner, Canon Purdy, and Daniel T. Blumstein. Contrasting attitudes and perceptions of California sea lions by recreational anglers and the media. Marine Policy, 109(??):Article 103710, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309138.

Salayo:2012:MDL

[SPGP12] Nerissa D. Salayo, Maripaz L. Perez, Len R. Garces, and Michael D. Pido. Mariculture development and livelihood diversification in the Philippines. Marine Policy, 36(4):867–881, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001990.

Silos:2012:TGM

[SPMW12] J. M. Silos, F. Piniella, J. Monedero, and J. Walliser. Trends in the global market for crews: a case study. Marine Policy, 36(4):845–858, July 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11002004.

Shabtay:2018:UEM

[SPO^{+}18] Ateret Shabtay, Michelle E. Portman, Eyal Ofir, Yohay Carmel, and Gideon Gal. Using ecological modelling in marine spatial planning to enhance ecosystem-based management. Marine Policy, 95(??):14–23, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300812.

Spence:2018:UPP

[SPP18] Elspeth Spence, Nick Pidgeon, and Paul Pearson. UK public perceptions of Ocean Acidification — The importance of place and environmental identity. Marine Policy, 97(??):287–293, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301921.

Skirtun:2019:TOS

[SPRH19] Maggie Skirtun, Graham M. Pilling, Chris Reid, and John Hampton. Trade-offs for the southern longline fishery in achieving a candidate South Pacific albacore target reference point. Marine Policy, 100(??):66–75, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303671.

Speir:2014:PLF

[SPS14] Cameron Speir, Caroline Pomeroy, and Jon G. Sutinen. Port level fishing dynamics: Assessing changes in the distribution of fishing activity over time. Marine Policy, 46(??):171–191, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000268.

Sarah:2015:CAC

[SRA15] Simons Laura Sarah, Döring Ralf, and Temming Axel. Combining area closures with catch regulations in fisheries with spatio-temporal variation: Bio-economic implications for the North Sea saithe fishery. Marine Policy, 51 (??):281–292, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002334.

Squires:2018:FBR

[SRGD18] Dale Squires, Victor Restrepo, Serge Garcia, and Peter Dutton. Fisheries bycatch reduction within the leastcost biodiversity mitigation hierarchy: Conservatory offsets with an application to sea turtles. Marine Policy, 93(??):55–61, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303020.

Suris-Regueiro:2013:MEP

[SRGGVL13] Juan C. Surís-Regueiro, M. Dolores Garza-Gil, and Manuel M. Varela-Lafuente. Marine economy: a proposal for its definition in the European Union. Marine Policy, 42(??):111–124, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000535.

Suris-Regueiro:2014:SEQ

[SRGGVL14] Juan C. Surís-Regueiro, M. Dolores Garza-Gil, and Manuel M. Varela-Lafuente. Socio-economic quantification of fishing in a European urban area: the case of Vigo. Marine Policy, 43(??):347–358, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001474.

Sethi:2014:AFC

[SRK14] Suresh Andrew Sethi, Matthew Reimer, and Gunnar Knapp. Alaskan fishing community revenues and the stabilizing role of fishing portfolios. Marine Policy, 48(??): 134–141, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000992.

Suris-Regueiro:2014:CFD

[SRS14] Juan C. Surís-Regueiro and Jose L. Santiago. Characterization of fisheries dependence in Galicia (Spain). Marine Policy, 47(??):99–109, July 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1400044X.

Shabudin:2017:AHT

[SRSN17] Ahmad Firdaus Ahmad Shabudin, Rashidah Abdul Rahim, Suzyrman Sibly, and Norizan Md Nor. From ad hoc towards the institutionalisation: an assessment of Malaysia’s policy evolution on Antarctica and the Southern Ocean. Marine Policy, 78(??):1–10, April 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630656X.

Scott:2011:SLT

[SS11] Robert D. Scott and David B. Sampson. The sensitivity of long-term yield targets to changes in fishery ageselectivity. Marine Policy, 35(1):79–84, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001454.

Sydnes:2013:NRC

[SS13] Are Kristoffer Sydnes and Maria Sydnes. Norwegian– Russian cooperation on oil-spill response in the Barents Sea. Marine Policy, 39(??):257–264, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002424.

Shucksmith:2015:RBA

[SS15a] Rachel J. Shucksmith and Richard L. Shelmerdine. A risk based approach to non-native species management and biosecurity planning. Marine Policy, 59(??):32–43, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001050.

Standal:2015:GLN

[SS15b] Dag Standal and Signe Annie Sønvisen. Gear liberalization in the Northeast Arctic cod fisheries — Implications for sustainability, efficiency and legitimacy. Marine Policy, 53(??):141–148, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003297.

Standal:2015:SIB

[SS15c] Dag Standal and Signe Annie Sønvisen. Into the scrap iron business: Transaction costs for fleet sustainability in Norway. Marine Policy, 62(??):213–217, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002729.

Simpson:2016:CEF

[SS16] Samantha J. Simpson and David W. Sims. Are critically endangered fish back on the menu? Analysis of U.K. fisheries data suggest post-ban landings of prohibited skates in European waters. Marine Policy, 69(??):42–51, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300926.

Schutz:2019:SMP

[SS19a] Sigrid Eskeland Schütz and Anne-Michelle Slater. From strategic marine planning to project licences — Striking a balance between predictability and adaptability in the management of aquaculture and offshore wind farms. Marine Policy, 110(??):Article 103556, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303215.

Song:2019:SHR

[SS19b] Andrew M. Song and Adam Soliman. Situating human rights in the context of fishing rights — Contributions and contradictions. Marine Policy, 103(??):19–26, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307759.

Sonvisen:2019:CBA

[SS19c] Signe A. Sønvisen and Dag Standal. Catch-based aquaculture in Norway — Institutional challenges in the development of a new marine industry. Marine Policy, 104 (??):118–124, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302252.

Standal:2016:FDW

[SSA16] Dag Standal, Signe Annie Sønvisen, and Frank Asche. Fishing in deep waters: the development of a deep-sea fishing coastal fleet in Norway. Marine Policy, 63(??):1–7, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002687.

Schumacher:2018:MSC

[SSB^{+}18] Johanna Schumacher, Gerald Schernewski, Malgorzata Bielecka, Michael I. Loizides, and Xenia I. Loizidou. Methodologies to support coastal management — a stakeholder preference and planning tool and its application. Marine Policy, 94(??):150–157, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306607.

Stuart-Smith:2015:PTB

[SSBL^{+}15] Rick D. Stuart-Smith, Amanda E. Bates, Jonathan S. Lefcheck, J. Emmett Duffy, Susan C. Baker, Russell J. Thomson, Jemina F. Stuart-Smith, Nicole A. Hill, Stuart J. Kininmonth, Laura Airoldi, Mikel A. Becerro, Stuart J. Campbell, Terence P. Dawson, Sergio A. Navarrete, German Soler, Elisabeth M. A. Strain, Trevor J. Willis, and Graham J. Edgar. The potential of traitbased approaches to contribute to marine conservation. Marine Policy, 51(??):148–150, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001766.

Sigurethardottir:2015:HCD

[SSC^{+}15] Sigrídur Sigurdardóttir, Elísabet Kemp Stefánsdóttir, Harriet Condie, Sveinn Margeirsson, Thomas L. Catchpole, Jose M. Bellido, Søren Qvist Eliasen, Raquel Goñi, Niels Madsen, Andreas Palialexis, Sebastian S. Uhlmann, Vassiliki Vassilopoulou, Jordan Feekings, and Marie-Joëlle Rochet. How can discards in European fisheries be mitigated? Strengths, weaknesses, opportunities and threats of potential mitigation methods. Marine Policy, 51(??):366–374, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002516.

Sangha:2019:SWE

[SSCC19] Kamaljit K. Sangha, Natalie Stoeckl, Neville Crossman, and Robert Costanza. A state-wide economic assessment of coastal and marine ecosystem services to inform sustainable development policies in the Northern Territory, Australia. Marine Policy, 107(??):Article 103595, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309837.

Sheaves:2016:POC

[SSD^{+}16] Marcus Sheaves, Ilva Sporne, Catherine M. Dichmont, Rodrigo Bustamante, Pat Dale, Roy Deng, Leo X. C. Dutra, Ingrid van Putten, Marie Savina-Rolland, and Anne Swinbourne. Principles for operationalizing climate change adaptation strategies to support the resilience of estuarine and coastal ecosystems: an Australian perspective. Marine Policy, 68(??):229–240, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300975.

Shaffril:2017:ATC

[SSD17a] Hayrol Azril Mohamed Shaffril, Asnarulkhadi Abu Samah, and Jeffrey Lawrence D’Silva. Adapting towards climate change impacts: Strategies for small-scale fishermen in Malaysia. Marine Policy, 81(??):196–201, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308466.

Shaffril:2017:CCS

[SSD17b] Hayrol Azril Mohamed Shaffril, Asnarulkhadi Abu Samah, and Jeffrey Lawrence D’Silva. Climate change: Social adaptation strategies for fishermen. Marine Policy, 81 (??):256–261, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301427.

Stilwell:2010:SDC

[SSFL10] Jonathan Stilwell, Alassane Samba, Pierre Failler, and Francis Laloë. Sustainable development consequences of European Union participation in Senegal’s Marine Fishery. Marine Policy, 34(3):616–623, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001857.

Stevenson:2012:EAC

[SSG12] Charlotte Stevenson, Sarah Abramson Sikich, and Mark Gold. Engaging Los Angeles County subsistence anglers in the California marine protected area planning process. Marine Policy, 36(2):559–563, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001369.

Stevens:2014:MBB

[SSG^{+}14] T. F. Stevens, E. V. Sheehan, S. C. Gall, S. C. Fowell, and M. J. Attrill. Monitoring benthic biodiversity restoration in Lyme Bay marine protected area: Design, sampling and analysis. Marine Policy, 45(??):310–317, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002030.

Svanberg:2019:AMR

[SSH^{+}19] Martin Svanberg, Vendela Santén, Axel Hörteborn, Henrik Holm, and Christian Finnsgård. AIS in maritime research. Marine Policy, 106(??):Article 103520, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309667.

Stuart-Smith:2017:PTN

[SSJ17] Sarah J. Stuart-Smith and Paul D. Jepson. Persistent threats need persistent counteraction: Responding to PCB pollution in marine mammals. Marine Policy, 84 (??):69–75, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630848X.

Siders:2016:DOM

[SSL16] Anne Siders, Rose Stanley, and Kate M. Lewis. A dynamic ocean management proposal for the Bering Strait region. Marine Policy, 74(??):177–185, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305371.

Shelmerdine:2014:IDF

[SSLR14] Richard L. Shelmerdine, Daniel Stone, Beth Leslie, and Martin Robinson. Implications of defining fisheries closed areas based on predicted habitats in Shetland: a proactive and precautionary approach. Marine Policy, 43(??): 184–199, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001267.

Spalding:2015:NEM

[SSM15] Ana K. Spalding, Daniel O. Suman, and Maria Eugenia Mellado. Navigating the evolution of marine policy in Panama: Current policies and community responses in the Pearl Islands and Bocas del Toro Archipelagos of Panama. Marine Policy, 62(??):161–168, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002717.

Sowman:2014:FEP

[SSRS14] Merle Sowman, Jackie Sunde, Serge Raemaekers, and Oliver Schultz. Fishing for equality: Policy for poverty alleviation for South Africa’s small-scale fisheries. Marine Policy, 46(??):31–42, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13002881.

Song:2017:TRF

[SSS^{+}17] Andrew M. Song, Joeri Scholtens, Johny Stephen, Maarten Bavinck, and Ratana Chuenpagdee. Transboundary research in fisheries. Marine Policy, 76(??):8–18, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304924.

Swartz:2017:SMB

[SSSO17] Wilf Swartz, Laurenne Schiller, U. Rashid Sumaila, and Yoshitaka Ota. Searching for market-based sustainability pathways: Challenges and opportunities for seafood certification programs in Japan. Marine Policy, 76(??): 185–191, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302573.

Sunde:2018:VDS

[ŠST^{+}18] Charlotte Šunde, Jim Sinner, Marc Tadaki, Janet Stephenson, Bruce Glavovic, Shaun Awatere, Annabelle Giorgetti, Nick Lewis, Aneika Young, and Kai Chan. Valuation as destruction? The social effects of valuation processes in contested marine spaces. Marine Policy, 97(??):170–178, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301222.

Spranz:2018:MMM

[SSV18] Roger Spranz, Achim Schlüter, and Björn Vollan. Morals, money or the master: the adoption of eco-friendly reusable bags. Marine Policy, 96(??):270–277, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304852.

Satria:2018:CTS

[SSW^{+}18] Fayakun Satria, Lilis Sadiyah, Agustinus Anung Widodo, Chris Wilcox, Jessica H. Ford, and Britta Denise Hardesty. Characterizing transhipment at-sea activities by longline and purse seine fisheries in response to recent policy changes in Indonesia. Marine Policy, 95(??):8–13, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17308023.

Swartz:2010:SST

[SSWP10] Wilf Swartz, U. Rashid Sumaila, Reg Watson, and D. Pauly. Sourcing seafood for the three major markets: the EU, Japan and the USA. Marine Policy, 34 (6):1366–1373, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1000117X.

Stevenson:2013:EMP

[ST13] Todd C. Stevenson and Brian N. Tissot. Evaluating marine protected areas for managing marine resource conflict in Hawaii. Marine Policy, 39(??):215–223, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002321.

Shea:2017:BBP

[ST17] Kwok Ho Shea and Allen Wai Lun To. From boat to bowl: Patterns and dynamics of shark fin trade in Hong Kong — implications for monitoring and management. Marine Policy, 81(??):330–339, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302543.

Sumaila:2016:OFT

[STB16] U. Rashid Sumaila, Alice Tipping, and Christophe Bellmann. Oceans, Fisheries and the Trade System. Marine Policy, 69(??):171–172, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300471.

Schofield:2011:CWA

[STG11] Clive H. Schofield and Ian Townsend-Gault. Choppy waters ahead in “a sea of peace cooperation and friendship”?: Slow progress towards the application of maritime joint development to the East China Sea. Marine Policy, 35 (1):25–33, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10001387.

Salas:2019:TMB

[STIC19] Silvia Salas, Edgar Torres-Irineo, and Eva Coronado. Towards a métier-based assessment and management approach for mixed fisheries in Southeastern Mexico. Marine Policy, 103(??):148–159, May 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302999.

Santos:2012:EMS

[STJ^{+}12] Catarina Frazão Santos, Zara Gonçalves Teixeira, João Janeiro, Renata Sousa Gonçalves, Rhema Bjorkland, and Michael Orbach. The European Marine Strategy: Contribution and challenges from a Portuguese perspective. Marine Policy, 36(5):963–968, September 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000188.

Said:2016:BTF

[STM16] Alicia Said, Joseph Tzanopoulos, and Douglas MacMillan. Bluefin tuna fishery policy in Malta: the plight of artisanal fishermen caught in the capitalist net. Marine Policy, 73(??):27–34, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302299.

Stolsvik:2019:DFC

[Stø19] Gunnar Stølsvik. The development of the fisheries crime concept and processes to address it in the international arena. Marine Policy, 105(??):123–128, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309576.

Standal:2011:HCS

[SU11] Dag Standal and Ingrid Bouwer Utne. The hard choices of sustainability. Marine Policy, 35(4):519–527, July 2011.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000029.

Sundstrom:2012:CRC

[Sun12] Aksel Sundström. Corruption and regulatory compliance: Experimental findings from South African small-scale fisheries. Marine Policy, 36(6):1255–1264, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000577.

Squires:2013:TCF

[SV13] Dale Squires and Niels Vestergaard. Technical change in fisheries. Marine Policy, 42(??):286–292, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300095X.

Squires:2015:PGC

[SV15] Dale Squires and Niels Vestergaard. Productivity growth, catchability, stock assessments, and optimum renewable resource use. Marine Policy, 62(??):309–317, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002079.

Silber:2012:RIM

[SVA^{+}12] Gregory K. Silber, Angelia S. M. Vanderlaan, Ana Tejedor Arceredillo, Lindy Johnson, Christopher T. Taggart, Moira W. Brown, Shannon Bettridge, and Ricardo Sagarminaga. The role of the International Maritime Organization in reducing vessel threat to whales: Process, options, action and effectiveness. Marine Policy, 36(6):1221–1233, November 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000528.

Soma:2018:SIF

[SvdBH^{+}18] Katrine Soma, Sander W. K. van den Burg, Ellen W. J. Hoefnagel, Marian Stuiver, and C. Martijn van der Heide. Social innovation — a future pathway for Blue growth? Marine Policy, 87(??):363–370, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305870.

Shephard:2015:MPT

[SvHdB^{+}15] Samuel Shephard, Ralf van Hal, Ingeborg de Boois, Silvana N. R. Birchenough, Jo Foden, Joey O’Connor, Steve C. V. Geelhoed, Gert Van Hoey, Francisco Marco-Rius, David G. Reid, and Matthias Schaber. Making progress towards integration of existing sampling activities to establish Joint Monitoring Programmes in support of the MSFD. Marine Policy, 59(??):105–111, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001633.

Selden:2016:ESE

[SVL^{+}16] Rebecca L. Selden, Sarah R. Valencia, Ashley E. Larsen, Jorge Cornejo-Donoso, and Amanda A. Wasserman. Evaluating seafood eco-labeling as a mechanism to reduce collateral impacts of fisheries in an ecosystem-based fisheries management context. Marine Policy, 64(??):102–115, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003334.

Sabau:2015:UUG

[SvZdJ15] Gabriela Sabau and Michael van Zyll de Jong. From unjust uneconomic growth to sustainable fisheries in Newfoundland: the true costs of closing the inshore fishery for groundfish. Marine Policy, 61(??):376–389, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14003133.

Sinan:2010:WBF

[SW10] Hussain Sinan and David Whitmarsh. Wealth-based fisheries management and resource rent capture: an application to the Maldives marine fisheries. Marine Policy, 34(3):389–394, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001250.

Stanford:2014:ILF

[SWB^{+}14] Richard J. Stanford, Budy Wiryawan, Dietriech G. Bengen, Rudi Febriamansyah, and John Haluan. Improving livelihoods in fishing communities of West Sumatra: More than just boats and machines. Marine Policy, 45(??):16–25, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002686.

Smallhorn-West:2018:TRM

[SWG18] Patrick Smallhorn-West and Hugh Govan. Towards reducing misrepresentation of national achievements in marine protected area targets. Marine Policy, 97(??):127– 129, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303646.

Soomai:2011:MSP

[SWM11] Suzuette S. Soomai, Peter G. Wells, and Bertrum H. Mac-Donald. Multi-stakeholder perspectives on the use and influence of “grey” scientific information in fisheries management. Marine Policy, 35(1):50–62, January 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001417.

Small:2013:JDI

[SWP13] Cleo Small, Susan M. Waugh, and Richard A. Phillips. The justification, design and implementation of Ecological Risk Assessments of the effects of fishing on seabirds. Marine Policy, 37(??):192–199, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000887.

Srinivasan:2012:GFL

[SWS12] U. Thara Srinivasan, Reg Watson, and U. Rashid Sumaila. Global fisheries losses at the exclusive economic zone level, 1950 to present. Marine Policy, 36(2):544–549, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001515.

Su:2018:EDM

[SY18] Meng Su and Ying Yang. Evolution of district marine policies in China: the case of Shandong Province. Marine Policy, 89(??):124–131, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308436.

Sybersma:2015:RSL

[Syb15] Stacie Sybersma. Review of shark legislation in Canada as a conservation tool. Marine Policy, 61(??):121–126, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002092.

Tingley:2010:RIP

[TÁB^{+}10] Diana Tingley, Jóhann Ásmundsson, Edward Borodzicz, Alexis Conides, Ben Drakeford, Ingi Rúnar Edvardsson, Dennis Holm, Kostas Kapiris, Sakari Kuikka, and Bogi Mortensen. Risk identification and perception in the fisheries sector: Comparisons between the Faroes, Greece, Iceland and UK. Marine Policy, 34(6):1249–1260, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000989.

Tupper:2015:EME

[TAGP15] Mark Tupper, Furqan Asif, Len R. Garces, and Michael D. Pido. Evaluating the management effectiveness of marine protected areas at seven selected sites in the Philippines. Marine Policy, 56(??):33–42, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000433.

Tam:2015:TEC

[Tam15] Chui-Ling Tam. Timing exclusion and communicating time: a spatial analysis of participation failure in an Indonesian MPA. Marine Policy, 54(??):122–129, April 2015.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000032.

Tang:2017:CEE

[Tan17] Jianye Tang. China’s engagement in the establishment of marine protected areas in the Southern Ocean: From reactive to active. Marine Policy, 75(??):68–74, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630567X.

Talloni-Alvarez:2019:CCI

[TÁSLC19] Nicolás E. Talloni-Álvarez, U. Rashid Sumaila, Philippe Le Billon, and William W. L. Cheung. Climate change impact on Canada’s Pacific marine ecosystem: the current state of knowledge. Marine Policy, 104(??):163–176, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306547.

Tissot:2010:HUO

[TBB^{+}10] Brian N. Tissot, Barbara A. Best, Eric H. Borneman, Andrew W. Bruckner, Cara H. Cooper, Heather D’Agnes, Timothy P. Fitzgerald, Amanda Leland, Susan Lieberman, Amy Mathews Amos, Rashid Sumaila, Teresa M. Telecky, Frazer McGilvray, Brian J. Plankis, Andrew L. Rhyne, Glynnis G. Roberts, Benjamin Starkhouse, and Todd C. Stevenson. How U.S. ocean policy and market power can reform the coral reef wildlife trade. Marine Policy, 34(6):1385–1388, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001089.

Tiller:2012:NAE

[TBB12] Rachel Tiller, Tove Brekken, and Jennifer Bailey. Norwegian aquaculture expansion and Integrated Coastal Zone Management (ICZM): Simmering conflicts and competing claims. Marine Policy, 36(5):1086–1095, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000401.

Tezzo:2018:MFT

[TBJC18] Xavier Tezzo, Ben Belton, Gareth Johnstone, and Martin Callow. Myanmar’s fisheries in transition: Current status and opportunities for policy reform. Marine Policy, 97(??):91–100, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304469.

Taylor:2013:SMP

[TBKH13] Elizabeth Taylor, Mark Baine, Annette Killmer, and Marion Howard. Seaflower marine protected area: Governance for sustainable development. Marine Policy, 41 (??):57–64, September 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002643.

Tengberg:2013:LLI

[TC13] Anna Tengberg and Annadel S. Cabanban. Lessons learned from investing in marine and coastal management initiatives in the East Asian Seas. Marine Policy, 38(??):355–364, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12001522.

Techera:2015:OID

[TC15a] Erika J. Techera and John Chandler. Offshore installations, decommissioning and artificial reefs: Do current legal frameworks best serve the marine environment? Marine Policy, 59(??):53–60, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001037.

Thunberg:2015:MFF

[TC15b] Eric M. Thunberg and Steven J. Correia. Measures of fishing fleet diversity in the New England groundfish fishery. Marine Policy, 58(??):6–14, August 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000871.

Tran:2019:FSD

[TCC^{+}19] Nhuong Tran, Long Chu, Chin Yee Chan, Sven Genschick, Michael John Phillips, and Alexander Shula Kefi. Fish supply and demand for food security in Sub-Saharan Africa: an analysis of the Zambian fish sector. Marine Policy, 99(??):343–350, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303798.

Trujillo:2016:CRU

[TCCV16] Juan C. Trujillo, Bladimir Carrillo, Carlos A. Charris, and Raul A. Velilla. Coral reefs under threat in a Caribbean marine protected area: Assessing divers’ willingness to pay toward conservation. Marine Policy, 68(??):146–154, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16300835.

Thyresson:2013:TVC

[TCN^{+}13] M. Thyresson, B. Crona, M. Nyström, M. de la Torre-Castro, and N. Jiddawi. Tracing value chains to understand effects of trade on coral reef fish in Zanzibar, Tanzania. Marine Policy, 38(??):246–256, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200139X.

Tam:2018:GFI

[TCS^{+}18] J. Tam, K. M. A. Chan, T. Satterfield, G. G. Singh, and S. Gelcich. Gone fishing? Intergenerational cultural shifts can undermine common property co-managed fisheries. Marine Policy, 90(??):1–5, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306152.

Torres:2016:EFA

[TdCdSR16] Paulo Torres, Regina Tristão da Cunha, and Armindo dos Santos Rodrigues. The elasmobranch fisheries of the Azores. Marine Policy, 73(??):108–118, November 2016.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304870.

Thebaud:2014:BEE

[TDI^{+}14] Olivier Thébaud, Luc Doyen, James Innes, Michel Lample, Claire Macher, Stephanie Mahévas, Christian Mullon, Benjamin Planque, Martin Quaas, Tony Smith, and Youen Vermard. Building ecological–economic models and scenarios of marine resource systems: Workshop report. Marine Policy, 43(??):382–386, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001243.

Tiller:2019:OFT

[TDMN19] Rachel Tiller, Elizabeth De Santo, Elizabeth Mendenhall, and Elizabeth Nyman. The once and future treaty: Towards a new regime for biodiversity in areas beyond national jurisdiction. Marine Policy, 99(??):239–242, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18307048.

Turner:2016:ITM

[TE16] Jonathan Turner and Stephen Essex. Integrated terrestrial and marine planning in England’s coastal inter-tidal zone: Assessing the operational effectiveness of the Coastal Concordat. Marine Policy, 72(??):166–175, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304584.

Techera:2019:BRP

[Tec19] Erika Techera. Book review: Petroleum Resource Management: How governments manage their offshore petroleum resources, John A.P. Chandler (Ed.). Edward Elgar, Cheltenham, UK (2018), 379 pp. ISBN: 978-1-78643-220-9 (hardback). Marine Policy, 105(??):10–11, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302301.

Thogersen:2015:EGI

[TEF^{+}15] Thomas Thøgersen, Ole Ritzau Eigaard, Mike Fitzpatrick, Simon Mardle, Jesper Levring Andersen, and Gunnar Haraldsson. Economic gains from introducing international ITQs — The case of the mackerel and herring fisheries in the Northeast Atlantic. Marine Policy, 59(??):85–93, September 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001062.

Teijo:2018:ALP

[Tei18] Carlos Teijo. An analysis of the legal problems arising from the application of exclusivity clauses in the European Union’s dormant fisheries agreements. Marine Policy, 95(??):174–179, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830040X.

Tolvanen:2019:SKS

[TEVN19] Harri Tolvanen, Anne Erkkilä-Välimäki, and Tua Nylén. From silent knowledge to spatial information — Mapping blue growth scenarios for maritime spatial planning. Marine Policy, 107(??):Article 103598, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830856X.

Taiarui:2019:SEB

[TFBS19] Marguerite Taiarui, Simon Foale, Tamatoa Bambridge, and Marcus Sheaves. Is stock enhancement the best option to manage fisheries? A case study from Taiarapu (French Polynesia). Marine Policy, 104(??):1–11, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308497.

Trouillet:2011:KCM

[TGBL11] B. Trouillet, T. Guineberteau, M. Bernardon, and S. Le Roux. Key challenges for maritime governance in West Africa: Fishery-based lessons from Guinea and Mauritania. Marine Policy, 35(2):155–162, March 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001545.

Trouillet:2011:PSF

[TGdCR11] Brice Trouillet, Thierry Guineberteau, Mathilde de Cacqueray, and Julien Rochette. Planning the sea: the French experience. Contribution to marine spatial planning perspectives. Marine Policy, 35(3):324–334, May 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001806.

Tweddle:2018:SPK

[TGS18] Jacqueline F. Tweddle, Matthew Gubbins, and Beth E. Scott. Should phytoplankton be a key consideration for marine management? Marine Policy, 97(??):1–9, November 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300854.

Taglioni:2019:SEA

[TGT^{+}19] François Taglioni, Sébastien Guiltat, Magali Teurlai, Mathieu Delsaut, and Denis Payet. A spatial and environmental analysis of shark attacks on Reunion Island (1980– 2017). Marine Policy, 101(??):51–62, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302951.

Thornton:2015:NNC

[TH15] Thomas F. Thornton and Jamie Hebert. Neoliberal and neo-communal herring fisheries in Southeast Alaska: Reframing sustainability in marine ecosystems. Marine Policy, 61(??):366–375, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003169.

Torres:2017:CRE

[TH17] Cati Torres and Nick Hanley. Communicating research on the economic valuation of coastal and marine ecosystem services. Marine Policy, 75(??):99–107, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304912.

Tolley:2015:TSP

[THA15] B. Tolley and M. Hall-Arber. Tipping the scale away from privatization and toward community-based fisheries: Policy and market alternatives in New England. Marine Policy, 61(??):401–409, November 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003042.

Tilley:2018:PVM

[THE^{+}18] Alexander Tilley, Pilar Herrón, Silvana Espinosa, Juliana López Angarita, and Stephen Box. Predicting vulnerability to management changes in data-limited, smallscale fisheries. Marine Policy, 94(??):39–45, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302835.

Trenkel:2015:IMP

[THF^{+}15] Verena M. Trenkel, Niels T. Hintzen, Keith D. Farnsworth, Christian Olesen, David Reid, Anna Rindorf, Samuel Shephard, and Mark Dickey-Collas. Identifying marine pelagic ecosystem management objectives and indicators. Marine Policy, 55(??):23–32, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000044.

Thiel:2013:ESR

[Thi13a] Andreas Thiel. Erratum to “Scalar reorganisation of marine governance in Europe? The implementation of the marine strategy framework directive in Spain, Portugal and Germany” [Marine Policy 39 (2013) 322– 332]. Marine Policy, 42(??):177, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300047X.

Thiel:2013:SRM

[Thi13b] Andreas Thiel. Scalar reorganisation of marine governance in Europe? The implementation of the marine strategy framework directive in Spain, Portugal and Germany. Marine Policy, 39(??):322–332, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002217.

Tsurita:2018:MPA

[THK^{+}18] Izumi Tsurita, Juri Hori, Toshihiro Kunieda, Masakazu Hori, and Mitsutaku Makino. Marine protected areas, Satoumi, and territorial use rights for fisheries: a case study from Hinase, Japan. Marine Policy, 91(??):41–48, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17305079.

Thorvaldsen:2015:EFN

[THM15] Trine Thorvaldsen, Ingunn M. Holmen, and Helene K. Moe. The escape of fish from Norwegian fish farms: Causes, risks and the influence of organisational aspects. Marine Policy, 55(??):33–38, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000196.

Thurstan:2012:MRN

[THNR12] Ruth H. Thurstan, Julie P. Hawkins, Luiza Neves, and Callum M. Roberts. Are marine reserves and non-consumptive activities compatible? A global analysis of marine reserve regulations. Marine Policy, 36(5):1096–1104, September 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12000504.

Thorvaldsen:2013:ICS

[Tho13] Trine Thorvaldsen. The importance of common sense: How Norwegian coastal fishermen deal with occupational risk. Marine Policy, 42(??):85–90, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300050X.

Thompson:2017:KCS

[THPJ17] Katie R. Thompson, William D. Heyman, S. Hoyt Peckham, and Lekelia D. Jenkins. Key characteristics of successful fisheries learning exchanges. Marine Policy, 77 (??):205–213, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300069.

Tiller:2015:WSN

[THRS15] Rachel Gjelsvik Tiller, Lillian Hansen, Russell Richards, and Hillevi Strand. Work segmentation in the Norwegian salmon industry: the application of segmented labor market theory to work migrants on the island community of Frøya, Norway. Marine Policy, 51(??):563–572, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002632.

Thur:2010:UFS

[Thu10] Steven M. Thur. User fees as sustainable financing mechanisms for marine protected areas: an application to the Bonaire National Marine Park. Marine Policy, 34 (1):63–69, January 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000657.

Tisdell:2013:FQA

[TI13] J. G. Tisdell and Md Sayed Iftekhar. Fisheries quota allocation: Laboratory experiments on simultaneous and combinatorial auctions. Marine Policy, 38(??):228–234, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001376.

Tokunaga:2019:AOU

[TII^{+}19] Kanae Tokunaga, Gakushi Ishimura, Shigehide Iwata, Keita Abe, Kazuhiko Otsuka, Kristin Kleisner, and Rod Fujita. Alternative outcomes under different fisheries management policies: a bioeconomic analysis of Japanese fisheries. Marine Policy, 108(??):Article 103646, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309333.

Thebaud:2014:MED

[TINL^{+}14] Olivier Thébaud, James Innes, Ana Norman-López, Stephanie Slade, Darren Cameron, Toni Cannard, Sharon Tickell, John Kung, Brigid Kerrigan, Lew Williams, and L. Richard Little. Micro-economic drivers of profitability in an ITQ-managed fishery: an analysis of the Queensland Coral Reef Fin-Fish Fishery. Marine Policy, 43(??):200–207, January 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001279.

Tipping:2016:BPF

[Tip16] Alice Tipping. Building on progress in fisheries subsidies disciplines. Marine Policy, 69(??):202–208, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003784.

Tirrell:2017:SIN

[Tir17] Andrew Tirrell. Sociocultural institutions in Norwegian fisheries management. Marine Policy, 77(??):37–43, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304341.

Trimble:2013:AFU

[TJ13] Micaela Trimble and Derek Johnson. Artisanal fishing as an undesirable way of life? The implications for governance of fishers’ wellbeing aspirations in coastal Uruguay and southeastern Brazil. Marine Policy, 37(??):37–44, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000632.

Tatar:2018:GKC

[TJ18] Bradley Tatar and Changkuk Jung. Getting to know the consumer: Toward mitigation of illegal whale meat consumption in South Korea. Marine Policy, 89(??): 116–123, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305729.

Todorovic:2019:PEO

[TJJAM19] Sara Todorović, Maria José Juan-Jordá, Haritz Arrizabalaga, and Hilario Murua. Pelagic ecoregions: Operationalizing an ecosystem approach to fisheries management in the Atlantic Ocean. Marine Policy, 109(??):Article 103700, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300806.

Techera:2011:FGR

[TK11] Erika J. Techera and Natalie Klein. Fragmented governance: Reconciling legal strategies for shark conservation and management. Marine Policy, 35(1):73–78, January 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001430.

Techera:2013:RLS

[TK13] Erika J. Techera and Natalie Klein. The role of law in shark-based eco-tourism: Lessons from Australia. Marine Policy, 39(??):21–28, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002011.

Tammi:2014:SMM

[TK14] Ilpo Tammi and Risto Kalliola. Spatial MCDA in marine planning: Experiences from the Mediterranean and Baltic Seas. Marine Policy, 48(??):73–83, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000876.

Thoya:2019:TED

[TKAO^{+}19] Pascal Thoya, Boaz Kaunda-Arara, Johnstone Omukoto, Cosmas Munga, Edward Kimani, and Arthur Omondi Tuda. Trawling effort distribution and influence of vessel monitoring system (VMS) in Malindi–Ungwana Bay: Implications for resource management and marine spatial planning in Kenya. Marine Policy, 109(??):Article 103677, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19303793.

Thorvaldsen:2018:WWR

[TKM^{+}18] Trine Thorvaldsen, Kim O. Kaustell, Tiina E. A. Mattila, Annbjørg á Høvdanum, Jørgen Møller Christiansen, Sarah Hovmand, Hilmar Snorrason, Kristinn Tomasson, and Ingunn M. Holmen. What works? Results of a Nordic survey on fishers’ perceptions of safety measures. Marine Policy, 95(??):95–101, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301489.

Tuda:2019:EPA

[TKN19] Arthur Omondi Tuda, Salit Kark, and Alice Newton. Exploring the prospects for adaptive governance in marine transboundary conservation in East Africa. Marine Policy, 104(??):75–84, June 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305372.

Thorsell:2016:EOP

[TL16] Devon Emily Thorsell and Thomas M. Leschine. An evaluation of oil pollution prevention strategies in the Arctic: a comparison of Canadian and U.S. approaches. Marine Policy, 72(??):255–262, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302366.

Tilley:2016:RPV

[TLA16] Alexander Tilley and Juliana López-Angarita. RE-TRACTED: Predicting vulnerability to management changes in data-limited, small-scale fisheries. Marine Policy, 72(??):211–218, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X16300938.

Tilley:2018:RNP

[TLA18] Alexander Tilley and Juliana López-Angarita. Retraction notice to “Predicting vulnerability to management changes in data-limited, small-scale fisheries”. Marine Policy, 95(??):401, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305785.

Terry:2019:IMC

[TLB19] Alan Terry, Kerry Lewis, and Blaise Bullimore. The impact of the Marine and Coastal Access Act (2009) on Welsh inshore fisheries and marine management. Marine Policy, 99(??):359–368, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830438X.

TuckerWilliams:2018:PDN

[TLH18] Ellary TuckerWilliams, Christopher A. Lepczyk, and Christopher T. Hawkins. Perspectives on developing a non-commercial saltwater fishing license program in Hawai’i. Marine Policy, 94(??):174–179, August 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305365.

Tallis:2010:MFE

[TLR^{+}10] Heather Tallis, Phillip S. Levin, Mary Ruckelshaus, Sarah E. Lester, Karen L. McLeod, David L. Fluharty, and Benjamin S. Halpern. The many faces of ecosystembased management: Making the process work today in real places. Marine Policy, 34(2):340–348, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001110.

Tallis:2012:NMM

[TLR^{+}12] Heather Tallis, Sarah E. Lester, Mary Ruckelshaus, Mark Plummer, Karen McLeod, Anne Guerry, Sandy Andelman, Margaret R. Caldwell, Marc Conte, Stephen Copps, David Fox, Rod Fujita, Steven D. Gaines, Guy Gelfenbaum, Barry Gold, Peter Kareiva, Choong ki Kim, Kai Lee, Michael Papenfus, Scott Redman, Brian Silliman, Lisa Wainger, and Crow White. New metrics for managing and sustaining the ocean’s bounty. Marine Policy, 36(1):303–306, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000728.

Toonen:2013:PSF

[TM13] Hilde M. Toonen and Arthur P. J. Mol. Putting sustainable fisheries on the map? Establishing no-take zones for North Sea plaice fisheries through MSC certification. Marine Policy, 37(??):294–304, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001005.

Tinlin-Mackenzie:2019:SMS

[TMDSF19] Ashleigh Tinlin-Mackenzie, Jane Delany, Catherine L. Scott, and Clare Fitzsimmons. Spatially modelling the suitability, sensitivity, and vulnerability of data poor fisheries with GIS: a case study of the Northumberland lugworm fishery. Marine Policy, 109(??):Article 103707, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301733.

Thomas:2015:WDF

[TMG15] Alyssa S. Thomas, Taciano L. Milfont, and Michael C. Gavin. What determines fishers’ knowledge of and attitudes towards regulations? A case study from the Marlborough Sounds, New Zealand. Marine Policy, 51(??): 547–554, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14002577.

Thiault:2018:STM

[TMG^{+}18] Lauric Thiault, Paul Marshall, Stefan Gelcich, Antoine Collin, Frédérique Chlous, and Joachim Claudet. Space and time matter in social-ecological vulnerability assessments. Marine Policy, 88(??):213–221, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304785.

Tokotch:2012:SPN

[TMHJ12] Britni N. Tokotch, Christopher F. Meindl, Armando Hoare, and Michael E. Jepson. Stakeholder perceptions of the northern Gulf of Mexico grouper and tilefish individual fishing quota program. Marine Policy, 36(1):34–41, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11000650.

Torres:2015:WUN

[TMKK^{+}15] H. Torres, F. Muller-Karger, D. Keys, H. Thornton, M. Luther, and K. Alsharif. Whither the U.S. National Ocean Policy Implementation Plan? Marine Policy, 53 (??):198–212, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003145.

Tosic:2019:PMS

[TML^{+}19] Marko Tosic, Flávio Martins, Serguei Lonin, Alfredo Izquierdo, and Juan Darío Restrepo. A practical method for setting coastal water quality targets: Harmonization of land-based discharge limits with marine ecosystem thresholds. Marine Policy, 108(??):Article 103641, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307425.

Tiller:2014:SFA

[TMR^{+}14] Rachel Gjelsvik Tiller, Jarle Mork, Russell Richards, Lionel Eisenhauer, Yajie Liu, Jens-Fredrik Nakken, and Åshild.L Borgersen. Something fishy: Assessing stakeholder resilience to increasing jellyfish (Periphylla periphylla) in Trondheimsfjord, Norway. Marine Policy, 46(??):72–83, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002893.

Tiller:2015:HCE

[TN15] Rachel Tiller and Elizabeth Nyman. Having the cake and eating it too: To manage or own the Svalbard Fisheries Protection Zone. Marine Policy, 60(??):141–148, October 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001700.

Tseng:2010:TCU

[TO10] Huan-Sheng Tseng and Ching-Hsiewn Ou. Taiwan and China: a unique fisheries relationship. Marine Policy, 34(6):1156–1162, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000679.

Todd:2012:MRE

[Tod12]  Paul Todd.  Marine renewable energy and public

rights. Marine Policy, 36(3):667–672, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001710.

Ting:2012:MDW

[TOL12] Kuo-Huan Ting, Ching-Hsiewn Ou, and Wen-Hong Liu. The management of the distant water tuna fishery in Taiwan. Marine Policy, 36(6):1234–1241, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200053X.

Torkington:2016:NZQ

[Tor16] Barry Torkington. New Zealand’s quota management system — incoherent and conflicted. Marine Policy, 63 (??):180–183, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15000664.

Tosun:2011:PPM

[Tos11] Jale Tosun. Political parties and marine pollution policy: Exploring the case of Germany. Marine Policy, 35(4):536–541, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000169.

Tedesco:2016:CBS

[TPKL16] Letizia Tedesco, Chiara Piroddi, Maria Kämäri, and Christopher Lynam. Capabilities of Baltic Sea models to assess environmental status for marine biodiversity. Marine Policy, 70(??):1–12, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301932.

Turner:2015:MIF

[TPS15] Rachel A. Turner, Nicholas V. C. Polunin, and Selina M. Stead. Mapping inshore fisheries: Comparing observed and perceived distributions of pot fishing activity in Northumberland. Marine Policy, 51(??):173–181, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002218.

Tiller:2018:OFE

[TR18] Rachel Tiller and Russell Richards. Ocean futures: Exploring stakeholders’ perceptions of adaptive capacity to changing marine environments in Northern Norway. Marine Policy, 95(??):227–238, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307030.

Tsounis:2013:MHS

[TRBS13] Georgios Tsounis, Sergio Rossi, Lorenzo Bramanti, and Giovanni Santangelo. Management hurdles for sustainable harvesting of Corallium rubrum. Marine Policy, 39 (??):361–364, May 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12002515.

Tran:2017:IAF

[TRC^{+}17] Nhuong Tran, U.-Primo Rodriguez, Chin Yee Chan, Michael John Phillips, Chadag Vishnumurthy Mohan, Patrik John Gustav Henriksson, Sonny Koeshendrajana, Sharon Suri, and Stephen Hall. Indonesian aquaculture futures: an analysis of fish supply and demand in Indonesia to 2030 and role of aquaculture using the AsiaFish model. Marine Policy, 79(??):25–32, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307205.

Thiebot:2019:APE

[TRCR^{+}19] Jean-Baptiste Thiebot, Yan Ropert-Coudert, Thierry Raclot, Timothée Poupart, Akiko Kato, and Akinori Takahashi. Adélie penguins’ extensive seasonal migration supports dynamic Marine Protected Area planning in Antarctica. Marine Policy, 109(??):Article 103692, November 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303471.

Tallman:2019:GAS

[TRM19] Ross F. Tallman, Marie-Julie Roux, and Zoya A. Martin. Governance and assessment of small-scale data-limited Arctic Charr fisheries using productivity-susceptibility analysis coupled with life history invariant models. Marine Policy, 101(??):187–197, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307467.

Thorvaldsen:2014:MCN

[TS14] Trine Thorvaldsen and Signe Annie Sønvisen. Multilingual crews on Norwegian fishing vessels: Implications for communication and safety on board. Marine Policy, 43 (??):301–306, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001395.

Tai:2019:EPF

[TSH^{+}19] Travis C. Tai, Nadja S. Steiner, Carie Hoover, William W. L. Cheung, and U. Rashid Sumaila. Evaluating present and future potential of Arctic fisheries in Canada. Marine Policy, 108(??):Article 103637, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300995.

Teneva:2018:SPA

[TSK18] Lida T. Teneva, Eva Schemmel, and John N. Kittinger. State of the plate: Assessing present and future contribution of fisheries and aquaculture to Hawai‘i’s food security. Marine Policy, 94(??):28–38, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308801.

Turner:2019:HCF

[TSW19] Rachel A. Turner, Nigel C. Sainsbury, and Benedict W. Wheeler. The health of commercial fishers in England and Wales: Analysis of the 2011 census. Marine Policy, 106(??):Article 103548, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308510.

Turner:2019:DSH

[TTFC19] Phillip J. Turner, Andrew D. Thaler, Amy Freitag, and Patrick Colman Collins. Deep-sea hydrothermal vent ecosystem principles: Identification of ecosystem processes, services and communication of value. Marine Policy, 101(??):118–124, March 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302811.

Teh:2018:EAM

[TTJ18] Louise S. L. Teh, Lydia C. L. Teh, and Gavin Jolis. An economic approach to marine megafauna conservation in the coral triangle: Marine turtles in Sabah, Malaysia. Marine Policy, 89(??):1–10, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304967.

Taufa:2018:APR

[TTM18a] Salome V. Taufa, Mele Tupou, and Siola’a Malimali. An analysis of property rights in the Special Management Area (SMA) in Tonga. Marine Policy, 95(??):267–272, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17309016.

Tebet:2018:UOP

[TTM18b] Guilherme Tebet, Micaela Trimble, and Rodrigo Pereira Medeiros. Using Ostrom’s principles to assess institutional dynamics of conservation: Lessons from a marine protected area in Brazil. Marine Policy, 88(??):174–181, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17305134.

Teh:2012:TSP

[TTP12] Lydia C. L. Teh, Louise S. L. Teh, and Tony J. Pitcher. A tool for site prioritisation of marine protected areas under data poor conditions. Marine Policy, 36(6): 1290–1300, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000711.

Teh:2014:TPS

[TTS14] Louise S. L. Teh, Lydia C. L. Teh, and U. Rashid Sumaila. Time preference of small-scale fishers in open access and traditionally managed reef fisheries. Marine Policy, 44(??):222–231, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001942.

Topfer:2014:CPC

[TTUR14] Klaus Töpfer, Laurence Tubiana, Sebastian Unger, and Julien Rochette. Charting pragmatic courses for global ocean governance. Marine Policy, 49(??):85–86, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1300287X.

Turris:2010:RPA

[Tur10] Bruce R. Turris. A rejoinder to E. Pinkerton et al., The elephant in the room: the hidden costs of leasing individual transferable fishing quotas. Marine Policy, 34(3):431–436, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001444.

Taljaard:2013:HSE

[TvN13] Susan Taljaard and Lara van Niekerk. How supportive are existing national legal regimes for multiuse marine spatial planning? — the South African case. Marine Policy, 38(??):72–79, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001194.

Thunberg:2015:MCM

[TWA^{+}15] Eric Thunberg, John Walden, Juan Agar, Ron Felthoven, Abigail Harley, Stephen Kasperski, Jean Lee, Todd Lee, Aaron Mamula, Jessica Stephen, and Andy Strelcheck. Measuring changes in multi-factor productivity in U.S. catch share fisheries. Marine Policy, 62(??):294–301, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001414.

Thompson:2017:GOF

[TWJ^{+}17] Katie R. Thompson, Amy Hudson Weaver, Lekelia Danielle Jenkins, Nathalie Zenny, Nicolas J. Pilcher, S. Hoyt Peckham, and William D. Heyman. Guidelines for organizing a fisheries learning exchange. Marine Policy, 77(??):214–218, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16303608.

Tynkkynen:2017:BSE

[Tyn17] Nina Tynkkynen. The Baltic Sea environment and the European Union: Analysis of governance barriers. Marine Policy, 81(??):124–131, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306674.

Tynkkynen:2018:RIT

[Tyn18] Nina Tynkkynen. The “Russian Issue” in transnational governance of the Baltic Sea environment: Analysis of drivers and constraints of Russia’s participation. Marine Policy, 98(??):220–226, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306821.

Tolentino-Zondervan:2018:FVF

[TZBBL18] Frazen Tolentino-Zondervan, Paul Berentsen, Simon R. Bush, and Alfons Oude Lansink. FAD vs. free school: Effort allocation by Marine Stewardship Council compliant Filipino tuna purse seiners in the PNA. Marine Policy, 90(??):137–145, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303950.

Urquhart:2013:CSF

[UA13] Julie Urquhart and Tim Acott. Constructing ‘The Stade’: Fishers’ and non-fishers’ identity and place attachment in Hastings, south-east England. Marine Policy, 37 (??):45–54, January 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000656.

Urquhart:2013:ISC

[UAZ13] Julie Urquhart, Tim Acott, and Minghua Zhao. Introduction: Social and cultural impacts of marine fisheries. Marine Policy, 37(??):1–2, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000681.

Uc-Espadas:2018:FAG

[UEMRG^{+}18] Martha Uc-Espadas, Dolores Molina-Rosales, Francisco D. Gurri, Juan Carlos Pérez-Jiménez, and Verónica Vázquez-García. Fishing activities by gender and reproductive stage in Isla Arena, Campeche, Mexico. Marine Policy, 89(??):34–39, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17306127.

Uhrin:2016:TCD

[Uhr16] Amy V. Uhrin. Tropical cyclones, derelict traps, and the future of the Florida Keys commercial spiny lobster fishery. Marine Policy, 69(??):84–91, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301762.

Uhre:2017:WOF

[UL17] Andreas Nordang Uhre and Einar Leknes. When the oil and fishing industries live side by side. Marine Policy, 75 (??):108–115, January 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1630207X.

Utizi:2018:IAE

[UNS^{+}18] Kizzi Utizi, Emilio Notti, Antonello Sala, Alessandro Buzzi, Ilaria Rodella, Umberto Simeoni, and Corinne Corbau. Impact assessment of EMFF measures on Good Environmental Status (GES) as defined by Italy. Marine Policy, 88(??):248–260, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308059.

VanDover:2018:SRI

[VAHG^{+}18] C. L. Van Dover, S. Arnaud-Haond, M. Gianni, S. Helmreich, J. A. Huber, A. L. Jaeckel, A. Metaxas, L. H. Pendleton, S. Petersen, E. Ramirez-Llodra, P. E. Steinberg, V. Tunnicliffe, and H. Yamamoto. Scientific rationale and international obligations for protection of active hydrothermal vent ecosystems from deep-sea mining. Marine Policy, 90(??):20–28, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306061.

Valman:2013:ISC

[Val13] Matilda Valman. Institutional stability and change in the Baltic Sea: 30 years of issues, crises and solutions. Marine Policy, 38(??):54–64, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001170.

Varkey:2010:IUU

[VAP^{+}10] Divya A. Varkey, Cameron H. Ainsworth, Tony J. Pitcher, Yohanis Goram, and Rashid Sumaila. Illegal, unreported and unregulated fisheries catch in Raja Ampat Regency, Eastern Indonesia. Marine Policy, 34(2):228–236, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000980.

VanDover:2014:ERD

[VAP^{+}14] C. L. Van Dover, J. Aronson, L. Pendleton, S. Smith, S. Arnaud-Haond, D. Moreno-Mateos, E. Barbier, D. Billett, K. Bowers, R. Danovaro, A. Edwards, S. Kellert, T. Morato, E. Pollard, A. Rogers, and R. Warner. Ecological restoration in the deep sea: Desiderata. Marine Policy, 44(??):98–106, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001486.

Varjopuro:2011:CES

[Var11] Riku Varjopuro. Co-existence of seals and fisheries? Adaptation of a coastal fishery for recovery of the Baltic grey seal. Marine Policy, 35(4):450–456, July 2011. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10002101.

vanAs:2019:CPS

[vAS19] Hendrik J. van As and Philippus J. Snijman. Challenges and possible solutions concerning the inspection/ investigation dichotomy in the context of transnational organised fisheries crime: a South African perspective. Marine Policy, 105(??):140–150, July 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309436.

Vatsov:2017:EFA

[Vat17] Mihail Vatsov. The EU’s failed attempt to innovate with Regulation 1026/2012. Marine Policy, 84(??):300–305, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17302075.

Vaughan:2017:FED

[Vau17] Duncan Vaughan. Fishing effort displacement and the consequences of implementing Marine Protected Area management — an English perspective. Marine Policy, 84 (??):228–234, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16307588.

Velez:2014:FPF

[VAW14] Mariana Velez, Sara Adlerstein, and Julia Wondolleck. Fishers’ perceptions, facilitating factors and challenges of community-based no-take zones in the Sian Ka’an Biosphere Reserve, Quintana Roo, Mexico. Marine Policy, 45(??):171–181, March 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13002868.

Vernazzani:2017:MCI

[VBHC^{+}17] B. Galletti Vernazzani, P. Burkhardt-Holm, E. Cabrera, M. Iñíguez, F. Luna, E. C. M. Parsons, F. Ritter, J. Rodríguez-Fonseca, M. Sironi, and M. Stachowitsch. Management and conservation at the International Whaling Commission: a dichotomy sandwiched within a shifting baseline. Marine Policy, 83(??):164–171, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302464.

Voyer:2017:CCS

[VBMM17] Michelle Voyer, Kate Barclay, Alistair McIlgorm, and Nicole Mazur. Connections or conflict? A social and economic analysis of the interconnections between the professional fishing industry, recreational fishing and marine tourism in coastal communities in NSW, Australia.

Marine Policy, 76(??):114–121, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306376.

Vince:2017:OGS

[VBSD17] Joanna Vince, Elizabeth Brierley, Simone Stevenson, and Piers Dunstan. Ocean governance in the South Pacific region: Progress and plans for action. Marine Policy, 79(??):40–45, May 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16308089.

vanBets:2016:LNG

[vBvTM16] Linde K. J. van Bets, Jan P. M. van Tatenhove, and Arthur P. J. Mol. Liquefied natural gas production at Hammerfest: a transforming marine community. Marine Policy, 69(??):52–61, July 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300343.

Vaughan:2015:HPC

[VC15] Mehana Blaich Vaughan and Margaret R. Caldwell. Hana Pa’a: Challenges and lessons for early phases of comanagement. Marine Policy, 62(??):51–62, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002067.

Venerus:2017:RMR

[VC17] Leonardo A. Venerus and Paula V. Cedrola. Review of marine recreational fisheries regulations in Argentina. Marine Policy, 81(??):202–210, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307242.

Vilela:2018:IFS

[VCdR^{+}18] R. Vilela, D. Conesa, J. L. del Rio, A. López-Quílez, J. Portela, and J. M. Bellido. Integrating fishing spatial patterns and strategies to improve high seas fisheries management. Marine Policy, 94(??):132–142, August 2018.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301884.

vanDoorn:2016:EAM

[vD16] Erik van Doorn. Environmental aspects of the Mining code: Preserving humankind’s common heritage while opening Pardo’s box? Marine Policy, 70(??):192–197, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300562.

Valman:2016:ORC

[VDB16] Matilda Valman, Andreas Duit, and Thorsten Blenckner. Organizational responsiveness: the case of unfolding crises and problem detection within HELCOM. Marine Policy, 70(??):49–57, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16301889.

vandenBurg:2019:AGP

[vdBAMJT19] S. W. K. van den Burg, José Aguilar-Manjarrez, Jeff Jenness, and Melanie Torrie. Assessment of the geographical potential for co-use of marine space, based on operational boundaries for Blue Growth sectors. Marine Policy, 100(??):43–57, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304615.

vandenBurg:2017:BCM

[vdBKB^{+}17] S. W. K. van den Burg, P. Kamermans, M. Blanch, D. Pletsas, M. Poelman, K. Soma, and G. Dalton. Business case for mussel aquaculture in offshore wind farms in the North Sea. Marine Policy, 85(??):1–7, November 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301707.

vandeGeer:2013:IMB

[vdGMA^{+}13] Casper van de Geer, Morena Mills, Vanessa M. Adams, Robert L. Pressey, and Daryl McPhee. Impacts of the Moreton Bay Marine Park rezoning on commercial fishermen. Marine Policy, 39(??):248–256, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12002400.

vanderMarel:2017:ICG

[vdMCJ17] Eva R. van der Marel, Richard Caddell, and Elise Johansen. Innovating for change in global fisheries governance: an introduction. Marine Policy, 84(??):285–286, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304232.

vonderPorten:2016:RMH

[vdPLMS16] Suzanne von der Porten, Dana Lepofsky, Deborah McGregor, and Jennifer Silver. Recommendations for marine herring policy change in Canada: Aligning with Indigenous legal and inherent rights. Marine Policy, 74(??):68–76, December 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16301464.

vanderValk:2016:FTP

[vdVD16] Olga M. C. van der Valk and Birgit I. De Vos. Family ties, preconceived images and trust: How local community defines market collaboration in the Dutch fish chain. Marine Policy, 71(??):175–183, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301257.

Vazquez:2015:CSC

[VG15] Alfredo Macías Vázquez and Pablo Alonso González. Collective symbolic capital and sustainability: Governing fishing communities in a knowledge economy. Marine Policy, 53(??):21–26, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14003157.

Vespe:2015:DIP

[VGA15] Michele Vespe, Harm Greidanus, and Marlene Alvarez Alvarez. The declining impact of piracy on maritime transport in the Indian Ocean: Statistical analysis of 5-year vessel tracking data. Marine Policy, 59(??):9–15, September 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15001001.

Voyer:2015:PMU

[VGBG15] Michelle Voyer, Natalie Gollan, Kate Barclay, and William Gladstone. ‘It’s part of me’; understanding the values, images and principles of coastal users and their influence on the social acceptability of MPAs. Marine Policy, 52(??):93–102, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002899.

Villarrubia-Gomez:2018:MPP

[VGCF18] Patricia Villarrubia-Gómez, Sarah E. Cornell, and Joan Fabres. Marine plastic pollution as a planetary boundary threat — The drifting piece in the sustainability puzzle. Marine Policy, 96(??):213–220, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305456.

Voyer:2012:MSA

[VGG12] Michelle Voyer, William Gladstone, and Heather Goodall. Methods of social assessment in Marine Protected Area planning: Is public participation enough? Marine Policy, 36(2):432–439, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001370.

Voyer:2015:OSL

[VGG15] Michelle Voyer, William Gladstone, and Heather Goodall. Obtaining a social licence for MPAs — influences on social acceptability. Marine Policy, 51(??):260–266, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002371.

Viana:2017:RBV

[VGL^{+}17] Daniel Viana, Kiya Gornik, Ching-Cheng Lin, Gavin McDonald, Nathaniel S. R. Ng, Christine Quigley, and Matthew Potoski. Recreational boaters value biodiversity: the case of the California Channel Islands National Marine Sanctuary. Marine Policy, 81(??):91–97, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630197X.

vanHoof:2015:FME

[vH15] Luc van Hoof. Fisheries management, the ecosystem approach, regionalisation and the elephants in the room. Marine Policy, 60(??):20–26, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500144X.

Verna:2016:RBW

[VH16] Danielle E. Verna and Bradley P. Harris. Review of ballast water management policy and associated implications for Alaska. Marine Policy, 70(??):13–21, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302044.

vanHees:2019:IIB

[vH19] Sander van Hees. Increased integration between innovative ocean energy and the EU habitats, species and water protection rules through Maritime Spatial Planning. Marine Policy, 100(??):27–42, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303245.

Vieira:2019:PDO

[VHA^{+}19] Mário Vieira, Elsa Henriques, Miguel Amaral, Nuno Arantes-Oliveira, and Luís Reis. Path discussion for offshore wind in Portugal up to 2030. Marine Policy, 100 (??):122–131, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18303932.

vanHelmond:2016:CFB

[vHCT^{+}16] A. T. M. van Helmond, C. Chen, B. K. Trapman, M. Kraan, and J. J. Poos. Changes in fishing behaviour of two fleets under fully documented catch quota management: Same rules, different outcomes. Marine Policy, 67(??):118–129, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000427.

vanHoof:2019:FOT

[vHFJ^{+}19] Luc van Hoof, Gianna Fabi, Vegar Johansen, Josien Steenbergen, Xabier Irigoien, Sarah Smith, Dennis Lisbjerg, and Gerd Kraus. Food from the ocean; towards a research agenda for sustainable use of our oceans’ natural resources. Marine Policy, 105(??):44–51, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17308618.

vanHoof:2014:SYC

[vHHB14] Luc van Hoof, Astrid Hendriksen, and Helen J. Bloomfield. Sometimes you cannot make it on your own; drivers and scenarios for regional cooperation in implementing the EU Marine Strategy Framework Directive. Marine Policy, 50(??):339–346, December 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001110.

vanHoof:2017:TNN

[vHK17] Luc van Hoof and Gerd Kraus. Is there a need for a new governance model for regionalised Fisheries Management? Implications for science and advice. Marine Policy, 84(??):152–155, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301586.

VanPelt:2017:MMC

[VHRM17] T. I. Van Pelt, H. P. Huntington, O. V. Romanenko, and F. J. Mueter. The missing middle: Central Arctic Ocean gaps in fishery research and science coordination. Marine Policy, 85(??):79–86, November 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301367.

Villasante:2010:GAE

[Vil10] Sebastian Villasante. Global assessment of the European Union fishing fleet: an update. Marine Policy, 34(3):663–670, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001936.

Villasante:2012:MBW

[Vil12] Sebastián Villasante. The management of the blue whiting fishery as complex social-ecological system: the Galician case. Marine Policy, 36(6):1301–1308, November 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000206.

Vince:2013:MBP

[Vin13] Joanna Vince. Marine bioregional plans and implementation issues: Australia’s oceans policy process. Marine Policy, 38(??):325–329, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001480.

Virto:2018:PAI

[Vir18] Laura Recuero Virto. A preliminary assessment of the indicators for Sustainable Development Goal (SDG) 14 “Conserve and sustainably use the oceans, seas and marine resources for sustainable development”. Marine Policy, 98(??):47–57, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301131.

Virdin:2019:WAC

[VKA^{+}19] John Virdin, Mimako Kobayashi, Stephen Akester, Tibor Vegh, and Stephen Cunningham. West Africa’s coastal bottom trawl fishery: Initial examination of a trade in fishing services. Marine Policy, 100(??):288–297, February 2019.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830335X.

Visbeck:2014:SBW

[VKGN^{+}14a] Martin Visbeck, Ulrike Kronfeld-Goharani, Barbara Neumann, Wilfried Rickels, Jörn Schmidt, Erik van Doorn, Nele Matz-Lück, Konrad Ott, and Martin F. Quaas. Securing blue wealth: the need for a special sustainable development goal for the ocean and coasts. Marine Policy, 48(??):184–191, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000773.

Visbeck:2014:SDG

[VKGN^{+}14b] Martin Visbeck, Ulrike Kronfeld-Goharani, Barbara Neumann, Wilfried Rickels, Jörn Schmidt, Erik van Doorn, Nele Matz-Lück, and Alexander Proelss. A sustainable development goal for the ocean and coasts: Global ocean challenges benefit from regional initiatives supporting globally coordinated solutions. Marine Policy, 49 (??):87–89, November 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000487.

Voltaire:2017:IOW

[VLKN17] Louinord Voltaire, Maria L. Loureiro, Camilla Knudsen, and Paulo A. L. D. Nunes. The impact of offshore wind farms on beach recreation demand: Policy intake from an economic study on the Catalan coast. Marine Policy, 81(??):116–123, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303189.

Vaissiere:2014:BOO

[VLPC14] Anne-Charlotte Vaissière, Harold Levrel, Sylvain Pioch, and Antoine Carlier. Biodiversity offsets for offshore wind farm projects: the current situation in Europe. Marine Policy, 48(??):172–183, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000955.

vanLeeuwen:2014:IMS

[vLRvH^{+}14] Judith van Leeuwen, Jesper Raakjaer, Luc van Hoof, Jan van Tatenhove, Ronán Long, and Kristen Ounanian. Implementing the Marine Strategy Framework Directive: a policy perspective on regulatory, institutional and stakeholder impediments to effective implementation. Marine Policy, 50(??):325–330, December 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1400075X.

vanLeeuwen:2012:IAI

[vLvHvT12] Judith van Leeuwen, Luc van Hoof, and Jan van Tatenhove. Institutional ambiguity in implementing the European Union Marine Strategy Framework Directive. Marine Policy, 36(3):636–643, May 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001576.

vanLeeuwen:2010:TMG

[vLvT10] Judith van Leeuwen and Jan van Tatenhove. The triangle of marine governance in the environmental governance of Dutch offshore platforms. Marine Policy, 34(3):590–597, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001791.

Vlachopoulou:2018:SAR

[VM18] Eirini Ioanna Vlachopoulou and Darien Danielle Mizuta. Shellfish aquaculture and resilience: Leadership experiences from Kesennuma Bay, Japan. Marine Policy, 92 (??):111–119, June 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307868.

Vanderklift:2019:COM

[VMMB^{+}19] Mathew A. Vanderklift, Raymundo Marcos-Martinez, James R. A. Butler, Michael Coleman, Anissa Lawrence, Heidi Prislan, Andrew D. L. Steven, and Sebastian Thomas. Constraints and opportunities for marketbased finance for the restoration and protection of blue carbon ecosystems. Marine Policy, 107(??):Article 103429, September 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304846.

Vega:2014:EIS

[VMO14] Amaya Vega, Ana Corina Miller, and Cathal O’Donoghue. Economic impacts of seafood production growth targets in Ireland. Marine Policy, 47(??):39–45, July 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000414.

Vianna:2018:SDT

[VMR^{+}18] Gabriel M. S. Vianna, Mark G. Meekan, Abbie A. Rogers, Marit E. Kragt, James M. Alin, and Johanna S. Zimmerhackel. Shark-diving tourism as a financing mechanism for shark conservation strategies in Malaysia. Marine Policy, 94(??):220–226, August 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307807.

vanOsch:2017:EIP

[vOHO^{+}17] Suzanne van Osch, Stephen Hynes, Tim O’Higgins, Nick Hanley, Danny Campbell, and Shirra Freeman. Estimating the Irish public’s willingness to pay for more sustainable salmon produced by integrated multi-trophic aquaculture. Marine Policy, 84(??):220–227, October 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301951.

VanNguyen:2019:IRE

[VPC19] Quang Van Nguyen, Sean Pascoe, and Louisa Coglan. Implications of regional economic conditions on the distribution of technical efficiency: Examples from coastal trawl vessels in Vietnam. Marine Policy, 102(??):51–60, April 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304792.

vanPutten:2010:LQF

[vPG10] Ingrid van Putten and Caleb Gardner. Lease quota fishing in a changing rock lobster industry. Marine Policy, 34(5):859–867, September 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000096.

vanPutten:2017:RFT

[vPJH^{+}17] Ingrid E. van Putten, Sarah Jennings, Alistair J. Hobday, Rodrigo H. Bustamante, Leo X. C. Dutra, Stewart Frusher, Elizabeth A. Fulton, Marcus Haward, Éva Plagányi, Linda Thomas, and Gretta Pecl. Recreational fishing in a time of rapid ocean change. Marine Policy, 76(??):169–177, February 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16306212.

Veiga:2013:TOA

[VPL^{+}13] P. Veiga, C. Pita, L. Leite, J. Ribeiro, R. B. Ditton, J. M. S. Gonçalves, and K. Erzini. From a traditionally open access fishery to modern restrictions: Portuguese anglers’ perceptions about newly implemented recreational fishing regulations. Marine Policy, 40(??):53–63, July 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000080.

vanPutten:2013:BMF

[vPLB^{+}13] Ingrid van Putten, Annie Lalancette, Peter Bayliss, Darren Dennis, Trevor Hutton, Ana Norman-López, Sean Pascoe, Eva Plagányi, and Tim Skewes. A Bayesian model of factors influencing indigenous participation in the Torres Strait tropical rocklobster fishery. Marine Policy, 37(??):96–105, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000620.

Villasante:2016:LLH

[VPP^{+}16] Sebastian Villasante, Cristina Pita, Graham J. Pierce, César Pazos Guimeráns, João Garcia Rodrigues, Manel Antelo, José María Da Rocha, Javier García Cutrín, Lee Hastie, Ussif Rashid Sumaila, and Marta Coll. To land or not to land: How do stakeholders perceive the zero discard policy in European small-scale fisheries? Marine Policy, 71(??):166–174, September 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302780.

Veiga:2016:ELO

[VPR^{+}16] Pedro Veiga, Cristina Pita, Mafalda Rangel, Jorge M. S. Gonçalves, Aida Campos, Paul G. Fernandes, Antonello Sala, Massimo Virgili, Alessandro Lucchetti, Jure Brčić, Sebastian Villasante, Marta A. Ballesteros, Rosa Chapela, José L. Santiago, Sveinn Agnarsson, Ólafur Ögmundarson, and Karim Erzini. The EU landing obligation and European small-scale fisheries: What are the odds for success? Marine Policy, 64(??):64–71, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003279.

Villasante:2012:GSM

[VRA^{+}12] Sebastián Villasante, David Rodríguez, Manel Antelo, Martin Quaas, and Henrik Österblom. The Global Seafood Market Performance Index: a theoretical proposal and potential empirical applications. Marine Policy, 36(1): 142–152, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000790.

Vrancken:2019:SJI

[Vra19] Patrick Vrancken. State jurisdiction to investigate and try fisheries crime at sea. Marine Policy, 105(??):129–139, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309527.

Vazquez-Rowe:2014:UCP

[VRB14] Ian Vázquez-Rowe and Enrico Benetto. The use of a consequential perspective to upgrade the utility of Life Cycle Assessment for fishery managers and policy makers. Marine Policy, 48(??):14–17, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000633.

Vazquez-Rowe:2011:EGD

[VRMF11] Ian Vázquez-Rowe, M.a Teresa Moreira, and Gumersindo Feijoo. Estimating global discards and their potential reduction for the Galician fishing fleet (NW Spain). Marine Policy, 35(2):140–147, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X10001521.

Vazquez-Rowe:2012:EAF

[VRMF12] Ian Vázquez-Rowe, María Teresa Moreira, and Gumersindo Feijoo. Environmental assessment of frozen common octopus (Octopus vulgaris) captured by Spanish fishing vessels in the Mauritanian EEZ. Marine Policy, 36(1): 180–188, January 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11000911.

Vazquez-Rowe:2013:IIS

[VRT13] Ian Vázquez-Rowe and Peter Tyedmers. Identifying the importance of the “skipper effect” within sources of measured inefficiency in fisheries through data envelopment analysis (DEA). Marine Policy, 38(??):387–396, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001571.

Villasante:2010:EET

[VS10] Sebastian Villasante and Ussif Rashid Sumaila. Estimating the effects of technological efficiency on the European fishing fleet. Marine Policy, 34(3):720–722, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900181X.

vanSittert:2015:FEF

[vS15] Lance van Sittert. The fire and the eye: Fishers knowledge, echo-sounding and the invention of the skipper in the St. Helena Bay pelagic fishery ca. 1930– 1960. Marine Policy, 60(??):300–308, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002024.

Vince:2018:PCI

[VS18] Joanna Vince and Peter Stoett. From problem to crisis to interdisciplinary solutions: Plastic marine debris. Marine Policy, 96(??):200–203, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303294.

VanDover:2012:DNC

[VSA^{+}12] C. L. Van Dover, C. R. Smith, J. Ardron, D. Dunn, K. Gjerde, L. Levin, and S. Smith. Designating networks of chemosynthetic ecosystem reserves in the deep sea. Marine Policy, 36(2):378–381, March 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11001175.

Vann-Sander:2016:CCS

[VSCH16] Sarah Vann-Sander, Julian Clifton, and Euan Harvey. Can citizen science work? Perceptions of the role and utility of citizen science in a marine policy and management context. Marine Policy, 72(??):82–93, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304079.

Valmonte-Santos:2016:FSU

[VSRD16] Rowena Valmonte-Santos, Mark W. Rosegrant, and Madan Mohan Dey. Fisheries sector under climate change in the coral triangle countries of Pacific Islands: Current status and policy issues. Marine Policy, 67(??):148–155, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003929.

Vince:2015:AOP

[VSS^{+}15] Joanna Vince, Anthony D. M. Smith, Keith. J. Sainsbury, Ian David Cresswell, David C. Smith, and Marcus Haward. Australia’s Oceans Policy: Past, present and future. Marine Policy, 57(??):1–8, July 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000494.

vanTatenhove:2014:RCE

[vTRvLvH14] Jan van Tatenhove, Jesper Raakjaer, Judith van Leeuwen, and Luc van Hoof. Regional cooperation for European seas: Governance models in support of the implementation of the MSFD. Marine Policy, 50(??):364–372, December 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000700.

Verweij:2010:DCR

[VvD10] M. C. Verweij and W. L. T. van Densen. Differences in causal reasoning about resource dynamics and consequences for the participatory debate on North Sea fisheries. Marine Policy, 34(6):1144–1155, November 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10000667.

Verweij:2010:TBD

[VvDM10] M. C. Verweij, W. L. T. van Densen, and A. J. P. Mol. The tower of Babel: Different perceptions and controversies on change and status of North Sea fish stocks in multi-stakeholder settings. Marine Policy, 34(3):522–533, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001705.

VanNijen:2018:STE

[VVS18] Kris Van Nijen, Steven Van Passel, and Dale Squires. A stochastic techno-economic assessment of seabed mining of polymetallic nodules in the Clarion Clipperton Fracture Zone. Marine Policy, 95(??):133–141, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730903X.

Vrancken:2019:IOT

[VWG19] Patrick Vrancken, Emma Witbooi, and Jan Glazewski. Introduction and overview: Transnational organised fisheries crime. Marine Policy, 105(??):116–122, July 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309448.

Wright:2015:AMB

[WAG^{+}15] Glen Wright, Jeff Ardron, Kristina Gjerde, Duncan Currie, and Julien Rochette. Advancing marine biodiversity protection through regional fisheries management: a review of bottom fisheries closures in areas beyond national jurisdiction. Marine Policy, 61(??):134–148, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001955.

Weitzman:2018:PAE

[WB18] Jenny Weitzman and Megan Bailey. Perceptions of aquaculture ecolabels: a multi-stakeholder approach in Nova Scotia, Canada. Marine Policy, 87(??):12–22, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303810.

Wabnitz:2019:RCW

[WB19] Colette C. C. Wabnitz and Robert Blasiak. The rapidly changing world of ocean finance. Marine Policy, 107(??):Article 103526, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303343.

Wilewska-Bien:2018:RSB

[WBA18] Magda Wilewska-Bien and Stefan Anderberg. Reception of sewage in the Baltic Sea — The port’s role in the sustainable management of ship wastes. Marine Policy, 93(??):207–213, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17308680.

Walker:2015:HDC

[WBB^{+}15] Tony R. Walker, Meagan Bernier, Brenden Blotnicky, Peter G. Golden, Emma Hoffman, Jeffrey Janes, Allison Kader, Rachel Kovacs-Da Costa, Shauna Pettipas, and Sarah Vermeulen. Harbour divestiture in Canada: Implications of changing governance. Marine Policy, 62 (??):1–8, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X15002377.

Wamukota:2014:MIR

[WBC14] A. Wamukota, T. D. Brewer, and B. Crona. Market integration and its relation to income distribution and inequality among fishers and traders: the case of two small-scale Kenyan reef fisheries. Marine Policy, 48(??):93–101, September 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14000852.

Willsteed:2018:SCE

[WBGJ18] Edward A. Willsteed, Silvana N. R. Birchenough, Andrew B. Gill, and Simon Jude. Structuring cumulative effects assessments to support regional and local marine management and planning obligations. Marine Policy, 98 (??):23–32, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18305256.

Waldo:2013:SCH

[WBH^{+}13] Staffan Waldo, Kim Berndt, Cecilia Hammarlund, Martin Lindegren, Anders Nilsson, and Anders Persson. Swedish coastal herring fisheries in the wake of an ITQ system. Marine Policy, 38(??):321–324, March 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001479.

Woods:2015:CQB

[WBH^{+}15] Pamela J. Woods, Caroline Bouchard, Daniel S. Holland, André E. Punt, and Gudrun Marteinsdóttir. Catch-quota balancing mechanisms in the Icelandic multi-species demersal fishery: Are all species equal? Marine Policy, 55(??):1– 10, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000068.

Wood:2013:SFO

[WBSW13] Apanie L. Wood, James R. A. Butler, Marcus Sheaves, and Jacob Wani. Sport fisheries: Opportunities and challenges for diversifying coastal livelihoods in the Pacific. Marine Policy, 42(??):305–314, November 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300081X.

Wang:2016:SLP

[WC16] Xiu Wei Wang and Yen-Chiang Chang. ‘Sea Law Phalanx’ initiative: Conference report. Marine Policy, 72 (??):11–13, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16302834.

Wang:2017:MEL

[WC17] Yu-Shan Wang and Shui-Kai Chang. Is MSC ecolabelling workable in Taiwan? Responses from various sectors of the Taiwanese sergestid shrimp fishery. Marine Policy, 77(??):164–170, March 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630478X.

Williams:2018:WGF

[WCCO18] Chris Williams, Griffin Carpenter, Robert Clark, and Bethan C. O’Leary. Who gets to fish for sea bass? Using social, economic, and environmental criteria to determine access to the English sea bass fishery. Marine Policy, 95(??):199–208, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17307650.

Wood:2017:UMA

[WCK^{+}17] Daniel Wood, Elisa Capuzzo, Damien Kirby, Karen Mooney-McAuley, and Philip Kerrison. UK macroalgae aquaculture: What are the key environmental and licensing considerations? Marine Policy, 83(??):29–39, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16305577.

Wamukota:2012:CMC

[WCM12] A. W. Wamukota, J. E. Cinner, and T. R. McClanahan. Co-management of coral reef fisheries: a critical evaluation of the literature. Marine Policy, 36(2):481–488, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001461.

Wabnitz:2018:ECC

[WCMHO18] Colette C. C. Wabnitz, Andrés M. Cisneros-Montemayor, Quentin Hanich, and Yoshitaka Ota. Ecotourism, climate change and reef fish consumption in Palau: Benefits, trade-offs and adaptation strategies. Marine Policy, 88(??):323–332, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302063.

Wan:2018:IBD

[WCS18] Zheng Wan, Jihong Chen, and Dan Sperling. Institutional barriers to the development of a comprehensive ballastwater management scheme in China: Perspective from a multi-stream policy model. Marine Policy, 91(??):142–149, May 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300599.

Worm:2013:GCE

[WDK^{+}13] Boris Worm, Brendal Davis, Lisa Kettemer, Christine A. Ward-Paige, Demian Chapman, Michael R. Heithaus, Steven T. Kessel, and Samuel H. Gruber. Global catches, exploitation rates, and rebuilding options for sharks. Marine Policy, 40(??):194–204, July 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13000055.

Wear:2016:MBC

[Wea16] Stephanie L. Wear. Missing the boat: Critical threats to coral reefs are neglected at global scale. Marine Policy, 74(??):153–157, December 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630358X.

Webster:2011:IEA

[Web11] D. G. Webster. The irony and the exclusivity of Atlantic bluefin tuna management. Marine Policy, 35(2):249–251, March 2011. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10001442.

Wegge:2011:SSM

[Weg11] Njord Wegge. Small state, maritime great power? Norway’s strategies for influencing the maritime policy of the European Union. Marine Policy, 35(3):335–342, May 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001818.

Wegge:2015:EPA

[Weg15] Njord Wegge. The emerging politics of the Arctic Ocean. Future management of the living marine resources. Marine Policy, 51(??):331–338, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002486.

Walters:2019:MDB

[WEKH19] Carl Walters, Karl English, Josh Korman, and Ray Hilborn. The managed decline of British Columbia’s commercial salmon fishery. Marine Policy, 101(??):25–32, March 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18303361.

Wenzel:2011:PBM

[Wen11] George W. Wenzel. Polar bear management, sport hunting and Inuit subsistence at Clyde River, Nunavut. Marine Policy, 35(4):457–465, July 2011. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X10002071.

Wenzel:2016:OCP

[Wen16] Bertolt Wenzel. Organizing coordination in a public marine research and management advice organization: the case of the Norwegian Institute of Marine Research. Marine Policy, 64(??):159–167, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003619.

Wenzel:2017:OCE

[Wen17] Bertolt Wenzel. Organizing coordination for an ecosystem approach to marine research and management advice: the case of ICES. Marine Policy, 82(??):138–146, August 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16308156.

Westholm:2018:ASL

[Wes18] Aron Westholm. Appropriate scale and level in marine spatial planning — Management perspectives in the Baltic Sea. Marine Policy, 98(??):264–270, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306869.

Wilson:2018:RNS

[WF18] A. Meriwether W. Wilson and Catherine Forsyth. Restoring near-shore marine ecosystems to enhance climate security for island ocean states: Aligning international processes and local practices. Marine Policy, 93(??):284–294, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17304682.

Waugh:2012:ERA

[WFK^{+}12a] Susan M. Waugh, Dominique P. Filippi, David S. Kirby, Edward Abraham, and Nathan Walker. Ecological Risk Assessment for seabird interactions in Western and Central Pacific longline fisheries. Marine Policy, 36(4):933–946, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001874.

Westhead:2012:FGO

[WFK^{+}12b] Maxine C. Westhead, Derek G. Fenton, Tanya A. Koropatnick, Paul A. Macnab, and Hilary B. Moors. Filling the gaps one at a time: the Gully Marine Protected Area in Eastern Canada. A response to Agardy, Notarbartolo di Sciara and Christie. Marine Policy, 36(3):713–715, May 2012. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X11001886.

Walden:2015:PCC

[WFSV15] John Walden, Ben Fissel, Dale Squires, and Niels Vestergaard. Productivity change in commercial fisheries: an introduction to the special issue. Marine Policy, 62(??):289–293, December 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001785.

Whitney:2016:IWA

[WGB^{+}16] Charlotte K. Whitney, Julie Gardner, Natalie C. Ban, Chantal Vis, Soonya Quon, and Suzan Dionne. Imprecise and weakly assessed: Evaluating voluntary measures for management of marine protected areas. Marine Policy, 69(??):92–101, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16300239.

Weijerman:2016:TEB

[WGMG^{+}16] Mariska Weijerman, Cynthia Grace-McCaskey, Shanna L. Grafeld, Dawn M. Kotowicz, Kirsten L. L. Oleson, and Ingrid E. van Putten. Towards an ecosystem-based approach of Guam’s coral reefs: the human dimension. Marine Policy, 63(??):8–17, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002791.

Watson:2018:SPS

[WH18] Maryann S. Watson and Stephanie M. Hewson. Securing protection standards for Canada’s marine protected areas. Marine Policy, 95(??):117–122, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302318.

Wakefield:2018:UAT

[WHB^{+}18] Corey B. Wakefield, Sybrand A. Hesp, Stuart Blight, Brett W. Molony, Stephen J. Newman, and Norman G. Hall. Uncertainty associated with total bycatch estimates for rarely-encountered species varies substantially with observer coverage levels: Informing minimum requirements for statutory logbook validation. Marine Policy, 95(??):273–282, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301994.

Wallner-Hahn:2016:DGU

[WHMG^{+}16] Sieglind Wallner-Hahn, Fia Molander, Gloria Gallardo, Sebastian Villasante, Johan S. Eklöf, Narriman S. Jiddawi, and Maricela de la Torre-Castro. Destructive gear use in a tropical fishery: Institutional factors influencing the willingness-and capacity to change. Marine Policy, 72(??):199–210, October 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304225.

Wang:2016:IMM

[WHP^{+}16] Ying Wang, Jianfeng Hu, Haoran Pan, Shiyu Li, and Pierre Failler. An integrated model for marine fishery management in the Pearl River Estuary: Linking socio-economic systems and ecosystems.

Marine Policy, 64(??):135–147, February 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15003553.

Ward:2018:CPR

[WIC18] Timothy Mark Ward, Alex Ivey, and Jonathan Carroll. Code of practice for reducing accidental mortality of dolphins in purse-seine fisheries. Marine Policy, 87(??): 203–211, January 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17305900.

Wise:2014:LTE

[Wis14] Sarah P. Wise. Learning through experience: Nonimplementation and the challenges of protected area conservation in The Bahamas. Marine Policy, 46(??):111–118, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000128.

Witbooi:2015:RTN

[Wit15a] Emma Witbooi. RETRACTED: Towards a new ‘fisheries crime’ paradigm: Challenges and opportunities with reference to South Africa as an illustrative African example. Marine Policy, 55(??):39–46, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000172.

Witbooi:2015:RNT

[Wit15b] Emma Witbooi. Retraction notice to: “Towards a new ‘fisheries crime’ paradigm: Challenges and opportunities with reference to South Africa as an illustrative African example” [Mar. Policy 55 (2015) 39– 46]. Marine Policy, 55(??):R1, May 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000949.

Witteveen:2019:EIS

[Wit19] Ana Witteveen. Estimating the impact of size dependent pricing on fisher revenues under the EU landing obligation.

Marine Policy, 110(??):Article 103570, December 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305827.

Walden:2014:MFP

[WK14] John B. Walden and Nolan Kitts. Measuring fishery profitability: an index number approach. Marine Policy, 43 (??):321–326, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001449.

Wallmo:2017:UCM

[WK17] Kristy Wallmo and Rosemary Kosaka. Using choice models to inform large marine protected area design. Marine Policy, 83(??):111–117, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307990.

Weir:2019:PPP

[WK19] Stephanie Weir and Sandy Kerr. Property, power and planning: Attitudes to spatial enclosure in Scottish seas. Marine Policy, 108(??):Article 103633, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301319.

Wever:2015:LSD

[WKB15] Lara Wever, Gesche Krause, and Bela H. Buck. Lessons from stakeholder dialogues on marine aquaculture in offshore wind farms: Perceived potentials, constraints and research gaps. Marine Policy, 51(??):251–259, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002310.

Waiti:2018:SMD

[WL18] Divine Waiti and Rebecca Lorrenij. Sustainable management of deep sea mineral activities: a case study of the development of national regulatory frameworks for the Republic of the Marshall Islands. Marine Policy, 95(??):

388–393, September 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17301252.

Williams:2010:RPS

[WLP10] Gregory D. Williams, Phillip S. Levin, and Wayne A. Palsson. Rockfish in Puget Sound: an ecological history of exploitation. Marine Policy, 34(5):1010–1020, September 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X10000515.

Walsh:2019:TTO

[WLT^{+}19a] Conor Walsh, Nicholas-Joseph Lazarou, Michael Traut, James Price, Carlo Raucci, Maria Sharmina, Paolo Agnolucci, Sarah Mander, Paul Gilbert, Kevin Anderson, Alice Larkin, and Tristan Smith. Trade and trade-offs: Shipping in changing climates. Marine Policy, 106(??):Article 103537, August 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306638.

Warner:2019:SSH

[WLT^{+}19b] Kimberly A. Warner, Beth Lowell, Walker Timme, Emily Shaftel, and Robert H. Hanner. Seafood sleuthing: How citizen science contributed to the largest market study of seafood mislabeling in the U.S. and informed policy. Marine Policy, 99(??):304–311, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302689.

Wang:2015:SBI

[WLV15] Yamin Wang, Wei Li, and Koen Van Waerebeek. Strandings, bycatches and injuries of aquatic mammals in China, 2000–2006, as reviewed from official documents: a compelling argument for a nationwide strandings programme. Marine Policy, 51(??):242–250, January 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001900.

Wang:2019:MME

[WLX19] Grace Wang, Kevin X. Li, and Yi Xiao. Measuring marine environmental efficiency of a cruise shipping company considering corporate social responsibility. Marine Policy, 99(??):140–147, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304184.

Walsh:2017:CDT

[WM17] Conor Walsh and Sarah Mander. Contextualising the drivers for trade: Some lessons from historical case studies. Marine Policy, 75(??):290–299, January 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300781.

Wakefield:2018:SCB

[WM18a] Jeffrey R. Wakefield and Kelley Myers. Social cost benefit analysis for deep sea minerals mining. Marine Policy, 95(??):346–355, September 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303694.

Warnken:2018:CII

[WM18b] Jan Warnken and Razieh Mosadeghi. Challenges of implementing integrated coastal zone management into local planning policies, a case study of Queensland, Australia. Marine Policy, 91(??):75–84, May 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306358.

Wakamatsu:2019:EPS

[WM19] Mihoko Wakamatsu and Shunsuke Managi. Examining public support for international agreements on tuna management and conservation. Marine Policy, 100(??): 298–306, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18300976.

Wade:2016:IDU

[WMJF16] H. M. Wade, E. A. Masden, A. C. Jackson, and R. W. Furness. Incorporating data uncertainty when estimating potential vulnerability of Scottish seabirds to marine renewable energy developments. Marine Policy, 70(??):108–113, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1630241X.

Wilson:2015:RMI

[WMK15] A. Meriwether W. Wilson, Robert Mugerauer, and Terrie Klinger. Rethinking marine infrastructure policy and practice: Insights from three large-scale marina developments in Seattle. Marine Policy, 53(??):67–82, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14003212.

Wilcox:2016:UEE

[WML^{+}16] Chris Wilcox, Nicholas J. Mallos, George H. Leonard, Alba Rodriguez, and Britta Denise Hardesty. Using expert elicitation to estimate the impacts of plastic pollution on marine wildlife. Marine Policy, 65(??):107–114, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002985.

Walsh:2017:CLC

[WML17] Conor Walsh, Sarah Mander, and Alice Larkin. Charting a low carbon future for shipping: a UK perspective. Marine Policy, 82(??):32–40, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307813.

Watson:2015:SLM

[WMSB15] G. J. Watson, J. M. Murray, M. Schaefer, and A. Bonner. Successful local marine conservation requires appropriate educational methods and adequate enforcement. Marine Policy, 52(??):59–67, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002784.

Worcester:2018:RMB

[WMVR18] Peter F. Worcester, James H. Miller, and Kathleen J. Vigness-Raposa. Rejoinder to: M. B. Kaplan, S. Solomon, A coming boom in commercial shipping? The potential for rapid growth of noise from commercial ships by 2030 [Mar. Policy 73 (Suppl. C) (2016) S119–S121]. Marine Policy, 98(??):174–175, December 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830112X.

Willis:2018:HSW

[WMWH18] Kathryn Willis, Clémentine Maureaud, Chris Wilcox, and Britta Denise Hardesty. How successful are waste abatement campaigns and government policies at reducing plastic waste into the marine environment? Marine Policy, 96(??):243–249, October 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305171.

Welch:2015:IDA

[WNB^{+}15] David J. Welch, Stephen J. Newman, Rik C. Buckworth, Jennifer R. Ovenden, Damien Broderick, Robert J. G. Lester, Neil A. Gribble, Aaron C. Ballagh, Robbie A. Charters, Jason Stapley, Raewyn Street, Rod N. Garrett, and Gavin A. Begg. Integrating different approaches in the definition of biological stocks: a northern Australian multi-jurisdictional fisheries example using grey mackerel, Scomberomorus semifasciatus. Marine Policy, 55(??):73– 80, May 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000214.

Watson:2017:GST

[WNLS17] Reg A. Watson, R. Nichols, V. W. Y. Lam, and U. R. Sumaila. Global seafood trade flows and developing economies: Insights from linking trade and production. Marine Policy, 82(??):41–49, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301604.

Wright:2016:ELR

[WOdG^{+}16] Glen Wright, Anne Marie O’Hagan, Jiska de Groot, Yannick Leroy, Niko Soininen, Rachael Salcido, Montserrat Abad Castelos, Simon Jude, Julien Rochette, and Sandy Kerr. Establishing a legal research agenda for ocean energy. Marine Policy, 63(??):126–134, January 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500281X.

Wondirad:2019:RPC

[Won19] Amare Wondirad. Retracing the past, comprehending the present and contemplating the future of cruise tourism through a meta-analysis of journal publications. Marine Policy, 108(??):Article 103618, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303045.

Wosu:2019:AIS

[Wos19] Adaoma Wosu. Access and institutions in a smallscale octopus fishery: a gendered perspective. Marine Policy, 108(??):Article 103649, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301150.

Warner:2012:CCC

[WP12] Tammy E. Warner and Robert S. Pomeroy. Creating compliance: a cross-sectional study of the factors associated with marine protected area outcomes. Marine Policy, 36(4):922–932, July 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X12000048.

Watson:2013:CFG

[WP13] Reg A. Watson and Daniel Pauly. The changing face of global fisheries — The 1950s vs. the 2000s. Marine Policy, 42(??):1–4, November 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13000390.

Ward-Paige:2017:GOS

[WP17] Christine A. Ward-Paige. A global overview of shark sanctuary regulations and their impact on shark fisheries. Marine Policy, 82(??):87–97, August 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305899.

Wielgus:2014:FCC

[WPdR^{+}14] Jeffrey Wielgus, Sarah Poon, Esperanza Carballal del Río, Daylin Muñoz, Daniel Whittle, and Rod Fujita. Fishery cooperatives in Cuba: Potential benefits, legal feasibility, and governance pre-conditions. Marine Policy, 45(??):128–137, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002789.

Waldo:2010:NES

[WPJ10] Staffan Waldo, Anton Paulrud, and Anna Jonsson. A note on the economics of Swedish Baltic Sea fisheries. Marine Policy, 34(3):716–719, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001808.

Willis:2018:HAB

[WPRA18] Cheryl Willis, Eleni Papathanasopoulou, Duncan Russel, and Yuri Artioli. Harmful algal blooms: the impacts on cultural ecosystem services and human wellbeing in a case study setting, Cornwall, UK. Marine Policy, 97(??):232–238, November 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1830397X.

Wright:2015:MGI

[Wri15] Glen Wright. Marine governance in an industrialised ocean: a case study of the emerging marine renewable energy industry. Marine Policy, 52(??):77–84, February 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002838.

Wright:2016:RWT

[Wri16] Glen Wright. Regulating wave and tidal energy: an industry perspective on the Scottish marine governance framework. Marine Policy, 65(??):115–126, March 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1500384X.

Wiber:2010:CMC

[WRP^{+}10] Melanie G. Wiber, Murray A. Rudd, Evelyn Pinkerton, Anthony T. Charles, and Arthur Bull. Coastal management challenges from a community perspective: the problem of ‘stealth privatization’ in a Canadian fishery. Marine Policy, 34(3):598–605, May 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09001833.

Waggitt:2017:CSR

[WRW^{+}17] James J. Waggitt, Alexandra M. C. Robbins, Helen M. Wade, Elizabeth A. Masden, Robert W. Furness, Angus C. Jackson, and Beth E. Scott. Comparative studies reveal variability in the use of tidal stream environments by seabirds. Marine Policy, 81(??):143–152, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305024.

Waggitt:2014:USO

[WS14] J. J. Waggitt and B. E. Scott. Using a spatial overlap approach to estimate the risk of collisions between deep diving seabirds and tidal stream turbines: a review of potential methods and approaches. Marine Policy, 44(??):90–97, February 2014. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X13001498.

Waltham:2015:ECU

[WS15a] Nathan J. Waltham and Marcus Sheaves. Expanding coastal urban and industrial seascape in the Great Barrier Reef World Heritage Area: Critical need for coordinated planning and policy. Marine Policy, 57(??):78–84, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000792.

Wynveen:2015:EPC

[WS15b] Christopher J. Wynveen and Stephen G. Sutton. Engaging the public in climate change-related pro-environmental behaviors to protect coral reefs: the role of public trust in the management agency. Marine Policy, 53(??):131–140, March 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14002929.

Witter:2017:PRA

[WS17] Allison Witter and Joshua Stoll. Participation and resistance: Alternative seafood marketing in a neoliberalera. Marine Policy, 80(??):130–140, June 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16305863.

Wasson:2015:LLE

[WSA^{+}15] Kerstin Wasson, Becky Suarez, Antonia Akhavan, Erin Mc-Carthy, Judith Kildow, Kenneth S. Johnson, Monique C. Fountain, Andrea Woolfolk, Mark Silberstein, Linwood Pendleton, and Dave Feliz. Lessons learned from an ecosystem-based management approach to restoration of a California estuary. Marine Policy, 58(??):60–70, August 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000846.

Wade:2019:IPR

[WSB19] Eric Wade, Ana K. Spalding, and Kelly Biedenweg. Integrating property rights into fisheries management: the case of Belize’s journey to managed access. Marine Policy, 108(??):Article 103631, October 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304858.

Wang:2019:MVC

[WSC19] Ou Wang, Simon Somogyi, and Sylvain Charlebois. Mapping the value chain of imported shellfish in China. Marine Policy, 99(??):69–75, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302628.

Warlick:2018:HWC

[WSG18] Amanda Warlick, Erin Steiner, and Marie Guldin. History of the West Coast groundfish trawl fishery: Tracking socioeconomic characteristics across different management policies in a multispecies fishery. Marine Policy, 93(??):9– 21, July 2018. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307911.

Wellman:2017:CWC

[WSGID17] Emory Wellman, Ariana Sutton-Grier, Meg Imholt, and Adam Domanski. Catching a wave? A case study on incorporating storm protection benefits into Habitat Equivalency Analysis. Marine Policy, 83(??):118–125, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17301616.

Wylie:2016:KSB

[WSGM16] Lindsay Wylie, Ariana E. Sutton-Grier, and Amber Moore. Keys to successful blue carbon projects: Lessons learned from global case studies. Marine Policy, 65(??):76–84, March 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15003905.

Wakita:2014:HUM

[WSO^{+}14] Kazumi Wakita, Zhonghua Shen, Taro Oishi, Nobuyuki Yagi, Hisashi Kurokura, and Ken Furuya. Human utility of marine ecosystem services and behavioural intentions for marine conservation in Japan. Marine Policy, 46(??):53– 60, May 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X14000025.

Wakamatsu:2017:CBR

[WSWM17] Mihoko Wakamatsu, Kong Joo Shin, Clevo Wilson, and Shunsuke Managi. Can bargaining resolve the international conflict over whaling? Marine Policy, 81(??):312–321, July 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17300556.

Watson:2018:MNC

[WT18] Reg A. Watson and A. Tidd. Mapping nearly a century and a half of global marine fishing: 1869– 2015. Marine Policy, 93(??):171–177, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300605.

Williams:2016:GAH

[WTA^{+}16] Rob Williams, Len Thomas, Erin Ashe, Christopher W. Clark, and Philip S. Hammond. Gauging allowable harm limits to cumulative, sub-lethal effects of human activities on wildlife: a case-study approach using two whale populations. Marine Policy, 70(??):58–64, August 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16301956.

Wallace:2015:CPG

[WTD^{+}15] Scott Wallace, Bruce Turris, John Driscoll, Karin Bodtker, Brian Mose, and Gordon Munro. Canada’s Pacific groundfish trawl habitat agreement: a global first in an ecosystem approach to bottom trawl impacts. Marine Policy, 60(??):240–248, October 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15001931.

Wuisan:2012:GIS

[WvLvK12] Lindsey Wuisan, Judith van Leeuwen, and C. S. A. (Kris) van Koppen. Greening international shipping through private governance: a case study of the Clean Shipping Project. Marine Policy, 36(1):165–173, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X11000819.

Wakamatsu:2017:CSS

[WW17] Mihoko Wakamatsu and Hiroki Wakamatsu. The certification of small-scale fisheries. Marine Policy, 77(??):97–103, March 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16308260.

Wang:2019:RMI

[WW19] Yixuan Wang and Nuo Wang. The role of the marine industry in China’s national economy: an input–output analysis. Marine Policy, 99(??):42–49, January 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304226.

White:2017:MWI

[WWH^{+}17] Mathew P. White, Abigail Weeks, Tom Hooper, Luke Bleakley, Deborah Cracknell, Rebecca Lovell, and Rebecca L. Jefferson. Marine wildlife as an important component of coastal visits: the role of perceived biodiversity and species behaviour. Marine Policy, 78(??):80–89, April 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306686.

Waldman:2019:FPI

[WWM^{+}19] S. Waldman, S. Weir, R. B. O’Hara Murray, D. K. Woolf, and S. Kerr. Future policy implications of tidal energy array interactions. Marine Policy, 108(??):Article 103611, October 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X19302994.

Wanyonyi:2016:MFP

[WWT^{+}16] Innocent Ngao Wanyonyi, Andrew Wamukota, Paul Tuda, Victor Alati Mwakha, and Lydia Mukii Nguti. Migrant fishers of Pemba: Drivers, impacts and mediating factors. Marine Policy, 71(??):242–255, September 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300215.

Wade:2010:CIR

[WWW10] Lucie Wade, Hal Whitehead, and Linda Weilgart. Conflict of interest in research on anthropogenic noise and marine mammals: Does funding bias conclusions? Marine Policy, 34(2):320–327, March 2010. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09001092.

Wielgus:2010:EFR

[WZCHS10] Jeffrey Wielgus, Dirk Zeller, Dalila Caicedo-Herrera, and Rashid Sumaila. Estimation of fisheries removals and primary economic impact of the small-scale and industrial marine fisheries in Colombia. Marine Policy, 34(3):506–513, May 2010. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001687.

Wang:2018:DRP

[WZF^{+}18] Yebao Wang, Chunchang Zhang, Long Feng, Cheng Tang, Xiang Yu, Xiangyang Zheng, Xin Liu, and Robert Costanza. Determination of refuge places for oil tankers in emergencies in the Chinese Bohai Sea. Marine Policy, 90(??):95–104, April 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1730800X.

Xiong:2016:USL

[XDG^{+}16] Xiong Xiong, Priscilla D’Amico, Lisa Guardone, Lorenzo Castigliego, Alessandra Guidi, Daniela Gianfaldoni, and Andrea Armani. The uncertainty of seafood labeling in China: a case study on cod, salmon and tuna. Marine Policy, 68(??):123–135, June 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300653.

Xinmin:2019:CAP

[Xin19] M. A. Xinmin. China’s Arctic policy on the basis of international law: Identification, goals, principles and positions. Marine Policy, 100(??):265–276, February 2019.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301982.

Xavier:2018:AWT

[XJT18] Luciana Yokoyama Xavier, Pedro Roberto Jacobi, and Alexander Turra. On the advantages of working together: Social Learning and knowledge integration in the management of marine areas. Marine Policy, 88(??): 139–150, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17303858.

Yates:2014:VWP

[Yat14] K. L. Yates. View from the wheelhouse: Perceptions on marine management from the fishing community and suggestions for improvement. Marine Policy, 48(??): 39–50, September 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X14000736.

Ye:2017:FSD

[YBB^{+}17] Yimin Ye, Manuel Barange, Malcolm Beveridge, Luca Garibaldi, Nicolas Gutierrez, Alejandro Anganuzzi, and Marc Taconet. FAO’s statistic data and sustainability of fisheries and aquaculture: Comments on Pauly and Zeller (2017). Marine Policy, 81(??):401–405, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17300921.

Young:2019:LGS

[YBD^{+}19] Nathan Young, Camilla Brattland, Celeste Digiovanni, Bjorn Hersoug, Jahn Petter Johnsen, Kine Mari Karlsen, Ingrid Kvalvik, Erik Olofsson, Knud Simonsen, Ann-Magnhild Solås, and Helgi Thorarensen. Limitations to growth: Social-ecological challenges to aquaculture development in five wealthy nations. Marine Policy, 104 (??):216–224, June 2019. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18302781.

Yeeting:2016:INE

[YBRBB16] Agnes D. Yeeting, Simon R. Bush, Vina Ram-Bidesi, and Megan Bailey. Implications of new economic policy instruments for tuna management in the Western and Central Pacific. Marine Policy, 63(??):45–52, January 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002870.

Yu:2018:OBO

[YC18] Yaodong Yu and Yen-Chiang Chang. The ‘One Belt One Road’ Initiative and its impact on shipping law in China. Marine Policy, 87(??):291–294, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17303883.

Yagi:2012:AIT

[YCA^{+}12] Nobuyuki Yagi, Michael L. Clark, Lee G. Anderson, Ragnar Arnason, and Rebecca Metzner. Applicability of Individual Transferable Quotas (ITQs) in Japanese fisheries: a comparison of rights-based fisheries management in Iceland, Japan, and United States. Marine Policy, 36(1):241–245, January 2012. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1100100X.

Yang:2014:ESG

[YCHM14] Yuwen Yang, Ross Cullen, Edward Hearnshaw, and Ian A. Macdonald. An evaluation of self-governance in the New Zealand Bluff oyster fishery — The indicator system approach. Marine Policy, 43(??):273–282, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1300136X.

Ye:2015:ICM

[YCY^{+}15] Guanqiong Ye, Loke Ming Chou, Shengyun Yang, Jiaping Wu, Pei Liu, and Changwei Jin. Is integrated coastal management an effective framework for promoting coastal sustainability in China’s coastal cities? Marine Policy, 56(??):48–55, June 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15000445.

Young:2016:WDF

[YFB16] Matthew A. L. Young, Simon Foale, and David R. Bellwood. Why do fishers fish? A cross-cultural examination of the motivations for fishing. Marine Policy, 66(??):114–123, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000257.

Yang:2013:ESG

[YFC13] Yu Wen Yang, Dan Fu, and Ross Cullen. An evaluation of self-governance in the New Zealand Bluff oyster fishery — a modelling approach. Marine Policy, 38(??):163–173, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001303.

Yang:2010:SGW

[YFR10] Yu Wen Yang, Allen Frazer, and Eugene Rees. Selfgovernance within a QMS framework — The evolution of self-governance in the New Zealand Bluff oyster fishery. Marine Policy, 34(2):261–267, March 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900102X.

Yiallourides:2018:ITF

[Yia18] Constantinos Yiallourides. It takes four to tango: Quadrilateral boundary negotiations in the North-East Atlantic. Marine Policy, 87(??):78–83, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17305183.

Yang:2019:CSM

[YLZ^{+}19] Xiangbo Yang, Ning Liu, Peidong Zhang, Zhen Guo, Changlei Ma, Peipei Hu, and Xuanzhao Zhang. The current state of marine renewable energy policy in China.

Marine Policy, 100(??):334–341, February 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304718.

Yoshino:2017:TPQ

[Yos17] Keiko Yoshino. TURFs in the post-quake recovery: Case studies in Sanriku fishing communities, Japan. Marine Policy, 86(??):47–55, December 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730115X.

Young:2016:ITL

[You16a] Margaret A. Young. International trade law compatibility of market-related measures to combat illegal, unreported and unregulated (IUU) fishing. Marine Policy, 69(??):209–219, July 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X16000385.

Young:2016:GAO

[You16b] Oran. R. Young. Governing the Arctic Ocean. Marine Policy, 72(??):271–277, October 2016. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0308597X1630224X.

Yang:2017:ICB

[YP17] Di Yang and Robert Pomeroy. The impact of communitybased fisheries management (CBFM) on equity and sustainability of small-scale coastal fisheries in the Philippines. Marine Policy, 86(??):173–181, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304670.

Yates:2013:IRN

[YPS13] K. L. Yates, A. Payo Payo, and D. S. Schoeman. International, regional and national commitments meet local implementation: a case study of marine conservation in Northern Ireland. Marine Policy, 38(??):140–150, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X12001285.

Yamakita:2017:IIM

[YSJU^{+}17] Takehisa Yamakita, Kenji Sudo, Yoshie Jintsu-Uchifune, Hiroyuki Yamamoto, and Yoshihisa Shirayama. Identification of important marine areas using ecologically or biologically significant areas (EBSAs) criteria in the East to Southeast Asia region and comparison with existing registered areas for the purpose of conservation. Marine Policy, 81(??):273–284, July 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16303761.

Yeh:2015:TJC

[YTSO15] Yun-Hu Yeh, Huan-Sheng Tseng, Dong-Taur Su, and Ching-Hsiewn Ou. Taiwan and Japan: a complex fisheries relationship. Marine Policy, 51(??):293–301, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002504.

Yagi:2010:MPA

[YTTK10] Nobuyuki Yagi, Akira P. Takagi, Yukiko Takada, and Hisashi Kurokura. Marine protected areas in Japan: Institutional background and management framework. Marine Policy, 34(6):1300–1306, November 2010. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X10001077.

Yoskowitz:2016:GMO

[YWC^{+}16] David W. Yoskowitz, Sandra R. Werner, Cristina Carollo, Carlota Santos, Travis Washburn, and Gary H. Isaksen. Gulf of Mexico offshore ecosystem services: Relative valuation by stakeholders. Marine Policy, 66(??):132–136, April 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000809.

Yu:2019:ESE

[YY19] Jinkai Yu and Wang Yin. Exploring stakeholder engagement in mariculture development: Challenges and prospects for China. Marine Policy, 103(??):84–90, May 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306559.

Yamakita:2015:IIM

[YYN^{+}15] Takehisa Yamakita, Hiroyuki Yamamoto, Masahiro Nakaoka, Hiroya Yamano, Katsunori Fujikura, Kiyotaka Hidaka, Yuichi Hirota, Tadafumi Ichikawa, Shigeho Kakehi, Takahiko Kameda, Satoshi Kitajima, Kazuhiro Kogure, Teruhisa Komatsu, Naoki H. Kumagai, Hiroomi Miyamoto, Kazushi Miyashita, Haruyuki Morimoto, Ryota Nakajima, Shuhei Nishida, Kou Nishiuchi, Shingo Sakamoto, Masayoshi Sano, Kenji Sudo, Hiroya Sugisaki, Kazuaki Tadokoro, Katsuhiko Tanaka, Yoshie Jintsu-Uchifune, Kentaro Watanabe, Hiromi Watanabe, Yumiko Yara, Norishige Yotsukura, and Yoshihisa Shirayama. Identification of important marine areas around the Japanese Archipelago: Establishment of a protocol for evaluating a broad area using ecologically and biologically significant areas selection criteria. Marine Policy, 51(??):136–147, January 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14001833.

Yan:2015:MIC

[YYYL15] Xiao Yan, Liang Yan, Xi-Long Yao, and Ming Liao. The marine industrial competitiveness of blue economic regions in China. Marine Policy, 62(??):153–160, December 2015. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X15002663.

Zagonari:2018:RIE

[Zag18] Fabio Zagonari. Responsibility, inequality, efficiency, and equity in four sustainability paradigms: Policies for a shared sea from a multi-country analytical model. Marine Policy, 87(??):123–134, January 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302737.

Zanoli:2015:VUM

[ZCD^{+}15] Raffaele Zanoli, Lorenzo Carlesi, Roberto Danovaro, Serena Mandolesi, and Simona Naspetti. Valuing unfamiliar Mediterranean deep-sea ecosystems using visual Qmethodology. Marine Policy, 61(??):227–236, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002286.

Ziegler:2019:APG

[ZEP^{+}19] Friederike Ziegler, Ole Ritzau Eigaard, Robert W. R. Parker, Peter H. Tyedmers, Erik Skontorp Hognes, and Sepideh Jafarzadeh. Adding perspectives to: “Global trends in carbon dioxide (CO_{2}) emissions from fuel combustion in marine fisheries from 1950–2016”. Marine Policy, 107(??):Article 103488, September 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309643.

Zervaki:2018:HSC

[Zer18] Antonia Zervaki. Human security and climate change mitigation: the case of ocean governance. Marine Policy, 98 (??):286–294, December 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18306973.

Zhang:2012:FPC

[ZFG12] Junjie Zhang, Jason Fleming, and Ralf Goericke. fishermen’s perspectives on climate variability. Marine Policy, 36(2):466–472, March 2012. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X11001035.

Ziegler:2014:SSM

[ZH14] Friederike Ziegler and Sara Hornborg. Stock size matters more than vessel size: the fuel efficiency of Swedish demersal trawl fisheries 2002–2010. Marine Policy, 44 (??):72–81, February 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X13001413.

Zou:2018:ICC

[ZH18] Leilei Zou and Henry P. Huntington. Implications of the Convention on the Conservation and Management of Pollock Resources in the Central Bering Sea for the management of fisheries in the Central Arctic Ocean. Marine Policy, 88(??):132–138, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730550X.

Zhang:2016:CFD

[Zha16] Hongzhou Zhang. Chinese fishermen in disputed waters: Not quite a “people’s war”. Marine Policy, 68(??):65–73, June 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000701.

Zhang:2018:FCS

[Zha18] Hongzhou Zhang. Fisheries cooperation in the South China Sea: Evaluating the options. Marine Policy, 89 (??):67–76, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X17307145.

Zhao:2014:DQC

[ZHH14] Rui Zhao, Stephen Hynes, and Guang Shun He. Defining and quantifying China’s ocean economy. Marine Policy, 43(??):164–173, January 2014. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X1300122X.

Zimmerhackel:2019:EIE

[ZKR^{+}19] Johanna S. Zimmerhackel, Marit E. Kragt, Abbie A. Rogers, Khadeeja Ali, and Mark G. Meekan. Evidence of increased economic benefits from shark-diving tourism in the Maldives. Marine Policy, 100(??):21–26, February 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18305116.

Zheng:2019:GSP

[ZLL19] Sining Zheng, Qiao Liang, and Qiang Liu. Governance structure and performance of mariculture Sci-Tech parks: Evidence from Zhejiang Province, China. Marine Policy, 109(??):Article 103670, November 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307966.

Zhu:2017:IPR

[ZLSL17] Mo Zhu, Kevin X. Li, Wenming Shi, and Jasmine Siu Lee Lam. Incentive policy for reduction of emission from ships: a case study of China. Marine Policy, 86(??):253–258, December 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17303457.

Zivin:2015:VBF

[ZM15] Joshua Graff Zivin and Jamie Mullins. Vessel buybacks in fisheries: the role of auction and financing structures. Marine Policy, 53(??):188–197, March 2015. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X14002668.

Zimmer:2014:IOP

[ZMM^{+}14] Beth Zimmer, Leslie Manzello, Keld Madsen, James Sinclair, and Rebecca E. Green. An innovative ocean planning tool for the Atlantic outer continental shelf: the EcoSpatial Information Database. Marine Policy, 45(??):60–68, March 2014. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X13002492.

Zhang:2016:NSR

[ZMZ16] Yiru Zhang, Qiang Meng, and Liye Zhang. Is the Northern Sea Route attractive to shipping companies? Some insights from recent ship traffic data. Marine Policy, 73(??):53–60, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302081.

Zheng:2018:EDA

[ZMZ18] Hui Zheng, Hairong Mu, and Xin Zhao. Evaluating the demand for aquaculture insurance: an investigation of fish farmers’ willingness to pay in central coastal areas in China. Marine Policy, 96(??):152–162, October 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301829.

Zhang:2016:SFR

[ZP16] Pengfei Zhang and Edward Phillips. Safety first: Reconstructing the concept of seaworthiness under the maritime labour convention 2006. Marine Policy, 67(??):54–59, May 2016. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16000440.

Zeller:2018:PBT

[ZP18] Dirk Zeller and Daniel Pauly. The ‘presentist bias’ in time-series data: Implications for fisheries science and policy. Marine Policy, 90(??):14–19, April 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306334.

Zampoukas:2013:MME

[ZPB^{+}13] Nikolaos Zampoukas, Henna Piha, Emanuele Bigagli, Nicolas Hoepffner, Georg Hanke, and Ana Cristina Cardoso. Marine monitoring in the European Union: How to fulfill the requirements for the marine strategy framework directive in an efficient and integrated way. Marine Policy, 39(??):349–351, May 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1200245X.

Zeller:2016:SCA

[ZPT^{+}16] D. Zeller, M. L. D. Palomares, A. Tavakolie, M. Ang, D. Belhabib, W. W. L. Cheung, V. W. Y. Lam, E. Sy, G. Tsui, K. Zylich, and D Pauly. Still catching attention: Sea Around Us reconstructed global catch data, their spatial expression and public accessibility.

Marine Policy, 70(??):145–152, August 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302421.

Zhang:2019:BME

[ZQZ19] Chao Zhang, Yongbo Quan, and Haiyue Zhong. The building of marine ecological civilization and sustainable development: Conference report. Marine Policy, 110(??):Article 103627, December 2019. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305287.

Zhang:2018:MOW

[ZS18] Jihua Zhang and Wenjing Sun. Measurement of the ocean wealth of nations in China: an inclusive wealth approach. Marine Policy, 89(??):85–99, February 2018. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17304827.

Zhang:2018:FFC

[ZSH18] Dengjun Zhang, Marius Sikveland, and Øystein Hermansen. Fishing fleet capacity and profitability. Marine Policy, 88(??):116–121, February 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17302373.

Zhang:2019:NSR

[ZSZPR19] Pengfei Zhang, Desai Shan, Minghua Zhao, and Nicola Pryce-Roberts. Navigating seafarer’s right to life across the shipping industry. Marine Policy, 99(??):80–86, January 2019. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18300927.

Zhang:2019:FWS

[ZT19] Dengjun Zhang and Ragnar Tveterås. A fish out of water? Survival of seafood products from developing countries in the EU market. Marine Policy, 103(??):50–58, May 2019.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306626.

Zhao:2013:WVI

[ZTAO13] Minghua Zhao, Marilyn Tyzack, Rodney Anderson, and Estera Onoakpovike. Women as visible and invisible workers in fisheries: a case study of Northern England. Marine Policy, 37(??):69–76, January 2013. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12000747.

Zappes:2013:CBS

[ZVP^{+}13] Camilah Antunes Zappes, Camila Ventura da Silva, Mônica Pontalti, Mônica Lauriano Danielski, and Ana Paula Madeira Di Beneditto. The conflict between the southern right whale and coastal fisheries on the southern coast of Brazil. Marine Policy, 38(??):428–437, March 2013. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X12001637.

Zhang:2015:MLC

[ZZ15a] Pengfei Zhang and Minghua Zhao. Maritime Labour Convention, 2006 and the Chinese seafarers: How far is China to ratification? Marine Policy, 61(??):54–65, November 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15002171.

Zhu:2015:FAA

[ZZ15b] Ling Zhu and Ya Chao Zhao. A feasibility assessment of the application of the Polluter-Pays Principle to shipsource pollution in Hong Kong. Marine Policy, 57(??):36– 44, July 2015. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X15000597.

Zhang:2016:NSF

[ZZ16] Guobin Zhang and Pai Zheng. A new step forward: Review of China’s 2016 legislation on international seabed area exploration and exploitation. Marine Policy, 73(??):244–255, November 2016. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302378.

Zhang:2017:MHC

[ZZ17a] Pengfei Zhang and Minghua Zhao. Maritime health of Chinese seafarers. Marine Policy, 83(??):259–267, September 2017. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17304074.

Zhu:2017:SLP

[ZZ17b] Ling Zhu and Ruixuan Zhuo. A survey of the legal and policy framework for controlling, compensating and criminalizing ship-source pollution in Hong Kong. Marine Policy, 76(??):38–47, February 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16304948.

Zou:2017:LCC

[ZZ17c] Keyuan Zou and Lei Zhang. Legal challenges of carbon capture and storage in the South China Sea region. Marine Policy, 84(??):167–177, October 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1730043X.

Zhao:2017:CWS

[ZZZ^{+}17] Minghua Zhao, Lijuan Zhao, Pengfei Zhang, Jianjun Wu, Kate Pike, and Emma Broadhurst. Chinese women seafarers: a case study of the women cadets in Shanghai. Marine Policy, 83(??):40–47, September 2017. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301082.

Zhang:2018:SKT

[ZZZL18] Huiming Zhang, Yu Zheng, Dequn Zhou, and Xingle Long. Selection of key technology policies for Chinese offshore wind power: a perspective on patent maps. Marine Policy, 93(??):47–53, July 2018. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302562.